A Complete Bibliography of Publications in Software—Practice and Experience

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 May 2024 Version 3.37

Title word cross-reference

0 [GW96]. 1 [GW96]. $1.50 [Bar78d]. $11 [Bar84a]. $12.00 [Rob72]. $13 [Bar84a]. $13.00 [Rob72]. $18.50 [Jon74]. $185 [Bar79b]. $19.30 [Lan74a]. $19.50 [Dav78]. $25.00 [Pet77, And78]. 3 [BE02, FMA02]. $31-25 [Pet77]. $31.35 [Bri82]. 32 [VED06]. $35.00 [Inc86]. $39.50 [Sim83]. 5 [CPMAH^{+}20]. $58.50 [Wal81a]. $6.95 [Tho74]. 64 [AM10, VED06]. 68 [Ear76, Hol77]. $68.25 [Pit82]. $7.00 [Bar72a]. $7.50 [Bar78d]. $7.95 [Bar76a, Lav77]. $78.50 [Sim83]. 8 [Plu74, SF85]. $8.95 [Bar82a, Bar82c, Bar84b]. $9.75 [Bar77e, Mul76]. $9.80 [Atk79a]. $9.95 [Bar82a, Bar82c, Bar84b]. < [SMGMOFM07a, SMGMOFM07b]. > [SMGMOFM07a, SMGMOFM07b]. ^{+} [NN24]. ^{2} [MST13, MDB19]. ^{3} [DS09]. ^{4} [MSR^{+}07]. ^{♮} [PK04]. ^{T M} [MZB00, Win02]. B [CSD23]. ∀ [DB21b]. k [AW93, Mer93]. κ [MG94]. μ [BS90c, BDS^{+}92, SMNB21]. N [MS98, Coh98, KST94, YAVHC21]. P^{3} [DC03]. P M_{2.5} [CLD^{+}17]. q [GSR17]. τ [TSZ14, UDS^{+}07].

-ary [MS98]. -aware [MSWC23]. -bit [AM10, SF85, VED06]. -gram [Coh98, KST94, YAVHC21]. -grams [GSR17]. -level [FM77]. -queens [Plu74]. -R [Ear76, Hol77]. -shortest-paths [MG94]. -System [BS90c].

. [Bis81b]. .NET [Coo04, Han04].

0 [Bar81, Edw98a, Edw98b, Gru83, Llo82, Val77a, Val78, Wal83b]. 0-12-380230-X [Dea86]. 0-13-751975-3 [Edw98a, Edw98b]. 0-201-03822-6 [Llo82]. 0-201-08571-2 [Wal83b]. 0-201-13787-9 [Gru83]. 0-387-10256-6 [Cav83a]. 0-387-9642 [Mee87]. 0-471-19161-2 [Cor99a, Cor99b]. 0-471-87834-0 [Ano87a]. 0-471-88311-5 [Ano88c]. 0-471-89052-9 [Ree84b]. 0-471-91128-3 [Ano88a]. 0-471-91754-0 [CO88]. 0-521-21965-5 [Bar81]. 0-521-30689-2 [Lon88]. 0-521-31423-2 [Lon88]. 0-632-01309-5 [Rop88a]. 0-632-01609-4 [Bow88]. 0-7204-0437-1 [Val78]. 0-7204-2802-5 [Val77a]. 0-7458-0092-0 [Mar88]. 0-7458-0110-2 [Vel88]. 0-85012-612-6 [Rop88b]. 0-85312-846-4 [Sto88]. 0.90 [Bar75a].

1 [BH87, Col72b, Hus86, Les72, Pal74, SHR80, Val78, Yuv77a, Wis74, Nic72]. 1-85058-035-9 [Ano88b]. 1-like [EBD^{+}74]. 1.50 [Bar72b]. 1.60 [Eve73]. 1.65 [Bar74c]. 1.75 [Wel72]. 1.9 [Kur78]. 1.90 [Bar74f]. 10.50 [Rob82b]. 10.75 [Ree82, Wal83c]. 10/LSD [Les72]. 103 [Whi87]. 10646 [Wu01]. 10th [WCK11]. 11’ [Hay80, BLLP04, DM84, GKM90, HGWBS75, Mat80]. 11-11 [BLLP04]. 11.20 [Val80]. 11.70 [Bar79a]. 11.80 [Bry77]. 11.95 [Jac84]. 12 [CST75]. 122 [SFS97c]. 13.50 [Rop88a]. 13.95 [CO88, Wil84b]. 130 [Hal82]. 14 [Val79]. 14.00 [Bra80]. 14.20 [Wal84b]. 14.95 [Bra75]. 15 [KP94]. 15.00 [For72, Roh77a]. 15.50 [Ree84b]. 15.95 [Cam85]. 1517 [Ano23m]. 16.50 [Vel88]. 18.50 [Mar88, Sha83]. 19 [ACKJ22, GVBG22, KRK21, MVM^{+}22, MAA^{+}22, OHD22, PPPP23, SKD^{+}22]. 19.50 [Dea86, Sto88, Wal86b]. 190 [Her84]. 1900 [Far74, NY78, REC75, WQ72]. 1969 [Rob72]. 1971 [Bar73e]. 1973 [Val77a].

1975 [Val78]. 1980 [Bis81a]. 1982

[Bar83a, Bar83a]. 1998 [Wri98]. 1A [Wit83]. 1D [MV95]. 1D-FFT [MV95].

2 [Bar74a, Bar74b, Bar80b, Bud85, Cor88b, Dun75, Fos86, Gut87, Hop80b, MS74a, OSW92, Pro92, Rai81, RCH^{+}23, Tag88, Ter86, UGBW91, Val76b, Wal83b, Woo86, Hun72, Mee87, Ano87a, Bow88, Bar78b]. 2-75 [Lav78]. 2.0 [Hsu12]. 2.0-based [Hsu12]. 2.25 [Atk83, Cou84b]. 2.50 [Bar73d, Bar75c, Jac71]. 2.55 [Bar73b]. 2.6 [TCM07]. 2.65 [Bul72a]. 2.95 [Cou84a, Tho77]. 20 [Liv75]. 20.00 [Gar86, Wil87]. 2010 [Obe11]. 2011 [BN13, PL14]. 205 [LS84]. 22.50 [Rob82a]. 25 [Val77b]. 25p [Bar73a]. 2900 [EP79, Iza80]. 2D [BBGP01]. 2nd [Bar75a, Bul72a, Cou84a, Llo82, McD71].

3 [Bar72c, Edw98a, Edw98b, Hun72, Oes71]. 3-540-10256-6 [Cav83a]. 3.10 [Bar71]. 3.15 [McD71]. 3.30 [Bar76e]. 3.50 [Bar78c, Bis79a, Nic72, Wis74]. 3.75 [Bar76d, Hut74, Mil72]. 3.95 [Han78b]. 30 [XZ01, XZ03]. 30.00 [Bar83a]. 315 [Pra96a, Pra96b]. 32.10 [Ano88c]. 336 [Fen94a]. 360 [Haz74, RS76, Haz72, Lar71]. 370 [BB75]. 39.95/$32.08 [Cor99a, Cor99b]. 3D [MTT83, Woo71, XAN07].

4 [Lar75a]. 4.0 [ACK^{+}22, APR22, CJTK22, DAC^{+}21, MGJ^{+}24, SSKG22, ZWQC22]. 4.00 [Bar74d, Han78a, Rog74]. 4.20 [Bar75f]. 4.25 [Eme84]. 4.40 [Haz72]. 4.50 [Haz71, HW77, Rop88b]. 4.80 [Rog71]. 4.85 [Bul73]. 4.95 [Atk79b, Col77b, Ree76]. 40 [Bar76b, Bul72b, Rog73, Wil74a]. 450 [Bis86].

5 [Bar77c, Bar81, LG73, Val77a]. 5-50 [Bar77c]. 5.00 [Wic72a]. 5.25 [Ano73a, Lan75, Ros74]. 5.50 [Bar73e, Bar78b, Wan82]. 5.75 [Han72, Mer74]. 5.90 [Ken77]. 5.95 [Bar80d, Edm82]. 50 [Bar77c]. 51 [Ano23m]. 512 [CL23]. 56th [Bar82a, Bar82c, Bar84b]. 5G [BSDF20, CVGCGCFF23, XCZ^{+}21].

6 [Lar71, Llo82]. 6.00 [Bar75b, Ree73]. 6.25 [Bar73c, Bar75d]. 6.50 [Bar75e, Hop74]. 6.75 [Sha72, Wil72]. 6.95 [Bis84]. 60 [HSW75, Hut76, Wic72b, WJ76, Wil76]. 6000 [Bak72, Rob79, Yuv77a, Yuv77c]. 6000-Series [Bak72]. 6000/7000 [Has77, Rob79, Yuv77a]. 653 [DKM11]. 68 [DV85, FM78, IR80, Inc81, PH86, ST79, She75, Woo72, Woo84, Wyv77, Bar74e, Bra80]. 68-R [Bar75a, FM78]. 68K [Poh81].

7 [HCD84, WK06a, Bar76b]. 7.00 [Bar82b, Lar75a, Ree75]. 7.30 [Flo74]. 7.35 [Woo74]. 7.50 [Bis79b]. 7.60 [Ald72]. 7.80 [Bar76c]. 7.85 [Bar77b]. 7.95 [Ano88b]. 7000 [Rob79]. 70’s [Spo71]. 750 [HJ88b]. 77 [HWS^{+}88, Lar81, Edm86, RB82].

8 [Ell72, Har71b, KL21]. 8.00 [Ear77, Hop73]. 8.20 [Bux78]. 8.25 [Edm86, How76]. 8.50 [Dav74, Han77a]. 8.75 [Flo79]. 8.80 [Bar77d]. 8.95 [Cou85a].

9 [Gru83]. 9.45 [Bar80e, Val76a, Wal82]. 9.70 [Edw77]. 9.80/$16.60 [Lav77]. 9.95 [Ano88a, Cou85b]. 90 [SM90].

= [Edw77].

AAOP [JZ10]. ABACUS [JT00]. Abbreviations [New86, MT84a]. ABCD [KAS^{+}16]. Abecedarian [Bar76d]. Ability [YH97]. Ablego [ZA07]. Abmash [OMM15]. Abnormal [BMZ92, XLZ^{+}20]. abort [PDPQ22]. Abowd [Wri98]. Abstract [AD87, BCHR81, CFL84, Die97, ELRV93, Fle82, FH82a, Gri80, GH84, HOS85, Ian90, Jal87, Lar90, NPW72, Pow87, AG06, CFC15, MGG^{+}09]. Abstraction [BR95, Fel81, GR79, LHC97, Sal79a, SL78, CLSE05, GAB^{+}23, WZLN08]. Abstractions [Kat83a, KS87, Mor80, AYdS^{+}06, CPD13, SM01, VGF21]. academia [CEMS24, JPM23]. Academic [Bar75f, Bux78, Dav78, Dea86, Hop74, Inc86, Jon74, Rob72, Sha72, SFB13, Whi87, Wic72a, Wil72, Wil87, Bar77d, BLC19, Han77a]. Academics [Ano71e]. accelerated [NPHJ18]. Accelerating [TT82, WLZ^{+}22]. acceleration [LFH^{+}22, ZKW21]. accelerator [ZZJ21]. Acceptance [Mat83b, WWB03]. Access [BMY03, Coh73, CFL84, Cow87, Day83, DPS03, Hun81, LN71, PSR83, Poo71b, Ree71, Sil81, SY79, SY86, SL87, Sti79, Tag88, TB72, Wil73, WMG94, WP96, BSC^{+}05, CKL^{+}02, CYJ^{+}22, Gay80, GAB^{+}23, HNW^{+}01, HLW08, KKN04, MLC02, NH03, WJC^{+}14, KT01b, SROAdM^{+}08]. Access-control [Sil81]. Accessed [SW87, HJC00]. Accesses [Har92, PF97]. Accessing [Ker80]. accessor [GAQO23]. accident [JH03]. Accommodating [Not90]. Accounting [CW82b, Sre76]. accounts [BLNU15]. Acculock [XXZ13]. accumulator [CRT80]. accuracy [GAQO23, MKA^{+}22, PKvdWB17]. Accurate [Oli83, Bin06, Spi04, WC08, XXZ13, YMH16]. ACET [LPF^{+}11]. Achieve [Nee77c]. achieved [BLC19]. Achieving [CW97, WW09, WC08]. ACID [FZ98]. Ackermann [Wic77]. Acknowledgements [Ano17m]. Acos [SH17]. acquire [LZLL22]. Acquiring [Ard87, Ano80b, Jos79, Jos80]. Acquisition [Har80a, SDF^{+}21, WPL^{+}21, ZWAL22]. acronyms [CK15]. across [DGRB15, DW91, ZWML14]. action [ST12, TBF^{+}22]. Actions [Mös88, Set84, TE90, FZS^{+}17, OMM15, ZRX^{+}99]. activation [SSO13]. activator [SSO13].

Active [AN88, Car98, CC97, Cho96, MK96, RMC97, SCD^{+}23, TS02]. ActiveX [Lev01]. activities [ABBS22, NRUP21, SJK^{+}21]. Activity [FM78, HLR^{+}03, CmJHL18, HLR20, SH17, aSZP^{+}16, XMTL21]. actor [AMSG23]. actor-based [AMSG23]. actors [BPDL23]. ACTUS [PCM83]. acyclic [LSZ16]. Ada [GWA91, AB88, Ard87, Bar80a, BAP87, Bri84, Bru84, BK86b, BK87, DHGR92, FIL86, FFW96, Gau95, Geh85, Hol83, Ibs84, IM93, Jac85, KO86, LMSP92, LvLS84, Lun89, LF90, PCBE96, RA87, REMC81, SB93, WL81a, vK87, Wal83c, Wal84b]. Ada-based [LvLS84]. Adam [LvLS84]. ADAMS [DFOT10]. Adaptability [JZ10, Han77b, KKLL99]. Adaptable [Ell79a, VAP^{+}17, WN88, BHR15, BS89, ZKW21]. Adaptation [AE06b, AE06a, PA91, CLC09, CRGIP15, GBE^{+}09, GDH13, HBB22, HK06b, IHS^{+}14, KY05, NS01b, PDBG10]. Adapters [HL94]. Adapting [LLS06, MNW14, SSCdA^{+}03, HIR06, MA20a]. adaption [SH22, Wal81a]. Adaptive [AS97b, HMS^{+}95, LH82, LDI98, LXY^{+}11, NZH20, SXWL17, TP03, VBH^{+}98, VC90, AF99, AE06b, AE06a, dODP21, BDA20, BFNP08, CPCL10, FFF^{+}13, GLT08, GM22, HZWC22, HKC^{+}12, HML04, LKC12, MSC23, Mos06, NB19, PRA^{+}06, PCML09, PLPA22, PDPM^{+}16, SIC^{+}20, ST12, SLRS06, SJP^{+}09, SM15, Sta07, VGF21, YFC06, ZXT^{+}17, ZCN06, ZWX^{+}21]. adaptivity [MK03]. ADARC [JZLP20]. Ada(R) [GC84]. Added [Bro80]. Addendum [Sau88]. Adding [CD94, JGT95, Ron99, Ste92, Str83a, ZM95, ABL08, KGL06, MSB18]. Addison [Bar76e, Bis79b, Cam85, Cou85a, Ear77, Gru83, Jac84, Llo82, Wal83b, Wil84b]. Addison-Wesley [Bar76e, Bis79b, Cam85, Cou85a, Ear77, Gru83, Llo82, Wal83b, Wil84b]. Address [HEV^{+}98, Wil79, CRT80, GNSP12, LGZ^{+}08].

Addresses [Bel74, HP87]. Addressing [FS11, Har92, vK87, DFRR15, FS13, JK14]. Addyman [Bis79a]. ADIC [BRMO97]. adjusting [BG93, WZH01]. Administration [BR97, Bur98, CWD08, Ped86, FSS99]. Administrator [Gen81]. Administrators [GST92]. admission [MNH04]. admitted [LQL^{+}22]. Adopting [FBH^{+}23]. Adoption [KHGSS12, FBV22, PdSCJM22, YQL22]. ADT [CS91a]. Adtpp [NSM16]. Advanced [DB21b, FL75b, MVM^{+}22, BB10, HAM18, JC22, MMM18, RMSMML^{+}11, Sav04, For72]. advancements [JC19]. Advances [Jon74, Wil72]. Advantages [DRG11, HKW77]. adversarial [JFZ^{+}21]. adversary [McI99]. advertisement [HLH15]. advertising [LS16, LLJ^{+}23]. Advice [Ans86]. Advisory [AC80b, Ans86, DCA82, SB22]. AEDBAR [VL73]. aerial [Ano23m, SGA21]. Aesthetic [Blo93]. Aesthetics [CP96]. Aesthetics-based [CP96]. affecting [MCLL21]. affective [ZZC^{+}17]. affine [NS23]. After [Nee75]. AG [Car97]. Again [Hut78, Sal79b]. against [AGG06, Mid86]. age [MVM^{+}22, PMC22]. agency [ZWAL22]. Agent [BDMP17, BHR^{+}02, BPR01, BMM19, CCP06, CNAM^{+}10, DO99, DS09, GHM^{+}06, GCK^{+}02, HL02a, KT01a, LM02, MAJ15, NZPWR22, Pei02, RCH^{+}23, TKT^{+}07]. agent-based [BMM19, MAJ15, NZPWR22, TKT^{+}07]. Agent-oriented [BDMP17]. Agents [GdLC04, Lib97b, BLE^{+}08, CCP06, FZ12, HBM06, KY05, MKC11, PL08]. aggregate [Mid79]. Aggregating [dSRdSS^{+}21]. aggregation [FO10, FKL^{+}13, FSC08, WMSY12]. Agile [DPAG11, GEI^{+}11, GH11, Han11, dSdMSNO^{+}11, ABBS22, BBS11, FSR11, MdCGdC^{+}17, BBB^{+}22, GH09]. agility [VVG^{+}24]. Aglets [OT02]. agreement [ASP^{+}19, BS19, DTB12]. agreement-aware [BS19]. Ahab [VSID17]. ahead [DSD^{+}19, HKM^{+}09, PES^{+}20]. ahead-of-time [HKM^{+}09]. Aho [KAO23, NK07]. AHP [AHRR22b]. Aid [AC75, BCL^{+}94, CT90, CP76, Gri80, Gri82, RR85, Bud85]. AIDA [CC87]. aided [CGK89, FR78, KCS^{+}20, LPT82, SM15]. Aids [CL83, Fox78, Sco77b, Val76a]. AIMS [YSM95]. Air [DP85, MPN^{+}95]. aircraft [CGH^{+}15, MdCGdC^{+}17]. airplane [LLK04]. airplane-landing [LLK04]. Ajalon [PWI^{+}21]. Ajanta [KT01a]. Al [Ano23m]. ALADIN [FHS92]. ALCHEMIST [LTV96]. Alchemy [GZX^{+}21]. Alcock [Ree78]. ALEPH [Gru79]. Alex [Haz72]. Algebra [MV86, HBC15, SCPL23]. Algebraic [IR80, vHLB^{+}88, HM12, NSM16]. ALGOL [Bar74e, Woo74, Bra80, Cor82, AvdSGS80, BW71, BCP71, Bro74, Ear76, HSW75, Hud72, Kaw79, Mid79, PH86, She75, VV84, WJ76, Woo72, AM78, DV85, FM78, Hol77, IR80, Inc81, KA87, NSKK83, ST79, Sha77, Wal86a, Wic72b, Woo84, Hop74, Bar75a, Woo74, Fox79]. ALGOL-like [VV84, BW71, Kaw79]. Algorithm [Bul87, CCM96, Coh98, Coo83, Dro86, Fen96, Fis86b, Gor94, GT92, Gri86, Gru79, GF80, Hal86, Han95, Inn77, Kan97, KST94, KS89, Kur81, LDI98, MG94, Mac77a, MC91, McG89, MHB90, MN80, MS94, PK89, Poo71a, Rai92, SMFBB93, Smi91, Smi94, Thi96, VWB91, Wal90, dCV88, dV89, Abe07, Abe10, AHRR22a, AA20, AA21, AS87, BLM00, BMAV05, CWS07, Deo00, Deo02, Gai82a, GARSR18, GZW^{+}22, GJ22, GF78, HZWC22, HL02a, Hug82, IMKN12, IK15, JSPP21, JZLP20, KSBL22, KSH11, Kir07, KSDG24, LNhCW16, Maa06, MUE23, MGL19, OG16, Ple99, PA01, PP16, SB21, SAC06, SS07, SI10, SS23, Sta07, TRGA18, Tor23, TC07, WW00, XHP^{+}21, ZYW^{+}20, ZWQC22, RCC17]. Algorithm-oriented [MS94]. algorithmic [GVL10, OY10]. Algorithms [ACCM83, CRR94, CSR93, Cd91, CPHS83, DS86a, DS88, DB86, ELRV93, Gai82b, GK21, HJS89, Har80c, HSW75, IC85, Jar75, JTU96, Kob77, Kra97, Lec95, LES95, McG82, Mon96a, Mon96b, Mus97, Nic98, Nør91, Shr76, de 82, ASA22, BMY06, BST10, BG01, BGSG20, CRB^{+}11, CO88, CLCC15, CCT01, Col79, Deo02, DS03, FGK^{+}00, FCA12, Gol81b, HB18, JT00, KA22, KS01a, Man18, Mha05, MAW^{+}16, MCHN05, NS23, NLA15, RR05, SCL00, ST14, SA20, THG17, TRRK21, UCCPM19, VMY24, VDG^{+}00, Lin98a, Llo82, Edw77, Wil84b]. Alias [Boy01, MW93]. aliasing [Cor84, ZC01, NL01]. Alibaba [XYW^{+}24]. alignment [RJZ^{+}20]. All-in-one [Kat17]. all-layer [ZZJ21]. all-pairs [GK21, Tor23]. Allen [Ano73a, Val80]. alleviating [LB02]. Allison [Lon88]. Allocating [PH84]. Allocation [App89a, App89b, DF84, DDZ94, GM85a, Gom74, GW96, Han90, LH82, OLS89, QSA90, VSM87, AS87, BCF00, Bur16, CW08, GWZ^{+}20, KJB11, KSH11, MGT20, SS03, XMTL21, ZXT^{+}17]. allocation-aware [GWZ^{+}20]. Allocator [NP98, Vo96, JSC^{+}10, MRR^{+}08, MSK01]. Allocators [GZ93]. Allowing [Poo71a]. Allworth [Wan82]. Almost [SW86a, IIL17, ML20]. alone [Wil74b]. along [NM19]. Alpha [Wic72a, MDWD01]. Alphabet [TP97, Gu05]. Also [Bar74e, Wad85]. Alt [Jon74, Wil72]. altered [Wic81]. Alternative [And82a, BAP95, Pow95, CMF^{+}17, CW82a, SB03]. Alternatives [DO91, FH92a, AÇCT22, BKP^{+}22, HJ14]. alto [MDWD01]. Amazon [CCE^{+}21]. AmbientTalk [MVT^{+}09]. Ambiguities [WSH77]. ambiguity [Par85b]. Ambiguous [HP87, Sit79, MGP03]. ambulance [SM15]. American [Bar76a, Bar77e, Wel72]. AMGA [AKL^{+}09]. Ammeraal [Ano88a].

Amoeba [vRvST89]. Among [Han79b, CD15, SWY^{+}21]. Amorphous [Bot77]. Amsterdam [Ald72, Bar74e, Flo73, Lan74a, Mul76, Val78, Woo74]. Analgesic [Gar96]. analogical [YHGC20]. analogy [SBB22]. analogy-based [SBB22]. Analyser [Bha88]. Analysers [Gro89, Heu86]. analyses [BN00, BNS18, DZS09, LHB18, PMP^{+}16, vDD11]. Analysing [Hol83, RAN03, VL73]. Analysis [APS95, Aji95, AJT79, CLW90, CG93, DS88, FKV98, Fre78b, GBG^{+}14, GM85b, GS90, Har80c, Har95, HGW94, HJ88a, Hoa73, Hol88, HC93, KLLK98, KMSS98, MTdT93, MW93, MMN79, OW83, PMY97, RS93a, Rey87, RT77, SP88, SB93, SM20, SW91, Set79, SFIK80, ST77, Str95, SO77, TAJ81, WC81, Wai86, WIS^{+}97, WI85, YR92, Yoo96, AAB^{+}21, AKL^{+}09, ALKL19, ARCN^{+}06, ACSK23, dODP21, AZS19, ABA20, BCPL13, BDQR23, BFGS05, BGA20, BLS03, BWA82, BDM16, CW92, CS15, CBF23, CL82, CFC15, DFW^{+}12, DD21, DLWF17, DdB15, DP09, DDD16, DAC06, Ell72, FSRF^{+}22, GC20, GRA14, GZX^{+}21, HBB22, HAM18, HOY17, HCG^{+}16, IASC16, ISUG06, JXG^{+}21, JH03, KSSA23, KW09, KW17, Ker17, Kil19, KAYH^{+}99, KRR19, LCT^{+}21, LCA09, LCC14, LCZ08, LLLY19, LWZ^{+}21, MPC^{+}19, MMM18, MCCJ23]. analysis [MdCGdC^{+}17, NNLR17, NLA15, NZL19, NZTL^{+}23, NEP^{+}17, OY10, Ozk18, PPR^{+}21, hPmKgH15, PLR18, PLPA22, PNP20, Pit82, PVR99, PKvdWB17, QC17, RaMB^{+}24, Rec79, dSRdSS^{+}21, RGS^{+}20b, RJGH06, SDB^{+}22, SD75, SSV^{+}20, SPPH10, SR02, Söz15, SYXZ14, SLJ^{+}18, TK09, TSO19, UT19, WPL^{+}21, XXJS18, YLR23, ZZKA17, Zdu07, ZLTX18, ZCO13, ZWSS15, dOdO16, dAPMV10, vDV04, CF05]. Analysis/Synthesis [WC81]. Analysts [Wil82b]. Analytic [WS74, CP22]. Analytics [Ano13, JPG^{+}17, ANSK16, APR22, BBMN18, dCCCdAC20, JLBR22, NRUP21, SB21, VSID17, WSL^{+}20, YAFA19, YOH15, SZSB19]. Analytics-as-a-service [JPG^{+}17]. analyzability [RW12]. analyze [Cer18]. Analyzer [BF97, BPS00, Fer13, GN16]. Analyzing [dCCCdAC20, JK14, RD14, SDIP23, ACG^{+}21, FCYL18, SMNB21]. Anatomy [Joh84, KKA^{+}16, Val80]. Ancilla [She81b]. Anderson [Ald72, Rop88a]. Andra [GW84b]. Andre [Whi87]. Andrew [Fox79]. Android [DMC17, FB22, FZW19, HYH15, HTWS15, MPC^{+}19, CCK21, MTPC14, NZL19, NZTL^{+}23, RMM19, RMZ17, TSO19, YAVHC21, ZLTX18, ZSH^{+}23]. Angeles [Flo74, Tho74]. Angell [Edm82]. Anger [Bar73b]. Animating [JG89]. Animation [KN88, KS89, WSB96, KPJ^{+}17]. ANN [TMS18]. Annotated [AS78, vdBdJKO00]. annotation [DCEG23]. annotations [KPK^{+}18, WWGP10]. Announcement [Ano78a, Ano95n, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano16a]. Announcements [Ano95n]. Announcing [Ano16a]. Annual [Bar74c, Roh77a]. anomalies [GVG^{+}18, KSSJ^{+}23, DAP21]. Anomaly [CC87, KSSA23, AZS19, GDW^{+}23, JZLP20, LB02, SIC^{+}20, SIK^{+}16]. anonymization [PES^{+}20]. ANSI [BRMO97, FH91a, Ten85]. ANSI-C [BRMO97]. Answering [Har71b]. ant [KSK15]. anti [MV16]. anti-virus [MV16]. Anticipation [VH04]. Anticipation-based [VH04]. antiphishing [MGBO22]. Antivirus [MVTH14]. ANTLR [BP08, PQ95]. ANTLRWorks [BP08]. Apache [BHNDdMRG23, KBB^{+}20, SKI08]. APET [Bai73]. API [BBMG08, BB10, GK08, Hor21, LBP^{+}13, CCK21, RK15a, TWNH12]. API-usage [LBP^{+}13]. APIC [Inc86, Wic72a]. APIs [BBGP01, GLD^{+}21]. APISonar [Hor21]. APL [Ear77, BS74, Dun74, Gel75, Sam75, Str77, Bar71, Tho74].

App [LLLY19, CKRC20, GLD^{+}21, CCK21, NZTL^{+}23, ZGGL23]. append [SH82]. appliance [HKC^{+}12]. applicability [EGL18, Man01]. applicable [Glü12]. Application [AE06a, Bai73, BS88, CG93, CSIL93, DV84, DP85, Ell79a, Fje79, GLW82, Gon87, Gru79, JDJ^{+}06, KVG19, KT84, KS80, LL96, LCC97, MPN^{+}95, MFdlP12, MGS^{+}20, Pei02, Pfe84, Ric76, Sav06, Sre76, TCC^{+}94, WH98, WG92b, Woo84, vHE87, ASP^{+}19, AWNS18, BJ21, BFRFB22, CLC09, CZL21, DDP07, DSD^{+}05, DM15, EKM^{+}99, FRGPLF^{+}12, FRBRF19, FCBF^{+}21, FSRF^{+}22, FFRF19, FFRFS19, GBE^{+}09, GAH05, GB14, HK06a, HBD04, HLFS05, HPZ^{+}20, JLZ09, JSRM18, KGL06, KNT^{+}01, KAS^{+}16, KS10, LTK^{+}20, MAR^{+}16, MN18, MSWC23, NBOS99, NS23, PRTS06, PPSO17, RBB12, RMZ17, SI10, SSS^{+}02, Sny08, UFS99, VS20, VC21, WIYC20, YHGY06, ZB23, ZC03, ZPGHIA18, Dav74]. Application-customized [LCC97]. application-level [GBE^{+}09, MN18]. Applications [Ano13, ABBE98, BP97, BH92, CDG^{+}98, CSIL93, Dew93, Gar96, GH93, HUS^{+}91, HJC05, Hum76, Jaa95a, Kor92, LFW96, LKK19, LK93, Mar86, Mat94b, NHP81, NSM86, RS86, Sco73, TP92, Wai73a, WR95, WW95, Wit82, Yas94, ASA22, AGC10, ACG^{+}21, ALF01, Ano23m, AYdS^{+}06, BMM^{+}18, BFG^{+}11, BBMG08, BDL09, BB10, BDP02, BT21a, BSNB20, BdPGS14, BMM19, BRS18, BC13, BCRF22, BMAV05, CBR10, CNRB13, CGM^{+}03, CMCL03, CRC18, CV03, CPD13, CA18, CGIP15, CSZO23, CWZ17, CP07, CDF^{+}24, CB00b, CD15, CALL18, DDGP18, DP09, DAJ^{+}15, DO07, DM07, ET07, EC13, ESB^{+}17, FDN^{+}18, FJ03, FMC18, FFF^{+}13, FZW19, FPT07, GN00, GB13, GAF^{+}09, GCRD04, GFS^{+}05, GLL20, GLT08, HSRB24, HIR06, Hsu12, HTWS15, HCG^{+}16, IK15, JDPB08, JGSG^{+}21, KKR03, KY05, Kap13]. applications [KGAR18, KRZ02, KSDG24, KHC^{+}19, KKA^{+}17, wKJM18, LLM05, LKCC00, Li18, LHFL07, LD99, LQ99, LTW^{+}21b, MvSdL09, Mal17, Man01, MM02, MZC10, MMCF03, Mau05, MGT20, Mej03, NZL19, NZTL^{+}23, OMM15, PRA^{+}06, PLPA22, PHB21, PKC^{+}13, RJ09, RH78, RBS14, dRRGdC15, RSLAGCLB16, RGC^{+}21, RW17, RAN03, SGA21, SFK^{+}01, DAP21, SAYM24, SMLS23, SGCM11, TGAS22, TSO19, TAG^{+}10, UFR18, WJC^{+}14, WSN^{+}23, WMHL21, WKKK24, XWC^{+}17, XDVB24, YOH15, ZML13, ZKW21, ZLZ^{+}23, ZWC21, ZHO^{+}19, GCK^{+}02, Bar73d, Ear77, For72, Mer74, Nic72]. applications-Review [Ano23m, SGA21]. Applicative [KGP96, Tur79]. Applied [Kuh90, ACF13, CCM05, PGK^{+}10]. Applying [CGP^{+}06, CF05, DFRR15, GZW^{+}22, Hal86, Har84a, MdCGdC^{+}17, ST19, WHS^{+}00, Yi12]. Appraisal [LPT78]. appraiser [JSRM18]. Approach [Aji95, And82a, AZ97b, AZ97a, Bar97, Ber85b, BT76, CSR93, CFL84, CGWL80, Cro91, Ein88, FKV98, GW85, HO91, Hop86, HL94, HKV95, HM84, KFJS88, Knu84, Kop97, LM81a, LES95, MS98, MP82, Mid86, MXYQ86, NSM86, OCH91, STH97, SCGP92, Spo71, Tra79b, WP96, vHE87, APS^{+}11, Add80, AHRR22b, ÁSARSG09, dODP21, AWNS18, BHvR05, BMM^{+}18, BELS14, BB10, BBH^{+}22, BS99b, CCQ16, CMCL03, CCC^{+}16, CA08a, CLD^{+}17, Col77b, CS17, CM08, Cou85a, DTB12, DLWF17, DHA11, DSD^{+}19, FLPM20, FIÁSLSAR05, Flo79, FCBF^{+}21, GAAA22, GFS^{+}05, GMGL24, Gla82, GMDM17, IMG^{+}21, JSPP21, JSRM18, KGL06, KKR03, KH07, KMB^{+}21, KGAR18, KGAR19, KB06, KSKG12, KSK15, LFGCGCRP14, LKCC00, LJL^{+}10, LER17, LQ04, LHC15, LG19, LLH14, LZZZ18, LZWX22, LLW^{+}23, LMPR07, MR07, MMOD16, MvSdL09]. approach [MS22, MDH^{+}13, MGBO22, MS18, MAA^{+}22, MSZ^{+}24, Mus17, NZH20, PSTV10, PPR^{+}21, PKK12, Pit82, PPN^{+}23, Pre85, QTV23, RRK^{+}18, RL14, SIC^{+}20, SDB^{+}22, ST12, SGAS21, SSV^{+}20, SGA20, SYG^{+}18, SRS18, SM18, SBFF24, SZ09, aSZP^{+}16, aSJW^{+}24, TGAS22, TXHL18, TLB^{+}18, TBF^{+}22, UT19, WKG^{+}13, WAH^{+}12, XWC^{+}17, XZW^{+}22, YFC06, ZJY^{+}15, ZWAL22, ZGGL23, ZZ11, dAPMV10, vdWCB17, Ano87a, Eme84, Bar71]. approaches [DBOT23, EYG22, FBMA05, MZC08, NHB23, NRS13, SE11, SSGA20, UFR18, ZSAM23, ZYF20]. approaching [HLH15]. appropriate [CK15, KKPP20]. Approximate [JTU96, OM88, Wri94, ZD95, Cox76, Rön07]. Approximating [LPF^{+}11]. Approximation [Col77c, DCS23]. apps [BFGL20, MPC^{+}19, RMM19, WWCW19, ZLTX18]. AQuoSA [PCML09]. Ara [Pei02]. Arabic [ASAK03, Ber99, Kha86, RS93a, Sha05]. Arabization [ASAQ05]. Arbitrary [Pal80, ST79]. Arcademis [PVBB06]. Architecting [CMCL03, CBB17]. Architectural [CLW90, Ein88, MFdlP12, SDDD10, VS18, ACF13, BAPC23, LHB18, Mal17, Ozk18, PRTS06, RLB^{+}11, SGBR13, SL04, ZHZL22]. Architecture [ACC95, FKV98, GH84, Iza80, KWW81, LA90, MR96, MBGC21, NCFCFV12, SZ88, Spo71, TM95, UT19, VBS22, BKL^{+}02, Bla04, BB99b, BNS18, BR01b, BGG01, CKRC20, DMD^{+}06, DO99, DS09, EMRK20, GCARPC^{+}01, GCF15, GLT08, GVG^{+}18, HJ08, HB11, JPM17, JC19, JSRM18, KRK21, KCG^{+}12, KBPM^{+}20, LJL^{+}10, LHC15, LLJ^{+}23, LGRL08, MA20a, MK18, MGS^{+}20, MBG^{+}00, PRA^{+}06, PCML09, PKK12, PPPP23, PBGM18, PKvdWB17, RR22, Rei99, RBL14b, RSRCGC15, RMSMML^{+}11, SMGMOFM07a, SMGMOFM07b, SRC^{+}18b, SMM13, STT24, STA09, SOKE23, SROV06, SGCM11, TM14, TP03, TVCB15, TBF^{+}22, Vo00, WWJ07, WMJ04, WN06, WYAZ15, ZWQC22, ZSFY05, SM01, ZTC24]. Architectures [AL90, RB89, APS^{+}11, ABG^{+}22, ACF13, ACK^{+}22, BP02, BD14, CRGIP15, GB14, GHC^{+}07, HHMMG12, MVV12, MZ00, MOTG18, NS08, PNM^{+}20, QM13, RBB12, RJGH06, SFK^{+}01, SMR^{+}12, SMNB21, TV09, TVSG21, ZPGHIA18]. Archival [RRP95, JKW74]. Archive [CLLT98]. archived [PB03]. archives [ZPSC07]. Area [Her84, KG95a, LP86, Tag88, BBL02, YWN^{+}00]. areas [DDB^{+}18b]. Argument [Mid86]. Arguments [GF81, KYF^{+}22, FCG83]. Ariadne [MR96]. ARINC [DKM11]. ARINC-653 [DKM11]. Arithmetic [CM82, CT90, Hor78, Mac96a, Nee77b, Nee77c, Ric76, ST79, Van92, Rob79, Win02, Wu99]. Arjuna [LS22]. Arm [KS84, GYCL16, MMHB08]. Arms [Col77b]. Arne [Dav78]. Array [Bel74, Bro86a, Jen89, Kno81, PF97, PS81, SY86, Wet80, WB77, CS15, FK16, GR17, KFMF18, KAO23, MFYiA01, MN80, OAF^{+}03, SS07, YOM^{+}07, ZZJ21]. array-based [ZZJ21]. Arrays [And82a, GS90, BK23, GL05]. Art [Bar76b, Bul72b, Hut76, Bar79b, BFGL20, Llo82, Rog73, Wil74a, Wil76, XZX^{+}21, ZPSH21]. artefacts [DFOT10]. ARTful [SFC^{+}21]. articulation [FSR11]. Artificial [Cho96, GZW^{+}22, QTV23, Cam85]. ARTK [DHGR92]. ARTK-M2 [DHGR92]. ARX [PES^{+}20]. ary [MS98]. ASAB [LYL^{+}03]. ASADAL [KLLK98]. ASADAL/SIM [KLLK98]. ASDL [Han99a]. ask [Ros71]. ASN.1 [Gar96]. ASN EZE [Gar96]. ASP [Eva71]. Aspect [Bar15, DB09, CCC^{+}16, FG14, FSC^{+}21, GC20, LGRL08, MF08, PSD^{+}04, PPSO17]. aspect-based [GC20, PSD^{+}04]. Aspect-oriented [Bar15, DB09, CCC^{+}16, LGRL08, MF08, PPSO17]. AspectJ [PWBK07, SM18]. Aspects [Ano09, AvdSGS80, Bae73, BP09, Jeg83, NPJ79, SS95, CCF^{+}09, DCS23, KSKG12, SGBR13, SBL06]. aspectualizing [POM03]. Assembler [ACDP85, CNG^{+}83, Haz71, HC79, HGWBS75, NSW77]. Assemblers [DS86b, PVS79, Wil79, Bar75b]. assembling [UCCPM19]. Assembly [BW71, Fid82, Hay83, Jon83, PSV85, Wel78a, CMT17, DB09, Ham79, Lev80, Nie79]. Assert [Col88]. Assessing [RLB^{+}11, UFR18, JSRM18, MC23]. Assessment [CGHP79, Fid82, HW80, QC83, TF79a, AL21, CVGCGCFF23, Dod82, HL02b, KKR03, LT83, PPR^{+}21, SM20, Hop74]. Assessments [GRVC23, Liv75]. asset [LTW^{+}21b]. assets [TMGR^{+}22, FLPM20]. Assigning [AJ04]. Assignment [Bla92, BCHS98, BTZ94, Wel83, GHBH05, SBS20, VH04]. assignments [TGPS08]. assimilation [KHS^{+}20]. Assist [CMH91, MM80a, Wil82b, SS19]. assistance [LC12]. Assistant [Ram83, PPN^{+}23, WCsH16]. assistants [PLFIVRAR23, PWI^{+}21, SLBC^{+}23]. assisted [JSC^{+}10, LSF94, RGC^{+}21, TJB^{+}19, YYSG11, ZSL21]. Association [Han76e]. Associations [Han78d]. Associative [SRRFGC^{+}10]. assumption [SC14]. Assurance [LY92, Pra96a, Pra96b, KHMB17, WBN^{+}20]. Assuring [JTG^{+}11]. AsterixDB [KBB^{+}20]. astronomical [KCH08]. asymmetric [HL03, SGS08, Was12]. asymmetry [CYJ^{+}22]. Asynchronous [BMZ92, EBD^{+}74, Geh90, Yoo96, BMC^{+}24, LLJ12, SNL15, dAHCdAC18, CYJ^{+}22]. Atkinson [Bis82]. ATL [TSMGD^{+}11]. ATLAS [PALNGD^{+}06]. Atomic [MDP96, TE90, JEG99, Ren23, ZRX^{+}99]. atomic-broadcast [JEG99]. Attached [Pry85]. Attack [RS21, BSDF20]. attention [LCT^{+}21]. attention-deficit [LCT^{+}21]. Attribute [BV89, BPY90, Fro93, KR83, Pap79, SIN95, DS12, WRD99]. Attribute-based [BPY90]. Attributed [GF84]. attributes [BLC19]. attribution [BUT14, LLJ^{+}23]. auction [BBH^{+}22]. auction-inspired [BBH^{+}22]. auctions [MG09]. AuctionWhisk [BBH^{+}22]. audio [MA01]. audio/video [MA01]. audit [KS10]. Audlib [KS10]. Auerbach [Hop73]. augmentation [SO21]. Augmented [RS93a, Sav06]. August [Val77a]. Aula [SMGMOFM07a, SMGMOFM07b]. Auminaux [Bri82]. AUML [DFRR15]. authenticating [LFGCGCRP14]. Authentication [SW94, SSKG22]. Author [ABC^{+}21, Han79a]. Authoring [STH97, PWI^{+}21, SMGMOFM07a, SMGMOFM07b]. Authorities [HBJ05]. Authorizations [CFL84]. authorship [BUT14]. autism [KRK21]. auto [GJ22, AE06b, AE06a, MPC^{+}19]. Auto-adaptive [AE06b, AE06a]. auto-encoder [GJ22]. autocovariance [BCPSC18]. Automata [Car97, LK93, BT21b, KAO23, LSZ16, Lam20, NNK21, NWE99, NKW06, NK07, Wat04]. automata-driven [NWE99]. Automate [LT85]. Automated [Bur98, HA90, HS97, KBPM^{+}20, LBC^{+}11, MSZ^{+}24, Pet01, SO07, TNGT09, TCMM00, Wat89, ZC02, AWNS18, BCPL13, CBR10, Day00, DS09, FDHH04, GHM^{+}06, LLH14, OMM15, SATV22, XP23]. Automatic [AB95, BPK13, CSD23, CMT02, CMCH92, CA00, DF87, Heu86, HZ95, KL86, KY05, KAS^{+}16, KKPP20, KM94, Kra10, LL96, LD87, LES95, dSMLP22, MP02, MMM18, Mid79, MM86, OW89, RB75, Wal84a, YAVHC21, vdMF13, Bar74c, BFGS05, BRMO97, CDV88, CM08, CA08b, CS04, DE16, DHGR92, GLD^{+}21, GQ15, KHS^{+}24, KMY^{+}05, KCP22, LV01, LER17, LJS20, PTU03, Roh77a, RZ17, SBS13]. Automatically [CZL24, BT07]. automatically-tuned [BT07]. Automating [DAC06, HS85, WZF94, FL02, SSO13]. Automation [Cou92, Lib97a, Lin79, LOS83, Lor91, MN18, LM15, OOG19, PDPMM17, TL14]. automaton [CLS^{+}07, RK15b]. automotive [DHG^{+}19, XWHX21, XZX^{+}21]. Autonomic [GB23, MGT20, DAP21, SGWVP15, TKT^{+}07, BdPGS14, JZ10, KGAR18, SGAS21, SSK^{+}22]. autonomically [PT14]. Autonomous [Cho96, FZ12, ARK21, BHR15, GJ22, MMHB08, NNK21, RBD^{+}24]. autoregressive [XLZ^{+}20]. autoscaling [JJI22]. Autostereograms [Thi96]. AutoToolVCC [XWHX21]. Availability [Hun81, RGS^{+}20b, SAEMM21, DHWZ14, Fra99, KKR03, KS01b, LLH14, MRG^{+}19, dSRdSS^{+}21, SESK23, SC22, SMT^{+}18, Liv75]. availability-aware [SESK23]. Available [FGIS97, Bar74e]. Avatars [Gau95]. Avdeenko [Ano23m]. Ave [Bar82a, Bar82c, Bar84a]. Avenue [Bar78d, Bar84b]. average [XLZ^{+}20]. avionics [WYAZ15, HJ14]. AVOCLOUDY [SAL16]. avoidance [AHRR22a, WCT19]. Avoiding [Rai84]. AVX [CL23]. AVX-512 [CL23]. aware [AFNG20, AO12, BS19, BN00, BMM^{+}18, BSNB20, CLCC15, CSTL19, DTB12, FCYL18, FDN^{+}18, FR09, FFF^{+}13, GDBP23, GWZ^{+}20, HSRB24, HB18, HB11, Hsu12, HC12, HLH15, JSPP21, dSJCM16, KCH07, KSDG24, LLWB14, LCW07, LQS^{+}24, MLR19, MAR^{+}16, MKC20, MF18, OHD22, PKK12, RMdL12, SESK23, SHF16, SGWVP15, SA20, SSK^{+}22, WCT19, WK06a, WSN^{+}23, WTF^{+}22, XZW^{+}22, YRJ18, YZL^{+}22, ZML13, ZSRR22, DGRB15, MSWC23]. awareness [CDRV03, CDF^{+}24, OFRW10, YHGY06, ZXW^{+}17]. away [Bro76, Rob83a]. Awk [Bai85a, Van86, AKW79]. AWT [WWJ07, WW09]. AWT/Swing [WWJ07, WW09]. Axiomatic [Jal87]. Axioms [Pyl80].

B

[Bar74f, Bar77e, Bis84, Con77, Hop73, Hop74, Lav77, Mad82, Pet77, Ree84a, SPHB11]. B.V [Nee77a]. B6700 [Lak80]. Babel [Sco73]. Back [Wil83, Lon88, Rob81, Rus95]. Back-up [Wil83]. backend [BSM^{+}21]. backends [BPK13]. Backtalk [SG93]. Backtrack [Hel95, McG82]. Backtracking [ADS93, KH04]. backup [Fra99]. Backups [Dri93]. Backward [Shr79b, Shr79a]. Bad [KHMB17, Vel85, SBF19]. Bae [XZ01, XZ03]. Baker [Col77b]. BAL [Bar75b]. Balanced [FP82, IC85, ASTW03]. balancer [FB23]. Balancing [HC97b, Rin92, SZ88, ASTW03, BS19, BS85, CPCL10, CSTL19, FB23, HL02a, Kar21, PDPM^{+}16, SJA^{+}04, SA20]. Balaton [Val78]. Balfour [RB82]. Bandwidth [LLWB14]. Bandwidth-aware [LLWB14]. Bank [Sch72]. banking [MVM^{+}22]. Bar [VL73]. Barnes [Bar78b]. Barnett [Bul73]. Barrett [Bis81b]. barriers [DBOT23]. Barron [Atk78, CK13]. Bartee [Bar79a]. Base [Hut79a, Hut79b, KWW81, Ano81n, Flo74, Flo79, Wil74a, Wil76, WCsH16]. base62 [Wu01]. Base64 [ML20]. Based [AM86a, AS97a, ACDP85, AD87, AP84, AP91, BP84a, BP09, BE81, CW91, Cra76, EP79, EV89, HW88, Han90, HL94, IR80, Inc84, KKM80, LCW98, MTT81, MGW82, SZ88, SIN95, SFS97a, SFS97b, SFS97c, UFR18, VSB86, WPT95, Wat89, ACKJ22, ATO10, AF02, AF99, AFF02, AHRR22a, AML20, AÇCT22, AGC10, AMSG23, Al 13, ASEB09, ACF13, ALF01, AFFR08, ASA^{+}21, BS84, BAM^{+}20, BV89, Bar15, BP11, BELS14, BBMG08, BHNDdMRG23, BCL13, Ber20, BSNB20, BPY90, BGP17, BMM19, BNS18, BSC^{+}05, BD14, BCRF22, CLZ99, CDR13, CFLC14, CMT17, dCCCdAC20, CP22, CS18, CCR19, CCT01, CGH^{+}04, CSM^{+}16, CLD^{+}17, CYJ^{+}22, Che22, CI03, CP07, CF05, CCC96, CP96, CBF23, CS17, CW01, CRGIP15, CFD^{+}22, Čuk16, CW08, DFW^{+}12, DLX^{+}22, DD21, DTB12, DGRB15, DT96]. based [DFPT09, DW13, DFQ^{+}22, DBvG^{+}23, DO99, DGPT14, DC15, DE16, DHA11, DHWZ14, DZS09, DS12, Dun93, DFRR15, FB22, FLL^{+}22, FMA02, FKL^{+}13, FG08, FRBRF19, FCBF^{+}21, FSRF^{+}22, FZW19, GH03, GT00, GC20, GLMS18, GR79, GA12, GMNR20, GAAA22, GDW^{+}23, GM22, GAH05, GSR17, GK21, GQ15, GSS^{+}20, GLKZ21, GCZ^{+}24, Har84b, HvdH02, HSC21, HC13, HZWC22, HB18, HGK^{+}19, HMN11, HP11, HLGSW11, HATvdW99, Hsu12, HL20, HKC^{+}12, HM18, HYZ^{+}18, HTX^{+}22, Ier09, Inc85, IHS^{+}14, IS05, IAPC17, Iwa02, IH01, JPM17, JAkM^{+}21, JKH22, JZLP20, JSRM18, KSSA23, KRK21, KJB11, KLY20, KSBL22, KCH08, KMB^{+}21, KSSJ^{+}23, KM13, Kil19, KB06, Kim15, KSH^{+}15, KHH^{+}15, KO91, KW92, KT01b, KPGH02, KCCV05, KEL^{+}21, KSKG12, KIB09, LLJ12, LS03, LKCC00, LBC^{+}11, LSK^{+}18, LG19, LLZ20, LFHL22, LLH14, LC07]. based [LD99, LYX^{+}17, LZZZ18, LJS20, LZLL22, LLJ^{+}23, LLW^{+}23, LQS^{+}24, LQ99, LS97, LvLS84, dSMH13, MUE23, MAT94a, ML08, MS22, MVV12, MDH^{+}13, Mar79, MMM18, MKE18, Mau92, MGT20, MC23, MS18, MVS^{+}18, Mił10, MR05, MSWC23, MKW^{+}22, Mof89, MAJ15, Mos06, MVT^{+}09, MSZ^{+}24, MDB19, MSK^{+}24, MB97, MGJ^{+}24, NZH20, NMMS02, NNL^{+}14, NAU^{+}21, NNLR17, NNR18, NZTL^{+}23, NS01b, NP98, NZPWR22, OHD22, OCH91, OAZ19, PPR^{+}21, PKK12, PRTS06, PPPP23, PNP20, PSD^{+}04, PBGM18, PSRCC02, PGK^{+}10, PHB21, PPN^{+}23, PDPM^{+}16, PP98, RR22, RK89, RCH^{+}23, dRRGdC15, RCA^{+}19, RB19, RGC^{+}21, RZ17, RL23, RS21, ROFGFR^{+}16, ROFGFRM16, RdLFF05, RQL^{+}20, RMSMML^{+}11, RMMLSME14, SIC^{+}20, SDB^{+}22, SO21, SBS20, SCR94, SFK^{+}01, SSV^{+}20, San17, SAYM24, Sav11, SW86a, SK03, SCGP92, SMR^{+}12, SE11]. based [SRRFGC^{+}10, SGA20, SBB22, SBcC07, SAL^{+}04, SWY^{+}21, SMGMOFM07a, SSKG22, SGD05, Smi85, SGDA18, SMNB21, SBFF24, SNK21, SOKE23, SYXZ14, SYB04, aSZP^{+}16, SWBS17, SCLD21, SDC04, TGAS22, TT96, TJB^{+}19, TW16, TY14, TMS18, TKT^{+}07, Tur06, UT19, VH04, VSID17, WP00, WSL^{+}20, WHYC23, Wat04, WBK91, WR22, WKD96, WMHL21, WLS^{+}21, XHP^{+}21, XAN07, XXJS18, XLZ^{+}20, XCZ^{+}21, XDVB24, YZW^{+}12, YAFA19, Yas94, YB06, YFC06, YHK^{+}22, YC16, YLP^{+}11, YSBL22, ZZKA17, ZYW^{+}20, ZZJ21, ZLG08, ZYYC12, ZJY^{+}15, ZZC^{+}17, ZWQC22, ZGGL23, ZZT^{+}21, ZSH^{+}23, Zho03, ZSL21, ZWX^{+}21, ZHZ17, ZQWL24, dAKdGJ11, dAHCdAC18, dOED^{+}20, vdHW03, Fra79, GRVC23, KE85, Lib97b, SMGMOFM07b]. basedalias [IASC16]. baseline [Ber99]. bases [ZB23, Sha83]. basic [Fra99, BL83, Ell82b, Ham77, Heh76, Hop80a, Law78, Orm77, RO77, dB00, Cou84a, Cou84b, Rog74, Bul73, Ree78, Bul72a]. Basics [Key92]. Basis [Lor91, Ano23m, SGA21, vdRW79]. Batch [Gom78, RT77, Coh74]. batch-transmit [Coh74]. batched [WSWR23]. Batching [REC75, SS89]. battery [CCK21, SCLD21]. Bayes [ScG09, ZYW^{+}20]. Bayesian [SIC^{+}20, TRGA18]. Bayesian-network-based [SIC^{+}20]. BBC [RR85]. BCEdge [ZSL21]. BCOOPL [dB00]. BCPL [AC80a, AK83, AJ78, CW82c, Fis84, Lak80, MR80, Ric71, Ric76, Bar81]. BD [DS86a, DS88]. BDDs [CQC98]. be [Bro80, CM96, SCT02, TKF09]. Beach [Rob72]. Beale [Wri98]. Beats [TLB98, THG17]. Becker [Tho74].

Becoming [Kor92, CC02]. Bed [Bat74, DTJ89]. bee [GZW^{+}22]. Been [TB72, Wic81]. Beginning [Cou84a]. Begins [Cas92]. Behavior [CCC96, LB94, DFW^{+}12, GDH13, JAA^{+}20, KKS10, LvDDM06, SZ01, WXR16]. Behavior-based [CCC96, DFW^{+}12]. Behavioral [KKS10]. Behaviour [CC77, Hur80, LA90, Pat94, PS80, QSA88, AFFR08, BE02, KIB09, TKF09]. Behavioural [Bha88, BA98]. behind [CRGIP15]. Belady [Inc86]. belief [SNK21, ZDY^{+}17]. Bell [Ano73a, Val78, Wit83]. Bellcore [Lan90]. Bench [TSZ14]. Benchmark [BHL73, HRW73, TSZ14, HCG^{+}16, KSBW18, MMM18, RMM19, Sta05, Yuv79b]. Benchmarking [Gay80, LNW82, MKD^{+}22, ALG23, DLWF17, RW04, TBT23, ZZKA17, Bry77]. benchmarks [CS03, PF09, AGR^{+}23]. Benefit [Rin92]. Benefits [Com83, FAA^{+}22, MPPC22, Sny08]. Benjamin [Cou85b, Wal84b]. Benjamin/Cummings [Cou85b, Wal84b]. Bent [Ken77]. Benwell [Bry77]. Berkeley [MMS86, PSA87]. Berks [Hut76, Wil74a]. Berlin [Atk79a, Cav83a]. Bernstein [Lav77]. Berztiss [Sha72]. Best [RCMZ13]. Better [CLKG16, JSKB22, ScG09, KL24]. Between [FH74, Gen81, GJ93, Yan91, BBB^{+}22, BFGS05, BRS18, CZ04, CZL24, CDM^{+}16, EYG22, HI85, JB84, KCK^{+}23, KHH^{+}15, LD14, LPF^{+}11, NAGL10, PK04, RBD^{+}24, SXWL17, ZZH^{+}24]. beyond [AGR^{+}23, KL16]. bi [WHYC23]. bi-directional [WHYC23]. Bias [GC84]. Bibliographic [AC75, Lee80, ŠS08]. bibliographical [Jak04]. Bibliography [AS78]. Bid [MG09]. Big [Ano13, APR22, JGB15, QC17, SDG^{+}20, AI22, ACG^{+}21, BCPSC18, CGIP15, CP22, CHC^{+}17, GB23, HTWS15, KHH^{+}15, KKA^{+}16, LHC15, MAW^{+}16, VSID17, WSL^{+}20, XWC^{+}17, XDZ^{+}17, YOH15, AHH15, IAPC17, Kil19, KKA^{+}17, MGS^{+}20, RWJ^{+}17, SZSB19]. big-data [HTWS15]. BigDataSDNSim [ACG^{+}21]. Bilingual [LTL^{+}03]. billions [LB15, LM22]. binaries [MM06]. Binary [AW93, And91, AGG06, BG93, CT92, CG95b, FP82, IC85, Kil81, TD94, AF99, AFF02, ALG23, ASTW03, Ayc15, FHL^{+}18, GYCL16, Kat17, KHOY16, Mau05, MK18]. binding [LB02, NT84]. Bio [ARCN^{+}06]. Bio-Broker [ARCN^{+}06]. biological [ARCN^{+}06]. biology [PD00]. Biomac [HGWBS75]. biomedical [DP09]. BIP [LSK^{+}18]. Birds [Gre80]. Birrell [Gar86]. biseries [JZLP20]. Bit [Sla86, AM10, BLM00, SF85, VED06]. bit-counting [BLM00]. Bit-mapped [Sla86]. Bitmap [PLR85, CLKG16, KL16, PB03]. bitmaps [CLKG16, LSYKK16, LKK^{+}18]. Bitslice [Wit82]. Black [Yuv75, Car22]. Blackboard [DT96]. Blackboard-based [DT96]. Blackwell [Bow88, Rop88a]. Blair [Sau88]. blanks [Fra74]. BLAS [GAQO23, WP05]. Blink [LHGM15]. BLISS [Bre02]. Blit [Car85a, PLR85]. Block [AS97b, GG96, Han81e, HGWBS75, HJ88b, Mar85, Ten82, Wal81b, CPP12, Mor77]. block-sorting [CPP12]. Block-structured [GG96, Mar85, Wal81b]. Blockchain [CFS^{+}22, RS21, SWY^{+}21, FLL^{+}22, HSC21, LLJ^{+}23, LPGBD^{+}19, LTW^{+}21b, OHD22, RGSGHGCG21, WMHL21, XCZ^{+}21, ZWC21, dlRNGCC23, ACKJ22, ZSL21]. Blockchain-based [RS21, SWY^{+}21, FLL^{+}22, LLJ^{+}23, OHD22, WMHL21, XCZ^{+}21, ACKJ22, ZSL21]. blockchains [ASEF23, XHP^{+}21]. blocking [YYLJ24]. Blocks [Shr78]. Bloom [GSR17]. blueprints [BBRB12, GVG^{+}18]. board [MPC^{+}19, SDF^{+}21, VvK99, VC02]. Bolliet [Roh77a]. Bond [Bar75a]. Bonsai [DCW93]. Booch [Wal84b]. Book [AS73, Ald72, And78, Ano73a, Ano79a, Ano87a, Ano88c, Ano88b, Ano88a, Atk78, Atk79a, Atk79b, Atk82b, Atk83, Bar71, Bar72c, Bar72a, Bar72b, Bar73e, Bar73c, Bar73b, Bar73a, Bar73d, Bar74e, Bar74d, Bar74f, Bar74c, Bar75a, Bar75c, Bar75e, Bar75f, Bar75d, Bar75b, Bar76a, Bar76d, Bar76b, Bar76c, Bar76e, Bar77e, Bar77d, Bar77b, Bar77c, Bar78c, Bar78b, Bar78d, Bar79b, Bar79a, Bar80d, Bar80e, Bar81, Bar82b, Bar82a, Bar82c, Bar83a, Bar84b, Bar84a, Bis79a, Bis79b, Bis81a, Bis81b, Bis82, Bis84, Bis86, Bow88, Bra75, Bra80, Bri82, Bry77, Bul72a, Bul72b, Bul73, Bux78, Cam85, CO88, Cav83a, Cla98, Col77b, Con77, Cor82, Cor99a, Cou84a, Cou84b, Cou85a, Cou85b, Dav74, Dav78, Dea86, Ear77, Edm82, Edm86, Edw77, Edw98a, Edw98b, Ell72, Eme84]. Book [Eve73, Fin77, Flo73, Flo74, Flo79, For72, Fox79, Gar86, Gru83, Han72, Han78a, Han78b, Han77a, Haz71, Haz72, Her84, Hop73, Hop74, HW77, How76, Hun72, Hut74, Hut76, Inc86, Jac71, Jac84, Jon74, Ken77, Lan74a, Lan75, Lar71, Lar75a, Lav77, Lav78, Liv75, Llo82, Lon88, Mad82, Mar88, McD71, Mee87, Mer74, Mil72, Mul76, Nee77a, Nic72, Nic98, Pet77, Pit82, Pra96a, Pra96b, Ree78, Ree82, RB82, Ree84b, Ree84a, Ree73, Ree75, Ree76, Rob72, Rob81, Rob82b, Rob82a, Rog71, Rog73, Rog74, Roh77a, Rop88b, Rop88a, Ros74, Sha72, Sha83, Sim83, Sto88, Tho77, Tho74, Val76b, Val76a, Val77a, Val77b, Val78, Val79, Val80, Vel88, Wal83b, Wal81a, Wal82, Wal83c, Wal84b, Wal86b, Wan82, Wel72, Whi87, Wic72a, Wil72, Wil74a, Wil76, Wil84b, Wil87]. Book [Wis74, Woo74, Wri98]. Bookies [TLB98]. Books [Bar73e, Bar75e, RB82, PPBP06]. boolean [Sar77, Dod82, GR73, LM81b, ZZL^{+}21]. Boon [Bar76b, Hut76, Wil74a]. Boosting [YAVHC21, SB21]. boot [DBO^{+}18]. Bootstrap [GLN76]. Bootstrapping [LG73]. borderless [OHD22]. Bornat [Rob81]. both [Pag84]. Bottom [FH91b].

Bottom-up [FH91b]. Bound

[PK89, Wal86a, BM01, GAQO23, KJHG10, SSK^{+}17]. bounded [CPP12, KQZ^{+}11]. bounds [GvRN^{+}11]. Bourne [Jac84]. Bowles [Atk79a]. Box [BS98, SW90, Wit77a, NEFZ00]. Boyer [NT05, Rai92, Smi94]. BPMN [YSBL22]. Braille [ASAK03]. Brailsford [Cor82]. Brain [CHC^{+}17, MBO97, CP22, KCS^{+}20]. Brain-tumor [MBO97]. brainstorming [BDA20]. Branch [KWB^{+}05, BM01, JT00, MMK04, PdSCJM22]. branch-and-cut-and-price [JT00]. Branching [CK86]. Branded [Kot01]. Breach [Bar75c]. Breaking [KP81, Buy21]. Breeze [LHC15]. Brian [Lav78, Wal83c]. Brice [Ree75]. Bridge [HBJ05]. Bridging [CDM^{+}16, MGG^{+}09]. Brinch [Hor07c]. Bringing [BVB^{+}12, GMS20]. BRISK [BMR00]. British [Bar82b]. broadcast [JEG99, MA01, NH03]. broker [AMM10, RCA^{+}19, VNGB08, ARCN^{+}06, CMR07, KNC94]. broker-centric [AMM10]. brokerage [ZPSC07]. brokering [GB14]. Brook [CVV97]. Brooks [Bar76e]. brought [SCT02]. Brown [Lan75, Ree82, Hor07b]. Browser [FSO91, RDM^{+}87, SDKS16]. Browserbite [SDKS16]. browsing [TH01, NEFZ00]. Broy [Sim83]. Bruce [Val76a]. BSD [CV98]. bubbles [RBL14b]. Bucket [CS82]. Buckle [Bar78c]. Buddy [Cha88]. budget [BMAV05, TKF09]. Buffer [KH96, LC03, AGG06, KCH07, UWW^{+}05]. Buffering [Mer73]. Buffers [McC90]. Bug [ACSK23, PMG71, Phi99, SBS20, SO07]. bug-assignment [SBS20]. Bugs [Spa90, JWTG11]. Build [SWA^{+}97, CHT98]. builder [KAS^{+}16, Spi02]. Building [ABSS98, Bro81b, CBR10, CS91a, CMT17, Che04, CFP83, FL92, HBD04, MBO97, NM06, SBG^{+}05, VBH^{+}98, BM98, DPH16, FFF^{+}13, GA12, GF11, HPB^{+}00, PT14, PPR02, RPCS08, WWB03]. Builds [CW97].

Built [FD92, PŻZ13]. Built-ins [FD92]. Bunyan [Wil74a]. Burroughs [Lak80]. Burrows [Abe07, Abe10, Deo00, Deo02, Fen02, NT20]. Burstall [Hun72]. burying [Boy01]. Bus [WIS^{+}97, SNL15]. business [ASC^{+}01, BBMN18, HAM18, JC22, KKR03, LPGBD^{+}19, LTW^{+}21b, LLLY18, DCEG23, MTM22, PCdGPP12, SDIP23, SRRFGC^{+}10, TC03, ZZ11]. busy [TNGT09]. busy-wait [TNGT09]. Butterworths [Bar72b, Rog74]. Buttons [Hes91]. Buyya [Hor14]. BWT [Abe07, Abe10, FBMA05]. BWT-stage [Abe07]. Bycer [Con77]. byte [KL21, Wu99, YAVHC21]. byte-wise [Wu99]. Bytecode [Ler02, BDLM04, BMTA16, CMS07, MJ99, SS09, VDMW06, VB14]. bytes [WL72].

C [Bar73d, Bar74e, Bar75c, Bar75f, Bar76d, Bar76b, Bar77e, Bar77c, Bar79a, Bra80, BDS^{+}92, Ell72, Eve73, Fin77, GR88, Hut76, Jon74, Ken77, KL12, Rob82a, Rog74, Roh77a, SCL00, Val76a, Val78, Wil74a, ZB18, AE14, AM00, AFI98, BN00, Bai85b, BR95, BFGS05, BAFR96, BCV06, BDG93, BRMO97, BC17, Bou91, BB95, BDS^{+}92, CA18, CMCH92, Che04, CCP06, CQH^{+}13, CKW02, Cor88a, Čuk16, Dar00, DH88, DB21b, DP09, DDZ94, Dew87, Eng06, FYP93, FH91a, GM85a, GL05, GR86, Geh90, Geh92, GR92, Gor87, Han04, HM12, HL92, Hel95, Ian90, IASC16, Jaa95a, JM08, JPL03, Kat83a, Kat83b, KH97, KS95, LP83, Lee83, Lev95, Lev97, LS84, Lin98b, MP18, MFH10, Mes96, MSB18, MB97, NSM16, Nar94, NLA15, Nic08, NZPWR22, OM96, PK04, PCBE96]. C [PDC^{+}98, PZ00, PF97, Phi99, PR98, PPA20, Rin07, SH03, SS95, SHF16, Sav07, SATV22, SG97, SB13, SW12, Ste92, SAC^{+}92, Str83a, SB03, TEBK99, THS95, TAAT84, Van92, VP05, WC04, WH98, WW96, YHK^{+}22, ZWSS15, dR86, Ano88b, Ano88a, Mar88].

C# [HP04]. c-mean [ZB18]. C-strider [SHF16]. C. [HM84]. C.U.P [Fox79]. C/C [CCP06, SATV22]. C99 [She07]. CA [NH03]. CA-PK [NH03]. CA3 [JSRM18]. Cache [Dun93, MLR19, Wha93, WH97, CC18, KMB^{+}21]. Cache-aware [MLR19]. Cache-based [Dun93, KMB^{+}21]. Cached [Qui91]. Caching [KH97, LCC97, ASGA23, ADZF21, CLCC15, ET07, LFHL22, MN18, SAC06, WSWR23, SAC06]. CAD [BS90b, GB87, HKB72, Liu03, MR07, WCE^{+}72, Wol92]. CADAM [BS90b]. CADiZ [TM95]. Cadow [Lar71]. Cagan [Flo74]. Calculation [SP88, Vör84, Cox76]. Calculations [Bel74, DR90, RDC93]. Calendar [CSR93, NS23]. Calendars [Gau95, RDC93, UDS^{+}07]. Calendrical [DR90, RDC93]. Calibration [Gom78, Gom82]. California [Flo74]. Call [Ano09, Ano13, BP09, Cor08, CW82b, FS11, GH09, Sta82, Wic77, AGG06, Kan18, KF02, LLW^{+}23, Spi04, TN98]. calling [DDF16, MBV^{+}10]. Calls [CC84, DW91, Er83, FZ98, GG96, Har71b, LQ96, BBG04, Rin07, SNL15, Sto94]. CAM [FPT07]. CAM/DAOP [FPT07]. Cambridge [Atk78, Bar73d, Bar74f, Bar80d, Bar81, Bis81b, Bis84, Eve73, Fin77, Fox79, Gar86, Han78a, Han78b, Lon88, Mad82, Ree78, Sha83, Tho77, Bre82, Col82, LN71, LB81, She81b, VSB86, Wil73]. Camille [BFJ^{+}11]. Campus [EP79, Sno91, NCFCFV12]. can [Bro80, CM96, GC20, SCT02, TKF09]. Canary [TBPK20]. CAP [Her77]. Capabilities [NM78, Rue93, dSCdRS^{+}19, MLC02, PM12, SDIP23]. Capability [CFL84, CL95]. capping [KCP22]. Capture [MPN^{+}95]. carbon [WSN^{+}23]. carbon-aware [WSN^{+}23]. Cards [Coo83, Ler02, VR06]. Carl [Flo74]. Carnahan [Ree75]. carrier [CVGCGCFF23]. carrier-grade [CVGCGCFF23]. Case [Ben89, BTM81, Blu86, Byr91, CFP83, Dew93, DS86b, FIL86, Fle90, Fow90, Fre78a, Geh82, HS89, Hop96, Hop80b, Kat71, Kat83a, KIB09, Lai95, MBO97, MG76, Sal81b, UGBW91, WH97, Zel80, dSC16, AB88, ALG23, ACF13, Ano80b, ADH^{+}00, Atk82a, Bar74i, BLLP04, BTS09, CKRC20, CGH^{+}15, CFD^{+}22, DB09, DMC17, Eba18, Eba20, ENP^{+}24, Fen01b, FC98, FBH^{+}23, GLL20, GMGL24, GK08, GW04, GF78, HP11, HTX^{+}22, HBJ05, JHKS19, Jos79, Jos80, KP94, KCS^{+}20, KRZ02, KSK15, LF82, MS99, Man01, MGG^{+}09, OOG19, OMGDG14, PCdGPP12, PGK^{+}10, Pol01, RBR21, Ree73, RdOTF14, RLB^{+}11, SPPH10, SSS22, SRCP19, Sne78, STT24, UT19, VP05, WHS^{+}00, ZC02, ZRX^{+}99, BBC91, Ber85a, HM82]. Case-based [KIB09]. cases [GRVC23, MMOD16, RMM19]. CASiNO [BKL^{+}02]. Casting [HFPB98, GS06a]. Casual [TS81]. catalog [HC13]. Catalogue [Mus79, VR06]. Catastrophes [Bar84a]. Catching [CWD08, SJP^{+}09]. Categorical [Lin87, LT90]. Caterpillar [LPGBD^{+}19]. Causal [CG93, AKNJ21, MCCJ23]. causes [Eba18, ZHZL22]. Causeway [Čuk16]. causing [Kra10, ZPSH21]. cautionary [TBPK20]. CBack [Hel95]. CCC [KH97]. CCIGS [CT92]. CCL [Mad79]. CCM [DKM11]. CCNPASCAL [NPJ79]. CCR [OM96]. CD [Wai02]. CDC [Bak72, Rob79, Yuv77c]. CDC6000 [FH74]. CDT [Vo97]. CDTC [CZL24]. Celebrating [Buy21]. cell [IAA^{+}21, ZLZ^{+}23, ZRX^{+}99]. cell-centered [ZLZ^{+}23]. cells [SM24]. cellular [HLH15, LHK99]. Center [Pet77, EMRK20, FCYL18, FDN^{+}18, MS22, MRG^{+}19, dSRdSS^{+}21, RGS^{+}20b]. centered [BBS11, ZLZ^{+}23]. centers [ARA18, AHRR22b, ACG^{+}21, ARMMA18, DSD^{+}19, NZH20, PDCB17, RGS^{+}20b, ZWML14]. Centre [Bar72c, Bar83a, BIA24]. centric [AMM10, BGS20, BPP10, CWZ17, DAJ^{+}15, SROV06, TXHL18, WAH^{+}12]. Ceres [Wir90]. Ceres-Net [Wir90]. Certificate [HBJ05, BCL13, BL15, GJ00, LCZ08, RMSMML^{+}11]. certificates [ACKJ22]. Cetratus [MPJ20]. Cfengine [BR97]. CFGs [McK90]. CGAL [FGK^{+}00]. CGLIB [Zho03]. Chae [XZ01, XZ03]. chain [SESK23]. Chaining [WIS^{+}97, SHB19]. Challenges [FS11, BT21a, BFGL20, CEMS24, CBB17, CHC^{+}17, CJTK22, DAC^{+}21, FS13, GdCF^{+}18, HKA12, MF18, PCBR18, PES^{+}20, Sri24, XZX^{+}21, ASA22, KJVS21]. Chameleon [DF15]. Champagne [Ano93a]. Change [Aji95, Car85b, GHM96, GJ93, Als22, CC00, HYC19, Lin98b, MC23, ZNWS18]. change-prone [Lin98b]. changed [CBF23]. changes [CDM^{+}16, HYZ^{+}18, LPP09, CCK21]. Changing [Key92]. channels [DHS01]. Chaos [CCLN22]. Character [GS85, Lib97a, Mei80, Mei81, Par85a]. Character-graphic [Lib97a]. characteristic [ZZL^{+}21]. characteristics [SCLD21]. Characterization [NS74, RBRB23, SSB^{+}16]. Characterizing [VS20, MLV18]. Characters [Wai85, CL23, LM22, Mha05]. charge [TVSG21]. Charging [SCLD21]. Charles [Bar80e]. Charlotte [YF91]. Charlottesville [Liv75]. Charts [HW88, MBV^{+}10, Wor83, ZLWG11]. Cheap [BA81, TKF09]. check [GvRN^{+}11, MAT94a, XWHX21]. checker [Sha05]. Checkers [MM90]. Checking [BS74, BDLM04, CK86, PF97, Rad80, RS94, Ste92, ZZH^{+}24, AMSG23, CCQ16, CGH^{+}15, CCE^{+}21, DLX^{+}22, DS12, GMPL11, GS06b, MHN18, Pet01, PKvdWB17, PD78, Rya80, TVCB15]. checkout [Gla82]. checkpoint [HC99, JKH22]. checkpoint-based [JKH22]. Checkpointing [HC97a, LSF94, AF02, PCL^{+}99]. Checks [Wel78b]. CHEF [MP81]. Chen [Pit82].

Cheong [Her84]. chess [Mes80]. Chichester [Bis82, Bri82, CO88, Cor82, Edm86, Flo79, Lav78, Ree82, Ree84b, Ree84a, Rob82b, Rob82a, Wal82, Wal83c]. Chicon [WLL98]. child [BDD09]. Children [MER84, HBD04, YHGC20]. Chilton [TB72]. Chimera [WG92b]. China [SDF^{+}21, WPL^{+}21]. Chinese [CT92, Gu05, LYL^{+}03, Mei80, Mei81, PZ92, Thi03a, VZ98, WLL98, ZZC^{+}17]. chip [IBA^{+}21, LLJ12, QM13]. chipping [SO07]. choice [Loe07]. Choosing [GKWS11, Gru79]. Christian [Mee87, Ree84b]. Christine [Edw98a, Edw98b]. Christopher [Cav83a]. Chunk [ACC83]. CI [PCBR18]. Ciechanowicz [Ree84a]. Circles [Poo88]. circuit [LM81b]. circuits [Eve73]. Circular [All89, MTM22]. CIRL [CDGP93]. CIRL/PIWI [CDGP93]. cities [JGB15, XWC^{+}17]. city [CWZ17, LYX^{+}17, LZL^{+}17, SRC^{+}18b, XLZ^{+}20]. CL [AV84]. Clarifying [Mog04]. Class [AW93, CK78, GR88, Gor87, Gri86, HS97, HC98, Roh77b, Thi96, Al 13, CZL24, DM11, FGNZ00, HC10, KAS^{+}16, LD99, LLW^{+}23, NS01a, PZ00, SW14, ZJY^{+}15]. ClassBench [HS97]. Classes [Han76d, Str83a, CKB00, CKB01, CKB03, DHS02, Li18, Lin98b, XZ01, XZ03]. Classic [CMH91]. Classification [CT92, CCC96, LPT82, ABC^{+}21, GDW^{+}20, HC13, KCS^{+}20, KSK15, MGBO22, SDB^{+}22, SCD^{+}23, STH^{+}18, ScG09, ZNWS18]. classified [ZZT^{+}21]. classifier [AA20, MPPC22]. Classifying [Wij05]. Clean [Law78]. cleaning [CLC99]. cleanly [CLSE05]. Cleverbyte [Wir77a]. Client [HKM^{+}09, PCBE96, Wid90, ASC^{+}01, GHC^{+}07, LHFL07, Rei99, SFK^{+}01, ÜY22, BGS20]. Client-centric [BGS20]. client-server [LHFL07, Rei99, SFK^{+}01]. clients [CZ04]. Cliffs [Bar73c, Bar74d, Bar75d, Bar75b, Bar76c, Bar80e, Edw77, Ros74]. Climbing [Car22]. Clock [DO07, dCV88]. clone [AML20, LBC^{+}11]. clone-based [LBC^{+}11]. Cloning [RRR97]. closed [GRFFGC^{+}21, SC14]. closed-loop [GRFFGC^{+}21]. closed-world [SC14]. Closure [GL85]. Cloth [Nic72, Bis81b, Con77, Lav77, Lav78]. Cloud [JXG^{+}21, JSRM18, KCS^{+}20, Man18, RCA^{+}19, SWBS17, YAFA19, ARA18, AHRR22b, ACG^{+}21, AA23, ARMMA18, BS19, BGSG20, BSNB20, BGS20, BIA24, CRB^{+}11, CCLN22, CFLC14, CWC^{+}21, CKCG23, CBB17, CCR19, CD15, DGM^{+}22, DCS23, DAC^{+}21, DC15, DSD^{+}19, EMRK20, ESB^{+}17, FCYL18, FDN^{+}18, FCBF^{+}21, FZS^{+}17, GB13, GARSR18, GMGB22, GWZ^{+}20, HB18, HLRVB18, HIRB24, HL20, ISPB21, IB13, IK15, JHKS19, JPM17, JPG^{+}17, KKL17, Kar21, KSBL22, KHGSS12, KSSJ^{+}23, KGAR18, KGAR19, KCG^{+}12, KEL^{+}21, KHC^{+}19, KKA^{+}17, LG19, LLWB14, LYX^{+}17, LWZ^{+}19, LLLY18, MMOD16, MVOD19, MS22, MAG^{+}21, MGT20, MKW^{+}22, MA20b, MKM^{+}17, MOTG18, MRG^{+}19, NB19, NZH20, NNK21, NM19, PDCB17, QRD16, RBL^{+}14a, RGS^{+}20b, DAP21, SM20, SGA20, SRS18, SGDA18, SAA^{+}20, TBT23, TRGA18, TMGR^{+}22, VNLB20, VS20, VBS22, VMY24]. cloud [VSID17, WMSY12, WSL^{+}20, XYW^{+}24, YRJ18, YWT^{+}12, ZWKX17, ZXW^{+}17, ZYYC12, ZDY^{+}17, ZLX^{+}24, ZKB^{+}22b, ZHO^{+}19, ZSRR22, ZPGHIA18, CNRB13, DTB12, GB14, KHGSS12, MST13, MAJ15, RBB12, RWJ^{+}17, SZSB19, VAP^{+}17, XDZ^{+}17, KRK21]. Cloud-aided [KCS^{+}20]. Cloud-based [SWBS17, YAFA19, BSNB20, CCR19, FCBF^{+}21, LYX^{+}17, MKW^{+}22, VSID17, WSL^{+}20]. Cloud-cyber-physical [JXG^{+}21]. cloud-enabled [CBB17]. cloud-fog-edge [GMGB22]. cloud-native [VBS22]. CloudEyes [SWBS17]. cloudlet [MAR^{+}16]. CloudPick [DGRB15]. clouds [CD15, DGRB15, SCF^{+}17, SAL16, SAEMM21, SMLS23, VS18, WSYO11, ZB18, CMF^{+}17, GdCF^{+}18]. CloudSim [CRB^{+}11, HIRB24, JHKS19]. CloudSimSDN [SHB19]. CloudSimSDN-NFV [SHB19]. CloudsStorm [ZHO^{+}19]. CLP [BM01, KMSS98]. Cluster [BB99a, KSH11, RB19, XYW^{+}24, YB06]. clustered [NS08, PDPM^{+}16, PDPMM17, WSL03]. Clustering [PW97, CLC99, DB21a, FG08, HZWC22, MAW^{+}16, NT20, SI10, ST14, WR22]. Clusters [MC91, Buy00, DD21, EGCCM21, HMRZ20, LLS06, LCW07, SAL^{+}04, SC22, ZWKX17, ZLG08]. CMS [ACC83]. CNN [KCS^{+}20]. Co [Ear77, Flo74, Lar75a, Mac96b, Sim83, Val78, Vör84, ABC^{+}21, CHY^{+}22, CBF23, Hor14, ZKZ^{+}21]. co-author [ABC^{+}21]. co-changed [CBF23]. co-design [ZKZ^{+}21]. co-editor [Hor14]. co-editors [CHY^{+}22]. Co-operative [Mac96b]. Co-ordinates [Vör84]. Coal [TBPK20]. Coarse [Wis93]. Coarse-grain-parallel [Wis93]. Cobol [McD71, AJT79, Ano80a, Chv79, FS82, Har83, Jal82, LT83, TT96, TAJ81, Wya84, Ano76a, Pet76, VV06, WB77, Val76a]. Codasyl [Flo79, Ano80a, HT82, Ano76a]. CODD [KM83]. Code [AC80a, AL82, Amm77, AL90, ASME24, Bro72, Bro77, CCM96, CMH91, CH73, CCE^{+}21, Cla89, Cla86, CH90, FH91a, FH91b, GF84, GLS^{+}23, Han83c, Har95, HS85, Inc84, Joh78, Jon83, KP94, KPU04, KG95a, KKM80, LS76, Len90, LKL95, MK96, OMA96, PBW78, Sch89a, Ste80, UFR18, VSM87, WR79, vR92, ATO10, AML20, AL21, ACSK23, AvRAF09, AB20, AG06, BCPL13, BN00, BFGS05, BDLM04, Ber85a, BLS03, BTZ07, BUT14, CQH^{+}13, CZL24, CMM75, CNAM^{+}10, DFQ^{+}22, DC03, DWL^{+}15, EvG04, Eng06, GHBH05, GVG^{+}18, GCZ^{+}24, GLL^{+}24, HTJNL19, HIRB24, HATvdW99, HPZ^{+}20, HYZ^{+}18, HJS^{+}20, HTWS15, JM08, KKN04, LZK^{+}23, LGRL08, LPF^{+}11, MPBH13, MRZ15, MCCJ23, MR05, MK18, MF08, NSW77, PACK07, PKB23, PMP^{+}16, RBR21, RBL14b, RMM19, RMZ17, SDIP23, SO21, SATV22, SD18, SS19]. code [Söz15, Thi03b, TAFC00, WHYC23, WC08, XCG06, ZGG07, ZSH^{+}23, ZYF20, ZWSS15, Hal82, Sch89a]. Code-Based [UFR18]. code-first [MRZ15]. Code-level [CCE^{+}21]. codec [Was12]. coded [Vis76]. CodeQL [YLR23]. codes [Fen02, LQ04, LM06, OG16]. Coding [Con84, Con85, Pla97, DDMD20, FH91b, HC79, HNJ^{+}23, IMBB20, NT20, PD05, Wu99]. CoFeed [FKL^{+}13]. cognitive [GDW^{+}20, PWI^{+}21, Wal83b]. Cohen [Val76a]. Cohesion [RC92, Al 13, CKB00, CKB01, CKB03, RRK^{+}18, XZ01, XZ03]. COIVA [HB11]. Cold [BZD17]. Cole [Han78b]. COLIMATE [SCT02]. Colin [Bar80d, Bar81, Wel72]. collaborating [FZ12]. Collaboration [Bis90, CEMS24, YYLJ24, ZLX^{+}24]. Collaborative [BBB^{+}22, MBO97, ALF01, AGM17, Ber20, BFHR99, BMM19, DFPT09, FKL^{+}13, GH02, HBD04, KPJ^{+}17, LWJ^{+}21, LWW^{+}23, MR07, MCGS08, MMCF03, NM19, OFRW10, PK11, ZKB^{+}22b, dAHCdAC18]. collaborators [CBF23]. Collecting [BCLF^{+}07]. Collection [App89b, Ban71, BW88, BMA72, Chr84, CM96, FH92a, GT87, Nil88, RRR97, Wen90, Zor93, CS02, CS15, Hug82, MKD^{+}22, PDPM^{+}16]. collections [WZH01]. Collector [Ono93a, Wad87, NS01a]. Collins [Hun72]. collision [XAN07]. colony [GZW^{+}22, KSK15, Cho96]. Color [McC90]. Coloring [Duc11]. Colour [Rey87]. column [Bra99, KSSA23, RÁdMRGAM19]. column-gridded [Bra99]. column-oriented [RÁdMRGAM19].

Combination [Qui83, HSC21]. Combinations [WS94b]. Combinator [Har91, vDV04]. Combinatorial [HW94, LES95, JT00, MG09]. Combinatoric [Roh81]. Combinators [Lin87, LT90]. Combine [LLLY19]. Combining [Bud85, LMN91, LLN16, RSLAGCLB16, ZYF20]. coming [Bar82c]. COMLNK [vdBT77]. Command [BBM84, Bud89, Col81, McD87, MD88, Pfe84, SCT02, Wes83, Wil82a, Gai82a, Har82, Mad79, Mau05, SCT02]. Command-line [SCT02, Mau05]. Comment [Gro72a, Rai72, Sam71b, HJS^{+}20]. Comments [Bar74b, BCP71, CKB01, HL79, Pem80, SW74, XZ01, CKB03, Ham79, HLS73, Jos80, LKCW13, XZ03]. commerce [TP03]. Commercial [DJM97, Els76, JLR79, SLJ^{+}18, ACKT20, JDPB08, PVR99, RLB^{+}11]. commit [dSMH13]. commodity [BB99a, DSD^{+}05]. Common [Han04, KH04, McG82, Deo10, ESRI14, Maa06, MK90]. Communicating [Fid88, HD86, HC93, KS86, NAGL10, RS94, RB81, Wre88]. Communication [Ayc15, Bar80c, BMS83, CLKG16, DD90, FIL86, FH74, Han79b, HL98, HW15, KH96, LRMM93, LP86, Mar86, PR90, SG93, Sta82, Str81, WL81a, Wid90, WH84, WG92b, vdBT77, CMR07, DMR^{+}22, DF15, HPB^{+}00, HL02b, HL03, IBA^{+}21, KD13, KSST22, LC05, MR07, MAG^{+}21, MK04, NAU^{+}21, PVBB06, PGK^{+}10, RPCS08, RQL^{+}20, SMKZ06, SHIS99, Sch83b, SM01, VAP^{+}17, WAH^{+}12, vO03, Sau88, Bar73a]. communication-based [PGK^{+}10]. communication-oriented [HPB^{+}00]. Communications [AP91, GK86, KG95a, LBS78, PP80, Rai72, CZ04, LFGCGCRP14, Sam71b]. Community [BB81, CW80, WL81b, AFNG20, DWL^{+}15, GLS^{+}23, SSKG22]. community-based [SSKG22]. Compact [Con84, Con85, DCW93, Han85, JJL17, Jor78, MV20, PM18, Ric79, DGM19, Fra79, OAF^{+}03, SCPL23]. Compacting [CM96]. Compaction [AL90, HR77, LH86, HC87a, Vis76]. Company [Ald72, Cou85b, CW82b, Mul76, Wal81a, Wil84b]. Compaq [MDWD01]. Comparative [TRRK21, WL81a, WW89, Yoo96, HJ14, NEP^{+}17, SH03, ZXX24]. compare [AS08]. Comparing [BUT14, GK08, Lar08, Phi99, vGPB10]. Comparison [BdJ80, CSR93, DP95, DBH04, Fle90, HH79, HZ94, JTU96, LKBT92, LKC12, MM85, Pal74, QK78, SAN^{+}81, Slo93, de 82, ACKT20, Bar15, BFGS05, BLE^{+}08, FBMA05, GMP^{+}21, IS05, RJGH06, She07, Ten85, TCM07, WH06]. Comparisons [Liu86, PK89, Rön07]. Compatibility [DLX^{+}22, Ten78]. Compatible [BP98, MM06, Bar80d]. competent [LBC^{+}11]. compilable [PPA20]. Compilation [AS97a, AP94, CW97, Cro87, Die98, FFW96, Fos86, Gut87, HGW94, HM82, Ono93b, Hop74, KGSC01, LYM04, LCY07, SC14]. Compile [Cor84, Han76a, SGH93, LS15, Sav07]. Compile-time [Cor84, SGH93, LS15]. Compiled [Han79b, MAF91, SD18, vdWCB17]. Compiler [Amm77, BT75, Ber78, BB95, BD76, BP84b, CCRD^{+}80, CAFH94, CMH85, CRT80, CW82c, Far88, Fos86, FH92b, GMO01, Gra92, GH81, GS90, GLN76, Gru79, Gut87, HJ88a, HCD84, HS89, Hut79a, Ise90, Joh78, KH97, LS76, LSF94, MG76, MGW82, Oli83, PKH07, QC83, Rai81, Ree84a, Ric71, REC75, RS76, SIN95, SF88, SFIK80, Ste92, SAC^{+}92, Tse97, UWW^{+}05, Wai85, WG92a, War80, WQ72, WB78, Wir71, YYSG11, AI22, Bar76a, BC17, BRL^{+}15, BPK13, CGR00, DM77, FKR^{+}00, GRVA09, HP04, HKM^{+}09, HW77, JK14, KY77, Kul74, KKdC^{+}23, LvDDM06, LS84, LPF^{+}11, MS83, NBOS99, Pal78b, Sav07, She07, VB14, YC16, SSP11, ZC01, Bar77e, Bar81, Rob82a, Han72, Hop73]. Compiler-assisted [LSF94, YYSG11]. Compiler-Based [MGW82]. Compiler-Compiler [BB95]. compiler-only [KKdC^{+}23]. Compiler-provided [Oli83]. Compilers [Bro80, CLR84, DW89, HR77, LPT78, LHH^{+}91, Pag88, Pro92, PD78, Sco73, Vel85, WC81, WJ76, WB77, WKL76, Dod78, HCG^{+}16, LT83, LMK16, LKK19, PPA20, Ree82, SYXZ14, ZXX24, Rob81, Rob82b]. Compiling [BCP79, Bro76, Dew87, HMS^{+}95, LM81b, MJ99, Mös88, OE92, PJ76, Rob83a, SAC^{+}92, Wal81c, Wei72, LPT78]. Complete [Pag84]. Completely [CLCC15]. completenes [CD84]. Completion [Bla92]. Complex [BH94, Gri82, Lai95, TS91, WA77, WS94b, LMPR07, MvSdL09, MBG19b, SJK^{+}21, TKF09, dAKdGJ11]. Complexity [HG89, HL98, WH98, DBvG^{+}23, Har84a, ML08]. compliance [PKvdWB17]. compliant [BPR01, LK99, MBG^{+}00]. complier [Rei82]. Component [BSNB20, FZW19, LCZ08, Obe11, Sil81, Ste02, BKL^{+}02, BGP17, BCL^{+}06, CMT17, CP07, CRGIP15, DB09, DGR^{+}06, DAC06, DKM11, GH19, HP11, KSSA23, KCH08, KMY^{+}05, KSKG12, LSK^{+}18, ML08, NMMS02, NS01a, PRTS06, POM03, RGV14, RdLFF05, SMR^{+}12, SA02, TMS18, vdHW03]. Component-aware [BSNB20]. Component-based [FZW19, BGP17, CP07, CRGIP15, HP11, KCH08, KSKG12, ML08, NMMS02, PRTS06, RdLFF05, SMR^{+}12, TMS18, vdHW03]. component-oriented [DGR^{+}06]. Components [CS97, CSIL93, FFD96, PW93, ALF01, BHR15, BMSZ17, FT01, GH02, KH18, Lev01, Mau05, Spi02]. compose [RGS^{+}20b, vO03]. Composing [BA98, KPK^{+}18, CV08, RGN^{+}14]. Composite [CSIL93, CS18, ZHZ^{+}14].

CompositeCalls [BJP^{+}00]. Composition [MNM79, GDBP23, GARSR18, GDH13, HBC15, Mal17, Wis74, YHGC20, ZWAL22, ZHZ17]. compositional [Mej03]. compositions [BELS14, BZM^{+}17, GMS20, XLLY19]. Comprehension [STS83, VC21]. Comprehensive [CNG^{+}83, ASA22, GBE^{+}09, GMP^{+}21, KSST22, LQL^{+}22, RMM19, RCMZ13]. Compressed [KL16, ACM^{+}15, Fra06, LSYKK16, NT05]. Compressing [MIA94, ZG06, CHY^{+}22]. Compression [BK93, CW91, CT92, HC98, KPT86, Mof89, VZ98, Yu96, ZM95, Abe07, Abe10, AF99, AFF02, AM10, BGM99, Coo05, CBC00, Deo00, Deo02, Fen02, Fen12, Gu05, HATvdW99, HZ95, LBK16, PKB23, PM18, Ris05, SGD05, SGS08, Sta07, SS09, XWC^{+}17]. compressor [MR04, WSWR23]. Compressors [Fen98, BFNP08]. Computation [Cox85, Far88, LQ93, MV95, Nee77c, VS80, BDG^{+}00, BCPSC18, CCQ16, DFD23, LKK19, LNhCW16, Maa06, Pet01, SGA20, SSGA20, SF88, XCZ^{+}21, dMFÁE17, Bar73a]. Computational [FW78, ALKL19, FGK^{+}00, HHPSS19, MSK^{+}24, SAL^{+}04, TBF^{+}22, ZLZ^{+}23, dOED^{+}20]. Computationally [SV22]. Computations [QSA88, QSA90]. compute [SSK^{+}17]. compute-bound [SSK^{+}17]. Computer [AC80b, Ano71d, Ano71a, Ano71b, Ano71c, Ano72a, Ano72b, AS83, AP84, Ard87, AJ78, Bai73, Bar75c, Bee82, BW71, Bis79b, Bra75, BM72, CGK89, CMF^{+}98, Col87, Cou85a, CB72, DCA82, Ell72, FIL86, FR78, Foo72, Gal79, Gom78, Gom82, Gut87, Haá82, HHK90, Kin71, Lan76, LG73, LPT82, Len90, Les72, LOS83, Liv75, Mor82, NIEN85, NL76, Nut76, Pal79, Pal80, PH84, Pra96a, Pra96b, Pyl72, RS95, Sch78, Sre76, SNM80, Tan73, Tra79a, TV96, Van82, WSB96, WW91, Wir90, WS74, ZZWD93, AIB02, Ano76h, Bar74g, Bar79a, Bar83a, Cav83a, Edm82, Edw98a, Edw98b, EE90, Fel79, For72, Gru83, GF78, Her77, HJC00, Hug77, KRZ02, Lar08, Llo82, MR05, NSKK83, NSW77, Pet77, Pil75, Rei84, Rob72, SM15, Ste79, SYB04, Bar74f, Mad82, Bar73b]. Computer [Dav74, Dav78, Rog73, Val79, Wis74, Wri98, Eme84]. Computer-aided [CGK89, FR78, LPT82, SM15]. computer-based [MR05, SYB04]. Computer-to-Computer [CB72]. computerized [ASAK03, Mos73]. Computers [BS90c, FHJ94, Jal82, JB84, Kil71, Mor82, PBW78, Tho78, WOKT81, WQ72, Bul73, Knu11, LX04, Mer74, RAB^{+}79, Ano73a, Han72, Jon74, Lav78, Tho77, Wil72]. Computing [AC80b, Ans86, AMW91, Bar72c, Bar83a, Bar84b, Bar84a, BS99a, Cho98, EMVW83, JI80, KGP96, Mey78, Pet88, Ree75, SB83, TWNH12, WMG94, ASA22, ASGA23, ASC^{+}01, ARMMA18, ARK21, BB99a, BBL02, Bar78d, Bar82a, Bar82c, BGSG20, BHT^{+}23, BFHR99, BGS20, BC13, BIA24, CRB^{+}11, CNRB13, CVGCGCFF23, CWC^{+}21, CSZO23, CCE99, CHC^{+}17, CPMAH^{+}20, CMR07, DGM^{+}22, DDB^{+}18b, FLPM20, FR09, GB13, GB23, GARSR18, GMGB22, GLL20, GKS^{+}22, GWZ^{+}20, GDGB17, HSRB24, HLLZ21, HB18, HLRVB18, HIR06, HIRB24, HMRZ20, HL20, HC20, HBJ05, ISPB21, IB13, IK15, JAA^{+}20, JGC^{+}21, Kar76, Kar21, KMB^{+}21, KGAR18, KBM02, KDA20, KCP22, KKA^{+}16, KKA^{+}17, LLK04, LG19, LLWB14, Loe07, LZD20, LWW^{+}23, MUE23, MAG^{+}21, MKM^{+}17, MOTG18, NM19, NUB24, PT14, PL08, PGK^{+}10, PMC22, RRFD23, RBB12, RVS^{+}20, Rog74, SGAS21]. computing [SSV^{+}20, SFC^{+}21, SGA20, SSGA20, SGDA18, SHB19, Sri24, SGCM11, SSK^{+}22, TGAS22, TJB^{+}19, TMJ^{+}21, TRGA18, TMGR^{+}22, VNLB20, VS20, VMY24, VP05, WMSY12, WTF^{+}22, XCZ^{+}21, XZW^{+}22, YHGY06, YB06, YRJ18, ZDY^{+}17, ZLZ^{+}19, ZGL^{+}23, ZZT^{+}21, ZSL21, ZGB24, ZSRR22, dlRNGCC23, Col77b, Bar77b, Bar84a, Bul72b, Han78a]. computing-assisted [TJB^{+}19]. computing-based [SSV^{+}20]. CONA [AM78]. Concept [Ans86, Gen81, Pal82, Val84, CY01b, GHBH05]. Concepts [AH85, Bar72a, BY17, ZZL24, vGB01, Rog71]. concern [AKM17]. concern-oriented [AKM17]. concerning [SH82]. Concerns [GL85, CZL21, CEF02, MHN18, ZHZ^{+}14]. concolic [GMDM17]. concrete [MGG^{+}09]. Concurrency [AG95, AZ97b, BS90c, BDS^{+}92, BK87, Cor88a, KT84, Neh79, Rob84, DB21b, SM85, Sto88]. Concurrent [ABBE98, BA81, BNOW92, DS86b, Gai85, Gai86, GC84, GR88, Har85, HP83a, MM97, NPJ79, NW78, Nil90, Ols90, PF97, PR98, SW91, SR91, TBA89, WH84, BMSZ17, CGIP15, Coo04, DIS99, Hay80, Mat80, OW16, SM18, aSZP^{+}16, dB00, BAP87, BK87, CGHP79, DSW82, GR86, GR88, Geh90, GR92, GKLM79, Han76b, Ker82b, Kru82, Rav82, Shr79b, Shr79a, TAAT84]. Concurrently [Har80a]. condition [KWB^{+}05]. Conditional [AG95, CK94, NH03]. Conditioned [WZLN08, FDHH04]. conditions [CSD23, CCPY12, GC20, Mos73, TCMM00]. conduct [LHB18]. conduit [KSK15]. cone [CCQ16]. cone-of-influence [CCQ16]. Conference [Bar75e, BC13, CQH^{+}13, DDF16, DDDF17, DC15, EMD13, FBB^{+}14, GBG^{+}14, GB13, GMDM17, GQ15, HYH15, HCG^{+}16, LSZ16, LMK16, MMOD16, MDH^{+}13, PT14, POZ^{+}16, PDPM^{+}16, PKvdWB17, QM13, QL13, QRD16, SFB13, aSZP^{+}16, Val78, WCK11, AE14, Bar73e, BGS^{+}13, BPK13, DE16, Lan74a, USE85, Val77a, WJC^{+}14, Woo74, Flo73]. conferences [Val77b]. conferencing [CL09]. Confidence [AHRR22a]. confidentiality [FLL^{+}22]. configurability [DHS01]. Configurable [SZ20, CBR10, GRA14, KS10]. configuration [AW04, HLLZ21, KMY^{+}05, SYG^{+}18, SDC04, TKT^{+}07, dAKdGJ11]. configurations [RGS^{+}20b]. configuring [QRD16]. Confined [VB01]. Confirmably [Nee77c]. conflicts [Kra10, RBD^{+}24]. conformance [TVCB15, UT19, ZZH^{+}24]. Congestion [WIS^{+}97]. congruential [SV22]. conjugate [IB13]. conjunction [HOY17]. Connecting [Lib93]. Connection [SW86b, BMR82, LCW07]. Conquer [GM85c]. consensus [DW13, XHP^{+}21]. consensus-based [DW13]. Consequences [Wex81b]. Conservative [Ono93a, Wen90, Zor93]. Considerations [CPHS83, Er85, Mat83b, PS81, Wet80]. Considered [KW90, Vau79, HC20]. Consistency [CK86, CLLT98, GHM96, AA19, CHY^{+}22, IS05, Pet01, SXWL17, XWHX21]. consistent [WW09]. Consistently [LSYKK16]. Consisting [Com82]. ConSIT [FDHH04]. consolidation [ARA18, KS20, KJHG10, NZH20, NTF^{+}17, WCT19, YRJ18, ZSRR22]. consortium [XHP^{+}21]. CONST [MNEM21]. Constant [MV95, MRR^{+}08]. constant-time [MRR^{+}08]. Constant-valence [MV95]. Constantine [GRA14]. Constants [Ber86]. Constrained [Mon96b, BMAV05, EGL18, Ker17, KSST22, PCC^{+}12, SWBS17, dlRNGCC23]. Constraint [BV89, FMT04, KJB11, CFL^{+}98, DDP07, DBvG^{+}23, KAYH^{+}99, LQ99, ST01, TV09, Zho03]. Constraint-based [BV89, KJB11, DBvG^{+}23, Zho03]. Constraints [BA98, LY92, SMFBB93, Van92, AA19, AB20, BGSG20, NZH20, PLR13, RaMB^{+}24, VHM^{+}05, WJC^{+}14]. Constructing [CSIL93, HMS88, HL91, HW98, LLZ20, OG16, XLLY19]. Construction [BCHS98, BK86b, CNG^{+}83, CGWL80, DL23, FGNZ00, Lam81, LS76, PM81, Thi93, WI85, BB03, BST10, Han72, KB06, Nee77a, PTU03, RK15b, Rob82b, SS07, TEBK99, VvK99]. Constructive [Bow88, vHLB^{+}88]. Constructs [Coo96, MS90, Kra10, MGP03]. Consul [MPS93]. consumed [EYG22]. Consumer [MLR19]. consuming [RCMZ13]. Consumption [CP96, DLWF17, ROFGFRM16, WCsH16]. contact [ZYW^{+}20]. Container [RB19, Vo97, WTF^{+}22, PSRCC02, ZLX^{+}24]. Container-based [RB19, PSRCC02]. ContainerCloudSim [PDCB17]. containerized [SDG^{+}20]. containers [PDCB17]. Containing [GH72, Ram96]. content [ASGA23, ABA20, CI03, FIÁSLSAR05, ISUG06, LCW07, Mos06, SS19, UGK^{+}14, VR06]. content-aware [LCW07]. content-based [CI03, Mos06]. Contention [STB14, Smi80, SGWVP15, ZCCB22]. contention-aware [SGWVP15]. contents [ÁSARSG09, BFPAGS^{+}08, WCS^{+}17]. contest [LS03]. Context [AFF02, AP94, GCZ^{+}24, Kea91a, AF99, CPP12, EF13, FFF^{+}13, HIR06, HOY17, HB11, Hsu12, HLH15, MAR^{+}16, MBV^{+}10, SM18, ST19, SYXZ14, TSO19, WC08, ZML13, Rag86]. context-aware [FFF^{+}13, HB11, Hsu12, HLH15, MAR^{+}16]. Context-based [AFF02, GCZ^{+}24, AF99]. Context-sensitive [AP94, EF13, SM18, ST19, SYXZ14, WC08]. context-sensitivity [HOY17]. contexts [DDF16]. contextual [CL82, GDH13, RRK^{+}18]. Contextually [HD86]. Continuation [Slo93, Čuk16]. continuation-based [Čuk16]. Continuations [Cla89, HW94, CA14, GRR06]. continuity [QTV23]. Continuous [ACK^{+}22, Coh98, HC97a, MNH04, MGL19, MBGC21, NB19, PCBR18, RJZ^{+}20, MNEM21]. continuously [OM16]. Contract [Cra77, CLSE05, LWZ^{+}21, Sav06]. contracts [BLS03, DAC06, PRP^{+}23, XP23]. contradiction [BBK^{+}12]. Contributions [Buy21, AB20]. Control [BJ72, Bar75d, BT76, CC84, CK94, CG96, CE97, CK97, Fje79, HKB72, HS83, Inc84, KT84, Lic77, Mat94b, MPN^{+}95, Par75, Ray75, RS93b, Rob84, SDF^{+}21, SL87, Sti79, Thi93, TK72b, Tic85, Web87, WR84, Wol91, AIB02, Ano76c, BMY03, BSC^{+}05, CC00, CA00, DPS03, DB21b, DFRR15, EKM^{+}99, FO10, GT00, HKC^{+}12, HM18, HC20, HYH15, dSJCM16, Lar71, Lev80, MNH04, MLC02, PLL^{+}02, RH78, RAP21, SM85, Sil81, YCY03, vO03, Has77]. control-flow [DB21b]. Controlled [Han79b, NW85, AK15, KAZ13]. Controller [KS84, CGH^{+}15, FVF^{+}18, GMPL11, PT17]. Controlling [SLRS06, ZHO^{+}19, ZCCB22]. Convenient [Mös88]. Conventional [Mid86]. Conventions [Wid90, DC03]. convergence [VRC^{+}06]. Conversational [AM78, AN81, Coh75, Hum76, Rob83b, Wal82]. Conversationally [Ker80]. conversations [PLFIVRAR23]. Conversion [MS90, Par85a, RB75, Sam71a, Sam81, WZF94, CM08, Eng06]. Conversions [WR78, Wol82]. Convert [WR79, WR78]. Converter [MW91]. converters [Pyl84]. Converting [BR88, Mau05, Pag88, Roh77b, Sch89a, LP83, MM06]. convolutional [JP22, WLS^{+}21, YAFA19]. Conway [Fin77]. cooccurrence [SBF19]. Cook [McG89]. cookies [EGL18]. cooling [ARMMA18]. cooperation [CLZ99, FSR11]. Cooperative [CMF^{+}98, LFW96, YH97, GH03, LZL^{+}17, PTU03, YWN^{+}00, dAPMV10]. cooperatively [PT14]. coordinated [ZRX^{+}99]. Coordinating [FT01]. coordination [BMSZ17, DO99, ZCN06]. Coordinator [ABBE98]. COPAS [AN81]. Coping [ZC01]. Copy [SGH93, ML20, Ste98]. Copying [FM86, Ono93a, NS01a]. CORAL66 [Yip82]. Corasick [KAO23, NK07]. CORBA [AKS06, BM03, FJ03, GCARPC^{+}01, HL02a, HLR^{+}03, NMMS02, SFK^{+}01, UFS99]. CORBA-based [SFK^{+}01]. CORDIS [PT00a]. Core [REC75, BOPN12]. CoreASM [FG11]. Cork [JM10]. Coroutine [MR80, PS80]. Coroutines [Bai85b, Fis84, KS80, dR86, AK83, Ger82, HT86, Cav83a]. Corporation [Has77, Bry77]. Correct [All83a, Bor83, Con85, Nee77c, CY01b]. Correcting [Nør91]. Correction [Ano72c, Ano88d, Ano89a, Fen94a, FL76, HF76, KP94, EF13, Han79a, PD05]. Corrections [Mou72, Poo71a]. Correctness [MJ83, Rob84, Sav06]. correlates [PdSCJM22]. correlation [EYG22, IAA^{+}21, JZLP20]. Corrigendum [Ano79b, Ano81a, Ano83a, Ano84a]. CosmOpen [TKF09]. Cost [Bai85b, Del82, HB18, JXG^{+}21, Moh81, Mor82, QSA88, Rin92, WC85, Wai86, Zor93, BMAV05, FCYL18, FVF^{+}18, FIÁSLSAR05, GM22, HLLZ21, KY77, Li22, LCGS17, MVOD19, PF97, SRS18, Wir90, YRJ18, ZXW^{+}17, ASC^{+}01, MKC20]. cost-aware [FCYL18, YRJ18]. cost-effective [FIÁSLSAR05, ZXW^{+}17, ASC^{+}01]. Cost-efficient [HB18]. cost-performance [HLLZ21, Li22]. cost-time [BMAV05]. Cost/Benefit [Rin92]. Costing [Wol82]. Costs [Com83, DDZ94, QSA90, Hat73, KRB21, WP05]. CoT [RWJ^{+}17]. COTS [ZCCB22]. Count [Chr84, Abe07]. countermeasure [DAP21]. countermeasures [FLPM20]. counters [CMPQ23]. counting [BLM00, Hea81]. coupled [AFFR08, AP95]. Coupling [RC92, Str95, RRK^{+}18]. course [Ear77, Fox79]. Courses [vdRW79]. Cover [DDB^{+}18a, LTW^{+}21a, MBG19a, SRC^{+}18a, WY18a, ZKB^{+}22a, Atk78, Fox79, Lon88, Ree78]. coverage [Wij05]. COVID [ACKJ22, GVBG22, KRK21, MVM^{+}22, OHD22, PPPP23, SKD^{+}22, MAA^{+}22]. COVID-19 [ACKJ22, GVBG22, KRK21, MVM^{+}22, OHD22, PPPP23, SKD^{+}22, MAA^{+}22]. CppyABM [NZPWR22]. CPS [YQL22]. CPU [CFLC14, Cro91, JDBP04, TRGA18, WBV96]. crafting [Sav07]. Crash [AM86b, Wal83a]. crawler [BCSV04, GS08]. crawling [UGK^{+}14]. Cray [Fon85, Hus86, Yuv77a]. CRAY-1 [Hus86, Yuv77a]. Create [IC85]. created [Als22]. Creating [Lin86, SA02, Cer18, HC10, WBN^{+}20]. Creation [DV85, Hef82, RA95, Bro72, Bro77, GLD^{+}21, LER17, SJA^{+}11, SBS13]. creativity [VVG^{+}24]. credential [BHW05]. credit [HC20, TRRK21]. credit-considered [HC20]. crevasse [JC19]. CREW [MKC20, CFL^{+}98]. crimsonHex [QL13]. crisis [AKM17]. criteria [MST13, NEP^{+}17, VDMW06, ZZKA17]. criterion [TRGA18]. Critical [BdJ80, REMC81, TB72, BAM^{+}20, DFD23, GEI^{+}11, GVG^{+}18, MMS90, MPJ20, ZRX^{+}99]. Critique [BMM85, Thi99]. CRL [BDL09]. CRMF [LLW^{+}23]. Croatian [DRG11]. Croft [Eme84]. Cross [AS88, ACDP85, BSMV09, HYH15, DM15, DRH^{+}23, LZLL22, LQS^{+}24, MSZ^{+}24, SDKS16, WLTJ13, ZQWL24, SH22]. Cross-Assembler [ACDP85]. cross-browser [SDKS16]. cross-domain [LZLL22]. Cross-layer [HYH15]. cross-organizational [WLTJ13]. cross-platform [DM15, DRH^{+}23, LQS^{+}24, MSZ^{+}24]. Cross-profiling [BSMV09]. cross-project [ZQWL24]. crosscutting [CEF02, SGBR13, ZHZ^{+}14]. crossing [DNL^{+}20]. crosswalk [DNL^{+}20]. crossword [GK08]. CrowdRE [KYF^{+}22]. CrowdRE-VArg [KYF^{+}22]. crowdsourced [DMR^{+}22]. CRT [Coh74, Fra79]. CRT-based [Fra79]. cryptographic [ESRI14]. cryptography [KSST22]. crystal [LFH^{+}22]. CSP [AFFR08, Kou87, OM96, Wre88]. CSP-based [AFFR08]. CSP-i [Wre88]. CTW [HJC00]. CUA [UGBW91]. CUA-2 [UGBW91]. cube [LER17]. CUDA [AGR^{+}23]. CudaFilters [NPHJ18]. cultures [Bar74h]. Cummings [Cou85b, Wal84b]. Cumulative [Fen94b, Fen96, Fen94a, Mof99]. Curracurrong [KAS^{+}14]. Current [AH12, PES^{+}20, QM13]. curve [BG01, KIB09]. Curves [Col83, Fis86b, Gri86, Pal86, WW83, Ano71d, Gol81b]. CustoMalloc [GZ93]. customated [AWNS18]. customer [ASA^{+}21, BBS11]. customer-centered [BBS11]. Customisable [AFI98, BNS18]. customization [HHRS03, WLTJ13]. customized [CV08, JPM23, LCC97]. Customizing [HBM06]. cut [JT00]. CWSH [Wei85]. cxterm [PZ92]. CYBA [Art82]. CYBA-M [Art82]. Cyber [RS21, CWC^{+}21, CEMS24, DLX^{+}22, FBH^{+}23, GLKZ21, JXG^{+}21, JFZ^{+}21, WHH21, XZX^{+}21, YQL22, ZZJ21, ZZL^{+}21, ZKZ^{+}21, LS84]. cyber-physical [CWC^{+}21, DLX^{+}22, FBH^{+}23, GLKZ21, JFZ^{+}21, WHH21, XZX^{+}21, YQL22, ZZJ21, ZZL^{+}21, ZKZ^{+}21]. CYBER76 [AEH76]. cyberattacks [JLBR22]. Cybersecurity [XZX^{+}21, GMC^{+}21]. cycle [LLN16]. cycles [OY10]. Cyclic [Rad80, LD14]. CZT [Mal11].

D [Ano79a, Atk78, Atk79b, Bar76c, Bul72b, Cav83a, Cor82, Cou85b, Ear77, Fin77, For72, Fox79, Gar86, Gru83, Han77a, Ken77, Lav77, McD71, Mer74, Nee77a, Ree78, RB82, Ree73, Sau88, Sim83, Sto88, Whi87, BE02, FMA02, SNL15, Tse97, Wor83]. D-Bus [SNL15]. D-Charts [Wor83]. D.C [Bry77]. D2D [ZSL21]. D2D-assisted [ZSL21]. DaaS [CMF^{+}17]. DACA [HZWC22]. DAG [GNV88, XHP^{+}21]. DAG-based [XHP^{+}21]. D’Agents [GCK^{+}02]. Dahl [Bar75f]. DAI [SG93]. dairy [TJB^{+}19]. DAIS’10 [Kap13]. Dalvik [YC16]. dam [GMPL11]. DAMA [MGGS18]. Daniel [Ell72]. Danies [Rob82a]. DAOP [FPT07]. DAP [RT77]. DARTS [GWA91]. DASD [Ott82]. Data [Abb89, AS97a, ALD^{+}23, AD87, Ano13, Atk77, Bai85a, BCHR81, Ban71, Bar72a, Bot77, BMA72, BSRS85, BY90, Car85b, CC87, CS02, CT92, CJTK22, CK97, Coo86, CW82b, CGWL80, CB72, Des74, Dew91, Dew84, Edw77, Ell79b, Fel81, Fen94b, Fen96, Fle82, FGMM93, GR79, Har80a, Has77, HPC^{+}96, Hut78, Hut79a, Hut79b, Ian90, Inc86, JI21, JG89, Jal87, Kat83a, KS87, KCK^{+}23, KWW81, KG95a, Kow81, KK97, LCT^{+}21, LD87, MTdT93, MW81, Mau92, MS98, Mor80, Nil88, NSM86, O’N88, OPTZ96, PDC^{+}98, PP80, Per85, Pow87, Ree76, RA95, RMC97, SG79, SW86a, Sch76a, Sch72, SL78, SZSB19, Sre76, TB86, Tha84, TS91, Vo97, Wic72a, Wil84a, WR78, WZF94, Yu96, vR92, ARA18, AHRR22b, AI22, ALKL19, ADZF21, ARCN^{+}06, ACG^{+}21, ALG23, Ano81n]. data [App22, ARMMA18, APR22, BGM99, BM06, Bla04, BT21a, BCPSC18, BIA24, CRC18, CGIP15, Cer18, CP22, CLCC15, CCR19, CHC^{+}17, CWZ17, CFS^{+}22, CLC99, Dan82, DLWF17, DKS08, DP09, DHWZ14, DAJ^{+}15, DMC17, DSD^{+}19, EMRK20, Ell72, FCYL18, FLL^{+}22, FDN^{+}18, Fen94a, FCA12, Flo79, FSC08, FLSCC15, GMGB22, GKBK16, GP14, GDW^{+}20, GLKZ21, HM12, HSC21, HL20, HC20, HTX^{+}22, HTWS15, IMKN12, IAPC17, JGB15, JPG^{+}17, JLZ09, KVG19, KHS^{+}20, Kil19, KHH^{+}15, KHS^{+}24, KCCV05, KA87, KKA^{+}16, KKdC^{+}23, LHC15, LWJ^{+}21, MSB20, MVM^{+}22, MS22, MBG19b, MGGS18, MC02, Mof99, MAW^{+}16, MRG^{+}19, MSK^{+}24, NSM^{+}22, NZH20, NSM16, NRUP21, OJP99, PKN^{+}12, PJS^{+}23, PDCB17, PES^{+}20, QC17, Ren23, dSRdSS^{+}21, RT10, RGS^{+}20b, SJK^{+}21, SDG^{+}20, Sha77, Sha83, SWY^{+}21, SCPL23, SRC^{+}18b, SDF^{+}21, SXWL17, aSJW^{+}24, TGAS22, TTC^{+}13, TJB^{+}19, TS02]. data [TBT23, TK09, TCMM00, Vis76, VSID17, WSL^{+}20, WPL^{+}21, WSWR23, Was12, WH06, XWC^{+}17, XXZ13, XDZ^{+}17, YOH15, YZL^{+}22, ZZKA17, ZG06, ZWML14, ZNWS18, ZLY18, ZPSC07, dAHCdAC18, ALKL19, AHH15, Coo85, Hal82, Kil19, KKA^{+}17, MGS^{+}20, RWJ^{+}17, Flo74, Lav77, Sha72, Wil74a, Wil76]. data-analysis [WPL^{+}21]. data-based [IAPC17]. data-centric [CWZ17, DAJ^{+}15]. Data-directed [CGWL80]. data-driven [GDW^{+}20, SRC^{+}18b, TJB^{+}19]. Data-flow [FGMM93, RMC97]. data-handling [ZNWS18]. data-intensive [TGAS22, TBT23]. data-oriented [LHC15]. Data-structuring [Ell79b]. Database [BS81, Bul87, BIO94, CC97, Com82, Fri92, GT92, HHR93, HUS^{+}91, HKV95, HC87b, JKRS85, Joh84, LHS^{+}95, LD87, MTdT93, Mac96b, MRNL92, PSR83, RDC89, SW86a, Sil92, TS81, WOKT81, Wes83, WPN86, WMG94, dV89, BCF00, Bra99, DDPP02, FMA02, LLM05, LK99, LMPR07, MR07, Mes80, PPSO17, PT00a, Rei84, RÁdMRGAM19, TS02, WK06a]. Database-driven [Fri92]. Databases [Clo85, LMN91, MB96, SS93, Sha80, WP96, BHNDdMRG23, CDR13, CKL^{+}02, DS99, FO10, Fra99, LG19, Liu03, PTU03, SBS13]. Dataflow [GS90, OCH91]. Datagrams [LP86]. DataMill [POZ^{+}16]. datasets [BCLF^{+}07, SCPL23, SSS^{+}02]. dataView} [SSS^{+}02]. Datel [Har71b]. DAVE [OF76]. David [Atk83, Han72, Wal86b, CK13, SFS97a]. Day [Bar73d, Bar80d]. DBpedia [hPmKgH15]. Dbxtool [AM86a]. DC [Pet77]. DCAV [MRG^{+}19]. DCE [FJ03]. DDA [Bai85a]. DDS [KCK^{+}23]. deactivation [SSO13]. dead [XCG06]. Deadline [LSAF16, BMAV05]. Deadlock [HJS89, HC93, MFdlP12, DIS99]. Deallocation [AN88, Han90]. Deasington [Vel88]. debt [ACAB22, LQL^{+}22]. Debugger [AM86a, ASH73, Bov87, Car85a, Gai85, Gon87, GWM88, HR96, JKRS85, LF90, SW90, SS94, Smi85, SA97, Ell82b, GIF01, Han99b]. debuggers [WGM08]. Debugging [ADS93, AI80, DR92, Gel75, Gon87, Han78d, HHL84, Joh79, KS87, Lau79, Lea77, LHGM15, LG76, Lop89, MM80a, OCH91, PSV85, Rai73, Sat72, Ste84, Tra79b, WN88, Wit83, ACKS09, Bar76c, Cia07, DAJ^{+}15, GAF^{+}09, IMBB20, KM13, LKC12, NJGG12a, NJGG12b, NJG14, NNW13, PMC05, Tse13, Bar74d]. decade [Sri24]. decades [IMBB20]. December [Rob72]. Decentralized [RS21, FP15, WMHL21, XLLY19, ZWC21]. Decision [Chv83, DW73, Inc81, JI21, Lew83, AKIM24, GH03, HZWC22, KYF^{+}22, MTM22, PPN^{+}23, RCA^{+}19, SSS22, SRS18]. Decision-Making [JI21, AKIM24, KYF^{+}22, RCA^{+}19, SSS22]. decision-tree [PPN^{+}23]. decisions [HBB22, KHGSS12, MST13, SGBR13]. Decks [LS75]. declarations [vdWCB17]. Declarative [ACC95, BM06, EHV99, Fle90, Fos89, YLR23, RPP07]. Declarators [Set81]. declared [Sal81a]. decline [TBPK20]. Decoding [LB15, LM06, LS96b, CWS07, ML20]. Decompilation [CG95b]. decompiler [CQH^{+}13]. Decomposing [MS98, STA09]. Decomposition [CmJHL18, SPHB11, MBG19b, ZTC24]. Decorating [MGP03]. decoupled [LPA13]. DECSystem [GLN76]. DECSystem-10 [GLN76]. Dedicated [SB83, Val84]. deductive [Liu03]. Deep [BGA20, HYC19, JFZ^{+}21, ZZC^{+}17, ASA^{+}21, CBB20, Che22, DNL^{+}20, FB22, GDRV20, MA20a, MSC23, PPR^{+}21, PNK^{+}23, SSV^{+}20, SKD^{+}22, SNK21, ZDY^{+}17, ZZJZ24, KSBL22]. Deep-Q [KSBL22]. deeply [TMJ^{+}21]. DeepPlayer [PNK^{+}23]. defect [GKWS11, GWX^{+}23, ZYW^{+}22, ZZJZ24, ZQWL24]. defects [KRB21, MLV18, PdSCJM22]. defense [BSDF20, HLR20]. defensive [Jos80, Sav04]. deficit [LCT^{+}21]. Define [TDH97]. Defined [RS21, ACG^{+}21, Fis82, LWZ^{+}19, Pyl80, SFC^{+}21, Tur22, Wal81c, YQL22]. Defining [TP92, MTPC14]. Definition [ACDP85, Bai85a, BMC17, BS88, LBS78, SL78, SW74, TS81, dODP21, KA13, SAA^{+}20, ZSH^{+}23, Bar75d]. Definitions [Lor91]. delay [GRFFGC^{+}21, LWW^{+}23, XZW^{+}22]. delay-aware [XZW^{+}22]. delay-sensitive [LWW^{+}23]. Delayed [LQ96, PMG71]. delays [KQZ^{+}11, MKD^{+}22]. deleting [Fra74]. deletion [YOM^{+}07]. Delimiters [STS83]. Delivery [SWN94, JGC^{+}21]. DellEMC [OOG19]. DeltaBlue [SMFBB93, ST01]. deltas [Vis76]. DeltaUp [ST01]. demand [KL24, QM13, SSO13, TW16, WCS^{+}17]. demand-driven [QM13]. demanding [Man01]. Demands [PH84]. Demeter [KHS^{+}24]. Demonstrating [Col79]. Demonstration [Ric76, ZH91, CGR00]. DEMOS [MPP87]. DEMOS/MP [MPP87]. denotational [Lon88]. density [MS18]. dependable [RdLFF05]. Dependence [OE92]. Dependencies [LAD^{+}94, Wal84a, CALL18, LD14, Rai99]. Dependency [LA90, DTB12, LD14, PKvdWB17, TV09]. dependency-aware [DTB12]. deploy [SGCM11]. Deploying [GAAA22, DTB12, KCG^{+}12]. Deployment [SAA^{+}20, AÇCT22, DGRB15, ESB^{+}17, FV03, JSRM18, LTK^{+}20, MKE18, SDG^{+}20, Sav06, VS18, WSYO11, YSBL22]. DEPO [KCP22]. depression [KCS^{+}20]. Depth [Hua87]. Depth-First [Hua87]. Dereferee [AE14]. Derivation [Poo88]. derived [Geh85, GKBK16]. Deriving [AW96, HL98]. Descartes [KU97]. Descent [Kos90, Han85]. describe [GLL^{+}24]. Describing [Mou72, Ros77, AFFR08, RCMZ13, Sch72]. Description [ABBH^{+}79, BNOW92, CCPR91, GHM96, Hef82, Hut79b, Pat94, dSC16, EL05, ZB23]. Descriptions [Pag84, Wat86, LLLY19, WK06a]. Design [ARV77, AL82, AKS06, ASH73, AMW91, AZ97b, BGM99, BCC^{+}22, Bar80c, Bat74, BCL^{+}94, BA86, BS88, Bou71, BSM^{+}21, Bro81a, BP84b, Bud89, BCRF22, Buh93, BDM16, Cel82, CGK89, CW94, CS91b, CVV97, CF05, CDKK85, CPHS83, Col77a, CDH^{+}76, CE84, CK78, DGM80, DPK12, Die97, DO91, Ell82a, FT79a, Fre78a, Fre78b, GOQ16, GM85b, Gom82, GRVC23, Gon87, GT93, Ham84, HRS^{+}09, HS77, HKC^{+}12, Hug79, HP83a, Joh79, JW75, KS98, KCYY12, KMB98, Kim15, KM83, Kin93, KD83, KMY^{+}05, KNPS88, Kou87, Lea82, LFW96, Lei84, LHS^{+}95, LCZ08, LHC97, LQ93, Lor91, Mac77b, MWB95, MC91, Mat83b, Mau92, MM81, MM80b, Mei80, Mei81, MMN79, MOTG18, MW91, MNM79, Mul76, Nar94, NRUP21, NP98, Oes71, PU84, PLPA22, PDPQ22, PS81, PJ75, Pyl72, RS86, Rei99, RH77]. Design [Rob84, SS95, SWN94, Sch76a, SL78, SF98, SM01, SR88, Ste98, TH01, Thi87, TS81, TN98, TCC^{+}94, TAG^{+}10, Wal86b, WWB03, WBK91, Wet80, WS94a, WB78, Wir71, Wir77b, Woo71, ZWML14, vGB01, Al 13, And82b, BH01, Bar76c, Bar77e, Bar78b, Bar15, BMM^{+}18, BP02, BL15, BKP^{+}22, BGG01, CARB10, CMT17, CGH08, CYW^{+}15, CZL24, CLSE05, CdA12, Čuk16, DB09, DC03, DS03, DE16, DAC06, DZS09, DCA04, Eba20, EM12, Eve73, FGK^{+}00, FVF^{+}18, FSR11, FPAF18, GKLM79, Han81a, Har82, HBB22, HE82, Him00, HP11, Inc85, JDGCGA12, JC22, KF02, KA13, LSK^{+}18, LS16, LLJ^{+}23, MLR19, MHM01, MMCF03, MSR^{+}07, MG13, NHB23, NW84, OM16, PPBP06, PLR13, PMC05, PH14, PNM^{+}20, PGK^{+}10, PPN^{+}23, Pur76, RRK^{+}18, Rob82b, Rog71, RGS^{+}20b, RW12, SMKZ06]. design [SL04, Sne78, SBF19, STT24, TL14, Tur22, UFS99, WG04, WHS^{+}00, WYAZ15, XXJS18, YWN^{+}00, YCY03, YZW^{+}12, Zdu07, ZKZ^{+}21, ZRX^{+}99, dAPMV10, CPMAH^{+}20, Sav06, Bar77d, Pit82, Wan82, Jac71]. design-based [Al 13]. design-stage [CGH08]. design-time [CMT17]. Design-view [LFW96]. Designed [HG84, RS87]. designers [LYLY20]. Designing [BMY06, BY17, Cra76, Dew93, FS82, GM77, MER84, SM24, Sef97, SM15, SC90, TGCF08, VNGB08, VL73, WM20, Wal81b, ZML13, AYdS^{+}06, JJK^{+}12, PRTS06, Bar73c]. Designs [SC94, BK23, HL03]. DesignScrum [VVG^{+}24]. DESP [Dar00]. DESP-C [Dar00]. destination [MVS^{+}18]. Destruction [BCHS98]. destructive [Boy01]. Detail [Bul87]. Detailed [SD75, UCCPM19]. detect [MGBO22, MGJ^{+}24]. detectable [Thi12]. detected [TVCB15]. Detecting [JM10, KH18, LKCW13, PRP^{+}23, CDM^{+}16, GWX^{+}23, IASC16, Mha05, Par78, Sco77a, ZPSH21]. Detection [CC87, CL83, Cor08, FYP93, HC93, KW90, KSSJ^{+}23, OF76, RS21, WHLM98, AML20, AL21, ACSK23, AZS19, BBMG08, BTZ07, CP22, Cor84, DIS99, DDD16, FB22, FBB^{+}14, GDW^{+}23, GDW^{+}20, HSY^{+}20, HLH15, HLR20, JZLP20, KSSA23, Kra10, KH04, LMK16, LQL^{+}22, LZK^{+}23, LLN16, MBGC21, OAZ19, RMM19, RW17, SPR^{+}19, SIC^{+}20, SDB^{+}22, SIK^{+}16, ScG09, SSST15, ST19, SWBS17, TNGT09, VV06, WHYC23, XAN07, XXZ13, XLZ^{+}20, YAVHC21, ZSH^{+}23]. detector [SDF^{+}21]. determination [ZJY^{+}15]. Determining [RC92, MMK04]. determinism [Sel75]. Deterministic [PP98, GP01, KM13]. Determinization [LSZ16, Lam20]. detonation [ZGL^{+}23]. DevDocOps [RJZ^{+}20]. develop [CL09, Kim02, Wai02]. developed [PD00, PVR99]. developer [CC02, GLS^{+}23, SROAdM^{+}08]. developers [BMR14, CBF23, CCK21, WBN^{+}20]. Developing [ALF01, BDL^{+}11, BPR01, BFJ^{+}11, BN13, CPZ02, CI03, CR18, DFST08, GK14, GLL20, GB87, HHK90, Iwa02, Jac85, LC05, Man01, Mej03, Mił10, PL91, Poo71b, Sur13, Wai07, ZCO13, BLE^{+}08, CEMS24, GH03, GFS^{+}05, GKS^{+}11, GHC^{+}07, Haf13, LMPR07, TAG^{+}10, WSN^{+}23]. Development [ACC95, Ano87a, AJ78, AP91, BP84a, BE81, Blu86, BSC^{+}05, CC73, CMF^{+}98, CM83, Com79, CP76, DFPT09, DRL82, Dro85b, FR78, FL75b, Gri80, HHZ^{+}95, Haz80, HHMMG12, Jac85, JEG99, Key92, KR85, Lan71, LN71, LL91, LDG^{+}96, LY92, MPP87, NMH^{+}24, PZA87, QC83, Rin84, SCGP92, TLMP93, WA77, Wor83, ACKS09, ABBS22, Ano23m, AGM17, BBB^{+}22, BBMG08, BBS11, BP08, BV06, Che22, CCE^{+}21, CSS15, DDMD20, DGPT14, DM15, DFRR15, FRGPLF^{+}12, FSR11, FT01, FPT07, GRVC23, HPP22, JDGCGA12, Kar14, MvSdL09, MPPC22, MGL19, MVS^{+}18, MSZ^{+}24, NNL^{+}14, NDKL^{+}24, NW84, Pal78b, PVAHRG^{+}15, PVBB06, PW11, RBB12, RLB^{+}11, RdLFF05, SDIP23, SGA21, SSCdA^{+}03, SN01, STH^{+}18, SR02, SZ09, TWJ^{+}13, WWJ07, WWCW19, WP05, WKG^{+}13, XCL^{+}18, ZSAM23, ZC03, GH09, Ano76a, GH11]. Development [Gar86, Bow88, Ano81n]. Developments [Ray75, SRCP19, Her84]. Device [CF80, DMC17, HMRZ20, KBPM^{+}20, MM06, MGGS18, WCS^{+}17]. Devices [GF80, BCC^{+}22, BBMG08, CC01, CSM^{+}16, EGL18, IBA^{+}21, JSPP21, KY05, KJVS21, KSST22, LCT^{+}21, LC07, Lv21, PCC^{+}12, RMZ17, RMdL12, SWBS17, XMTL21]. DeViouS [RS95]. DevOps [AKIM24, FAA^{+}22, FBH^{+}23, RJZ^{+}20, ZHO^{+}19]. DEVS [Wai02]. Dew [HMRZ20]. DewSim [HMRZ20]. Dfl [Bar76a, Bar77e, Mul76]. diagnosis [GSPA^{+}11, JFZ^{+}21, PDPMM17, RW17]. Diagnostic [Gri75, HA72, HR77, CLS^{+}07, PPPP23]. Diagnostics [WB85a, WB85b, AE14, MPC^{+}19]. Diagram [BH94, SS93, GHC^{+}07, KAS^{+}16]. Diagrams [CCvKH95, FGMM93, KM94, Lan82, Thi97, CGH08, CmJHL18, CZL24, DE16, SW14, aSZP^{+}16]. DIALOG [NHP81]. Dialogue [AS83, KS82, Pfe84]. dialogues [BB99b]. DiaSim [BC13]. Dickson [Lav77]. dictations [TC07]. dictionaries [KFMF18]. Dictionary [CS82, LD87, BGA20, GJ22, Ris05, Rön07, SGD05]. dictionary-based [SGD05]. diff [CFD^{+}22]. Difference [GH72, LA11]. differences [Yan91]. Different [QK78, WW89, DM07, KY05]. Differential [Dun93, McK99, BK23, WLZ^{+}22]. differentiation [BRMO97]. DigiHome [RHT^{+}13]. Digital [Bar75c, BCC^{+}22, BLC19, BFPAGS^{+}08, BDMP17, BPP10, CR18, Eve73, GMGL24, Han72, SAYM24, SAY16, ZZKA17, Bar79a, Ree75]. Dijkstra [Bar75f]. DIKE [PTU03]. DILAF [AZS19]. Dimension [KK90]. Dimensional [BS88, MTT83, Wit77b, DW90, Gut76, LLJ12]. dimensionality [CJTK22]. Dimensions [Lyo85, Pet01, SBFF24, vD99]. Dining [Car82]. Direct [Coh73, Cow87, LZLL22, SY79, CZ04, CYJ^{+}22, Fra06, LKK19, PP84]. Direct-Memory-Access [Coh73]. Directed [All83b, RDM^{+}87, CGWL80, DB21a, FL76, FR91, GNV88, GJ00, GG08, HW88, KPT86, KU97, Nil90, PL91, SK96, aSJW^{+}24, Thi03a, WG83]. Directing [Sos95]. direction [WBB15]. directional [WHYC23]. directions [ASA22, BIA24, CKCG23, CJTK22, GKS^{+}22, HSRB24, MFB^{+}02, RB19, Sri24]. DirectJ [BBGP01]. directories [LAG00]. Directory [Han80a, Bar83a]. Dirty [Coo86]. Disassembler [DB83]. disaster [WR22]. Discipline [BS84, Nee76, Vo00]. disclosure [FO10]. disconnected [Tor23]. discover [EMD13]. discoverability [MRZ15]. Discovering [CT90, DS99, Kot96, RCMZ13]. discovery [AMM10, FZ12, HYT13, MCGS08, NEP^{+}17, WSWR23, XDZ^{+}17]. Discrete [GHM96, Hać84, Ols90, She75, Bru84, DPH16, DDP07, Dar00, DDDF17, MM02, RMMH^{+}23, The77, WM20, WW00]. discrete-event [Dar00, WM20]. discriminative [BGA20, GJ22]. Discussion [Nee77b]. disease [WHH21]. Disk [Han76a, QK78, TTH97, VC90, CLCC15, DD18, HC16]. disk-aware [CLCC15]. Diskless [Lob85]. disks [CLCC15]. dismantling [LW14]. disorder [LCT^{+}21]. dispatch [BCV06, SM15]. dispatchers [CV08]. dispatching [TEBK99]. dispensers [SAYM24]. Display [CF80, HKB72, Ham84, Jon71, LES95, Mac77b, VR06]. Displaying [EL96, Gri86]. Displays [Dew91, Dun93, Les72, Sla86]. dispute [LKCW13]. dissemination [BT21a, LZWX22, YZL^{+}22]. Dissimilar [FH74]. Dissimilarity [HTX^{+}22]. Dissimilarity-based [HTX^{+}22]. Distance [Ans86, Bur16, CIM22, CIM24, GK21, JZLP20, TC07]. distasteful [Spi76]. distinguished [Bis19]. Distribute [KG95a, DSD^{+}05]. Distributed [ARS^{+}94, AS97a, AP95, BBC91, BS85, BL85, BL90a, Bar83b, Ben90, BP97, Bro86a, Buh93, BR97, Car82, CS91b, CE84, DFD23, DR92, FP97, FHJ94, FGIS97, Gra96, GZX^{+}21, HJS89, Han87a, HMS88, HMS^{+}95, Jeg83, Kap13, KDP83, KNC94, LRMM93, LGC84, LLW98, LKBT92, LT91, LOBF88, MWB95, MCG^{+}88, MMS86, MPP87, MS96, Pet88, PŻZ13, RK91, Ram83, RB91, RA95, RS95, SZ88, SF98, Sha80, She81b, SS94, Sil92, SS89, SY86, TKWW85, TAAT84, TH86, TLMP93, Val84, WSB96, Whi83, Wis93, WMG94, WZF94, YSM95, YH97, ZZWD93, vRvST89, ACV10, And82b, AIB02, ASC^{+}01, AZS19, BMY03, BBL02, BMR00, BVGVEA11, BHR15, BDP02, BFHR99, Bla04, BCSV04, BdPGS14, CL09, CDRV03, CMCL03, CCCZ05, CGH^{+}04, CDF^{+}24, CMTCC^{+}17, CB00b, DD21, DO07, ET07, ESB^{+}17, FT01]. distributed [FMPR02, GH03, GMO01, GAF^{+}09, GFS^{+}05, GMDM17, GLKZ21, HTJNL19, HJC05, HHRS03, HMN11, HYT13, IHS^{+}14, IB13, IH01, JJK^{+}12, JKH22, KS01a, KRZ02, KSBW18, KMY^{+}05, KPGH02, KBM02, LLK04, LQ04, MUE23, MA20a, MK04, MMHB08, MZ00, MC02, NRS13, Obe11, OFRW10, PLL^{+}02, PSD^{+}04, PALNGD^{+}06, POZ^{+}16, PDBG10, QC17, RJGH06, STB14, SESK23, SLRS06, SIK^{+}16, SYB04, TN98, TKT^{+}07, VSID17, WLTJ13, XZX^{+}21, YZW^{+}12, ZWKX17, ZB18, ZLY18, GWA91, HZWC22, IS05, CO88, Sha83]. Distributed-memory [Gra96, HMS^{+}95]. Distributing [BAP87, CFL84, Wai75, WL81b]. Distribution [Fje79, SBL06, Yu96, CNAM^{+}10, HSY^{+}20, Li18, LLS06, LCW07, PJS^{+}23]. ditroff [Ber99, AB89]. ditroff/ffortid [Ber99]. Diversions [WBS82]. diversity [GBG^{+}14]. Divide [GM85c]. DIVINE [WN06]. divisible [IK15, LLWB14]. Division [Han94a, Han95, BLM00, Fro81, Jam80, Wic79]. Dix [Wri98]. DjangoChecker [ST19]. DJM [LLW98]. DLL [BB10]. DM21.40 [Atk79a]. DM29.50 [Cav83a]. DMERT [Wal83a]. DML [HT82]. DMT [ZH91]. DNA [AAB^{+}21, MR04, TP97]. do [CCK21, NHTT08, CA86]. DO-loops [CA86]. Docker [BSNB20, BNS18, BRS18]. Document [BPP10, CDH^{+}76, GW84b, HSM81, HCC96, Kin93, WBS82, Wol91, CFD^{+}22, LTL^{+}03, MS22, YLM^{+}05]. document-based [CFD^{+}22]. Document-centric [BPP10]. Documentation [BA86, Bro86b, CV84, Flo72, FF80, Kat71, NL76, OF76, Rag86, Sco77b, SWBT86, Hug77, RJZ^{+}20]. documenting [HK06a]. Documents [AB89, Cho98, CH06, FKD14, Iwa02, KL24, ZZLL24]. Dodo [Atk79d]. Does [BS74, HJS^{+}20, Str77, Bro82]. DOLAR [SSD11]. Dolotta [Lav77]. DOM [ÁSARSG09]. Domain [Iza80, Lea77, LCW98, LJS20, MBO97, BFG^{+}11, BFRFB22, CIM22, CA08b, FBH^{+}23, HOY17, JFZ^{+}21, LLLY19, LZLL22, LQ99, MPBH13, OJP99, SZ09, WGM08, WAH^{+}12, ZZL24, ZZH^{+}24]. Domain-independent [LCW98]. Domain-specific [LJS20, BFG^{+}11, BFRFB22, MPBH13, SZ09, WGM08, WAH^{+}12]. Domains [SHC74, CFC15]. DOMICO [ZZH^{+}24]. dominated [HKW77]. Donald [Llo82]. Doo [XZ01, XZ03]. Doo-Hwan [XZ03]. Dora [Wit77a]. Dorn [Ree73]. DOS/VSE [Ott82]. DOSE [KFJS88]. dot [Kha86]. Double [BCV06, KFMF18, KAO23, MFYiA01, OAF^{+}03, YOM^{+}07]. double-array [KFMF18, KAO23, MFYiA01, OAF^{+}03]. Douglas [Hor07a, VWB91]. Down [Lei84, Inc83, Rai84, Set79]. downtime [MPJ20]. Draffan [Sha83]. Draft [ABBH^{+}79]. Dragon [Gut87]. Dragonmail [Pet88]. DrawCAD [Liu03]. Drawing [BJL06, DDPP02, FR91, Hop71, Lan82, Pal86, Thi97, Thi96, vdP14, EBFK10, Ple99, VDG^{+}00]. Drawings [Geo77]. draws [GNV88]. drift [RLB^{+}11]. DRIP [ZZLL24]. drive [LFHL22]. Driven [UFR18, AA19, AMM10, AGRS11, BDMP17, CCC^{+}16, CM08, DB21a, DHS02, ENP^{+}24, FBLS12, Fri92, GMGB22, GLD^{+}21, GDW^{+}20, HMRZ20, JJK^{+}12, LT83, LJL^{+}10, LQS^{+}24, LGP^{+}11, LTW^{+}21b, MPPC22, MT94, MMCF03, MC23, MZ00, MGG^{+}09, MVS^{+}18, Mus17, NWE99, NZL19, QM13, RBR21, RRFD23, ST12, SNL15, SRC^{+}18b, TJB^{+}19, TL14, WLTJ13, WGM08, YB06, ZC03, ZZ11, FCBF^{+}21]. Driver [CF80, MK03]. drivers [MM06]. driving [TVSG21]. DRM [WCS^{+}17]. drone [CVGCGCFF23]. Drones [ZWQC22]. drought [ZLY18]. Drug [IAA^{+}21]. DSL [DBvG^{+}23]. DSM [KMB98, LLS06, NS01b]. DSOS [Fra75]. DSP [WJC^{+}14]. DTL [HP83a, HP83b]. Dual [CBF23, MS80a, WHH21, Web87]. Dual-grained [WHH21]. Dual-processor [Web87]. Duality [SMR93]. due [RBD^{+}24]. Dumb [McC90]. Dump [MM80a, NY78]. duplication [MK18]. during [ACCD01, JK14, MVTH14, ZHO^{+}19]. Dyadic [Fis82]. Dynamic [APS95, ADS93, Bro81b, CC87, Cro87, Des74, Dun91, FM86, GM85a, GT93, HK06b, HO91, IM93, JDBP04, KCH07, LH82, LGP^{+}11, PKB23, RT77, SG93, SM90, Sha78, SWA^{+}75, SM18, TAJ81, Whi83, ZYW^{+}20, ZPGHIA18, ARMMA18, Ber99, BGP17, BPS00, CFLC14, CSMML12, CALL18, DTJ89, EGCCM21, FHL^{+}18, GOQ16, GLL20, GS06a, GG08, GQ15, HJC05, HB18, JZ02, KFMF18, KGSC01, KCP22, LC05, LV20, MM02, MV20, MRR^{+}08, NB19, NZH20, NJGG12a, NJGG12b, NJG14, OJP99, OMGDG14, PSD^{+}04, RAP21, RGV14, Sav11, SI10, Sto05, TKF09, TNGT09, WXR16, XXJS18, YYSG11, ZML13]. dynamic-reconfigurable [LC05]. Dynamically [HH88, MW81, KCK^{+}23, PKK12, RGN^{+}14]. Dynamics [GDW^{+}23, LKCW13, LFH^{+}22, ZLZ^{+}23]. Dynamo [YWN^{+}00]. DYNIX [Bad98]. e-Aula [SMGMOFM07a, SMGMOFM07b]. e-business [KKR03]. e-government [PCdGPP12]. e-mail [BS99b, SN07, Kor92, HL94]. e-Scientists [BSC^{+}05]. E-whiteboard [CGH08]. E12.50 [Bis81b]. E7 [Fin77]. E7-95 [Fin77]. Eagle [MKC20]. earliest [KSBL22]. Early [BL90b, Han99a, CGH08, FMC18]. EASE [LL96]. Easily [LV20]. Easy [BF80, Car98, FGIS97, Wal86a, MP13, PD00, Rog23, SV22, Val76b]. EasyLocal [DS03]. Ebert [Wal81a]. ebIOP [TC03]. EC [Kat83b]. echo [ZWX^{+}21]. Eclipse [Hal82, SWPS89, SR91, ACF13, Coo85, GRA14, MSB^{+}13, Sur13]. Economic [Wel78b, CC01]. economical [HK84a]. economics [For72]. economy [MTM22, SAL^{+}04]. economy-based [SAL^{+}04]. ecosystems [ASP^{+}19, APR22, MGS^{+}20, dOED^{+}20]. Ed [Ald72, Bar72a, Bar74d, Bar76b, Bul72b, Jon74, Lan74a, Rob72, Wic72a, Wil72, Wil74a, Woo74, Hop73, Pit82, Rop88a]. EDDIE [TLB98]. Edge [ZKW21, ASA22, ASGA23, CVGCGCFF23, CWC^{+}21, CPMAH^{+}20, DGM^{+}22, DDB^{+}18b, GMGB22, GWZ^{+}20, HLLZ21, HC20, IBA^{+}21, ISPB21, JAA^{+}20, KJVS21, Lv21, PMC22, RRFD23, RVS^{+}20, RQL^{+}20, SSV^{+}20, SGA20, SSGA20, SMLS23, SHB19, TBT23, TMJ^{+}21, VS20, XLZ^{+}20, XCZ^{+}21, XZW^{+}22, ZLZ^{+}19, ZLZ^{+}21, ZLX^{+}24, ZZT^{+}21, ZSL21, ZGB24, PT90, GDGB17, JAA^{+}20]. Edge-adaptable [ZKW21]. edge-based [RQL^{+}20, XLZ^{+}20]. edge-cloud [CWC^{+}21, ZLX^{+}24]. EDI [LW04]. Edinburgh [Hun72, ACG78]. EDISON [ALKL19, DMW88, Han81a, Han81b, Han81c, KS84]. EDISON-DATA [ALKL19]. Edison-N [DMW88]. edit [TC07, HS77]. Edited [Bux78, Hut76, Liv75, Pra96a, Pra96b, Val78, Wal81a, Bry77, Han77a, Pet77, Roh77a, Val77a]. Editing [All83b, Car81, Lev83, Poo71a, SK96, Wol91, GHC^{+}07, Lev82b, NM19, Sne78]. Edition [Mad82, PR90, Cam85, Ken77, Llo82, Ree76, Wil87]. Editor [Bar77f, BH94, Bou71, BHZ85, CDH^{+}76, Ell82a, Fra80, Fra82, HW88, Haz74, Haz80, Hol89, HH82, KFJS88, KU97, Koo87, MP81, Mac77b, MT78, PT90, PM81, Pik87, TK72a, WBK91, AP85, Bar77a, BFJ^{+}11, Bro75, Bro78, Car79, Col72a, Ehr73, FS73, FC83, Fra79, Gos81, Han87c, Har77, Him00, Hor81, Hor14, JCL85, Lea81, LDH92, Mit73, MW82, MIR78, MTRC83, NL75, NM77, Pat83, PK82, Rai72, RR82, Sam71b, She81a, SFB13, ŠS08, SW82, Sur13, Vil80, Wag78, Wex75, Wex78, Wex81a, WKG^{+}13, Wul75, Ano80b, Bar73f, Bis80, BH94, Bud86, Dan82, Gal79, GW84a, HR90, Hay80, Her77, Jam80, JP79, Jos80, Lin98a, Mal80, Nie79, Rec79, Rei82, Rya80, SF88, Ste79]. Editorial [AE06b, AE06a, Ano71e, Ano71f, Ano75a, Ano76b, Ano89b, Bar74h, Bar74g, Bar84c, BP11, BN13, CC90, CM98a, CWZ17, CM05, DRZ13, FHB02, FS13, GK14, GCM11, Glu74, GH11, Gro72b, Gue03, Hal71, Han81d, Han84, Han88, Hoa72, HW10b, Hor12a, Hor12b, Hor14, HC00, Kap13, Ken90, KH12, KSRR17, Kri90, Kri04, Lam72, Lan74b, Lan76, LM02, Nee75, NL01, PL14, RBB12, RBL^{+}14a, RWJ^{+}17, Ros71, Rus95, SFB13, Tse13, TGC15, WW00, WCK11, Wir72, Wir77a, WL03, WK06b, YOH15, Zam03, ZWC21, Ano16a, CM98b, D’A73, SB22, Wai73b, HW10a]. Editorials [CJ73, D’A73, GF11, Hat73, HW10a, Obe11, Wai73b]. Editors [Dan90, Dav82, KW92, Sco81, CHY^{+}22, CW01, CL81]. Edn [Cou84a, Bar75a, Bra80, Bul72a, McD71]. Eds [For72, Sha83, Wil72, Ree84a, Sim83]. Education [Cou92, SWN94, PR16, dCGG13]. Educational [Joh84, RB82, RSRCGC15, VVG^{+}24, YMY17]. Eductive [DW90]. EEG [KCS^{+}20]. EESSI [DRH^{+}23]. Effect [Gai86, PMG71, STS83, WHLM98, CIM22, CIM24, IAA^{+}21, Mha05]. Effective [AJ78, AG06, BMR14, FDD20, NHP81, SGS08, ZYF20, AC13, ASC^{+}01, CYW^{+}15, FIÁSLSAR05, MAJ15, NT20, PDPMM17, UGK^{+}14, WKJ15, YYLJ24, ZXW^{+}17]. Effectively [UW99, SZ01, UWW^{+}05]. Effectiveness [How78, JDPB08, WHLM98, AvRAF09]. Effects [Thi93, Zel77, MM86]. Efficiency [Coh73, DCS23, Lin87, Str81, WW96, ADZF21, Bar76c, BIA24, FCR^{+}09, IMG^{+}21, PDROFRM13, TBSI18]. Efficient [AC75, AMSG23, AN95, AMS92, BT21b, Bot77, BTZ07, CK97, Dan90, DS94, FHL^{+}18, GKS03, GNSP12, Gro89, Gro90, GZ93, HA90, IC85, dSJCM16, Knu92b, KK97, LYM04, LCW07, LLN16, MVOD19, NWE99, NK07, PACK07, PHB21, PR98, QH21, Qui83, RR05, SC14, TD94, UN19, Vo96, YZL^{+}22, vdBdJKO00, AC80a, AÇCT22, AA23, AGG06, BHvR05, BHK^{+}04, Bra99, CCQ16, CK99, CKL^{+}02, CP22, CNAM^{+}10, CV08, EGKP02, FVF^{+}18, FDD20, FK16, Gai82a, GOQ16, GLKZ21, HB18, HPK^{+}12, IBA^{+}21, IIL17, JAkM^{+}21, Kar21, KHS^{+}20, KMB98, Ker17, KBB^{+}20, KR83, KEL^{+}21, KCP22, Maa06, MUE23, MAG^{+}21, MSK01, MFH10, MDB19, NSM^{+}22, NSM16, NM19, PA01, RCC17, SM24, SM01, SAY16, Ush77, UGK^{+}14, VMY24, WKJ15, WCsH16, WBB15, WR22, XHP^{+}21, XXZ13, XDVB24, YOM^{+}07]. Efficiently [Lar90, SSO13, BK23, LHB18, PD00, SZ01, SCT02]. Effort [BP98, KVG19, Loe07, MKA^{+}22, SBB22, MAG^{+}21]. eID [BLC19]. Eiffel [ZC01]. EJVM [CC01]. Elaboration [LMSP92]. Elastic [Cha88, KS98, KCG^{+}12, NTF^{+}17, XYW^{+}24, ZXT^{+}17, ISPB21]. electric [HHMMG12, SCLD21, TVSG21]. electric/electronic [HHMMG12]. electrified [CL19]. Electromagnetic [SDF^{+}21]. Electronic [Gro73, HP87, SS84, Geh83, HHMMG12, NRUP21, OHD22, TP03, Ree76]. Electrostatic [GF80]. Elek [Val76b]. Element [EE90, GSWZ95]. elements [OAF^{+}03]. Eliminate [Geo77]. Eliminating [Roh81]. Elimination [SGH93, GvRN^{+}11, HNW^{+}01, KKN04, KWB^{+}05, OAF^{+}03, VH04, XCG06]. elixir [Bar78d]. Elliot [For72]. Ellis [Atk82b, Bis86, Cor82, Cou85a, Lav78, Mar88, Rob82a, Sto88, Vel88, Wal86b]. Elmwood [LLCG^{+}89]. Elsevier [Bar76a, Bar77e, Mul76, Wel72]. elsewhere [Bar82a, Bar82c, Pet77]. elusive [New82]. ELXSI [Car86]. Emacs [HH88]. EMAS [Bro86a, RS82, SYRS80]. Embed [LQ99]. Embedded [BP97, LF90, Set84, TLMP93, WR95, AH12, BP02, BC17, BRL^{+}15, CC01, GJ22, HKM^{+}09, JJK^{+}12, LMK16, MdCGdC^{+}17, Obe11, PACK07, PK04, SLRS06, SJP^{+}09, Sto05, TMJ^{+}21, VvK99, VC02, YYSG11]. Embedding [GL78, Sel75, WHYC23, ZWQC22]. Emblem [PPBP06]. embodied [BLE^{+}08]. Emerald [RTL^{+}91]. emergency [RGSGHGCG21, TLB^{+}18]. emerging [CGM^{+}03]. Emery [Bar73a, Ree76]. emotion [LCT^{+}21, SNK21, ZZC^{+}17]. EMP [SSK^{+}17]. Empirical [AJT79, BBB^{+}11, CSR93, Hoa73, Knu71, MW93, SP88, TV96, WXR16, ACKT20, CCPY12, CMS07, CBF23, CIM22, CIM24, DHA11, FSRF^{+}22, HKA12, KSK15, Lin98b, CCK21, NLA15, RN00, XP23, ZNWS18]. employer [TW16]. employing [LC12]. empowering [MPPC22]. empty [OAF^{+}03]. Emulating [Fra93, SROAdM^{+}08]. emulation [CBR10, CNRB13, PR16]. Emulator [PZ92, ACG78]. EMUSIM [CNRB13]. enable [JAkM^{+}21, Knu11, TMJ^{+}21]. enabled [CPD13, CBB17, JPG^{+}17, MTM22, PPSS05]. enablers [GVL10]. Enabling [DDGP18, TY14, Han11, WKG^{+}13, DGM^{+}22, RJZ^{+}20]. encapsulation [KT01b]. encoder [GJ22]. Encoding [LS96b, CWS07, ML20]. encrypting [LFGCGCRP14]. End [BP84b, HR06, KPJ^{+}17, WKS^{+}98, ASP^{+}19, Bha88, FBV22, Mej03, RSW^{+}23, WAML12, ZZJZ24]. end-to-end [ASP^{+}19, WAML12, ZZJZ24]. End-user [HR06, KPJ^{+}17, WKS^{+}98, Mej03, RSW^{+}23]. endgame [Mes80]. endpoints [SROAdM^{+}08]. ends [MP19]. Energy [AA23, BIA24, DLWF17, HSRB24, JI21, MDB19, XMTL21, YRJ18, ARMMA18, FDN^{+}18, HPK^{+}12, IBA^{+}21, IMG^{+}21, KCP22, LW19, NRS13, SDF^{+}21, SSK^{+}22, TBSI18, ÜY22, VMY24, WCsH16, WR22, ZSRR22, MAG^{+}21]. Energy-aware [HSRB24, FDN^{+}18, SSK^{+}22, ZSRR22]. Energy-efficiency [BIA24]. Energy-efficient [AA23, MDB19, HPK^{+}12, VMY24, WCsH16]. energy-performance [KCP22]. Energy-Saving [JI21]. Enforcing [CZ04, DDMD20]. Engelwood [Edw77]. Engine [AMW91, KMSS98, AI22, BB03, CD15, FG08, JC22, LPGBD^{+}19, PKB23, dKM04]. Engineering [BP09, BM93, BMS21, BC21, BW95, Byr91, CFKT17, Cd91, FS81, GLW82, GH02, HD86, KAO23, LN71, Mar86, NR04, Rin84, SWN94, SAN^{+}81, VC02, Wal84b, ACK^{+}22, AGRS11, BP11, BCSW20, BCP19, Bud85, DBOT23, DdB15, DPAG11, DBH04, ENP^{+}24, GN00, GdLC04, Han11, KKLL99, KPJ^{+}17, LKCC00, LTW^{+}21b, MP19, MGG^{+}09, NAU^{+}21, NZL19, OFRW10, PHB21, RBR21, Rob72, Rop88b, SLBC^{+}23, SKM01, TKF09, TAFC00, UFS99, WW00, ZSAM23, BAJMT21, FCBF^{+}21, Bar76e, Bux78]. engineers [Cou85b, Ell72]. engines [PSTV10, VC21]. England [Hut76, Wil74a]. Englewood [Bar73c, Bar74d, Bar75d, Bar75b, Bar76c, Bar80e, Ros74]. English [Ayc15, CS82, Coo05, Gu05, KHH^{+}15]. enhance [AA19, NTF^{+}17, RRK^{+}18]. Enhanced [FYP93, Kat83a, Kat83b, AKS06, AM00, CY01a, CY01b, LB02, LMK16]. enhancement [SO21, SSS22]. Enhancements [Web87, PH14]. Enhancing [BVGVEA13, BM01, HC10, KS01b, DFPT09, Haf13, KB06]. enlarged [CIM24]. enough [Wit77a]. enriched [LD14]. ensemble [CFLC14, SBB22, TRGA18, WR22, VBH^{+}98]. ensuring [SB13, YHK^{+}22]. Enterprise [GB02, MFB^{+}02, BMC^{+}24, BFRFB22, BSNB20, CPZ02, FCBF^{+}21, FSRF^{+}22, FFRF19, FFRFS19, HvdH02, JAkM^{+}21, KHGSS12, KJHG10, WWCW19, dAKdGJ11, FHB02]. enterprises [NB19]. entitled [CY01b]. entity [DS99, PP84, Pit82]. entity-relationship [PP84]. entropy [SDB^{+}22]. Entry [DW73, HPC^{+}96, vdBT77, Mor77]. Enumerated [Cai99]. enumeration [TEBK99]. Environment [ACC95, AJ78, BW88, BS93, Car81, CMF^{+}98, Cho98, Cro87, EMVW83, FM86, GR91, Hal86, HH88, Hay87, HW98, HD86, Jor90, KDP83, KM79, LL96, LFW96, Lei84, LS97, Lop89, Moh77, Org81, PL91, RS94, Rei90, RT77, RS95, SS95, SS93, Tay83, Thi93, TLMP93, WSB96, Wil82a, WMG94, Yip82, ACKS09, ASGA23, AGC10, ASAK03, AA23, Art82, BMC^{+}24, BDA20, BHMV09, BP08, CNRB13, CFLC14, CSMML12, CLS^{+}07, CC01, CC00, CSS15, DDMD20, FT01, GB13, GCRD04, GMC00, IB13, IK15, IH01, JPG^{+}17, KAS^{+}14, KGAR19, KEL^{+}21, KHC^{+}19, LLK04, LHGM15, LG19, MR07, MUE23, dSMLP22, Man01, MCGS08, MSR^{+}07, MKM^{+}17, NSM^{+}22, NNK21, PDCB17, PVR99, RGN^{+}14, SGAS21, SGA20, ŠS08, Spi02, SKM01, TGAS22, TRGA18, ZDY^{+}17, ZKB^{+}22b, dMdLvS99]. Environmental [Spa90]. Environments [Bre86, CL95, FHS92, FGIS97, Lyo85, SF98, Sha78, ARA18, AA14, AO12, BE02, CRB^{+}11, DGM^{+}22, DMR^{+}22, FJ03, GMNR20, GDGB17, HJC05, HB18, HIRB24, HMRZ20, HL03, HC12, ISPB21, JAA^{+}20, KKL17, KGAR18, KKA^{+}17, LQ04, MKW^{+}22, MA20b, NUB24, NRS13, QRD16, RBB12, dRRGdC15, RSRCGC15, SSCdA^{+}03, SHB19, TM14, TMGR^{+}22, Wet77, ZLTX18, dlRNGCC23]. EPE [FMC18]. EPE-Mobile [FMC18]. Equality [Van92]. Equation [DV84, Rin92, JL81]. equational [NWE99]. Equations [CFP83, HOS85, Ram96, BK23, Ell72, WLZ^{+}22]. Equivalence [Thi96]. Equivalent [JP22]. Erik [Cor99a, Cor99b]. Eriksson [Cor99a, Cor99b]. Errata [Ano86a, Ano87b, SFS97a]. Erratum [Ano73b, Ano19a, NJGG12a, NJG14, SMGMOFM07a]. Error [CG96, CL83, DP95, KL86, Nør91, OF76, PG81, PD78, Shr79b, Shr79a, SMM^{+}84, Sti85, Vau79, Bro82, EF13, Gla82, JK83, Pem80, Rön07, Thi12]. Error-checking [PD78]. error-handling [JK83]. error-recovery [Pem80]. Errors [FL76, Knu88, BPS00, Knu89, LF82, Mau82]. Ershov [Bar82b, Roh77a]. ESA [JH03]. ESA/NASA [JH03]. escapes [Fen12]. Esperanto [CMR07]. Essays [Bar76e]. Essence [Edw98a, Edw98b, JSKB22]. Essential [KW90, Pat94]. Essentials [Edm86]. Establishing [VDMW06, CZL24]. Estelle [TL98]. estimate [Rön07]. Estimating [Bai73]. Estimation [Moh81, FMC18, GM22, KKL17, KVG19, LMK16, MKA^{+}22, MS18, SBB22, TVSG21]. eSystem [ASA^{+}21]. eSystem-using [ASA^{+}21]. eSystems [BAJMT21]. etcd [LTK^{+}20]. Eternal [NMMS02]. Ethereum [LPGBD^{+}19, LWZ^{+}21, XP23]. Etter [Cou85b]. Etudes [Bar80e]. Euclid [BK87, Cor84]. Euclidean [NS23]. Eugene [Bul73]. EURECA [KPJ^{+}17]. European [BL15, BLC19]. EUSO [FCO^{+}19].

EUSO-SPB [FCO^{+}19]. evaluate

[MRG^{+}19, SRCP19]. evaluated [OM16]. Evaluating [BFRFB22, CMF^{+}17, CDG^{+}98, GRFFGC^{+}21, GR73, HCG^{+}16, MvSdL09, MPPC22, MKE18, Oli83, Sre76, TVSG21, dV89, DTJ89, EP05, Lar08, SMLS23, SGDA18, SB03]. Evaluation [And89, BG93, BBG04, BKP^{+}22, BF75, Dun93, ELRV93, Fra99, Ham77, HK84b, How78, KS98, KW92, LHH^{+}91, MHN18, MFdlP12, NPW72, OPTZ96, PKN^{+}12, REMC81, Rob83a, Sar77, Ste98, TB72, VG85, WG83, Wha93, WS99, WBV96, AMOS19, ACSK23, BB75, BSM^{+}21, CRB^{+}11, CNRB13, CS03, CCPY12, Chi17, DFPT09, DM15, HGK^{+}19, IB13, IK15, KCS^{+}20, KMB98, KSK15, MKA^{+}22, MNH04, Man18, MG09, MCHN05, QTV23, SS03, SH03, STB14, SJA^{+}04, SSRAH15, SZ00, UFS99, WRD99, YWN^{+}00, ZZKA17, ZXX24, ZSFY05]. evaluator [Glü12]. evaluators [ZZKA17]. Evans [Ano88c]. Even [HW90]. Event [ABG^{+}22, CSR93, Hać84, Han78d, Hug97, Mar84b, Ols90, OCH91, SNL15, She75, Sin81, Bru84, BD14, DPH16, Dar00, HL02a, IHS^{+}14, KRZ02, LCC14, Mal17, MZ00, PPR^{+}21, RMMH^{+}23, SSP11, The77, TKT^{+}07, WM20, XLZ^{+}20, SPHB11]. Event-B [SPHB11]. Event-based [OCH91, IHS^{+}14]. Event-driven [SNL15, MZ00]. Event-sourced [ABG^{+}22]. event-triggered [SSP11]. Events [BMZ92, DD18, GMGDMB19, RGSGHGCG21, WS94b]. everything [NHTT08]. evidence [BBB^{+}11]. Evolution [BJ72, Gra92, HJ08, HL94, Kil71, SBFF24, SFS97a, SFS97b, SFS97c, SYRS80, Str83a, ACCD01, CS17, CSS15, EAB^{+}03, FMNW04, FRBRF19, JTG^{+}11, LS22, PLR13, PPSO17, PSRCC02, SPR^{+}19, SDDD10, The77, vGB01, Loe07, Inc86]. Evolutionary [FCA12, GMGDMB19, ÖS96, WSYO11, WH06, NLA15, SBB22]. evolutionism [GM22]. evolutions [DZS09]. evolvable [MVV12]. evolving [NGLL14, SMT^{+}18, TTJ^{+}09]. eWare [JJK^{+}12]. exact [THG17]. Example [FS81, CC97, DRG11, MF08, ZLY18]. Examples [Rea73, Shr79b, Ten82, Hor21]. exascale [PMC22]. Excel [Tur22]. excellent [Bro82]. Exception [Knu84, Lee83, RdLFF05, SB93, vHLB^{+}88, CCF^{+}09, LYM04, NT84, TCMM00]. Exceptional [Geh92]. Exceptions [Geh92, Rin07, ZH01]. Exchange [JP74, GLS^{+}23]. exclusion [PCL^{+}99]. Executable [BM97, FGMM93, LB94, Özc98, Wat86, GHBH05]. Executing [RS94, Slo93, Van82, PCC^{+}12]. Execution [AG95, AP95, BBRB12, CRR94, GS76, GKM83, GH93, Hol89, JG94, Lar90, LQ96, BHMV09, DFD23, DS12, FSRF^{+}22, GCARPC^{+}01, Har99, HPK^{+}12, HML04, JJL17, JWTG11, LPGBD^{+}19, MC02, PJJM21, RMZ17, RGV14, SPPH10, SSK^{+}17]. execution-based [DS12]. executions [KM13]. Executive [Daw77, Heh76]. Executives [Ham74]. EXecutor [KE85]. Exercise [BNOW92, CK78, Fai87, Gom74, HWS^{+}88, Pet88, Sno78a, Str83a, Jon85]. exercises [QL13]. exercising [AWNS18]. exergames [WIYC20]. Exhaustive [DF84, RS93a]. Existing [Bro80, HUS^{+}91, MW13]. exit [Har84a, Mor77]. Exogenous [BMSZ17]. expand [YQL22]. Expansion [CMCH92, CK15, HYZ^{+}18, HYC19, NGLL14, SSD11]. Expected [PK89, Bur16]. Exper [XZ03]. Exper. [XZ01]. Experience [Ano23m, Ard87, BVB^{+}12, BCHR81, Ben90, Ber78, CC84, Coh75, CSS15, Cor08, Doo92, DFRR15, DF15, FSS99, FL94, GKBK16, GWY^{+}11, HW78, Har95, KHMB17, MSK01, MVS^{+}18, MSWC23, MPS93, MNW14, MS96, OSW92, OM16, OM96, OW16, OE92, Pal76, Pow79, RMZ17, Sam81, San88, SMFBB93, SL04, SAL16, Ste84, Sur13, Tag88, TK09, Var93, WBB15, Wis93, Woo72, vdRW79, vdWCB17, ABG^{+}22, BBB^{+}22, BM98, BDMP17, CL09, CARB10, CdA12, DBvG^{+}23, FSR11, FFRF19, Geh83, GS08, GHM^{+}06, Han99a, JGB15, JDGCGA12, MAR^{+}16, Pei02, PBGM18, SM01, SMGMOFM07b, SM15, Spi76, SGCM11, TGCF08, WWCW19, ZCO13, SMGMOFM07a]. Experiences [AK83, BS81, BHK^{+}04, CEMS24, Cer18, CB00b, DGR^{+}06, FP97, GSWZ95, GKS^{+}11, GHC^{+}07, GEF^{+}00, GVG^{+}18, HHR93, HPB^{+}00, Jor90, KG95b, LNW82, Lio79, NW78, Pry85, RPCS08, SC94, SAC^{+}92, SC90, TY80, Bir99, GMO01, KPK^{+}18, LG99, Sab76, SMT^{+}18, VHM^{+}05, AE06b, AE06a]. Experiment [Coo96, CHT91, CE84, Die97, ISUG06, MM80b, PD81, RMC97, SW86b, Str82, FCO^{+}19, Han77b, KAZ13, Man01, WZLN08]. Experimental [Ber85b, ELRV93, Har83, Lec95, LAD^{+}94, Lun89, OPTZ96, RB91, RG89, SS03, SS95, SSRAH15, SNM80, VDG^{+}00, Wol92, CVGCGCFF23, CS03, EGL18, HKWZ00, MVOD19, MdCGdC^{+}17]. experimentation [POZ^{+}16, TBT23]. Experimenting [IM93, TB86]. Experiments [Ano76c, BP90, DJM97, GM85a, KV98, Lec98, Smi91, TP92, AK15, GWY^{+}11, MSB20, NMG11]. Expert [LL91, Men97]. experts [GSS^{+}20]. explained [Vel88]. Explaining [Thi03b]. Explication [Hug79]. explicit [CEF02, KL12, SM18]. Exploit [AG95, PJ76]. Exploiting [BL15, CS15, DWL^{+}17, Dro84, EMD13, FH82a, Inn77, Man88, SWA^{+}97, ZH01, BCL13, CALL18, LBP^{+}13, UW99, UWW^{+}05]. Exploration [Rue93, AI22]. exploratory [ASME24, SBF19]. Exploring [CWC^{+}21, GVG^{+}18, dlRNGCC23, dSCdRS^{+}19, MBV^{+}10]. explosion [BDSV99]. exported [KF02]. express [HIRB24]. Expression [Ber85b, Ier09, Ric79, SM99, BY17, Chi17, KS08, LLZ20, SCF^{+}17]. Expressions [GR73, Han85, Kea91a, Ram98, Set81, HNW^{+}01, KKN04, LM81b]. expressive [DFQ^{+}22]. Extendable [BT75]. Extended [AE14, BGS^{+}13, BMD^{+}98, BPK13, BC13, CCPR91, CQH^{+}13, DDF16, DW73, DDDF17, DC15, DE16, EMD13, FBB^{+}14, GBG^{+}14, GB13, GMDM17, GQ15, HS83, HYH15, HCG^{+}16, Kap13, LSZ16, LMK16, MMOD16, MDH^{+}13, Obe11, PT14, POZ^{+}16, PDPM^{+}16, PKvdWB17, QM13, QL13, QRD16, aSZP^{+}16, WJC^{+}14, BPDL23, HLR^{+}03, KA87, KKA^{+}17, ST19]. Extendible [Kno81, PT90]. Extending [BB10, CEF02, Hsu12, Kea91a, LPA13, Spa90, Tsi82, WR95, MLC02, JGC^{+}21]. Extensible [Fin97, HH88, HC97b, IdFF96, Ker80, Sco73, ALKL19, Bar74c, BBMN18, BRMO97, BR01b, DCA04, GA12, Ged14, GLT08, NHTT08, PNP20, SBG^{+}05, SMGMOFM07a, SMGMOFM07b, Sta05, TK09, TGPS08, TLC^{+}18, WMJ04]. Extension [BR95, BAFR96, BMS83, Bou91, FD92, GH72, Gri80, IdFF96, KS80, Liu86, MTT81, MTT83, MB97, Sau88, Sch89b, CH06, Ger82, HT82, Kir07, vD99]. Extensions [CMH85, DT96, FYP93, HTJNL19]. Extensive [DAC^{+}21, ACSK23]. External [Col88, MKD98, BST10, CS17, Tsi82, ZZKA17, ZWAL22]. extract [Wir77a]. Extracting [NMRW98, BLNU15, CLP^{+}09, JAJB04, LJS20]. Extraction [Kea91a, AML20, BT21b, DFQ^{+}22, DGPT14, GHBH05, PJJM21, RR22]. extractor [UGK^{+}14]. eXtreme [CCM05]. extremely [JLZ09].

F [Bar76e, Bar77b, Bra75, Bul72b, Cor82, Ell72, Jon74, Lan74a, MDB19, Nic72, Sha83, Whi87, Wil72]. f2c [Lev95, Lev97]. f2cl [BW96]. F99.50 [Flo73]. faas [RRFD23]. faas-sim} [RRFD23]. fable [Hen79]. Face [OAZ19, LCGS17]. faceted [BDQR23].

Facilitate [LD87, MGP03, WYAZ15]. Facilities [AH85, Cav83b, CV98, SWA^{+}75, Kur78]. Facility [Bai85a, BL78, BL79, Bow73, Bro80, DLP85, EE90, Gri75, Jon71, MG94, Mal83, Mil74, PSA87, SL78, ZZWD93, Ano81n, CW82a, JZ02, MBB^{+}86]. factors [AKIM24, DBOT23, Han11, MCLL21, SdLJMP21]. Fagan [Doo92]. FAHP [KGAR18]. failed [Bar78d, Bar82c]. Failover [MKM^{+}17]. Failure [SO77, Wha72, Eba20, WWGP10]. failures [MGJ^{+}24, SC22]. Fair [CLCC15]. Fairthorne [Lav78]. fallback [ML23]. false [JK14]. families [MPBH13, NGLL14, Wij05]. family [AKM17, BCFT95, JKB04, SL04]. farming [TJB^{+}19, ZLZ^{+}21]. Fast [AC13, App89b, ACM^{+}15, BP98, CM96, Col77c, CS82, CW08, DF87, Dri93, Fen01a, GS06a, Han90, HCDB19, Heu86, Hor80, HS91, KST94, KH96, Kur81, KKdC^{+}23, MZB00, McC90, McK89, MEP96, MFYiA01, ML23, OM88, RK15b, Smi91, Spi04, Wha93, YLP^{+}11, BCRF22, Cox76, DD10, DPDA14, LJL^{+}10, MR04, Nav01, OAF^{+}03, OG16, PP16, SS07, SAA^{+}20, Sta07, TL14, ZC03]. fast-prototyping [ZC03]. FastCGI [BCL13]. Faster [Gor94, HW90, KG18, LKK19, Yuv79a, BMS21, KAO23, LSYKK16, LNhCW16]. Faulkner [Edw98a, Edw98b]. Fault [BTM81, CD94, DJM97, EKM^{+}99, FYP93, GSAE14, dSMH13, Pla97, SF98, SMR93, Web87, WHLM98, APS^{+}11, AA20, AA21, CC13, Cla98, DW13, GSPA^{+}11, GWY^{+}11, HGK^{+}19, IBA^{+}21, JKH22, JFZ^{+}21, LLW^{+}23, MKM^{+}17, NMMS02, NNLR17, NNR18, NM06, WHS^{+}00]. fault-proneness [WHS^{+}00]. Fault-tolerance [Pla97, JKH22]. Fault-tolerant [CD94, EKM^{+}99, dSMH13, SMR93, Web87, NMMS02]. faulty [ZGG07]. FC [SM02]. FcgiOCSP [BCL13]. FE [MK03]. Feasible [Hal86]. Feature [DHWZ14, GWX^{+}23, KKLL99, LKCC00, SO21, YAVHC21, ASEF23, DFQ^{+}22, GKWS11, GDW^{+}20, KB06, MRBB19, NSM^{+}22, NGLL14, PdSCJM22, San17, Tur06]. Feature-based [DHWZ14, LKCC00, KB06, Tur06]. Feature-oriented [KKLL99]. feature-rich [ASEF23]. FeatureC [KPK^{+}18]. Featured [LTW^{+}21a, ZKB^{+}22a]. Features [GR79, Heh76, Shr79b, AML20, LYLY20, MSB18, OAZ19, SROAdM^{+}08, SNK21, TTJ^{+}09, WLTJ13]. Federated [DFQ^{+}22, LHS^{+}95, DS99, GLKZ21, KS24, STB14, ZGB24]. federates [ATO10]. feed [OAZ19]. feed-forward [OAZ19]. Feedback [Bur98, FKL^{+}13, SW14]. Feldman [Bar77b]. femtolet [MDB19]. femtolet-based [MDB19]. Fenton [Pra96a, Pra96b]. Fenwick [MV20]. few [CCPY12]. FFG [Com82]. ffortid [Ber99]. FFT [MV95, NPHJ18]. Fi [CdA12]. FIDEL [KS24]. fidelity [KS10]. Field [BP90, GSS^{+}20, TP92, WWCW19, ZZJ21, BK23, Rei90]. Fields [Ham84]. Fiendly [Lin86]. figures [Bre82]. File [ADM96, AM86b, Bar78a, BB81, Bar75d, Ben77, Car79, CE97, CS91b, Col77a, Com82, Del82, EV89, Flo73, HJS89, Han80a, Jeg83, JB84, KK90, LA90, Lun86, MNH04, MM85, MM86, MS96, OSW92, PSA87, Qui91, RS86, RH77, RB75, SZ88, TWL94, TKWW85, WR78, vdBT77, AML20, BGM99, BGS18, HC12, HC16, Jac71, LWZ^{+}19, MM82, Wal83a, Flo73]. File-processing [Col77a]. File-store [SZ88]. Files [Bre86, Cow87, EL96, HC98, KPT86, Kno81, LB94, Mou72, MT84b, Org81, Als22, Ayc15, CBF23, HZ95, JRGC20, PB03]. Filestore [MM81]. Filing [PGH^{+}98, Wha72]. Filling [Col83, Ano71d, Gol81b, Gri86, Pal86, WW83]. Film [BMA72]. Filter [JMM03, AAB^{+}21, GSR17, MAT94a]. filtering [NPHJ18, PDROFRM13, ROFGFRM16]. Filters [JVR97, App22, NJ11, ROFGFR^{+}16, TRGA18]. Financial [DV84, GZX^{+}21, MVM^{+}22]. Find [Spa90, CBF23]. Findally [Atk79b]. Finder [JGR89]. Finding [Coh98, HK06a, ZD95, BPS00, LBP^{+}13]. findphrases [AB89]. Fine [CW97, DFOT10, JR92, MT94, Day00, LBP^{+}13, SHIS99]. Fine-grain [JR92, MT94]. Fine-grained [CW97, DFOT10, LBP^{+}13, SHIS99]. fine-granularity [Day00]. ‘Fingerprinting’ [MM82]. finish [KSBL22]. Finite [EE90, GH72, GSWZ95, HC93, LK93, PJJM21, BT21b, LSZ16, Lam20, LQ99, NKW06, Wat04]. Finite-element [GSWZ95]. Finite-state [PJJM21]. Finlay [Wri98]. FIPA [BPR01]. FIPA-compliant [BPR01]. FIR [NPHJ18]. FIRE [KS08]. FIRE/J [KS08]. First [CS91a, Cas92, CZA83, Hua87, Rue93, Wex81b, Fox79, Gla82, MRZ15, NNL^{+}14, WMHL21]. First-Order [CZA83]. fitness [WH06]. Fitting [Ell72]. five [FRBRF19]. five-year [FRBRF19]. Fixing [Wad87, ZPSH21]. flaky [ZPSH21]. flame [GARSR18, RCH^{+}23]. flamingo [SS23]. flash [CSM^{+}16, CLC99, HC16, LFHL22, DWL^{+}15]. flash-based [LFHL22]. Flat [Com82]. flaws [ST19]. Fleming [SFS97a]. Flex [JJK^{+}12]. Flex-eWare [JJK^{+}12]. Flexible [BP97, Dew91, Dew87, GHM96, GS85, HC97b, KS20, LD95, LHC97, Pfe84, PR98, PES^{+}20, PKC^{+}13, SDC04, AHRR22b, ALKL19, dODP21, AV84, BMR00, BDL09, CARB10, CV08, DMD^{+}06, DS03, DFRR15, HCDB19, JJK^{+}12, KS01a, Nav01, PD00, RCH^{+}23, TGCF08, WCS^{+}17]. FLight [ZGB24]. Floating [Far88, Has77, NC75, Ume91, VS80, SF88, Ush77]. Floating-Point [NC75, VS80, Far88, Ume91, SF88, Ush77]. flood [GMPL11, WR22]. FLORA [STA09]. Flores [Bar75d, Bar75b]. Florida [Rob72, TWS^{+}22]. Flow [AS97a, CK94, CC87, HGW94, LMK16, Mat94b, OPTZ96, BDLM04, Ber82, CCvKH95, CS15, DB21b, FGMM93, KBH^{+}03, PW11, RAP21, RMC97]. Flow-sensitive [LMK16]. Flowchart [Geo77, Glü12]. Flowcharting [Wit77b, Con77]. Flowcharts [Cha74, Wen80]. Floyd [Tor23]. fluid [ZGL^{+}23, ZLZ^{+}23]. fluid-structure [ZGL^{+}23]. fly [BGM99]. FMEA [GWY^{+}11]. FOAM [GW85]. Focus [HW10a, HW10b, Tse13, TGC15]. focusing [FFRFS19]. Fog [KS24, ARK21, BAM^{+}20, BBH^{+}22, BFGL20, CKCG23, DGM^{+}22, GAAA22, GMGB22, GLL20, HSRB24, HSY^{+}20, JSPP21, KRK21, KA22, KMB^{+}21, KDA20, LWW^{+}23, MSWC23, MDB19, PHB21, PJS^{+}23, RVS^{+}20, SGAS21, SA20, Sri24, TGAS22, TJB^{+}19, VS20, ZGB24, dlRNGCC23, FLPM20, GDGB17, MKC20]. fog-based [BAM^{+}20, PHB21]. fog-to-things [KDA20]. folk [Bar82a]. folksonomies [EMD13]. Follow [Atk79d, Fai87, Sti85, RGSGHGCG21]. follow-up [RGSGHGCG21]. followers [Bar77b]. font [KNT^{+}01]. Fonts [CT92, Ber99, NHTT08, PB03]. Fooling [Plu77]. foot [WHH21]. footprint [MTPC14]. FORALL [Ker80]. Force [FR91]. Force-directed [FR91]. Forced [Dro85a]. Ford [Ano87a]. forecast [MAA^{+}22]. forecasting [CLD^{+}17, JLBR22, OM16]. FOREET [BA86]. forensic [QC17]. forest [SDB^{+}22, SB21]. forest-based [SDB^{+}22]. forests [TRRK21]. Form [BCHS98, Bro72, CH73, Fai87, AMR90, Geh83, LMPR07, MP02, VH04]. Formal [BS88, CG96, Die98, Geh82, HL98, LBS78, MMS90, NN24, Özc98, Pag84, PGK^{+}10, SL87, WB78, AGRS11, BR01a, BLLP04, GF11, JJI22, MKE18, RMMH^{+}23]. formalism [Pol01]. Formalization [Hug79, KHHG15]. Formalized [CCvKH95].

Formalizing [BNOW92]. Formally

[FCYL18]. Format [Cha74, Gra81, HKW77, OMA96, TK72a, LC03, Wu01, Wu02]. Format-dominated [HKW77]. formation [NN24]. Formatted [RW81, Woo86]. Formatting [BS84, BF80, GW85, Kin93, Noo83, SW87, Ber99]. forms [RGC^{+}21]. Formulae [Lev83]. formulas [RD14]. Formulating [SAY16]. Forsythe [Ald72]. FORTRAN [RB82, Ree73, Bar72c, Cou85a, Cou85b, Edm86, Ree75, AI80, ASH73, Coh74, CA86, Cra76, DH79, Ell82b, GH72, GM73, GF81, Gut76, HS83, HLS73, HT82, Hoa73, Ker82b, Knu71, Lar73a, Lar73b, Les72, Lev95, Lev97, LV73, LS75, MS74a, MP79, Nee75, NC75, NY78, REC75, Sab76, Sch72, TR77, VS80, Ano81n, BA86, Ben77, BW96, CT90, Fre81, HWS^{+}88, Ker75, Ker80, KO91, Lar81, LHH^{+}91, Moh77, Oni85, OF76, OE92, Pal86, Par78, PD81, RT77, Sch89b, SM90, Sco77a, SAC^{+}92, Tse97, TWI88, FCG83, Bar80d, Wil87, Bar73d, Bis81a]. Fortress [Ryu16]. forum [KYF^{+}22, Val77b]. Forward [AF99, Sal81a, OAZ19, Rus95]. Forward-adaptive [AF99]. Forward-declared [Sal81a]. FOSSES [AMOS19]. Fought [Pal78a]. Foundation [Kor92, KNC94]. Foundations [KS95, JC19, Sim83, Atk82b]. Four [Fle90, HZ94]. FPFTS [JSPP21]. FPGAs [TL14]. FPS [SAC^{+}92]. FRACTAL [BCL^{+}06]. fragment [BPP10]. Frame [Har92, McC90, KCH07]. Framework [AMOS19, AFI98, BS98, CCR19, Gan82, Gra92, HS97, JG94, LCW98, RA95, Sef97, AA19, AKIM24, AMM10, AZS19, BN00, BJ21, BHR15, BGS^{+}13, BPR01, BFG^{+}11, BFPAGS^{+}08, BSDF20, BOPN12, CLZ99, CDR13, CGP^{+}06, CC02, CV03, CP22, CYW^{+}15, CI03, CP07, Coo04, DHS02, DGRB15, DDDF17, DZ21, DP09, DFQ^{+}22, DMR^{+}22, DM15, DS03, DAJ^{+}15, DF15, EF13, Eng06, EC13, FG11, FRGPLF^{+}12, FMC18, FP15, FLSCC15, FMPR02, GH03, GT00, GA12, GMNR20, GMGB22, GMC^{+}21, GRVC23, GDH13, Har82, HvdH02, HSC21, HK06a, HIRB24, HLFS05, HML04, JAA^{+}20, JXG^{+}21, JPM23, Kat17, KCH08, KYF^{+}22, KTG20, Kil19, Kim02, KHS^{+}24, KDA20, KSK15, KS24, wKJM18, LBGA^{+}21, LSK^{+}18, LS15, LYX^{+}17, MAG^{+}21, MS99, MGBO22, Mej03, Mos06, MPJ20, MGJ^{+}24, NMMS02, NZL19]. framework [NZTL^{+}23, OOG19, OMGDG14, PNP20, PSD^{+}04, PALNGD^{+}06, PVBB06, PPSO17, PDROFRM13, PDPMM17, RRFD23, RCH^{+}23, RZ17, RL23, Ryu16, SN01, SESK23, SCL00, SM20, SMLS23, SIK^{+}16, STA09, SH22, TTC^{+}13, TMGR^{+}22, UCCPM19, VSID17, WY18b, WIYC20, XCL^{+}18, ZA07, ZXT^{+}17, ZHO^{+}19, ZGB24, vDV04, HLR^{+}03]. Frameworks [vdWCB17, CL09, CPZ02, FHB02, FBV22, FRBRF19, GMGL24, GB02, GVL10, MFB^{+}02, PRTS06, PLPA22, SBD15, TSZ14, vGB01]. frameworks-A [GMGL24]. France [Lav77]. FReD [PJS^{+}23]. Free [AMR90, Gra81, OMA96, SW86a]. Free-form [AMR90]. Free-Format [Gra81]. Freeman [Lar75a]. FreeRTOS [MNW14]. Freeze’nSense [KKL17]. Freiberger [Wil72]. Frequency [Fen94b, Fen96, Abe07, BLM00, CW08, Fen94a, LLW^{+}23, ZWSS15]. frequency-based [CW08]. Frequency-to-Symbol [Fen96]. frequent [CLP^{+}09]. Fresh [Fid82]. friendly [GJ88, MRG^{+}19]. friends [Ber20, MP18]. Front [Bha88, BP84b, FBV22, MP19]. Front-end [Bha88, FBV22]. front-ends [MP19]. fsh [McD87]. FT [EKM^{+}99, BM03]. FT-CORBA [BM03]. FT-RT-Mach [EKM^{+}99]. Fuel [DPDA14]. fulfilling [LC07]. Full [BdPGS14, Car81, LSF94, ZM95]. Full-screen [Car81]. Full-text [ZM95]. Fully [JL91, BCSV04, FDHH04, PDPQ22, SLJ^{+}18, YMH16]. Fully-lazy [JL91].

FUMBLR [McC83]. Function

[BM93, CQC98, Col77c, DH88, DW91, Fai87, FP82, JGC^{+}21, Kan18, Lic86, OLS89, Ric79, Sch76b, Wic77, BBH^{+}22, CH06, Che04, NNR18, SESK23, SHB19, SOKE23, WKKK24, XP23, YSBL22, ZA07, ZTC24]. Functional [BY90, Fai87, FFD96, GSWZ95, HGW94, Koo87, KvEP95, Lei84, McD87, MV86, Wad85, WR95, BVGVEA11, BCPSC18, Jon85, KAS^{+}16, KA87, MJ99, PRP^{+}23, RBD^{+}24, SGCM11, VP05, SM02]. Functionality [UFR18, CS22, SRGCPB^{+}09]. Functions [Hol89, Mid86, Oli83, Sch72, Sew82, ESRI14, HHMMG12, JPL03, NS23, Sar77, Tur22, WH06, ZHO^{+}19, ZLY18]. Fundamental [Tra79b]. fundamentals [Mog04, Bar79a, Bis86]. Fusion [JI21, DFQ^{+}22, GDW^{+}20, HTX^{+}22, LCT^{+}21, Man01, NSM^{+}22]. Future [Moh81, ASA22, AH12, BLC19, CJTK22, DH00, GKS^{+}22, HSRB24, OHD22, RB19, Sri24, XZX^{+}21, ZML13, ZPSH21]. future-context-aware [ZML13]. fuzzing [MSK^{+}24]. Fuzzy [Kop97, LL91, PW97, ASA^{+}21, GT00, JSPP21, KGAR19, KSK15, MSWC23, ZB18]. fuzzy-ant [KSK15]. fuzzy-based [MSWC23].

G [And78, Ano73a, Ano79a, Bar73a, Bar74e, Bar75a, Bar76d, Bar77c, Bar78b, Bra80, Bul72a, Eme84, Ken77, Ree76, Roh77a, Rop88b, Val76a, Val78, Wri98, CPMAH^{+}20]. G. [Sau88]. GA [LBC^{+}11]. Gabriel [Nic72]. Gaia [DFRR15]. gains [MS99]. Game [TT74, Kar21, SSGA20, VC21, WWJ07, GRVC23]. Game-based [GRVC23]. game-theoretic [SSGA20]. games [PV22, RSRCGC15, Ano73a]. gamification [PBGM18]. gap [CDM^{+}16]. Garbage [App89b, BW88, Chr84, CM96, FH92a, GT87, Nil88, RRR97, Wad87, Wen90, Zor93, CS15, Hug82, MKD^{+}22, PDPM^{+}16].

Garbassi [McD71]. Gary [Ano87a]. gate [BK23, ZZJ21, GLMS18]. GATE-based [GLMS18]. gated [WHYC23]. Gateway [Yas94, KCK^{+}23, LAG00]. gateways [VRC^{+}06, dOED^{+}20]. Gathering [Yuv75]. Gauthier [Bar73c]. GBAaaS [GRVC23]. GCC [KSK09, LC12]. GCI [GB87]. GCM [BHR15]. GCM/ProActive [BHR15]. GCOS [HCD84]. GCOS-7 [HCD84]. GDBApex [JAkM^{+}21]. Gecko [BH01]. Gem [Lev82a]. General [Coo85, Dew84, FL75a, Hal82, Haz74, HM84, LF74, Lew83, LTV96, Par85a, RTL^{+}91, Spo71, Vo96, Wal80, Wal90, AYdS^{+}06, BK77, CIM22, CIM24, DPDA14, JSC^{+}10, KNT^{+}01, KW17, MLR19, MK03, PM18, WPL^{+}21]. General-Purpose [FL75a, Haz74, RTL^{+}91, Lew83, AYdS^{+}06, JSC^{+}10, MLR19, PM18]. Generalizable [WWGP10]. Generalizations [AS87]. Generalized [Blo93, Bor86, Kil81, MJ98, SG93]. generate [CQH^{+}13, PKK12]. Generated [WC85, WSB96, GIF01, GMGDMB19, HCG^{+}16, Sto05]. Generating [AB89, BB95, Bri87, Coh75, FIÁSLSAR05, Fis86b, FP82, KM89, NSW77, TWI88, VR06, WP00, AÇCT22, GMPL11, HKWZ00, ZZ11]. Generation [AC80a, AL82, Amm77, BK23, BLLP04, Cla89, CH90, EV89, FH91a, FH91b, Gor94, Gro89, Heu86, KFJS88, KL86, KKM80, Len90, Les72, LT85, LD87, OMA96, Pet76, Pfe84, Rév85, RB75, Ste80, Wal84a, WW83, vHE87, ATO10, AB88, AWNS18, BM06, BFGS05, BPK13, BIA24, CCR19, CA08b, DDGP18, FCA12, GNSP12, GQ15, HTJNL19, HKA12, HLGSW11, KSK09, KH07, KAS^{+}16, MPBH13, MP02, Mid79, MRG^{+}19, OJP99, PACK07, RGS^{+}20b, RZ17, TCMM00, VNLB20, WH06, WGM08, ZC02, ZGGL23]. Generational [App89b, Ono93a]. generative [KS08]. Generator [CCRD^{+}80, Cla86, FHS92, Gro90, GJ88, GS85, HS89, Hun97, KS82, KNPS88, Kou87, LTV96, Mat83b, Mau92, SIN95, Sch89a, SG97, SN90, VSM87, vR92, Abb78, DHGR92, EGKP02, HL87, Lar09, MS83, PQ95]. Generators [Ber88, GF84, LS76, WG83, SV22]. Generic [ELRV93, Ged14, Ian90, IHS^{+}14, JHKS19, MS94, Wil89, BMY06, CP07, Fer13, FP15, GL05, RJ09, RCC17, SH03, Sav04, SH22, TLB^{+}18, TGPS08]. Genesis [WS94a]. Genetic [Kra97, Mon96a, Mon96b, AA20, Nic98]. genetics [AA21]. GenEx [MM01]. Genuine [HO91]. geo [HC20]. geodata [HM18]. geographic [BCLF^{+}07, CKL^{+}02]. geographically [ZB18]. Geometry [DNSG89, FGK^{+}00]. GEORGE [Oes71, Ano73a, BT74]. Geschke [Bar77e]. Gesture [BCRF22, KHHG15]. gesture-based [BCRF22]. Ghost [CV84]. gigabyte [Lem21]. Gildersleeve [Jac71]. Gilman [Bar71]. Ginga [SMM13]. Ginga-NCL [SMM13]. GINI [YMY17]. Ginkgo [GAQO23]. GINO [Woo71]. GISQAF [ANSK16]. GitHub [AGM17, SCD^{+}23]. GitRanking [SCD^{+}23]. gives [Bro82]. GLAL [ASAQ05]. Glass [Bar78d, Bar82a, Bar82c, Bar84b, Bar84a]. Global [Er85, Fis83, FL94, GW96, LLW98, SDB^{+}22, ZLY18, Bra99, BMAV05, HOY17, Loe07, NDKL^{+}24, GPR^{+}98]. GloudSim [DC15]. GLU [PK04]. Glue [Van86]. GLUnix [GPR^{+}98]. GMB [JG89]. GNU [HH88, ZC01, BGM99]. GNU/ [BGM99]. Goal [Nil90, WG83, New82, ST12]. Goal-directed [Nil90, WG83]. goal-driven [ST12]. Goecke [Wal81a]. Golden [Buy21, SB22]. Good [KP94, Vel85, Ber85a, KHMB17, SV22]. Google [DC15]. Google^{T M} [GK08]. Gopal [Haz71]. Gorbachev [Ano23m]. Gordon [Bar75c]. Gosling [Cou84a]. goto [Yuv79a]. Gould [Bar72a]. government [HL20, PCdGPP12]. GPGPU [KLY20, TY14]. GPROC [O’N88]. gprof [Var93]. GPRS [SBcC07]. GPS [XDZ^{+}17]. GPU [RCH^{+}23, Che22, dSJCM16, KLY20, NPHJ18]. GPU-accelerated [NPHJ18]. GPUs [LFH^{+}22, RCH^{+}23]. Graceful [SFS97a, SFS97b, SFS97c]. grade [CVGCGCFF23]. Graded [Gru83]. Gradient [YAVHC21, IB13, SB21]. Grady [Wal84b]. Graham [How76]. grain [JR92, MT94, Wis93]. grained [CW97, DFOT10, LBP^{+}13, SHIS99, WHH21]. gram [Coh98, KST94, YAVHC21]. GramCheck [Sha05]. Grammar [HLGSW11, MP19, Mau92, MSK^{+}24, SIN95, WGM08, BP08, GQ15, JAJB04, LV01, LHC15, LZZZ18, Sha05]. Grammar-based [HLGSW11, Mau92, MSK^{+}24, GQ15, LZZZ18]. Grammar-driven [WGM08]. Grammars [BV89, Fro93, KR83, Pap79, GN16, HHM92, Mer93, Zdu07, Ier09]. Grammatical [EF13]. Grammers [FSO91]. grams [GSR17]. granularity [AML20, Day00, HBK20, HBB22, NS01b]. GRAP [BK86a]. Graph [Cd91, CP96, Ear76, FR91, HV88, Har91, HGW94, Hop71, Hos98, JG89, PT90, RS93b, VMJ97, WHYC23, BDG^{+}00, Bha88, BS99a, CCQ16, CMCL03, CCCZ05, CCT01, CYJ^{+}22, CHT98, DPDA14, EBFK10, GN00, GAAA22, HZWC22, Him00, JAkM^{+}21, LHC15, LQL^{+}22, LZZZ18, MHM01, RR22, Spi04, VDG^{+}00, WLS^{+}21, YLP^{+}11, dMFÁE17]. graph-based [JAkM^{+}21, WLS^{+}21, YLP^{+}11]. Graph-embedding-based [WHYC23]. graph-labeling [CCQ16]. graph-oriented [CMCL03, CCCZ05]. Graphic [Gan82, Lan74a, Lib97a, Lan74a]. Graphical [Bov87, Dan90, Dun93, HG89, HM90, KKS88, KRO93, LD95, MTT81, MTT83, MB96, PN83, Ros77, SG97, Str83b, BB99b, BE02, Deo10, JCL85, KBBS05, MKD^{+}22, MRG^{+}19]. Graphics [ARS^{+}94, BV89, Bec91, BH87, BT74, Ham84, JGT95, Ker82a, Kil71, Les72, Mil74, Mor82, NM78, PLR85, Ric76, Sla86, Van82, Woo71, BBGP01, For72, GRS74, Lar08, LD99, MBB^{+}86, SCT02, Yip84, Zho03, Bra75, Edm82]. Graphlet [Him00]. GraphRedex [SS21]. Graphs [BK86a, CFP83, MD88, OE92, RDM^{+}87, BS99a, GNV88, HB18, Ple99, RAP21, Tor23]. GraphSET [EBFK10]. GRASP [Wor83]. gray [CZL21, SGAS21]. Green [For72, DGM^{+}22]. grep [Nav01]. Greps [Hum88]. Grey [Ear77]. grid [CBR10, EHV99, HZWC22, KBM02, ZLZ^{+}23, ASEB09, BBL02, GAH05, HBJ05, MZC10, McN05, PPSS05, SROAdM^{+}08]. Grid-based [GAH05]. Grid-enabled [PPSS05]. gridded [Bra99]. gridification [MZC08]. Gridifying [MZC10]. Grids [BBL02, HML04, SGCM11, VNGB08, BMAV05]. GridSite [McN05]. Gries [Fin77, Han72]. Griswold [Lar75a]. Grogono [Bis79b]. Groner [Nic72]. Ground [Buy21, Coo08]. Ground-breaking [Buy21]. Group [Rin84, TP92, DF15, GLD^{+}21, GEF^{+}00, HM18, MMHB08, PK11, RPCS08, SAEGF11, ZZKA17]. group-based [HM18]. group-oriented [SAEGF11]. grouping [Nic98]. Groups [BIO94]. Groupware [YH97]. Growing [Cou92]. Growth [Tal71]. GRUMPS [EAB^{+}03]. GSL [WKS^{+}98]. GSM [BLLP04]. GSQL [MWB95]. Guaraná [BFRFB22]. Guard [SA97]. Guarded [Fis84, Fro93]. ‘Guardian [SJ79]. guards [Rai99]. Guest [Ano71f, Ano76b, CM98a, CM98b, CJ73, D’A73, Gro72b, Hal71, Han81d, Hat73, Hoa72, Jon74, Lam72, Ros71, SFB13, Wai73b, Wil72, Wir72, Wir77a, ZWC21]. GUI [CDGP93, Spi02, SA02]. GUI-builder [Spi02]. Guide [Bar72a, Cou84b, Mee87, Atk83, HvdH02, McD71, Mil72, Lev98, Bar75a]. guided [ANSK16, CMCH92]. Guidelines [RBS14, TKB78, DDMD20, vGB01]. Guides [Cou84b]. guilders [Flo73, Nee77a]. GUIs [AWNS18]. Gunther [Sim83].

H [Bar72a, Bar74e, Bar76a, Bis81b, Bra80, Bul73, Bux78, Cam85, Han77a, Ken77, Lar75a, Liv75, Mer74, Nee77a, RB82]. H.M.S.O [Bar75a]. Haar [OAZ19]. habits [CS15]. HACKERS [Yuv77a]. HADES [Wil82a]. Hadoop [dCCCdAC20, CP22, LCC14, hPmKgH15, RGS^{+}20a, TTC^{+}13]. Hadoop-based [CP22]. half [Has77]. half-word [Has77]. Hall [Bar73c, Bar74d, Bar75d, Bar75b, Bar80e, Edw77, Edw98a, Edw98b, Lar71, Ros74, Wri98, Bar76c]. Halpern [Roh77a]. Halstead [Bar76a, Woo84]. Halsted [Bry77]. Halting [Sch86]. Hamming [GK21]. hand [WHH21]. Handbook [Gar86]. HANDIN [CM85]. Handler [KWW81, NT84]. handlers [Han83a]. Handling [BPM93, BMZ92, DP95, EBD^{+}74, Hug97, Knu84, Lee83, SB93, Wal81c, WB77, vHLB^{+}88, CCF^{+}09, JK83, LYM04, RA87, RdLFF05, ZNWS18, Bar78d]. handoff [HM18, SBcC07]. handoffs [CLC09]. handwritten [BFGS05, RBR21]. Hans [Cor99a, Cor99b]. Hans-Erik [Cor99a, Cor99b]. Hansen [Hor07c]. Hard [ABRW94, BW95, FH91b, Hal86, Atk78, DKM11, Fox79, Lon88, Ree78, Rob81]. Hard-coding [FH91b]. Hardback [Atk82b, Bis82, Ano79a, Bis84, Cor82, Mad82, Mee87, Sim83, Ree84a]. hardcoded [NKW06]. Hardened [PF09]. hardening [NJ11]. Hardware [CK86, CPHS83, NC75, Pal78a, PLR85, RK89, SD18, Bar83a, CMPQ23, DSD^{+}05, Has77, Mer74, ZKZ^{+}21, Han78a]. Hardware/Software [PLR85]. Harland [Sto88, Wal86b]. harnessing [DDB^{+}18b]. Harrison [Ano88b]. Harry [Lar71]. HARTEX [AIB02]. Hartmann [Pem80]. harvesting [XMTL21]. Hash [Coh98, CS82, ESRI14, Rön07]. Hash-Bucket [CS82]. Hashing [BT89, CW91, GT93, Har71a, HC87a, MHB90, Qui83, DM11, IIL17, ZWF^{+}22]. HASKELL [JL91, LV20, SC94, Thi97]. Hatching [Vör84]. Having [LL91, SJK^{+}21]. Hayes [Tho74]. hazard [Thi12]. HDFS [FB23, KKA^{+}17]. Head [Mil72]. Headers [Lit93]. healing [SBD15, SMT^{+}18]. Health [CFS^{+}22, JI21, OHD22, WHH21]. Healthcare [Eba20, BCC^{+}22, BJ21, DGM^{+}22, KCS^{+}20, PPSS05, SAYM24, MGJ^{+}24]. Heap [ACCM83, Mar79, Sch80, PNM^{+}20, SHF16, SZ01, ZG06]. Heap-based [Mar79]. heap-object [SZ01]. heaps [Kat17]. Heart [Kah95, MGJ^{+}24, BGSG20]. heaven [Wir77a]. heavyweight [MK18]. Hebrew [Ber99]. Heedless [Thi12]. Heidelberg [Cav83a]. Heindel [Mul76]. Heinemann [RB82]. held [Bar73e, Rob72, Val77a, Val78]. Helenos [KSBW18]. help [BR88, CW80]. Helping [CM85, CBF23]. Hemisphere [Bry77]. Herman [Whi87]. Hermes [KG95b]. Heterogeneity [Not90]. Heterogeneous [Col87, MWB95, MS80a, SH98, WZF94, ZZWD93, AF02, CS02, EGCCM21, GCARPC^{+}01, HZ95, IHS^{+}14, JGC^{+}21, KSBL22, KA22, KTG20, KSH^{+}15, Li18, MUE23, PTU03, PMC05, POZ^{+}16, QC17, dRRGdC15, SSD11, VNGB08, ZLG08, ZCCB22]. Heung [XZ01, XZ03]. Heung-Seok [XZ03]. Heuristic [And89, Coo05, Mon96b, NGLL14, Wil74b, BGSG20, Bur16, RL14]. Heuristics [ARMMA18, LMK16, ROFGFRM16, SSRAH15, UCCPM19]. Heyden [Bar77c, Bar78b, Bar82b, Val79]. Hidden [BDG93]. Hierarchical [AS83, BE81, LCW98, LOS83, LS77, MSWC23, FG08, JPG^{+}17, KA22, LLJ12, NT84, SSKG22]. hierarchies [CA08a, FGNZ00, PZ00]. Hierarchy [AR93]. High [ACDP85, BPDL23, Cav83b, CG96, CDG^{+}98, CDFV12, CB72, FIL86, FM77, FN77, GH84, Har80a, HF73, JKRS85, JGT95, JZ93, KSH^{+}15, LQ93, Mer73, MW91, NM78, Nil90, Par75, Ped86, Pyl79, Rön07, RW04, SRS98, Sat72, SW86a, SR91, BGS18, Bra99, CSZO23, CCE99, CQH^{+}13, DHWZ14, EMRK20, Ell82b, FIÁSLSAR05, FMT04, Fra99, GA12, GIF01, GVL10, GZX^{+}21, HK84a, IMKN12, KCP22, KS10, Lev80, LZ10, Mad79, Mor77, NM06, PKN^{+}12, PGK^{+}10, SFC^{+}21, SC22, ScG09, SDF^{+}21, SAA^{+}20, VGF21, WW09, WSL03, ZWF^{+}22, Bar76b]. high-availability [DHWZ14]. high-definition [SAA^{+}20]. high-energy [SDF^{+}21]. High-error [Rön07]. high-fidelity [KS10]. high-latency [BGS18]. High-Level [Cav83b, Par75, FN77, JKRS85, JGT95, LQ93, MW91, NM78, Nil90, SW86a, EMRK20, Ell82b, FMT04, GIF01, GVL10, GZX^{+}21, Lev80, Mad79, VGF21]. High-performance [BPDL23, JZ93, RW04, CSZO23, IMKN12, KCP22, LZ10, PGK^{+}10, SFC^{+}21, WSL03, ZWF^{+}22]. high-precision [ScG09]. High-quality [CDFV12, NM06]. high-resolution [Bra99]. High-speed [KSH^{+}15, SRS98]. high-volume [PKN^{+}12]. Higher [BB95, JBCB79, Kat83a, GHBH05, Val77a]. higher-level [GHBH05]. Higher-order [BB95]. Highly [Bar78a, CLZ98, MM81, Pag79, ALF01, CARB10, DAJ^{+}15, NM19, SMGMOFM07a, SMGMOFM07b, TGCF08, UCCPM19, ZCN06]. Highly-structured [Pag79]. Hilbert [BG01, CWS07, Fis86b, Lin98a, LS96b]. Hill [Bar77b, Bar79a, Bra75, Ken77, Rog71, Val80]. Hints [Wai75]. HiP [APS^{+}11]. HiP-HOPS [APS^{+}11]. Hipec [LCC97]. Hirscheim [Her84]. Historical [RDC93]. History [LQ93, Bre02, GF78, TMM82]. HITAC [Hay87]. HLA [ATO10, KCK^{+}23]. HLH [CJ88]. hMod [UCCPM19]. HMRF [WY18b]. Hoare [Bar75f]. Hobbs [Bar77e]. Hobby [Wil80]. Hodder [Eme84]. hole [Car22]. holistic [BELS14]. Holland [Ald72, Bar72a, Bar74e, Bra80, Lan74a, Pit82, Val77a, Val78, Wal81a, Woo74]. Holography [DFW^{+}12]. Holt [Haz72]. home [HKC^{+}12, IS05, LM15, NDKL^{+}24, TLB^{+}18, Loe07]. home-based [IS05]. homology [HHPSS19]. homonyms [EMD13]. Honvood [Vel88]. hooking [BB10]. Hopfield [BL90a]. HOPS [APS^{+}11]. Horizontal [vO03]. Horn [WSWR23]. Horowitz [Bis86]. horror [SD75]. Horspool [Rai92, Smi94]. Horwood [Atk82b, Cor82, Lav78, Mar88, Rob82a, Sto88, Wal86b]. hospitals [SH22]. Hostile [Car81]. hosting [YMY17]. hot [ADZF21, DDF16, LMK16, OMGDG14]. hot-reprogramming [OMGDG14]. hotline [XDZ^{+}17]. HPC [BBK^{+}12, dSCdRS^{+}19]. HRV [GDW^{+}20]. HRV-PRV [GDW^{+}20]. HTEL [SM99]. HTML [JRGC20, MLV18]. HTTP [Mog04]. Hull [Ken77]. Human [CP96, Edw98a, Edw98b, HHK90, LBGA^{+}21, LBS78, Lin86, Pal79, Tra79a, Gal79, KJB11, NRUP21, OMM15, RBR21, SNK21, ZYW^{+}20, Wri98]. Human-Computer [Pal79, Edw98a, Edw98b, Wri98]. Human/Machine [LBS78]. humans [LBGA^{+}21]. Hundreds [Str95]. Hungary [Val78]. hunks [ZYW^{+}22]. Hunt [Cou84b]. Hunter [Rob82b]. hurricane [CGH^{+}04, TWS^{+}22]. Hutty [Bis81a]. Huxtable [Han77a]. Hwan [XZ01, XZ03]. Hybrid [BP97, Gom78, Kra97, Mon96a, Ono93a, RT91, XAN07, AA23, CCLN22, CLCC15, CLD^{+}17, FR09, GOQ16, HC16, JP22, LG19, MGBO22, SMLS23, VS18, ZZJZ24]. hybrid-cloud [CCLN22]. hybridized [ASA^{+}21]. hylomorphisms [LV20]. hyperactivity [LCT^{+}21]. hyperanimation [Hun00]. hyperbolic [NNR18]. Hypermedia [WW95]. hyperparameter [RaMB^{+}24]. Hypertext [SCGP92, BR88, SM99]. HyperTree [STH97]. Hypervideo [Hun97]. hypervisor [RSLAGCLB16]. Hyphenation [MMN79]. hypothetical [NSW77].

I-like [Neh79]. I.A.G [Flo73]. I.E.E.E [Mer74]. I/O [KJHG10, WBB15, Yoo96, ZCCB22]. I/O-related [ZCCB22]. Ian [Edm82]. iAPX286 [Le 88]. IAs [HLW08]. Iava [Ric00]. IBFET [AML20]. IBM [BB75, GA12, JDBP04, PKN^{+}12, RS76, UGBW91, Haz71]. IBM(R) [OM16]. ICARE [KMB98]. ICC [CDG^{+}98]. ICCCN [WL03]. ICE [PT17]. Icecream [Lin86]. ICL [Bar78c, EP79, Far74, Iza80, MBB^{+}86, Oes71, REC75, WQ72]. Icon [FH92a, GT93, Han80b, JG94, LC86, Nil90, PT00b, WG92a, WG83]. Iconic [RS93b]. iDARE [TM14]. IDE [AÇCT22]. Ideal [Des92, GMM90]. ideas [CBC00]. Identification [CS22, Hug93, WBPR20, BZD17, DB21a, GH19, JP22, MM82, WY18b, vdMF13]. identifier [WSWR23]. Identifiers [LV73, Sit79, Par78, Sco77a, Wu01]. Identifying [CCM96, CK15, CS17, IAA^{+}21, aSJW^{+}24, Yan91, ZHZ^{+}14, ADZF21, ZZL24]. identity [BLC19]. IDEs [ZCO13]. idioms [PZ00]. IDL [Atk77]. IDMS [Wya84]. If [Gre80, Wil74b]. IFCnCov [PPPP23]. IFIP [Lan74a, Val77a, Val78, Wic72b, Bar72a]. iFogSim [GDGB17]. ifthenelse [Atk79d]. IGES [Kah95]. ignoring [Thi12]. II [GH84, MPC^{+}19, Pur76, RDC93]. III [Rue93]. IKBS [Lei85]. ILDJIT [CARB10]. Ilem [Wal86a]. Illiac [Kar76]. Illustrate [Ric76]. Illustrating [PCBE96, Ree78]. illustration [LWJ^{+}21]. illustrative [MF08]. ILP [MM01]. Image [DDB^{+}18a, MBG19a, SRC^{+}18a, VS88, WY18a, ABA20, CI03, GSS^{+}20, IAA^{+}21, dSJCM16, KBBS05, KEL^{+}21, KKA^{+}17, LCT^{+}21, SDKS16, SAY16, Sta07, XAN07, YHGC20]. image-aware [dSJCM16]. image-based [XAN07]. Image-understanding [VS88]. Images [CT92, AF99, AFF02, BNS18, SAA^{+}20]. imaging [GDRV20, KCH08]. imbalance [ZNWS18]. imbalanced [LZK^{+}23]. IMDAC [ZZJZ24]. imitation [OMM15]. Immediate [Lar78, MT84b, New82]. immersive [WIYC20]. Impact [Aji95, Buy21, LTK^{+}20, GRFFGC^{+}21, GVBG22, HJ08, LPP09, MAA^{+}22, TTC^{+}13, UFR18, WAML12, ZSH^{+}23]. Impacts [ZHZL22, NDKL^{+}24]. imperative [CIM22, CIM24]. implantable [BCC^{+}22]. Implement [BF80, OM96, UGBW91, GKLM79, HIR06, ZXT^{+}17]. Implementation [ARV77, AL82, AN95, AMS92, AP84, AvdSGS80, Bai85b, Bat74, BH87, BCP71, Car85a, CGK89, CS91b, CVV97, CG95a, CDKK85, CDV88, Clo85, Com78, CL95, CDH^{+}76, Day00, Deb93, DO91, DW90, DMW88, EE90, Fen98, Fid88, Fis84, Fis86a, FHJ94, Fos89, GR91, GRS74, GT93, GF78, Han87b, Han89b, Han77c, HHR93, HHZ^{+}95, Har71a, Har84b, HA90, HS77, HOS85, Hop86, Hud72, HP83a, HP83b, HC87b, HH82, IB13, IK15, Jia97, KS98, KM83, Kin93, Koo87, Kos90, KH96, LL96, Lar75a, LPT78, LPT82, LFW96, LLK04, Lei85, LKK^{+}18, LHS^{+}95, LM76, Lit93, LHC97, LQ93, Mac79, Mac77b, MWB95, Mal83, Man88, Mar79, MRR^{+}08, Mat80, Mau92, MW93, MW91, MS96, NS79, Nee77b, Neh79, NW85, NP98, OW83, PCBE96, Pas87, PS80, Pik90, Poo71a, RK91, Rei84]. Implementation [RS90, RH77, RC89, RB81, Ros77, RT91, RS76, Sal81b, SS95, SW90, SK03, SWN94, SL78, SF98, Shr79a, SHC74, Ste98, SO77, TT74, TM95, TBA89, TTH97, Tur79, VWB91, WG83, Wan79, WW95, WS94a, Wir77b, Wol92, Woo71, Wre88, Yip82, Zel72, AKS06, And82b, Ano23m, BGM99, BCC^{+}22, BH01, Bea78, BL15, Col72b, DPK12, DHGR92, DCA04, DM11, DSW82, DFRR15, Eba20, GOQ16, GKS03, GP01, HJ14, HK84a, HE82, Him00, HP11, Hol77, HC99, HKC^{+}12, IS05, IIL17, JZ10, KCYY12, Kat17, KF02, KMB98, Ker82b, KMY^{+}05, LG99, LS15, LCZ08, LS16, LLJ^{+}23, Man18, NSKK83, NK07, Par85b, PDPQ22, PNM^{+}20, PT00b, Rai84, RCC17, RR05, Rei99, Rob82a, SGA21, STB14, Sav04, Sav11, SE11, SM01, SS09, TH01, UFS99, WWB03, Wet77]. implementation [Woo74, YWN^{+}00, YCY03, YZYL07, ZC01, ZWML14, ZXX24, vGB01, Hay80, Bar76a, Wal86b, Woo74]. implementation-based [SE11]. Implementations [BdJ80, DJM97, FL92, Jal87, LS97, ÖS96, SC94, TV96, WW89, Yas94, Bri84, KSH^{+}15, RT78, SSM11, SZ00, ZZH^{+}24]. implemented [PKN^{+}12, Zel72]. Implementing [BCHR81, BM98, Bis79c, BRL^{+}15, CK99, Cav83b, CP07, Dew93, Dun91, FP97, Fil98, FN77, GR79, GR92, Ham95, HUS^{+}91, HMPT89, Jaa95a, JB84, KRO93, KA87, Lak80, LS84, LT90, MGW82, MJ98, MG13, MDP96, PLFIVRAR23, PDC^{+}98, PH86, Sal79b, SZ01, Bas00, BHK^{+}04, CMCL03, CB00b, Duc11, JJK^{+}12, Mor77, PMP^{+}16, Sav07]. Implications [LS96a, CKRC20]. Implicit [Per85]. Importance [YAVHC21]. imprecise [WM20]. Improve [DCA82, BJP^{+}00, BLS03, CZ04, CSM^{+}16, CLC99, CMTCC^{+}17, DW13, MRZ15, MC02]. Improved [BY89, CCM96, CLP^{+}09, Com78, Ein88, Fen96, Hol88, LDI98, Ayc15, CBB20, GMDM17, IAA^{+}21, KSSA23, KMB^{+}21, Mof99, SMT^{+}18, ZG06]. Improvement [Fre78b, MT78, CGP^{+}06, GW04, JTG^{+}11]. Improvements [BCHS98, Deo00, Ree71, SS03]. Improving [BCPL13, BR95, CGZ^{+}20, Coh73, CALL18, FCR^{+}09, Han83a, HL02a, Lev95, LNhCW16, MZ00, NNLR17, QM13, RSLAGCLB16, RMZ17, SRGCPB^{+}09, SH17, Str81, WKJ15, You81, CCLN22, CSTL19, CDF^{+}24, DSD^{+}05, GAQO23, HC12, HYH15, PDROFRM13, ROFGFR^{+}16, ST14, ZGGL23]. IMS [SMGMOFM07a, SMGMOFM07b]. IMS-based [SMGMOFM07a, SMGMOFM07b]. In-Core [REC75]. In-memory [CMTCC^{+}17, ACM^{+}15]. In-Situ [RGK99]. IN-Tune [CGR00, RGK99]. in-vehicle [XZX^{+}21]. Inaccuracies [PF88]. inclusion [SYXZ14]. inclusion-based [SYXZ14]. incomplete [ZLTX18]. incomplex [SS07]. inconsistency [FBB^{+}14]. Incorporate [Mös88]. Incorporating [Al 13]. increasing [ROFGFRM16]. Incremental [Abe07, BS90a, CAFH94, CW01, CW97, Dan90, Dun93, FBB^{+}14, FHS92, Hol89, KLLK98, KW92, SN90, Wil83, GB23, Hug82, LSZ16, RO77]. Incrementally [MRNL92]. indent [KY77]. Indenting [MJ83, Mat83a]. independence [Knu11]. Independent [Bla92, FH82b, HR96, HS89, Kob77, Nee77b, Ray75, RRP95, SMM^{+}84, Thi87, AvRAF09, Atk77, BHMV09, CP76, Eng06, FR09, Han99b, Hum76, Jok89, JGSG^{+}21, LCW98, MP82, SP79, vdWCB17]. Index [AML20, AM10, CH88, Qui83, BCF00, DGM19, GSR17, PM18]. Index-based [AML20]. Indexes [AB89, ACM^{+}15, KL16]. Indexing [CRR94, Vis76, ALG23, KEL^{+}21, Mos06]. IndianaMAS [BDMP17]. indicator [LCY07]. Indicators [Atk79c, WLS^{+}21]. indirect [UW99]. individual [ZZLL24]. Individuals [Car85b]. indoor [NAU^{+}21]. induced [ZLZ^{+}19]. Inductive [Dro85b, FCR^{+}09]. Industrial [SFB13, SMT^{+}18, Web87, DBvG^{+}23, DMR^{+}22, FVF^{+}18, KBPM^{+}20, MKE18, MGS^{+}20, RSW^{+}23, RGC^{+}21, WZLN08, WYAZ15, ZHZL22]. Industry [Cou92, Kot96, BCPL13, CEMS24, DBOT23, Eba18, ENP^{+}24, GMC^{+}21, JPM23, KRB21, ACK^{+}22, APR22, CJTK22, DAC^{+}21, SSKG22, ZWQC22]. indx [AB89]. inexperienced [The77]. infection [SKD^{+}22]. infer [CA18]. Inference [APS95, DF87, MK90, PNK^{+}23, RaMB^{+}24]. Infinite [Har80b, MH05]. Inflected [RS93a]. Influence [CPHS83, BAPC23, CCQ16, SdLJMP21]. influencing [Eba18]. Info [Ano16j, Ano16k, Ano16l, Ano16m]. Informal [Geh82, Bar74e, Bra80]. Informatics [vdRW79]. Information [Ano16j, Ano16k, Ano16l, Ano16m, Ano16q, Ano16n, Ano16o, Ano16p, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l]. Information [Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Bar76b, Blu86, BK86b, CMH91, HL98, Hut76, Mac96b, MBO97, Mar86, Pet77, Wal81c, Wil74a, Wil76, WBPR20, YAVHC21, BVGVEA11, BDLM04, BLNU15, HB11, KBH^{+}03, LLN16, LQS^{+}24, LGP^{+}11, MR07, NMH^{+}24, NR04, PTU03, Pol01, Rob72, ROFGFRM16, RMdL12, SSD11, SI10, TRGA18, ZLTX18, ZYF20, dMFÁE17, Ald72, Bar74f, Bis84, Mad82]. informed [WLZ^{+}22]. InfoSphere [GA12, OM16]. infotainment [MPC^{+}19].

Infotech [Bar76b, Bul72b, Hut76, Rog73, Wil74a, Wil76, Bar79b]. infrastructure [BAM^{+}20, CPMAH^{+}20, GAAA22, HPK^{+}12, HKWZ00, KDA20, LZD20, MAR^{+}16, MST13, POZ^{+}16, RCA^{+}19, SGCM11, ZHO^{+}19]. infrastructures [BM03, BLC19, JAkM^{+}21, SDG^{+}20]. Ingest [Sil92]. Inheritance [APS95, Dew91, JR92]. initialization [FK16]. Initializing [McC83]. Initiation [RMMLSME14]. Injection [DJM97, BB10, GWY^{+}11, HLR20, TDDE15, ZYYC12, Cla98]. Inline [CMCH92, CHT91, HSY^{+}20]. Inliner [DH88]. inlining [ZA07]. Innovative [GVBG22]. Input [Com83, Dew84, DS94, Gra81, Kur78, MT78, Pyl79, TR77, AV84, AWNS18, Han83b, JHKS19, Wic72b]. input-output [Wic72b]. Input/Output [Pyl79, TR77, Kur78, Han83b]. inputs [GKBK16]. Inquiring [CC77]. ins [BN13, FD92, GK14, Kar14, SMM13, ZCO13]. Insecurities [CA86, WSH77]. Insert [Thi89]. insertion [MFYiA01]. Inspection [Doo92, Ber82, dSMLP22, SAA^{+}20]. inspired [ASA22, BBH^{+}22, TMGR^{+}22]. Installable [CE97, OSW92]. Installation [Gri82]. Installing [Eva71]. instance [JXG^{+}21, KHC^{+}19, LW14, WKG^{+}13, XLZ^{+}20, ZQWL24]. instance-based [ZQWL24]. instances [SBS13]. Instruction [AG95, ABSS98, CAFH94, MAF91, Pas87, Wha93, CW08, Fra06, HW15, IMKN12, KL21, Was12, YLP^{+}11]. Instruction-level [Pas87]. instructions [CL23, GYCL16, LM22, PACK07, PKH07, YLP^{+}11]. instructive [SD75]. instrumentation [BMR00, BMTA16, CCC^{+}16, YMH16]. instrumentations [MK18]. Instrumenting [LS75, SSS^{+}02, AE14]. insurance [MVM^{+}22, TWS^{+}22]. INSWF [AAB^{+}21]. Integer [Ber86, Fro81, GW96, Jam80, Nee77c, Par85a, Wic79, CIM22, Fen02, JT00, PM17, Win02]. Integers [Sam71a, LB15, LBK16]. integrate [NZPWR22]. Integrated [HL20, HW98, LD87, LTW^{+}21b, MXYQ86, O’N88, PL91, Sav11, Söz15, Tay83, dCGG13, ACKS09, BMM^{+}18, BIA24, CNRB13, CSS15, CW08, DGM^{+}22, DDMD20, FT01, FPT07, HJC00, KS24, LS15, NS08, SB21, SMGMOFM07a, SMGMOFM07b, TM14, XLZ^{+}20, HJ14]. Integrating [ADDM84, BS90b, Bro86a, CFL^{+}98, UDS^{+}07, vDD11, BRTT09, BDL09, dSCdRS^{+}19, KAZ13, LHFL07, MCGS08, RBR21, SDB^{+}22]. Integration [BH92, CMF^{+}98, CSIL93, LC86, Lob85, SZSB19, YCY03, ARCN^{+}06, BFRFB22, FRBRF19, FCBF^{+}21, FSRF^{+}22, FFRF19, FFRFS19, FLSCC15, KS01a, KBPM^{+}20, LG19, MVM^{+}22, MP13, MGL19, MBGC21, MSK^{+}24, Mus17, NR04, PCBR18, SH22, SGCM11, ZSAM23, ZJY^{+}15, vGPB10]. integration-oriented [vGPB10]. Integrity [Sha80, AA19, CR18]. integrity-related [CR18]. IntegrityCatalog [CR18]. Intel [HK84a]. intelligence [GLMS18, LW19, MS18, QTV23, SRRFGC^{+}10, Cam85]. Intelligent [AAB^{+}21, Ano13, BS90b, MGBO22, SSV^{+}20, Sef97, YOH15, BFPAGS^{+}08, DDB^{+}18b, JCL85, PKK12, ZZJZ24]. Intelligibility [WKS^{+}98]. intensive [TGAS22, TBT23]. Inter [Bar80c, Mar86, RNS^{+}16, Str81, Val76a, Wid90, NZTL^{+}23, GB14]. inter-app [NZTL^{+}23]. Inter-Client [Wid90]. Inter-Cloud [GB14]. Inter-JVM [RNS^{+}16]. Inter-module [Str81]. Inter-process [Mar86]. Inter-task [Bar80c]. Interacting [Daw77, Rei90]. Interaction [Edw98b, ZGL^{+}23, Edw98a, Wri98]. Interactions [AP95, Gan82, JK14]. Interactive [AS83, ASH73, Bat74, Bec91, Bra75, Bro86b, CW94, CS97, Com83, CDH^{+}76, CSIL93, GB87, Ham84, HS77, Jaa95a, Jen89, Jon71, Kin71, Koo87, Kue95, LNW82, LFW96, Les72, Lib93, Mil74, Mul76, NHP81, ORT81, Org81, PSV85, Par79, PN83, SB83, SW86a, SN90, Tha84, Thi93, WW95, WOKT81, WR77, vdRW79, AP85, ALF01, Bar71, FKD14, Har82, HL87, JAJB04, NW84, Ree82, RSW^{+}23, VV84, XXJS18, Rog74]. interactivity [HYH15, MA01, TCM07]. interception [AGG06, Kan18]. Interchangeability [Str82]. Interchanging [ÖS96]. interclass [SJK^{+}21]. Interconnecting [CS97, Col87]. interconnection [SDG^{+}20]. interest [FKL^{+}13, YZL^{+}22]. interest-aware [YZL^{+}22]. Interface [AC80b, Bad98, CD82, Cha88, FH91a, Han76c, HUS^{+}91, HHK90, Hof89, HM90, Hug88, KRO93, LDG^{+}96, LD95, Lop89, Pal79, Pal80, PA91, RDC89, SHR80, SM01, SWPS89, Sos95, Str83b, TS81, UGBW91, WC81, WN88, WG92b, BCL13, CYW^{+}15, CHS^{+}05, FT79b, HK06a, KBBS05, KV98, Kot01, KKA^{+}17, MM02, MRG^{+}19, Sne78, VC21, BM98, PZ00]. Interface-Application [WG92b]. Interfaces [GB87, Hol93, Jaa95b, JI80, Lin86, Pow87, SMFBB93, BB99b, CZL21, CRGIP15, MKD^{+}22, SCT02]. Interfacing [vMC77]. interference [CHT98]. interim [CLP^{+}09]. interim-support [CLP^{+}09]. Intermediate [GF84, HW78, Han04, KKM80, MFH10, SHGG16, SBS13, YHK^{+}22]. Internal [AW93, Jon72, Oes71, CPW73, SBFF24]. International [Bar79b, Cou85a, Pra96a, Pra96b, WCK11, YLM^{+}05, PL14]. Internet [JGSG^{+}21, LFGCGCRP14, ALD^{+}23, BJ21, BSM^{+}21, CTLL07, CHCC07, DMR^{+}22, GDGB17, IBA^{+}21, JSPP21, JAA^{+}20, KPGH02, KSST22, LCT^{+}21, LLW98, Lv21, MA01, MTM22, NSM^{+}22, NRUP21, PHB21, RS21, SGAS21, SRCP19, SZSB19, SWBS17, TH01, VSID17, WBPR20, YCY03, ZKW21, ZWQC22, ZZT^{+}21]. Internet-based [KPGH02]. internode [CSTL19]. interoperability [BGS20, ENP^{+}24, KCK^{+}23, MCGS08, PBGM18, SH17]. Interoperable [MPBH13, BCC^{+}22, GRVC23, Kap13, RMMH^{+}23]. interplay [ALD^{+}23]. Interpolants [FR78]. Interpolation [WJ93]. interpolative [NT20]. Interpretation [CST75, DF87, ELRV93, Hef82, Kli81, KKM80, Lic77, Fra06]. interpreted [BJP^{+}00, SS09]. Interpreter [ARV77, BBM84, Bro81a, Bud89, CJ88, Hal82, HOS85, Jen89, LOBF88, McD87, MD88, MM80a, Pag79, Sch89a, Gai82a, GMO01, Ric00, Sny08, Yuv79b]. interpreter-minded [Yuv79b]. Interpreters [Pag88, EGKP02, HATvdW99, Ree82]. Interpreting [MR05, AA14]. Interpretive [KFJS88, TR77]. Interprocedural [AS97a, MW93, RG89, OY10]. Interprocess [BMS83, KH96, PR90, Sau88]. interrupt [RA87]. interruption [JH03]. Interrupts [EBD^{+}74, Hun80]. Interscience [Dav74, Jac71, Nic72, Wis74]. intersection [LBK16]. intersections [KL16]. Interval [CM82, WS94b, AHRR22a]. interval-based [AHRR22a]. Interventions [WBN^{+}20]. Intraprogram [Flo72]. intrinsic [KKdC^{+}23]. introduce [STA09]. Introducing [AA19, BBS11, CDRV03, NM78]. Introduction [BCSW20, BC21, BCP19, Coo96, Die98, Kat83b, TMM82, WCK11, Ano79a, Atk78, Atk79b, Bar74e, Bar77b, Bis79a, Bra80, Coo08, Edm82, Eve73, HW77, Lon88, SFB13, Tho77, Ros74, Bar75c, Bar77c, Wan82, Wel72]. Introductory [vdRW79, Cor82]. introspection [CKW02]. Introspective [Mus97, Val00]. intrusion [GBG^{+}14, PRA^{+}06, RPCS08, WWB03]. intrusion-tolerant [PRA^{+}06, RPCS08]. intrusive [CKW02, CGR00]. invaded [LZWX22]. Invariants [CK78, Sav06]. invasive [JSC^{+}10, RGK99]. Inventing [Har80c]. inventor [CY01b]. inverted [PM18]. Investigating [BLS03, WBB07]. Investigation [RB91, SW91, GKWS11, HKA12, IAA^{+}21, Lin98b, ZHZL22]. Invited [dSMH13]. Invocation [LT91, RK89, DMD^{+}06, AV05]. invocations [IH01]. involvement [BR01a]. Involving [BH94]. iOS [WWCW19]. IoT [SWBS17, ASP^{+}19, ARK21, BAM^{+}20, BT21a, CFS^{+}22, dSCdRS^{+}19, DGM^{+}22, DDB^{+}18b, FVF^{+}18, GAAA22, GMGDMB19, HSRB24, IMG^{+}21, KRK21, KBPM^{+}20, KJVS21, LSK^{+}18, NEP^{+}17, PPPP23, SAYM24, SWY^{+}21, SRC^{+}18b, TGAS22, TJB^{+}19]. IoT-based [PPPP23]. IoT-device [KBPM^{+}20]. IoT-edge-fog-cloud [DGM^{+}22]. IoT-fog-cloud [KRK21]. IoTSim [JAA^{+}20, SSK^{+}22]. IoTSim-Edge [JAA^{+}20]. IoTSim-Osmosis-RES [SSK^{+}22]. IP [SBcC07]. IP-based [SBcC07]. iPhone [BVB^{+}12]. IPIP [Woo74]. IPTV [RSRCGC15]. iQuantum [NUB24]. IRONMAN [Wan79]. irreducibility [SW12]. Irregular [CDG^{+}98, HMS^{+}95, HY20]. ISA [AW93]. ISAMadapt [AYdS^{+}06]. ISBN [Ano87a, Ano88c, Ano88b, Ano88a, Bar81, Bow88, CO88, Cav83a, Cor99a, Cor99b, Dea86, Edw98a, Edw98b, Gru83, Llo82, Lon88, Mar88, Mee87, Ree84b, Rop88b, Rop88a, Sto88, Val77a, Val78, Vel88, Wal83b]. ISDMS [BSRS85]. ISO [Ten85, Wu01, Wu02]. Isolating [JWTG11]. Isolation [HC79, CBB20, KKL17, SO07]. isomorphism [KH04]. ISORC [Obe11]. Issue [Ano16j, Ano16k, Ano16l, Ano16m, Ano16q, Ano16n, Ano16o, Ano16p, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l]. Issue [Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, BC21, Cor08, DDB^{+}18a, KH12, MBG19a, TBT23, WCK11, WY18a, WBPR20, BP11, BN13, BCSW20, BCP19, CSZO23, GK14, ISPB21, KSRR17, PL14, RWJ^{+}17, WY18b, XDVB24, ZWC21, DAC^{+}21]. Issues [FP97, HHZ^{+}95, HL92, Her84, KGP96, KH96, LT91, Mad95, NS79, RC89, Wol92, AW04, Bas00, DFRR15, GW04, MFB^{+}02, PRP^{+}23]. Istio [LTK^{+}20]. Italian [BCC^{+}22]. items [BMS21]. itemsets [CLP^{+}09]. Iteration [NW85, XXJS18]. iteration-based [XXJS18]. Iterative [Blu86, MAW^{+}16, LLLY19, SRS18]. Iterators [Gra96, Ian90, Mes96]. ITSs [LCGS17]. IUP [LDG^{+}96]. IUP/LED [LDG^{+}96]. IV [Ree73, Ree75, Ben77, Kar76].

J [Ald72, And78, Bar71, Bar72b, Bar75f, Bar77b, Bar78c, Bar78b, Bis81b, Col77b, Han78b, HW77, Hun72, Hut74, Ken77, Lan75, Nee77a, Ree82, Ree84a, Roh77a, Sau88, Tho74, Val76b, Val78, Vel88, Wal81a, Wel72, Whi87, Wil74a, Woo74, Wri98, KS08]. J.-D [Nee77a]. J.-P [Whi87]. J2EE [LLM05]. J2EE^{T M} [JDJ^{+}06]. J3DV [FMA02]. J9 [WKJ15]. JAC [KT01b, PSD^{+}04]. Jackson [Bar77d, Hug79, Rya80]. James [Mer74, Ree75, DDDF17]. Jane [HW77]. Janson [Dea86]. Janus [CPW74, Deb93, HW78]. JaRec [GCRD04].

JAS [KS01b]. JATS [CFD^{+}22]. Java [CY01b, ACSK23, ABL08, AV05, BMR14, BVGVEA11, BVGVEA13, BAF03, BBGP01, BBG04, BDLM04, BDP02, BKP^{+}22, Bin06, BSMV09, BHK^{+}04, BS98, BS00, BZD17, BCL^{+}06, BE02, Cai99, CV03, CC01, CCT01, CY01a, CMS07, CS04, DHS02, DDDF17, DIS99, DC03, Die98, DCA04, ET07, EvG04, FMA02, FRGPLF^{+}12, Fer13, FKR^{+}00, FDD20, GvRN^{+}11, GCRD04, GCARPC^{+}01, HKM^{+}09, HC98, HC10, IH01, JJL17, JMM03, KMS98, KS01b, KT01b, LYM04, LC05, LMK16, Ler02, LHFL07, LQ99, LPA13, Man01, MZB00, MZC10, MJ99, OW16, PJJM21, PZ00, PMP^{+}16, PNM^{+}20, Phi99, PDPM^{+}16, PKC^{+}13, Ric00, RBRB23, RPP07, RJGH06, SH03, SPPH10, ST04, SZ00, SKM01, Thi99, TWNH12, VED06, VDMW06, VB01, VP05, WWJ07, WW09, Win02, XAN07, YME05, ZLG08, ZWSS15, vD99]. Java-based [CCT01, FMA02]. Java-type [FDD20]. Java/CORBA [GCARPC^{+}01]. JavaBIP [BMSZ17]. JavaCC [GN16]. JavAdaptor [PKC^{+}13]. JavaOS [HPB^{+}00]. JavaScript [FBV22, HRM00, JGSG^{+}21, KRR19, PLR18, PKB23, RW17, Ryu16, VB14, WXR16]. JavaServer [DBH04]. Java^{T M} [CHS^{+}05, DM07]. Jay [SMLS23]. JCrasher [CS04]. JDAS [XCL^{+}18]. JDB [WN88]. Jenkins [Ken77]. Jerry [Mul76]. Jersey [Lar71]. JFSL [CCG14]. JGAP [CCT01]. JGRIM [MZC10]. Jgroup [MMHB08]. Jgroup/ARM [MMHB08]. Jim [Ham79]. Jini [GH03, YCY03]. Jini-based [GH03]. JiST [BHvR05]. JIT [ZXX24]. JJTraveler [vDV04]. JnJVM [TGCF08]. Job [BJ72, Han76c, Lar71, Par75, Ray75, Sch78, vdBT77, FCYL18, GF78, SAL^{+}04, Bar75d]. jobshop [BDA20]. John [Atk83, Bul72a, Bul73, Edm86, Hut74, Ree75, Val80, Wis74]. Johnson [Val76a]. Johnsson [Bar77e]. join [PP16]. Joining [HC97a]. Joint [SGA20, JSPP21, JPM23]. Jointgraph [XHP^{+}21]. Journal [Ano76a, Buy21, SB22, Ano81n]. journaling [HC12]. Journey [CKCG23]. Joyce [Han87b, Han87a, Han89a, Han89b]. Jr [Bar76e, Haz72, Lav77, OW16]. JR-like [OW16]. JRebel [KV14]. Js of ocaml [VB14]. JSetL [RPP07]. JSON [BMC17, KL24]. JSP [EV89, RS87]. Jubilee [Buy21, SB22]. Julian [Dav74]. Juliet [ACSK23]. Julius [Rob72]. July [Bar73e, Woo74]. JUMBO [Ric76]. jumps [UW99]. June [USE85]. Just [ZYW^{+}22, Gre80, KRB21, LYM04, LMK16, ZQWL24]. Just-in-time [ZYW^{+}22, KRB21, LYM04, LMK16, ZQWL24]. JVM [RNS^{+}16, SSB^{+}16].

K-MING [HM18]. K-nearest [MGT20]. K6 [Le 88]. Kaare [Mee87, Ree84b]. KAL [JSC^{+}10]. Kapur [Haz71]. KDF9 [BCP71]. Keith [Mar88]. Kemeny [Bul72a]. Kenneth [Atk79a]. Kent [Bro71]. Kernel [CS91a, CTLL07, Cor88a, FHJ94, HWS^{+}88, Hug93, LGC84, NS79, RS82, Str83b, TY14, AIB02, BV06, CSTL19, DD21, DD10, DHGR92, HBC15, JGS^{+}08, JSC^{+}10, KGL06, Le 88, LSAF16, MSK01, MS18, NJ11, NAGL10, RLPA18, Ter86, TXHL18, WAML12, dOdO16, KM13]. kernel-assisted [JSC^{+}10]. Kernel-based [TY14, KM13]. Kernels [Kue95]. Key [Tal71, BPM93, LFHL22]. keyboard [Gai82b]. Keys [FP82]. Keyword [Gra81, GK08]. Keywords [Coh98, GLD^{+}21]. Keywords-driven [GLD^{+}21]. Kibitz [Lib93]. KidPad [HBD04]. killer [McI99]. Kim [McG89]. KIND [MGGS18]. KIND-DAMA [MGGS18]. Kinect [MGGS18, PT17]. Kinect-like [MGGS18]. Kinect-type [PT17]. kinetic [XMTL21]. Kitrace [Kue95]. Klava [BDP02]. KNIME [JGB15]. know [NHTT08]. Knowledge [WBPR20, Bas00, CMT17, CSMML12, GT00, GLS^{+}23, LHB18, WSWR23, ZB23]. Knuth [Llo82]. Knuth’s [Hoa73]. Kolmogorov [Cox76]. Korean [KHH^{+}15]. Kroese [Nee77a]. Kronos [Ano76c]. KubCG [EGCCM21]. Kubernetes [BSNB20, EGCCM21, LTK^{+}20]. Kupka [Wal82]. Kurtz [Bul72a]. Kwon [XZ01, XZ03]. Kwyjibo [CA08b].

L [Atk79a, Bar71, Bar73b, Bar74f, Bar78d, Bar82a, Bar82c, Bar84b, Bar84a, Bis84, Ear77, Inc86, Jon74, Lav78, Mad82, Roh77a, Rop88b, Val77a, Wil72, Woo74]. Lab [PT17]. labeling [BG01, CCQ16]. laboratories [MCGS08]. Laboratory [Lin79, LOS83, Orm77, PBW78, Bar76a]. Lake [Val78]. lakes [KHS^{+}24]. Lakewood [Bar78d]. Lalr [Gro90]. Lambda [JL91, JPL03]. LAN [SBcC07, Yas94]. landing [LLK04]. Lang [Mul76]. Lang-Pak [Mul76]. Language [Abb89, ACDP85, AP84, AO88, Atk77, Bar75d, BR95, BW71, BCL^{+}94, BK86a, BE81, BdJ80, BDS^{+}92, BY90, CCPR91, CC73, Cav83b, CC77, Col81, Coo96, Cor88b, CE84, CP76, EG84, Ell79a, EBD^{+}74, FL92, FM77, FN77, FYP93, Fox87, FF80, GM85a, GR79, GC84, HW78, Han87a, Han89a, Han94b, Han80b, HHR93, HG84, Har85, Hay83, HG89, HP83a, HC87b, HMS^{+}95, IdFF96, JGT95, Jen89, Jok89, Jon83, Ker82a, KKS88, KGP96, KO91, Kin93, KW92, KD83, Koo87, KvEP95, KG95b, KNPS88, Kos90, Lea77, LPT82, LOBF88, MS74a, Mac79, MS74b, Mar79, Mei80, Mei81, MW91, Mul76, MB97, NS79, Neh79, Pag84, Pal76, Par75, PJ76, PSR83, Ped86, PCM83, Plu77, RTL^{+}91, Rey87, RC89, Rob83b, RB81, RT91, RW12, SW86a, Shr79b, SMM^{+}84, Sti78]. Language [Str83a, TS81, TDH97, TBA89, TAAT84, Wad85, WG92a, Wal81c, WOKT81, WB79, WBK91, Wes83, Wex81b, WKS^{+}98, Wir77c, Wir88b, Wit82, WBS82, WR78, WLL98, Zel72, dSC16, AKW79, And82b, Ano76c, Ano80a, AM00, AFFR08, Bar81, Bla04, BFRFB22, Bre02, BFNP08, CL19, CW01, Day00, DGPT14, DM07, EL05, FG14, GOQ16, GMO01, GA12, GN02, Gou86, Haf13, Han81b, HMS88, Haz71, HK84a, HRM00, HGWBS75, Hoh04, Inc85, JB07, JP79, KA13, Lev80, LvLS84, Mad79, MGP03, Mor77, MSB18, PSTV10, PL08, PPA20, PT00b, Rei84, RZ17, Sny08, SHGG16, Sto05, TV09, The77, VV84, Wal86b, WGM08, Yi12, YHK^{+}22, Zdu07, ZCN06, ZWSS15, dB00, vdWCB17, Han04, KU97, SM99, Bar73b, Lar71, Wal83b]. Language-based [KW92, WBK91, CW01, DGPT14]. Language-independent [CP76, Jok89, vdWCB17]. Language-Sensitive [Rob83b]. Languages [AH85, BJ72, Bar76b, Bec91, Bee82, BT76, FIL86, FS11, Fle90, Gel75, GG96, HGW94, HZ94, Kaw80, KV98, KKM80, NM78, OW89, Ono93a, Par79, PS81, Pra80, Pyl79, Ray75, RW81, Sat72, SW74, SAN^{+}81, Tur79, Wal81b, Wan79, Wel78a, Wet80, BAPC23, Bar74c, Duc11, FS13, Glü12, GS06b, GP01, Ham79, Har82, JM10, KW17, Lan74a, Lan74b, MPBH13, MGG^{+}09, MSZ^{+}24, Mus79, Nie79, OMGDG14, Ozk18, PVAHRG^{+}15, PMC05, Ron99, SSB^{+}16, Sav07, SHIS99, SK03, SC14, SS09, SZ09, UN19, Val77a, Wu00, Atk78, Bis86, Lan74a, Sto88, Wal82]. LANSF [GR91]. LARA [CCC^{+}16]. Large [BT89, BCP71, Coh98, Com79, DLP85, DDZ94, Fin97, Fit77, HWS^{+}88, HG81, HP88, Hos98, Jal82, LP86, LK93, MN80, REC75, RBRB23, SATV22, ST77, Van82, You81, ZZWD93, ZD95, AF99, AKL^{+}09, AZS19, Bar74d, BCLF^{+}07, BTZ07, CRC18, Deu99, FMNW04, GLL20, Gu05, HB18, HGK^{+}19, HCG^{+}16, KEL^{+}21, Lin98b, Mos06, OY10, PK11, SYG^{+}18, SSS^{+}02, WWCW19, WZH01, WHS^{+}00, XP23, XYW^{+}24, ZZKA17, ZGL^{+}23, vGPB10]. large-alphabet-oriented [Gu05].

Large-array [MN80]. Large-Scale [HWS^{+}88, RBRB23, SATV22, AKL^{+}09, AZS19, CRC18, Deu99, FMNW04, HB18, HGK^{+}19, KEL^{+}21, PK11, WHS^{+}00, XP23, XYW^{+}24, ZZKA17]. Laski [Roh77a]. latencies [WAML12]. latency [BGS18, CWC^{+}21, DDD16, PKN^{+}12, RAN03]. Lattice [Kaw79]. lattices [DDP07]. Laurence [Bis82]. Laver [Tho77]. Law [LG76]. Lawrie [Atk82b]. Layer [BA98, GPR^{+}98, AS08, HYH15, RSLAGCLB16, SDDD10, SBS13, ZZJ21, ACF13]. layered [BB99b, DMD^{+}06, Hun00, KKdC^{+}23, vdP14]. Layout [Blo93, CP96, LES95, AP85, CMT02, LZZZ18]. layouts [SB03]. Lazy [Com83, GT87, Har91, Kos90, GKS03, IS05, JL91, MJ99, SH82, BM97]. lcc [Han99a]. lcc.NET [Han04]. LCCD [Mei80, Mei81]. LCD [KCH07]. LDAP [LAG00, LCZ08]. LDMBL [MK18]. leading [WLS^{+}21]. leak [JSC^{+}10, RMM19, SSST15]. leakage [HSY^{+}20]. Leaks [Wad87, JM10, RW17, TSO19]. LeakSpot [RW17]. lean [PW11]. Learned [BMD^{+}98, ALG23, CC02, FL02, MVS^{+}18, VHM^{+}05]. learner [GDW^{+}20]. Learning [GLL^{+}24, TMS18, ABC^{+}21, BGA20, Che22, DFPT08, DFPT09, DFQ^{+}22, DNL^{+}20, FB22, GMP^{+}21, GB23, GJ22, GDRV20, GLKZ21, GWX^{+}23, GCZ^{+}24, HvdH02, HL20, HYZ^{+}18, HYC19, KSBL22, KGAR19, KS24, LQL^{+}22, LZK^{+}23, MKA^{+}22, MSC23, MPPC22, MGL19, MCGS08, MKD^{+}22, MG09, MR05, MAA^{+}22, MGJ^{+}24, NNK21, PK11, PPR^{+}21, PALNGD^{+}06, PBGM18, PNK^{+}23, QL13, RGS^{+}20a, SB21, SSV^{+}20, SGA20, SKD^{+}22, SH17, TRRK21, Val76b, XLZ^{+}20, ZZC^{+}17, ZKW21, ZHZ^{+}14, ZGB24, ZZJZ24]. learning-based [Che22, GLKZ21, KSBL22, MGJ^{+}24, SGA20]. learnt [BL15, DdB15]. Least [Inn77]. Leave [Thi80, Wil74b]. LECTOR [SM24]. Lecture [Cav83a]. lectures [DFPT08, Bar82b]. LED [LDG^{+}96]. Lee [Mul76]. Leendert [Ano88a]. legacy [BBS11, DFST08, DFPT08, LQ04, MMOD16, MBG19b, OMM15, SFK^{+}01, SATV22, SJA^{+}11, TL14]. legal [LTL^{+}03]. Lego [Hug93, Hug97]. Lehman [Inc86]. Leiden [Nee77a]. LEKTOR [Hum76]. Lempel [BK93, NT05]. Length [AW93, Cow87, New86, Fen02, Han94a, JL81, MT84a]. less [CB00a, KL21, LM15]. Lessons [BMD^{+}98, CC02, FL02, Men97, VHM^{+}05, BL15, DdB15, MVS^{+}18]. LETOS [Har99]. Letter [Ano80b, Bis80, Bro75, Bro78, Bud86, CW91, Ehr73, FC83, Gal79, Gos81, Han87c, HR90, Hor81, Jam80, Lea81, LDH92, Lin98a, Mit73, MW82, MIR78, NL75, NM77, Nie79, Pat83, RR82, Rec79, She81a, SF88, Ste79, Vil80, Wag78, Wex78, Wex81a, Ber99]. Letter-oriented [CW91]. Letters [Bar77a, Col72a, Dan82, FS73, GW84a, Har77, Hay80, Her77, JP79, Jos80, Mal80, MTRC83, PK82, Rai72, Rei82, Rya80, Sam71b, SW82, Wex75, Wul75]. Level [AG95, AE06b, AE06a, ACDP85, Bar76b, Cav83b, CDG^{+}98, FIL86, GW85, GH84, HF73, JBCB79, Kat83a, LOS83, PSV85, Par75, Ped86, Pyl79, Sat72, AML20, Al 13, ASP^{+}19, BS19, BA78, CCE^{+}21, Cia07, DD18, DTB12, EMRK20, Ell82b, FMT04, FM77, FN77, GXN10, GBE^{+}09, GIF01, GHBH05, GRR06, GVL10, GAB^{+}23, GZX^{+}21, HK84a, JKRS85, JGT95, KKR03, KLLK98, Kaw80, Lev80, LQ93, Mad79, MK04, MN18, Mor77, MW91, NM78, Nil90, PLR13, Pas87, PDBG10, SW86a, Spi09, Tag88, TKF09, TK09, Val77a, VGF21, XP23, YZYL07]. levels [KKPP20, ZJY^{+}15]. Levenberg [RCC17]. Leveraging [CGM^{+}03, GMP^{+}21, LQ04, ADZF21, MW13]. LexAGen [SN90]. Lexical [BF97, Gro89, GN16, Heu86, RS93a, Wai86, ZYF20]. lexicon [CD01]. Lexicons [ZD95]. Libra [SAL^{+}04]. Libraries [Cox85, Ker80, MS94, BDMP17, GS06b, LKK19, MBBS21, PLR18, PM18, Vo00].

Library

[ARS^{+}94, DV85, FBDH79, Gor87, Nar94, PR98, Pry85, RH77, Sch76b, Vo97, ADDM84, Ano76h, BT07, Bri84, Che04, CS17, Čuk16, DKS08, FGK^{+}00, GL05, GCF15, KS20, KL12, LKK^{+}18, LD99, MSZ^{+}24, NPHJ18, NZPWR22, PMP^{+}16, RPP07, VR06, Zho03, ASAQ05, JPL03, PPBP06]. LibVM [GCF15]. Life [Cho96, CK13, DFPT09]. lifecycle [ABBS22, TC03, WTF^{+}22, ZHO^{+}19]. lifecycle-aware [WTF^{+}22]. lifetime [SdLJMP21]. Lifetimes [Han90]. Lift [GR95]. Lifter [JL91]. lifting [GS06b]. Light [BS90c, RS91, CDR13, NAU^{+}21]. Light-weight [BS90c, RS91]. Lightweight [GN02, wKJM18, SCR94, TEGF08, YME05, GLT08, Har99, KCS^{+}20, LNhCW16, MSC23, NSM16, Pol01, RMMLSME14, TMJ^{+}21, WBN^{+}20, WCS^{+}17, ZGB24]. Like [Ham74, BW71, EBD^{+}74, HY20, Kaw79, MGGS18, Neh79, Pla97, HCC96, OW16, VV84]. Lilith [GW84b, Rei84]. Lime [BH94]. Limitations [Lav77, Var93, BLC19]. Limited [Bar72c, Mos73]. Limits [Gut87]. limp [Ree78]. LINDA [CD94, CLZ98]. Lindsey [Bar74e, Bra80]. Line [Ban71, BMA72, Bro71, Pan72, VWB91, Als22, BMR03, BBS11, Car79, DPAG11, FV03, GJ93, Han11, LJ99, MRBB19, Mau05, Rag86, SCT02, TDH97, WPL^{+}21, CCLN22]. Linear [GF84, Lic77, Ram96, ABBS22, Ber82, BJL06, HBC15, PM17, SS03, vdP14]. linearizability [AMSG23]. Lines [KP81, ADH^{+}00, CL19, SYG^{+}18, TAFC00, dSdMSNO^{+}11, vGPB10]. Linger [Han95]. Lingo [FMT04]. Linguistic [ALBN81, Gri80, KD13, KMS98]. Link [CB72, vdBT77, KH07, MDWD01, BDG^{+}00]. link-time [MDWD01]. Linkage [MT78, YR92]. Linked [Kil71, Nil88]. Linker [FH82b]. Linking [AEH76, HO91, IM93]. links [AC13, ACCD01, CZL24, SBcC07]. Linux [BGM99, BTS09, BV06, CGR00, JGS^{+}08, LSAF16, MM06, NJ11, NAGL10, RLPA18, RGK99, SJP^{+}09, TCM07, dOdO16]. LIS [HCD84]. LISP [Bai85c, Fid88, FN77, GH81, Kur78, Lic86, Rei82, Ume91, BW96, Iwa02, MK90, Val80]. Lisp-based [Iwa02]. LISP/PROLOG [Bai85c]. List [Bae73, Hum76, LH86, Mes96, Pal74, TT96, BL15, Coo05, Gru79, MUE23]. List-based [TT96, MUE23]. List-oriented [Hum76]. Lists [Jor78, McG89, Sti79, Har81, Sal81a]. Literate [Knu92a, RM91]. Literature [Ano09, ARA18, BBB^{+}11, CZL21, DPAG11, FAA^{+}22, MCLL21, NHB23, PRP^{+}23, SPR^{+}19]. Little [Bec91, BP98, Mar83, Hoh04]. Live [FK90, DSD^{+}19, IMBB20]. Lizuka [Pra96a, Pra96b]. LKMs [TXHL18]. LL [GJ88, PQ95, SMM^{+}84]. LLAMA [GAB^{+}23]. Lloyd [Lon88]. Llun [GIF01]. LMA [RCC17]. lmbench [Sta05]. Load [BS85, HC97b, SZ88, SA20, ZZWD93, BS19, CFLC14, CPCL10, CSTL19, DTJ89, GDW^{+}20, HL02a, IK15, Kar21, Li18, PACK07, PDPM^{+}16, SJA^{+}04, TDDE15, TRGA18]. Load-balancing [BS85, SJA^{+}04]. load-sharing [DTJ89]. load/store [PACK07]. loader [MT78]. loading [DGPT14]. Local [ABSS98, BP90, Er85, FIL86, Fis83, LP86, NIEN85, Poo88, Tag88, TP92, DDDF17, DS03, LQ96, SCL00, STA09, YWN^{+}00, SCL00, Her84]. local-search [DS03]. locale [Eng06]. Locality [Bae73, BGA20, CDF^{+}24]. locality-sensitive [BGA20]. localization [CC13, DW13, HGK^{+}19, LM15, LLW^{+}23, NAU^{+}21, NNLR17, NNR18]. Localizing [CT90]. locally [SC22]. Locating [ZGG07]. Location [GDBP23, Smi89, FR09, NSM^{+}22, SO21]. Location-aware [GDBP23, FR09]. Lock [BPM93, UN19]. Lock-and-key [BPM93]. lock-step [UN19]. Locking [App89a, Day00, PGK^{+}10]. log [KKPP20]. Logic [CZA83, KP90, LL91, Sch83b, TY80, War80, ASC^{+}01, BK23, CFL^{+}98, FCR^{+}09, RBL^{+}16, Sav06, SRRFGC^{+}10]. Logic-programming [Sch83b]. Logical [Har95, TTH97, AA19, Eve73, Nee77a]. Logicon [LC86]. logics [ZB23]. logs [AZS19, KSSA23]. London [Ano73a, Bar72a, Bar72b, Bar73e, Bar75e, Bar75f, Bar77d, Bar77c, Bar78c, Bar78b, Bar82b, Bis81a, Bry77, Bul72a, Bux78, Col77b, Edm82, For72, Han77a, Haz72, Hop74, HW77, Jac84, RB82, Ree73, Ree76, Rob72, Rob81, Rog74, Wel72, Wic72a]. Long [Han95, MS96, Wil79, DWL^{+}17, SOKE23, WBN^{+}20, Str81]. long-term [DWL^{+}17, WBN^{+}20]. Long/Short [Wil79]. Longest [BK93, Deo10]. Longest-match [BK93]. Look [Ten78, WMHL21, SS21]. lookahead [Abb78]. Looking [Rus95]. Lookup [Sew82]. Loop [GXN10, Hoa73, WJC^{+}14, WW91, GRFFGC^{+}21, PLR18, RBR21, UWW^{+}05, WSN^{+}23]. Loops [DH79, Dro85a, WW91, CA86]. Loose [FH74]. Loosely [AP95]. Loosely-coupled [AP95]. LORETO [BDSV99]. loss [CTLL07, CHCC07, TWS^{+}22]. Lossless [Was12, Sta07]. LOTOS [BDSV99, JEG99, LOBF88, VSC93]. Lout [Kin93]. Low [Bai85b, Del82, Kaw80, Mor82, PF97, Tag88, Wir90, Al 13, DD18, FBB^{+}14, GAB^{+}23, GAQO23, GCZ^{+}24, HIRB24, IBA^{+}21, LCGS17, Loe07, MVOD19, PKN^{+}12, SDIP23, TK09, XP23]. low-code [SDIP23]. Low-Cost [Bai85b, PF97, Wir90, LCGS17, MVOD19]. low-effort [Loe07]. low-latency [PKN^{+}12]. Low-level [Kaw80, Tag88, Al 13, DD18, GAB^{+}23, TK09, XP23]. lowering [KKdC^{+}23]. LR [AHS86, DP95, GL78, HHM92, HC87a, HW90, McK90, Mer93, SSM11, SK96, WRD99]. LR-WPAN [SSM11]. LSD-1 [Les72]. LSE [CLD^{+}17]. LSI [Hay80, Mat80]. LSI-11’ [Hay80, Mat80]. LSM [CGZ^{+}20]. LSM-trie [CGZ^{+}20]. LTAP [LAG00]. Ltd [Bar76b, Bar79b, Cou84a, Sto88, Wal86b, Wil76]. LTPL [KRTW81]. LTPL-E [KRTW81]. LTTng [WKJ15]. Lua [IdFF96]. Luegger [Wal81a]. LZ [Ris05]. LZ77 [Fra06, LNhCW16]. LZ77-compressed [Fra06]. LZgrep [NT05].

M [Ald72, Ano79a, Bar75a, Bar76a, Bar76d, Bar77e, Bar77d, Bar77b, Bis79a, Bra75, Bri82, Col77b, Cou85a, Cou85b, Eme84, Eve73, Fen94a, Gar86, Han78a, Han77a, How76, Hun72, Hut74, Inc86, Jon74, Lav77, Rob82a, Roh77a, Sto88, Val76a, Val79, Wal86b, Wil72, Art82, DS09, Joh78, MZC10]. m-JGRIM [MZC10]. M2 [DHGR92]. Maarssen [Val77a]. MAC [SSM11]. Macdonald [HW77, Wel72]. Mach [EKM^{+}99, EKM^{+}99]. Machine [Atk77, BA78, Bar74a, CD82, Die97, FBDH79, FH82a, FH82b, Gob71, GM73, Gri80, GM85c, GH84, HR96, Hum76, JDJ^{+}06, KvEP95, Lar75b, LBS78, LA90, LLW98, MP82, NPW72, Ray75, REC75, San88, SHR80, Sch76b, TT96, TY14, TTH97, AF02, AHRR22a, AHRR22b, AA20, AvRAF09, CARB10, CHCC07, Dun75, EF13, EGKP02, GMP^{+}21, GCARPC^{+}01, Ham81, Han99b, HL20, Ibs84, MKA^{+}22, Man18, MPPC22, MAA^{+}22, MGJ^{+}24, NZH20, NNK21, RGS^{+}20a, SB21, SGA20, TRRK21, WKJ15, YME05, YC16, YRJ18, ZLZ^{+}19, ZKW21, BZD17, DCA04, KM13, PNM^{+}20, Val77a]. Machine-Independent [FH82b, HR96, Ray75, Atk77, Hum76, MP82, AvRAF09, Han99b]. Machine-level [BA78]. Machine-Specific [FH82a]. Machines [Bow73, FH82a, HC93, HMS^{+}95, KM94, LF74, RS94, ABL08, BDQR23, BHvR05, BGSG20, DC15, IMBB20, LPP09, PMC05, Rob79, TRRK21, TGCF08, ZSRR22, VED06]. Macmillan [Bar78c, Bis79a, Bis81a, Cou84a, Edm82, Rob81, Wan82].

Macro [ADM96, Bro80, BO83, Com79, DM77, Hay83, KS87, Lar75a, Nie79, Rév85, Wel78a, Zel72, Ham79, Sas79, TC19, Jon71, Han78b, Lan75]. Macro-implemented [Zel72]. macro-optimizations [TC19]. Macro-Oriented [KS87]. Macroprocessor [BP84a]. Macros [Bro79]. MaD [ACV10]. MaD-WiSe [ACV10]. Made [Car98, MP13]. madness [Ano72b]. MaDViWorld [FMPR02]. Magic [Yuv75]. magnetic [HC16, VP05]. Magnus [Cor99a, Cor99b]. Maidenhead [Bar76b, Bar79b, Bul72b, Hut76, Rog73, Wil74a, Wil76]. Mail [Lib97b, BS99b, HL94, Pre85, SN07, Kor92]. Mainframe [Ben89, DSW82]. Maintain [IC85]. Maintainability [Ein88, FRBRF19, KB06]. Maintained [MRNL92]. Maintaining [AS88, ACCD01, CLLT98, Fra80, Fel79]. Maintenance [Aji95, Har95, RDLK90, WI85, CHY^{+}22, Car79, Inc85, MM82, PLR13, PPR02, RQL^{+}20, WP05, Wal81a]. Major [GM73, Ber82, SKI08]. Majuscules [Sal79c]. Make [Fel79, LS81, Wal84a, Fow90]. makes [MCCJ23]. makespan [FSRF^{+}22]. Making [AHH15, BDG93, Fai87, JI21, QTV23, SYXZ14, YLM^{+}05, AKIM24, KYF^{+}22, KY77, RCA^{+}19, SSS22]. malpractice [Spi76]. Malus [MS74b]. malware [DFW^{+}12, FB22, MV16, SWBS17, YAVHC21]. Man [AC80b, Bar76e, CD82, Pap79, SHR80]. Man-Machine [CD82, SHR80]. Man-Month [Bar76e]. manage [GMC^{+}21, TV09]. manageable [SC22]. managed [JM10]. Management [ALBN81, AD87, ACC83, AFI98, BMD^{+}98, Bre86, BSRS85, BK86b, CAC^{+}84, Coo86, CL95, GHM96, Hal86, Han77c, Han80b, HUS^{+}91, Hos98, Hut79a, Kat71, KP90, KH96, LCC97, LQ93, Mar85, NIEN85, PH84, REMC81, Sin81, SWA^{+}97, SWBT86, SMR89, TT74, Wal81b, Wat89, WG92b, YH97, AKIM24, AKM17, ASEB09, ACV10, ACAB22, AMR90, AA23, ARMMA18, BGS^{+}13, Bla04, BIA24, CMPQ23, CPCL10, CHS^{+}05, DFOT10, DAC^{+}21, FIÁSLSAR05, Flo74, FP15, FZW19, GMPL11, GRFFGC^{+}21, GB23, GMGB22, GB02, GDGB17, HSRB24, KCH07, KBB^{+}20, KBM02, KJVS21, Li22, LZ10, LGP^{+}11, LTW^{+}21b, LTL^{+}03, MVOD19, MM02, MMHB08, MGGS18, MVS^{+}18, NRUP21, NRS13, PK11, QC17, SDIP23, STB14, San17, TJB^{+}19, TW16, TLB^{+}18, XDVB24, YWN^{+}00, YYSG11, YB06, ZSL21, dAKdGJ11, dOED^{+}20, vdHW03, Ano88c, Flo79, Tho74]. management [Wil74a, Hut74, Hut76]. Manager [ORT81, RS90, SF98, Sil81, CC18, Rei99]. Managing [BMC^{+}24, CB00a, Cho98, Kno81, MH05, Mac96b, PSRCC02, PJS^{+}23, PW93, SY79, TC03, ADZF21, BB99b, CR18, FSR11, GVBG22]. Manchester [Bar72c]. mandatory [RdOTF14]. MANET [KHS^{+}20]. Manfred [Sim83]. Manipulate [TDH97]. Manipulating [BY90, Car97, CdA12, JG89, TS91, KRR19]. Manipulation [Bis84, CQC98, Car85b, IR80, Lee80, MN80, SW86a, Vau89, WLL98, Bar74f, CS15, Mad82]. ManPy [DPH16]. MANTIS [ASH73]. manual [Bar76a, Wid90, Bar72c]. Manufacturers [GM73]. Manufacturing [BH92, DPH16, DS09, DFRR15, GMC^{+}21]. Manuscripts [AS88]. many [BOPN12]. many-core [BOPN12]. Map [SCPL23, Com79, WY18b]. Maple [Car97]. MAPLIB [Sch72]. mapped [Sla86]. Mapping [Des74, Des92, Jak04, MRNL92, RB89, SHC74, BGM17, BOPN12, CCC^{+}16, HYY23, HBK20, HAM18, MSZ^{+}24, PP84, SLBC^{+}23, SYB04, dSdMSNO^{+}11]. Mappings [Hut78, DS99, NGLL14]. MapReduce [ANSK16, KKA^{+}17, TBSI18, ZXT^{+}17, ZLY18]. MARC [Sur13]. Marcus [Bar76d]. Mariani [Sau88]. maritime [KRB21]. Mark [Ano88b]. marker [LM15]. marker-less [LM15]. Market [GL97, TMGR^{+}22, PKN^{+}12, YB06]. market-based [YB06]. Market-inspired [TMGR^{+}22]. Marking [Kur81, TC07, TGPS08]. Markov [BF75, NZH20]. Markov-based [NZH20]. markup [YLM^{+}05]. Marlin [Cav83a]. Marmot [FKR^{+}00]. Marquardt [RCC17]. marriage [PK04]. Mars [Bra99]. Marshall [Bow88]. Marshalling [Bar97]. Martin [Bar81]. Marwick [RB82]. MARY [Rai81]. MARY/2 [Rai81]. Mary2 [Rai84]. Maryland [Wei85]. MASCOT [Bud85]. MASH [MP13]. mashing [OMM15]. mashup [PVAHRG^{+}15]. Mask [DW73]. masking [GSAE14]. Mass [Bar76e, Ear77, Fin77, Llo82, PMY97]. Massive [RB89, GP14, ZWML14]. Massively [ABBE98, CHC^{+}17, FMPR02]. Master [Bul87, BK87, RH77]. Master-Detail [Bul87]. Master/Slave [BK87]. Mastering [SGBR13]. Masthead [Ano71g, Ano71h, Ano71i, Ano71j, Ano72d, Ano72e, Ano72f, Ano72g, Ano73c, Ano73d, Ano73e, Ano73f, Ano74a, Ano74b, Ano74c, Ano74d, Ano75b, Ano75c, Ano75d, Ano75e, Ano76d, Ano76e, Ano76f, Ano76g, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l, Ano79m, Ano79n, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano80m, Ano80n, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82l, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i]. Masthead [Ano83j, Ano83k, Ano83l, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano88e, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano88m, Ano88n, Ano88o, Ano88p, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano90l, Ano90m, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g]. Masthead [Ano91h, Ano91i, Ano91j, Ano91k, Ano91l, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j, Ano92k, Ano92l, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano93m, Ano94l, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94j, Ano94k, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano95k, Ano95l, Ano95m, Ano96k, Ano96l, Ano96m, Ano96n, Ano96o, Ano96p, Ano96q, Ano96r, Ano96s, Ano96t, Ano96u, Ano96v]. Match [DS88, BK93]. Matches [ZD95, Mha05]. Matching [AC75, DB86, JTU96, KST94, Lec95, Lec98, Liu86, Maa06, OM88, PB87, Ric79, Som82, TP97, VSM87, Wri94, de 82, AG06, CFKT17, DGM19, Fen01a, Fen01b, FBMA05, Ier09, Nav01, NT05, NWE99, NK07, Sas79, THG17, ZWAL22, LCZ08]. MATE [SCT02]. Material [Sch72]. materialization [RL14]. materialized [SS23]. materializing [RÁdMRGAM19]. materials [MSK^{+}24]. Mathematical [Cox85, Lev83, MM02, SRS18].

Mathematics [Day83, Glu74]. MATLAB [BC17, JB07]. matrices [Dod82]. Matrix [HP88, Mat94b, RB91, Kha86, KKdC^{+}23, LD14]. maturity [CGP^{+}06]. Maximal [McG82]. Maximizing [MAR^{+}16, DSD^{+}05]. maximum [Rog23]. Maynard [Bar72b]. Mayoh [Wal83c]. MC [MST13]. MC68000 [Poh81]. Mccabe [Har84a]. McCracken [McD71, Ree73]. McGee [Roh77a]. McGettrick [Fox79, Gru83]. McGraw [Bar77b, Bar79a, Bra75, Ken77, Rog71, Val80]. McGraw-Hill [Bar77b, Bar79a, Bra75, Ken77, Rog71]. McKeag [Han77a]. McKeever [Hut74]. MCL [ZCN06]. mCRL2 [GKS^{+}11]. MDA [LER17]. MEADOW [CKL^{+}02]. mean [ZB18]. meaningful [AE14]. Means [BTZ94, MMOD16, PLFIVRAR23]. Measure [LB94, CKB00, CKB01, CKB03, Geh85, Har84a, ML08, XZ01, XZ03]. Measured [Zor93, Cer18]. Measurement [BMA72, Cro91, FL75a, HG89, Kue95, Pra96a, Pra96b, RK89, YSM95, Al 13, CGR00, HL02b, MBBS21, SMT^{+}18, SSK^{+}17, TSMGD^{+}11, WMJ04]. Measurement-based [RK89]. Measurements [DD90, WS94b, Pal78b, dSRdSS^{+}21]. measures [RK15a]. Measuring [DP85, Lop89, PW11, WAML12, WH98, ASA^{+}21, AHH15]. Mechanics [Liv75]. Mechanism [LF74, MR80, Sil81, WBV96, CSTL19, CY01a, CY01b, DHWZ14, KS10, LCC14, NT84, Tsi82, WCT19, WR22]. Mechanisms [ALBN81, AO88, BAF03, ET07, GST92, Kow81, LMN91, PT14, VL73, WH84, And82b, JZ10, MF08, SKI08, Wij05, Dea86]. Media [MNH04, DO07, JKW74, WK06a, ZSFY05]. Median [CMR92]. mediator [NR04]. Medical [ABA20, BJ21, Ald72, BCC^{+}22, GDRV20, GLKZ21, MMOD16]. medicine [SAYM24]. Medium [Lea82]. Meek [Lav78]. meet [CW01]. meets [ZXX24]. Meglos [GK86]. Melville [Flo74]. member [Pil75]. memoization [LV20]. Memory [AS97a, AFI98, Bae73, BH82, BAFR96, BMD^{+}98, BF75, BS90c, CQC98, Cha88, CSTL19, Coh73, DDZ94, FHJ94, GZ93, Han90, HC97a, KLY20, Lec98, LKBT92, McC83, PCL^{+}99, RK91, Rey90, Sch83a, Smi80, SJKL94, SSST15, TA91, Vo96, WZF94, ACM^{+}15, BST10, CYJ^{+}22, CLC99, CMTCC^{+}17, FBB^{+}14, GT92, Gra96, GAB^{+}23, GAQO23, HC99, HC16, HMS^{+}95, JSC^{+}10, JM10, KBB^{+}20, KSBW18, LCC97, LX04, MM02, MSK01, Mos73, ML20, PDPQ22, PNM^{+}20, Poh81, RLPA18, RW17, SB13, SOKE23, SB03, WJC^{+}14, Wat04, WS99, YYSG11, YHK^{+}22, ZWKX17, ZG06, ZCCB22, ZWF^{+}22, IS05]. Memory-aware [CSTL19]. memory-based [SOKE23]. memory-bound [GAQO23]. Memoryless [GS76]. MemSafe [SB13]. Menu [Hef82]. merge [Har81]. merges [PdSCJM22]. Merging [Fra80, Jon72, CPW73, KWB^{+}05, SZ20]. Message [CCvKH95, Fje79, Geh90, Gen81, HI85, JVR97, LZWX22, LB81, MT94, NJ11, Pat94, Smi85, Sta82, TA91, ZZT^{+}21, Bre82, GB13, PŻZ13, SNL15, SZ00, TEBK99]. Message-based [Smi85]. Message-driven [MT94]. Message-flow [CCvKH95]. message-oriented [PŻZ13]. Message-passing [TA91, Bre82, GB13, SZ00]. Message-sensing [ZZT^{+}21]. Message-state [Pat94]. Messages [HA72, HR77, KL86, Bro82, Ker17]. Messaging [WC87, KQZ^{+}11]. Messer [Bow88]. Meta [Fid82, FL94, BGSG20, HP11, Mar83, Wit82]. Meta-Assembly [Fid82]. meta-heuristic [BGSG20]. Meta-model [FL94]. meta-modeling [HP11]. metadata [CR18, JRGC20, MC23]. metaheuristic [IMG^{+}21, TGAS22]. metaheuristic-based [TGAS22]. Metaheuristics [HLRVB18, ARMMA18, DDDF17]. metalog [Sch83b]. metamodel [PLR13]. metamodel-level [PLR13]. metamorphic [LLZ20, aSJW^{+}24]. metaprogramming [LS15, Rin07]. MetaSockets [SMKZ06]. Metastructures [SG79]. Metcalf [Wil87]. Meteorological [Cra76, Ham84]. Method [AV05, CK97, Col87, Doo92, Dri93, EE90, HI85, Hos98, Hug79, Hun81, KT84, LH86, MPN^{+}95, MM88, MIA94, Par85a, RS87, Sew82, SMM^{+}84, SY79, vHLB^{+}88, AF99, ADZF21, AGRS11, BBG04, Cox76, CV08, DB21a, GW04, GM22, HHM92, HLH15, IB13, IH01, JAJB04, KSH^{+}15, LHB18, LLZ20, LFH^{+}22, LC07, LZLL22, LLW^{+}23, MSC23, MBG19b, MCCJ23, Mor77, OAF^{+}03, PPR02, RRK^{+}18, SNL15, SJ79, ST01, Vo00, Wu99, XXJS18, XDZ^{+}17, XCZ^{+}21, YOM^{+}07, YWT^{+}12, ZLX^{+}24, ZZL^{+}21, ZWX^{+}21, ZQWL24, ZSRR22, Jac85]. methodical [Atk79b]. methodological [DFRR15]. Methodologies [DRL82, PVAHRG^{+}15]. Methodology [BP84b, Cel82, HL91, Mac79, MXYQ86, OLS89, PU84, She92, SBS13, ARK21, CA18, CSMML12, CCM05, IHS^{+}14, LC12, PPR02, WYAZ15, YYLJ24, ZZKA17, ZC03, Sim83, Val79]. Methods [AI80, DW91, Ham77, QK78, Rai73, Ree75, ST14, Thi93, BAJMT21, BR01a, CLP^{+}09, Dav78, DFST08, Fra99, GEI^{+}11, GMP^{+}21, GF11, KFMF18, KVG19, LW14, MKE18, MdCGdC^{+}17, MOTG18, MFYiA01, PGK^{+}10, Ree73, ZNWS18]. metric [Mar22]. Metrics [BP90, HK84b, Poo88, RCC91, CS17, CIM22, CIM24, DB21a, DD18, FRBRF19, GKWS11, KSKG12, MCLL21, SBF19, WS99, WHS^{+}00]. metrics-based [CS17]. Metropolitan [DDB^{+}18b]. Meulen [Bar74e, Bra80]. Mexico [KDP83]. MFHS [KTG20]. MHP [BFPAGS^{+}08, PALNGD^{+}06, VRC^{+}06]. MHP-OSGi [VRC^{+}06]. Miami [Rob72].

Mianyang [WPL^{+}21]. Michael [Ano88c, Val76a, Wil87, Hug79]. Michel [CO88]. Micro [CW92, Cor88a, Mor82, TC19, WOKT81, FO10, Hen79, Sta05, AÇCT22]. Micro- [TC19]. micro-aggregation [FO10]. Micro-analysis [CW92]. micro-benchmark [Sta05]. Micro-Computers [WOKT81]. Micro-IDE [AÇCT22]. Micro-Kernel [Cor88a]. Microbenchmarks [MMK04]. microcloud [AFNG20]. Microcode [CLL91, Ise90]. Microcoded [CMH85]. Microcomputer [CW82c, EE90, GW85, GLW82, HH79, MV86, OW83, RR85, SW86a, SB82, Atk79a]. Microcomputer-based [SW86a]. Microcomputers [Ben89, Del82, DMW88, JI80, Oni85, PVS79, HK84a]. Microcontrollers [KR85]. microkernel [FC98]. MicroMAIS [PCC^{+}12]. Microprocessor [CM83, Gon87]. Microprocessors [SF85, Bri82]. Microprogram [MP82]. Microprogrammed [CJ88, Hal82]. Microprogramming [FM77]. Microservice [HBK20, LQS^{+}24, STT24, BMC^{+}24, BNS18, CZL21, JC19, KSSJ^{+}23, KSDG24, MA20a, SMNB21, XDVB24]. microservice-based [BNS18, KSSJ^{+}23, SMNB21, XDVB24]. Microservice-driven [LQS^{+}24]. Microservices [BHJ^{+}18, AÇCT22, CDF^{+}24, CS22, DLX^{+}22, DZ21, FSC^{+}21, FBH^{+}23, HBB22, JJI22, LBGA^{+}21, RR22, TJB^{+}19, XYW^{+}24, ZHZL22]. microservices-based [TJB^{+}19]. Microsoft [Tur22]. MidCloud [MAJ15]. Middleware [BFHR99, BR01b, CKL^{+}02, CPCL10, dSCdRS^{+}19, GA12, GFS^{+}05, KBH^{+}03, KHC^{+}19, MZC10, MGGS18, MAJ15, NRS13, OEA05, PKN^{+}12, PVBB06, PŻZ13, SLRS06, SMR^{+}12, ZLG08, ZCN06, Gue03]. middleware-transparent [GFS^{+}05]. midsummer [Ano72b]. Migrating [DFPT08, FSC^{+}21, MMOD16, SFK^{+}01, WKKK24, SSP11]. Migration [CLL91, DO91, FJ03, HKV95, MR96, SH98, BHJ^{+}18, CS02, CS22, DFST08, DSD^{+}19, FV03, JPM17, JGSG^{+}21, KLY20, KHS^{+}24, wKJM18, MKC11, MBG19b, NB19, RCA^{+}19, RGC^{+}21, SATV22, SM20, SRS18, XZW^{+}22, ZLG08, ZLZ^{+}19]. Miklos [Tho74]. million [TAFC00, WWCW19]. million-user [WWCW19]. millions [NL24]. MILLIPEDE [FGIS97]. Mills [Han95]. MILS [DPK12]. MIMD [GM85c]. minded [Yuv79b]. Mine [TBPK20, GC20]. Minefield [Han94a, VV06]. MING [HM18]. Mini [EE90, Jal82, RAB^{+}79, Ste79, Mer74]. Mini-computer [EE90, Ste79]. Mini-computers [RAB^{+}79]. Minicomputer [CRT80, Fra75, Hal82, PJ75, Pur76, Val77b, Val78]. Minicomputers [BD76, Hor78, PP80]. Minimal [AW93, Dun91, YOM^{+}07]. Minimal-prefix [Dun91]. Minimization [WHLM98, LSZ16, PM17]. Minimize [Bla92, QSA90]. minimizers [GR17]. Minimizing [WP05, ST01]. MINIMOP [Ree71]. minimum [KG18]. Mining [DDF16, Hor21, JLZ09, SMNB21, HPZ^{+}20, KVG19, LYX^{+}17, LLLY19, MRZ15, NSM^{+}22, RT10]. Miniscules [Sal79c]. Minivital [MG76]. MIRA [MTT83]. MIRA-3D [MTT83]. Mirage [FHJ94]. Mirror [MVT^{+}09]. Mirror-based [MVT^{+}09]. Miscellaneous [Ano81n]. misconfigurations [RGS^{+}20a]. mishap [JH03]. Misleading [Vau79]. missing [ZZL24]. mission [JH03, SDF^{+}21]. misspelt [Par78, Sco77a]. MISTRESS [AS73]. Misuse [FS81, LP78]. mitigation [IBA^{+}21, OY10]. Mixed [EG84, HMS88, MS74a, BB99b, LHGM15]. mixed-environment [LHGM15]. Mixed-language [HMS88]. mixed-strategy [BB99b]. MK1 [Wyv77].

Mkscan [HL87]. ML [BM97]. MLFQ [TCM07]. MM [SHR80]. MM/1 [SHR80]. MMLT [GMNR20]. MMRUC3 [RRK^{+}18]. MOAManager [MSB20]. MobiGATE [ZCN06]. Mobile [CPW74, AvRAF09, AWNS18, BHR^{+}02, BBMG08, BDP02, BCRF22, CKRC20, CPD13, CCP06, CSM^{+}16, CPMAH^{+}20, DM15, FMC18, FFF^{+}13, GCK^{+}02, GWZ^{+}20, HSC21, HLLZ21, HMRZ20, HSY^{+}20, HLH15, HM18, HC20, ISUG06, KY05, KT01a, LC07, LS16, MH05, MAR^{+}16, MKC11, MAG^{+}21, MZC10, MSZ^{+}24, PL08, Pei02, PCC^{+}12, RMZ17, RMMLSME14, RMdL12, SSGA20, TKT^{+}07, WWCW19, XZW^{+}22, YZL^{+}22, ZYW^{+}20, ZLZ^{+}19, ZZT^{+}21, ZCN06, ZSL21, CCP06, FMC18, LM02, SBcC07]. mobile-agent [GCK^{+}02]. Mobile-C [CCP06]. MobileRMI [AV05]. mobility [AV05, BHK^{+}04, JSPP21, LGRL08, XZW^{+}22, ZLZ^{+}19]. mobility-aware [JSPP21, XZW^{+}22]. mobility-induced [ZLZ^{+}19]. Mobolic [AWNS18]. Mock [Tho74]. Mockup [ZC03]. Mockup-driven [ZC03]. mockups [DDGP18]. mode [GG08, Le 88]. mode-directed [GG08]. Model [ATO10, CS91a, CGH^{+}15, CLSE05, Cho96, CH90, Des92, Dew91, ENP^{+}24, Fid88, FBLS12, FF80, Gom78, Gom82, Hut79a, LSK^{+}18, LGZ^{+}08, MMCF03, Mat94b, RGC^{+}21, SAYM24, SW90, SCGP92, She81b, SROV06, TL14, UFR18, WPT95, WW95, Wol82, WS74, AA19, AS08, ALD^{+}23, AGRS11, BELS14, BCL^{+}06, CCQ16, CFLC14, CA18, CLD^{+}17, CCE^{+}21, CEF02, CM08, CRGIP15, CA14, Čuk16, DB21a, DS12, DKM11, FSRF^{+}22, FL94, GMPL11, GA12, GZX^{+}21, GQ15, GDW^{+}20, HAM18, Hsu12, HY20, HNJ^{+}23, JJK^{+}12, JTG^{+}11, KSSA23, KH18, Kim15, KKS10, KEL^{+}21, KA87, LB02, LW04, LTW^{+}21b, MK01, MDH^{+}13, MPPC22, MCGS08, MTM22, MGG^{+}09, MVS^{+}18, MA20b, MAA^{+}22, Mus17, NNK21, NNL^{+}14, NZL19, NZTL^{+}23, PJJM21, PM17, PP84, RBR21, RN00, RZ17, RL23, SFC^{+}21, SZ20, SRS18, TWS^{+}22, TMJ^{+}21, UT19, VRC^{+}06]. model [WP00, XLZ^{+}20, ZZL24, ZHZ17, ZZJZ24, dAKdGJ11, dAHCdAC18, vDV04, FCBF^{+}21]. Model-based [ATO10, LSK^{+}18, RGC^{+}21, SAYM24, SCGP92, BELS14, CLD^{+}17, GA12, MDH^{+}13, NNL^{+}14, NZTL^{+}23, RZ17, UT19, WP00, dAKdGJ11]. Model-centric [SROV06]. model-checking [CCQ16]. Model-Driven [UFR18, ENP^{+}24, FBLS12, MMCF03, TL14, AA19, AGRS11, CM08, LTW^{+}21b, MPPC22, MGG^{+}09, MVS^{+}18, Mus17, NZL19, RBR21, FCBF^{+}21]. model-to-model [CA14]. Modelica [CL19]. Modeling [AZ97a, CGIP15, EYG22, IAPC17, LD95, Sef97, SHB19, YSM95, ZHZ17, dODP21, CRB^{+}11, CNRB13, CA08a, DHG^{+}19, FCYL18, FG11, GB13, GDGB17, HP11, JAA^{+}20, KKR03, LHC15, NUB24, NZPWR22, PDCB17, RMMH^{+}23, SAEMM21, VS18, Wai07, WAH^{+}12, WYAZ15, YSBL22, dAPMV10]. Modelling [AKM17, BBC91, BZM^{+}17, CD82, DV84, Gan82, GR91, Gri80, KR83, LL91, NPW72, NSM86, SM79]. Models [AR93, BF75, HHK90, MFdlP12, SRCP19, TV96, Wat89, ABBS22, AFFR08, CSD23, DPH16, ENP^{+}24, FB22, HTJNL19, DCEG23, POM03, San17, SE11, TSMGD^{+}11, Wai02, ZZH^{+}24, dMFÁE17]. Modern [HZ94, FG14, KW17, MSB18, ZCO13]. Modes [Har92]. modest [SL04]. Modification [CG93, CRT80]. Modified [SNK21, Wen80]. MODULA [Bud85, BE81, BK87, Cor88b, DP85, Fos86, Gut87, HW80, Hop80b, Pro92, RH78, Rei84, Tag88, Ter86, Wir77b, Wir77c, Wir77d, Wir88a, Woo86, Mee87, Ano87a, Bow88]. MODULA-2 [Bud85, Cor88b, Fos86, Gut87, Hop80b, Pro92, Tag88, Ter86, Woo86, Mee87, Ano87a, Bow88]. Modula/R [Rei84]. Modular [CFP83, FWS74, GKM83, HJ14, HC87b, Hus86, JL91, Kos90, Mal17, OW89, SR88, SM81, WB79, Wir77c, BAF03, DCA04, KY05, KTG20, MGGS18, Mos06, SMGMOFM07a, SMGMOFM07b, Tur22, Bar72b]. Modularity [Bee82, MPS93, Tal71, Add80, BTS09, Mos73]. Modularization [HG81, CCF^{+}09]. modularized [Bra99]. Modularizing [PPSO17, Hoh04]. Module [GL85, PA91, CW82a, KNT^{+}01, KV17, Str81]. Modules [ABBE98, Han79b, Ian90, LT91, Wis93, ADDM84, BTS09, KW09, Mal17, ZZ11]. MOLE [BHR^{+}02]. Molecular [Str95, LFH^{+}22, PD00]. MOLP [ZB18]. Mondrian [SRGCPB^{+}09]. Monitor [JKRS85, MMS86, OM96, Rei72, SC90, Tho78, TTH97, VSB86, Wai73a, Wit83, WS74, CY01b, Gai82b, LX04, WWB03]. Monitoring [CLW90, Cum71, DR92, Fin97, FM78, GL97, JI21, JG94, ZLWG11, BJ21, Buy00, CYW^{+}15, DTB12, GMGB22, IHS^{+}14, KRK21, KCH07, LC07, LCC14, LZD20, MA00, MKW^{+}22, PM12, RBL^{+}14a, SGCM11, TLB^{+}18, TKT^{+}07, WR22, ZLY18]. Monitors [Han76d, LM76, LS77, PU84, Str82, HL79, Han78c, Ter86, YME05]. monolith [CS22, RR22]. monolithic [FSC^{+}21, WKKK24]. Month [Bar76e]. Moo [Her77, Gro72a]. Moore [Atk82b, Rai92, Smi94, NT05]. Mooshak [LS03]. Mortem [NY78]. MOS [BL85]. Mosaic [MWB95]. most [CK15, ESRI14]. mostly [NS01a]. mostly-copying [NS01a]. moth [GARSR18]. moth-flame [GARSR18]. motion [KKN04]. Mount [Smi89]. MOUSE4 [Com78]. mouth [WHH21]. Move [Gor94, RRK^{+}18]. movies [KSH11]. Moving [ASC^{+}01, XLZ^{+}20]. MP [MPP87]. MpAssign [BOPN12]. MPEG [WK06a]. MPEG-7 [WK06a]. MPI [PGK^{+}10]. MPL1700 [FM77]. MPLOT3 [SP79]. MPMD [CCE99]. MRI [JKB04]. MROS [Poh81]. MROS-68K [Poh81].

MRPC [CCE99]. MS [LHFL07].

MS-Windows [LHFL07]. MTA [HJ08]. mTags [RdOTF14]. Multi [AO88, BS93, Cho98, Day83, Dew93, Fis86a, Gay80, Gut76, HRW73, JDJ^{+}06, KKR03, KS98, KLLK98, KRO93, LOS83, LT90, NEP^{+}17, Poo71b, Pyl72, Ree71, SMFBB93, Sch76b, Sno91, SS23, SY86, TB72, WLZ^{+}22, WCE^{+}72, BDQR23, BMC^{+}24, BPR01, BB99b, DFQ^{+}22, DO99, FCA12, GCRD04, GHM^{+}06, GM22, HL02a, JPM17, JPG^{+}17, Kru82, LLJ12, LS03, MKW^{+}22, QH21, RBS14, RGK99, SIK^{+}16, STT24, SOKE23, TKF09, XLZ^{+}20, YLP^{+}11, ZZKA17, ZZJZ24]. Multi-Access [Day83, Poo71b, Ree71, TB72, Gay80]. multi-agent [BPR01, DO99, GHM^{+}06, HL02a]. multi-case [STT24]. multi-cloud [JPM17, JPG^{+}17]. Multi-combinators [LT90]. Multi-Computer [Pyl72]. Multi-criteria [NEP^{+}17, ZZKA17]. Multi-dimensional [Gut76]. multi-faceted [BDQR23]. multi-instance [XLZ^{+}20]. multi-layered [BB99b]. Multi-Level [LOS83, KKR03, KLLK98, TKF09]. Multi-Machine [Sch76b]. Multi-Net [WLZ^{+}22]. Multi-objective [SS23, FCA12, GM22, ZZJZ24]. multi-output [YLP^{+}11]. Multi-party [Cho98]. Multi-processor [Fis86a, LLJ12]. Multi-protocol [Sno91]. Multi-Purpose [WCE^{+}72]. multi-resource [SOKE23]. multi-site [LS03]. multi-source [SIK^{+}16]. multi-task [DFQ^{+}22]. Multi-Tasking [JDJ^{+}06, AO88]. multi-tenant [BMC^{+}24]. Multi-Terminal [HRW73]. multi-threaded [GCRD04, RGK99]. multi-touch [RBS14]. Multi-User [SY86, BS93, Dew93, KRO93, Kru82, QH21]. multi-vendor [MKW^{+}22]. Multi-way [SMFBB93]. Multi-window [KS98]. multiagent [BGS^{+}13, DFRR15, KCYY12, STH^{+}18, SAEGF11]. multiagent-based [DFRR15]. MultiArray [GL05]. Multicast [Hug88, Jia97, KG95a, LRMM93, Bir99, MA00, SR02]. multicloud [BDA20, GMS20, GMNR20, MA20b]. multicomponent [BRS18]. Multicomputer [BS85, BL85]. Multicomputers [MT94, MV95]. multicore [AKNJ21, BP02, GXN10, IMKN12, Knu11, LJL^{+}10]. multicores [MNW14, VGF21]. multicriteria [JSRM18, MA20b, RCA^{+}19]. Multics [Col81]. Multidatabases [FZ98, XCL^{+}18]. multidepot [GZW^{+}22]. multidimensional [PK04]. multidimensionality [Ron99]. multidomain [ZWQC22]. multies [Gro72a]. multifaceted [KSK15]. multifactor [Ell72]. multilayer [MGT20, OAZ19]. Multilevel [MR92, GMNR20]. MultiLex [BF97]. multilingual [KNT^{+}01, NHTT08, Wu00, Wu01, YLR23]. Multilinked [BY90]. Multimedia [HL94, HCC96, MWB95, TL98, WBV96, WP96, WKD96, WRR97, BFHR99, CGM^{+}03, CB00b, DFPT08, QC17, RSRCGC15, WCS^{+}17, ZC02]. multiobjective [ZWQC22]. multiphase [GvRN^{+}11]. multiphysics [DWL^{+}15]. multiplatform [NCFCFV12]. multiplayer [PV22]. Multiple [APS95, AM00, CAFH94, Han94a, LN71, Lib93, Mey78, MY87, OEA05, VS80, Wil73, AS08, CCQ16, CKL^{+}02, Fen01a, Har84a, IMKN12, JDPB08, Li18, LWW^{+}23, Maa06, Mal17, MP19, Mha05, MP00, PACK07, SFC^{+}21, UDS^{+}07, WW09, Was12, WCS^{+}17, ZGG07, ZWML14]. Multiple-Access [LN71, Wil73]. multiple-data [IMKN12]. multiple-exit [Har84a]. Multiple-length [Han94a]. Multiple-type [AM00]. Multiplication [Ber86, RB91, KKdC^{+}23]. multipliers [SV22]. Multiprocessing [Bar78a, HC87b, Rey90, Art82, DBO^{+}18, LvLS84, RGK99]. Multiprocessor [AP84, BS90c, GST92, GT92, Hal86, Han89b, LLCG^{+}89, Lun89, SNM80, TRO17, TAAT84, CM98a, CM98b, Han81b, LX04, QM13, RR05]. Multiprocessors [REMC81]. Multiprogrammed [Sch78]. Multiprogramming [Han73, Sch74, SWA^{+}75, Smi80, SB82, WB79, Wir77c, Bea78]. MULTISAFE [Har84b]. multiscale [BCLF^{+}07, SNK21]. multiseed [LJS20]. Multisensor [JI21]. Multistage [CRR94]. multiswarm [CS18]. multitarget [GJ22]. Multitasking [Cav83b]. multitenant [KHC^{+}19, LWZ^{+}19]. Multithreaded [SRS98, Yoo96, Fer13, GRR06, JWTG11, YZYL07]. multitier [KGAR18]. multitime [DSD^{+}19]. multitime-steps-ahead [DSD^{+}19]. Multitouch [KHHG15]. Multiuser [LWJ^{+}21, PALNGD^{+}06]. MultiView [NS01b]. MultiView-based [NS01b]. MUMPS [Bro81a, WOKT81]. Mungi [HEV^{+}98]. Munich [Woo74]. Murray [Tho77]. Music [BH94, Fil98, Fox87, Gro73, Lan90, BB03, Gou86, FFF^{+}13, Fox87]. musical [TC07]. MUSS [BCP79, FT79a, FT79b, TF79a, TF79b]. MUSYS [Gro73]. muSystem [BMZ92]. mutant [JP22]. mutating [Lam20]. Mutation [KO91, OPTZ96, Spa90, DdB15, GMGDMB19, aSJW^{+}24]. Mutation-based [KO91]. mutations [NNW13]. Mutual [YAVHC21, GMNR20]. MVS [Mar86]. MVS/XA [Mar86]. MVT [BL78, BL79]. My [Bro82]. Myers [And78]. MyNeutronDAQ [WPL^{+}21]. MyProxy [BHW05]. Myriad [LHS^{+}95]. Myth [Sch74, Sch83a]. Mythical [Bar76e]. Mythology [Ros75].

N [Ano23m, Bar74c, Bar76d, Cor82, Edw77, Gar86, Ken77, Lav77, Ros74, Wal82, DMW88, MDB19]. N.J [Bar73c, Bar74d, Bar75d, Bar75b, Bar76c, Edw77, Ros74]. NaaVRE [ZKB^{+}22b]. Naftaly [Val76a].

NAG [DV85, FBDH79, RH77]. naive [ZYW^{+}20, ScG09]. Nake [Lan74a]. naked [MVTH14]. Name [BPY90, KW17, CA08b]. namespaces [SDG^{+}20]. NAND [CSM^{+}16]. NAND-flash-based [CSM^{+}16]. Naplus [ZWKX17]. NAS [AGR^{+}23]. NASA [Coo08, JH03]. Nassi [HW88]. Nassi-Shneiderman [HW88]. National [Bar72c, Bar83a, Wu00, SWN94]. Native [KS95, PZ00, AGC10, ŠS08, VBS22]. natural [BFNP08, GN02, Har81]. nature [ASA22]. nature-inspired [ASA22]. navigating [SSS^{+}02]. navigation [Ano23m, SGA21]. navigations [KH07]. Navigator [MB96]. NCC [Rop88b, Bar83a, Bar83a]. NCL [SMM13]. Near [AW93, BT89, GW96, MY87]. Near-optimal [GW96]. Near-perfect [BT89]. nearest [MGT20]. Nearly [FP82, OG16]. Necessary [Han81e, Bar74g, CVGCGCFF23, Yuv77b]. Necessity [Oli83]. Need [BS74, HJS^{+}20, Str77]. NEEDS [SWN94, CW01, CJ73, Ozk18]. negotiation [EL05, MS18]. negotiation-based [MS18]. Negra [GS08]. neighbor [MGT20]. Neon [GYCL16]. Nested [Jen89, TE90, HY20, KS20]. Nesting [Gre80]. Net [HL91, WLZ^{+}22, HAM18, dMFÁE17, Wir90]. Netkit [PR16]. Netlink [NAGL10]. Nets [Inc84, Wen80]. Network [BNOW95, Cho98, DLP85, Daw77, Del82, DMW88, EP79, FIL86, Fje79, FHJ94, GPR^{+}98, Gom82, HS77, HH82, HMPT89, Joh84, LOS83, LP86, LD87, MRNL92, NIEN85, RS93a, RS21, SM79, SC90, Tag88, VSB86, Wir90, ZWX^{+}21, ASA^{+}21, ABA20, BGS18, BKL^{+}02, BSDF20, CGM^{+}03, CDR13, FB22, HB18, JFZ^{+}21, JGC^{+}21, KPU04, KCCV05, LZWX22, MA20a, MS22, MDB19, OAZ19, PR16, Pre85, RQL^{+}20, SIC^{+}20, SDG^{+}20, SBG^{+}05, SHB19, SNK21, SOKE23, WHYC23, WMJ04, WR22, WLS^{+}21, YWN^{+}00, YFC06, ZDY^{+}17, ZLZ^{+}19, ZWQC22, BLNU15]. network-aware [HB18]. network-based [MS22, YFC06]. networked [BV06, EGL18, SSS^{+}02]. networking [HYT13, WN06]. Networks [BL90a, Col87, Her84, HP83a, JI80, WC87, dCV88, ABC^{+}21, ACV10, AFNG20, BGS18, Ber20, CVGCGCFF23, CBB20, CLS^{+}07, EC13, GCARPC^{+}01, HPK^{+}12, HLH15, IMG^{+}21, KAS^{+}14, KA22, KS24, LLJ12, LQL^{+}22, MTPC14, NH03, SIC^{+}20, SDB^{+}22, SA20, WLZ^{+}22, WAML12, XZX^{+}21, XCZ^{+}21, YAFA19, YMY17, YZL^{+}22, ZYW^{+}20, ZZJ21, ZZJZ24, dAKdGJ11, KG95a, Rog73, Vel88]. networks-on-chip-based [LLJ12]. NeuCheck [LWZ^{+}21]. Neural [BL90a, ASA^{+}21, CBB20, JFZ^{+}21, KS24, LQL^{+}22, MS22, OAZ19, WLZ^{+}22, WHYC23, WLS^{+}21, YAFA19, YFC06, ZZJ21]. neuroimages [VP05]. neutron [WPL^{+}21]. Newcastle [BMR82, SW86b]. NEWLONG [Car85b]. newly [Als22]. Newman [Bra75]. News [Lib97b, KHMB17]. Newsqueak [Pik90]. next [MRG^{+}19, RGS^{+}20b]. next-generation [MRG^{+}19, RGS^{+}20b]. nexus [JPM23]. NFS [BH01]. NFV [KDA20, SHB19]. NFVLearn [SOKE23]. Nial [Jen89]. NICE [WS94b]. Nicholas [Bry77]. Nie [Ken77]. NIL [Lic86]. Nim [Bar82c]. Ninth [PR90]. NMFECC [Fon85]. No [Ald72, And78, Ano73a, Ano79a, Ano87a, Ano88c, Ano88b, Ano88a, Atk78, Atk79a, Atk79b, Atk82b, Atk83, Bar71, Bar72c, Bar72a, Bar72b, Bar73e, Bar73c, Bar73b, Bar73a, Bar73d, Bar74e, Bar74d, Bar74f, Bar74c, Bar75a, Bar75c, Bar75e, Bar75f, Bar75d, Bar75b, Bar76a, Bar76d, Bar76b, Bar76c, Bar76e, Bar77e, Bar77d, Bar77c, Bar78c, Bar78b, Bar78d, Bar79b, Bar79a, Bar80d, Bar80e, Bar81, Bar82b, Bar82a, Bar82c, Bar84b, Bar84a, Bis79a, Bis79b, Bis81b, Bis82, Bis84, Bow88, Bra75, Bra80, Bri82, Bul72a, Bul72b, Bul73, Bux78, CO88, Col77b, Cor82, Cou85b, Dav74, Dav78, Dea86, Ear77, Edm82, Ell72, Eme84, Eve73, Fen94a, Fin77, Flo73, Flo74, Flo79, For72, Fox79, Gar86, Gru83, Han72, Han78a, Han78b, Haz71, Haz72, Her84, Hop73, Hop74]. No [HW77, How76, Hun72, Hut74, Hut76, Inc86, Jac71, Jac84, Jon74, Ken77, Lan74a, Lan75, Lar71, Lar75a, Lav77, Lav78, Liv75, Llo82, Lon88, Mad82, Mar88, McD71, Mee87, Mer74, Mil72, Mul76, Nee77a, Nic72, Pet77, Pit82, Pra96a, Pra96b, Ree78, RB82, Ree84b, Ree84a, Ree73, Ree75, Ree76, Rob72, Rob81, Rob82b, Rob82a, Rog71, Rog74, Rop88b, Rop88a, Ros74, Sha72, Sim83, SFS97a, SFS97c, Sto88, Tho77, Tho74, Val76b, Val76a, Val78, Val79, Val80, Vel88, Wal83b, Wal81a, Wal82, Wal84b, Wan82, Wel72, Whi87, Wic72a, Wil72, Wil74a, Wil76, Wil84b, Wil87, Wis74, Woo74, Bar77b, Bis86, Bry77, Cav83a, Cav83b, Con77, Cou85a, Edm86, Edw77, Han77a, Rog73, Roh77a, Sha83, Val77a, Val77b, Wal83c, Wal86b]. NOAH [AFFR08]. Node [Wal90]. Node-positioning [Wal90]. Nodes [Wal80, FVF^{+}18, HSY^{+}20, MSWC23]. noise [AAB^{+}21, SIC^{+}20, YYLJ24, ZYYC12]. noisy [GSS^{+}20]. nomadic [AO12, CMR07]. Non [BVGVEA11, BK77, CDH77, CKW02, Cla86, Fin88, FP82, LQ96, Mer93, Pal79, Roh77b, Sel75, Set79, Bar73d, Bas00, CGR00, EYG22, ESRI14, GP01, HHM92, HC16, JSC^{+}10, KM13, Kil81, RGK99, RBD^{+}24, SGCM11, vdP14, TR77]. Non-computer [Pal79]. non-cryptographic [ESRI14]. Non-determinism [Sel75]. non-deterministic [GP01, KM13]. Non-functional [BVGVEA11, RBD^{+}24, SGCM11]. Non-general-purpose [BK77]. Non-‘Interpretive’ [TR77]. Non-intrusive [CKW02, CGR00]. non-invasive [JSC^{+}10, RGK99]. non-layered [vdP14]. Non-local [LQ96]. Non-LR [Mer93, HHM92]. non-numerical [Bar73d]. non-parametric [EYG22]. Non-random [FP82]. Non-recursive [CDH77, Roh77b, Set79, Kil81]. Non-sequential [Fin88]. Non-specialist [Cla86]. non-technical [Bas00]. non-volatile [HC16]. nondeterministic [BT21b]. nonfunctional [JSRM18]. Nonintrusive [RRR97]. Nonlinear [Ram96]. Nonscalars [Coo96]. nonvolatile [PNM^{+}20]. Noosphere [BV06]. NORA [BHNDdMRG23]. normalization [Wu99]. Norman [Pra96a, Pra96b]. North [Ald72, Bar72a, Bar74e, Bra80, Lan74a, Pit82, Val77a, Val78, Wal81a, Whi87, Woo74]. North-Holland [Bar72a, Bar74e, Lan74a, Pit82, Wal81a, Woo74]. Norway [Val77a]. NoSQL [BHNDdMRG23, MNEM21]. Notation [Abb89, BP84b, Mou72, Ros77, WB78, WI85, BB03, WS94b]. Notations [Buh93]. Note [And91, Bar77f, CH73, Col83, Col81, Jon72, Mey78, PD05, Rév85, Sam71a, Bar73f, CY01b, CPW73, Fra74, Ger82, Hug77, JL81, RBL^{+}14a, RH78, SH82, Wic72b, WL72]. Notebook [ZKB^{+}22b]. Notebook-as-a-VRE [ZKB^{+}22b]. notebooks [ZKB^{+}22b]. Notes [Bar74i, Cav83a]. Notification [Lib97b]. NovAtel [Cro91]. Novel [Bar97, Cro91, Add80, AI22, CC18, HIRB24, HLH15, KGL06, LFH^{+}22, LCC14, MZC10, MV16, Mus17, PDROFRM13, SIC^{+}20, SDB^{+}22, TLB^{+}18, YWT^{+}12, ZQWL24]. Novice [Nut76, MR05]. NovidChain [ACKJ22]. NPP [BF80]. NR [Nav01]. NR-grep [Nav01]. NS [SGDA18]. NSEDIT [HW88]. nuclear [DGR^{+}06, Man01]. Nucleus [Hop80b, SWA^{+}75, SM85]. NUMA [CSTL19]. Number [Lem21, PK89, Pra80, Ree82, Ume91, ML23, SV22, ST01].

Numbering [BCS97, Ano76h, DM11]. Numbers [Coh98, CMR92]. Numeric [Lev01]. Numerical [CLL91, Els76, HPC^{+}96, O’N88, Ree75, Bar73d, EP05, Hoh04, MM02, PMC22, Ree73]. NVM [CSM^{+}16]. NVMRA [CSM^{+}16].

O [Bar75f, Bar77e, Edm82, Ree75, KJHG10, WBB15, Yoo96]. O-related [ZCCB22]. O.-J [Bar75f]. Oberon [BCFT95, Wir88a, Wir88b, WG89]. obfuscated [SLJ^{+}18]. obfuscating [WWCW19]. Obfuscation [LW14, FDD20, SLJ^{+}18]. Obituaries [CK13, Hor07a, Hor07b]. Obituary [Hor07c]. Object [AD87, AN88, AZ97b, AZ97a, BBC91, BLL88, BS93, Bud89, BDS^{+}92, BGG01, CCC96, CAC^{+}84, DNSG89, EvG04, Gor87, Gra92, Han90, HUS^{+}91, HZ94, HKV95, Hug93, Jaa95b, JGS^{+}08, Jon83, JVR97, Kan97, KMSS98, Kuh90, LK99, LT91, LD99, Mac96b, Mad95, Men97, Ono93a, PMC05, Pow95, RK89, Rus95, San88, Sef97, SFS97a, SFS97b, SFS97c, SMR89, Thi93, TBA89, Wol92, WP96, YH97, vHE87, ACCD01, BBMG08, BCF00, BLS03, BZD17, CHY^{+}22, CPZ02, CA18, CKB00, CKB01, CKB03, CI03, CKW02, CEF02, CB00b, DDDF17, DS03, DPDA14, Duc11, DM11, ET07, GdLC04, GEF^{+}00, HHRS03, HC00, HLFS05, HKWZ00, JDGCGA12, KRR19, LKCC00, LW14, Liu03, MS99, MM02, MHM01, MMHB08, MF08, NL01, NR04, PLL^{+}02, PK04, PVBB06, PVR99, PA01, SPR^{+}19, Sav11, SZ01]. object [SM02, SC14, SBFF24, TV09, TN98, WIYC20, XWC^{+}17, XZ01, XZ03, YWN^{+}00, dB00, vGB01, HRM00, KNC94, MG13]. Object-Based [SFS97a, SFS97b, SFS97c, Sav11]. Object-JavaScript [HRM00]. object-manipulating [KRR19]. Object-orientation [Rus95].

Object-Oriented [Gor87, KMSS98, Men97, AD87, AN88, AZ97b, AZ97a, BBC91, BLL88, Bud89, BDS^{+}92, BGG01, CCC96, DNSG89, EvG04, Gra92, HUS^{+}91, HZ94, HKV95, Jaa95b, JGS^{+}08, JVR97, Kuh90, Mad95, Ono93a, PMC05, San88, Sef97, Thi93, TBA89, Wol92, YH97, vHE87, ACCD01, BCF00, BLS03, CPZ02, CA18, CKB00, CKB01, CKB03, CI03, DDDF17, DS03, Duc11, DM11, GdLC04, HLFS05, JDGCGA12, LKCC00, LW14, MS99, MM02, MHM01, MF08, NL01, NR04, PLL^{+}02, PK04, PVBB06, PVR99, SPR^{+}19, SM02, SC14, SBFF24, TV09, TN98, WIYC20, XWC^{+}17, XZ01, XZ03, dB00]. Object-process [LD99]. object-relational [Liu03]. objective [FCA12, GM22, SS23, ZZJZ24]. Objects [APS95, Aji95, AN88, BDG93, BNOW95, BTZ94, CCM96, Car98, Cho96, CFL84, LT91, MKD98, TTH97, AM00, BKL^{+}02, DFPT08, IH01, JMM03, MZ00, MP00, NEFZ00, QL13, dRRGdC15, WXR16, vK87]. observable [ABG^{+}22]. observation [TKF09]. Observations [New86, Loe07]. observed [Phi99]. obstacles [Ber82]. Occam [WW89, Bor86, CJ88, KS84, SAC^{+}92, Fis86a, Wil89]. occurrence [CGH^{+}04]. OCL [SW14]. OCSP [BDL09, BCL13]. OCSP-based [BCL13]. octagon [HOY17]. October [KP94]. ODA [HCC96]. ODA-like [HCC96]. Ode [GL97, LG99]. oDect [BBMG08]. ODMG [LK99]. ODMG-compliant [LK99]. Odyssey [WSL03]. off [LPF^{+}11, SXWL17, TS02]. off-the-shelf [TS02]. Office [Bar83b, CW82b]. offline [SM20]. offload [MAG^{+}21]. offloading [CVGCGCFF23, HSC21, HTWS15, SGA20, SSGA20, SMLS23, XCZ^{+}21, ZLZ^{+}21]. offs [PV21, PLR85, RJ09]. Oiled [She92]. OIntEd [WKG^{+}13]. OLAP [LER17, SRGCPB^{+}09]. old [CBC00, SJ79]. Olle [Flo79]. OmniCon [SBcC07]. on-board [MPC^{+}19, VvK99, VC02]. on-chip [IBA^{+}21]. On-demand [KL24, SSO13]. On-Line [Ban71, BMA72, Bro71, Pan72, GJ93, Rag86, TDH97, BMR03, LJ99, WPL^{+}21]. on-the-fly [BGM99]. onboard [FCO^{+}19]. One [Cla89, CRT80, Gut87, Joh78, SMFBB93, SIN95, Wex81b, CL81, Kat17, KL21, KR83, LM81b, PGK^{+}10, VHM^{+}05, FWS74]. One-address [CRT80]. One-pass [Cla89, Gut87, Joh78, KR83]. One-pass-type [SIN95]. one-sided [PGK^{+}10]. One-way [SMFBB93, VHM^{+}05]. Ones [Roh77b]. Ongoing [DWL^{+}15]. Online [GJ22, Poo71b, SIK^{+}16, SY79, Val77b, AWNS18, BHW05, DRG11, Glü12, KCS^{+}20, KYF^{+}22, LKCW13, LLN16, MNEM21, NJGG12a, NJGG12b, NJG14, SH17, WKG^{+}13, YFC06]. only [KKdC^{+}23]. onto [KSSA23, RB89]. ontological [IAPC17]. ontologies [GHM^{+}06, RÁdMRGAM19]. Ontology [ASEB09, GLMS18, MVS^{+}18, STH^{+}18, TW16, TWJ^{+}13, AHH15, BDMP17, DTB12, DGRB15, HB11, hPmKgH15, PNP20, SBS13, WKG^{+}13, ZSAM23]. Ontology-based [ASEB09, MVS^{+}18, TW16, DTB12, DGRB15, PNP20]. ontology-driven [BDMP17]. ontology-powered [HB11]. OntoSuSD [ZSAM23]. OO [TDH97]. OPAL [PNP20]. OPC [GNSP12]. Open [Cas92, Mad95, AMOS19, ACKT20, BV06, BFGL20, CCLN22, DPH16, DP09, EC13, FRBRF19, GLMS18, GN00, GEI^{+}11, GMGL24, HL20, KHS^{+}24, CCK21, Mił10, NMG11, NZPWR22, PNK^{+}23, PdSCJM22, SBS20, SRGCPB^{+}09, SSS22, VRC^{+}06, ZZL24, vGPB10]. open-source [AMOS19, ACKT20, DPH16, GLMS18, GMGL24, Mił10, NMG11, NZPWR22, PNK^{+}23, PdSCJM22, SBS20, SRGCPB^{+}09]. OpenACC [HY20, LFH^{+}22].

OpenACC-like [HY20]. OpenCL [TY14]. OpenGIS [CKL^{+}02]. OpenGL [ASAQ05]. OpenPnP [KBPM^{+}20]. OpenStack [CMF^{+}17, KSSA23, SAEMM21]. OpenTracing [KSSJ^{+}23]. Operating [AMW91, Bad98, BL85, BK77, DH00, FWS74, Fra75, FT79a, FH74, Fra93, Han76b, Han76c, Han76d, HF80, HEV^{+}98, Hus86, JLR79, Kue95, LLCG^{+}89, Lin79, MCG^{+}88, MPP87, Oes71, PU84, Pow79, PJ75, Ree71, RS82, Ree84b, RAB^{+}79, RRR97, RRP95, SF98, Sno78a, SYRS80, TF79a, TF79b, TH86, Val84, Web87, WR84, WR77, vRvST89, BJP^{+}00, Bar76a, CM98a, CM98b, Col79, DD10, EC13, GBG^{+}14, Han77a, KGL06, KS20, Kru82, Lan71, Poh81, Pur76, SJ79, Spi09, Wel72, WAML12, Wu00, ZL84, Dea86]. Operating-system [Web87]. Operation [Cum71, ROFGFR^{+}16, SMKZ06, TRRK21]. Operational [KvEP95, Lor91, Dav78, Har99]. Operations [Coh73, Coo08, FH82a, KS98, Sil92, BMY06, CFL^{+}98, CSM^{+}16, FL02, FZS^{+}17, Wat04, Wet77]. operative [Mac96b]. Operator [De 96, MJ98, CVGCGCFF23, Dun74, Fav07, Sam75, Sav11]. Operators [Fis82, GH72, Kea91a, Pyl80, Ram96, Ram98, MM02, Mid79]. opportunistic [LZWX22, YZL^{+}22]. opportunistically [KV17]. opportunities [BT21a, CHC^{+}17]. optical [BB03]. OPTIMA [WS83]. Optimal [GW96, Li18, QSA90, Vör84, ZB18, FPAF18, LPF^{+}11, OG16, PKK12, ZLX^{+}24]. optimisation [KSK15]. optimised [DRH^{+}23]. optimiser [MKC20]. Optimising [Chi17]. Optimistic [KT84]. Optimistically [PGH^{+}98]. Optimization [Ber85b, CQC98, DF84, DF87, DW89, EM90, ELRV93, GP01, Hoa73, LES95, McK89, Pan72, RG89, Wil79, WW96, WH97, ZB74, APS^{+}11, AI22, AKL^{+}09, AFNG20, BBG04, BMAV05, CWC^{+}21, CS18, CGR00, DDDF17, DHA11, GARSR18, GCARPC^{+}01, GM22, GSS^{+}20, HC79, HLLZ21, IMG^{+}21, JP22, JK14, JSPP21, JT00, Kar21, KSDG24, Li22, LZL^{+}17, LPF^{+}11, CCK21, DCEG23, MG09, MA20b, OEA05, SGAS21, SYG^{+}18, SS23, WSYO11, WC08, ZWQC22, ZZJZ24, dAPMV10, TMS18, Wil87]. optimization-based [GSS^{+}20]. Optimizations [AS97a, CMH91, Han83c, AA14, AvRAF09, KPU04, LvDDM06, PKH07, TC19]. optimize [ALD^{+}23, CS15]. Optimized [GP14, MG94, TWI88, AA23, BBGP01, GDW^{+}20, LKK^{+}18, RK15b, RGS^{+}20b, WP05, YMH16]. Optimizer [Lam81, Ste80, Wes83, WS83, KCP22, MDWD01]. optimizers [KSK09]. Optimizing [Atk82a, CRC18, Er83, GG96, GS90, Har92, JKH22, LQ96, OKN04, TBSI18, WG92a, dOED^{+}20, Bar77e, Dod78, FVF^{+}18, FKR^{+}00, KS08, MGL19, NNK21, PCL^{+}99, UGK^{+}14]. Optional [GF81, FCG83]. options [Ano23m, JSRM18, SGA21]. oracle [RGC^{+}21]. orchestrating [BRS18, PCC^{+}12]. orchestration [BSNB20, GRFFGC^{+}21, RB19]. orchestrations [YSBL22]. Order [BIO94, CZA83, LMSP92, LS96b, PMG71, BB95, CWS07, CFKT17, DGM19, Dro84, Lin98a, ZJY^{+}15]. order-preserving [CFKT17, DGM19]. ordering [JK14]. ordinary [BK23]. ordinates [Vör84]. Oregon [USE85]. organisation [Flo73]. organization [MMK04, NRUP21, TTC^{+}13]. organizational [WLTJ13]. Organizing [Hut79b]. Orientation [Kan97, Rus95]. Oriented [ARV77, BT76, Ell79a, FF80, Gor87, KS87, KMSS98, MTdT93, Mac77b, Men97, Rei72, RHT^{+}13, WP96, AKM17, AD87, ACCD01, AN88, AZ97b, AZ97a, BBC91, Bar15, BGS^{+}13, BLL88, BCF00, Bla04, BLS03, BDMP17, Bud89, BDS^{+}92, BGG01, CPZ02, CMCL03, CCC^{+}16, CA18, CKB00, CKB01, CKB03, CCCZ05, CW91, CA08a, CI03, CCC96, DB09, DDDF17, DS03, DGR^{+}06, DNSG89, Duc11, DM11, EvG04, FSC^{+}21, GdLC04, Gra92, Gu05, GH93, HUS^{+}91, HPB^{+}00, HZ94, HLFS05, HKV95, Hum76, Ise90, Jaa95b, JGS^{+}08, JCL85, JDGCGA12, JVR97, KKLL99, Kuh90, LKCC00, LHC15, LW14, LGRL08, LGP^{+}11, Mad95, MvSdL09, MS99, MM02, MHM01, MF08, MOTG18, MS94, Mus79, NB19, NL01, NR04, Obe11, Ono93a, PLL^{+}02, PK04, PMC05, PLPA22, PL08, PVBB06, PPSO17, PVR99, PŻZ13, RR22]. oriented [RÁdMRGAM19, RT91, RMdL12, SPR^{+}19, San88, Sef97, SMR^{+}12, SRRFGC^{+}10, SM02, SC14, SBFF24, Sti78, SAEGF11, SCLD21, TV09, Thi93, TBA89, TN98, TWJ^{+}13, Val77a, WSYO11, WIYC20, Wol92, WBB07, WYAZ15, XWC^{+}17, XXJS18, XZ01, XZ03, YH97, ZLZ^{+}21, dB00, dOED^{+}20, vGB01, vGPB10, vHE87]. Orion [CJ88]. Orthogonal [CH90, GH84, PNM^{+}20, PPSO17]. Orthogonality [GL85]. orthogonally [MZB00]. OS/2 [OSW92]. OS/360 [Haz74, Lar71]. OS/MVT [BL78, BL79]. OSGi [BVGVEA13, PF09, PŻZ13, VRC^{+}06]. OSI [CDV88]. Osmosis [SSK^{+}22]. osmotic [LZD20, SSK^{+}22]. Other [Gel75, Bar78d]. Oto [TGPS08]. Ould [Gar86]. Our [GMM90, Bis19]. OUSAF [AHH15]. outage [WCT19]. outlier [JZLP20]. outliers [SDB^{+}22]. outline [PB03]. outlining [ZA07]. Output [Coh73, DS94, HKW77, Lev95, Lev97, Pyl79, TR77, TWI88, GRS74, Han83b, Kur78, She07, Wic72b, YLP^{+}11]. Outward [Wal86a]. overall [RGC^{+}21]. overflow [LC03, SS19]. overflows [AGG06]. Overhead [MP79, FBB^{+}14, IBA^{+}21, KGSC01, OKN04, SB03, UWW^{+}05, ZLZ^{+}19]. Overlapping [Coo83, YYSG11]. Overlay [GM77, AFNG20, Han83a]. overload [AHRR22a]. Overloading [MJ98, Sav11].

Overview [RB75, Bar80a, Lev82a]. OWL [BLR^{+}17, BHNDdMRG23, RÁdMRGAM19]. Own [LS81]. ownership [YHK^{+}22]. ownership-based [YHK^{+}22]. Oxford [Bar74c, Roh77a, Whi87].

P [Bar75a, Bar76e, Bar78b, Bar82b, Bow88, Cam85, Cou84a, Gru83, Lan75, Ree82, Rog74, Roh77a, Whi87, Wic72a, AV84, Ber78, CRT80, Hal82, HM84, Hur80, Lin79, Sch89a]. P# [Coo04]. P-CODE [Hal82, Sch89a]. P-Compiler [Ber78]. P/CL [AV84]. P2P [BMM19]. P4 [Rob82a]. Package [Gau95, HKB72, HH80, Mar84b, RC92, Sin81, Thi87, Woo71, BDP02, Dar00, HHPSS19, JK83, Ken77, OW16, SP79]. packaged [Mil72]. Packages [Car97, Val76a, LD14]. Packaging [GW04]. packet [CdA12, Vel88, WAML12]. packing [Has77, WL72]. pad [YYSG11, BM98]. Page [Ano16k, Ano16l, Ano16m, Bar74f, Bis84, Inn77, Mad82, MN80, Ano16j, JDPB08, Wu02, Wis74]. Page-1 [Wis74]. page-shift [Wu02]. pageable [JDPB08]. Paged [Jor78]. Pages [Ano88b, Ano88a, How76, Mar88, Ald72, And78, Ano73a, Ano79a, Ano87a, Ano88c, Atk78, Atk79a, Atk79b, Atk82b, Atk83, Bar71, Bar72c, Bar72a, Bar72b, Bar73e, Bar73c, Bar73b, Bar73a, Bar73d, Bar74e, Bar74d, Bar74f, Bar74c, Bar75a, Bar75c, Bar75e, Bar75f, Bar75d, Bar75b, Bar76a, Bar76d, Bar76b, Bar76c, Bar76e, Bar77e, Bar77d, Bar77b, Bar77c, Bar78c, Bar78b, Bar78d, Bar79b, Bar79a, Bar80d, Bar80e, Bar81, Bar82b, Bar82a, Bar82c, Bar84b, Bar84a, Bis79a, Bis79b, Bis81b, Bis82, Bis84, Bis86, Bow88, Bra75, Bra80, Bri82, Bry77, Bul72a, Bul72b, Bul73, Bux78, CO88, Cav83a, Col77b, Con77, Cor82, Cou85a, Cou85b, Dav74, Dav78, Dea86, Ear77, Edm82, Edm86, Edw77, Ell72, Eme84, Eve73, FMNW04, Fin77, Flo73, Flo74, Flo79, For72, Fox79]. pages [Gar86, Gru83, Han72, Han78a, Han78b, Han77a, Haz71, Haz72, Her84, Hop73, Hop74, HW77, Hun72, Hut74, Hut76, Inc86, Jac71, Jac84, Jon74, Ken77, Lan74a, Lan75, Lar71, Lar75a, Lav77, Lav78, Liv75, Llo82, Lon88, Mad82, McD71, Mee87, Mer74, Mil72, Mul76, Nee77a, Nic72, Pet77, Pit82, Pra96a, Pra96b, Ree78, Ree82, RB82, Ree84b, Ree84a, Ree73, Ree75, Ree76, Rob72, Rob81, Rob82b, Rob82a, Rog71, Rog73, Rog74, Roh77a, Rop88b, Rop88a, Ros74, Sha72, Sha83, Sim83, Sto88, Sto05, Tho77, Tho74, ÜY22, Val76b, Val76a, Val77a, Val77b, Val78, Val79, Val80, Vel88, Wal83b, Wal81a, Wal82, Wal83c, Wal84b, Wal86b, Wan82, Wel72, Whi87, Wic72a, Wil72, Wil74a, Wil76, Wil84b, Wil87, Wis74, Woo74, DBH04]. pagination [CDFV12]. Paging [CMM75, HC97a, Wei72]. paid [Bar82a, Bar82c, Bar84b]. pairs [GK21, Tor23]. Pairwise [GKBK16]. Pak [Mul76]. pandemic [GMGB22, GVBG22]. pandemics [OHD22]. Pao [Bar75c]. Paper [BMC17, CBB17, DDF16, DDDF17, EMD13, FBB^{+}14, GBG^{+}14, HCG^{+}16, dSMH13, Nic72, NRS13, PT14, PH14, POZ^{+}16, PDPM^{+}16, QM13, SBD15, BGS^{+}13, CY01b, Ham79, Lav77, Lav78, Lon88, Rob81, Con77]. paper-back [Lon88]. paperback [Ano87a, Atk82b, Bis82, Bis84, Bow88, Cor82, Fin77, Fox79, Inc86, Lon88, Mad82, Ree84b, Sim83, Bar74e]. Papers [Ano09, Ano13, BP09, BC13, CQH^{+}13, Cor08, DC15, FS11, GB13, GMDM17, GH09, GQ15, HYH15, KKA^{+}16, LSZ16, LMK16, MMOD16, MDH^{+}13, MAW^{+}16, PKvdWB17, QL13, QRD16, aSZP^{+}16, AE14, BPK13, DE16, Flo73, Kap13, Obe11, WJC^{+}14, Wit77a]. paradigm [DCS23, dOED^{+}20]. Paradigms [LKBT92, BLE^{+}08, DDB^{+}18b, MOTG18]. Paragraphs [KP81]. Parallel [AL90, AGR^{+}23, AP84, AP94, AMW91, Bad98, BL90a, BAFR96, BLL88, CS97, CDG^{+}98, CPHS83, CLZ98, EM90, Ell79a, Fri92, FGIS97, GT92, GWM88, Kar76, KS86, KGP96, LS97, Lun89, LKL95, MC91, MV95, McG89, MM80b, Mes80, NC75, QSA88, QSA90, SS89, Str95, THS95, VWB91, Wel83, Whi87, YSM95, dMFÁE17, AFF02, BGSG20, CARB10, CCCZ05, CCE99, CHC^{+}17, CGZ^{+}20, CA00, DB09, DAJ^{+}15, GB13, GVL10, GP01, IK15, KSH^{+}15, KPGH02, KS80, LLLY18, MSK01, SHIS99, SYG^{+}18, SYB04, UN19, Wis93, ZGL^{+}23, ZLZ^{+}23]. parallel/distributed [CCCZ05, KPGH02]. Parallelising [GSWZ95]. Parallelism [CT90, Gra96, RB89, Wri94, CFKT17, HY20, Knu11, Rog23, VGF21]. Parallelization [SI10, DDP07]. parallelize [LPA13]. parallelizing [ZZL^{+}21]. Parameter [Kow81, Sal81a, BMAV05]. Parameter-lists [Sal81a]. parameterised [SYXZ14]. parameterized [Yi12]. Parameters [HW94, Pra80, SKI08, Wil89, Sto94]. parametric [EYG22, HE82]. PARC [CC18, THS95, BAFR96]. PARCIV [SJK^{+}21]. Pareto [LPF^{+}11]. PARMON [Buy00]. Parse [FSO91, Kea91a, KL24]. PARSEC [HHZ^{+}95]. Parser [Coh75, De 96, GL78, GJ88, HHZ^{+}95, KM89, SK96, WC85, Fav07, HC87a, PQ95]. Parsers [BP98, BB95, DP95, Gro90, SMM^{+}84, GIF01, MSK^{+}24]. Parsing [AHS86, Han85, HT82, HP87, HW90, Kop97, Kos90, McK90, Mer93, NL24, CW01, GRVA09, HHM92, Lem21, MFH10, ML23, Ryu16, ST19, Str77, WRD99, Ier09]. Parslow [For72]. Part [Bar74c, Lar73b, PJ75, CK99, Pur76, SFB13, Spo71]. part-of-speech [CK99]. Partial [DS88, HNW^{+}01, KLY20, KKN04, XCG06, Dro84, Glü12, PDPQ22, VH04, WLZ^{+}22, ZA07]. Partial-Match [DS88]. partially [Har81]. participant [Loe07]. participants [KAZ13]. particle [AA20, JSPP21, SDF^{+}21]. particular [CCPY12]. parties [GMNR20]. partition [YZW^{+}12, Che08]. partition-based [YZW^{+}12]. Partitioned [Hun81, EHV99, TRO17]. partitioned-grid [EHV99]. Partitioning [LFW96, GAAA22, HJ14, VS18, XLLY19]. partitioning-based [GAAA22]. ParTransgrid [ZLZ^{+}23]. Parts [WC04]. party [Cho98]. Pascal [Hay80, Ten85, Amm77, BD76, GLN76, Haá82, HE82, LP83, MS83, NW78, Tsi82, WQ72, Wir71, Ano80b, ABBH^{+}79, ADDM84, AP84, Atk79c, AN81, Atk82a, BS84, Ber78, Bis79c, Bis79d, BWA82, BO83, CC87, CD84, CGHP79, Com79, CW82a, Com83, CL82, CMH85, CRT80, DS86b, DSW82, FM86, Fre81, Ger82, GKLM79, Han76b, HM82, HT86, Hur80, JCL85, Jos79, Jos80, KE85, Ker82b, KS84, KS80, Kru82, LF82, Liu86, Mac79, MTT81, MTT83, MS90, Mar79, Mar84b, Mat80, McC83, Moh77, NW84, ORT81, OW83, Par85b, PV84, PD81, Rav82, Rob83b, RS76, Sal79b, Sal79d, Sal79a, Sal81b, Sch80, Sch89a, SFIK80, Shr78, Shr79b, Shr79a, SM81, Ten78, WC81, Wal86a, WSH77, Wel78b, WB79, WB85a, WB85b, Wil80, Yip84, You81, Ano79a, Atk79a]. Pascal [Atk82b, Atk83, Bar77c, Bis79a, Bis79b, Bis81b, Bis84, Fin77, Ree84a, Atk79b, Bis82, Rob82a]. Pascal-Again [Sal79b]. PASCAL-Compiler [GLN76]. Pascal-P [CRT80, Hur80]. Pascal-Plus [KS84, WB79]. pass [Cla89, Gut87, Joh78, KR83, LM81b, Mös88, SIN95]. Passing [Geh90, Gen81, HI85, Kow81, Sta82, Bre82, GB13, SZ00, TA91]. past [DH00, YMH16]. Patching [Bis87]. Path [AW93, PSR83, SW86a, WW91, DFD23, HNW^{+}01, KCCV05, DS86b]. Path-free [SW86a]. pathfinder [FCO^{+}19]. paths [MG94, Tor23]. patients [SKD^{+}22]. Pattern [DB86, FS13, Har80c, JPM17, Liu86, PJ76, Ric79, Som82, VSM87, Abb78, AKW79, ACF13, Ano23m, AG06, BKP^{+}22, BD14, CFKT17, DGM19, Fen01b, FBMA05, Haf13, Ier09, KAZ13, KA13, Kim15, NHB23, Nav01, NWE99, NK07, OM16, PLR13, PRTS06, PH14, PPN^{+}23, RZ17, RL23, SGA21, Sas79, SK03, STH^{+}18, SSO13, WC04, Zdu07, vdMF13, FS11]. Pattern-based [JPM17, BD14, RL23, SK03]. pattern-matching [Ier09, Nav01, NWE99]. Patterns [Kot96, Men97, WW91, AG06, BHJ^{+}18, Bar15, BVGVEA11, CS17, DE16, DZS09, EM12, HRS^{+}09, HC13, KAZ13, MG13, PMC05, SN07, SBF19, TWJ^{+}13, WWGP10]. PAxSPL [MRBB19]. PBASIC [Hop80a]. PBX [KKLL99]. PC [Yu96]. PCODE [WS83]. PDB [MTdT93]. PDE [Hoh04]. PDF [PB03]. PDG [NP98]. PDG-based [NP98]. PDMA [XZW^{+}22]. PDP [BD76, DM84, Har71b, HGWBS75, Les72]. PDP-10 [Les72]. PDP-10/LSD-1 [Les72]. PDP-11 [DM84, HGWBS75]. PDP-8 [Har71b]. PDP11 [JB84]. PEARL [HK84a]. Peck [Woo74]. PEDANT [Mos73]. pedestrian [DNL^{+}20]. Pedro [RC10]. Peephole [DF84, DF87, DW89, Lam81, McK89]. Peer [PGH^{+}98, HYT13, MR07, WN06]. Peer-to-peer [PGH^{+}98, HYT13, MR07, WN06]. Pemberton [Rob82a]. Pengelly [Col77b]. Penker [Cor99a, Cor99b]. Pentium [RGK99]. Peonage [Bar84b]. People [TB73, AJ04, LM15]. PEP [VV84]. perceive [CCK21]. perceptions [NDKL^{+}24]. Peregrine [JZ93]. Perfect [CW91, DM11, BT89, KSH^{+}15]. Performance [AKL^{+}09, AR93, AO88, AW96, Bai73, Bar74a, BP02, Bre82, CCC^{+}16, CG96, CSR93, DS88, DD21, DD90, ESRI14, HJS89, Har91, HJ88a, Jal82, JKRS85, KGP96, LD95, MNH04, MY87, NKW06, Pal78b, PMY97, PV84, RT77, SF98, Sre76, SFS97a, SFS97b, SFS97c, ST77, SR91, Tse97, TA91, Wha93, WS94b, WH97, YSM95, vRvST89, BMR03, BDQR23, BPDL23, CNRB13, CMPQ23, CLKG16, CSZO23, CCE99, CZ04, CGZ^{+}20, CLC99, CGR00, CMTCC^{+}17, DSD^{+}05, DHMS11, Dun75, Fen01b, FMC18, FSRF^{+}22, FFRF19, FFRFS19, GA12, GCK^{+}02, HYY23, HLLZ21, HC12, HHPSS19, IMKN12, IB13, IK15, JZ93, KKL17, KMY^{+}05, KCP22, LTK^{+}20, Li22, LCC14, LCZ08, LZ10, MKA^{+}22, Mar22, PGK^{+}10, PCL^{+}99, PDPMM17, QM13, RGK99, RGS^{+}20b, RW04, RD14, SJA^{+}04, SFC^{+}21, SRGCPB^{+}09, SXWL17, TTC^{+}13, WW09, WY18b]. performance [WBGS22, WMJ04, WBB07, WSL03, ZG06, ZWF^{+}22]. Performance-driven [CCC^{+}16]. performance-optimized [RGS^{+}20b]. performance-portable [WBGS22]. Performance-steered [BP02]. Performances [Yas94, CGIP15]. performant [RCH^{+}23]. Performing [Sil92]. Pergamon [Bar74c, Roh77a]. Perils [KV17, Fen01b]. Periodic [Jor90]. Perl [KBBS05]. Perlman [Ano79a]. Perly [Hes91]. permission [FZW19]. Permissive [TK72b]. PERQ [Coo83, MBB^{+}86]. Perrott [Bux78]. Perseus [ZL84]. Persian [Ber99]. Persistence [RC89, SWA^{+}97, BHK^{+}04, PPSO17, PNM^{+}20, SBL06]. Persistent [ACCM83, BDG93, CAC^{+}84, HHR93, HC00, PDC^{+}98, SMR89, AM00, Day00, DH00, HKWZ00, HHPSS19, LCW07, MZB00, MBB^{+}86, MBG^{+}00, PSRCC02, PNM^{+}20, PA01, YWN^{+}00, ZWF^{+}22]. persistently [WP05]. Personal [Cra77, Del82, FHJ94, GW84b, Rei84]. personalized [BFPAGS^{+}08, FG08, HLH15]. Perspective [Mad95, dCCCdAC20, DBH04, FFRF19, HLLZ21, JJI22, LYLY20, SSGA20, Pra96a, Pra96b]. Perspectives [PGH^{+}98, Pow95, BGS20]. perturbation [SJKL94]. Perturbing [EP05]. pervasive [AYdS^{+}06, BC13, HIR06]. pervasiveness [EAB^{+}03]. Peter [Bis79b, Pit82, Hor07b]. Petri [HAM18, Wen80, dMFÁE17].

Peucker [VWB91]. Pewss [TLC^{+}18]. PFACC [HY20]. PFORT [Ryd74]. PGS [KM89]. PHANtom [FG14]. Pharo [FG14]. phase [CZL24, JK14, LJL^{+}10, LC07, dSMH13]. Phases [HS89]. PhD [Ano09]. Phenomenon [WIS^{+}97]. Philippe [Dea86]. Philosophers [Car82]. philosophy [Bar78b]. phishing [MGBO22]. PHOEBE [PDPMM17]. phones [LHK99]. Phrase [Hud72]. PHS [PHS84]. physical [CWC^{+}21, CEMS24, CB00a, CM08, DLX^{+}22, FBH^{+}23, GLKZ21, JXG^{+}21, JFZ^{+}21, Pet01, SJK^{+}21, SDC04, TL14, WHH21, XZX^{+}21, YQL22, ZZJ21, ZZL^{+}21, ZKZ^{+}21]. physical-design [TL14]. physics [WLZ^{+}22]. physics-informed [WLZ^{+}22]. physiological [GDW^{+}20]. Pi [MVOD19]. PIC [Ker82a]. PICASSO [KKS88]. Pictorial [MTdT93]. Pictures [Buh93, Wyv77]. PICTURES-68 [Wyv77]. piece [LZWX22]. Pierson [Wis74]. PIGS [PN83]. Pilkey [Liv75, Pet77]. Pilot [RT78]. Pin [Pit82]. Pin-Shan [Pit82]. pint [Coo85]. Pipeline [WSL^{+}20, CDF^{+}24, LPA13, MBGC21, SAA^{+}20]. Pipelined [BF97, HATvdW99]. pipelines [QH21]. PIT [Zel72]. Pitfalls [Wen90]. Pitman [Atk79b, Atk83, Cou84b, Ree76]. PIWI [CDGP93]. PK [NH03]. PKI [BDL09, LCZ08]. PL [Bar73b, Bar76d, Ell82b, EBD^{+}74, Els76, Hać84, Joh78, Mar84a, Neh79, Pal74, WA77, WKL76, Zel77, Val76b, Nic72]. PL/1 [EBD^{+}74, Pal74, Nic72]. PL/1-like [EBD^{+}74]. PL/C [Bar76d]. PL/I [Bar73b, Ell82b, Els76, Hać84, Mar84a, Neh79, WA77, WKL76, Zel77, Val76b]. PL/I-like [Neh79]. PL/M [Joh78]. PLAC [FS11]. place [BFGL20]. Placement [FR91, AHRR22a, ARK21, BBH^{+}22, CS18, FVF^{+}18, GWZ^{+}20, HSY^{+}20, Man18, MKC20, MSWC23, SGAS21, SESK23, TGAS22, ZLZ^{+}21]. placements [ZXW^{+}17]. plagiarism [BTZ07]. PLAIN [KWW81]. plan [NB19]. plan-oriented [NB19]. Planar [Hop71, VL73, Ple99]. Planning [Cra77, GHM^{+}06, TVSG21]. plant [DGR^{+}06]. Platform [BHMV09, Eng06, JGSG^{+}21, ACKJ22, ALKL19, AÇCT22, ALG23, ACAB22, BAM^{+}20, BBMN18, BOPN12, BMM19, CCLN22, CCT01, CCP06, CMR07, CB00b, CHS^{+}05, DM15, DRH^{+}23, FCBF^{+}21, HPP22, HLW08, HKC^{+}12, HYT13, LW19, LQS^{+}24, MMHB08, MTPC14, MSZ^{+}24, QRD16, Ric00, SMR^{+}12, SGDA18, SAEGF11, TJB^{+}19, TRO17, TLC^{+}18, Wij05, YMY17, ZGL^{+}23, RHT^{+}13]. Platform-independent [BHMV09, Eng06, JGSG^{+}21]. platforms [BBH^{+}22, BR01b, Che22, FSRF^{+}22, FFRF19, FFRFS19, HKM^{+}09, JGC^{+}21, PT14, PF09, RRFD23, SDIP23, SSM11, WW09, WCS^{+}17, ZCCB22]. PLaTHEA [LM15]. Play [Buh93, JDGCGA12]. Playing [Coo83, FFF^{+}13, Ano73a]. PLEIADES [KPGH02]. Plenum [For72]. Plotter [GF80, Ano71d, Ano72b]. Plotting [HF73, Ear76, SP79, Bar75c]. PLT [Ste02]. plug [ACF13, BN13, GK14, GH02, Kar14, MSB^{+}13, NNL^{+}14, SMM13, ZCO13]. plug-in [GH02, MSB^{+}13, NNL^{+}14]. plug-in-based [ACF13]. plug-ins [BN13, GK14, Kar14, SMM13, ZCO13]. pluggable [Bar15]. Plugging [RBL14b]. plugin [PNK^{+}23, San17]. plugin-based [San17]. Plumb [QH21]. Plumbr [SSST15]. Plus [AP91, KS84, WB79]. PMS [Wis93]. PMS-Prolog [Wis93]. Pocket [Atk83, Cou84b, Atk83, Cou84b]. Poel [Val77a]. POET [Yi12]. Point [NC75, VS80, Far88, Has77, SF88, Ume91, Ush77]. Pointer [FH82a, PF97, Rav82, TD94, SYXZ14]. Pointers [BDG93, Rad80, AE14, Sav04]. Points [Smi89, LCA09, LLN16, ZGG07]. points-to [LCA09, LLN16]. policies [CSTL19, DTJ89]. Policy [BS85, GJ00, Rob84, WLTJ13, BSC^{+}05, FB23, LGP^{+}11, TXHL18]. policy-centric [TXHL18]. Policy-directed [GJ00]. Policy-driven [WLTJ13, LGP^{+}11]. policy-writing [BSC^{+}05]. Polish [Bro77, Bro72, CH73]. Pollack [Hop73]. Polling [GC84]. Pollution [DP85, LZD20]. Polygraph [RW04]. Polymorphic [BMTA16, GH84, HG84, Man88, NSM16, Wal86b]. Polymorphism [BR95]. Polynomial [Shn73, ZZL^{+}21]. PolyRec [BCRF22]. Ponto [Bar73c]. Pool [SY79]. Poole [Sha83]. Poor [Pap79]. POP [Bar74a, Dun75, MS74a, Hun72]. POP-2 [Bar74a, Dun75, MS74a, Hun72]. Popplestone [Hun72]. popularity [MBBS21]. population [GLMS18]. Populations [Car85b]. Portability [BCP71, HF76, HWS^{+}88, Lar81, LPT78, NW78, PV84, PD78, Ric71, Sab76, SSM11, TKB78, Mar22, Wir72]. Portable [Ard87, BCP79, Bin06, CC73, DM84, GH81, Gri75, Han80a, Han80b, HC87b, JKRS85, Jen89, JGR89, Lan75, LDG^{+}96, MP81, MR96, MS80b, NPW72, Oni85, PM81, RS90, RAB^{+}79, SF85, Sch80, Sch89a, Ste84, Ste92, TF79b, Val84, Wai75, WS83, BMR00, Buy00, EC13, Fra79, GCRD04, Han83b, Ibs84, MZB00, MFH10, Rei82, WWJ07, WBGS22, WCS^{+}17, ZL84]. portal [MVS^{+}18]. portfolio [SYG^{+}18]. Porting [HL92, CD15]. Portland [USE85]. PORTREP [JGR89]. ports [BSM^{+}21, HMN11]. posing [Sny08]. Positioning [Wal80, FR09, Wal90]. Posits [PMC22]. POSIX [BT21b, TEGF08]. Possibilities [Sch89b]. possible [FLPM20]. Post [Abe10, KSST22, NY78, SB22, Abe07, Bar82a, Bar82c, Bar84b, KRB21, NT20]. Post-quantum [KSST22]. post-release [KRB21]. post-transformation [NT20]. postage [Bar78d]. postage/handling [Bar78d]. Postfix [Ram98]. Postlude [GLN76]. Postman [Thi03a]. postpaid [Bar84a]. postprocess [Cer18]. PostScript [Ber99, NMRW98]. pot [Coo85]. Potential [Lav77, Tal71]. Pound [Mac96a]. Power [Bar84b, Har80b, McI90, QSA90, WBS82, BIO94, DGR^{+}06, EYG22, KCH07, KCP22, LZ10, LZL^{+}17, SM24, Tsi82]. power-aware [KCH07]. power-efficient [SM24]. powered [HB11, VVG^{+}24]. powerful [BHT^{+}23, Fav07, YQL22]. pp [Fen94a, SFS97a, SFS97c]. PPM [Fen12]. Pract [XZ01, XZ03]. Practical [Ano09, AZ97b, AZ97a, BP09, BMS83, BCHS98, BH92, CCG14, CDV88, Eme84, Er85, Gar86, HP88, Hof89, Hor80, HKV95, KFMF18, KMSS98, LM81a, LS96a, MEP96, NSM86, OSW92, PK89, PV21, RAP21, Sau88, She07, TSO19, Var93, Woo72, dV89, BST10, Bis79a, Col77b, Edm82, GMPL11, Lon88, LWZ^{+}21, Maa06, PBGM18, RaMB^{+}24, Rog23, SYXZ14, KPK^{+}18, Ano88b]. Practicality [TT82]. Practically [FK16]. Practice [Ano23m, BW95, Cor08, SFB13, XYW^{+}24, vdRW79, BCSW20, BCP19, GMP^{+}21, JKH22, MGL19, OOG19, WZH01, BC21, Sha72, Wal81a]. practices [NDKL^{+}24, PCBR18, RCMZ13, TBT23, vGPB10, And78]. Practicing [Fel81]. practitioners [Ozk18]. Pragmatic [CL83, NS08, MW13]. praise [Dod78]. PRBP [FB23]. Pre [DW73]. Pre-processor [DW73]. Precedence [De 96, BGSG20, Dun74, Fav07, Sam75]. Precise [Kue95, HOY17, PLR18]. precisely [WCsH16]. Precision [ST79, VS80, GAQO23, OKN04, Rob79, ScG09]. Precompiled [Lit93, LLM05]. precondition [RaMB^{+}24]. Predicate [Har84b, HL91]. Predicate-based [Har84b]. predicated [PQ95, XCG06]. predicated- [PQ95]. Predicates [PH86]. predict [GM22, SKD^{+}22, WHS^{+}00]. predictable [VvK99]. Predicting [Als22, KRB21, LLLY18, RGS^{+}20a, SC22, ABC^{+}21].

Prediction [HF76, TMS18, WJ93, AA20, AA21, CFLC14, DDP07, DSD^{+}19, Fen01b, GKWS11, HBC15, Kil19, KIB09, LQS^{+}24, MKA^{+}22, RBL^{+}14a, SZ01, SOKE23, TRGA18, WHH21, WLS^{+}21, ZML13, ZYW^{+}20, ZDY^{+}17, ZWX^{+}21, ZYW^{+}22, ZZJZ24, ZQWL24]. predictor [MMK04]. predictors [NM06]. PREEMPT [dOdO16]. preferences [DWL^{+}17, HIR06]. Prefix [Ram98, Dun91, LM06, OG16, PV21, YOM^{+}07]. prefix-sum [PV21]. Prentice [Bar73c, Bar74d, Bar75d, Bar75b, Bar76c, Bar80e, Edw77, Edw98a, Edw98b, Lar71, Ros74, Wri98]. Prentice-Hall [Bar73c, Bar74d, Bar75d, Bar75b, Bar80e, Edw77, Ros74, Bar76c]. Preparation [CH88, GW84b, HSM81, WBS82]. Preprocessing [Set84, ZLZ^{+}23]. Preprocessor [BF80, Com78, Com79, Dew86, Hay83, Ker75, MS80b, OM96, TY80, BN00, DC03, Iwa02, Wya84]. preprocessor-aware [BN00]. Preprocessors [LHH^{+}91, MP79, OM96, TWI88]. Presence [CK94]. Present [Moh81, DH00]. Presentation [RR85, WRR97]. Presentations [WKD96]. Preservation [ADM96]. preserving [ACKJ22, CFS^{+}22, CFKT17, DGM19, FKL^{+}13, LS16, LLJ^{+}23, WMSY12]. Presorted [McG89]. Press [Ano88b, Atk78, Bar73a, Bar73d, Bar74f, Bar75f, Bar80d, Bar81, Bis79a, Bis81b, Bis84, Bux78, Cou84a, Dav78, Dea86, Eve73, For72, Gar86, Han78a, Han78b, Hop74, Hun72, Inc86, Jon74, Liv75, Lon88, Mad82, Mer74, Pra96a, Pra96b, Ree78, Rob72, Sha72, Sha83, Tho77, Wic72a, Wil72, Wil87, Bar77d, Bry77, Han77a]. pressure [SSRAH15]. PRESTO [BLL88]. PREttier [BB95]. Pretty [Vau80]. Pretty-Printing [Vau80]. Prettyprinter [Jok89]. Prettyprinting [BS89]. Prevent [KLY20]. Prevention [HJS89]. previews [Chi17].

PRICE

[Atk83, Ald72, And78, Ano73a, Ano79a, Ano87a, Ano88c, Ano88b, Ano88a, Atk78, Atk79a, Atk79b, Atk82b, Bar71, Bar72c, Bar72a, Bar72b, Bar73e, Bar73c, Bar73b, Bar73a, Bar73d, Bar74e, Bar74d, Bar74f, Bar74c, Bar75a, Bar75c, Bar75e, Bar75f, Bar75d, Bar75b, Bar76a, Bar76d, Bar76b, Bar76c, Bar76e, Bar77e, Bar77d, Bar77c, Bar78c, Bar78b, Bar78d, Bar79b, Bar79a, Bar80d, Bar80e, Bar81, Bar82b, Bar82a, Bar82c, Bar83a, Bar84b, Bar84a, Bis79a, Bis79b, Bis82, Bis84, Bow88, Bra75, Bra80, Bri82, Bul72a, Bul72b, Bul73, Bux78, CO88, Cav83a, Col77b, Cor82, Cou84a, Cou84b, Cou85b, Dav74, Dav78, Dea86, Ear77, Edm82, Ell72, Eme84, Eve73, Flo73, Flo74, Flo79, For72, Fox79, Gar86, Gru83, Han72, Han78a, Han78b, Haz71, Haz72, Hop73, Hop74]. Price [How76, Hun72, Hut74, Hut76, Inc86, Jac71, Jac84, Jon74, Lan74a, Lan75, Lar71, Lar75a, Lav77, Lav78, Liv75, Llo82, Lon88, Mad82, Mar88, McD71, Mee87, Mer74, Mil72, Mul76, Nee77a, Nic72, Pit82, Pra96a, Pra96b, Ree78, Ree82, RB82, Ree84b, Ree84a, Ree73, Ree75, Ree76, Rob72, Rob81, Rob82b, Rob82a, Rog71, Rog73, Rog74, Rop88b, Rop88a, Ros74, Sha72, Sha83, Sim83, Sto88, Tho77, Tho74, Val76b, Val76a, Val77b, Val78, Val79, Val80, Vel88, Wal81a, Wal82, Wal83c, Wal84b, Wan82, Wel72, Whi87, Wic72a, Wil72, Wil74a, Wil76, Wil84b, Wil87, Wis74, Woo74, Bar77b, Bry77, Cam85, CL81, Con77, Cou85a, Edm86, Edw77, Han77a, HW77, JT00, Ken77, Pet77, Roh77a, Val77a, Wal86b, WLS^{+}21]. Prime [BIO94, JB84]. Prime-power [BIO94]. primer [Fin77]. Primitive [Gen81]. Primitives [Com82, Hop86, Thi80]. principal [KSSA23]. principle [BLM00]. Principled [RaMB^{+}24]. Principles [And78, HG84, DPS03, LD99, TAG^{+}10, Bar77d, Bra75, How76]. printers [Kha86].

Printing [Kha86, Vau80, DBvG^{+}23, Gou86]. printouts [FIÁSLSAR05]. prioritization [HTX^{+}22]. Prioritized [Hun81, FB23]. Prioritizing [GSPA^{+}11, SJA^{+}11, GVG^{+}18]. Priority [Per85, MGT20]. priority-based [MGT20]. Privacy [AB20, AO12, LS16, WMSY12, ACKJ22, BMM^{+}18, CFS^{+}22, FKL^{+}13, Haf13, LLJ^{+}23, LQS^{+}24, MHN18, OHD22, ZYYC12]. Privacy-aware [AO12, BMM^{+}18, LQS^{+}24, OHD22]. Privacy-preserving [LS16, ACKJ22, CFS^{+}22, FKL^{+}13, LLJ^{+}23]. PrivAPP [BMM^{+}18]. private [CMF^{+}17, SAEMM21, ZKB^{+}22b]. pro [CJ73]. Proactive [ASGA23, KGAR19, BHR15]. Probabilistic [RBL^{+}16, BLR^{+}17, FPAF18, ZB23, XZW^{+}22, ZLWG11]. probabilities [WP00]. Probability [Fen96, Mof99]. Probe [Gai86, WMJ04]. Problem [Car82, Dro86, Kra97, LMSP92, McG82, Mon96a, Sch86, SO77, TDH97, YH97, Atk79a, BGSG20, BOPN12, EM12, FCA12, GZW^{+}22, HBK20, Kil12, LQ04, Maa06, MSR^{+}07, Mus79, NBOS99, Par85b, PV21, Plu74, CFL^{+}98, Thi03a, Wal83c]. problem-oriented [Mus79]. problem-solving [LQ04, MSR^{+}07]. Problems [Cor88b, GSWZ95, RM75, RC92, Sha80, Ano79a, BM01, CCQ16, Deo10, EHV99, Gru83, LV20, Nic98]. Procedural [HW94, Sos95, Thi80, dODP21, Ron99]. Procedure [CC84, Er83, FZ98, GG96, LQ96, MMN79, Pal86, Sta82, Sti78, Sto94, TN98, Bar77b, DE16, KF02, Mor77, OJP99, Rin07, Tsi82]. Procedure-oriented [Sti78]. Procedures [HKW77, Kno81, Man88, Mid86, Roh77b, Roh81, Sal81a, Wic77, Wil83, YL95, Bar77b, Wic72b, Wil74b]. Proceedings [Bar73e, Val77a, Val78, Lan74a, Rob72, USE85, Woo74, Bar75e]. Process [Bha88, CS91a, CC00, CG93, DO91, DF95, FF80, Har80a, LL91, LS97, Pal82, Ped86, RA95, RB81, RT91, SH98, Sti78, Tra79a, TP92, Web87, Wil84a, AGRS11, BBMN18, CGP^{+}06, CS02, FPT07, Gal79, GLD^{+}21, GW04, GMC00, HAM18, JTG^{+}11, LBP^{+}13, LD99, LPGBD^{+}19, MKC11, MRBB19, Mar86, DCEG23, MC23, PCdGPP12, RH78, RGC^{+}21, RMdL12, SDIP23, UGK^{+}14, VvK99, Wal83b, XLLY19, ZZ11]. process-aware [RMdL12]. Process-based [LS97]. process-driven [ZZ11]. Process-graph [Bha88]. Process-oriented [RT91, Sti78]. Process/ADT [CS91a]. ProcessAtlas [BBMN18]. Processe [Fid88]. Processes [Col88, Gen81, GWM88, GJ93, Han76d, Har85, HD86, KS86, MS90, MD88, SCR94, Smi85, Str82, Wis93, Wre88, YR92, HC99, LTW^{+}21b, SSCdA^{+}03, SSS22, YZYL07]. Processing [Bar83b, BAFR96, Ben77, Bro86a, Bul87, CD94, CH88, Coo96, CW82b, EM90, Ell79a, EV89, Fil98, Ham77, Inc86, Mar86, MT84a, NC75, New86, Nil90, O’N88, PS81, QSA90, RS86, SS89, WSB96, Wet80, Wic72a, ALKL19, AKW79, ANSK16, Ald72, BCPSC18, BD14, CRC18, CCCZ05, CHC^{+}17, CYJ^{+}22, Col77a, DLWF17, Deo10, DHWZ14, DHMS11, EvG04, GAF^{+}09, GA12, Ged14, HL03, HTWS15, JPG^{+}17, JKH22, KBBS05, KPU04, Kru82, KKA^{+}17, Lav77, LS22, PKN^{+}12, PP16, QH21, SDKS16, SAY16, SHGG16, TAG^{+}10, ZWML14, ZLY18, Bar72a, Ree76]. Processor [BO83, Ell79a, Ise90, Jor78, KNPS88, MS80a, MV86, Pas87, Pry85, Wit83, AV84, DW73, Fis86a, KCCV05, LLJ12, LJL^{+}10, Sas79, SPPH10, Web87]. processor-based [KCCV05]. Processors [BS80, Har92, Lan75, SY86, BSMV09, GXN10, IMKN12, OKN04, PKH07, SBG^{+}05, Han78b]. Produce [BS90b, NPW72, Wit77a]. Producer [MLR19, AvRAF09]. producer-side [AvRAF09]. Producing [Ber85a, KP94].

Product [CCLN22, Als22, ADH^{+}00, BBS11, DPAG11, FV03, Han11, MRBB19, SL04, SYG^{+}18, Wij05, dSdMSNO^{+}11, vGPB10]. Production [Cd91, LPT82, NHP81, Sch82, Sch76b, FSC^{+}21, NSW77, Sch83b, ZRX^{+}99]. Productive [Ano88c]. Productivity [PVR99, Val76a, KV14, MS99, Phi99, vDD11]. Products [Her84]. professional [Mar88]. Professor [Wir77a]. proficient [WSWR23]. Profile [BA78, CCPR91, CMH91, CMCH92, Els76, Yuv78]. Profile-guided [CMCH92]. Profiler [GKM83, GH93, AKNJ21, DFW^{+}12]. Profilers [PF88]. profiles [HRS^{+}09, KKS10, LXY^{+}11, MBV^{+}10]. Profiling [Bis87, Car86, Deb88, Fit77, Mat94b, PWBK07, RCC91, SHIS99, ZCCB22, BBRB12, Bin06, BSMV09, BHMV09, HSD10, MMM18, McK99, Spi04]. Profit [CLCC15]. Program [AB88, All83b, AJ78, BF75, Bou91, BCP71, Bro81b, Car85b, CLW90, Cd91, Com79, CGWL80, CK78, Daw77, DV85, Dro85b, Ein88, Fit77, Fra80, Han76b, Han78d, Hay83, HI85, Hoa73, Hop71, Hug79, Hur80, KPT86, KS89, Lan82, LB94, LT85, LAD^{+}94, Lop89, MJ83, Mat83b, McC83, MM80a, MM80b, MM85, OE92, PZA87, PF88, Ric76, RT77, Sco77b, Sil92, Sil81, Sos95, STS83, VMJ97, Wil84a, WR78, Bar77d, Ber82, BRL^{+}15, Bow88, Bri84, BWA82, CC00, CGR00, Fel79, FDHH04, GNV88, GHBH05, HCG^{+}16, Inc85, JJL17, KKS10, LBP^{+}13, LCY07, dSMLP22, Mos73, NW84, NGLL14, PJJM21, PNP20, SD75, SO07, SLJ^{+}18, Tse13, Ush77, WPL^{+}21, WBN^{+}20, Wic81, Yi12, YMH16, dCGG13, vDD11, Bar76c, Inc86]. Programmability [KGP96]. Programmable [Fra82, Lev82b, BK23, ZZJ21]. Programmatically [MTPC14]. programmed [Val76b]. Programmer [Fel81, GS76, GJ88, VHM^{+}05, vDD11]. Programmer-friendly [GJ88].

Programmers [Chv79, MR05, RSW^{+}23, Zel77, Ano88a, Bar80e, Mar88]. Programming [AH85, AO88, Bad98, Bar76d, BHR15, BCL^{+}94, BA81, BLL88, Ber88, BdJ80, CDG^{+}98, CV84, CPW74, Cou84b, CM85, CFP83, DNSG89, EG84, EMVW83, Fai87, Fel81, FHS92, FYP93, Fle90, Fox78, FGIS97, GC84, GR88, GW96, GM85c, GF80, GH84, HH88, Han87a, Han94b, Han80b, HHR93, HG84, Hel95, HZ94, HG89, HW98, Hua87, HC87b, Hum76, Ian90, Inc83, JGT95, JP79, Kat83a, KPH76, KM79, KD83, Knu92a, Knu92b, KvEP95, KP90, KCCV05, KS80, Kuh90, Lan74b, LGC84, LT91, Lev98, Lew83, LS97, Lyo85, Mad95, MS74b, Mar79, MT94, MM97, Mor80, NPJ79, Nic72, Nut76, OW89, Ols90, Pag84, Pal76, PP80, PCM83, PL91, Plu77, PR98, PN83, Pyl79, RTL^{+}91, Ram83, RM91, Ree75, RW81, RT91, SB83, SS95, SW74, Sha78, SAN^{+}81, Shr76, SM81]. Programming [Tag88, Thi80, Thi93, Tra79a, TBA89, TAAT84, Val76a, WG92a, WR95, War80, Wei72, Wel78a, Wex81b, Wir88b, Yip82, vdRW79, And82b, Ano76h, Ano79a, AM00, Atk78, Atk82b, Atk83, BVB^{+}12, BMR14, Bar72c, Bar74c, Bar79b, Bar15, BAF03, Bis86, Bre02, BPS00, CDRV03, CFL^{+}98, CCC^{+}16, CCCZ05, CZL21, Coh74, Čuk16, Day00, FMT04, FCR^{+}09, FSC^{+}21, Gal79, GL05, GMO01, GA12, Glu74, GVL10, GZX^{+}21, GG08, HR06, Ham79, HGWBS75, HY20, Jon85, JT00, KAS^{+}14, KS08, LS03, LV20, Llo82, LQ99, Mes80, MSB18, MSZ^{+}24, Nee76, OW16, PM17, PK04, PL08, Pei02, QL13, Rei84, RBL^{+}16, Ron99, Ros74, RPP07, SH03, Sav04, Sch83b, Sim83, Spi02, Sto88, Thi12, TGPS08, TN98, Val79, VC21, VV84, Wal86b, Wic72a, Wit77a, Wu00, Zel77, ZHO^{+}19]. programming [ZWSS15, dB00, Bar75f, Haz72, How76, Atk79b, Bar72b, Bar74c, Bis82, Bul72a, Cor82, Cou85a, Haz71, McD71, Roh77a, Bis79b, Bis81b, Bul73, Hun72, RB82]. Programs [Abb89, AJT79, All89, BA86, BAP87, Bri87, CC87, CMCH92, CG95b, CV84, CC77, CW92, Col77a, Con85, CT90, CP76, Deb88, DDZ94, DR92, Els76, EV89, Fin88, FM78, FKD14, Gai85, Gai86, Gor87, GKM83, Ham77, Han81c, HV88, HMS88, HG81, Hol83, HP83a, HMS^{+}95, Jac85, Jal82, JBCB79, KS87, Kaw79, Knu71, Koo87, Lar90, Lee83, Lib93, Lib97a, Liv75, Mar85, Mat83a, Mat94b, MMS86, Oni85, OF76, Pal80, PF97, Pet76, RB75, RS87, Sch76a, SFIK80, SS94, SJKL94, TAJ81, Van82, Van86, Wai73a, WW91, Wil84a, Woo84, WW96, WH97, YSM95, Yan91, You81, ZB74, All83a, ADDM84, BDSV99, Bor83, CMS07, CL82, Cor84, CIM22, CIM24, DIS99, EP05, Fel79, Fer13, FS82, Fra06, Har84a, JAJB04, JWTG11, JL80, KNT^{+}01, KRR19, Lan74b, LF82, LHGM15]. programs [LW14, LPA13, Mal80, MK01, MJ99, NWE99, NLA15, Pet01, Pet77, Pil75, SJ79, SM18, SW12, SSK^{+}17, aSZP^{+}16, Wen80, YLR23, Edw77, Whi87, Bar73c, Nee77a]. progress [LCY07, Lav77]. progressive [TBF^{+}22]. Project [Kat71, MCG^{+}88, QC83, RM91, Sno78b, AKIM24, Als22, Eba20, JPM23, NDKL^{+}24, ZQWL24, Wal86a]. projection [CGH^{+}04, KSSA23, TWS^{+}22]. projects [AJ04, Bar78d, Bar82c, BBB^{+}22, DHA11, KJB11, KVG19, PdSCJM22, SBS20, XWHX21, vGPB10]. PROLOG [Bai85c, BA98, BS90b, CRR94, Clo85, Col88, Coo04, DT96, De 96, Deb88, Deb93, ELRV93, FD92, Knu92b, Kou87, LMN91, LC86, LQ93, Mat94b, Pas87, RC10, Rue93, SW90, TCC^{+}94, Vau89, Wis93]. promotion [PA01]. prompting [Gai82b]. prone [Lin98b]. proneness [WHS^{+}00]. Proof [MJ83]. proofreading [Mił10]. Propagation [GHM96]. Propagations [FZ98]. Properties [AB95, FZ98, Sch72, CCQ16]. Property [ZLWG11, AKS06, WG04]. proportion [Bis80]. proposal [RCA^{+}19]. Proposals [KRTW81, SSS22]. Proposed [Sch89b]. PROSIT [FPAF18]. prospects [Ano23m, SGA21]. PROTEAN [Lai95]. protect [FLL^{+}22]. protected [Le 88]. protecting [TXHL18]. Protection [Har84b, AGG06, JZ02, MV16, XZX^{+}21, YWT^{+}12, ZYYC12]. Protective [JI80, SHC74]. protein [DDP07]. ProTest [SW90]. proto [CPZ02, OM16]. proto-frameworks [CPZ02]. proto-pattern [OM16]. PROTOB [BBC91]. Protocol [AP91, Bor86, CG96, CDV88, DD90, EP79, Fri92, GM85b, GR91, HA90, Hol88, Hol93, HL98, Jia97, JB84, Lai95, LL96, LQ96, PHS84, Ste98, BGP17, CLC09, HL02b, JEG99, JTG^{+}11, Ker17, LBP^{+}13, LC05, dSMH13, NN24, SSM11, SR02, Sno91, SSK^{+}17, WMSY12, WMJ04, LFGCGCRP14, RMMLSME14, SW86b]. protocol-finding [LBP^{+}13]. Protocols [CW94, CLZ98, HMPT89, VSC93, GRR06, KD13, RSLAGCLB16, Vel88, CO88]. Prototype [Fri92, GR95, Ham95, Kuh90, LHS^{+}95, Tse97, Liu01, LS16, LLJ^{+}23, MST13]. Prototypes [BK86b, KRK21]. Prototyping [BBC91, ÖS96, RS94, VSC93, Zel80, BFG^{+}11, BCRF22, CPMAH^{+}20, FBLS12, Geh83, LHK99, SMLS23, TL14, ZC03]. Prototyping’10 [KH12]. Provably [DMR^{+}22]. provenance [WSL^{+}20, dAHCdAC18]. provide [BFPAGS^{+}08, CEF02, PALNGD^{+}06]. Provided [GM73, Oli83]. provider [BGS20, GAH05, MA20b]. provider-centric [BGS20]. providers [LBGA^{+}21, SM20]. Providing [BLC19, BS90c, MP00, SY86, VGF21, OW16]. provisioning [CRB^{+}11, FDN^{+}18, GdCF^{+}18, KGAR18, KGAR19, LWW^{+}23, MGT20, SGA20, VBS22, XYW^{+}24]. Proxies [Not90, HJC05]. Proximate [HC20, HM18]. proxy [BH01, BS99b, CLZ99, CZ04, HM18]. proxy-based [CLZ99]. PRTDS [WB85a]. PRV [GDW^{+}20]. PS [KA87]. PS-Algol [KA87]. PSAMS [HL02b]. pSeries [JDBP04]. pseudo [CJ73]. pseudo-pro [CJ73]. pseudorandom [SV22]. PSO [CS18]. PSSCC [DMR^{+}22]. pSystem [LS97]. PT [MS83]. PTDOM [WK06a]. Public [SY79, CMF^{+}17, FZS^{+}17, TWS^{+}22, WHH21]. publication [Thi03b, Bis79d]. Publications [Bow88, Rop88b, Rop88a]. publish [RC10]. publish/subscribe [RC10]. Publisher [Ano95n, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j]. Publishers [Fin77]. Publishing [Ald72, Cou85b, Flo74, Mul76, Sim83, Val78, Wal81a, Wil84b, BCLF^{+}07, Bry77, CDFV12, Mal80]. pull [ASME24, SdLJMP21]. Pulsar [Fin97]. PULSE [TKWW85]. Pun [Wit77a]. Pun-Dora [Wit77a]. Pure [BY90, CS91a, HGK^{+}19]. Purpose [FL75a, Haz74, LF74, LTV96, RTL^{+}91, WCE^{+}72, AYdS^{+}06, ABA20, BK77, CIM22, CIM24, DPDA14, JSC^{+}10, KNT^{+}01, KD83, Lew83, Mac79, MLR19, MK03, PM18]. Purposes [Gob71]. puzzles [GK08]. PyScribe [GLL^{+}24]. Pythia [PMY97]. Pythia/WK [PMY97]. Python [GLL^{+}24, MP19, NZPWR22, OMGDG14, RaMB^{+}24, Ren23, SSS22, ZXX24].

Q [KSBL22, MA20a]. QD [Deb93]. QD-Janus [Deb93]. Q’Nial [Jen89]. QoE [MSWC23]. QoS [CDRV03, CMT17, DGRB15, HHRS03, HKC^{+}12, KSDG24, ZXW^{+}17]. QoS-aware [DGRB15, KSDG24]. QoS-awareness [ZXW^{+}17]. QoS-based [CMT17, HKC^{+}12]. QS [Dew84]. QSIC [WCK11]. QTcl [CDRV03]. Qualitative [RJGH06, PPR02]. Quality [BP90, KSKG12, LY92, Poo88, Pra96a, Pra96b, TP92, WCK11, AMM10, CPZ02, CDFV12, DW13, DS12, FIÁSLSAR05, Hoa72, KRB21, LHB18, LGP^{+}11, MAR^{+}16, MS99, NZH20, NM06, PCML09, SLRS06, TMS18, TGC15, YZW^{+}12, MSWC23]. quality-driven [AMM10]. quality-of-service [LHB18, NZH20, SLRS06]. quantification [ZSH^{+}23]. Quantifying [SB03]. quantile [DHWZ14]. Quantitative [HK84b, AL21, GZX^{+}21, RJGH06]. quantization [ZZJ21]. Quantum [GKS^{+}22, MGJ^{+}24, BHT^{+}23, HPP22, KSST22, NUB24]. QuantumPath [HPP22]. quart [Coo85]. Quartiles [CMR92]. Quasi [KS80]. Quasi-parallel [KS80]. Quasiparallel [KPH76, PR77]. queens [Plu74]. Queries [dV89, BRTT09, KG18]. Query [HYZ^{+}18, KKS88, PSR83, SS93, SRRFGC^{+}10, ANSK16, HYC19, OEA05, PSTV10, PP16]. Querying [SS93]. quest [CC13]. Queueing [LF74, SM79]. Queues [Per85, TK72b, MLR19]. Queuing [HM84, CLCC15]. Quick [DW89, DL23, LS76, NHP81]. quicksort [McI99, WBGS22, Dro84, Mot81]. Quill [Wol91].

R [Bar73c, Bar74d, Bar75a, Bar75f, Bar76d, Bar77e, Bar78d, Bis79a, Bra75, Bux78, Col77b, Cou85a, Ear76, Fin77, For72, Han77a, Her84, Hol77, Hun72, Jac71, Jac84, Lav77, MDB19, Val78, Val80, Vel88, Wal81a, Wri98, BSC^{+}05, FM78, Rei84]. R-F [MDB19]. R-What [BSC^{+}05]. R2D2 [GDRV20]. rabbits [SJP^{+}09]. race [ZML22]. races [XXZ13]. Radar [KK97]. RADF [ZTC24]. radio [KCH08]. radio-astronomical [KCH08]. Radix [Sam71a, KEL^{+}21]. Rae [XZ01, XZ03]. Rafting [PV22]. RAID [BV06, Pla97]. RAID-like [Pla97]. Raiding [BV06]. railway [CL19]. rain [JP22]. rainbow [KSH^{+}15, WCE^{+}72]. Raj [Hor14]. Ralph [Lar75a]. RAM [HC16]. Random [Lar09, Plu74, CSM^{+}16, FP82, HCG^{+}16, LXY^{+}11, LKC12, MSC23, SDB^{+}22, SB21, TRRK21]. Randomly [SW87]. Range [Wel78b, KG18, TVSG21]. Ranked [Kir07]. Ranking [Fen98, FRBRF19, FFRF19, BZM^{+}17, MS22, MV20, SCD^{+}23]. RAP [NP98]. Rapid [Bel74, KH12, Sew82, SR02, Smi89, VSC93, Zel80, BFG^{+}11, FBLS12, FBH^{+}23, HIRB24]. Rapidly [Sav07]. Raspberry [MVOD19]. raster [SCPL23]. Rate [WJ93, Ker17, KCH07]. rates [Phi99]. RATFOR [Com78, Ker75]. ratio [Li22]. Rational [Hor78, Ker75]. Rationale [WBK91]. RATMAC [MS80b]. RATs [CWD08]. RATSNO [Han77b]. Raynal [CO88]. RBE [CC97]. RCS [Tic85]. RDB2RDF [MSB^{+}13]. RDF [AHH15, MSB^{+}13]. Re [Bro72, Bro77, JKW74, TDH97]. Re-creation [Bro72, Bro77]. Re-use [JKW74]. Reachability [Hol88, HC93, Wat04, dMFÁE17]. reachability-based [Wat04]. Reactive [Bou91, BS98, MM97, RMC97]. Reactor [WPL^{+}21]. Readability [PCBE96, CQH^{+}13]. readable [MCCJ23]. reader [LYL^{+}03]. readiness [HNJ^{+}23]. Reading [Bar76e, Ear77, Llo82]. reads [Boy01]. ready [DRH^{+}23]. ready-to-use [DRH^{+}23]. Real [ABRW94, Buh93, BL83, BW95, CS91a, CC84, Des92, DR92, Fra75, Gla82, Hal86, Heh76, HHL84, Jor90, KLLK98, LY92, LHC97, LF90, MA00, Nil88, Orm77, PJ75, QSA88, RS94, RA87, Ric76, REMC81, SF85, TH86, WC87, WR22, Wit83, AIB02, BVGVEA11, BVGVEA13, BSDF20, Bud85, BDM16, CBB20, CY01b, DHS01, DSD^{+}05, DHWZ14, DKM11, EKM^{+}99, FDN^{+}18, FPAF18, GKBK16, HK84a, HLFS05, JGB15, JLBR22, Kil19, KQZ^{+}11, LLK04, LCGS17, MvSdL09, Obe11, PLL^{+}02, PPA20, Pur76, RBS14, SLRS06, SM85, SJP^{+}09, SRCP19, SAA^{+}20, TRO17, VvK99, VC02, WM20, Wan82, SSP11]. Real-Time [Fra75, Hal86, HHL84, PJ75, RS94, SF85, TH86, Wit83, ABRW94, Buh93, BL83, BW95, CS91a, CC84, DR92, Gla82, Heh76, Jor90, KLLK98, LY92, LHC97, LF90, MA00, Orm77, RA87, Ric76, REMC81, WC87, AIB02, BVGVEA11, BVGVEA13, BSDF20, Bud85, BDM16, CBB20, DHWZ14, DKM11, EKM^{+}99, FDN^{+}18, FPAF18, HK84a, HLFS05, JLBR22, Kil19, KQZ^{+}11, LLK04, LCGS17, Obe11, PLL^{+}02, PPA20, Pur76, RBS14, SLRS06, SM85, SJP^{+}09, SAA^{+}20, TRO17, VvK99, VC02, WM20, Wan82, SSP11]. real-valued [GKBK16]. real-world [DSD^{+}05]. Realising [FL94]. realistic [BR01a, KSBW18]. reality [NMH^{+}24, WIYC20]. Realization [HS83, HTJNL19, Pap79, SvGB05]. Realizing [TS02, GHC^{+}07, WAH^{+}12]. Reallocation [BS90a]. Really [BS74, Bar74g, CVGCGCFF23, Str77, Yuv77b]. Realtime [Har80a]. Rearrangement [AS97b, VC90, KFMF18]. reasoner [BHNDdMRG23]. reasoning [BLR^{+}17, YHGC20, ZB23]. reasons [Kul74]. Recall [Thi80]. Recently [Inn77]. Recognising [Bis19]. recognition [Ano23m, BB03, DE16, DNL^{+}20, LCT^{+}21, LD99, MGP03, SGA21, SNK21, TBF^{+}22, WC04, XMTL21, ZZC^{+}17, vdMF13]. Recognizing [BHZ85, SJK^{+}21]. recommendation [Ber20, DWL^{+}17, FKL^{+}13, GLD^{+}21, HL20, JXG^{+}21, KKPP20, MF18, RRK^{+}18, TVCB15]. recommendations [BFPAGS^{+}08, LJS20]. recommender [CMTCC^{+}17, SBS20]. Recommending [LYLY20]. Recompilation [OW89, EP05, Rai84]. reconciliation [dAHCdAC18]. reconfigurable [HR06, KCK^{+}23, LC05, SMKZ06, SMR^{+}12, ZCN06, AE06b, AE06a]. reconfiguration [BGP17, JDBP04, ZPGHIA18]. Reconsidered [Rey90, dR86]. reconstruct [TKF09]. Reconstructing [Lam20]. reconstruction [RAP21, SD18, SROV06]. Record [HKW77, Vau89, GCRD04]. record/replay [GCRD04]. recorded [GMGDMB19]. Records [Bul87, Cow87, Rea73, OHD22, ŠS08, Sur13, Ald72]. Recoverability [Jeg83]. recoverable [KMB98]. Recovering [DD18]. Recovery [HJ88b, PG81, Shr78, Shr79b, Shr79a, SMM^{+}84, Sti85, Vau79, CHCC07, FZS^{+}17, LV01, MvSdL09, Pem80, PCdGPP12, SDDD10, STA09, TKT^{+}07, YZYL07]. recovery-oriented [MvSdL09]. Recreating [CH73]. recreation [GXN10]. recreations [Ano71d, Ano71a, Ano71b, Ano71c, Ano72a, Ano72b, BM72, Foo72, Tan73]. Rectangles [Coo83]. recurrent [MA20a]. Recursion [CDH77, Gol81a, Roh81]. Recursive [AI80, Han85, Kil81, Kos90, Roh77b, Ste80, YL95, vR92, CDH77, Glü12, RK15b, Set79, Wen80]. Recursive-descent [Han85]. Redesigning [CV98]. Redisplay [Dan90]. redocumentation [GMP^{+}21]. REDOM [TDH97]. reduce [KRB21, Kra10]. Reducing [BS93, KGSC01, Kur99, Ono93b, TS91, Wat04, ZLZ^{+}19, BDSV99, MK18, WAML12, LFHL22]. Reduction [HV88, Har91, CJTK22, LC07, LLW^{+}23, OJP99, SSRAH15, SH82, WJC^{+}14]. redundancy [HNW^{+}01, KKN04, VH04]. redundant [SAEMM21]. Reed [Pla97, PD05]. reengineering [MRBB19]. reentrant [DD10]. Reeves [Eve73]. refactor [CA18]. refactoring [CSD23, LBC^{+}11, MF08, RRK^{+}18, SMNB21]. refactorings [CS22, RMZ17]. referees [Bis19]. Reference [Bae73, Chr84, MS96, Bar73d, KBPM^{+}20, Mha05, MGS^{+}20, RN00, STT24, TM14]. References [AS88]. referential [All89]. Refined [SW90, CQH^{+}13]. Refinement [Dro85b, Mor80]. Reflection [KMS98, LMN91, ZLTX18, MVT^{+}09]. Reflections [BTS09]. reflective [CV03]. Reformulation [Mar22]. refresh [KCH07]. refresh-rate [KCH07]. region [YC16]. region-based [YC16]. Regions [Rey90, XCG06]. Register [Bak72, BS90a, Bur16, DF84, DW91, FH92b, GW96, NP98, VSM87, AS87, CW08, Ham81, SS03, SSRAH15, WJC^{+}14, Yuv77b]. register-pressure-reduction [SSRAH15]. registers [Yuv77c]. registry [BCC^{+}22]. Regression [Gom78, PM17, AA19, JTG^{+}11]. regret [KPK^{+}18]. Regular [IIL17, Kea91a, Ric79, Chi17, KS08, SCF^{+}17]. regular-expression [SCF^{+}17]. Regulation [Bur98, KP90]. Regulations [TDH97]. Rehabilitation [BDD09]. Reidel [Sim83]. reinforcement [Che22, GWX^{+}23]. Reitman [Dav74]. rejuvenation [DAP21]. related [CR18, Deo10, HHMMG12, JH03, KH18, CCK21, PRP^{+}23, RGSGHGCG21, ZCCB22]. Relational [Bul87, FKV98, Hut79a, Hut79b, MRNL92, MV86, MXYQ86, PSR83, Sil92, TS81, Wes83, FSC08, Liu03, MSB^{+}13, SBS13]. relational-XML [FSC08]. Relationlog [Liu01]. relations [Al 13, LLZ20, aSJW^{+}24]. relationship [LQS^{+}24, MLC02, PP84, Pit82]. Relationships [MS98, MKD98, Sha78, KAZ13]. Relative [ACKS09, SA97, SNM80, JZLP20, LZK^{+}23]. relay [LZL^{+}17]. release [IS05, KRB21, vdHW03]. Relevance [Sri24]. Reliability [MKC20, TV96, And78, SGDA18, WCsH16]. Reliable [AS78, Any85, Bar78a, ESB^{+}17, Jia97, MM81, YZYL07, Bir99, CWC^{+}21, IBA^{+}21, KQZ^{+}11, RT78]. ReliaCloud [SGDA18]. ReliaCloud-NS [SGDA18]. Relocatable [MT78]. remainder [LKK19]. Remapping [TA91]. Remarks [Hay80, Ano80b]. Remote [ACG78, Ans86, AV05, CC84, CS97, FZ98, GKC87, LT91, WMG94, WP96, CYJ^{+}22, Fra99, GCARPC^{+}01, HC99, IH01, MCGS08, MKW^{+}22, Rin07, TLB^{+}18, ZWML14, ZLY18, CWD08, Sto94, TN98]. removal [Ber82, UFR18]. rendering [NHTT08, WW09]. Rendezvous [GKC87, GST92]. renewable [SSK^{+}22]. Renumber [Law78]. Reorganization [WPN86, KKdC^{+}23]. repair [BdPGS14]. repairing [TVCB15]. repartitioning [HB18]. repartitioning-based [HB18]. repeatable [AK15, CGR00]. Repeated [JGR89]. Repeating [Poo71a]. Replacement [CK94, Inn77, MN80, CS03, Kra10, MM88, dRRGdC15, RGV14, XP23]. Replay [KM13, GCRD04]. replica [FB23, TGAS22, ZXW^{+}17]. Replicated [Bre86, CLLT98, PGH^{+}98, IH01, MAJ15]. Replication [PSA87, BM03, Fra99, LGZ^{+}08, MMHB08, PJS^{+}23, SXWL17]. Reply [Gen81]. replying [LKCW13]. Report [Bar76b, Bul72b, DFRR15, GKBK16, Han89a, Hut76, Kot96, KHMB17, LD87, MNW14, SAL16, Sur13, Wil76, ABG^{+}22, CL09, CSS15, DBvG^{+}23, FFRF19, GS08, JPM23, PH14, Rog73, SM15, Wil74a, ZCO13, Bar79b]. Reporting [Mau82, PG81, Hut74]. Reports [DF15, OM16, OW16, Pet76, RMZ17, WBB15, vdWCB17]. repositories [BTZ07, CR18, HPZ^{+}20, RBR21]. repository [BHW05, HC10, LCZ08, QL13]. Representation [Bis84, DCW93, Fre78a, Fre78b, HHK90, Lic77, RS93b, Bar74f, Dod82, Mad82, RJ09, WHH21]. Representations [GF84, MFH10, PMC22]. represented [SCPL23]. Representing [JKB04, LK93, Wil84a]. Reproducibility [WBA^{+}23]. Reproducible [Han78c, HL79, TBT23]. reprogramming [OMGDG14]. Request [KNC94, LCW07]. requests [ASME24, SdLJMP21, ZDY^{+}17]. requirement [Kur99, ZZLL24]. Requirements [BS93, GdCF^{+}18, KN88, LLJ^{+}23, Lor91, MPN^{+}95, Nut76, WKS^{+}98, BAPC23, DHGR92, DS12, GN02, JSRM18, KAS^{+}16, KYF^{+}22, KPJ^{+}17, LPP09, LS16, LZLL22, MST13, Rop88a, RBD^{+}24, Ste79, SGCM11, Wat04, YZW^{+}12, ZWAL22, ZZLL24, ZZL24]. Requiring [Ric76]. reranking [YAVHC21]. RES [SSK^{+}22]. rescue [CFD^{+}22, DGM^{+}22, LW04]. Research [CJTK22, Cra77, MBO97, SFB13, VS88, WPL^{+}21, BMY06, CFL^{+}98, CKCG23, Che22, CCM05, GH19, HYY23, HP04, JPM23, LZ10, MFB^{+}02, PPR02, SS21, SB22, Sri24, TLC^{+}18, ZKB^{+}22b, BT21a, Dav78]. Researchers [MBO97]. Researching [CCM05]. RESeED [SCF^{+}17]. Reserved [Hun81, Sal79d]. reservoir [Kir07, ZWX^{+}21]. resident [Poh81]. residential [VRC^{+}06]. residual [JFZ^{+}21]. resiliency [CCLN22]. Resistant [AM86b, Wal83a]. resolution [Bra99]. Resolving [LD14, Sit79]. resonance [VP05]. Resource [ALBN81, BR97, GdCF^{+}18, Gom74, HJ14, KJVS21, LWW^{+}23, Nut76, PU84, Rei72, SWA^{+}75, TDH97, ZDY^{+}17, ADZF21, ASEB09, AA23, CRB^{+}11, CHS^{+}05, DAC^{+}21, FDN^{+}18, GCZ^{+}24, GDGB17, HSRB24, HYH15, KJB11, KGAR18, KGAR19, KBM02, KSDG24, KSST22, MVOD19, MGT20, MKD^{+}22, NRUP21, NEP^{+}17, PKK12, RMM19, ROFGFRM16, SGA20, SGWVP15, SC22, SOKE23, SWBS17, VNGB08, VVG^{+}24, XYW^{+}24, YB06, ZXT^{+}17, ZSL21, ZB18, dOED^{+}20, dlRNGCC23]. resource-aware [PKK12]. resource-constrained [KSST22, SWBS17, dlRNGCC23]. Resource-Oriented [Rei72]. resource-restricted [MKD^{+}22]. Resources [PH84, BIA24, VS20]. Response [CKB01, CKB03, HBC15, ZGGL23]. responses [ZGL^{+}23]. Responsive [Str83b]. responsiveness [CALL18]. rest [Ano71e, BMC17]. restart [CTLL07]. RESTful [FLSCC15, dSMH13]. Restoration [MG94, CS02]. Restore [Bak72]. Restores [Dri93]. Restoring [DW91]. Restricted [Har92, TA91, MKD^{+}22]. Restrictions [McK90]. Restructuring [Har83, Hop96, Kob77, Zim90, Lam20]. Results [BLC19, Lec95, MW93, RG89, CJTK22]. resurrecting [CBC00]. Retail [Ban71]. retailing [MDB19]. Retargetability [CDGP93]. Retargetable [ABSS98, FH92b, GF84, FHL^{+}18]. Retargeting [Ard87, LC12]. Retract [Col88]. Retraction [Ano23m]. Retrieval [CCC96, FFD96, TS81, ZM95, ABA20, CI03, GRS74, GJ00, GSS^{+}20, KEL^{+}21, LTL^{+}03, MRBB19, Mos06, SI10]. Retrospect [Wil73]. Retrospective [KFJS88, Mal83, JLvR^{+}02, Mal11, RW12, ZL84]. Retry [CAFH94]. Return [Str81]. Returns [Er83]. Reusability [JR92, MCLL21, PW97, Wie96, MC23]. Reusable [ABBE98, FFD96, KW09, PW93, HC10, PM12, SA02, Vo00]. Reuse [CCC96, LCW98, PA91, AKM17, BGM17, CCF^{+}09, CS17, DSD^{+}05, JLZ09, Kim02, KSRR17, LKCC00, MW13, RGN^{+}14, RN00, SB21, STH^{+}18, TL14, VC02, vGPB10]. reuseability [KKLL99]. Reusing [ÁSARSG09, DCS23, KV17]. Reverse [Bro72, Bro77, Byr91, CH73, Cd91, HC93, TAFC00, NZL19, SKM01, TKF09, WBB15]. Reversible [Bri87, SWBS17]. Review [Ald72, And78, Ano73a, Ano79a, Ano87a, Ano88c, Ano88b, Ano88a, Ano23m, Atk78, Atk79a, Atk79b, Atk82b, Atk83, Bar71, Bar72c, Bar72a, Bar72b, Bar73e, Bar73c, Bar73b, Bar73a, Bar73d, Bar74e, Bar74d, Bar74f, Bar74c, Bar75a, Bar75c, Bar75e, Bar75f, Bar75d, Bar75b, Bar76a, Bar76d, Bar76b, Bar76c, Bar76e, Bar77e, Bar77d, Bar77b, Bar77c, Bar78c, Bar78b, Bar78d, Bar79b, Bar79a, Bar80d, Bar80e, Bar81, Bar82b, Bar82a, Bar82c, Bar83a, Bar84b, Bar84a, Bis79a, Bis79b, Bis81a, Bis81b, Bis82, Bis84, Bis86, Bow88, Bra75, Bra80, Bri82, Bry77, Bul72a, Bul72b, Bul73, Bux78, Cam85, CO88, Cav83a, Cla98, Col77b, Con77, Cor82, Cor99a, Cou84a, Cou84b, Cou85a, Cou85b, Dav74, Dav78, Dea86, Ear77, Edm82, Edm86, Edw77, Edw98a, Edw98b, Ell72, Eme84]. Review [Eve73, Fen98, Fin77, Flo73, Flo74, Flo79, For72, Fox79, Gar86, Gru83, Han72, Han78a, Han78b, Han77a, Haz71, Haz72, Her84, Hop73, Hop74, HW77, How76, Hun72, Hut74, Hut76, Inc86, Jac71, Jac84, Jon74, Ken77, Lan74a, Lan75, Lar71, Lar75a, Lav77, Lav78, LPT82, Liv75, Llo82, Lon88, Mad82, Mar88, McD71, Mee87, Mer74, Mil72, Mul76, Nee77a, NPW72, Nic72, Nic98, Pet77, Pit82, Pra96a, Pra96b, Ree78, Ree82, RB82, Ree84b, Ree84a, Ree73, Ree75, Ree76, Rob72, Rob81, Rob82b, Rob82a, Rog71, Rog73, Rog74, Roh77a, Rop88b, Rop88a, Ros74, Sch76a, Sha72, Sha83, Sim83, Sto88, Tho77, Tho74, Val76b, Val76a, Val77a, Val77b, Val78, Val79, Val80, Vel88, Wal83b, Wal81a, Wal82, Wal83c, Wal84b, Wal86b, Wan82, Wel72, Whi87, Wic72a, Wil72]. Review [Wil74a, Wil76, Wil84b, Wil87, Wis74, Woo74, Wri98, ARA18, ASGA23, Bir99, BBB^{+}11, CZL21, DAC^{+}21, DPAG11, FAA^{+}22, FRBRF19, FBH^{+}23, GH19, GKS^{+}22, HSRB24, JJI22, JC19, MCLL21, MF18, MAW^{+}16, NHB23, NRS13, OFRW10, PVAHRG^{+}15, PRP^{+}23, Pol01, SPR^{+}19, SSGA20, SS19, ZGGL23, vdMF13, SGA21]. Reviewer [Ano17m]. Reviews [Ano09, Pet77, LLLY19]. revised [Bra80, Wil87]. Revisited [Han94a, Lun86, Wel83, Han99b, Ros75, Val00]. Revisiting [CPP12, SGD05]. revocation [NM19]. Revoking [CFL84]. Rewriting [LB94, Lin87, AGG06]. RFID [BBMG08, CPD13, CPCL10, DWL^{+}17]. RFID-based [BBMG08]. RFID-enabled [CPD13]. RGraph [CYJ^{+}22]. rich [ASEF23, RSLAGCLB16]. Richard [Ano87a, Rob81, Rog71]. Richards [Bar81]. Rides [WIYC20]. Rie [SIN95]. RIFLING [Che22]. right [KT01b]. Rinehart [Haz72]. Ring [KDP83, WC87, Bre82, Col82, MBV^{+}10, LB81, VSB86]. Ring-Star [KDP83]. RIOS [WBB15]. Ripple [ZLTX18]. RISC [Ise90]. Risk [JI21, MFdlP12, BGS^{+}13, JXG^{+}21, MC23, PPR^{+}21, SM20, SRS18]. River [SHGG16]. RJ [OW16]. RMI [ET07]. RMX86 [HK84a]. RMX86-PEARL [HK84a]. RNA [AS08]. RNFREE [Gra81]. Road [Cas92]. roadmap [ZPSH21]. Roaring [LKK^{+}18, CLKG16, LSYKK16]. Robert [Bar82a, Bar82c, Bar84b, Bar84a, How76, Mil72, Wil84b]. Roberto [Mul76]. Robin [Pra96a, Pra96b, Rob82b]. Robot [KS84, RMC97]. robotic [RBD^{+}24]. robotics [OMGDG14]. robots [LZWX22, PKK12]. Robust [Car82, KBB^{+}20, LZL^{+}17, MM81, NW85, BGP17, FGNZ00, KSSA23, POZ^{+}16, ZZJZ24]. robustness [CS04]. Rodin [BFJ^{+}11]. Roger [Bis81a, Cou84b]. Rohl [Bis81b, HW77]. Role [JDGCGA12, SE11, BSC^{+}05, DFPT09, FZ12, HvdH02, LB02, YHGC20]. role-based [BSC^{+}05, DFPT09, HvdH02]. role-binding [LB02]. Role-play [JDGCGA12]. Roles [Bis90, AJ04, ST04, SE11]. Roll [Bow73]. Roll-out-Roll-in [Bow73]. rollback [YZYL07]. Rollforward [MDP96]. ROME4EU [RMdL12]. Ronald [Mer74]. Root [ZPSH21]. rooted [BJL06]. Roper [Val76b]. Ropes [BAP95]. Rose [Bar71]. Rosenblatt [Lav77]. Rosenfeld [Lan74a]. Ross [Hor07a]. Rostering [CFL^{+}98, Mon96a]. Round [Mac96a, dRRGdC15]. round-based [dRRGdC15]. Round-Pound [Mac96a]. router [LLJ12]. routers [KCCV05, SBG^{+}05]. Routines [CLL91, GF81, Mid86, Oli83, Sch76b, FCG83]. routing [AK15, GZW^{+}22, KRZ02]. Row [MM88]. Row-replacement [MM88]. RPC [CCE99, JZ93, KLY20, LT91, SM01, Sto94, Yas94]. RPC-based [KLY20, Yas94]. RT [EKM^{+}99, dOdO16, EKM^{+}99, KHC^{+}19]. RT-Mach [EKM^{+}99]. RTAG [HA90]. RTD [DHS01]. RTL [Bar74b, Bar80b, Bar78b]. RTL/2 [Bar74b, Bar80b, Bar78b]. RTS [QSA88]. Rubinoff [Jon74, Wil72]. RUE [ADZF21]. RUGRAT [HCG^{+}16]. Rule [CC97, DW73, MB97, DE16, JC22, LLH14, MGG^{+}09, Mił10, ROFGFR^{+}16, ROFGFRM16, WSWR23]. Rule-based [MB97, DE16, LLH14, Mił10, ROFGFR^{+}16, ROFGFRM16]. Rule-by-example [CC97]. Rules [DF87, BRL^{+}15, SH82]. RuleSIM [ROFGFR^{+}16]. Run [BS74, CC77, Dan82, FM78, GWA91, Hol83, Joh79, KW90, Kow81, Ste92, WB85a, WB85b, Yuv75, CMT17, CC01, FFRFS19, LF82, Str77]. Run-Time [WB85a, Yuv75, BS74, CC77, FM78, GWA91, Hol83, Joh79, KW90, Ste92, WB85b, CMT17, CC01, FFRFS19, LF82, Str77]. Runabout [Gro08]. Running [AK15, BS90c, Har80a, HJ88b, LNhCW16, SJ79]. Runtime [DDD16, FZS^{+}17, HMS^{+}95, AGC10, AE14, AGG06, LMK16, PKC^{+}13, SMKZ06, SFC^{+}21, SD18, SB13, Söz15]. Rustin [Bar74d].

S [Ano79a, Ano87a, Bar73c, Bar74e, Bar74f, Bar75a, Bar77e, Bar82a, Bar82c, Bar84a, Bis81b, Bis84, Bra80, Ell72, HW77, HM84, Hun72, Jac84, Lav77, Mad82, Ree73, Rob82a, Sau88, Val76b, Wan82, BB75, Hal82, MSR^{+}07]. S/130 [Hal82]. S/370 [BB75]. SaaS [CS18, WY18b]. SaaT [TC19]. Saczalski [Liv75]. Safe [GvRN^{+}11, HFPB98, Kur81, Nar94, AIB02, NSM16, Win02]. SafeMan [GMC^{+}21]. SafeOSL [YHK^{+}22]. Safety [MMS90, GEI^{+}11, GMC^{+}21, HHMMG12, KH18, MPJ20, RLPA18, SB13, WWGP10, WYAZ15, YHK^{+}22, ZRX^{+}99]. Safety-critical [MMS90, GEI^{+}11, MPJ20, ZRX^{+}99]. safety-oriented [WYAZ15]. safety-related [HHMMG12, KH18]. SafeType [IASC16]. SAHAYOG [DTJ89]. Said [Ano23m]. Salford [Bai85c]. SALOON [QRD16]. sam [Pik87]. Sample [AKDN90]. Sampled [GR17]. samples [GLS^{+}23]. Sampling [JI21, Wai73a, Bin06, Kir07, SCD^{+}23]. sandboxing [GCF15]. Sanderson [Rog74]. Sangrah [PG81]. SASL [HV88, Jon85, LT90]. Satellite [BS80, FL75b, SDF^{+}21]. SATHE [AvdSGS80]. satisfaction [ASA^{+}21]. Satisfy [PH84]. Satisfying [YZW^{+}12]. Saturation [MY87]. Save [Bak72, FH91b]. Save/Restore [Bak72]. Saving [DW91, JI21, ÜY22]. SCADA [BAM^{+}20]. SCADA-based [BAM^{+}20]. Scalability [LKL95, ZSFY05, HBB22, HL02a, LGZ^{+}08, MZ00]. Scalable [AMW91, BHNDdMRG23, BCPSC18, hPmKgH15, Ryu16, WN06, dAKdGJ11, AML20, AL21, BJ21, BBMN18, BCL13, BCSV04, BSM^{+}21, Buy00, EMRK20, GDBP23, GDRV20, HOY17, Nic08, PT14, PLR18, SGDA18, SSO13, TDDE15, ZLZ^{+}23, ZWF^{+}22]. Scalar [CK94, CS03]. Scalars [Atk79c]. Scale [DLP85, HWS^{+}88, AKL^{+}09, AZS19, CGM^{+}03, CRC18, Deu99, FMNW04, GLL20, HB18, HGK^{+}19, KEL^{+}21, LC07, Mos06, PK11, RBRB23, SATV22, WHS^{+}00, XP23, XYW^{+}24, ZZKA17, ZGL^{+}23]. scale-based [LC07]. Scaled [Ric76]. scaler [LS84]. Scaling [JDJ^{+}06, JPM23, KCH07, MGT20]. scam [MGBO22]. scan [PP16, SS03]. Scandinavian [Mad95]. Scanner [DGM80, FHS92, SN90, HL87]. scanners [ACKT20, JKB04]. scanning [AKW79].

SCARS [RBD^{+}24]. Scatter [LV73].

scattering [WPL^{+}21]. scenario [MGS^{+}20]. Scenarios [HMN11, TL98, LKC12, SESK23, Sin81]. Scenarios-based [HMN11]. Schaeffer [Liv75]. SchedSP [GAH05]. Schedulability [Ker17]. Schedule [LT85, DHA11]. Scheduler [ABSS98, SRS98, Che22, EGCCM21, TCM07]. Schedulers [Gra96, Rog23, SFC^{+}21]. Scheduling [BMAV05, CA14, DF95, Hal86, Han76a, KA22, Lar75b, Lar78, LHC97, RGV14, Sch78, Shr76, WBV96, AA23, BS19, BGSG20, BDA20, BM01, CBB20, CLCC15, CW08, DSD^{+}19, FCYL18, GHM^{+}06, GAH05, GF78, HB18, HYH15, IK15, JSPP21, KSBL22, Ker17, KMB^{+}21, KTG20, KSDG24, Lan71, LBC^{+}11, LSAF16, LLWB14, MUE23, MAR^{+}16, NS08, RR05, ROFGFRM16, SGWVP15, SAL^{+}04, SA20, TLC^{+}18, VS20, VMY24, WJC^{+}14, WTF^{+}22, ZWML14, ZB18, TDH97]. Schema [Mat83a, BMC17, MNEM21, PSRCC02, WK06a]. schema-aware [WK06a]. schemas [DDPP02, GP01, LMPR07]. schemata [GRVA09]. Schematic [TY80]. Scheme [AC80a, Bar80c, Bec91, CW91, HJ88b, NHP81, RS21, Ano76h, FLL^{+}22, FDN^{+}18, Gu05, GLKZ21, HC13, HM18, HC20, HTWS15, KQZ^{+}11, LZL^{+}17, NSM^{+}22, NM19, OT02, Pem80, SSKG22, Ste02, ZZT^{+}21, ZWF^{+}22, BVB^{+}12]. Schemes [Mös88, Wal81b]. Schmidt [Sim83]. Schneider [Ano79a]. Schoeffler [Mer74]. School [Wil80]. Schooner [CS97]. Schrödinger [AH01]. Science [Bis79b, Cou85a, Val76b, Woo84, ALKL19, Bar73b, Cav83a, CC13, JL80, JL81, MSK^{+}24, Val76a, GdCF^{+}18, Gru83]. sciences [Rob72, Ken77]. Scientific [Bow88, BSRS85, DRL82, KD83, Lew83, Mar86, MS80b, O’N88, Rop88a, VP05, BBB^{+}22, CS03, DRH^{+}23, FRGPLF^{+}12, KSBL22, LWJ^{+}21, MM02, Pet01, RMZ17, SFK^{+}01, SSS^{+}02, dOED^{+}20]. Scientists [BSC^{+}05, Ell72, Cou85b]. Scope [Sal81a, STS83]. Scopira [DP09]. Scores [Fox87, Hoa73]. SCORM [HC10]. SCOZ [AKNJ21]. Scrabble [Gor94]. scratch [YYSG11]. scratch-pad [YYSG11]. Screen [Ell82a, HH82, RS90, Car81, LYL^{+}03, NMH^{+}24]. Screening [SJKL94, AKL^{+}09]. scriptable [LBP^{+}13]. scripted [PLFIVRAR23]. Scripting [KV98, DM07, Ric00, Sto05, Yi12]. Scripts [Fra80, ÜY22]. SCRUB [Law78]. SCRUB-Systematically [Law78]. scrum [JPM23, BBB^{+}22, JSKB22]. SDASs [LCGS17]. SDFS [LWZ^{+}19]. SDFunc [Tur22]. SDK4ED [ACAB22]. SDL [BFGS05]. SDN [FVF^{+}18, KDA20]. SDN/NFV [KDA20]. SDSN [KHC^{+}19]. seamless [MN18, Mus17]. Seamlessly [BRTT09, ZHO^{+}19]. Search [AW93, AC75, And91, BP09, BG93, CS82, DS88, FP82, IC85, McG82, Mon96b, RS93a, Shn73, Smi91, ALG23, ACM^{+}15, ASTW03, BP11, CGZ^{+}20, CIM24, DDDF17, DS03, DHA11, FKL^{+}13, FG08, GK08, HPZ^{+}20, KHOY16, KH04, PSTV10, Plu74, Rai99, Rön07, SCF^{+}17, SCL00, SS23, TC19, dKM04, PSTV10]. Search-Based [BP09, BP11, DHA11]. Searches [HW94, Fen01a, KS08]. Searching [And91, BY89, BK93, CS82, Dav82, Hor80, HS91, LDI98, QK78, Rai92, Smi94, TT82, ALG23, ASTW03, Ayc15, MS22, Mha05, PT00a]. Seattle [Bar78d, Bar82a, Bar82c, Bar84b, Bar84a]. Second [Deo02, LG76, Mad82, PMG71, Wic77, Bar82c, Cam85, CIM24, Fox79, Ken77, LB15, Lem21, LM22, NMH^{+}24, NL24, Ree76, DFPT09]. Second-Order [PMG71]. second-screen [NMH^{+}24]. secondary [AS08]. Secretary [SS84]. Section [HW10b, RBB12, SFB13, Tse13, TGC15, HW10a]. sector [LW19]. Secure [BT21a, JW75, KHS^{+}20, ABBS22, BAM^{+}20, BAF03, BDLM04, BZM^{+}17, CH06, CNAM^{+}10, DDMD20, DMR^{+}22, DMC17, FO10, HNJ^{+}23, KSST22, LJ99, MS22, MKC11, MPJ20, NSM^{+}22, OHD22, PPSS05, SCF^{+}17, SAEGF11, TP03, VAP^{+}17]. secured [NM19]. securing [TMGR^{+}22]. Security [JLBR22, KT01a, Lv21, MR92, PF09, ZSH^{+}23, AL21, AB20, BCPL13, BGS^{+}13, BTS09, BS99b, CV03, CZ04, FLPM20, HJ08, KD13, KDA20, LWZ^{+}21, MDH^{+}13, MLC02, McN05, MVTH14, MGS^{+}20, OT02, PRP^{+}23, RdOTF14, SRS18, WBN^{+}20, XWC^{+}17, dAKdGJ11, CF05, Zam03]. Security-based [ZSH^{+}23]. security-related [PRP^{+}23]. Sedgewick [Wil84b]. Seen [KO86]. SEFT [dKM04]. Segmentation [Kaw79, GDRV20]. Segmented [BH82]. Segmenting [ZZLL24]. Segments [Sla86]. Seismic [HWS^{+}88]. Seismo [SDF^{+}21]. Seismo-Electromagnetic [SDF^{+}21]. selected [Flo73, Ano09]. Selecting [CMR92, DdB15, HBC15, MHB90, QRD16, RL14, ST14]. Selection [And89, Dro86, HS85, LNW82, Mus97, NS74, PK89, AMOS19, AHRR22b, FZS^{+}17, GKWS11, GH19, KSK15, MV20, MS18, MA20b, NHB23, NNK21, ST12, SS23, Val00, YLP^{+}11, Zdu07, ZB18]. selections [ST01]. Selective [AS78, CMES05, HOY17, MP18]. Self [All89, BG93, BDA20, CHY^{+}22, CK86, LPT78, LQL^{+}22, PDBG10, SAC06, WZH01, dODP21, CBR10, ESB^{+}17, FFF^{+}13, Gai82b, Glü12, GM22, HIR06, KGAR19, MA20a, MV16, NNK21, PLPA22, ST12, SBD15, SMT^{+}18, TDDE15, VGF21, APS95]. Self-adaptation [PDBG10]. self-adapting [HIR06, MA20a]. Self-adaptive [BDA20, dODP21, FFF^{+}13, GM22, PLPA22, ST12, VGF21]. Self-adjusting [BG93, WZH01]. Self-admitted [LQL^{+}22]. self-applicable [Glü12]. Self-compiling [LPT78]. Self-compressing [CHY^{+}22]. self-configurable [CBR10]. self-deployment [ESB^{+}17]. self-healing [SBD15, SMT^{+}18]. self-learning [KGAR19]. self-optimizing [NNK21]. self-prompting [Gai82b]. self-protection [MV16]. Self-referential [All89]. self-scalable [TDDE15]. Self-tuning [SAC06]. SELFNET [CPMAH^{+}20]. Semantic [FZ98, HG84, Inc84, KH07, KW92, Mös88, Sch89b, SW91, Wat86, BGA20, Ber20, BFRFB22, CD15, DFQ^{+}22, FLSCC15, GK08, KEL^{+}21, DCEG23, SJK^{+}21, WZLN08, dMFÁE17, BAJMT21, HL20]. Semantically [BS84, JPG^{+}17]. Semantically-based [BS84]. semantically-enabled [JPG^{+}17]. Semantics [ARV77, CFD^{+}22, GL78, Slo93, WB78, GMS20, Har99, HYC19, HLR20, LQL^{+}22, Lon88]. Semaphores [DF95, RM75]. Semi [CDV88, LV01, BDD09, GSR17, Hug82, PTU03, SATV22, ZHZ^{+}14]. semi-automated [SATV22]. Semi-automatic [CDV88, LV01, PTU03]. semi-incremental [Hug82]. semi-index [GSR17]. semi-splaying [BDD09]. semi-supervised [ZHZ^{+}14]. Semiblock [Kaw80]. Semigroups [Car97]. Seminumerical [Llo82]. sense [AHH15, Bis80]. Sensei [DDMD20]. sensing [ZWML14, ZZT^{+}21, ZLY18]. Sensitive [Rob83b, AP94, BGA20, BDM16, CALL18, EF13, KRR19, LMK16, LWW^{+}23, SM18, ST19, SYXZ14, VBS22, WC08, XWC^{+}17]. sensitivity [HOY17, PLR18]. sensor [ACV10, CDR13, EC13, HPK^{+}12, KAS^{+}14, LZWX22, MTPC14, SIC^{+}20, SDB^{+}22, WR22]. sensors [HSY^{+}20]. Sentence [CCRD^{+}80, MS83]. Senti [ASA^{+}21]. Senti-eSystem [ASA^{+}21]. sentiment [ASA^{+}21, GC20, Kil19, PPR^{+}21, SSV^{+}20]. sentiment-based [ASA^{+}21]. Sentinels [Thi89]. Sentry [CG95a]. Seok [XZ01, XZ03]. Separate [Fos86].

Separately [Han79b]. Separating [Rob84]. separation [wKJM18]. September [Val78]. Sequence [NW85, PP98, Sal79a, Str95, Vau79, ZLWG11, CHY^{+}22, Rya80]. Sequence-based [PP98]. Sequence-Controlled [NW85]. SequenceL [Coo96, CA00]. Sequences [MDP96, BLLP04, HYC19, LPF^{+}11, SNK21]. Sequencing [Mac77a]. Sequential [Ben77, Cow87, Deb93, Fid88, Gen81, HD86, RB81, Shr78, Wre88, FB22, Fin88, IS05, Jac71, LQS^{+}24]. serialization [BHK^{+}04]. serializer [DPDA14]. Serializing [MFH10]. Series [Bak72, Bis79b, Cou85a, EP79, Har80b, Iza80, McI90, SAC^{+}92, WQ72, SIC^{+}20, SDB^{+}22, SBFF24, Has77, Bar78c]. Serious [Lar73a, Lar73b]. Server [ARA18, AKDN90, BPY90, CGK89, Del82, HLLZ21, HM90, Ono93b, She81b, Sno91, AW04, Bas00, EYG22, GNSP12, GLT08, HC20, IH01, KS01b, Li22, LHFL07, NTF^{+}17, Rei99, RC10, SFK^{+}01, SJA^{+}04, SH17, ST19, WSL03, ZLZ^{+}21, CVV97, MNH04]. serverless [BBH^{+}22, DCS23, JGC^{+}21, RRFD23, WTF^{+}22, YSBL22, ZKW21]. Servers [CLZ98, JDJ^{+}06, McC90, YF91, CZ04, JDBP04, KSH11, Li18, SKI08]. Service [ASP^{+}19, BS19, GRVC23, HS77, HLR^{+}03, RHT^{+}13, ZLZ^{+}21, AGC10, AMM10, AKS06, AFNG20, ARK21, BELS14, BL15, Bla04, BGS20, BZM^{+}17, CWC^{+}21, CTLL07, CHCC07, CCR19, CF05, CNAM^{+}10, DB21a, DTB12, DGRB15, DMD^{+}06, GDBP23, GMS20, GARSR18, GLD^{+}21, GAH05, GSAE14, HK06b, HL02a, HKC^{+}12, HYT13, JPG^{+}17, Kar14, KRZ02, KMY^{+}05, KS20, LBGA^{+}21, LHB18, LLH14, LC07, LGP^{+}11, MKC20, MS18, MF18, MA20b, MOTG18, NZH20, Obe11, PCML09, PKK12, PLPA22, PL08, PDBG10, QTV23, RR22, RMSMML^{+}11, RMMLSME14, RMdL12, SPR^{+}19, SGAS21, SESK23, SLRS06, SMR^{+}12, SM20, SHB19, TDDE15, TWJ^{+}13, TLC^{+}18, WSYO11, WLTJ13, WMSY12, WBA^{+}23, WBB07, WKKK24, XXJS18, XZW^{+}22, XLLY19, ZTC24, dOED^{+}20]. service-based [AGC10, CF05]. Service-level [BS19]. Service-Oriented [RHT^{+}13, Bla04, LGP^{+}11, MOTG18, Obe11, PLPA22, PL08, RR22, RMdL12, SPR^{+}19, SMR^{+}12, TWJ^{+}13, WSYO11, WBB07, XXJS18, dOED^{+}20]. Services [DCA82, HP87, Hun81, WL81b, BMY03, BJP^{+}00, BMC17, CGM^{+}03, DGM^{+}22, DTB12, GB23, GAAA22, KCG^{+}12, KJHG10, LQ04, dSMH13, MZC10, MRZ15, MAJ15, PT14, PALNGD^{+}06, PDROFRM13, PCC^{+}12, RR22, RBL^{+}14a, RCMZ13, SMKZ06, SSO13, ZZC^{+}17, ZHZ17, dAKdGJ11, AC80b, CCE^{+}21]. SESAG [HLFS05]. Session [Hol89, SZ88, CA08a, RMMLSME14]. Session-Based [SZ88]. session-oriented [CA08a]. Set [Abb89, CQC98, Car97, CMR92, Kob77, MAF91, Sti85, WW89, WHLM98, Thi93, ZZL^{+}21]. Sethi [AS87]. Sethu [SFS97c]. Sets [BT89, FP82, GT93, BMS21, DKS08, HW15, JLZ09]. setting [BCPL13, Li22, RGC^{+}21]. seven [Kar21]. Several [BdJ80, NM78, CCPY12]. SEWMS [RQL^{+}20]. SGFinFETs [SM24]. SGOS [Coo08]. Shan [Pit82]. Share [Lar75b, BA79]. Shared [BAFR96, BS90c, EMVW83, FHJ94, GT92, IS05, LKBT92, RK91, Rey90, RA95, SJKL94, WZF94, AO12, Bul73, GCF15, Har80a, LX04, PT14, Ren23, ZWKX17]. Shared-Memory [BS90c, GT92, LX04]. Sharing [Fon85, HI85, LLM05, NMG11, Rei72, RNS^{+}16, TB73, WR84, ZZWD93, CFS^{+}22, DTJ89, GKLM79, HM18, HC20, HKWZ00, Lio79, NMH^{+}24, NS01b, Ott82, Rog71, SWY^{+}21]. sheet [Tur22]. sheet-defined [Tur22]. shelf [TS02]. Shell [RDC89, YH97, Wei85]. Shelley [Atk83, Edm86]. Shepherd [Sau88]. Sherwood [Bul72b]. shift [Kra10, Wu02]. shift/reduce [Kra10]. Shimba [SKM01]. Shneiderman [HW88]. Shock [Pet77, ZGL^{+}23, Pet77]. shop [DWL^{+}17, LP83]. Short [Ayc15, CLKG16, Gol81b, HW15, Rai72, Sam71b, Sch83b, Wil79, CY01b, DWL^{+}17, LM81b, SH82, SOKE23]. short-circuit [LM81b]. short-term [DWL^{+}17, SOKE23]. shortest [MG94, Tor23]. shorthand [Wya84]. Should [Atk79d, TB72, BA79]. shrinkage [TBF^{+}22]. Shuttle [Coo08]. SID [BCP71]. Side [MM86, AvRAF09, ST19]. Side-effects [MM86]. sided [PGK^{+}10]. Sigma [Ano88b, LG73]. signal [AAB^{+}21]. Signalling [Rey90]. SignalPlant [NPHJ18, PNK^{+}23]. Signals [GRR06]. Signature [MAT94a, GDW^{+}23, RMMLSME14]. Signature-check [MAT94a]. Signatures [BR95, TT82, BPP10]. silicon [LFH^{+}22]. sim} [RRFD23, KLLK98]. SIMD [CFKT17, FHL^{+}18, LBK16, LM22, PL91, PKH07, RB89]. similarities [EMD13]. Similarity [FFD96, PT00a, BRTT09, GK21, RRK^{+}18]. Simon [Lav78]. Simons [Rop88b]. Simple [App89b, CM96, DV84, Dew86, Ell82a, FH92b, Han79b, Han83c, HM12, HMS88, Hel95, Hop80b, Jar75, LS75, MM81, MM88, Mil74, OW89, Ram96, Sil92, SW94, Sta07, Wad85, WW91, WPN86, dCV88, BHT^{+}23, Fav07, LP83, MR04, Plu74, ZXX24, Daw77]. simpler [Pre85]. Simplicity [NNL^{+}14]. Simplicity-first [NNL^{+}14]. Simplification [Joh78, Kan97, PB87, VWB91]. Simplifying [GG08, PWI^{+}21]. SIMULA [CK78, LT85, Pal74, Pal76, Pal82, KO86, Pal78b]. Simulate [QSA88, CL19, WCsH16]. simulated [Cer18]. Simulating [Bad98, Gob71, Lev80, BKL^{+}02, HMRZ20, PLL^{+}02, ROFGFR^{+}16, ZGL^{+}23]. Simulation [BL90a, Cav83b, CW94, FF80, GR95, GARSR18, Gom78, Gom82, Hać84, KLLK98, KS80, KNC94, LL91, LS81, LB81, Mac77a, MS90, Mar84b, MAF91, Ols90, RB89, RT91, SCR94, SR88, She75, Sti78, SR91, TWL94, BHvR05, Bru84, CRB^{+}11, CNRB13, DPH16, Dar00, Dav74, EMRK20, ENP^{+}24, GB13, GDGB17, Haá82, HIRB24, IMBB20, JAA^{+}20, KS01a, LJL^{+}10, MVOD19, NUB24, PDCB17, RRFD23, RR05, RCH^{+}23, RMMH^{+}23, Sha77, SGDA18, SHB19, SYB04, The77, TLC^{+}18, WSN^{+}23, ZSRR22]. Simulation/Regression [Gom78]. Simulations [Ben89, SYB04, SDC04]. Simulator [ABRW94, Coo08, DM84, HHL84, Pas87, SRS98, SG97, ACG^{+}21, BHT^{+}23, BC13, DC15, Ham81, LLK04, SAL16, SR02]. simulators [DGR^{+}06, JHKS19, Man18]. simultaneous [EBFK10]. SINA [TBA89]. Sine [Col77c]. Single [BCHS98, CRT80, HEV^{+}98, MLR19, PP16, Ste98, IMKN12, KHC^{+}19, VH04, Was12, MLR19]. Single-accumulator [CRT80]. Single-Address-Space [HEV^{+}98]. Single-copy [Ste98]. single-instance [KHC^{+}19]. single-instruction [IMKN12]. Single-Producer-Single-Consumer [MLR19]. Single-scan [PP16]. sinks [FVF^{+}18]. SIPmsign [RMMLSME14]. SIRSALE [Mos06]. Sisal [KGP96]. site [LS03]. Sites [Fin97]. Situ [RGK99]. situation [YHGY06]. situations [RGSGHGCG21]. Six [DJM97, WKL76]. size [LPF^{+}11]. sizes [JDPB08]. Skeletal [Fra75]. skeleton [GVL10]. sketch [SWBS17]. sketching [CGH08]. Skip [Coh98]. SLA [PM12]. slanted [Ber99]. slanted-baseline [Ber99]. SLATEC [JK83]. Slave [BK87]. slicer [FDHH04]. Slicing [ADS93, BSDF20, GHBH05, GMC00, KH18, NJGG12b, SM18, WZLN08, ZGG07, NJGG12a, NJG14]. Slide [RR85]. Slisp [BP97]. SLP [Jor78]. Small [AJ78, Bar74a, Bar83b, BW71, Bow73, Gob71, Gol81a, Kin71, LF74, Lyo85, Tho78, Van82, BMS21, DDF16, Dun75, GKLM79, JLZ09, NSKK83]. smaller [LSYKK16]. Smalltalk [PL14, Ben90, FG14, SMR89]. SMArDT [DHG^{+}19]. Smart [TEBK99, BSM^{+}21, CWZ17, CRGIP15, JGB15, KRK21, KH07, Ler02, LYX^{+}17, LZL^{+}17, LWZ^{+}21, PLFIVRAR23, PPPP23, PRP^{+}23, RQL^{+}20, Sav04, SRC^{+}18b, TJB^{+}19, XP23, XWC^{+}17, ZZC^{+}17, ZLZ^{+}21, XLZ^{+}20]. SmartCLIDE [PPN^{+}23]. smarter [QTV23]. SmartHerd [TJB^{+}19]. smartphone [SJK^{+}21]. smartphones [DF15]. SmartSantander [JGB15]. smartwatch [DMC17]. SMD [MCG^{+}88]. smell [LZK^{+}23, ZSH^{+}23]. smells [ASME24, SPR^{+}19, SBF19, ZHZL22]. Smironov [Cox76]. Smith [Bar75c, Gru83, Lav77]. smooth [TRGA18]. SMP [KGL06, ZLG08]. SNIPE [Daw77]. snippet [FG08]. Snobol [Lar75a]. SNOBOL4 [Abb78, DM77, Fle82, Gri75, Han76e, Han77c, Han78d, Liu86, Pag79]. SOAP [FJ03, Sco73]. SOBS [RO77]. social [ABA20, Ber20, Ken77, LQS^{+}24, XWC^{+}17, ZYW^{+}20, BLNU15]. social-based [Ber20]. Socially [AFNG20]. Society [TK72b]. sockets [NAGL10, SM01]. SOFA [HP11]. Soft [CGL76, AC13, Atk78, FPAF18]. softback [RB82]. softerware [LX04]. SOFTLIB [SWBT86]. Softw [XZ01, XZ03]. Software [Aji95, AA20, AA21, ACC95, AR93, AS78, And89, AKDN90, Ano87a, Ano93a, Ano09, ADH^{+}00, BA78, BP84a, Bar76e, BP09, BH82, BP90, BC21, BTM81, BL78, BL79, BP97, Bro74, Buy21, Byr91, CK86, CPD13, CMF^{+}98, CM83, CLW90, CLL91, CLLT98, CWZ17, CPHS83, CW92, CG93, Cor08, Cra77, Cum71, CZA83, DJM97, DRL82, DP85, FV03, FKV98, FL75a, FS81, Fre78a, Fre78b, Gar86, GH19, GLW82, GHM^{+}06, GH09, Gri80, Gri82, Gro73, GS85, GJ93, HH80, Har95, HL92, HC13, Hat73, HK84b, Hop96, Hos98, HHL84, HD86, Inc83, IS05, JKRS85, JL80, JP74, Jor90, KLLK98, Kat71, Key92, KO91, KR85, LL96, LN71, Lea82, LM81a, LL91, LCW98, Lin86, LF90, MKA^{+}22, MK01, MER84, Moh81, MM97, MNM79, MS80b, NHB23, NHP81]. Software [NPW72, Not90, OLS89, ORT81, Pal78a, PW97, PL91, PLR85, PW93, Poo88, PP98, Pry85, Pyl72, Rai73, RDLK90, RBB12, RVS^{+}20, Rin84, RCC91, RS21, Sam81, SB21, SM79, SLBC^{+}23, SF85, Sch82, SM85, SAN^{+}81, SKD^{+}22, Sno78b, Spi76, Spo71, TKB78, TP92, TV96, TLMP93, VL73, Wai75, WPT95, WCK11, Wat89, WA77, WRD99, Wie96, WSN^{+}23, WH98, Wir72, Wol82, WS74, WI85, Woo71, Woo84, Wor83, Yu96, ZKZ^{+}21, vdHW03, AJ04, AMOS19, ABBS22, ABG^{+}22, Als22, ACG^{+}21, ALG23, ALF01, Ano88c, ACCD01, AGM17, BCPL13, Bar83a, Bar15, BP11, BBB^{+}22, BP02, BCSW20, BMSZ17, BGM17, BBS11, BCP19, BGP17, CK13, CGP^{+}06, CKRC20, CMPQ23, Car22, Cer18, CCR19, CGH^{+}15, CCM05, CR18, CCE^{+}21, CSS15, CMTCC^{+}17, DPH16, DB09, DBOT23, DSD^{+}05, DFOT10, DdB15]. software [Deu99, DHA11, DHG^{+}19, DRH^{+}23, DBH04, DFRR15, Eba18, Eba20, EAB^{+}03, FRGPLF^{+}12, FCO^{+}19, FMPR02, GH03, GN00, GKWS11, GdLC04, GEI^{+}11, GMP^{+}21, GVBG22, GSPA^{+}11, GW04, GM22, GH02, GWX^{+}23, HYY23, Han77b, Han11, HGK^{+}19, HPP22, Hoa72, HPZ^{+}20, HL03, HNJ^{+}23, Inc85, JLZ09, JTG^{+}11, JH03, JL81, JC19, KKLL99, KJB11, KRB21, KVG19, KCH08, KSSJ^{+}23, KB06, KSRR17, KV17, KHMB17, Lar08, LKK^{+}18, LKK19, LHC15, LHFL07, LLS06, LWZ^{+}19, LYLY20, LGRL08, LPA13, MH05, MMOD16, MVM^{+}22, MVV12, MRBB19, MCLL21, MC23, MST13, Mer74, MdCGdC^{+}17, MTPC14, MOTG18, MRG^{+}19, MPJ20, MK03, MCHN05, NB19, NMG11, NDKL^{+}24, NM06, OFRW10, PKK12, PLR13, PH14, PDPQ22, PGK^{+}10, PW11, PPR02, PVR99, RRK^{+}18, RBL^{+}14a, RN00, RSRCGC15, Rop88b, RLB^{+}11, ST12, SSCdA^{+}03, DAP21, SCD^{+}23, SDDD10, SSM11]. software [SBB22, SAY16, SYG^{+}18, SJA^{+}11, ST14, SMLS23, SRC^{+}18b, Sny08, SDF^{+}21, SBF19, SBFF24, STA09, SB22, SROV06, SKM01, SGCM11, TM14, TP03, TV09, TMS18, TWJ^{+}13, TGC15, TTJ^{+}09, UCCPM19, Val78, VvK99, VC02, VBS22, Wai07, Wal81a, WP00, Wan82, WY18b, WBN^{+}20, WHS^{+}00, WYAZ15, XCL^{+}18, XWHX21, YHGY06, YWT^{+}12, YQL22, Yuv78, ZSAM23, ZWKX17, ZWAL22, ZZ11, ZZLL24, ZZL24, ZNWS18, ZYW^{+}22, ZZJZ24, dSdMSNO^{+}11, dAPMV10, vGPB10, vO03, CCLN22, CSZO23, GH11, TWS^{+}22, ZSAM23, Zam03, Lan75, And78, Bar73e, Bar75e, Bux78, Cla98, Pra96a, Pra96b, Rob72, Rop88a, Wal84b]. Software-Defined [RS21, ACG^{+}21, LWZ^{+}19, YQL22]. Software-in-the-loop [WSN^{+}23]. SOHO [JH03]. solar [ZPSC07]. Solaris [MM06]. Sole [BTZ94]. solid [LFHL22]. Solnsteff [Bar74c]. SOLO [Pan72, Pow79, Han76b, Han76c, Han76d]. Solomon [Pla97, PD05]. Solution [Car82, HP88, BDL09, dSCdRS^{+}19, CMTCC^{+}17, GLMS18, GBE^{+}09, JJK^{+}12, KW17, OHD22, Plu74, SRRFGC^{+}10, SAY16]. Solutions [KS84, CMF^{+}17, CBB17, DPS03, EHV99, GAH05, KJVS21, RJ09, ZHZL22]. Solver [DV84, Ram96, DDP07, ST01]. solvers [GCARPC^{+}01, Hoh04]. Solving [Deo10, Kra97, RM75, SO77, YH97, ZZL^{+}21, Ano79a, Atk79a, BK23, BOPN12, GK08, Kil12, LQ04, LV20, MSR^{+}07, Wal83c, WLZ^{+}22]. Some [Ano80b, AvdSGS80, Bas00, BCP71, Fen01b, GM73, HLS73, Heh76, Jos80, Kul74, Liu86, NPJ79, New86, Pal86, Pyl72, RK15a, Ree71, Sco77b, Vel85, Ham77, LQ99, Sab76, Sco81, Wad87]. Sophisticated [SC90]. Sort [BM93, Thi89, ZML22, Che04, Har81, Che08]. sorted [ALG23, Har81, LBK16]. sorters [BMS21]. Sorting [Har81, Mus97, BT07, CPP12, Hea81, IMKN12, Val00]. Source [ADM96, BAP87, Bro72, Bro77, CH73, Con85, Inc84, MK96, OMA96, Pet76, WR79, vDV04, AMOS19, AL21, ACKT20, AG06, BN00, BUT14, CZL24, Cia07, DPH16, DP09, EvG04, FRBRF19, GLMS18, GEI^{+}11, GMGL24, Gla82, GHBH05, JM08, CCK21, Mił10, MF08, NMG11, NZPWR22, PMP^{+}16, PNK^{+}23, PdSCJM22, RMM19, SO21, SBS20, SRGCPB^{+}09, SSS22, SIK^{+}16, Yi12, ZWSS15, vGPB10]. Source-to-source [ADM96, Yi12]. sourced [ABG^{+}22]. SourceForge [TBPK20]. sources [ARCN^{+}06]. South [Bar78d, Bar84b]. SP&E [CY01b]. Space [AC80a, Col83, FH91b, Gri86, HEV^{+}98, KR83, Pal86, RA95, SY79, Wad87, WW83, AI22, DDF16, ENP^{+}24, GNSP12, Gol81b, KSSA23, Kur99, NAGL10, RK15b, SDF^{+}21, SB03, YYSG11, Zdu07, Ano71d]. Space-efficient [AC80a, KR83]. Space-filling [Gri86, Pal86, WW83, Gol81b, Ano71d]. space-optimized [RK15b]. Spacecraft [SRS98]. spacefilling [BG01]. spaces [CIM24, SSD11]. spam [PDROFRM13, ROFGFR^{+}16, ROFGFRM16, ScG09, Cor08]. SPARE [WC04]. Spark [BHNDdMRG23, dCCCdAC20, Kil19]. Spark-based [Kil19]. Sparse [HP88, MM02, Tor23, CW91]. Spatial [NSM86, ANSK16, dCCCdAC20, SB13]. Spatio [GMGB22]. Spatio-temporal [GMGB22]. Spatiotemporal [PPR^{+}21]. Spatiotemporal-based [PPR^{+}21]. SPB [FCO^{+}19]. SPE [Cor08, KP94, BL90b]. Special [CSZO23, Cor08, Gru79, ISPB21, KSRR17, KD83, Mac79, Oli83, RBB12, RWJ^{+}17, Sch76b, TBT23, WCK11, WBPR20, XDVB24, Bar73d, BP11, BN13, BCSW20, BCP19, GK14, PL14, SFB13, BC21, KH12, ZWC21]. Special-purpose [KD83, Mac79]. specialist [Cla86]. Specialists [Pal79]. specialization [HK06a]. Specific [FH82a, Lea77, BFG^{+}11, BFRFB22, EC13, LJS20, MPBH13, SZ09, WGM08, WAH^{+}12]. Specification [ACC95, CCC96, FF80, Hof89, HL98, Jaa95b, KU97, KvEP95, LY92, LOBF88, Lop89, Mat83a, MXYQ86, OMA96, PP98, SL87, TWNH12, TL98, WKS^{+}98, WKD96, ASP^{+}19, dODP21, Bla04, CSMML12, HL02b, KW09, KAYH^{+}99, ML08, NN24, Pol01, Rop88a, WWGP10]. Specification-based [WKD96]. Specifications [BM97, FGMM93, Geh82, HL91, Jal87, KLLK98, KN88, ÖS96, Özc98, Par85b, SG97, VSC93, vHLB^{+}88, Ano80a, BLLP04, JTG^{+}11, LPP09, SK03, Tur06, WM20]. Specify [Abb89]. Specifying [HvdH02, RS94, TTJ^{+}09]. Spectral [NNR18, NNLR17]. Spectral-based [NNR18, NNLR17]. spectrally [SV22]. spectrometers [WPL^{+}21]. spectrum [HGK^{+}19]. spectrum-based [HGK^{+}19]. Speculation [Nee77b, GXN10]. Speculative [AA14, KKN04]. speech [CK99, ZZC^{+}17]. Speed [CB72, Har80a, Mer73, Sch86, KSH^{+}15, Li22, ML20, SRS98]. Speed-Up [Sch86]. Speeding [CRR94]. speeds [Li18]. spellchecking [DRG11]. Spelling [CS82, MM90]. SPHT [ZWF^{+}22]. SPIDER [SESK23, dMdLvS99]. Spilling [FH92b, Bur16]. SPIN [CG96]. SPINE [BFG^{+}11]. SPITBOL [DM77]. splay [LM07]. splaying [BDD09]. Splaysort [MEP96]. SPOON [PMP^{+}16]. SPOPB [LFHL22]. sporadic [FZS^{+}17]. spot [LMK16]. Spotting [LA11]. spreading [KHS^{+}20]. Spreadsheet [DW90, SP88, Tur22]. Springer [Atk79a, Bis86, Cav83a, Mee87]. Springer-Verlag [Bis86, Cav83a, Mee87]. Sprite [DO91]. Sproull [Bra75]. SPSS [LP78]. spurious [OY10]. spy [LZWX22]. spy-robots [LZWX22]. SQL [BRTT09, FSC08, LG19]. squeeze [CD01]. squeezing [Coo85]. Squinting [McI90]. SQVDT [AL21]. SR [And82b, AO88, Ols90, OM96]. SRAM [SM24]. SRE [BHZ85]. stab [CMM75, Art82, CST75, Col72b]. stab-1 [CMM75, Col72b]. STAB-12 [CST75]. Stabdump [MM80a]. stability [MVV12, MC23]. Stable [Any85, Mot81]. Stack [Cia07, EE90, GR79, Har92, MY87, Ste98, DRH^{+}23, SS19]. Stack-based [GR79]. stacks [LC05]. staff [DHA11]. stage [Abe07, CGH08]. Stages [Wal86a, Abe10, Val76b]. stakeholder [NDKL^{+}24]. stakeholders [SWY^{+}21]. STAMP [JH03]. Standalone [SIC^{+}20]. Standard [ALG23, De 96, GM85b, REC75, BLLP04, BDLM04, DKS08, PBGM18, RB82, Mar84b, Han04, Bar72c]. standard-based [PBGM18]. Standardization [Bar80b, Pal76, TWNH12]. Standardized [Hol93]. Standards [YSBL22, Jak04, JP79, Wu00, Ten85]. Standards-based [YSBL22]. Stanley [Val76a]. Star [Gom82, KDP83, PP16, SF98]. star-join [PP16]. StarMod [LGC84]. stars [ABC^{+}21]. Starvation [KLY20]. State [Atk79c, AZ97a, Bar76b, Bul72b, CLR84, Fos89, GJ93, HC93, Hut76, KDP83, KM94, RS94, Rog73, Wil76, ABL08, Atk82a, BDSV99, Bar79b, CSD23, GN16, LFHL22, LPP09, dSMLP22, MKE18, PJJM21, Pat94, PMC22, Wil74a, ZWX^{+}21, BFGL20, XZX^{+}21, ZPSH21]. state-based [MKE18]. State-of-the-art [XZX^{+}21]. State-transition [Fos89]. statecharts [CMT02, KH18]. stateful [JGSG^{+}21]. Statement [Bar74i, KP94, Ber85a, HM82, ZWSS15]. Statements [Sal81b, Van92, Atk82a, LLM05]. Static [BCHS98, GMC00, HAM18, JM08, Knu84, MPC^{+}19, PLR18, SB93, WB78, ACSK23, ALG23, BCPL13, BFGS05, BWA82, BPS00, CFC15, Fer13, GOQ16, GRA14, GS06b, HOY17, KSH11, NNLR17, OY10, PKvdWB17, RaMB^{+}24, Söz15, TVCB15, TSO19, VH04, YC16, YLR23]. Station [BB81]. stations [SCLD21]. stations-oriented [SCLD21]. statistic [Cox76]. Statistical [WPT95, CC13, EF13, FO10, HYZ^{+}18, Ken77]. Statistics [Cra76, HV88, LV73, Yuv75, Kul74, Maa06]. Status [BS81, BL15, MHN18, PES^{+}20]. stdio.h [Lev97]. Steady [CLR84]. Steady-State [CLR84]. Steel [Lav77]. Steensgaard [LLN16]. steered [BP02]. Steinbrenner [Ken77]. Stenfert [Nee77a]. Step [Cas92, Deo02, UN19]. Steps [CS91a, Ush77, BLC19, DSD^{+}19]. Stepwise [Dro85b, MBG19b]. STLlint [GS06b]. stochastic [EYG22, GQ15]. Stock [GL97, RRR97, KCYY12, WLS^{+}21, YZW^{+}12]. STOIC [SB83]. stone [Kar21]. Stony [CVV97]. StopGap [NTF^{+}17]. STOPPAGE [GMGB22]. Storage [AHS86, Any85, Bot77, BS93, CDKK85, CL95, DLP85, Far74, GM85a, Gol81a, Hal86, Han77c, Han80b, KK97, LH82, LV73, PMY97, SHC74, Wal81b, DD18, DD21, HBM06, JKW74, LWZ^{+}19, MRR^{+}08, PM18, SCF^{+}17, WCsH16, ZXW^{+}17]. Store [Pow87, WR84, LLLY19, PACK07, SZ88]. stored [SBS13]. STORK [BL15]. story [KV14, SD75]. storytelling [HBD04]. Stoughton [Eme84]. Straddling [JC19]. strategic [BMR14]. Strategies [ALBN81, BPM93, CMPQ23, CLZ98, Wei72, CCC^{+}16, CCPY12, GAF^{+}09, Lan71, SJA^{+}04, ZWML14]. Strategy [Hua87, Kob77, BB99b, DW13, MKM^{+}17, PDPM^{+}16, SC14, SCLD21, WLZ^{+}22, ZYYC12]. Strategy-Independent [Kob77]. stratigraphic [LJS20]. Stream [HKW77, ACV10, CRC18, DLWF17, DHWZ14, GAF^{+}09, GA12, Ged14, JKH22, KAS^{+}14, MSB20, QH21, SHGG16, TAG^{+}10, VGF21, SM01]. streaming [Kil19, RSLAGCLB16, SIK^{+}16, SAA^{+}20, ZSFY05]. streamlined [NM19]. Streams [Coh98, Wis93, CA08a, RBRB23, AP91, GA12, OM16]. STREAMS-Based [AP91]. strengthening [SB22]. Stress [Pro92, ZC02, ABRW94]. Stretching [App22, Ber99]. Strevens [Bar81]. strider [SHF16]. Strides [WH97]. String [ARV77, AC75, BY89, BK93, Dav82, HS91, JTU96, JGR89, KST94, Lec95, Lec98, Liu86, LDI98, Nar94, OM88, Rai92, Smi94, TP97, TT82, Wri94, de 82, Ayc15, CFC15, Fen01a, FBMA05, LC03, Mha05, NT05, THG17, WC04]. string-searching [Mha05]. Stringlish [Ayc15]. Strings [Bis79c, BAP95, Hor80, Nil88, Sal79b, Sal79a, SM90, Bar74b]. Strongly [Pow87]. Structural [Lyo85, Pil75, RS87, STH97, Sha78, Wat89, BLNU15, LD14, RK15a, VDMW06, ZGL^{+}23, Liv75]. Structure [ACG78, ADM96, CK97, Dan90, Des74, Fen94b, Fen96, Han81e, HK84b, Hur80, KFJS88, Kaw79, Kaw80, Not90, Oes71, Rai81, Sti79, Web87, Wil82a, You81, Ano16a, Bra99, DDP07, Deu99, Fen94a, LBP^{+}13, Mof99, MFYiA01, OAF^{+}03, Sha72, ZGL^{+}23]. Structured [AI80, CP76, Fel81, GS90, Ham79, HP83a, Lea77, MW81, Noo83, TCC^{+}94, TWI88, Wel78a, WA77, WI85, ZB74, Bea78, Cou85a, FS82, GVL10, GG96, HGWBS75, LLK04, Mar85, Mor77, Pag79, Wal81b, Wit77a, ZML13, Zel77, Bar75f, Bar76d, Bar79b, Cou85b]. Structures [All89, AMS92, AS83, Bae73, BY90, CLW90, Dea86, Dew91, Dew87, Dun93, Edw77, FM86, FW78, GM77, Hal86, HS83, Hud72, JG89, Kow81, Lec98, MIA94, Nil88, Pal74, PDC^{+}98, Per85, SMR93, TB86, TD94, Wil84a, vR92, AS08, BWA82, CA00, Dan82, GP14, Lev80, LJS20, SCPL23]. Structuring [Hay83, Jor90, MK96, Ten82, Val84, Ell79b]. student [JL80]. Students [Nut76, Bis81a]. Studies [Eme84, Inc86, Wic72a, WH97, RN00, SRCP19, VDG^{+}00, Ree73, Han77a].

Studio [Gro73]. Study

[AC80b, BA78, Ben89, BTM81, Blu86, Byr91, CDV88, CFP83, DH88, Dew93, DS86b, FIL86, Fle90, Fre78a, Geh82, HJS89, Ham77, Hoa73, Hop96, Hop80b, Kat71, Kat83a, Knu71, Lai95, Lav77, LAD^{+}94, LB81, MBO97, MG76, Ols90, RK89, SNM80, TV96, UGBW91, WL81a, Zel80, AB88, ADH^{+}00, Atk78, ASME24, BLLP04, BTS09, BLE^{+}08, BGM17, CKRC20, CGH^{+}15, CMS07, CBF23, CIM22, CIM24, CFD^{+}22, DB09, DHA11, DMC17, Eba18, EGL18, ENP^{+}24, Fen01b, FMNW04, FC98, FSRF^{+}22, FBH^{+}23, GLL20, GMGL24, GK08, GW04, HJ14, HYY23, HBK20, HP11, JHKS19, KCS^{+}20, KRZ02, LF82, MS99, CCK21, NDKL^{+}24, OOG19, OMGDG14, PCdGPP12, PGK^{+}10, Pol01, RBR21, RdOTF14, RLB^{+}11, SLBC^{+}23, SN07, SSS22, Sne78, SBF19, STT24, SW12, TRRK21, UT19, VP05, WXR16, WHS^{+}00, WBB07]. study [XP23, ZNWS18, ZRX^{+}99, dSdMSNO^{+}11]. Stxxl [DKS08]. Style [Fai87, GSWZ95, UGBW91, Wol91, Zim90, Bar76c, KPU04, LHFL07, MA01, vO03]. Styles [KS95]. stylesheets [ÜY22]. Stylistics [Sal79d]. Subclassing [Man88]. SubCollaboration [PK11]. Subgraph [McG82, KH04]. Subject [Car85b, WJC^{+}14]. submatch [BT21b]. submission [LJ99]. SUbprogram [Sto94]. Subroutine [Ker80]. Subroutines [JBCB79]. subscribe [RC10]. Subscripted [Bel74]. Subsegment [WJ93]. subsequence [Deo10]. Subset [DL23, Pag79, BC17, MS83]. Substituting [PB03]. Substitution [CHT91, LLH14]. substitutions [CSD23]. Substring [Har71a, Smi91, Maa06, Rai99]. subsumption [BGG01]. Subsystem [AP91]. Subtype [BR95]. subversion [MV16]. Succeeded [Pal78a]. Success [SO77, WJ93, AKIM24, DBOT23].

Successive [Mor80]. succinct [GP14]. Sue [Bar82c]. Suffix [AN95, BST10, GKS03, GR17, Kur99, SS07]. SugarCubes [BS98, BS00]. Suitability [BK87, MKE18, MBBS21, OMGDG14, RH78]. Suitable [Hal86]. suite [CD84, CFC15, PM17, Sta05]. suites [MW13, ACSK23]. sum [PV21]. SUMLOW [CGH08]. summaries [Pet77]. summarisation [SYXZ14]. summarization [GLS^{+}23, GCZ^{+}24, ZYF20]. summation [Ush77]. Summations [Gut76]. Summer [USE85]. Sums [Mey78]. Sun [AM86a]. Supercomputer [PZA87, PL91, BB99a]. superior [YHGC20]. Superlinear [Sch86]. SUPERMAC [Bro80, BO83]. superoptimization [HW15]. superpaging [QM13]. SuperPascal [Han94b]. supervised [MAA^{+}22, ZHZ^{+}14]. Support [CLW90, CDG^{+}98, Far88, FKV98, HMS^{+}95, Joh79, KJHG10, MD88, Par79, PN83, RK89, RRR97, Val76a, WR84, YHGY06, AA20, BVGVEA13, BBMG08, BFHR99, Bla04, BV06, BCL^{+}06, CLZ99, CMPQ23, CCCZ05, CTLL07, CHCC07, CLP^{+}09, CEF02, DFPT09, DH00, FL02, GH03, Ged14, GRVC23, GH02, GVG^{+}18, HRS^{+}09, KGL06, Kim15, LCW07, MSB20, MTM22, Mos73, PDPQ22, PBGM18, SSD11, SJP^{+}09, SF88, Ste02, TRRK21, TY14, WP00, Wu00, ZLG08, ZWAL22, vD99]. Supportable [Hua87]. Supported [CMF^{+}98]. Supporting [AGM17, BE81, CDGP93, DHS01, Dew91, FPT07, GHM96, LP86, MR96, WA77, CLSE05, GDH13, HLR^{+}03, KHGSS12, PTU03, RBS14, RPP07, Ter86, WP05]. Supports [Bar78a, Wil82a, CLC09]. suppression [AAB^{+}21, JWTG11]. SUPRA [Sto94]. SUPRA-RPC [Sto94]. Surface [FR78]. surveillance [DDB^{+}18b, XWC^{+}17, XLZ^{+}20]. Survery [BMC17, CBB17, KKA^{+}16, MAW^{+}16]. Survey [AH85, FFRFS19, GMGL24, NRS13, PH14, SBD15, Sco81, Wil82b, ASA22, Bar74c, BST10, BGS20, CBB17, DBOT23, DCS23, FO10, GVL10, GB14, KBM02, KSST22, LZ10, MZC08, PCBR18, VMY24, CKCG23]. Surveys [Liv75]. sustainability [BIA24]. sustainable [CWC^{+}21, JXG^{+}21, JFZ^{+}21, ZSAM23, ZZJ21, ZKZ^{+}21]. Sustained [Mer73]. Suturing [RBD^{+}24]. SVD [SI10]. SVM [ZZC^{+}17]. SVM-based [ZZC^{+}17]. Swap [CBC00]. swarm [AA20, JSPP21, MS18, HBM06]. sweep [BMAV05]. sweeping [CMES05]. Swets [Flo73]. Swing [ABL08, WWJ07, WW09]. SwingStates [ABL08]. switched [WAML12]. switching [HC20, Vel88]. symbiosis [AWNS18]. Symbol [Dew87, Fen96, Fen98, Rai99]. Symbolic [AM86a, FR78, Fil98, How78, KE85, LF90]. Symmetric [DBO^{+}18, RGK99]. Symmetry [Che08]. symposium [Rob72]. SYN [EGL18]. Synchronization [AO88, DD10, Hop86, Hos98, JLR79, RM75, TE90, TL98, WH84, WKD96, dCV88, CY01a, CY01b, DZ21, DO07, UN19, Whi87]. synchronizations [TNGT09]. synchronizing [Wet77]. Synchronous [BMZ92, CPHS83, Geh90, MM97, CLZ99, DFPT09, WAML12]. Synchronous/Reactive [MM97]. Synergies [BGM17]. synergy [BRS18]. synonyms [EMD13, SO21]. Syntactic [DP95, Yan91, Kra10]. syntactical [ZYF20]. Syntactically [Con85]. Syntax [All83b, Ber85b, BHZ85, Con84, CFP83, FL76, Fis82, HW88, KL86, KPT86, KU97, LT83, Mar84a, PL91, Rec79, Rey87, Set81, Set79, SK96, Thi97, AG06, Har82, Mau82, LZLL22, Wal83b]. Syntax-directed [FL76, HW88, KPT86, KU97, PL91, SK96]. Syntaxes [Woo86, MGG^{+}09]. Synthesis [Bha88, CW94, KM94, MP82, WC81, GMPL11, HZ95, JRGC20]. Synthesized [GZ93, WRD99]. Synthesizer [Cla86, CW82b, CIM22, CIM24]. Synthesizing [ABBS22, Jal87]. Synthetic [SJKL94, BM06]. Synthetic-perturbation [SJKL94]. System [AB89, ARS^{+}94, AE06b, AMR90, ACDP85, Any85, AM86b, AP95, AM78, AN81, ACC83, AMW91, Bad98, Ban71, BL85, BP84a, Bar78a, BK77, BLL88, BS90b, BM97, BMA72, Bro71, BSRS85, Bur98, BK86b, BW95, BNOW92, CC73, CC84, CC87, CT92, CC97, CLC09, CS91b, CG95a, CAC^{+}84, Coh75, CM82, CPW74, CGL76, Com82, Coo86, CM85, CW80, Cra76, Cum71, CP76, DNSG89, DP85, Eva71, FR78, Fil98, FWS74, Fos89, FL75b, Fra75, FT79a, Fra93, FL94, Fri92, GMM90, GW85, Gay80, GKM90, Gom78, GLW82, GWA91, GW84b, HJS89, Ham84, Han73, Han76b, Han76c, Han76d, HF80, Han80a, Han80b, HHR93, Har83, HUS^{+}91, Har80a, HMS88, HF73, Hef82, HEV^{+}98, HK84b, Hol77, Hol83, HCC96, HL03, Hug97, Hum76, Hun97, Hus86, Hut79a, IR80, Inc84, Jeg83]. System [JLR79, Joh84, JZ93, KDP83, KH12, Ker80, Kil71, KO91, Kin71, KM89, KK90, Kue95, LNW82, LRMM93, LLCG^{+}89, LCC97, LA90, Les72, LL91, LH82, Lev82b, Lin79, LS81, Lin87, LP86, Lio79, LQ93, Lor91, Lun89, MK90, MS74a, Mac96b, MWB95, MBO97, MCG^{+}88, Mar83, MR96, MT94, MPP87, MM97, NY78, NS74, Nut76, O’N88, Oes71, OF76, PSV85, Pan72, Par79, Pat94, PZA87, PN83, Poo71b, PR90, PJ75, Pyl72, QSA88, Qui91, Rag86, Rai73, Ree71, RS82, RAB^{+}79, RH77, RB75, Rob83b, RRR97, Ros77, RT91, RRP95, SB83, SG93, SW86a, SWN94, SMR93, SS89, SB82, SH98, Sno78a, Som82, SWBT86, Sre76, SNM80, SYRS80, SL87, SMR89, SR91, SO77, Tal71, TB73, Tha84, TF79a, TF79b, TWL94, TB72, TS81, Tic85, TKWW85]. System [TH86, VS88, VL73, VC90, WR95, WC87, Wha72, WB85a, Wil82b, WPN86, WR84, WG89, WCE^{+}72, WR77, Wit83, Wit82, Wol92, WS74, Wor83, ZM95, vdRW79, AH12, AKNJ21, ANSK16, ACV10, Ano23m, AZS19, BGM99, Bai85c, BMR00, BJP^{+}00, BGS18, Bar76a, BHR^{+}02, BGS^{+}13, BLR^{+}17, BCL13, BDG^{+}00, BCFT95, Bro82, BLNU15, Buy00, CL09, CCE99, CGH^{+}04, CFS^{+}22, CF05, CR18, DFST08, DFPT09, DFOT10, DH00, DD10, Deu99, DGPT14, DHMS11, DNL^{+}20, EC13, FL02, FR09, FSS99, GC20, GN00, GBG^{+}14, Geh83, GRS74, GHM^{+}06, GCK^{+}02, GZX^{+}21, GJ22, Haá82, Han83b, HBM06, HTJNL19, HATvdW99, HJC00, HL02b, HC12, HYT13, HLH15, HC16, Hun00, JZ10, JAkM^{+}21, JZ02, JXG^{+}21, JFZ^{+}21, JB07, JT00, KCYY12, KT01a, KCS^{+}20, KTG20, KSH^{+}15, KPGH02, Kru82, LLJ12, Lan71, Lan74b, LS03]. system [LK99, LM15, Lev82a, LCC14, LS22, Liu01, LCGS17, LZL^{+}17, LWZ^{+}19, LJS20, LJ99, ML08, MK04, MTM22, McN05, MR05, MSR^{+}07, MRG^{+}19, Mos06, NJGG12a, NJGG12b, NJG14, NMH^{+}24, NAU^{+}21, NHTT08, NW84, PTU03, PKN^{+}12, Pei02, PCdGPP12, PSRCC02, Poh81, Pol01, Pow79, Pur76, RPCS08, RO77, Rog71, RQL^{+}20, RMdL12, SGA21, SBS20, SNL15, SDDD10, SP79, SBcC07, SAL^{+}04, SAY16, STH^{+}18, SRCP19, SMGMOFM07a, SMGMOFM07b, SM15, SKD^{+}22, Spi09, SH22, TWS^{+}22, TH01, TVCB15, TVSG21, TN98, TKT^{+}07, TTJ^{+}09, VBS22, VV84, WM20, Web87, WAML12, WK06a, WS99, WHS^{+}00, WBB07, WCS^{+}17, YCY03, YZW^{+}12, Yip84, YQL22, ZPSC07, ZL84, vRvST89, BS90c, CE97, CDKK85, DD90, Fon85, HWS^{+}88, LN71, PZ92, SG90, She81b, Wei85, Wil73, Wol91, WG92b, Jac84, Mul76, Ree84b, Wic72a, Haz72]. system-administration [FSS99]. system-independent [SP79]. system-level [MK04]. system-on-chip [LLJ12]. system-specific [EC13]. system-wide [AKNJ21]. System/360 [Haz72]. Systematic [Col77a, HBB22, Kop97, Shr76, Zdu07, ARA18, BGM17, BBB^{+}11, CBB17, CZL21, DPAG11, FLPM20, FAA^{+}22, GKS^{+}22, HYY23, HBK20, LC12, MvSdL09, MKA^{+}22, MCLL21, NHB23, PVAHRG^{+}15, PLPA22, PRP^{+}23, SPR^{+}19, SLBC^{+}23, SZ09, TSMGD^{+}11, WBB07, dSdMSNO^{+}11, Ros74]. Systematically [Law78]. Systems [AE06b, AE06a, AR93, AH85, AS83, AB95, AN88, ABRW94, AZ97b, BBC91, BV89, BCP79, Blu86, BF75, Bou91, Buh93, Cas92, Cha88, CE84, Day83, Dea86, FH74, Gri82, HKB72, Han87a, HSM81, How76, HKV95, JW75, JVR97, LPT82, LY92, LOBF88, MS74b, Men97, MMS90, OSW92, PU84, PP80, PSR83, PMY97, Pfe84, Pla97, PP98, SM79, SSP11, Sch78, Sef97, ST77, TAAT84, Val84, VBH^{+}98, Wan79, Wei72, Whi83, WA77, WBV96, ZZWD93, AKNJ21, AMSG23, AKM17, AIB02, ACK^{+}22, dODP21, Bar73c, BP02, BPR01, BB75, BCF00, BC17, BGP17, BR88, BD14, Bud85, BDM16, CCLN22, CWC^{+}21, CBB20, dCCCdAC20, CEMS24, CPCL10, CSZO23, CM98a, CM98b, CBB17, CWZ17, CSTL19, Col79, CMTCC^{+}17, CMR07, DPH16, DLX^{+}22, DDB^{+}18b, DH00, DPK12, DBvG^{+}23, DO99]. systems [Deu99, DKM11, DFRR15, FVF^{+}18, FIÁSLSAR05, FCR^{+}09, FFRFS19, FSC^{+}21, FPAF18, FBH^{+}23, GH03, Ged14, GVBG22, GB02, GKBK16, GKLM79, GEF^{+}00, GLKZ21, GP01, HR06, Han78a, HLS73, HHRS03, HGK^{+}19, HMN11, HP11, HC00, HLFS05, HSY^{+}20, HKWZ00, IHS^{+}14, IAPC17, JJK^{+}12, JKH22, dSJCM16, KGL06, KRK21, Kap13, KCH08, KMY^{+}05, KS20, KBM02, KSKG12, LM02, LSK^{+}18, LHC15, LHFL07, LZ10, LGP^{+}11, MK04, MVV12, MSC23, MC23, MC02, MPJ20, NS01b, NL01, Obe11, PLL^{+}02, PTU03, PDBG10, Pit82, PCL^{+}99, PDPM^{+}16, PDPMM17, PA01, QC17, RT78, RMMH^{+}23, RB19, RGV14, RGS^{+}20b, RBD^{+}24, ROFGFRM16, RdLFF05, SPR^{+}19, STB14, SJA^{+}04, SFC^{+}21, San17, SJ79, SLRS06, SBD15, Sch83b, SM85, SRGCPB^{+}09, SJA^{+}11, SGDA18, SMT^{+}18, SYB04, SKM01, TBT23, TMJ^{+}21, TRO17, TMS18, VvK99, VC02]. systems [WM20, Wal83a, WLTJ13, WWB03, WHH21, WCsH16, Wu00, XXJS18, XZX^{+}21, XDVB24, YYSG11, YB06, YFC06, ZXW^{+}17, ZZJ21, ZC02, ZZL^{+}21, ZKZ^{+}21, ZRX^{+}99, dAPMV10, Hut76, Bar74d, Flo74, Han77a, Hut74, Jac71, Mil72, Wel72, Wil76]. Systolic [Len90].

T [Bar75c, Cou85a, Flo79, Hor07a, Lav77, Mul76, Rob72, Sha72, Wan82, Wel72, PALNGD^{+}06, Lic86, SAC^{+}92]. t-learning [PALNGD^{+}06]. T-Series [SAC^{+}92]. Tab [Wai85]. Table [CW91, Dew87, FD92, OMA96, Sew82, DHS02, Sha77]. Tables [AHS86, Chv83, Con84, DW73, Dew86, Fen94b, Fen96, FH91b, GL78, GT93, Inc81, Joh78, Lew83, Pat94, Qui83, Fen94a, HC87a, KSH^{+}15, Mof99, TEBK99]. TableSpec [OMA96]. tabling [GG08]. Tabular [LLZ20]. Tabular-expression-based [LLZ20]. TACO [CCG14]. TACOMA [JLvR^{+}02]. taffy [App22]. Tag [Gru79, EMD13]. Tag-list [Gru79]. Tagged [GH84]. tagger [CK99]. Tagging [VG85]. tags [JRGC20]. TAIC [SFB13]. TAIC-PART [SFB13]. Tailorable [AKDN90, Mej03]. Tailored [HC98]. taint [ST19]. Tale [Hum88, MM90, TBPK20]. Talents [RGN^{+}14]. Tales [Bar82a]. TALK [Kin71]. Taming [DBvG^{+}23, MLC02, Mar84a]. Tape [Dri93]. target [BDM16]. target-sensitive [BDM16]. targeted [LS16]. targeting [BC17, Han04, SFC^{+}21]. Task [LF74, REMC81, WL81a, AA23, Bar80c, BGSG20, CALL18, Čuk16, DFQ^{+}22, FSRF^{+}22, JSPP21, LLWB14, MUE23, PKK12, UN19, ZWML14, ZB18]. task-based [FSRF^{+}22, PKK12]. task-parallel [UN19]. Tasking [JDJ^{+}06, KRTW81, RMC97, AO88, DHGR92]. Tasks [Bla92, BS19, DHWZ14, KLY20, MW13, SM02]. Tassel [Bar76c]. taxi [LYX^{+}17, XDZ^{+}17]. Taxonomy [MF18, GKS^{+}22, GB14, HSRB24, KBM02, Rei99, RB19, SYB04, SvGB05, YB06]. TBFLP [Dew86]. Tcl [Lib97b, PD00]. Tcl/Tk [Lib97b, PD00]. Tcl/Tk-based [Lib97b]. TCP [DJM97]. Teaching [CM83, CM85, Fox78, Gob71, JDGCGA12, TMS18]. Team [RM91, NN24, SB22]. Teams [MG13]. teamwork [OEA05]. Tears [Bro79]. technical [ACAB22, Bas00, KHH^{+}15, LQL^{+}22]. Technique [AHS86, CCC96, CS82, Cow87, Dun93, Ell79b, Fje79, Han79b, Hol88, HC93, Lar90, Man88, OW89, Pfe84, SHC74, Str81, Tur79, AML20, AL21, AWNS18, BB75, CPCL10, Dod82, Duc11, HC87a, IBA^{+}21, JH03, KLY20, LP83, LLN16, MM82, NAU^{+}21, NZL19, SSV^{+}20, SM24, SW14, SLJ^{+}18, Vis76, XLLY19, LZLL22]. Techniques [BG93, CT92, CM83, Chv79, Clo85, DW73, EM90, ELRV93, Gon87, HHK90, Kli81, LN71, Lan75, Lau79, LV73, McK89, PJ76, Pyl72, RB91, Sch76a, SJKL94, TWL94, VZ98, Wha93, ARA18, AH12, BAJMT21, Bar73d, Bar74d, BT21a, BM01, BUT14, CFL^{+}98, CJTK22, DAC^{+}21, DHA11, DNL^{+}20, FO10, For72, GKWS11, GDGB17, HGK^{+}19, HZ95, Kan18, KJVS21, KSST22, LSZ16, LZ10, MKA^{+}22, MA01, MRZ15, RBL^{+}14a, RVS^{+}20, SHIS99, SvGB05, TKF09, WBN^{+}20, SFB13, Hop73]. technological [Nic72]. Technologies [Ano13, PL14, BBL02, CZL21, DGR^{+}06, Haf13, YOH15]. Technology [Pow95, THG17, AA23, BMR03, CHC^{+}17, DBOT23, DFST08, FR09, LHFL07, NBOS99, NR04, RC10, TS02, VR06, YCY03, Ano09]. technology-independent [FR09]. telecontrol [CP07]. telematics [HYT13]. Telephone [CW82b, Har71b, HJC00]. telephone-accessed [HJC00]. telephony [KRZ02]. Teletext [WL81b]. Teletype [JP74]. television [MA01]. Template [RS86, DKS08, JHKS19, LHB18, Rin07]. Templates [HS85, BY17, NS01a]. templating [LHB18]. Temple [Mer74]. Temporal [CCPR91, CCvKH95, HSD10, Lam20, GMGB22, LLLY18, RD14, SB13]. tenant [BMC^{+}24]. Term [MS96, DWL^{+}17, SOKE23, WBN^{+}20]. Terminal [ACG78, HRW73, PZ92, Thi87, Coh74, MH05]. Terminals [CF80, WR77, CGL76, Bul72b]. Termination [Dro85a]. terminologies [KHH^{+}15]. Terms [Bar72a, BBK^{+}12, SO21, vdBdJKO00]. terrain [Bra99]. Terry [Wal83b]. Test [ACSK23, Bat74, CW82b, Har71a, HS89, LKL95, MGW82, WHLM98, ACKJ22, Ano88c, BLLP04, CCPY12, CCR19, CD84, DTJ89, FCA12, GQ15, HLGSW11, HTX^{+}22, KSK15, LXY^{+}11, MW13, Man01, OOG19, OJP99, PM17, RMM19, SATV22, TCMM00, WH06, ZC02, ZJY^{+}15]. test-a-few [CCPY12]. test-data [TCMM00]. test/vaccine [ACKJ22]. testability [BLS03]. Testbed [SCR94, CBR10, JGB15, MVOD19, RR05, SJA^{+}04]. tester [CS04]. Testing [ASEF23, AW96, CCRD^{+}80, HW88, Ham95, Han73, HS97, HS89, How78, HHL84, KO91, Lib97a, OPTZ96, Pro92, RS87, SFB13, Spa90, Tay83, WPT95, WW91, WJ76, AA19, AWNS18, BELS14, CCPY12, DHS02, DHG^{+}19, GKBK16, GMDM17, GMGDMB19, HL79, Han78c, HMN11, HCG^{+}16, JTG^{+}11, KD13, LXY^{+}11, LKC12, MK01, MSC23, MDH^{+}13, MGL19, MM01, NNW13, PDPMM17, SDKS16, She07, aSZP^{+}16, UT19, VDMW06, WP00, WSN^{+}23, ZC02, ZCO13, Bar76c, Rop88a]. tests [FL02, GSPA^{+}11, SJA^{+}11, ZPSH21]. Text [AMR90, BF80, Bou71, Coh98, Dav82, De 96, Fen98, Fra82, FK90, GW85, Haz74, Haz80, Lev82b, MP81, Mac77b, Mof89, MK96, MNM79, NMRW98, Noo83, Pik87, Sco81, TT82, VZ98, WLL98, BFJ^{+}11, BFNP08, CK15, Fra79, GRS74, Gu05, Ier09, KD13, Kha86, MRZ15, NT05, NHTT08, PT00a, Sne78, WZH01, ZM95, dKM04]. Text-editing [Lev82b]. Text-management [AMR90]. textbook [Val76b]. Texts [SW87]. textual [KHH^{+}15]. TGMS [DNSG89]. Theatrical [Thi93]. Theatrical-set [Thi93]. Their [Gon87, ASA22, ALG23, ELRV93, IH01, LPT78, MHN18, MBBS21, NS23, SPR^{+}19, SSD11]. them [CW01, Wil74b]. Then’s [Gre80]. Theodore [Tho74]. theoretic [MVV12, SSGA20]. Theoretical [Hos98, Sim83]. Theory [BW95, Sch82, Sha72, Woo84, JKH22]. thermal [LCT^{+}21, WCT19]. thermal-aware [WCT19]. Thesaurus [LCW98]. Thesen [Dav78]. thin [GHC^{+}07]. thin-client [GHC^{+}07]. Thing [IBA^{+}21]. Things [RWJ^{+}17, SWBS17, CKCG23, KDA20, SKD^{+}22, ALD^{+}23, BJ21, BSM^{+}21, DMR^{+}22, GDGB17, JSPP21, JAA^{+}20, JGSG^{+}21, KSST22, LCT^{+}21, Lv21, MTM22, NSM^{+}22, NRUP21, PHB21, RS21, SGAS21, SRCP19, SZSB19, VAP^{+}17, VSID17, WBPR20, XDZ^{+}17, ZKW21]. Things-enabled [MTM22]. ThingsMigrate [JGSG^{+}21]. third [GMNR20, Rob72]. Thomas [Bar79a, Bul72a, Haz72, Jac71]. Thomson [Pra96a, Pra96b]. Thought [Tra79a, Gal79]. Thoughts [Wic77]. thousand [KV14]. thrashing [JZ02]. Thread [KBH^{+}03, LS97, MR96, BHK^{+}04, CY01a, CY01b, GXN10, ZLG08]. Thread- [LS97]. thread-based [ZLG08]. thread-level [GXN10]. Threaded [IC85, PBW78, GCRD04, RGK99]. Threads [MR96, BS00]. Threat [BGS^{+}13, LW19, MDH^{+}13]. Threat-oriented [BGS^{+}13]. Threats [HLR20]. Three [BM03, CK86, DW90, KS84, MTT83, MM90, RDC93, RN00, Rog23, WW89, de 82, KSK15, LLJ12, SZ20, ACF13]. Three-Dimensional [MTT83, DW90, LLJ12]. Three-Layer [ACF13]. Three-tier [BM03, KSK15]. three-way [SZ20]. thresholds [KHOY16]. Throughput [SNM80, ROFGFR^{+}16, ROFGFRM16]. Throw [Bro76, Rob83a]. Throw-away [Bro76, Rob83a]. Tice [PPA20]. TICL [MK90]. tidy [vdP14]. tier [ASC^{+}01, BM03, KSK15]. Time [Bla92, Cel82, EMVW83, Fon85, Fra75, FH91b, Hal86, Han76a, Har80a, HHL84, Kow81, Lio79, MF18, Nil88, Ono93b, PJ75, QSA88, QSA90, RS94, Rei72, SF85, Sno91, TB73, TH86, WB85a, Wit83, Yuv75, AIB02, Ano71d, Ano72b, ABRW94, BVGVEA11, BVGVEA13, BS74, BA79, BSDF20, BJL06, Bud85, Buh93, Bul73, BL83, BW95, BDM16, BMAV05, CS91a, CMT17, CC84, CBB20, CC01, CC77, Cor84, CALL18, Dan82, DHS01, DHWZ14, DR92, DKM11, EKM^{+}99, FDN^{+}18, FM78, FFRFS19, FPAF18, Gla82, GWA91, GKLM79, Heh76, HK84a, Hol83, HKM^{+}09, HLFS05, HBC15, JLBR22, Joh79, Jor90, KLLK98, KRB21, KSBL22, KW90, Kil19, KQZ^{+}11, LF82, LYM04, LLK04, LMK16, LY92, LS15, LCGS17, LHC97, LF90, MA00, MRR^{+}08, MDWD01, NLA15, Obe11, Orm77, PLL^{+}02, PPA20, Pur76, RA87, Ric76]. time [RBS14, REMC81, Ros71, SIC^{+}20, SDB^{+}22, SBS20, SLRS06, SSP11, SGH93, SPPH10, SM85, SJP^{+}09, SBFF24, SAA^{+}20, Ste92, Str77, SSK^{+}17, TRO17, VvK99, VC02, VBS22, WM20, Wan82, WC87, WB85b, WR22, ZYW^{+}22, ZQWL24, vdP14, SSP11, TL98, Rog71]. Time-aware [MF18]. Time-Estelle [TL98]. time-sensitive [CALL18, VBS22]. time-series [SIC^{+}20, SDB^{+}22]. time-share [BA79]. Time-Shared [EMVW83, Har80a, Bul73]. Time-Sharing [Fon85, Rei72, Lio79, GKLM79]. time-triggered [SSP11]. Timed [ZLWG11]. timely [RGV14]. Timer [CV98]. timers [GRR06]. Timesharing [Hun81, Lin79, NS74]. Timestamp [DS94, dSMH13]. timestamp-based [dSMH13]. Timetabling [Kra97, Mon96b]. Timing [CBB20, Kar76, KV98, KAYH^{+}99, LY92, dOdO16, WC08]. Timings [WW89]. TinyVM [HPK^{+}12]. titan [Hen79, Lan71]. Tizzard [Mar88]. Tk [PD00]. Tk-based [Lib97b]. TLA [NN24]. TLB [QM13]. TLex [Kea91b]. Tm [vR92]. TMO [LLK04]. TMO-structured [LLK04]. TMS [AMR90]. TOC [Ano16q, Ano16n, Ano16o, Ano16p]. Together [Lib93]. toggles [PdSCJM22]. Token [Cel82, SK96, WC87, AH01]. Token-by-token [SK96]. tokens [MGP03]. tolerance [GBG^{+}14, JKH22, JSC^{+}10, MKM^{+}17, Pla97]. Tolerant [BTM81, Wha72, APS^{+}11, CD94, EKM^{+}99, dSMH13, NMMS02, PRA^{+}06, RPCS08, SMR93, WWB03, Web87]. Tom [Rop88a]. Tonge [Bar77b]. Tool [AL82, AP95, Bai73, BBC91, BA86, Bha88, BS98, Cav83b, CW94, DJM97, Dew84, FL75a, Fin97, GF11, Gri82, GB87, Hać84, HW88, HUS^{+}91, Har80c, Hel95, Hua87, Inc83, JG89, KLLK98, KS01a, LDG^{+}96, MGW82, PW93, QSA90, RDLK90, RÁdMRGAM19, Ste84, VSB86, WW91, WI85, ZH91, AAB^{+}21, AÇCT22, ARCN^{+}06, BDSV99, BCF00, BRMO97, BSC^{+}05, BCRF22, Cer18, DPH16, DdB15, DIS99, EBFK10, FMA02, FSS99, FPAF18, GRA14, GVG^{+}18, HRS^{+}09, Har99, Ier09, Inc85, IAA^{+}21, JSRM18, Kim15, KCP22, LC12, LWZ^{+}21, MSB20, MP13, DCEG23, MRZ15, MM06, Mił10, MM01, Nav01, NT05, PLL^{+}02, RGK99, SCF^{+}17, SPHB11, TGPS08, WCsH16, XWHX21, ZLZ^{+}23, BCRF22]. Tool-Supportable [Hua87]. toolchain [SMNB21]. Toolkit [BP97, Cor99b, CDGP93, FL92, KRO93, WRR97, YSM95, ABL08, AO12, CRB^{+}11, CV08, GDRV20, GDGB17, HMRZ20, KBBS05, NUB24, RMMH^{+}23, ROFGFR^{+}16, Wai02, Wai07, WC04, KHGSS12, Cor99a]. toolkits [Kot01]. Toolpack [BH87]. Toolpack/1 [BH87]. Tools [CM83, CW92, CNG^{+}83, CT90, CZA83, GAF^{+}09, Ham95, HJ88a, HPC^{+}96, HMPT89, KR85, KS89, Lan90, PMY97, Sat72, Sno78b, TM95, UGBW91, ARCN^{+}06, ACSK23, AYdS^{+}06, BAJMT21, BN13, DFST08, DM15, GK14, GHC^{+}07, HCG^{+}16, KHMB17, MA01, PVAHRG^{+}15, PDPMM17, RBL^{+}14a, RVS^{+}20, SM02, Spi02, SYB04, TBT23, dCGG13, vDD11, EMVW83]. toolset [AGRS11, GKS^{+}11, RCMZ13]. Top [BA98, Fra93, Inc83, Lei84, Set79]. Top-Down [Lei84, Inc83, Set79]. TOPI [BN13, GK14]. Topic [Cox85]. topics [SCD^{+}23]. Torii [GC20]. TOSCA [BSNB20, BRS18, SMNB21, YSBL22]. TOSI [ARV77]. TosKer [BRS18]. Tou [Rob72]. touch [RBS14]. Tour [Han94a]. TPDL* [CCPR91]. TPF [JZ02]. TPTS [LJL^{+}10]. Trace [BL78, BL79, EL96, KM94, Kou87, Sch80, TS91, CZL24, DD18, DC15, HMRZ20, KSK09, LJL^{+}10, LYX^{+}17, MMM18, MC02, RRFD23, SD18]. trace-driven [HMRZ20, LJL^{+}10, RRFD23]. Traceability [LS96a, ACCD01, BAPC23, FLL^{+}22, KH18]. TraceAnalyzer [DHMS11]. TraceChain [FLL^{+}22]. tracepoints [HCDB19]. traces [CdA12, DD21, DFD23, DHMS11, RD14]. Tracing [Lar90, Mal83, MK04, MS96, PR77, CDF^{+}24, DD10, NJGG12a, NJGG12b, NJG14, TEGF08, WKJ15]. tracking [CDM^{+}16, GJ22, LM15, LCGS17, ST19]. Trade [PLR85, LPF^{+}11, PV21, RJ09, SXWL17]. trade-off [LPF^{+}11, SXWL17]. Trade-offs [PLR85, PV21, RJ09]. tradeoff [HLLZ21]. Tradeoffs [PCBE96, BGM17]. trading [IAPC17, KCYY12, YZW^{+}12]. Tradition [MR92]. traditional [SSCdA^{+}03]. Traffic [MPN^{+}95, WIS^{+}97, FB22, LFHL22, Wai07, XDZ^{+}17, ZWX^{+}21]. trailing [Fra74]. train [EKM^{+}99, KS24]. Training [RSW^{+}23, Sef97]. trains [CL19]. trajectories [DWL^{+}17]. Transaction [CD94, HL03, Kru82, LS22]. transactional [KSBW18, PDPQ22]. transactions [HLR^{+}03, Spi09]. transceiver [SSM11]. Transcoding [CL23, LM22, ÜY22]. Transcripta [Bar73e, Bar75e]. Transducers [Pyl84]. Transfer [CW82c, GJ93, JB84, TD94, ABC^{+}21, DBOT23, DFST08, GWX^{+}23, GCZ^{+}24, KHS^{+}20]. Transferability [LM81a]. Transfers [Mer73]. Transform [WR79]. Transformation [Abb89, HI85, LTV96, BDLM04, BRL^{+}15, CRGIP15, DGPT14, HAM18, ISUG06, JAkM^{+}21, Kim15, NT20, RL23, aSZP^{+}16, TSMGD^{+}11, Wu01, Wu02, ZHZ17]. transformation-based [aSZP^{+}16]. Transformations [BH94, CAFH94, BS99a, CA14, CPP12, LGZ^{+}08, PMP^{+}16, UWW^{+}05, Yi12]. Transformer [GDH13, ZGGL23]. transformer-based [ZGGL23]. Transforms [LS75]. Transition [AB95, HL91, RS93a, Wu00, CSD23, Fos89, HBK20, WP00]. transitive [Al 13]. Translating [BM97, GYCL16, SC94, EC13]. Translation [ADM96, BAP87, BW96, CST75, KHH^{+}15, KR83, EF13, FHL^{+}18, WS99]. Translator [Fre81, Joh79, Lev97, Mar83, FHL^{+}18, GYCL16, JB07, Lev95]. transmission [BVGVEA11, ZZT^{+}21]. Transmissions [HKW77]. transmit [Coh74]. transparency [BFRFB22, KBH^{+}03]. Transparent [DO91, NS01b, CSMML12, GFS^{+}05, NMMS02, PDPQ22]. transparently [SSO13]. Transport [GM85b, LB81, vdBT77]. Transportable [BT75, HH80, Lin86]. Transportation [QC83, Sno78a]. Transporting [Hay87, Pow79]. Transputer [dCV88].

Traps [WBS82]. travel [OHD22]. Traversal [Kil81, SHF16]. Travis [PCBR18]. Treatment [Wai85]. Tree [ARV77, And91, BG93, CK97, Lic77, PB87, BST10, MA00, PPN^{+}23, PSTV10]. Trees [AW93, AN95, Blo93, DS86a, DS88, DCW93, FP82, IC85, Kil81, Vau80, Wal80, Wal90, Wil84a, ASTW03, BJL06, CLP^{+}09, GKS03, Kur99, LM07, MV20, WZH01, ZWAL22, vdP14]. Trends [Bar78d, Bar82a, Bar82c, Bar84b, Bar84a, AH12, CKCG23, CSZO23, SB22]. Trials [KV98]. trickle [Rai84]. trickle-down [Rai84]. TridentFS [HC16]. Trie [AMS92, MIA94, CGZ^{+}20, KEL^{+}21, Ris05]. trie-based [KEL^{+}21]. Tries [Dun91]. trigger [LCC14, LAG00]. triggered [SSP11]. Triggers [GL97]. Trigonometric [Sew82]. Trilateration [NAU^{+}21]. Trilateration-based [NAU^{+}21]. TRINI [PDPM^{+}16]. Trio [HF80]. Triplex [CM82]. TRIPOS [RAB^{+}79]. Trojans [CWD08]. Trondheim [Val77a]. truly [KA22]. trust [BMY03, FP15, GMNR20, ZYYC12]. trust-based [ZYYC12]. Trusted [TWNH12, BL15, GMNR20, SWY^{+}21]. trustworthy [YHGY06]. TTEthernet [Ker17]. Tui [SH98]. tumor [CP22, MBO97]. Tune [CGR00, RGK99]. tuned [BT07]. Tuning [GT92, HHPSS19, Rai92, Smi94, SKI08, YL95, CSMML12, MNEM21, RGK99, SAC06, SSS^{+}02]. tuple [DO99]. tuple-based [DO99]. Tuplespace [FP97]. turbine [JFZ^{+}21]. Turing [AP91]. turning [LBGA^{+}21]. Turnround [Lar78, New82]. Turski [Val79]. Tutorial [Pla97, PD05]. tutorials [RSW^{+}23]. tutoring [BB99b]. TV [BFPAGS^{+}08]. tweets [PPR^{+}21]. twin [GMGL24]. twins [SAYM24]. Two [Bri84, CL81, CMR92, ELRV93, GW85, Hum88, IMBB20, Jar75, LJL^{+}10, LKBT92, LC07, Mös88, Rai73, Rea73, Ten82, Yan91, Yas94, Atk82a, Bar74h, Ber82, dSMH13, MMOD16, MCHN05, SJP^{+}09, SKI08]. Two-Level [GW85]. Two-pass [Mös88]. Two-phase [LJL^{+}10, LC07, dSMH13]. two-state [Atk82a]. Type [APS95, BR95, GF80, HFPB98, MK90, Pyl84, Set81, Ten78, Vo97, Wal81c, AM00, CS15, FDD20, IASC16, KW09, PT17, Par85b, SIN95, SHF16, Sha77]. type-ambiguity [Par85b]. type-aware [SHF16]. type-basedalias [IASC16]. Type-converters [Pyl84]. Type-Safe [HFPB98]. Typed [Pow87, Pra80]. Types [AD87, BCHR81, Fle82, Ian90, Jal87, MTT81, Wal81c, Cai99, Geh85, HM12, HE82, LMPR07, NSM16, Ren23, VB01, vK87]. Typesetting [BK86a, Day83, Fox87, Ker82a, Lan76]. typing [GOQ16].

U [McD71]. U.S [Atk79a]. U.S.A [Bar78d, Bar82a, Bar82c, Bar84b]. UA [GNSP12]. UbiCrawler [BCSV04]. ubiquitous [HLW08, YHGY06]. UCSD [PV84]. UI [AO12]. UIAP [HLW08]. UIMS [RS91]. UKI [PT17]. Ullman [AS87]. UML [BBB^{+}11, CGH08, CmJHL18, DE16, HRS^{+}09, Hsu12, KAS^{+}16, ML08, PLR13, aSZP^{+}16, Cor99b, SW14, Cor99a]. UML/ [SW14]. Umple [FBLS12]. un- [FBV22]. unavailability [Eba18]. Unbounded [FW78]. Uncooperative [BW88, dSMLP22]. Undefined [BPM93, KW90]. Understandable [Pag84]. Understanding [AW04, EM12, FL94, LvDDM06, MK96, SDDD10, VS88, Rob81]. underwater [LZWX22]. Undo [Dan90]. Unearthing [SSS22]. Unguarded [Fis84]. Unicode [Chi17, CL23, LM22, NK07, Wu00]. Unification [Nør91, MAT94a]. Unified [Sch82, BDL09, GMC^{+}21, HRS^{+}09]. Uniform [LS76, Set81, ALG23]. Unifying [GHBH05]. Union [BL15]. unions [KL16]. UniPDM [Kim02]. uniprocessor [KGL06].

Uniprocessors [MDP96]. Unique [Boy01]. UNISEX [KE85]. Unit [MBBS21, WH97, KPU04, Loe07, SJA^{+}11]. unite [BMR82]. United [Lob85]. units [Bar15, CM08, Deo10, Geh85, Pet01, RGN^{+}14]. Univers [BPY90]. Universal [BHL73, HW78, Bar78d, IIL17, PT17, SAC06]. University [Atk78, Bar73a, Bar73d, Bar74f, Bar80d, Bar81, Bis81b, Bis84, Eve73, Gar86, Han78a, Han78b, Hun72, Liv75, Lon88, Mad82, Ree78, Sha83, Tho77, AC80b, Bai85c, FWS74, KDP83]. UNIX [Sau88, Jac84, Ree84b, AS97b, Any85, AM86b, Bad98, Bai85a, BS80, Bis87, BMS83, BMM85, Bre86, BBM84, BS90c, Car86, CE97, Coo85, DF95, Har80a, Hes91, HM90, Hug88, KDP83, KE85, KM79, LA90, Lio79, Lob85, McD87, MR92, MMS86, Yoo96, Col82, Cro87, Fin97, FSS99, GPR^{+}98, GMC00, HJ88b, Lan90, PW93, PR90, PSA87, Spi02]. UNIX-based [KE85]. UNIXes [BMR82]. Unix(R) [KK90]. unlimited [Ham81]. unloved [BDD09]. unmanned [Ano23m, SGA21]. unnamed [JPL03]. unobtrusive [MKW^{+}22]. unpacking [WL72]. Unparsing [Ram98]. Unrelated [BGSG20]. Unrolling [DH79]. unsafe [Win02]. unsatisfiable [SW14]. Unscrambling [Fin88]. unsorted [Har81]. unstructured [ZLZ^{+}23]. unstructured-grid [ZLZ^{+}23]. Unsupervised [HPZ^{+}20]. Untangling [ASTW03]. UnThemida [SLJ^{+}18]. untyped [Sav11]. Unusual [Rai73]. Unwin [Ano73a]. Upas [Pre85]. Update [Dan90, Dun93, FCG83, FZ98, BGP17, ZLX^{+}24]. Updates [Hos98, MVTH14, MPJ20, PKC^{+}13]. Updating [BTZ94, Lun86, MM86]. upgrade [CHCC07]. upgrading [AV05]. UPnP [HLW08]. upon [CW91]. Uppaal [BDL^{+}11]. Upper [PK89]. urban [DDB^{+}18b, LZD20, Wai07]. URLs [NL24].

USA [Bar84a, Pet77, USE85]. Usability [CKRC20, RK15a, VC21]. Usage [Cro91, WPT95, ALD^{+}23, AHH15, Hor21, LBP^{+}13, PDPMM17, SOKE23, TK09]. Use [BH87, CL19, CV84, GS90, Kou87, LP78, Nee77c, Orm77, Özc98, PJ76, Rey87, Ric76, RCC91, Sti78, WB78, Wil80, Wir77d, WW83, WS74, dSC16, BMY03, BLS03, Bri82, DHA11, DRH^{+}23, GMPL11, GRVC23, JKW74, KAZ13, KS87, Kul74, LM07, MMOD16, MPN^{+}95, NNW13, PD00, RdOTF14, Sha77, UFS99, XP23, dlRNGCC23]. Used [Inn77, CK15, LN71, TKF09]. Useful [Ell79b, KL86]. Usefulness [ZZL24]. User [AS73, BS89, BT76, CW80, FSR11, Fis82, GB87, GWZ^{+}20, Ham74, HUS^{+}91, HM90, Jaa95b, KV98, LDG^{+}96, LD95, Lop89, Mat83b, Özc98, Plu77, PH84, Pow87, Pyl80, SMFBB93, Sne78, SWPS89, Spi09, SY86, Str83b, UGBW91, Wal81c, WL81b, WG92b, vMC77, BB75, BS93, dCCCdAC20, CYW^{+}15, CW01, CRGIP15, Dew93, FT79b, GRR06, HR06, KY05, KBBS05, KYF^{+}22, KPJ^{+}17, KRO93, Kru82, LLLY19, MHN18, MKD^{+}22, Mej03, NM19, NAGL10, QH21, RSW^{+}23, SFC^{+}21, SSKG22, WWCW19, WKS^{+}98, WAH^{+}12, WCS^{+}17, YZYL07, ZLZ^{+}19, BM98]. User-adaptable [BS89]. user-centric [WAH^{+}12]. User-defined [Fis82, Pyl80, Wal81c, SFC^{+}21]. user-demand [WCS^{+}17]. User-interface [KV98, Sne78, KBBS05]. User-level [Spi09, GRR06, YZYL07]. User-Like [Ham74]. User-Oriented [BT76]. user-space [NAGL10]. Users [Bar75a, Law78, NL76, TS81, Hug77, LYLY20, LWW^{+}23, PCBR18, The77]. userspace [DD21]. Uses [ACG78, Pal82, Tho77]. Using [AG95, Bai85a, BJP^{+}00, BCL^{+}94, Bis84, Bru84, CL09, CG96, CMH91, CT92, CS97, CLC99, CCvKH95, Cla89, CH90, CK78, DW89, DJM97, FIL86, FZ98, GM85a, GW96, GAQO23, GJ93, HHR93, HUS^{+}91, HIR06, HK84b, HA90, HP11, HS89, HT86, HW94, JI21, Jac85, Kat83a, KS84, Kil12, KG95a, Knu11, Lai95, LS76, Lea77, LL91, Lev97, LCA09, Lic77, LES95, MC02, Moh77, MDP96, Ols90, OM96, PDC^{+}98, PPR02, Pow79, RGSGHGCG21, ROFGFRM16, ST04, SCGP92, SS19, SK96, Thi89, TL98, TA91, UGBW91, VBH^{+}98, WW91, WR79, Wor83, ZWAL22, ZXT^{+}17, ZRX^{+}99, ABC^{+}21, APS^{+}11, AA20, AA21, ACSK23, AM10, ALF01, AA23, AWNS18, ACM^{+}15, ASA^{+}21, Atk79a, AG06, BJ21, BHvR05, BBB^{+}22, BFGS05, BGS^{+}13, BB10, BLM00, BGSG20, Bis90, BRL^{+}15, BW96]. using [BR97, Bur16, CPZ02, CMCL03, CCC^{+}16, Car79, CCRD^{+}80, CP22, CCG14, CGR00, CIM24, DGM^{+}22, DW73, DDDF17, DS99, DBvG^{+}23, DNL^{+}20, ET07, EF13, EYG22, EMRK20, FB22, FFD96, FSC^{+}21, GK08, GHM^{+}06, GKLM79, GJ22, GDW^{+}20, Har81, HGW94, HBM06, HTJNL19, HGWBS75, HC99, HYT13, HLH15, HC20, HBJ05, HBC15, IAA^{+}21, ISUG06, JEG99, JP22, JWTG11, JH03, dSJCM16, KY05, Kar21, KHS^{+}20, Kha86, KST94, KR83, KBPM^{+}20, Kra97, KPK^{+}18, KA87, KSK15, KSDG24, KHMB17, LC05, LBC^{+}11, Les72, LER17, LFH^{+}22, LCZ08, LHFL07, LT90, Liu03, LTL^{+}03, LMPR07, Mac79, MVOD19, MWB95, MS99, MK01, MLC02, Mej03, MS18, MAA^{+}22, MTPC14, MDB19, NSM^{+}22, NNLR17, NNR18, NAGL10, NMG11, NK07, NZL19, NN24, NSW77, NR04, OEA05, PM17, PPR^{+}21, hPmKgH15, PLR18, Pat94, PCdGPP12]. using [PM18, Pol01, PES^{+}20, PPA20, RM75, RS93a, RRK^{+}18, Rin07, RGS^{+}20a, RC10, Rön07, RS21, RD14, SGAS21, SB21, SSV^{+}20, SM24, San17, SJK^{+}21, SCD^{+}23, SK03, SE11, SBG^{+}05, She07, SKD^{+}22, SIK^{+}16, SBF19, SS23, SAA^{+}20, SYXZ14, aSZP^{+}16, Sur13, TC03, TRGA18, TMS18, TC07, UDS^{+}07, UT19, VR06, VP05, WM20, WY18b, Wij05, Wis74, Wri94, XDZ^{+}17, YSM95, YCY03, YAFA19, YLR23, YSBL22, ZZKA17, Zdu07, ZB18, vDV04, vdMF13, BBH^{+}22, Lav78]. UTF [KL21]. UTF-8 [KL21]. utilities [SRC^{+}18b]. Utility [YF91, vdBT77, YB06, Yuv77c]. utility-driven [YB06]. utilization [ADZF21, MAJ15]. Utilizing [DSD^{+}05, CSM^{+}16]. Utopia [ZZWD93].

V [Her84, Mil72]. vaccine [ACKJ22]. valence [MV95]. validate [KKS10]. Validating [CS91a, CD82, DS12, KL21, MW13, Tur06, BZM^{+}17]. Validation [CG96, Gom78, KSK09, OF76, Özc98, Sto05, TP92, WS74, BCL13, BL15, CNRB13, GN02, MVOD19, DCEG23, RMSMML^{+}11, YFC06, Ree84a]. Valuating [KYF^{+}22]. Value [BCS97, LFHL22, LZK^{+}23]. valued [GKBK16]. Values [DW91, CFC15]. VAnDroid [NZL19]. VArg [KYF^{+}22]. Variability [San17, BDQR23, JKB04, MH05, SvGB05]. Variable [Cow87, Han76e, HI85, New86, Pra80, Rob79, Fen02, GKBK16, Li18, MT84a, XWHX21]. variable-length [MT84a]. Variables [BPM93, CV84, Er85, Fis83, KW90, Rav82, Boy01, CLSE05, TMM82]. variant [Win02]. Variants [Fra80, ALG23]. variate [Lar09]. variation [Loe07, Wij05]. variations [SJK^{+}21]. various [LKC12, SSRAH15, Wet77]. Vartalaap [LRMM93]. Vasarhelyi [Tho74]. VAX [PV84]. VAX11 [HJ88b]. VAX11/750 [HJ88b]. VBR [MNH04]. Vc [KL12]. VCluster [ZLG08]. VCR [MA01]. VCR-style [MA01]. VDM [BM97, Jac85]. Vector [LHH^{+}91, RB89, SAC^{+}92, AA20, TRRK21]. vectorization [KL12, LB15]. Vectorized [WBGS22]. vegetation [ZLY18]. vehicle [GZW^{+}22, HHMMG12, SCLD21, XZX^{+}21]. vehicles [Ano23m, GJ22, SGA21, TVSG21, ZZT^{+}21]. vendor [MKW^{+}22]. Verification [AB95, BE81, CW94, CV84, Fri92, MMS90, Tay83, TL98, BDSV99, BDLM04, CCG14, CSS15, DPK12, DWL^{+}15, GARSR18, JJI22, Ler02, MMOD16, NN24, Söz15, ZHZ17, ZSRR22, dCGG13]. Verifier [Hop80a, Ryd74, RZ17]. Verifying [Fle82, Lai95]. Verjus [Whi87]. Verlag [Atk79a, Bis86, Cav83a, Mee87]. Versatile [Bai73, MP81, MEP96, Duc11]. Version [CGK89, CK97, GKM90, GJ93, McG89, Tic85, dSJCM16, CE97]. versioning [LK99]. Versions [LG99, Wie96]. Versus [Fis83, Geh90, Sta82, ALG23, Ano71e, Dun74, Sam75, SMFBB93, VED06]. Vertex [BG01]. Vertex-labeling [BG01]. Vertical [CLL91, CLC09, HSD10, SBcC07]. Very [BP98, SW86a, Smi91, Tor23]. VFP [GYCL16]. VHDL [Bha88]. via [Bis81b, DS12, GHM96, GG08, KCH07, KHOY16, KKdC^{+}23, LQL^{+}22, NJGG12a, NJGG12b, NJG14, RMZ17, SDKS16, SLRS06, SO07, Thi93, ÜY22, YHK^{+}22, ZZJZ24, IS05]. VIA-based [IS05]. viable [Deu99, HW15]. vibration [Pet77]. Video [CVV97, CGL76, BGA20, DFPT08, KSH11, MA01, Mos06, PV22, SNK21, TH01, WSL03, YAFA19, RR85]. Video-Slide [RR85]. Video-terminals [CGL76]. videogames [SN01]. VidNet [WSB96]. Vienna [Jac85]. view [LFW96, SROAdM^{+}08, SS23]. Viewdata [Pal80, WL81b]. Viewing [MB96]. Viewpoint [GS76]. views [MLC02, MP00, RL14]. VIFOR [RDLK90]. viii [Bri82, Wal83c]. VILE [AP85]. violations [IASC16, LLLY18, MBGC21, TVCB15]. Virginia [Liv75]. ViRPlay [JDGCGA12]. Virtual [AF02, Bae73, BH82, BF75, BZD17, CHCC07, CV84, DV85, DCA04, EE90, Far74, Hal86, HC97a, JDJ^{+}06, LCC97, LLW98, LHK99, PNM^{+}20, Rea73, San88, Sch83a, TT96, TY14, TA91, VS88, AK15, AHRR22a, AHRR22b, AKL^{+}09, BDQR23, BHvR05, BHMV09, CARB10, CB00a, DC15, EGKP02, FMPR02, GCARPC^{+}01, HSC21, Ibs84, IMBB20, JDGCGA12, Man18, NZH20, NNK21, NMH^{+}24, NCFCFV12, PMC05, RCA^{+}19, RGS^{+}20b, SSCdA^{+}03, SOKE23, TEBK99, TGCF08, WKJ15, WIYC20, YME05, YC16, YRJ18, YMY17, ZWKX17, ZLZ^{+}19, ZWQC22, ZKB^{+}22b, ZHO^{+}19, ZSRR22, KM13, PPBP06, SM01, VED06]. Virtual-machine-based [AF02]. Virtual-memory [LCC97]. virtuality [Nic08]. Virtualization [KLY20, AH12, Cia07, DBO^{+}18, HC12, JM08, KS20, RSLAGCLB16, SHB19, TMJ^{+}21]. Virtualized [BBK^{+}12, ARA18, OOG19, RSRCGC15, STB14]. virtualizing [SSD11]. virtues [BTS09]. virus [MV16]. VIS [VS88]. Viscom [CGL76]. visible [NAU^{+}21]. Vision [Buy21, RS95, RMC97, BIA24]. visit [Wir77a]. visitor [PH14, vDV04]. Visual [CCCZ05, HPC^{+}96, HW98, KU97, RDC89, SS93, dSC16, DGPT14, FKD14, KEL^{+}21, PSTV10, SK03, Spi02, SAA^{+}20, XLZ^{+}20]. Visualization [Tha84, VMJ97, YSM95, BMR03, FMA02, GN00, HB11, JJL17, Lar08, LJS20, LWJ^{+}21, MA00, MHM01, MBGC21, PJJM21, hPmKgH15, WRD99]. Visualizing [AR93, AP95, LAD^{+}94, MBV^{+}10, LKCW13]. Viúva [GS08]. VLIW [KPU04, NS08]. VLIW-style [KPU04]. VLSI [AP85]. VM [HC12, HW15, WCT19]. Vmalloc [Vo96]. Vmgen [EGKP02]. VMs [NTF^{+}17, ZB18]. VMware [SIK^{+}16]. VNC [RSLAGCLB16]. Vocabularies [LK93]. Vocabulary [SBS20]. Voice [LFGCGCRP14]. Vol [Fen94a, Rob72, SFS97a, SFS97c, Wil72, Cav83a, Llo82, Wal83b]. Volatile [UFR18, HC16]. Volume [Bar74c, DDB^{+}18a, Jon74, MBG19a, WY18a, PKN^{+}12, Roh77a]. volumetric [LWJ^{+}21]. volunteer [SAL16]. VOS3 [Hay87]. voting [HSC21]. VR [WIYC20].

VR-Rides [WIYC20]. VRE [ZKB^{+}22b]. vs [Ben89, Jal82, UFR18]. VSE [Ott82]. VulGraB [WHYC23]. vulnerabilities [FLPM20, LC03]. vulnerability [AL21, ACKT20, NZL19, NZTL^{+}23, WHYC23]. vulnerable [TXHL18].

W [Ano79a, Ano88c, Atk78, Atk79b, Bar75f, Bar77e, Bra75, Col77b, Hop73, Lar71, Lar75a, Liv75, Pet77, Ree73, Rog71, Roh77a, Sau88, Sha83, Val77a, Val79, Wil72, MSR^{+}07, Sha77]. WA [Bar82c, Bar84b, Bar82a]. WA98118 [Bar84a]. wait [TNGT09]. Wales [Bar73a]. Walker [Cor82]. Walkthrough [BCL^{+}94]. warehouse [FSC08]. warehouses [TS02]. Warnier [Nee77a]. Warshall [Tor23]. Warwick [FWS74]. Was [TB72]. Washington [Bar78d, Bry77, Pet77]. waste [Ano71d, Ano72b]. Watch [Ise90]. Watch-oriented [Ise90]. water [HSY^{+}20]. watermarking [MCHN05, YWT^{+}12, ZZKA17]. Watfiv [Ree75]. Watfor [Ree75]. Watson [Rog71]. Watt [Atk79b, Atk83]. wavelet [JP22]. Way [CK86, Mös88, DM07, EP05, KL24, KKS10, MN18, PLR18, SMFBB93, SZ20, VHM^{+}05]. WBSN [BFG^{+}11]. WBSNs [JI21]. WCET [LPF^{+}11]. Weakest [RaMB^{+}24]. Weakly [KRR19]. Wear [DMC17]. wearable [DMC17, PWI^{+}21, RQL^{+}20, XMTL21]. Weather [KK97]. Web [CCE^{+}21, ACKT20, BMC17, BLR^{+}17, CSS15, DDGP18, FT01, GARSR18, GLD^{+}21, wKJM18, RJ09, RBL^{+}16, SKD^{+}22, SGDA18, ÜY22, UFR18, XLLY19, ZB23, AMM10, ÁSARSG09, AW04, BH01, BELS14, BCSV04, CLZ99, CMCL03, CZ04, CGH^{+}04, DMD^{+}06, FKL^{+}13, FG08, FMNW04, HKA12, Hsu12, HL20, KH07, LS03, Lev98, LQ04, LCW07, LQ99, MRZ15, MMCF03, McN05, Mej03, MLV18, OMM15, PCC^{+}12, RCMZ13, RW04, RAN03, STH97, SJA^{+}04, SROAdM^{+}08, SRRFGC^{+}10, Sto05, SKI08, UGK^{+}14, WRR97, ZC03, ZHZ17, vdMF13]. web-based [SGDA18, CGH^{+}04, LS03, LQ99]. web-integrated [CSS15]. Web-oriented [SRRFGC^{+}10]. Web-snippet [FG08]. Web/ [McN05]. Web2.0 [HKC^{+}12]. WebDAV [WG04]. WebSphere [JDBP04]. Webster [Bar77c]. weight [BS90c, RS91]. Weinberg [Bar76d]. Weingart [Ano79a]. Weinstock [Bar77e]. Welch [Ham79]. welcome [Hor14]. Well [She92]. Well-Oiled [She92]. Wells [Han78a]. Wesley [Bar76e, Bis79b, Cam85, Cou85a, Ear77, Gru83, Jac84, Llo82, Wal83b, Wil84b]. Wetherell [Bar80e]. whale [KSDG24, AA21, MKC20]. WhatsApp [MHN18]. Wheeler [Abe07, Abe10, Deo00, Deo02, Fen02, NT20]. which [Bar78a, Bar78d, Bar82c]. while [DSD^{+}05]. Whitby [Bar81]. Whitby-Strevens [Bar81]. whiteboard [CGH08]. Whitty [Pra96a, Pra96b]. Wi [CdA12]. Wi-Fi [CdA12]. Wichmann [Hop74, Ree84a]. Wide [FL94, RAN03, WRR97, AKNJ21, BBL02, KG95a]. wide-area [BBL02]. widely [BMY03]. Widening [KHOY16]. WIDES [The77]. Widespread [Nør91, Thi12]. Wiener [Ano87a]. Wikipedia [CK15]. Wiley [And78, Ano79a, Ano87a, Ano88c, Ano88a, Bar71, Bar73b, Bar76d, Bis82, Bri82, Bul72a, Bul73, CO88, Col77b, Con77, Cor99a, Cor99b, Dav74, Edm86, Ell72, Flo79, Han72, Haz71, Her84, How76, Hut74, Jac71, Lan75, Lav77, McD71, Mil72, Nic72, Ree82, Ree84b, Ree84a, Ree73, Ree75, Rob82b, Tho74, Val76a, Wal82, Wal83c, Wis74]. Wiley-Inter-science [Val76a]. Wiley-Interscience [Dav74, Nic72]. Wilkes [Ree75]. William [Flo79]. Wilsing [Wal82]. Wilson [Bar74f, Bis79a, Bis84, Han77a, Mad82]. wind [JFZ^{+}21]. Window [AM86a, DD90, GMM90, GKM90, GH93, PZ92, SG90, Wei85, AAB^{+}21, KS98]. Window-Based [AM86a]. Window-oriented [GH93]. Windowing [Jar75, Ged14, Wei85]. Windows [KS98, YZYL07, LHFL07]. Winograd [Wal83b]. Winston [Cam85, Haz72]. Winthrop [Fin77]. WiPal [CdA12]. Wire [LT85]. Wire-wrap [LT85]. Wirebrush4SPAM [PDROFRM13]. Wireless [YMY17, ACV10, CDR13, EC13, HSY^{+}20, KAS^{+}14, LZL^{+}17, MSR^{+}07, SIC^{+}20, SDB^{+}22, SBcC07, WR22]. Wirth [Edw77, Ros74]. Wisconsin [CDKK85]. wise [Wu99, ACV10]. Within [Lei84, Wri94, Eng06, Ric00, RA95, WHH21]. Within-Word [Wri94]. Without [Yuv75, App89a, Boy01, Bro79, EP05, Fen12, KPK^{+}18, ML23, Sav07]. Witt [Han95]. WK [PMY97]. wolf [SGAS21]. wonglediff [EP05]. Wood [Ell72]. Woodward [Bar75a]. Word [BT89, Mof89, Wri94, Coo05, Has77]. Word-based [Mof89]. Words [RS93a, Sal79d, AM10]. Work [CMF^{+}98, NDKL^{+}24, PCBR18, FSR11, GH02, HvdH02, Loe07, Wil74b]. Work-from-home [NDKL^{+}24]. workbench [PD00]. workflow [CCE^{+}21, GB02, MMM18, Rog23, TLC^{+}18, VMY24]. workflow-based [MMM18]. workflows [BPP10, KSBL22, LLLY18]. Working [Val78, Lan74a, Val77a, Woo74]. Workload [Li22, NS74, SSB^{+}16, EYG22, HKA12, LLS06]. Workloads [AW96, AI22, BMC^{+}24, BIA24, HKWZ00]. Workplace [FC98]. Workshop [PL14]. workspace [TK09]. Workstation [Bov87, BH87, GW84b, Kuh90, MC91, JCL85]. Workstations [AM86a, GPR^{+}98, Lob85, Str95, RGK99]. world [BMR82, DSD^{+}05, SC14, RAN03, WRR97]. worlds [FMPR02, JDGCGA12].

Worldwide [Pra96a, Pra96b]. WORM [Qui91, RRP95]. Worst [SPPH10]. Worst-case [SPPH10]. wounds [RBD^{+}24]. WPAN [SSM11]. wrap [LT85, Wit77a]. wrapper [MCGS08]. wrappers [CKL^{+}02, JGS^{+}08]. Writable [WR84]. write [CSM^{+}16, LFHL22]. Writers [Gar96]. Writing [AKDN90, Bow73, Bro74, DS86b, HSW75, Mar83, McC90, MS80b, Oni85, Pry85, SFIK80, War80, BSC^{+}05, DM07, Lan74b, Rob81, HW77, Ree82]. Written [Hop80b, Kru82]. WSB [AMM10]. WSH [BBM84]. Wulf [Bar77e]. WWW [ALF01, ASC^{+}01, HJC00]. WWW-based [ALF01]. WYSIWIB [LBP^{+}13].

X [Dea86, DD90, GMM90, GKM90, McC90, PZ92, SG90]. X.25 [Vel88]. X11 [AKDN90, Wid90]. x86 [HCDB19, OKN04]. XA [Mar86]. XBDDs [LCA09]. xDEVS [RMMH^{+}23]. XDS [LG73]. Xerox [Gut87]. XERROR [JK83]. XGBoost [GDW^{+}20]. XHAMI [KKA^{+}17]. xii [Lon88]. xiv [Llo82]. xix [Mee87, Wal84b]. XKMS [RMSMML^{+}11]. XML [BM06, BPP10, CH06, CDM^{+}16, CFD^{+}22, FSC08, Jak04, Kar14, LW04, LTL^{+}03, NEFZ00, RT10, SGS08, ŠS08, SDC04, VR06, WK06a, YLM^{+}05]. XML-based [SDC04]. XML/ [LW04]. XMM [BMD^{+}98]. XPath [ACM^{+}15]. XPL [LG73]. XQuery [FSC08, Kil12]. XR [QC83]. XSLT [CH06, DZS09, DBH04, ISUG06]. XSLT-based [DZS09]. XSS [ST19]. Xtext [Sur13]. xvi [Ano87a]. XXL [DKS08].

YACC [BP98, Jon85, FSO91, Mer93]. YACC-Compatible [BP98]. Yale [Lev82a]. Yasukawa [Bar76d]. Year [Buy21, FRBRF19, NBOS99, SB22]. Years [BL90b, BDL^{+}11, KV14]. Yershov [Wic72a]. Yezerki [Bar74c]. yfx [Fav07]. Yima [ZSFY05]. Yoix(R) [DM07]. Yong [XZ01, XZ03]. Yong-Rae [XZ03]. York [And78, Ano79a, Bar71, Bar73b, Bar75c, Bar76a, Bar76d, Bar77e, Bar77b, Bar79a, Bra75, Bul73, Con77, Dav74, Dav78, Ell72, For72, Han72, Haz71, Hop73, How76, Hut74, Jon74, Ken77, Lan75, Lav77, McD71, Mil72, Nic72, Rob72, Rog71, Sha72, Val76a, Wil72, Wis74]. Yoshinori [Pra96a, Pra96b]. YouGen [HLGSW11]. Young [CW82a]. Yovits [Jon74]. Ytracc [FSO91].

Z [Ree84a, SC94, TM95]. Z80 [CW82c]. Z8000 [DB83]. Zaman [UDS^{+}07]. Zanshin [SM15]. ZBDDs [LCA09]. zebra [DNL^{+}20]. zebra-crossing [DNL^{+}20]. ZED [Haz80]. Zeitlinger [Flo73]. zero [CTLL07, CHCC07, MPJ20, PHB21, UWW^{+}05]. zero-loss [CTLL07, CHCC07]. Zilog [DB83]. Zimmerman [Fin77]. ZipfAllocation [KSH11]. Ziv [BK93, NT05]. zooming [BM98]. ZX [BGS18].

References

Adler:2014:SOI

[AA14] Philipp Adler and Wolfram Amme. Speculative optimizations for interpreting environments. Software—Practice and Experience, 44(10):1223–1249, October 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abadeh:2019:MDF

[AA19] Maryam Nooraei Abadeh and Shohreh Ajoudanian. A model-driven framework to enhance the consistency of logical integrity constraints: Introducing integrity regression testing. Software—Practice and Experience, 49(2):274– 300, February 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Alsghaier:2020:SFP

[AA20] Hiba Alsghaier and Mohammed Akour. Software fault prediction using particle swarm algorithm with genetic algorithm and support vector machine classifier. Software—Practice and Experience, 50(4):407– 427, April 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Alsghaier:2021:SFP

[AA21] Hiba Alsghaier and Mohammed Akour. Software fault prediction using Whale algorithm with genetics algorithm. Software—Practice and Experience, 51(5):1121–1146, May 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Arasan:2023:EET

[AA23] K. Kalai Arasan and P. Anandhakumar. Energyefficient task scheduling and resource management in a cloud environment using optimized hybrid technology. Software—Practice and Experience, 53(7): 1572–1593, July 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ahmad:2021:IDS

[AAB^{+}21] Muneer Ahmad, Iftikhar Ahmad, Muhammad Bilal, Alireza Jolfaei, and Raja Majid Mehmood. IN-SWF DNA signal analysis tool: Intelligent noise suppression window filter. Software—Practice and Experience, 51(3):670– 685, March 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Allen:1988:PGA

[AB88] Pat Allen and Alan Burns. Program generation for Ada — a case study. Software—Practice and Experience, 18(12):1125–1138, December 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abe:1989:IFS

[AB89] Kris K. Abe and Daniel M.

                                        Berry. indx and findphrases,

a system for generating indexes for ditroff documents. Software—Practice and Experience, 19(1):1– 34, January 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Arnold:1995:AVP

[AB95] A. Arnold and S. Brlek.

Automatic verification of properties in transition systems. Software—Practice and Experience, 25(6):579– 596, June 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Andrade:2020:PSC

[AB20] Rodrigo Andrade and Paulo Borba. Privacy and security constraints for code contributions. Software—Practice and Experience, 50(10):1905–1929, October 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ayadi:2020:MSN

[ABA20] Mouhamed Gaith Ayadi, Riadh Bouslimi, and Jalel Akaichi. Medical social network content analysis for medical image retrieval purpose. Software—Practice and Experience, 50(1): 65–86, January 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abbott:1978:LPG

[Abb78] Russell J. Abbott. A lookahead pattern generator for SNOBOL4. Software— Practice and Experience, 8 (5):641, September 1978.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abbott:1989:SNL

[Abb89] Russell J. Abbott. Set notation as a language to specify data transformation programs. Software —Practice and Experience, 19(6):593–606, June 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Arbab:1998:RCM

[ABBE98] F. Arbab, C. L. Blom, F. J. Burger, and C. T. H. Everaars. Reusable coordinator modules for massively concurrent applications. Software—Practice and Experience, 28(7):703– 735, July 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1805;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1805&
PLACEBO=IE.pdf.

Addyman:1979:DDP

[ABBH^{+}79] A. M. Addyman, R. Brewer, D. G. Burnett-Hall, R. M. de Morgan, W. Findlay, M. I. Jackson, D. A. Joslin, M. J. Rees, D. A. Watt, J. Welsh, and B. A. Wichmann. A draft description of Pascal. Software—Practice and Experience, 9(5): 381–424, May 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Alenezi:2022:SSS

[ABBS22] Mamdouh Alenezi, Hamid Abdul Basit, Maham Anwar Beg, and Muhammad Saad Shaukat. Synthesizing secure software development activities for linear and agile lifecycle models. Software—Practice and Experience, 52(6): 1426–1453, June 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abbasi:2021:ACU

[ABC^{+}21] Rashid Abbasi, Ali Kashif Bashir, Jianwen Chen, Abdul Mateen, Jalil Piran, Farhan Amin, and Bin Luo. Author classification using transfer learning and predicting stars in co-author networks. Software—Practice and Experience, 51(3): 645–669, March 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abel:2007:IFC

[Abe07] Jürgen Abel. Incremental frequency count — a post BWT-stage for the Burrows–Wheeler compression algorithm. Software—Practice and Experience, 37(3):247–265, March 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abel:2010:PBS

[Abe10] Jürgen Abel. Post BWT stages of the Burrows– Wheeler compression algorithm. Software—Practice and Experience, 40(9):751– 777, August 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Alongi:2022:ESO

[ABG^{+}22] Francesco Alongi, Marcello M. Bersani, Nicolò Ghielmetti, Raffaela Mirandola, and Damian A. Tamburri. Event-sourced, observable software architectures: an experience report. Software—Practice and Experience, 52(10): 2127–2151, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Appert:2008:SAS

[ABL08] C. Appert and M. Beaudouin-Lafon. SwingStates: adding state machines to Java and the Swing toolkit. Software—Practice and Experience, 38 (11):1149–1182, September ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Audsley:1994:SSH

[ABRW94] N. C. Audsley, A. Burns, M. F. Richardson, and A. J. Wellings. STRESS: a simulator for hard realtime systems. Software —Practice and Experience, 24(6):543–564, June 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Allan:1998:BRL

[ABSS98] Vicki Allan, Steven J.

Beaty, Bogong Su, and Philip H. Sweany. Building a retargetable local instruction scheduler. Software—Practice and Experience, 28(3):249–283, March 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1780;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1780&
PLACEBO=IE.pdf.

Aho:1975:ESM

[AC75] Alfred V. Aho and Margaret J. Corasick. Efficient string matching: an aid to bibliographic search. Communications of the Association for Computing Machinery, 18(6):333– 340, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Agarwal:1980:SEC

[AC80a] Rakesh K. Agarwal and Samuel T. Chanson. A space-efficient code generation scheme for BCPL. Software—Practice and Experience, 10(2):77–95, February 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Alty:1980:UCA

[AC80b] J. L. Alty and M. J. Coombs. University Computing Advisory Services: The study of the man– computer interface. Software—Practice and Experience, 10(11):919–934, November 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Akinyemi:2013:FES

[AC13] John A. Akinyemi and Charles L. A. Clarke. Fast and effective soft links. Software—Practice and Experience, 43(5):577– 593, May 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ampatzoglou:2022:SPT

[ACAB22] Apostolos Ampatzoglou, Alexander Chatzigeorgiou, Elvira Maria Arvanitou, and Stamatia Bibi. SDK4ED: a platform for technical debt management. Software—Practice and Experience, 52(8):1879–1902, August 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1983:CCM

[ACC83] Malcolm P. Atkinson, Ken J. Chisholm, and W. Paul Cockshott. CMS: a chunk management system. Software—Practice and Experience, 13(3):273– 285, March 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ambriola:1995:DSA

[ACC95] V. Ambriola, Paolo Ciancarini, and Andrea Corradini. Declarative specification of the architecture of a software development environment. Software—Practice and Experience, 25(2): 143–174, February 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Antoniol:2001:MTL

[ACCD01] G. Antoniol, G. Canfora, G. Casazza, and A. De Lucia. Maintaining traceability links during objectoriented software evolution. Software—Practice and Experience, 31(4):331– 355, April 10, 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        77004440/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=77004440&PLACEBO=IE.
                                        pdf.

Atkinson:1983:APH

[ACCM83] Malcolm P. Atkinson, Ken J. Chisholm, W. Paul Cockshott, and Richard Marshall. Algorithms for a persistent heap. Software—Practice and Experience, 13(3):259–271, March 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aksakalli:2022:MIT

[AÇCT22] Isil Karabey Aksakalli, Turgay Çelik, Ahmet Burak Can, and Bedir Tekinerdogan. Micro-IDE: a tool platform for generating efficient deployment alternatives based on microservices. Software—Practice and Experience, 52(7): 1756–1782, July 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ancona:1985:HLL

[ACDP85] M. Ancona, A. Clematis, G. Dodero, and A. Pino. A high level language based system for cross-assembler definition. Software—Practice and Experience, 15 (12):1159–1184, December 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ameller:2013:TLA

[ACF13] David Ameller, Oriol Collell, and Xavier Franch. The Three-Layer architectural pattern applied to plug-in-based architectures: the Eclipse case. Software—Practice and Experience, 43(4):391–402, April 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Adams:1978:SUE

[ACG78] J. C. Adams, W. S. Currie, and B. A. C. Gilmore. The structure and uses of the Edinburgh Remote Terminal Emulator. Software—Practice and Experience, 8(4):451–459, July/ August 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Alwasel:2021:BSA

[ACG^{+}21] Khaled Alwasel, Rodrigo N.

Calheiros, Saurabh Garg, Rajkumar Buyya, Mukaddim Pathan, Dimitrios Georgakopoulos, and Rajiv Ranjan. BigDataSDNSim: a simulator for analyzing big data applications in software-defined cloud data centers. Software—Practice and Experience, 51(5): 893–920, May 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Antonino:2022:CEI

[ACK^{+}22] Pablo Oliveira Antonino, Rafael Capilla, Rick Kazman, Thomas Kuhn, Frank Schnicke, Tagline Treichel, Adam Bachorek, Zai Müller-Zhang, and Victor Salamanca. Continuous engineering for Industry 4.0 architectures and systems. Software—Practice and Experience, 52(10): 2241–2262, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abid:2022:NBB

[ACKJ22] Amal Abid, Saoussen Cheikhrouhou, Slim Kallel, and Mohamed Jmaiel. NovidChain: Blockchainbased privacy-preserving platform for COVID-19 test/vaccine certificates. Software—Practice and Experience, 52(4):841–867, April 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abramson:2009:RDI

[ACKS09] David Abramson, Clement Chu, Donny Kurniawan, and Aaron Searle. Relative debugging in an integrated development environment. Software— Practice and Experience, 39(14):1157–1183, September 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Amankwah:2020:ECC

[ACKT20] Richard Amankwah, Jinfu Chen, Patrick Kwaku Kudjo, and Dave Towey. An empirical comparison of commercial and opensource web vulnerability scanners. Software—Practice and Experience, 50 (9):1842–1857, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Arroyuelo:2015:FMX

[ACM^{+}15] Diego Arroyuelo, Francisco Claude, Sebastian Maneth, Veli Mäkinen, Gonzalo Navarro, Kim Nguy˜ ên, Jouni Sirén, and Niko Välimäki. Fast in-memory XPath search using compressed indexes. Software—Practice and Experience, 45(3):399–434, March 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Amankwah:2023:BDJ

[ACSK23] Richard Amankwah, Jinfu Chen, Heping Song, and Patrick Kwaku Kudjo. Bug detection in Java code: an extensive evaluation of static analysis tools using Juliet Test Suites. Software—Practice and Experience, 53(5):1125–1143, May 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Amato:2010:MWD

[ACV10] Giuseppe Amato, Stefano Chessa, and Claudio Vairo. MaD-WiSe: a distributed stream management system for wireless sensor networks. Software—Practice and Experience, 40(5):431– 451, April 25, 2010. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Annevelink:1987:OOD

[AD87] Jurgen Annevelink and P. Dewilde. Objectoriented data management based on abstract data types. Software—Practice and Experience, 17(11): 757–782, November 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Addyman:1980:NAM

[Add80]  A. M. Addyman.  A

novel approach to modularity. Software—Practice and Experience, 10(8):685, August 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ancona:1984:ILM

[ADDM84] M. Ancona, Leila De Floriani, G. Dodero, and S. Mancosu. Integrating library modules into Pascal programs. Software—Practice and Experience, 14(5): 401–412, May 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ardis:2000:SPL

[ADH^{+}00] Mark Ardis, Nigel Daley, Daniel Hoffman, Harvey Siy, and David Weiss. Software product lines: a case study. Software—Practice and Experience, 30(7):825– 847, June 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72001832/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72001832&PLACEBO=IE.
pdf.

Andrews:1996:MFS

[ADM96] Kristy Andrews, Paul Del Vigna, and Mark Molloy. Macro and file structure preservation in source-to-source translation. Software—Practice and Experience, 26(3):281–292, March 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16794.

Agrawal:1993:DDS

[ADS93] Hiralal Agrawal, Richard A. DeMillo, and Eugene H. Spafford. Debugging with dynamic slicing and backtracking. Software —Practice and Experience, 23(6):589–616, June 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ai:2021:RCM

[ADZF21] Liang Ai, Yuhui Deng, Yi Zhou, and Hao Feng. RUE: a caching method for identifying and managing hot data by leveraging resource utilization efficiency. Software—Practice and Experience, 51 (11):2252–2273, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aksit:2006:EEAb

[AE06a] Mehmet Aksit and Tzilla Elrad. Editorial: Experiences with Auto-adaptive and Reconfigurable Systems (Application Level Adaptation). Software— Practice and Experience, 36(11–12):1227–1229, ???? 2006. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Aksit:2006:EEAa

[AE06b] Mehmet Aksit and Tzilla Elrad. Editorial: Experiences with Auto-adaptive and Reconfigurable Systems (System Level Adaptation). Software—Practice and Experience, 36(11– 12):1113–1114, ???? 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Allevato:2014:ECP

[AE14] Anthony Allevato and Stephen H. Edwards. Extended conference papers: Dereferee: instrumenting C++ pointers with meaningful runtime diagnostics. Software—Practice and Experience, 44(8):973– 997, August 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Austin:1976:LC

[AEH76] B. J. Austin, P. M. Ewens, and R. H. Hudson. Linking a CYBER76 to a 3600. Software—Practice and Experience, 6(4):541– 553, October/December 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ageenko:1999:FAM

[AF99] Eugene I. Ageenko and Pasi Fränti. Forwardadaptive method for contextbased compression of large binary images. Software—Practice and Experience, 29(11):943–952, September 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=63501201;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=63501201&PLACEBO=IE.
pdf.

Agbaria:2002:VMB

[AF02] Adnan Agbaria and Roy Friedman. Virtual-machine-based heterogeneous checkpointing. Software— Practice and Experience, 32(12):1175–1192, October 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
98516186/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=98516186&PLACEBO=IE.
pdf.

Ageenko:2002:CBC

[AFF02] Eugene Ageenko, Martti Forsell, and Pasi Fränti. Context-based compression of binary images in parallel. Software—Practice and Experience, 32(13):1223– 1237, November 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
98517573/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=98517573&PLACEBO=IE.
pdf.

Armstrong:2008:NCB

[AFFR08] C. W. Armstrong, R. W.

Ford, T. L. Freeman, and G. D. Riley. NOAH: a CSP-based language for describing the behaviour of coupled models. Software—Practice and Experience, 38(2):135–159, February ??, 2008. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Attardi:1998:CMM

[AFI98] Giuseppe Attardi, Tito Flagella, and Pietro Iglio. A customisable memory management framework for C++. Software—Practice and Experience, 28(11):1143–1184, September 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?

                                        ID=10007361;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=10007361&PLACEBO=IE.
                                        pdf.

Apolonia:2020:SAM

[AFNG20] Nuno Apolónia, Felix Freitag, Leandro Navarro, and Sarunas Girdzijauskas. Socially aware microcloud service overlay optimization in community networks. Software—Practice and Experience, 50(5):675– 687, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Adams:1995:UCE

[AG95] Rod Adams and Sue Gray. Using conditional execution to exploit instruction level concurrency. Software—Practice and Experience, 25(9):1003–1020, September 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:2006:EPM

[AG06] Darren C. Atkinson and William G. Griswold. Effective pattern matching of source code using abstract syntax patterns. Software—Practice and Experience, 36(4):413–447, April 10, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www.
cse.scu.edu/~atkinson/
papers/spe-06.html.

Al-Gahmi:2010:SBR

[AGC10] Abdulmalik Al-Gahmi and Jonathan Cook. A servicebased runtime environment for native applications. Software—Practice and Experience, 40(1):73– 100, January ??, 2010. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Avijit:2006:BRC

[AGG06] Kumar Avijit, Prateek Gupta, and Deepak Gupta. Binary rewriting and call interception for efficient runtime protection against buffer overflows. Software —Practice and Experience, 36(9):971–998, July 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Arora:2017:SCS

[AGM17] Ritu Arora, Sanjay Goel, and Ravi Kant Mittal. Supporting collaborative software development over GitHub. Software—Practice and Experience, 47 (10):1393–1416, October 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Araujo:2023:NPB

[AGR^{+}23] Gabriell Araujo, Dalvan Griebler, Dinei A. Rockenbach, Marco Danelutto, and Luiz G. Fernandes. NAS Parallel Benchmarks with CUDA and beyond. Software—Practice and Experience, 53(1):53– 80, January 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Arcaini:2011:MDP

[AGRS11] Paolo Arcaini, Angelo Gargantini, Elvinia Riccobene, and Patrizia Scandurra. A model-driven process for engineering a toolset for a formal method. Software —Practice and Experience, 41(2):155–166, February 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Appelbe:1985:SSP

[AH85] William F. Appelbe and Klaus Hansen. A survey of systems programming languages: Concepts and facilities. Software—Practice and Experience, 15(2):169– 190, February 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aycock:2001:ST

[AH01] John Aycock and R. Nigel Horspool. Schrödinger’s token. Software—Practice and Experience, 31(8):803– 814, July 10, 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
78502907/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=78502907&PLACEBO=IE.
pdf.

Aguiar:2012:CTF

[AH12] Alexandra Aguiar and Fabiano Hessel. Current techniques and future trends in embedded system’s virtualization. Software—Practice and Experience, 42(7):917–944, July 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ashraf:2015:MSB

[AHH15] Jamshaid Ashraf, Omar Khadeer Hussain, and Farookh Khadeer Hussain. Making sense from Big RDF Data: OUSAF for measuring ontology usage. Software— Practice and Experience, 45(8):1051–1071, August 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ahmadi:2022:CIB

[AHRR22a] Javad Ahmadi, Abolfazl Toroghi Haghighat, Amir Masoud Rahmani, and Reza Ravanmehr. Confidence interval-based overload avoidance algorithm for virtual machine placement. Software—Practice and Experience, 52(10): 2288–2311, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ahmadi:2022:FAV

[AHRR22b] Javad Ahmadi, Abolfazl Toroghi Haghighat, Amir Masoud Rahmani, and Reza Ravanmehr. A flexible approach for virtual machine selection in cloud data centers with AHP. Software—Practice and Experience, 52(5): 1216–1241, May 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Al-Hussaini:1986:YAS

[AHS86] A. M. M. Al-Hussaini and R. G. Stone. Yet another storage technique for LR parsing tables. Software —Practice and Experience, 16(4):389–401, April 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Arisawa:1980:DMR

[AI80] Makato Arisawa and Minoru Iuchi. Debugging methods in recursive structured FORTRAN. Software—Practice and Experience, 10(1):29–43, January 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ahmed:2022:TNE

[AI22] Hameeza Ahmed and Muhammad Ali Ismail. Toward a novel engine for compiler optimization space exploration of big data workloads. Software—Practice and Experience, 52(5): 1262–1293, May 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Angelov:2002:HSR

[AIB02] C. K. Angelov, I. E. Ivanov, and A. Burns. HARTEX — a safe real-time kernel for distributed computer control systems. Software—Practice and Experience, 32(3):209–232, March 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
89011425/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=89011425&PLACEBO=IE.
pdf.

Atkinson:1978:EPD

[AJ78] Malcolm P. Atkinson and M. J. Jordan. An effective program development environment for BCPL on a small computer. Software —Practice and Experience, 8(3):265–275, May/June 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Acuna:2004:APR

[AJ04] Silvia T. Acuña and Natalia Juristo. Assigning people to roles in software projects. Software—Practice and Experience, 34(7): 675–696, June 2004. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ajila:1995:SMA

[Aji95] Samuel Ajila. Software maintenance: An approach to impact analysis of objects change. Software— Practice and Experience, 25(10):1155–1181, October 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Al-Jarrah:1979:EAC

[AJT79] M. M. Al-Jarrah and I. S. Torsun. An empirical analysis of COBOL programs. Software—Practice and Experience, 9(5):341– 359, May 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkins:1983:ECB

[AK83] Stella Atkins and Brian Knight. Experiences with coroutines in BCPL. Software—Practice and Experience, 13(8):765–768, August 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abeni:2015:RRC

[AK15] Luca Abeni and Csaba Kiraly. Running repeatable and controlled virtual routing experiments. Software —Practice and Experience, 45(4):455–471, April 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Angebranndt:1990:WTS

[AKDN90] Susan Angebranndt, Philip L.

Karlton, Raymond Drewry, and Todd Newman. Writing tailorable software: The X11 sample server. Software—Practice and Experience, 20(S2):S69– S81, October 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Akbar:2024:DPM

[AKIM24] Muhammad Azeem Akbar, Arif Ali Khan, Najmul Islam, and Sajjad Mahmood. DevOps project management success factors: a decision-making framework. Software— Practice and Experience, 54(2):257–280, February 2024. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Ahn:2009:PAO

[AKL^{+}09] Sunil Ahn, Namgyu Kim, Seehoon Lee, Dukyun Nam, Soonwook Hwang, Birger Koblitz, Vincent Breton, and Sangyong Han. Performance analysis and optimization of AMGA for the large-scale virtual screening. Software—Practice and Experience, 39 (12):1055–1072, August 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Alam:2017:MFS

[AKM17] Omar Alam, Jörg Kienzle, and Gunter Mussbacher. Modelling a family of systems for crisis management with concern-oriented reuse. Software—Practice and Experience, 47(7):985– 999, July 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ahn:2021:SSW

[AKNJ21] Minwoo Ahn, Donghyun Kim, Taekeun Nam, and Jinkyu Jeong. SCOZ: a system-wide causal profiler for multicore systems. Software—Practice and Experience, 51(5): 1043–1058, May 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aleksy:2006:DIE

[AKS06] M. Aleksy, A. Korthaus, and N. Schmoigl. Design and implementation of an enhanced CORBA property service. Software—Practice and Experience, 36(10):1029–1050, August 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aho:1979:APS

[AKW79] Alfred V. Aho, Brian W.

Kernighan, and Peter J. Weinberger. Awk — a pattern scanning and processing language. Software —Practice and Experience, 9(4):267–279, April 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

AlDallal:2013:ITR

[Al 13]  Jehad Al Dallal.  In-

corporating transitive relations in low-level designbased class cohesion measurement. Software—Practice and Experience, 43(6): 685–704, June 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Akin:1982:DIC

[AL82] T. Allen Akin and Richard J. Leblanc. The design and implementation of a code generation tool. Software—Practice and Experience, 12(11):1027–1041, November 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anantha:1990:CCP

[AL90] Kasi Anantha and Fred Long. Code compaction for parallel architectures. Software—Practice and Experience, 20(6):537–554, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Akram:2021:SSQ

[AL21] Junaid Akram and Ping Luo. SQVDT: a scalable quantitative vulnerability detection technique for source code security assessment. Software—Practice and Experience, 51(2): 294–318, February 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ancilotti:1981:LMR

[ALBN81] P. Ancilotti, N. Lijtmaer, M. Boari, and A. Natali. Linguistic mechanisms for resource management strategies. Software—Practice and Experience, 11(3): 291–314, March 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Alderson:1972:BRB

[Ald72] M. R. Alderson. Book review: Information processing of medical records, J. Anderson and J. M. Forsythe (Ed.), North Holland Publishing Company, Amsterdam, 1970. No. of pages: 419+23. Price: £7.60. Software—Practice and Experience, 2(3):307– 308, July 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Andrade:2023:DIM

[ALD^{+}23] Leandro Andrade, Cleber Jorge Lira De Santana, Brenno De Mello Alencar, Claudio Silva, Jr., and Cássio Prazeres. Data interplay: a model to optimize data usage in the Internet of Things. Software—Practice and Experience, 53(6):1410–1437, June 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anido:2001:DWB

[ALF01] Luis Anido, Martin Llamas, and Manuel J. Fernández. Developing WWW-based highly interactive and collaborative applications using software components. Software—Practice and Experience, 31(9):845–867, July 25, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        78505027/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=78505027&PLACEBO=IE.
                                        pdf.

Amato:2023:SVU

[ALG23] Domenico Amato, Giosué Lo Bosco, and Raffaele Giancarlo. Standard versus uniform binary search and their variants in learned static indexing: the case of the searching on sorted data benchmarking software platform. Software —Practice and Experience, 53(2):318–346, February 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ahn:2019:EDF

[ALKL19] Sunil Ahn, Jeongcheol Lee, Jaesung Kim, and Jong-Suk R. Lee. EDISON-DATA: a flexible and extensible platform for processing and analysis of computational science data. Software—Practice and Experience, 49(10): 1509–1530, October 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Allan:1983:CP

[All83a] G. J. B. Allan. Correct programs? Software—Practice and Experience, 13(8):768– 769, August 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Allison:1983:SDP

[All83b]  Lloyd Allison.  Syntax

directed program editing. Software—Practice and Experience, 13(5):453– 465, May 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Allison:1989:CPS

[All89] Lloyd Allison. Circular programs and self-referential structures. Software— Practice and Experience, 19(2):99–109, February 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Alshehri:2022:PCN

[Als22] Yasser Ali Alshehri. Predicting change in newly created files in a software product line project. Software—Practice and Experience, 52(12):2499–2512, December 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1978:CCA

[AM78] L. V. Atkinson and J. J. McGregor. CONA: a conversational Algol system. Software—Practice and Experience, 8(6):699– 708, November/December 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Adams:1986:DWB

[AM86a] Evan Adams and Steven S. Muchnick. Dbxtool: a window-based symbolic debugger for Sun workstations. Software—Practice and Experience, 16(7):653– 669, July 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anyanwu:1986:CRU

[AM86b] J. A. Anyanwu and L. F. Marshall. A crash resistant UNIX file system. Software —Practice and Experience, 16(2):107–118, February 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aritsugi:2000:MTO

[AM00] Masayoshi Aritsugi and Akifumi Makinouchi. Multipletype objects in an enhanced C++ persistent programming language. Software—Practice and Experience, 30(2):151–174, February 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
69503458/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=69503458&PLACEBO=IE.
pdf.

Anh:2010:ICU

[AM10] Vo Ngoc Anh and Alistair Moffat. Index compression using 64-bit words. Software—Practice and Experience, 40(2):131–147, February ??, 2010. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Akram:2020:IIB

[AML20] Junaid Akram, Majid Mumtaz, and Ping Luo. IBFET: Index-based features extraction technique for scalable code clone detection at file level granularity. Software—Practice and Experience, 50(1):22– 46, January 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ammann:1977:CGP

[Amm77] Urs Ammann. On code generation in a PASCAL compiler. Software—Practice and Experience, 7(3): 391–423, May/June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Al-Masri:2010:WBC

[AMM10] Eyhab Al-Masri and Qusay H. Mahmoud. WSB: a broker-centric framework for quality-driven Web service discovery. Software —Practice and Experience, 40(10):917–941, September 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Adewumi:2019:FFO

[AMOS19] Adewole Adewumi, Sanjay Misra, Nicholas Omoregbe, and Luis Fernandez Sanz. FOSSES: Framework for open-source software evaluation and selection. Software—Practice and Experience, 49(5):780–812, May 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Amur:1990:TFF

[AMR90] Rekha Amur, K. Ananda Mohan, and M. D. Ramaswami. TMS: a freeform text-management system. Software—Practice and Experience, 20(3):321– 324, March 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aoe:1992:EIT

[AMS92] Jun-Ichi Aoe, Katsushi Morimoto, and Takashi Sato. An efficient implementation of trie structures. Software—Practice and Experience, 22(9):695– 721, September 1992. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Al-Mahfoudh:2023:ELC

[AMSG23] Mohammed S. Al-Mahfoudh, Ryan Stutsman, and Ganesh Gopalakrishnan. Efficient linearizability checking for actor-based systems. Software—Practice and Experience, 53(11):2163–2199, November 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Austin:1991:DOS

[AMW91] Paul Austin, Kevin Murray, and Andy J. Wellings. The design of an operating system for a scalable parallel computing engine. Software—Practice and Experience, 21(10):989–1013, October 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1981:CCP

[AN81] L. V. Atkinson and S. D.

North. COPAS: a conversational Pascal system. Software—Practice and Experience, 11(8):819–829, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkins:1988:ADO

[AN88] Martin C. Atkins and Lee R. Nackman. The active deallocation of objects in object-oriented systems. Software—Practice and Experience, 18 (11):1073–1089, November 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Andersson:1995:EIS

[AN95] Arne Andersson and Stefan Nilsson. Efficient implementation of suffix trees. Software—Practice and Experience, 25(2):129– 141, February 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anderson:1978:BRB

[And78] T. Anderson. Book review: Software reliability: Principles and practices, G. J. Myers, Wiley, New York, 1976. No. of pages: 360. Price: £15.8.5, $25.00. Software—Practice and Experience, 8(1):110, January 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Andrews:1982:AAA

[And82a] Gregory R. Andrews. An alternative approach to arrays. Software—Practice and Experience, 12(5):475– 485, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Andrews:1982:DPL

[And82b] Gregory R. Andrews. The distributed programming language SR: mechanisms, design and implementation. Software—Practice and Experience, 12(8):719– 753, August 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anderson:1989:HSE

[And89] Evan E. Anderson. A heuristic for software evaluation and selection. Software—Practice and Experience, 19(8):707–717, August 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Andersson:1991:NSB

[And91] Arne Andersson. A note on searching in a binary search tree. Software— Practice and Experience, 21(10):1125–1128, October 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1971:CRa

[Ano71a]  Anonymous.  Computer

recreations. Software— Practice and Experience, 1 (1):97–99, January/March 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1971:CRb

[Ano71b]  Anonymous.  Computer

recreations. Software— Practice and Experience, 1(2):201–204, April/June 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1971:CRc

[Ano71c]  Anonymous.  Computer

recreations. Software —Practice and Experience, 1(3):297–300, July/ September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1971:CRS

[Ano71d]  Anonymous.  Computer

recreations. Space-filling curves, or how to waste time with a plotter. Software—Practice and Experience, 1(4):403–410, October/December 1971. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1971:EAV

[Ano71e]  Anonymous.  Editorial:

Academics versus the rest. Software—Practice and Experience, 1(4):307–308, October/December 1971.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1971:EGE

[Ano71f]  Anonymous.  Editorial:

Guest editorial. Software—Practice and Experience, 1(1):1–4, January/ March 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1971:Ma

[Ano71g] Anonymous. Masthead.

Software—Practice and Experience, 1(1):fmi, January/March 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1971:Mb

[Ano71h] Anonymous. Masthead.

Software—Practice and Experience, 1(2):fmi, April/ June 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1971:Mc

[Ano71i] Anonymous. Masthead.

Software—Practice and Experience, 1(3):fmi, July/ September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1971:Md

[Ano71j] Anonymous. Masthead. Software—Practice and Experience, 1(4):fmi, October/December 1971. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1972:CR

[Ano72a]  Anonymous.  Computer

recreations. Software— Practice and Experience, 2 (1):93–96, January 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1972:CRM

[Ano72b]  Anonymous.  Computer

recreations. by midsummer madness, or how to waste more time with a plotter. Software—Practice and Experience, 2(3):293– 301, July 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1972:C

[Ano72c]  Anonymous.  Correc-

tion. Software—Practice and Experience, 2(2): 192, April 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1972:Ma

[Ano72d] Anonymous. Masthead. Software—Practice and Experience, 2(1):fmi, January 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1972:Mb

[Ano72e] Anonymous. Masthead.

Software—Practice and Experience, 2(2):fmi, April 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1972:Mc

[Ano72f] Anonymous. Masthead.

Software—Practice and Experience, 2(3):fmi, July 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1972:Md

[Ano72g] Anonymous. Masthead.

Software—Practice and Experience, 2(4):fmi, October 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1973:BRB

[Ano73a] Anonymous. Book review:

Games playing with computers, A. G. Bell, George Allen and Unwin, London, 1972. No. of pages: 204. Price: £5.25. Software— Practice and Experience, 3 (2):191, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1973:E

[Ano73b]  Anonymous.  Erratum.

Software—Practice and Experience, 3(4):401, October/December 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1973:Ma

[Ano73c] Anonymous. Masthead. Software—Practice and Experience, 3(1):fmi, January/March 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1973:Mb

[Ano73d] Anonymous. Masthead. Software—Practice and Experience, 3(2):fmi, April/ June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1973:Mc

[Ano73e] Anonymous. Masthead. Software—Practice and Experience, 3(3):fmi, July/ September 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1973:Md

[Ano73f] Anonymous. Masthead. Software—Practice and Experience, 3(4):fmi, October/December 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1974:Ma

[Ano74a] Anonymous. Masthead.

Software—Practice and Experience, 4(1):fmi, January/March 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1974:Mb

[Ano74b] Anonymous. Masthead.

Software—Practice and Experience, 4(2):fmi, April/ June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1974:Mc

[Ano74c] Anonymous. Masthead.

Software—Practice and Experience, 4(3):fmi, July/ September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1974:Md

[Ano74d] Anonymous. Masthead.

Software—Practice and Experience, 4(4):fmi, October/December 1974. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1975:E

[Ano75a]  Anonymous.  Editorial.

Software—Practice and Experience, 5(4):319, October 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1975:Ma

[Ano75b] Anonymous. Masthead. Software—Practice and Experience, 5(1):fmi, January 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1975:Mb

[Ano75c] Anonymous. Masthead. Software—Practice and Experience, 5(2):fmi, April 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1975:Mc

[Ano75d] Anonymous. Masthead. Software—Practice and Experience, 5(3):fmi, July 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1975:Md

[Ano75e] Anonymous. Masthead. Software—Practice and Experience, 5(4):fmi, October 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1976:CCJ

[Ano76a]  Anonymous.  Codasyl

Cobol Journal of Development, 1976. Software— Practice and Experience, 6 (4):580–581, October 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1976:EGE

[Ano76b]  Anonymous.  Editorial:

Guest editorial. Software— Practice and Experience, 6 (4):443–446, October 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1976:EKC

[Ano76c] Anonymous. Experiments with the Kronos control language. Software—Practice and Experience, 6(1): 133–135, January 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1976:Ma

[Ano76d] Anonymous. Masthead.

Software—Practice and Experience, 6(1):fmi, January 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1976:Mb

[Ano76e] Anonymous. Masthead.

Software—Practice and Experience, 6(2):fmi, April 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1976:Mc

[Ano76f] Anonymous. Masthead. Software—Practice and Experience, 6(3):fmi, July 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1976:Md

[Ano76g] Anonymous. Masthead. Software—Practice and Experience, 6(4):fmi, October 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1976:NSC

[Ano76h] Anonymous. A numbering scheme for a computer programming library. Software —Practice and Experience, 6(3):435–437, July 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1977:Ma

[Ano77a] Anonymous. Masthead. Software—Practice and Experience, 7(1):fmi, January 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1977:Mb

[Ano77b] Anonymous. Masthead. Software—Practice and Experience, 7(2):fmi, March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1977:Mc

[Ano77c] Anonymous. Masthead.

Software—Practice and Experience, 7(3):fmi, June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1977:Md

[Ano77d] Anonymous. Masthead.

Software—Practice and Experience, 7(4):fmi, July 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1977:Me

[Ano77e] Anonymous. Masthead.

Software—Practice and Experience, 7(5):fmi, September 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1977:Mf

[Ano77f] Anonymous. Masthead.

Software—Practice and Experience, 7(6):fmi, November 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1978:A

[Ano78a] Anonymous. Announcement. Software—Practice and Experience, 8(3): 377, May 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1978:Ma

[Ano78b] Anonymous. Masthead. Software—Practice and Experience, 8(1):fmi, January 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1978:Mb

[Ano78c] Anonymous. Masthead. Software—Practice and Experience, 8(2):fmi, March 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1978:Mc

[Ano78d] Anonymous. Masthead. Software—Practice and Experience, 8(3):fmi, May 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1978:Md

[Ano78e] Anonymous. Masthead. Software—Practice and Experience, 8(4):fmi, July 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1978:Me

[Ano78f] Anonymous. Masthead. Software—Practice and Experience, 8(5):fmi, September 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1978:Mf

[Ano78g] Anonymous. Masthead.

Software—Practice and Experience, 8(6):fmi, November 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1979:BRB

[Ano79a]  Anonymous.  Book re-

view: An introduction to programming and problems solving with Pascal, G. M. Schneider, S. W. Weingart and D. M. Perlman, Wiley, New York, 1978. No. of pages: 394. Price: £9.20 (hardback). Software —Practice and Experience, 9(9):792–794, September 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1979:CC

[Ano79b] Anonymous. Corrigendum:

Corrigendum. Software— Practice and Experience, 9 (2):i, February 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1979:Ma

[Ano79c] Anonymous. Masthead.

Software—Practice and Experience, 9(1):fmi, January 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1979:Mb

[Ano79d] Anonymous. Masthead. Software—Practice and Experience, 9(2):fmi, February 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1979:Mc

[Ano79e] Anonymous. Masthead. Software—Practice and Experience, 9(3):fmi, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1979:Md

[Ano79f] Anonymous. Masthead. Software—Practice and Experience, 9(4):fmi, April 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1979:Me

[Ano79g] Anonymous. Masthead. Software—Practice and Experience, 9(5):fmi, May 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1979:Mf

[Ano79h] Anonymous. Masthead. Software—Practice and Experience, 9(6):fmi, June 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1979:Mg

[Ano79i] Anonymous. Masthead.

Software—Practice and Experience, 9(7):fmi, July 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1979:Mh

[Ano79j] Anonymous. Masthead.

Software—Practice and Experience, 9(8):fmi, August 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1979:Mi

[Ano79k] Anonymous. Masthead.

Software—Practice and Experience, 9(9):fmi, September 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1979:Mj

[Ano79l] Anonymous. Masthead.

Software—Practice and Experience, 9(10):fmi, October 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1979:Mk

[Ano79m] Anonymous. Masthead.

Software—Practice and Experience, 9(11):fmi, November 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1979:Ml

[Ano79n] Anonymous. Masthead. Software—Practice and Experience, 9(12):fmi, December 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1980:CCL

[Ano80a] Anonymous. The 1978 CO-DASYL COBOL language specifications. Software —Practice and Experience, 10(7):589–590, July 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1980:LES

[Ano80b] Anonymous. Letter to the Editor: Some remarks on ‘A case for acquiring Pascal’. Software—Practice and Experience, 10(3):247– 248, March 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [Jos79].

Anonymous:1980:Ma

[Ano80c] Anonymous. Masthead. Software—Practice and Experience, 10(1):fmi, January 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1980:Mb

[Ano80d] Anonymous. Masthead. Software—Practice and Experience, 10(2):fmi, February 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1980:Mc

[Ano80e] Anonymous. Masthead.

Software—Practice and Experience, 10(3):fmi, March 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1980:Md

[Ano80f] Anonymous. Masthead.

Software—Practice and Experience, 10(4):fmi, April 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1980:Me

[Ano80g] Anonymous. Masthead.

Software—Practice and Experience, 10(5):fmi, May 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1980:Mf

[Ano80h] Anonymous. Masthead.

Software—Practice and Experience, 10(6):fmi, June 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1980:Mg

[Ano80i] Anonymous. Masthead.

Software—Practice and Experience, 10(7):fmi, July 1980. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1980:Mh

[Ano80j] Anonymous. Masthead. Software—Practice and Experience, 10(8):fmi, August 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1980:Mi

[Ano80k] Anonymous. Masthead. Software—Practice and Experience, 10(9):fmi, September 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1980:Mj

[Ano80l] Anonymous. Masthead. Software—Practice and Experience, 10(10):fmi, October 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1980:Mk

[Ano80m] Anonymous. Masthead. Software—Practice and Experience, 10(11):fmi, November 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1980:Ml

[Ano80n] Anonymous. Masthead. Software—Practice and Experience, 10(12):fmi, December 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:CC

[Ano81a] Anonymous. Corrigendum:

Corrigendum. Software —Practice and Experience, 11(5):i, May 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:Ma

[Ano81b] Anonymous. Masthead.

Software—Practice and Experience, 11(1):fmi, January 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:Mb

[Ano81c] Anonymous. Masthead.

Software—Practice and Experience, 11(2):fmi, February 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:Mc

[Ano81d] Anonymous. Masthead.

Software—Practice and Experience, 11(3):fmi, March 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:Md

[Ano81e] Anonymous. Masthead. Software—Practice and Experience, 11(4):fmi, April 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:Me

[Ano81f] Anonymous. Masthead. Software—Practice and Experience, 11(5):fmi, May 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:Mf

[Ano81g] Anonymous. Masthead. Software—Practice and Experience, 11(6):fmi, June 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:Mg

[Ano81h] Anonymous. Masthead. Software—Practice and Experience, 11(7):fmi, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:Mh

[Ano81i] Anonymous. Masthead. Software—Practice and Experience, 11(8):fmi, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:Mi

[Ano81j] Anonymous. Masthead.

Software—Practice and Experience, 11(9):fmi, September 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:Mj

[Ano81k] Anonymous. Masthead.

Software—Practice and Experience, 11(10):fmi, October 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:Mk

[Ano81l] Anonymous. Masthead.

Software—Practice and Experience, 11(11):fmi, November 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:Ml

[Ano81m] Anonymous. Masthead.

Software—Practice and Experience, 11(12):fmi, December 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:MFD

[Ano81n]  Anonymous.  Miscella-

neous: Fortran data base facility (Journal of development 1980). Software —Practice and Experience, 11(1):100, January 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1982:Ma

[Ano82a] Anonymous. Masthead. Software—Practice and Experience, 12(1):fmi, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1982:Mb

[Ano82b] Anonymous. Masthead. Software—Practice and Experience, 12(2):fmi, February 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1982:Mc

[Ano82c] Anonymous. Masthead. Software—Practice and Experience, 12(3):fmi, March 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1982:Md

[Ano82d] Anonymous. Masthead. Software—Practice and Experience, 12(4):fmi, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1982:Me

[Ano82e] Anonymous. Masthead. Software—Practice and Experience, 12(5):fmi, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1982:Mf

[Ano82f] Anonymous. Masthead.

Software—Practice and Experience, 12(6):fmi, June 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1982:Mg

[Ano82g] Anonymous. Masthead.

Software—Practice and Experience, 12(7):fmi, July 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1982:Mh

[Ano82h] Anonymous. Masthead.

Software—Practice and Experience, 12(8):fmi, August 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1982:Mi

[Ano82i] Anonymous. Masthead.

Software—Practice and Experience, 12(9):fmi, September 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1982:Mj

[Ano82j] Anonymous. Masthead.

Software—Practice and Experience, 12(10):fmi, October 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1982:Mk

[Ano82k] Anonymous. Masthead. Software—Practice and Experience, 12(11):fmi, November 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1982:Ml

[Ano82l] Anonymous. Masthead. Software—Practice and Experience, 12(12):fmi, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1983:CC

[Ano83a] Anonymous. Corrigendum: Corrigendum. Software —Practice and Experience, 13(9):i, September 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1983:Ma

[Ano83b] Anonymous. Masthead. Software—Practice and Experience, 13(2):fmi, February 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1983:Mb

[Ano83c] Anonymous. Masthead. Software—Practice and Experience, 13(3):fmi, March 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1983:Mc

[Ano83d] Anonymous. Masthead.

Software—Practice and Experience, 13(4):fmi, April 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1983:Md

[Ano83e] Anonymous. Masthead.

Software—Practice and Experience, 13(5):fmi, May 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1983:Me

[Ano83f] Anonymous. Masthead.

Software—Practice and Experience, 13(6):fmi, June 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1983:Mf

[Ano83g] Anonymous. Masthead.

Software—Practice and Experience, 13(7):fmi, July 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1983:Mg

[Ano83h] Anonymous. Masthead.

Software—Practice and Experience, 13(8):fmi, August 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1983:Mh

[Ano83i] Anonymous. Masthead. Software—Practice and Experience, 13(9):fmi, September 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1983:Mi

[Ano83j] Anonymous. Masthead. Software—Practice and Experience, 13(10):fmi, October 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1983:Mj

[Ano83k] Anonymous. Masthead. Software—Practice and Experience, 13(11):fmi, November 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1983:Mk

[Ano83l] Anonymous. Masthead. Software—Practice and Experience, 13(12):fmi, December 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1984:CC

[Ano84a] Anonymous. Corrigendum: Corrigendum. Software —Practice and Experience, 14(7):704, July 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1984:Ma

[Ano84b] Anonymous. Masthead.

Software—Practice and Experience, 14(1):fmi, January 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1984:Mb

[Ano84c] Anonymous. Masthead.

Software—Practice and Experience, 14(2):fmi, February 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1984:Mc

[Ano84d] Anonymous. Masthead.

Software—Practice and Experience, 14(3):fmi, March 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1984:Md

[Ano84e] Anonymous. Masthead.

Software—Practice and Experience, 14(4):fmi, April 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1984:Me

[Ano84f] Anonymous. Masthead.

Software—Practice and Experience, 14(5):fmi, May 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1984:Mf

[Ano84g] Anonymous. Masthead. Software—Practice and Experience, 14(6):fmi, June 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1984:Mg

[Ano84h] Anonymous. Masthead. Software—Practice and Experience, 14(7):fmi, July 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1984:Mh

[Ano84i] Anonymous. Masthead. Software—Practice and Experience, 14(8):fmi, August 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1984:Mi

[Ano84j] Anonymous. Masthead. Software—Practice and Experience, 14(9):fmi, September 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1984:Mj

[Ano84k] Anonymous. Masthead. Software—Practice and Experience, 14(10):fmi, October 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1984:Mk

[Ano84l] Anonymous. Masthead.

Software—Practice and Experience, 14(11):fmi, November 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1984:Ml

[Ano84m] Anonymous. Masthead.

Software—Practice and Experience, 14(12):fmi, December 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1985:Ma

[Ano85a] Anonymous. Masthead.

Software—Practice and Experience, 15(1):fmi, January 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1985:Mb

[Ano85b] Anonymous. Masthead.

Software—Practice and Experience, 15(2):fmi, February 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1985:Mc

[Ano85c] Anonymous. Masthead.

Software—Practice and Experience, 15(3):fmi, March 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1985:Md

[Ano85d] Anonymous. Masthead. Software—Practice and Experience, 15(4):fmi, April 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1985:Me

[Ano85e] Anonymous. Masthead. Software—Practice and Experience, 15(5):fmi, May 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1985:Mf

[Ano85f] Anonymous. Masthead. Software—Practice and Experience, 15(6):fmi, June 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1985:Mg

[Ano85g] Anonymous. Masthead. Software—Practice and Experience, 15(7):fmi, July 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1985:Mh

[Ano85h] Anonymous. Masthead. Software—Practice and Experience, 15(8):fmi, August 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1985:Mi

[Ano85i] Anonymous. Masthead.

Software—Practice and Experience, 15(9):fmi, September 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1985:Mj

[Ano85j] Anonymous. Masthead.

Software—Practice and Experience, 15(10):fmi, October 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1985:Mk

[Ano85k] Anonymous. Masthead.

Software—Practice and Experience, 15(11):fmi, November 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1985:Ml

[Ano85l] Anonymous. Masthead.

Software—Practice and Experience, 15(12):fmi, December 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1986:E

[Ano86a] Anonymous. Errata. Software—Practice and Experience, 16(5):i, May 1986.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1986:Ma

[Ano86b] Anonymous. Masthead. Software—Practice and Experience, 16(1):fmi, January 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1986:Mb

[Ano86c] Anonymous. Masthead. Software—Practice and Experience, 16(2):fmi, February 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1986:Mc

[Ano86d] Anonymous. Masthead. Software—Practice and Experience, 16(3):fmi, March 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1986:Md

[Ano86e] Anonymous. Masthead. Software—Practice and Experience, 16(4):fmi, April 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1986:Me

[Ano86f] Anonymous. Masthead. Software—Practice and Experience, 16(5):fmi, May 1986. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1986:Mf

[Ano86g] Anonymous. Masthead.

Software—Practice and Experience, 16(6):fmi, June 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1986:Mg

[Ano86h] Anonymous. Masthead.

Software—Practice and Experience, 16(7):fmi, July 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1986:Mh

[Ano86i] Anonymous. Masthead.

Software—Practice and Experience, 16(10):fmi, October 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1986:Mi

[Ano86j] Anonymous. Masthead.

Software—Practice and Experience, 16(11):fmi, November 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1986:Mj

[Ano86k] Anonymous. Masthead.

Software—Practice and Experience, 16(12):fmi, December 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1987:BRB

[Ano87a]  Anonymous.  Book re-

view: Modula-2: A Software Development Approach, Gary A. Ford and Richard S. Wiener, Wiley, 1985. ISBN 0-471-87834-0. No. of pages: 400 + xvi. Price: £21.45 (paperback). Software—Practice and Experience, 17(2):167, February 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1987:E

[Ano87b] Anonymous. Errata. Software—Practice and Experience, 17(4):i, April 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1987:Ma

[Ano87c] Anonymous. Masthead. Software—Practice and Experience, 17(1):fmi, January 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1987:Mb

[Ano87d] Anonymous. Masthead. Software—Practice and Experience, 17(2):fmi, February 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1987:Mc

[Ano87e] Anonymous. Masthead.

Software—Practice and Experience, 17(3):fmi, March 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1987:Md

[Ano87f] Anonymous. Masthead.

Software—Practice and Experience, 17(4):fmi, April 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1987:Me

[Ano87g] Anonymous. Masthead.

Software—Practice and Experience, 17(5):fmi, May 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1987:Mf

[Ano87h] Anonymous. Masthead.

Software—Practice and Experience, 17(6):fmi, June 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1987:Mg

[Ano87i] Anonymous. Masthead.

Software—Practice and Experience, 17(7):fmi, July 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1987:Mh

[Ano87j] Anonymous. Masthead. Software—Practice and Experience, 17(8):fmi, August 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1987:Mi

[Ano87k] Anonymous. Masthead. Software—Practice and Experience, 17(9):fmi, September 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1987:Mj

[Ano87l] Anonymous. Masthead. Software—Practice and Experience, 17(10):fmi, October 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1987:Mk

[Ano87m] Anonymous. Masthead. Software—Practice and Experience, 17(11):fmi, November 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1987:Ml

[Ano87n] Anonymous. Masthead. Software—Practice and Experience, 17(12):fmi, December 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:BRBc

[Ano88a]  Anonymous.  Book re-

view: C for programmers, Leendert Ammeraal, Wiley, 1986. ISBN: 0-471-91128-3. No. of Pages: 118. Price: £9.95. Software —Practice and Experience, 18(8):805, August 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:BRBb

[Ano88b]  Anonymous.  Book re-

view: Practical C, Mark Harrison, Sigma Press, 1985. ISBN: 1-85058-035-9. No. of Pages: 173. Price: £7.95. Software—Practice and Experience, 18(8): 805, August 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:BRBa

[Ano88c]  Anonymous.  Book re-

view: Productive software test management, Michael W. Evans, Wiley, 1984. ISBN 0-471-88311-5. No. of pages: 218. Price: £32.10. Software—Practice and Experience, 18(3): 285, March 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:C

[Ano88d] Anonymous. Correction. Software—Practice and Experience, 18(7):i, July 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:Ma

[Ano88e] Anonymous. Masthead. Software—Practice and Experience, 18(1):fmi, January 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:Mb

[Ano88f] Anonymous. Masthead. Software—Practice and Experience, 18(2):fmi, February 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:Mc

[Ano88g] Anonymous. Masthead. Software—Practice and Experience, 18(3):fmi, March 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:Md

[Ano88h] Anonymous. Masthead. Software—Practice and Experience, 18(4):fmi, April 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:Me

[Ano88i] Anonymous. Masthead.

Software—Practice and Experience, 18(5):fmi, May 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:Mf

[Ano88j] Anonymous. Masthead.

Software—Practice and Experience, 18(6):fmi, June 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:Mg

[Ano88k] Anonymous. Masthead.

Software—Practice and Experience, 18(7):fmi, July 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:Mh

[Ano88l] Anonymous. Masthead.

Software—Practice and Experience, 18(8):fmi, August 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:Mi

[Ano88m] Anonymous. Masthead.

Software—Practice and Experience, 18(9):fmi, September 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:Mj

[Ano88n] Anonymous. Masthead. Software—Practice and Experience, 18(10):fmi, October 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:Mk

[Ano88o] Anonymous. Masthead. Software—Practice and Experience, 18(11):fmi, November 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:Ml

[Ano88p] Anonymous. Masthead. Software—Practice and Experience, 18(12):fmi, December 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1989:C

[Ano89a]  Anonymous.  Correc-

tion. Software—Practice and Experience, 19(1):i, January 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1989:E

[Ano89b]  Anonymous.  Editorial.

Software—Practice and Experience, 19(1):iii, January 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1989:Ma

[Ano89c] Anonymous. Masthead.

Software—Practice and Experience, 19(1):fmi, January 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1989:Mb

[Ano89d] Anonymous. Masthead.

Software—Practice and Experience, 19(2):fmi, February 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1989:Mc

[Ano89e] Anonymous. Masthead.

Software—Practice and Experience, 19(3):fmi, March 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1989:Md

[Ano89f] Anonymous. Masthead.

Software—Practice and Experience, 19(4):fmi, April 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1989:Me

[Ano89g] Anonymous. Masthead.

Software—Practice and Experience, 19(5):fmi, May 1989. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1989:Mf

[Ano89h] Anonymous. Masthead. Software—Practice and Experience, 19(6):fmi, June 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1989:Mg

[Ano89i] Anonymous. Masthead. Software—Practice and Experience, 19(7):fmi, July 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1989:Mh

[Ano89j] Anonymous. Masthead. Software—Practice and Experience, 19(8):fmi, August 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1989:Mi

[Ano89k] Anonymous. Masthead. Software—Practice and Experience, 19(9):fmi, September 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1989:Mj

[Ano89l] Anonymous. Masthead. Software—Practice and Experience, 19(10):fmi, October 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1989:Mk

[Ano89m] Anonymous. Masthead.

Software—Practice and Experience, 19(11):fmi, November 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1990:Ma

[Ano90a] Anonymous. Masthead.

Software—Practice and Experience, 20(S1):fmi, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1990:Mb

[Ano90b] Anonymous. Masthead.

Software—Practice and Experience, 20(S2):fmi, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1990:Mc

[Ano90c] Anonymous. Masthead.

Software—Practice and Experience, 20(1):fmi, January 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1990:Md

[Ano90d] Anonymous. Masthead.

Software—Practice and Experience, 20(2):fmi, February 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1990:Me

[Ano90e] Anonymous. Masthead. Software—Practice and Experience, 20(3):fmi, March 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1990:Mf

[Ano90f] Anonymous. Masthead. Software—Practice and Experience, 20(4):fmi, April 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1990:Mg

[Ano90g] Anonymous. Masthead. Software—Practice and Experience, 20(5):fmi, May 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1990:Mh

[Ano90h] Anonymous. Masthead. Software—Practice and Experience, 20(6):fmi, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1990:Mi

[Ano90i] Anonymous. Masthead. Software—Practice and Experience, 20(7):fmi, July 1990. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1990:Mj

[Ano90j] Anonymous. Masthead.

Software—Practice and Experience, 20(8):fmi, August 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1990:Mk

[Ano90k] Anonymous. Masthead.

Software—Practice and Experience, 20(9):fmi, September 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1990:Ml

[Ano90l] Anonymous. Masthead.

Software—Practice and Experience, 20(10):fmi, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1990:Mm

[Ano90m] Anonymous. Masthead.

Software—Practice and Experience, 20(12):fmi, December 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1991:Ma

[Ano91a] Anonymous. Masthead.

Software—Practice and Experience, 21(1):fmi, January 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1991:Mb

[Ano91b] Anonymous. Masthead. Software—Practice and Experience, 21(2):fmi, February 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1991:Mc

[Ano91c] Anonymous. Masthead. Software—Practice and Experience, 21(3):fmi, March 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1991:Md

[Ano91d] Anonymous. Masthead. Software—Practice and Experience, 21(4):fmi, April 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1991:Me

[Ano91e] Anonymous. Masthead. Software—Practice and Experience, 21(5):fmi, May 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1991:Mf

[Ano91f] Anonymous. Masthead. Software—Practice and Experience, 21(6):fmi, June 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1991:Mg

[Ano91g] Anonymous. Masthead.

Software—Practice and Experience, 21(7):fmi, July 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1991:Mh

[Ano91h] Anonymous. Masthead.

Software—Practice and Experience, 21(8):fmi, August 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1991:Mi

[Ano91i] Anonymous. Masthead.

Software—Practice and Experience, 21(9):fmi, September 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1991:Mj

[Ano91j] Anonymous. Masthead.

Software—Practice and Experience, 21(10):fmi, October 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1991:Mk

[Ano91k] Anonymous. Masthead.

Software—Practice and Experience, 21(11):fmi, November 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1991:Ml

[Ano91l] Anonymous. Masthead. Software—Practice and Experience, 21(12):fmi, December 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1992:Ma

[Ano92a] Anonymous. Masthead. Software—Practice and Experience, 22(1):fmi, January 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1992:Mb

[Ano92b] Anonymous. Masthead. Software—Practice and Experience, 22(2):fmi, February 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1992:Mc

[Ano92c] Anonymous. Masthead. Software—Practice and Experience, 22(3):fmi, March 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1992:Md

[Ano92d] Anonymous. Masthead. Software—Practice and Experience, 22(4):fmi, April 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1992:Me

[Ano92e] Anonymous. Masthead.

Software—Practice and Experience, 22(5):fmi, May 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1992:Mf

[Ano92f] Anonymous. Masthead.

Software—Practice and Experience, 22(6):fmi, June 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1992:Mg

[Ano92g] Anonymous. Masthead.

Software—Practice and Experience, 22(7):fmi, July 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1992:Mh

[Ano92h] Anonymous. Masthead.

Software—Practice and Experience, 22(8):fmi, August 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1992:Mi

[Ano92i] Anonymous. Masthead.

Software—Practice and Experience, 22(9):fmi, September 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1992:Mj

[Ano92j] Anonymous. Masthead. Software—Practice and Experience, 22(10):fmi, October 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1992:Mk

[Ano92k] Anonymous. Masthead. Software—Practice and Experience, 22(11):fmi, November 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1992:Ml

[Ano92l] Anonymous. Masthead. Software—Practice and Experience, 22(12):fmi, December 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1993:CS

[Ano93a] Anonymous. Champagne and software. Software —Practice and Experience, 23(5):ii, May 1993. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1993:Ma

[Ano93b] Anonymous. Masthead. Software—Practice and Experience, 23(1):fmi, January 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1993:Mb

[Ano93c] Anonymous. Masthead.

Software—Practice and Experience, 23(2):fmi, February 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1993:Mc

[Ano93d] Anonymous. Masthead.

Software—Practice and Experience, 23(3):fmi, March 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1993:Md

[Ano93e] Anonymous. Masthead.

Software—Practice and Experience, 23(4):fmi, April 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1993:Me

[Ano93f] Anonymous. Masthead.

Software—Practice and Experience, 23(5):fmi, May 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1993:Mf

[Ano93g] Anonymous. Masthead.

Software—Practice and Experience, 23(6):fmi, June 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1993:Mg

[Ano93h] Anonymous. Masthead. Software—Practice and Experience, 23(7):fmi, July 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1993:Mh

[Ano93i] Anonymous. Masthead. Software—Practice and Experience, 23(8):fmi, August 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1993:Mi

[Ano93j] Anonymous. Masthead. Software—Practice and Experience, 23(9):fmi, September 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1993:Mj

[Ano93k] Anonymous. Masthead. Software—Practice and Experience, 23(10):fmi, October 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1993:Mk

[Ano93l] Anonymous. Masthead. Software—Practice and Experience, 23(11):fmi, November 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1993:Ml

[Ano93m] Anonymous. Masthead.

Software—Practice and Experience, 23(12):fmi, December 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1994:Ma

[Ano94a] Anonymous. Masthead.

Software—Practice and Experience, 24(1):fmi, January 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1994:Mb

[Ano94b] Anonymous. Masthead.

Software—Practice and Experience, 24(2):fmi, February 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1994:Mc

[Ano94c] Anonymous. Masthead.

Software—Practice and Experience, 24(3):fmi, March 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1994:Md

[Ano94d] Anonymous. Masthead.

Software—Practice and Experience, 24(4):fmi, April 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1994:Me

[Ano94e] Anonymous. Masthead. Software—Practice and Experience, 24(5):fmi, May 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1994:Mf

[Ano94f] Anonymous. Masthead. Software—Practice and Experience, 24(6):fmi, June 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1994:Mg

[Ano94g] Anonymous. Masthead. Software—Practice and Experience, 24(7):fmi, July 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1994:Mh

[Ano94h] Anonymous. Masthead. Software—Practice and Experience, 24(8):fmi, August 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1994:Mi

[Ano94i] Anonymous. Masthead. Software—Practice and Experience, 24(9):fmi, September 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1994:Mj

[Ano94j] Anonymous. Masthead.

Software—Practice and Experience, 24(10):fmi, October 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1994:Mk

[Ano94k] Anonymous. Masthead.

Software—Practice and Experience, 24(11):fmi, November 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1994:M

[Ano94l] Anonymous. Masthead.

Software—Practice and Experience, 24(12):fmi, December 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1995:Ma

[Ano95a] Anonymous. Masthead.

Software—Practice and Experience, 25(S4):fmi, December 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1995:Mb

[Ano95b] Anonymous. Masthead.

Software—Practice and Experience, 25(1):fmi, January 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1995:Mc

[Ano95c] Anonymous. Masthead. Software—Practice and Experience, 25(2):fmi, February 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1995:Md

[Ano95d] Anonymous. Masthead. Software—Practice and Experience, 25(3):fmi, March 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1995:Me

[Ano95e] Anonymous. Masthead. Software—Practice and Experience, 25(4):fmi, April 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1995:Mf

[Ano95f] Anonymous. Masthead. Software—Practice and Experience, 25(5):fmi, May 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1995:Mg

[Ano95g] Anonymous. Masthead. Software—Practice and Experience, 25(6):fmi, June 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1995:Mh

[Ano95h] Anonymous. Masthead.

Software—Practice and Experience, 25(7):fmi, July 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1995:Mi

[Ano95i] Anonymous. Masthead.

Software—Practice and Experience, 25(8):fmi, August 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1995:Mj

[Ano95j] Anonymous. Masthead.

Software—Practice and Experience, 25(9):fmi, September 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1995:Mk

[Ano95k] Anonymous. Masthead.

Software—Practice and Experience, 25(10):fmi, October 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1995:Ml

[Ano95l] Anonymous. Masthead.

Software—Practice and Experience, 25(11):fmi, November 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1995:Mm

[Ano95m] Anonymous. Masthead. Software—Practice and Experience, 25(12):fmi, December 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1995:PAP

[Ano95n] Anonymous. Publisher’s announcements: Publisher’s announcement. Software—Practice and Experience, 25(11):i, November 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:APAa

[Ano96a] Anonymous. Announcement: Publisher’s announcement. Software— Practice and Experience, 26(1):i, January 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:APAb

[Ano96b] Anonymous. Announcement: Publisher’s announcement. Software— Practice and Experience, 26(3):i, March 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:APAc

[Ano96c] Anonymous. Announcement: Publisher’s announcement. Software— Practice and Experience, 26(4):i, April 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:APAd

[Ano96d] Anonymous. Announcement: Publisher’s announcement. Software— Practice and Experience, 26(5):i, May 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:APAe

[Ano96e] Anonymous. Announcement: Publisher’s announcement. Software— Practice and Experience, 26(6):i, June 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:APAf

[Ano96f] Anonymous. Announcement: Publisher’s announcement. Software— Practice and Experience, 26(8):i, August 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:APAg

[Ano96g] Anonymous. Announcement: Publisher’s announcement. Software— Practice and Experience, 26(9):i, September 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:APAh

[Ano96h] Anonymous. Announcement: Publisher’s announcement. Software— Practice and Experience, 26(10):i, October 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:APAi

[Ano96i] Anonymous. Announcement: Publisher’s announcement. Software— Practice and Experience, 26(11):i, November 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:APAj

[Ano96j] Anonymous. Announcement: Publisher’s announcement. Software— Practice and Experience, 26(12):i, December 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:Ma

[Ano96k] Anonymous. Masthead. Software—Practice and Experience, 26(1):fmi, January 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:Mb

[Ano96l] Anonymous. Masthead.

Software—Practice and Experience, 26(2):fmi, February 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:Mc

[Ano96m] Anonymous. Masthead.

Software—Practice and Experience, 26(3):fmi, March 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:Md

[Ano96n] Anonymous. Masthead.

Software—Practice and Experience, 26(4):fmi, April 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:Me

[Ano96o] Anonymous. Masthead.

Software—Practice and Experience, 26(5):fmi, May 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:Mf

[Ano96p] Anonymous. Masthead.

Software—Practice and Experience, 26(6):fmi, June 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:Mg

[Ano96q] Anonymous. Masthead. Software—Practice and Experience, 26(7):fmi, July 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:Mh

[Ano96r] Anonymous. Masthead. Software—Practice and Experience, 26(8):fmi, August 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:Mi

[Ano96s] Anonymous. Masthead. Software—Practice and Experience, 26(9):fmi, September 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:Mj

[Ano96t] Anonymous. Masthead. Software—Practice and Experience, 26(10):fmi, October 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:Mk

[Ano96u] Anonymous. Masthead. Software—Practice and Experience, 26(11):fmi, November 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1996:Ml

[Ano96v] Anonymous. Masthead.

Software—Practice and Experience, 26(12):fmi, December 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2009:CPS

[Ano09] Anonymous. Call for papers: ‘Selected PhD Literature Reviews in the Practical Aspects of Software Technology’. Software— Practice and Experience, 39(7):737, May ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2013:CPI

[Ano13] Anonymous. Call for papers: Intelligent technologies and applications for big data analytics. Software—Practice and Experience, 43(6):731–732, June 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:AAN

[Ano16a] Anonymous. Announcement: Announcing a new editorial structure. Software—Practice and Experience, 46(5):579, May 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IIa

[Ano16b] Anonymous. Issue information. Software—Practice and Experience, 46(5):577– 578, May 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IIb

[Ano16c] Anonymous. Issue information. Software—Practice and Experience, 46(6):721– 722, June 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IIc

[Ano16d] Anonymous. Issue information. Software—Practice and Experience, 46(7):865– 866, July 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IId

[Ano16e] Anonymous. Issue information. Software—Practice and Experience, 46(8): 1009–1010, August 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IIe

[Ano16f] Anonymous. Issue information. Software—Practice and Experience, 46 (9):1153–1154, September 2016. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IIf

[Ano16g] Anonymous. Issue information. Software—Practice and Experience, 46(10): 1297–1298, October 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IIg

[Ano16h] Anonymous. Issue information. Software—Practice and Experience, 46 (11):1441–1442, November 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IIh

[Ano16i] Anonymous. Issue information. Software—Practice and Experience, 46 (12):1589–1590, December 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IIIa

[Ano16j] Anonymous. Issue information — info page. Software —Practice and Experience, 46(1):2, January 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IIIb

[Ano16k] Anonymous. Issue information — info page. Software —Practice and Experience, 46(2):132, February 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IIIc

[Ano16l] Anonymous. Issue information — info page. Software—Practice and Experience, 46(3):290, March 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IIId

[Ano16m] Anonymous. Issue information — info page. Software—Practice and Experience, 46(4):434, April 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IITa

[Ano16n] Anonymous. Issue information — TOC. Software —Practice and Experience, 46(1):1, January 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IITb

[Ano16o] Anonymous. Issue information — TOC. Software —Practice and Experience, 46(2):131, February 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IITc

[Ano16p] Anonymous. Issue information — TOC. Software—Practice and Experience, 46(3):289, March 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2016:IIT

[Ano16q] Anonymous. Issue information - TOC. Software —Practice and Experience, 46(4):433, April 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2017:IIa

[Ano17a] Anonymous. Issue information. Software—Practice and Experience, 47(1):1–2, January 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2017:IIb

[Ano17b] Anonymous. Issue information. Software—Practice and Experience, 47(2):161– 162, February 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2017:IIc

[Ano17c] Anonymous. Issue information. Software—Practice and Experience, 47(3):343– 344, March 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2017:IId

[Ano17d] Anonymous. Issue information. Software—Practice and Experience, 47(4):503– 504, April 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2017:IIe

[Ano17e] Anonymous. Issue information. Software—Practice and Experience, 47(5):645– 646, May 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2017:IIf

[Ano17f] Anonymous. Issue information. Software—Practice and Experience, 47(6):797– 798, June 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2017:IIg

[Ano17g] Anonymous. Issue information. Software—Practice and Experience, 47(7):939– 940, July 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2017:IIh

[Ano17h] Anonymous. Issue information. Software—Practice and Experience, 47(8):

1041–1042, August 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2017:IIi

[Ano17i] Anonymous. Issue information. Software—Practice and Experience, 47 (9):1157–1158, September 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2017:IIj

[Ano17j] Anonymous. Issue information. Software—Practice and Experience, 47(10): 1297–1298, October 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2017:IIk

[Ano17k] Anonymous. Issue information. Software—Practice and Experience, 47 (11):1463–1464, November 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2017:IIl

[Ano17l] Anonymous. Issue information. Software—Practice and Experience, 47 (12):1875–1876, December 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2017:RA

[Ano17m] Anonymous. Reviewer acknowledgements. Software—Practice and Experience, 47(1):157–159, January 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2018:IIa

[Ano18a] Anonymous. Issue information. Software—Practice and Experience, 48(1):1–2, January 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2018:IIb

[Ano18b] Anonymous. Issue information. Software—Practice and Experience, 48(2): 261–262, February 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2530.

Anonymous:2018:IIc

[Ano18c] Anonymous. Issue information. Software—Practice and Experience, 48 (3):383–384, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2531.

Anonymous:2018:IId

[Ano18d] Anonymous. Issue information. Software—Practice and Experience, 48 (4):759–760, April 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2532.

Anonymous:2018:IIe

[Ano18e] Anonymous. Issue information. Software—Practice and Experience, 48(5): 983–984, May 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2533.

Anonymous:2018:IIf

[Ano18f] Anonymous. Issue information. Software—Practice and Experience, 48 (6):1165–1166, June 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2534.

Anonymous:2018:IIg

[Ano18g] Anonymous. Issue information. Software—Practice and Experience, 48 (7):1331–1332, July 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2535.

Anonymous:2018:IIh

[Ano18h] Anonymous. Issue information. Software—Practice and Experience, 48(8): 1417–1418, August 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2536.

Anonymous:2018:IIi

[Ano18i] Anonymous. Issue information. Software—Practice and Experience, 48(10): 1727–1728, October 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2018:IIj

[Ano18j] Anonymous. Issue information. Software—Practice and Experience, 48 (11):1921–1922, November 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2018:IIk

[Ano18k] Anonymous. Issue information. Software—Practice and Experience, 48 (12):2109–2110, December 2018. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2018:IIl

[Ano18l] Anonymous. Issue information. Software—Practice and Experience, 48 (9):1521–1522, September 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2019:E

[Ano19a]  Anonymous.  Erratum.

Software—Practice and Experience, 49(4):694, April 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Correction of image caption in [MBG19b].

Anonymous:2019:IIa

[Ano19b] Anonymous. Issue information. Software—Practice and Experience, 49(2):149– 150, February 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2019:IIb

[Ano19c] Anonymous. Issue information. Software—Practice and Experience, 49(3):339– 340, March 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2019:IIc

[Ano19d] Anonymous. Issue information. Software—Practice and Experience, 49(4):559– 560, April 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2019:IId

[Ano19e] Anonymous. Issue information. Software—Practice and Experience, 49(5):695– 696, May 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2019:IIe

[Ano19f] Anonymous. Issue information. Software—Practice and Experience, 49(6):951– 952, June 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2019:IIf

[Ano19g] Anonymous. Issue information. Software—Practice and Experience, 49(7): 1053–1054, July 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2019:IIg

[Ano19h] Anonymous. Issue information. Software—Practice and Experience, 49(8): 1195–1196, August 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2019:IIh

[Ano19i] Anonymous. Issue information. Software—Practice and Experience, 49 (9):1329–1330, September 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2019:IIi

[Ano19j] Anonymous. Issue information. Software—Practice and Experience, 49(10): 1419–1420, October 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2019:IIj

[Ano19k] Anonymous. Issue information. Software—Practice and Experience, 49 (11):1571–1572, November 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2019:IIk

[Ano19l] Anonymous. Issue information. Software—Practice and Experience, 49 (12):1687–1688, December 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2020:IIa

[Ano20a] Anonymous. Issue information. Software—Practice and Experience, 50(1):1–2, January 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2020:IIb

[Ano20b] Anonymous. Issue information. Software—Practice and Experience, 50(2): 87–88, February 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2020:IIc

[Ano20c] Anonymous. Issue information. Software—Practice and Experience, 50(3):185– 186, March 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2020:IId

[Ano20d] Anonymous. Issue information. Software—Practice and Experience, 50(4):323– 324, April 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2020:IIe

[Ano20e] Anonymous. Issue information. Software—Practice and Experience, 50(5):471– 472, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2020:IIf

[Ano20f] Anonymous. Issue information. Software—Practice and Experience, 50(6):801– 802, June 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2020:IIg

[Ano20g] Anonymous. Issue information. Software—Practice and Experience, 50(7): 1023–1024, July 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2020:IIh

[Ano20h] Anonymous. Issue information. Software—Practice and Experience, 50(8): 1325–1326, August 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2020:IIi

[Ano20i] Anonymous. Issue information. Software—Practice and Experience, 50 (9):1627–1628, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2020:IIj

[Ano20j] Anonymous. Issue information. Software—Practice and Experience, 50(10): 1875–1876, October 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2020:IIk

[Ano20k] Anonymous. Issue information. Software—Practice and Experience, 50 (11):2009–2010, November 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2020:IIl

[Ano20l] Anonymous. Issue information. Software—Practice and Experience, 50 (12):2119–2120, December 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2021:IIa

[Ano21a] Anonymous. Issue information. Software—Practice and Experience, 51(1):1–2, January 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2021:IIb

[Ano21b] Anonymous. Issue information. Software—Practice and Experience, 51(2):157– 158, February 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2021:IIc

[Ano21c] Anonymous. Issue information. Software—Practice and Experience, 51(3):485– 486, March 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2021:IId

[Ano21d] Anonymous. Issue information. Software—Practice and Experience, 51(4):687– 688, April 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2021:IIe

[Ano21e] Anonymous. Issue information. Software—Practice and Experience, 51(5):849– 850, May 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2021:IIf

[Ano21f] Anonymous. Issue information. Software—Practice and Experience, 51(6): 1171–1172, June 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2021:IIg

[Ano21g] Anonymous. Issue information. Software—Practice and Experience, 51(7): 1481–1482, July 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2021:IIh

[Ano21h] Anonymous. Issue information. Software—Practice and Experience, 51(8):

1639–1640, August 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2021:IIi

[Ano21i] Anonymous. Issue information. Software—Practice and Experience, 51 (9):1847–1848, September 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2021:IIj

[Ano21j] Anonymous. Issue information. Software—Practice and Experience, 51(10): 1983–1984, October 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2021:IIk

[Ano21k] Anonymous. Issue information. Software—Practice and Experience, 51 (11):2103–2104, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2021:IIl

[Ano21l] Anonymous. Issue information. Software—Practice and Experience, 51 (12):2355–2356, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2022:IIa

[Ano22a] Anonymous. Issue information. Software—Practice and Experience, 52(1):1–2, January 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2022:IIb

[Ano22b] Anonymous. Issue information. Software—Practice and Experience, 52(2):317– 318, February 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2022:IIc

[Ano22c] Anonymous. Issue information. Software—Practice and Experience, 52(3):637– 638, March 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2022:IId

[Ano22d] Anonymous. Issue information. Software—Practice and Experience, 52(4):819– 820, April 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2022:IIe

[Ano22e] Anonymous. Issue information. Software—Practice and Experience, 52(5): 1093–1094, May 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2022:IIf

[Ano22f] Anonymous. Issue information. Software—Practice and Experience, 52(6): 1317–1318, June 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2022:IIg

[Ano22g] Anonymous. Issue information. Software—Practice and Experience, 52(7): 1541–1542, July 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2022:IIh

[Ano22h] Anonymous. Issue information. Software—Practice and Experience, 52(8): 1783–1784, August 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2022:IIi

[Ano22i] Anonymous. Issue information. Software—Practice and Experience, 52 (9):1903–1904, September 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2022:IIj

[Ano22j] Anonymous. Issue information. Software—Practice and Experience, 52(10):

2045–2046, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2022:IIk

[Ano22k] Anonymous. Issue information. Software—Practice and Experience, 52 (11):2347–2348, November 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2022:IIl

[Ano22l] Anonymous. Issue information. Software—Practice and Experience, 52 (12):2497–2498, December 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2023:IIa

[Ano23a] Anonymous. Issue information. Software—Practice and Experience, 53(1):1–2, January 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2023:IIb

[Ano23b] Anonymous. Issue information. Software—Practice and Experience, 53(2):241– 242, February 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2023:IIc

[Ano23c] Anonymous. Issue information. Software—Practice and Experience, 53(3):553– 554, March 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2023:IId

[Ano23d] Anonymous. Issue information. Software—Practice and Experience, 53(4):893– 894, April 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2023:IIe

[Ano23e] Anonymous. Issue information. Software—Practice and Experience, 53(5): 1123–1124, May 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2023:IIf

[Ano23f] Anonymous. Issue information. Software—Practice and Experience, 53(6): 1285–1286, June 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2023:IIg

[Ano23g] Anonymous. Issue information. Software—Practice and Experience, 53(7): 1465–1466, July 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2023:IIh

[Ano23h] Anonymous. Issue information. Software—Practice and Experience, 53(8): 1629–1630, August 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2023:IIi

[Ano23i] Anonymous. Issue information. Software—Practice and Experience, 53 (9):1739–1740, September 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2023:IIj

[Ano23j] Anonymous. Issue information. Software—Practice and Experience, 53(10): 1851–1852, October 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2023:IIk

[Ano23k] Anonymous. Issue information. Software—Practice and Experience, 53 (11):2065–2066, November 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2023:IIl

[Ano23l] Anonymous. Issue information. Special focus section:

Benchmarking, Experimentation Tools, and Reproducible Practices for Data-Intensive Systems from Edge to Cloud: Software: Practice and Experience, 53(12):2323–2324, December 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2023:RAS

[Ano23m] Anonymous. Retraction: Al Said, N, Gorbachev, Y, Avdeenko, A. An unmanned aerial vehicles navigation system on the basis of pattern recognition applications-Review of implementation options and prospects for development. : Practice and Experience 2021; 51: 1509– 1517. Software—Practice and Experience, 53(9): 1850, September 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [SGA21].

Anonymous:2024:IIa

[Ano24a] Anonymous. Issue information. Software—Practice and Experience, 54(1):1–2, January 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2024:IIb

[Ano24b] Anonymous. Issue information. Software—Practice and Experience, 54(2):107– 108, February 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2024:IIc

[Ano24c] Anonymous. Issue information. Software—Practice and Experience, 54(3):361– 362, March 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2024:IId

[Ano24d] Anonymous. Issue information. Efficient Management of Microservicebased Systems and Applications: Software: Practice and Experience, 54(4):541– 542, April 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2024:IIe

[Ano24e] Anonymous. Issue information. Software—Practice and Experience, 54(5):635– 636, May 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:2024:IIf

[Ano24f] Anonymous. Issue information. Software—Practice and Experience, 54(6):927– 928, June 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anstey:1986:CAD

[Ans86]  P. Anstey.  Comput-

ing advice at a distance: the ‘remote advisory’ concept. Software—Practice and Experience, 16 (11):1045–1052, November 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Al-Naami:2016:GMG

[ANSK16] Khaled Mohammed Al-Naami, Sadi Evren Seker, and Latifur Khan. GISQAF: MapReduce guided spatial query processing and analytics system. Software —Practice and Experience, 46(10):1329–1349, October 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anyanwu:1985:RSS

[Any85] J. A. Anyanwu. A reliable stable storage system for UNIX. Software —Practice and Experience, 15(10):973–990, October 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkins:1988:PMT

[AO88] M. Stella Atkins and Ronald A. Olsson. Performance of multi-tasking and synchronization mechanisms in the programming language SR. Software— Practice and Experience, 18(9):879–895, September 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Arthur:2012:PAS

[AO12] Richard B. Arthur and Dan R. Olsen, Jr. Privacyaware shared UI toolkit for nomadic environments. Software—Practice and Experience, 42(5):601–628, May 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ardo:1984:IPB

[AP84] Anders Ardö and Lars Philipson. Implementation of a Pascal based parallel language for a multiprocessor computer. Software —Practice and Experience, 14(7):643–657, July 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Allen:1985:VIL

[AP85] G. H. Allen and S. Perry.

A VLSI interactive layout editor (VILE). Software—Practice and Experience, 15(8):795–806, August 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atwood:1991:SBC

[AP91] J. William Atwood and Walter Prager. A STREAMSbased communications subsystem in Turing Plus for protocol development. Software—Practice and Experience, 21(12):1365– 1380, December 1991. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Asthagiri:1994:PCS

[AP94] Chandra R. Asthagiri and Jerry L. Potter. Parallel context-sensitive compilation. Software—Practice and Experience, 24(9): 801–822, September 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ashton:1995:TVE

[AP95] Paul Ashton and John Penny. A tool for visualizing the execution of interactions on a looselycoupled distributed system. Software—Practice and Experience, 25(10): 1117–1140, October 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Appel:1989:AL

[App89a] Andrew W. Appel. Allocation without locking. Software—Practice and Experience, 19(7):703–705, July 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Appel:1989:SGG

[App89b] Andrew W. Appel. Simple generational garbage collection and fast allocation. Software—Practice and Experience, 19(2):171– 183, February 1989. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Apple:2022:SYD

[App22] Jim Apple. Stretching your data with taffy filters. Software—Practice and Experience, 52(11):2349–2367, November 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aujla:2022:BDA

[APR22] Gagangeet Singh Aujla, Radu Prodan, and Danda B. Rawat. Big data analytics in Industry 4.0 ecosystems. Software—Practice and Experience, 52(3):639–641, March 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Agesen:1995:TIS

[APS95] Ole Agesen, Jens Palsberg, and Michael I. Schwartzbach. Type inference of SELF: Analysis of objects with dynamic and multiple inheritance. Software—Practice and Experience, 25(9):975– 995, September 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Adachi:2011:AOF

[APS^{+}11] Masakazu Adachi, Yiannis Papadopoulos, Septavera Sharvia, David Parker, and Tetsuya Tohdo. An approach to optimization of fault tolerant architectures using HiP-HOPS. Software—Practice and Experience, 41(11):1303–1327, October 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ammar:1993:VHP

[AR93] Reda A. Ammar and Carolyn Pe Rosiene. Visualizing a hierarchy of performance models for software systems. Software—Practice and Experience, 23(3): 293–315, March 1993. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abadi:2018:SCT

[ARA18] Reza Mohamadi Bahram Abadi, Amir Masoud Rahmani, and Sasan H. Alizadeh. Server consolidation techniques in virtualized data centers of cloud environments: a systematic literature review. Software—Practice and Experience, 48(9):1688–1726, September 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aldana:2006:BBT

[ARCN^{+}06] J. F. Aldana, M. Roldán-Castro, I. Navas, M. M. Roldán-García, M. Hidalgo-Conde, and O. Trelles. Bio-Broker: a tool for integration of biological data sources and data analysis tools. Software—Practice and Experience, 36(14): 1585–1604, November 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ardo:1987:EAR

[Ard87]  Anders Ardö.  Experi-

ence acquiring and retargeting a portable Ada computer. Software—Practice and Experience, 17(4):291– 307, April 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ayoubi:2021:AIS

[ARK21] Masoumeh Ayoubi, Mohammadreza Ramezanpour, and Reihaneh Khorsand. An autonomous IoT service placement methodology in fog computing. Software—Practice and Experience, 51(5):

1097–1120, May 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Arroba:2018:HMD

[ARMMA18] Patricia Arroba, José L. Risco-Martín, José M. Moya, and José L. Ayala. Heuristics and metaheuristics for dynamic management of computing and cooling energy in cloud data centers. Software— Practice and Experience, 48(10):1775–1804, October 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Adelstein:1994:DGL

[ARS^{+}94] Frank Adelstein, Golden Richard III, Loren Schwiebert, Rick Parent, and Mukesh Singhal. A distributed graphics library system. Software—Practice and Experience, 24(4):363–376, April 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Artym:1982:SME

[Art82] Richard Artym. The STAB multiprocessing environment for CYBA-M. Software—Practice and Experience, 12(4):323–329, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abramson:1977:TTO

[ARV77] Harvey Abramson, Tom Rushworth, and Ted Venema. TOSI: a tree oriented string interpreter for the design and implementation of semantics. Software—Practice and Experience, 7(6):663–670, November/December 1977. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Alcock:1973:MUB

[AS73] Donald Alcock and Brian Shearing. The MISTRESS user’s book. Software —Practice and Experience, 3(3):255–301, July/ September 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anderson:1978:RSS

[AS78] T. Anderson and S. K.

Shrivastava. Reliable software: a selective annotated bibliography. Software—Practice and Experience, 8(1):59–76, January/ February 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Apperley:1983:HDS

[AS83] Mark D. Apperley and R. Spence. Hierarchical dialogue structures in interactive computer systems. Software—Practice and Experience, 13(9):777– 790, September 1983. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Appel:1987:GSU

[AS87] Andrew W. Appel and Kenneth J. Supowit. Generalizations of the Sethi– Ullman algorithm for register allocation. Software —Practice and Experience, 17(6):417–421, June 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aho:1988:MCR

[AS88] Alfred V. Aho and Ravi Sethi. Maintaining cross references in manuscripts. Software—Practice and Experience, 18(1):1–13, January 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Agrawal:1997:IDF

[AS97a] Gagan Agrawal and Joel Saltz. Interprocedural data flow based optimizations for distributed memory compilation. Software—Practice and Experience, 27(5):519–545, May 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?

ID=7306;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7306&
PLACEBO=IE.pdf.

Akyurek:1997:ABR

[AS97b] Sedat Akyürek and Kenneth Salem. Adaptive block rearrangement under UNIX. Software—Practice and Experience, 27(1): 1–23, January 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7254;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7254&
PLACEBO=IE.pdf.

Allali:2008:MLM

[AS08] Julien Allali and Marie-France Sagot. A multiple layer model to compare RNA secondary structures. Software—Practice and Experience, 38(8):775– 792, July 10, 2008. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Asghar:2021:SES

[ASA^{+}21] Muhammad Zubair Asghar, Fazli Subhan, Hussain Ahmad, Wazir Zada Khan, Saqib Hakak, Thippa Reddy Gadekallu, and Mamoun Alazab. Senti-eSystem: a sentiment-based eSystemusing hybridized fuzzy and deep neural network for measuring customer satisfaction. Software—Practice and Experience, 51(3):571– 594, March 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Adhikari:2022:CSN

[ASA22] Mainak Adhikari, Satish Narayana Srirama, and Tarachand Amgoth. A comprehensive survey on nature-inspired algorithms and their applications in edge computing: Challenges and future directions. Software—Practice and Experience, 52 (4):1004–1034, April 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Al-Salman:2003:TCA

[ASAK03] AbdulMalik S. Al-Salman and Hend S. Al-Khalifa. Towards a computerized Arabic Braille environment. Software—Practice and Experience, 33(6):497– 508, May 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Al-Salman:2005:GOA

[ASAQ05] AbdulMalik S. Al-Salman and Fahad Al-Qahtani. GLAL: An OpenGL Arabization Library. Software —Practice and Experience, 35(14):1393–1407, November 25, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Alvarez-Sabucedo:2009:RWC

[ÁSARSG09] L. M. Álvarez-Sabucedo, L. E. Anido-Rifón, and J. M. Santos-Gago. Reusing Web contents: a DOM approach. Software—Practice and Experience, 39(3): 299–314, March 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anido:2001:MBL

[ASC^{+}01] Luis Anido, Juan Santos, Manuel Caeiro, Judith Rodríguez, Manuel J. Fernández, and Martín Llamas. Moving the business logic tier to the client. Cost-effective distributed computing for the WWW. Software—Practice and Experience, 31 (14):1331–1350, November 25, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
85512984/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=85512984&PLACEBO=IE.
pdf.

Amarnath:2009:OBG

[ASEB09] Balachandar R. Amarnath, Thamarai Selvi Somasundaram, Mahendran Ellappan, and Rajkumar Buyya. Ontology-based Grid resource management. Software—Practice and Experience, 39(17):1419–1438, December 10, 2009. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Arts:2023:TFR

[ASEF23] Thomas Arts, Hans Svensson, Clara Benac Earle, and Lars-Åke Fredlund. Testing feature-rich blockchains. Software—Practice and Experience, 53(5):1144– 1173, May 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aghazadeh:2023:PCC

[ASGA23] Rafat Aghazadeh, Ali Shahidinejad, and Mostafa Ghobaei-Arani. Proactive content caching in edge computing environment: a review. Software—Practice and Experience, 53(3):811– 855, March 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ashby:1973:DID

[ASH73] Gordon Ashby, Loren Salmonson, and Robert Heilman. Design of an interactive debugger for FORTRAN:MANTIS.

Software—Practice and Experience, 3(1):65–74, January/March 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sun:2024:IMR

[aSJW^{+}24] Chang ai Sun, Hui Jin, SiYi Wu, An Fu, ZuoYi Wang, and Wing Kwong Chan. Identifying metamorphic relations: a data mutation directed approach. Software—Practice and Experience, 54(3):394–418, March 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Azeem:2024:CSP

[ASME24] Muhammad Ilyas Azeem, Saad Shafiq, Atif Mashkoor, and Alexander Egyed. Code smells in pull requests: an exploratory study. Software—Practice and Experience, 54(3):419– 436, March 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Alqahtani:2019:SLA

[ASP^{+}19] Awatif Alqahtani, Ellis Solaiman, Pankesh Patel, Schahram Dustdar, and Rajiv Ranjan. Service level agreement specification for end-to-end IoT application ecosystems. Software—Practice and Experience, 49(12):1689–1711, December 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Austern:2003:UBS

[ASTW03] Matthew H. Austern, Bjarne Stroustrup, Mikkel Thorup, and John Wilkinson. Untangling the balancing and searching of balanced binary search trees. Software—Practice and Experience, 33(13): 1273–1298, November 10, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sun:2016:ECP

[aSZP^{+}16] Chang ai Sun, Yan Zhao, Lin Pan, Xiao He, and Dave Towey. Extended conference papers: A transformation-based approach to testing concurrent programs using UML activity diagrams. Software —Practice and Experience, 46(4):551–576, April 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1977:IMI

[Atk77] Malcolm P. Atkinson. IDL: a machine-independent data language. Software —Practice and Experience, 7(6):671–684, November/ December 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1978:BRB

[Atk78] M. P. Atkinson. Book review: An introduction to the study of programming languages, D. W. Barron, Cambridge University Press, 1977. No. of pages: 165. Price: £4.95 (hard cover), £2.50 (soft cover). Software—Practice and Experience, 8(4):497– 498, July 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1979:BRBa

[Atk79a] L. V. Atkinson. Book review: Microcomputer problem solving using Pascal, Kenneth L. Bowles, Springer Verlag, Berlin, 1977. No. of pages: 563. Price: DM21.40, U.S. $9.80. Software—Practice and Experience, 9(3):250– 252, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1979:BRBb

[Atk79b] L. V. Atkinson. Book review: Pascal: An introduction to methodical programming, W. Findally and D. W. Watt, Pitman, 1978. No. of pages: 306. Price: £4.95. Software—Practice and Experience, 9(6):511– 512, June 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1979:PSS

[Atk79c]  L. V. Atkinson.  Pas-

cal scalars as state indicators. Software—Practice and Experience, 9(6):427– 431, June 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1979:SIF

[Atk79d] L. V. Atkinson. Should if… then… else… follow the dodo? Software—Practice and Experience, 9(9): 693–700, September 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1982:OTS

[Atk82a] L. V. Atkinson. Optimizing two-state case statements in Pascal. Software —Practice and Experience, 12(6):571–581, June 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1982:BRB

[Atk82b] Laurence V. Atkinson. Book review: Foundations of programming with Pascal, Lawrie Moore, Ellis Horwood. No. of pages: 238. Price: £14.50 (Hardback) £5.00 (paperback). Software—Practice and Experience, 12(1):101– 102, January 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1983:BRB

[Atk83] L. V. Atkinson. Book review: Pocket guide to Pascal, David Watt, Pitman. No. of pages: 64. PRICE: £2.25. Pocket guide to programming, John Shelley, Pitman. No. of pages: 64. Price: £2.25. Software —Practice and Experience, 13(10):981, October 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Adak:2010:MBC

[ATO10] Mehmet Adak, Okan Topçu, and Halit Oguztüzün. Model-based code generation for HLA federates. Software—Practice and Experience, 40(2):149– 175, February ??, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Arciszewski:1984:PCF

[AV84] H. F. R. Arciszewski and E. M. Van Gasteren. P/CL: A flexible input processor. Software—Practice and Experience, 14 (12):1141–1148, December 1984. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Avvenuti:2005:MUJ

[AV05] Marco Avvenuti and Alessio Vecchio. MobileRMI: upgrading Java Remote Method Invocation towards mobility. Software—Practice and Experience, 35 (10):939–975, August 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aretz:1980:SSA

[AvdSGS80] F. E. J. Kruseman Aretz, Jan L. A. van de Snepscheut, H. Grasdijk, and J. M. H. Smeets. SATHE: Some aspects of an ALGOL implementation. Software —Practice and Experience, 10(7):563–573, July 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Amme:2009:EPS

[AvRAF09] Wolfram Amme, Jeffery von Ronne, Philipp Adler, and Michael Franz. The effectiveness of producerside machine-independent optimizations for mobile code. Software—Practice and Experience, 39(10): 923–946, July ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abuali:1993:ITC

[AW93] Faris N. Abuali and Roger L.

Wainwright. ISA[k] trees: a class of binary search trees with minimal or near minimal internal path length. Software—Practice and Experience, 23 (11):1267–1283, November 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Avritzer:1996:DWP

[AW96] Alberto Avritzer and Elaine J.

Weyuker. Deriving workloads for performance testing. Software—Practice and Experience, 26(6):613– 633, June 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16812.

Arlitt:2004:UWS

[AW04] Martin Arlitt and Carey Williamson. Understanding Web server configuration issues. Software— Practice and Experience, 34(2):163–186, February 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Arnatovich:2018:MAA

[AWNS18] Yauhen Leanidavich Arnatovich, Lipo Wang, Ngoc Minh Ngo, and Charlie Soh. Mobolic: an automated approach to exercising mobile application GUIs using symbiosis of online testing technique and customated input generation. Software—Practice and Experience, 48 (5):1107–1142, May 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2564.

Aycock:2015:SCS

[Ayc15] J. Aycock. Short communication: Stringlish: improved English string searching in binary files. Software—Practice and Experience, 45(11):1591– 1595, November 2015. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Augustin:2006:IAT

[AYdS^{+}06] I. Augustin, A. C. Yamin, L. C. da Silva, R. A. Real, G. Frainer, and C. F. R. Geyer. ISAMadapt: abstractions and tools for designing general-purpose pervasive applications. Software—Practice and Experience, 36(11–12):1231–1256, September/October 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Awad:1997:PAO

[AZ97a] Maher Awad and Jurgen Ziegler. A practical approach to object-oriented state modeling. Software—Practice and Experience, 27(3):311–328, March 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7293;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7293&
PLACEBO=IE.pdf.

Awad:1997:PAD

[AZ97b] Maher Awad and Jurgen Ziegler. A practical approach to the design of concurrency in objectoriented systems. Software—Practice and Experience, 27(9):1013–1034, September 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7331;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7331&
PLACEBO=IE.pdf.

Astekin:2019:DFD

[AZS19] Merve Astekin, Harun Zengin, and Hasan Sözer. DILAF: a framework for distributed analysis of large-scale system logs for anomaly detection. Software—Practice and Experience, 49(2):153–170, February 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barak:1978:SML

[BA78] Amnon B. Barak and Moshe Aharoni. A study of machine-level software profile. Software—Practice and Experience, 8(2):131– 136, March/April 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ben-Ari:1979:WYS

[BA79] Mordechai Ben-Ari. Why you should not timeshare. Software—Practice and Experience, 9(4):339– 340, April 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ben-Ari:1981:CCP

[BA81] Mordechai Ben-Ari. Cheap concurrent programming. Software—Practice and Experience, 11(12):1261– 1264, December 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ben-Ari:1986:FTD

[BA86] Mordechai Ben-Ari. FOREET: a tool for design and documentation of Fortran programs. Software—Practice and Experience, 16(10): 915–924, October 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bolognesi:1998:LTP

[BA98] Tommaso Bolognesi and Franco Accordino. A layer on top of Prolog for composing behavioural constraints. Software—Practice and Experience, 28(13):1415–1435, November 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1770;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1770&
PLACEBO=IE.pdf.

Badii:1998:SDO

[Bad98] Mehdi Badii. Simulating the DYNIX operating system parallel programming interface on a UNIX system. Software—Practice and Experience, 28(5):463– 480, May 25, 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1791;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1791&
PLACEBO=IE.pdf.

Baecker:1973:ARL

[Bae73] H. D. Baecker. Aspects of reference locality in list structures in virtual memory. Software—Practice and Experience, 3(3):245– 253, July/September 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bauer:2003:MSM

[BAF03] Lujo Bauer, Andrew W. Appel, and Edward W. Felten. Mechanisms for secure modular programming in Java. Software—Practice and Experience, 33(5): 461–480, April 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ben-Asher:1996:PEC

[BAFR96] Yosi Ben-Asher, Dror G. Feitelson, and Larry Rudolph. ParC — an extension of C for shared memory parallel processing. Software—Practice and Experience, 26(5):581–612, May 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16810.

Baird:1973:AVT

[Bai73] R. Baird. APET: a versatile tool for estimating computer application performance. Software —Practice and Experience, 3(4):385–395, October/December 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bailes:1985:DDD

[Bai85a] Paul A. Bailes. DDA — a data definition facility for UNIX using awk. Software—Practice and Experience, 15(10):1011–1020, October 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bailes:1985:LCI

[Bai85b] Paul A. Bailes. A low-cost implementation of coroutines for C. Software —Practice and Experience, 15(4):379–395, April 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bailey:1985:USL

[Bai85c] D. Bailey. The University of Salford LISP/PROLOG system. Software—Practice and Experience, 15(6):595– 609, June 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baker:2021:SEE

[BAJMT21] Thar Baker, Dhiya Al-Jumeily, Zakaria Maamar, and Zahir Tari. Semantic eSystems: Engineering methods, techniques, and tools. Software—Practice and Experience, 51(3):487– 488, March 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baker:1972:CSR

[Bak72]  John L. Baker.  CDC

6000-series register save/ restore. Software—Practice and Experience, 2(4):377– 387, October/December 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baker:2020:SFB

[BAM^{+}20] Thar Baker, Muhammad Asim, Áine MacDermott, Farkhund Iqbal, Faouzi Kamoun, Babar Shah, Omar Alfandi, and Mohammad Hammoudeh. A secure fog-based platform for SCADA-based IoT critical infrastructure. Software—Practice and Experience, 50(5):503–518, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Banton:1971:LRC

[Ban71] G. A. Banton. An online retail collection data system. Software—Practice and Experience, 1(3):235– 243, July/September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1987:DCA

[BAP87] Judy M. Bishop, Stephen R.

Adams, and David J. Pritchard. Distributing Concurrent Ada programs by source translation. Software—Practice and Experience, 17(12):859–884, December 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Boehm:1995:RAS

[BAP95] Hans-Juergen Boehm, Russ Atkinson, and Michael F. Plass. Ropes: An alternative to strings. Software—Practice and Experience, 25(12):1315–1330, December 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ballarin:2023:IAL

[BAPC23] Manuel Ballarín, Lorena Arcega, Vicente Pelechano, and Carlos Cetina. On the influence of architectural languages on requirements traceability. Software—Practice and Experience, 53(3):704–728, March 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1971:BRB

[Bar71] D. W. Barron. Book review: APL 360—An interactive approach, L. Gilman and A. J. Rose, Wiley, New York, 1970. No. of pages:

335. Price: £3.10. Software—Practice and Experience, 1(3):301–302, July/ September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1972:BRBb

[Bar72a] D. W. Barron. Book review: IFIP Guide to Concepts and Terms in Data Processing, I. H. Gould (Ed.), North-Holland, London, 1971. No. of pages:

161. Price: $7.00. Software—Practice and Experience, 2(3):306, July 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1972:BRBc

[Bar72b] D. W. Barron. Book review: Modular programming, J. Maynard, Butterworths, London, 1971. No. of pages: 100. Price: £1.50. Software—Practice and Experience, 2(4):406– 407, October 1972. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1972:BRBa

[Bar72c] D. W. Barron. Book review: Standard Fortran programming manual, The National Computing Centre Limited, Manchester. No. of pages: 152. Price: £3. Software—Practice and Experience, 2(1):98– 99, January 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1973:BRBd

[Bar73a] D. W. Barron. Book review: Computation and communication, G. Emery, University of Wales Press, 1972. No. of pages: 24. Price: 25p. Software— Practice and Experience, 3(3):304, July/September 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1973:BRBc

[Bar73b] D. W. Barron. Book review: Computer science: The PL/I language, A. L. Anger, Wiley, New York, 1972. No. of pages:

133. Price: £2.55. Software—Practice and Experience, 3(3):303–304, July/ September 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1973:BRBb

[Bar73c] D. W. Barron. Book review: Designing systems programs, R. Gauthier and S. Ponto, Prentice-Hall, Englewood Cliffs, N.J., 1971. No. of pages: 274. Price: £6.25. Software— Practice and Experience, 3(2):189–190, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1973:BRBe

[Bar73d] D. W. Barron. Book review: FORTRAN techniques with special reference to non-numerical applications, A. C. Day, Cambridge University Press, 1972. No. of pages: 96. Price: £2.50. Software —Practice and Experience, 3(4):404, October/ December 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1973:BRBa

[Bar73e] D. W. Barron. Book review: Software 71: Proceedings of a conference held in July 1971, Transcripta Books, London, 1972. No. of pages: 182. Price: £5.50. Software— Practice and Experience, 3 (2):189, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1973:EN

[Bar73f] D. W. Barron. Editor’s note. Software—Practice and Experience, 3(2):187, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barker:1974:PPS

[Bar74a] P. G. Barker. POP-2 performance on a small 1900 machine. Software—Practice and Experience, 4(3): 265–273, July/September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barnes:1974:CSR

[Bar74b] J. G. P. Barnes. Comments and strings in RTL/2. Software—Practice and Experience, 4(4):401–402, October/December 1974. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1974:BRBd

[Bar74c] D. W. Barron. Book review: Annual review in automatic programming, Volume 7,Part 5: A survey of extensible programming languages, N. Solnsteff and A Yezerki, Pergamon, Oxford, 1974. No. of pages:

46. Price: £1.65. Software—Practice and Experience, 4(4):407–408, October/December 1974. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1974:BRBb

[Bar74d] D. W. Barron. Book review: Debugging techniques in large systems, R. Rustin (Ed.), Prentice-Hall, Englewood Cliffs, N.J., 1971. No. of pages: 148. Price: £4.00. Software—Practice and Experience, 4(4):404– 405, October/December 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1974:BRBa

[Bar74e] D. W. Barron. Book review: Informal introduction to ALGOL 68, C. H. Lindsey and S. G. van der Meulen. North-Holland, Amsterdam, 1971. No. of pages: 382. Price: £7.00. Also available in paperback. Software—Practice and Experience, 4(3):304, July/September 1974. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1974:BRBc

[Bar74f]  D. W. Barron.  Book

review: Information representation and manipulation in a computer, E. S. Page and L. B. Wilson, Cambridge University Press, 1973. No. of pages:

244. Price: £1.90. Software—Practice and Experience, 4(4):407, October/ December 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1974:EYC

[Bar74g] D. W. Barron. Editorial: Is your computer really necessary? Software—Practice and Experience, 4(2):115, April/June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1974:ETC

[Bar74h] D. W. Barron. Editorial:

The two cultures. Software—Practice and Experience, 4(4):307, October/ December 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barth:1974:NCS

[Bar74i] C. Wrandle Barth. Notes on the case statement. Software—Practice and Experience, 4(3):289–298, July/ September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1975:BRBa

[Bar75a] D. W. Barron. Book review: Algol 68-R Users Guide (2nd edn), P. M. Woodward and S. G. Bond, H.M.S.O., 1974.

No. of pages: 99. Price: £0.90. Software—Practice and Experience, 5(2): 217, April 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1975:BRBf

[Bar75b]  D. W. Barron.  Book

review: Assemblers and BAL, I. Flores, Prentice-Hall, Englewood Cliffs, N.J., 1971, No. of pages:

420. Price £6.00. Software—Practice and Experience, 5(4):407–408, October 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1975:BRBb

[Bar75c] D. W. Barron. Book review: Introduction to Digital Computer Plotting, T. C. Smith and Y. C. Pao, Gordon and Breach, New York, 1973. No. of pages:

86. Price: £2.50. Software—Practice and Experience, 5(3):314, July 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1975:BRBe

[Bar75d] D. W. Barron. Book review: Job Control Language and File Definition, I. Flores, Prentice-Hall, Englewood Cliffs, N.J., 1971. No. of pages: 268.

Price: £6.25. Software— Practice and Experience, 5 (4):407, October 1975. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1975:BRBc

[Bar75e] D. W. Barron. Book review: Software 73 Conference Proceedings, Transcripta Books, London, 1974. No. of pages: 225. Price: £6.50. Software— Practice and Experience, 5 (3):314, July 1975. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1975:BRBd

[Bar75f] D. W. Barron. Book review: Structured Programming, O.-J. Dahl, E. W. Dijkstra and C. A. R. Hoare, Academic Press, London, 1972. No. of pages: 220. Price: £4.20. Software —Practice and Experience, 5(3):314–315, July 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1976:BRBa

[Bar76a] D. W. Barron. Book review: A laboratory manual for compiler and operating system implementation, M. H. Halstead, American Elsevier, New York, 1975. No. of pages: 108. Price: Dfl 19.00 (about $7.95). Software—Practice and Experience, 6(1):136, January 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1976:BRBc

[Bar76b]  D. W. Barron.  Book

review: High Level Languages: Infotech State of the Art Report No. 7, Ed. C. Boon, Infotech Information Ltd., Maidenhead, 1972. No. of pages: 500. Price: £40. Software —Practice and Experience, 6(2):285–286, April 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1976:BRBd

[Bar76c] D. W. Barron. Book review: Program style, design, efficiency, debugging and testing, D. van Tassel, Prentice-hall Inc., Englewood Cliffs, N.J., 1975. No. of pages: 256. Price: £7.80. Software—Practice and Experience, 6(4):582– 583, October 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1976:BRBb

[Bar76d]  D. W. Barron.  Book

review: Structured Programming in PL/C — an Abecedarian, G. M. Weinberg, N. Yasukawa and R.

Marcus, Wiley, New York, 1973. No. of pages: 220. Price £3.75. Software— Practice and Experience, 6(2):284–285, April 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1976:BRBe

[Bar76e] D. W. Barron. Book review: The Mythical Man-Month—Essays on Software Engineering, F. P. Brooks, Jr., Addison-Wesley, Reading, Mass., 1975. No. of pages: 195. Price: £3.30. Software— Practice and Experience, 6 (4):583–584, October 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barnes:1977:LE

[Bar77a] J. G. P. Barnes. Letters to the editor. Software— Practice and Experience, 7 (1):148–149, January 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1977:BRBc

[Bar77b]  D. W. Barron.  Book

review: Computing: An introduction to procedures and procedure followers, F. M. Tonge and J. Feldman, McGraw-Hill, New York, 1975. no. of pages: 302. price: £7.85. Software— Practice and Experience, 7 (2):304, March 1977. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1977:BRBd

[Bar77c] D. W. Barron. Book review: Introduction to Pascal, C. A. G. Webster, Heyden, London, 1976. No. of pages: 129. Price: £5-50. Software—Practice and Experience, 7(4):539– 541, July 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1977:BRBb

[Bar77d] D. W. Barron. Book review: Principles of program design, M. A. Jackson, academic press, London, 1975. No. of pages:

297. Price: £8.80. Software —Practice and Experience, 7(2):299–300, March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1977:BRBa

[Bar77e] D. W. Barron. Book review: The design of an optimizing compiler, W. Wulf, R. K. Johnsson, C. B. Weinstock, S. O. Hobbs and C. M. Geschke, American Elsevier, New York, 1975. No. of pages: 165. Price: Dfl 23.00 (about $9.75). Software—Practice and Experience, 7(1):152, January 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1977:ENE

[Bar77f] D. W. Barron. Editor’ note: Editor’ note. Software—Practice and Experience, 7(1):152, January 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barnett:1978:HRF

[Bar78a] J. K. R. Barnett. A highly reliable file system which supports multiprocessing. Software—Practice and Experience, 8(6):645– 672, November/December 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1978:BRBa

[Bar78b] D. W. Barron. Book review: RTL/2 design and philosophy, J. G. P. Barnes, Heyden, London, 1976. No. of pages: 176. Price: £5.50. Software—Practice and Experience, 8(1):110– 111, January 1978. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1978:BRB

[Bar78c] D. W. Barron. Book review: The ICL 2900 Series J. K. Buckle, Macmillan, London, 1978. No. of pages:

83. Price £3.50. Software —Practice and Experience, 8(6):759, November 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1978:BRBb

[Bar78d]  D. W. Barron.  Book

review: The universal elixir, and other computing projects which failed, R. L. Glass, Computing Trends, 3325 Lakewood Avenue South, Seattle, Washington 98144, U.S.A., 1977. No. of pages:

88. Price $7.50 + $1.50 postage/handling. Software—Practice and Experience, 8(5):642, September 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1979:BRBb

[Bar79a]  D. W. Barron.  Book

review: Digital computer fundamentals, Thomas C. Bartee, McGraw-Hill, New York, 1977. No. of pages:

563. Price: £11.70. Software—Practice and Experience, 9(3):250, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1979:BRBa

[Bar79b] D. W. Barron. Book review: Structured programming— infotech state of the art report, Infotech International Ltd., Maidenhead, 1976. No. of pages: 495. Price: $185. Software—Practice and Experience, 9(2):172, February 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barnes:1980:OA

[Bar80a] J. G. P. Barnes. An overview of Ada. Software —Practice and Experience, 10(11):851–887, November 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barnes:1980:SR

[Bar80b] J. G. P. Barnes. The standardization of RTL/2. Software—Practice and Experience, 10(9):707–719, September 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barnett:1980:DIT

[Bar80c] J. K. R. Barnett. The design of an inter-task communication scheme. Software—Practice and Experience, 10(10):801–816, October 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1980:BRBa

[Bar80d] D. W. Barron. Book review: Compatible FOR-TRAN, A. Colin Day. Cambridge University Press, 1979. No. of pages: 107. Price: £5.95. Software— Practice and Experience, 10(10):850, October 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1980:BRBb

[Bar80e] D. W. Barron. Book review: Etudes for programmers, Charles Wetherell, Prentice-Hall, Englewood Cliffs, 1978. No. of pages:

200. Price: £9.45. Software—Practice and Experience, 10(11):952, November 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1981:BRB

[Bar81] D. W. Barron. Book review: BCPL — The language and its compiler. Martin Richards and Colin Whitby-Strevens, Cambridge University Press 1980. No. of pages: 173. Price £9.50. ISBN 0-521-21965-5. Software—Practice and Experience, 11(5): 531–532, May 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1982:BRBb

[Bar82a] D. W. Barron. Book review: Tales of computing folk, Robert L. Glass, Computing Trends, 6925 56th Ave S., Seattle, Wa 98118, U.S.A., 1978. No of pages: 96, Price $8.95 post paid U.S.A. $9.95 post paid elsewhere. Software—Practice and Experience, 12(5): 495, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1982:BRBa

[Bar82b] D. W. Barron. Book review: The British lectures, A. P. Ershov, Heyden, London, 1980. No. of pages:

80. Price £7.00. Software—Practice and Experience, 12(3):295, March 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1982:BRBc

[Bar82c] D. W. Barron. Book review: The second coming: More computing projects which failed, Robert L. Glass and Sue de Nim, Computing Trends, 6925 56th Ave S., Seattle, WA 98118, U.S.A., 1980. No. of pages: 88. Price $8.95 post paid U.S.A., $9.95 post paid elsewhere. Software—Practice and Experience, 12(5):496, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1983:BRB

[Bar83a] D. W. Barron. Book review: NCC directory of computer software 1982. National Computing Centre, 1982. Price: £30.00. NCC directory of computer hardware 1982, National Computing Centre, 1982. price: £30.00. Software —Practice and Experience, 13(10):979–980, October 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barton:1983:DPS

[Bar83b] D. Barton. Distributed processing in a small office. Software—Practice and Experience, 13(6):527– 536, June 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1984:BRBb

[Bar84a] D. W. Barron. Book review: Computing Catastrophes, by Robert L. Glass, Computing Trends, 6925 56 Ave. S. Seattle WA98118, USA, 1983. No. of pages: 144. Price: $11 (postpaid), ($13 outside USA). Software—Practice and Experience, 14(2):196, February 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1984:BRBa

[Bar84b] D. W. Barron. Book review: The Power of the Peonage, Robert L. Glass, Computing Trends, 6925 56th Avenue South, Seattle, WA 98118, U.S.A. No. of pages: 101. Price: $8.95 ($9.95 outside U.S.A.) post paid. Software—Practice and Experience, 14(2):195, February 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1984:E

[Bar84c]  David Barron.  Edito-

rial. Software—Practice and Experience, 14(1):1, January 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bartoli:1997:NAM

[Bar97] Alberto Bartoli. A novel approach to marshalling. Software—Practice and Experience, 27(1):63–85, January 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7258;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7258&
PLACEBO=IE.pdf.

Barros:2015:AOP

[Bar15] Fernando J. Barros. Aspectoriented programming and pluggable software units: a comparison based on design patterns. Software—Practice and Experience, 45(3):289–314, March 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Basden:2000:STN

[Bas00] Andrew Basden. Some technical and non-technical issues in implementing a knowledge server. Software—Practice and Experience, 30(10):1127–1164, August 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72510351/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72510351&PLACEBO=IE.
pdf.

Bate:1974:DII

[Bat74] D. G. Bate. Design and implementation of an interactive test bed. Software—Practice and Experience, 4(1):91–109, January/March 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bergeron:1975:TEU

[BB75] R. Daniel Bergeron and Henri R. Bulterman. A technique for evaluation of user systems on an IBM S/370. Software—Practice and Experience, 5(1):83– 92, January/March 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barnett:1981:CFS

[BB81] Richard Barnett and Richard C. Beckwith. The community file station. Software—Practice and Experience, 11(10):1001–1008, October 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Breuer:1995:PCC

[BB95] Peter T. Breuer and Jonathan P. Bowen. A PREttier compiler-compiler: Generating higher-order parsers in C. Software—Practice and Experience, 25(11):1263–1297, November 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baker:1999:CCC

[BB99a] Mark Baker and Rajkumar Buyya. Cluster computing: the commodity supercomputer. Software—Practice and Experience, 29(6):

551–576, May 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61000492;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=61000492&PLACEBO=IE.
pdf.

Boufaida:1999:MLA

[BB99b] M. Boufaida and P. Barril. A multi-layered architecture for managing graphical interfaces and tutoring mixed-strategy dialogues. Software—Practice and Experience, 29(13):1223– 1239, November 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
66501671/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=66501671&PLACEBO=IE.
pdf.

Bainbridge:2003:MNC

[BB03] David Bainbridge and Tim Bell. A music notation construction engine for optical music recognition. Software—Practice and Experience, 33(2):173–200, February 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
102525918/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=102525918&PLACEBO=
IE.pdf.

Berdajs:2010:EAU

[BB10] J. Berdajs and Z. Bosnić.

Extending applications using an advanced approach to DLL injection and API hooking. Software—Practice and Experience, 40(7): 567–584, June 2010. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Budgen:2011:EEA

[BBB^{+}11] D. Budgen, A. J. Burn, O. P. Brereton, B. A. Kitchenham, and R. Pretorius. Empirical evidence about the UML: a systematic literature review. Software—Practice and Experience, 41(4):363– 392, April 10, 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baxter:2022:CEB

[BBB^{+}22] Amanda L. Baxter, Segev Y.

BenZvi, Walter Bonivento, Adam Brazier, Michael Clark, Alexis Coleiro, David Collom, Marta Colomer-Molla, Bryce Cousins, Aliwen Delgado Orellana, Damien Dornic, Vladislav Ekimtcov, Shereen El-Sayed, Andrea Gallo Rosso, Patrick Godwin, Spencer Griswold, Alec Habig, Remington Hill, Shunsaku Horiuchi, D. Andrew Howell, Margaret W. G. Johnson, Mario Juri’c, James P. Kneller, Abigail Kopec, Claudio Kopper, Vladimir Kulikovskiy, Mathieu Lamoureux, Rafael F. Lang, Shengchao Li, Massimiliano Lincetto, Lindy Lindstrom, Mark W. Linvill, Curtis McCully, Jost Migenda, Danny Milisavljevic, Spencer Nelson, Rita Novoseltseva, Erin O’Sullivan, Donald Petravick, Barry W. Pointon, Nirmal Raj, Andrew Renshaw, Janet Rumleskie, Tom Sonley, Ron Tapia, Jeffrey C. L. Tseng, Christopher D. Tunnell, Godefroy Vannoye, Carlo F. Vigorito, Clarence J. Virtue, Christopher Weaver, Kathryn E. Weil, Lindley Winslow, Rich Wolski, Xun-Jie Xu, Yiyang Xu, The Scimma, and Snews Collaborations. Collaborative experience between scientific software projects using Agile Scrum development. Software— Practice and Experience, 52(10):2077–2096, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baldassari:1991:POO

[BBC91] Marco Baldassari, Giorgio Bruno, and Andrea Castella. PROTOB: An object-oriented CASE tool for modelling and prototyping distributed systems. Software—Practice and Experience, 21(8):823– 844, August 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bellotti:2004:EOM

[BBG04] Francesco Bellotti, Riccardo Berta, and Alessandro De Gloria. Evaluation and optimization of method calls in Java. Software—Practice and Experience, 34(4):395–431, April 10, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bellotti:2001:DJA

[BBGP01] Francesco Bellotti, Riccardo Berta, Alessandro De Gloria, and Andrea Poggi. DirectJ: Java APIs for optimized 2D graphics. Software—Practice and Experience, 31(3):259–275, March 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
76507637/START;   http:
//www3.interscience.wiley.

                                        com/cgi-bin/fulltext?
                                        ID=76507637&PLACEBO=IE.
                                        pdf.

Bermbach:2022:AUA

[BBH^{+}22] David Bermbach, Jonathan Bader, Jonathan Hasenburg, Tobias Pfandzelter, and Lauritz Thamsen. AuctionWhisk: Using an auction-inspired approach for function placement in serverless fog platforms. Software—Practice and Experience, 52(5): 1143–1169, May 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Birkenheuer:2012:VHC

[BBK^{+}12] Georg Birkenheuer, André Brinkmann, Jürgen Kaiser, Axel Keller, Matthias Keller, Christoph Kleineweber, Christoph Konersmann, Oliver Niehörster, Thorsten Schäfer, Jens Simon, and Maximilian Wilhelm. Virtualized HPC: a contradiction in terms? Software—Practice and Experience, 42(4):485– 500, April 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baker:2002:GGT

[BBL02] Mark Baker, Rajkumar Buyya, and Domenico Laforenza. Grids and Grid technologies for widearea distributed computing. Software—Practice and Experience, 32 (15):1437–1466, December 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
98518912/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=98518912&PLACEBO=IE.
pdf.

Bresnahan:1984:WNC

[BBM84] J. Barrie Bresnahan, David T.

Barnard, and Ian A. Macleod. WSH: a new command interpreter for UNIX. Software—Practice and Experience, 14 (12):1197–1205, December 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bellotti:2008:ORB

[BBMG08] Francesco Bellotti, Riccardo Berta, Massimiliano Margarone, and Alessandro De Gloria. oDect: an RFID-based object detection API to support applications development on mobile devices. Software—Practice and Experience, 38(12):1241–1259, October ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Beheshti:2018:PSE

[BBMN18] Amin Beheshti, Boualem Benatallah, and Hamid Reza Motahari-Nezhad. ProcessAtlas: a scalable and extensible platform for business process analytics. Software—Practice and Experience, 48(4):842– 866, April 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2558.

Bergel:2012:EPB

[BBRB12] Alexandre Bergel, Felipe Bañados, Romain Robbes, and Walter Binder. Execution profiling blueprints. Software—Practice and Experience, 42(9):1165– 1192, September 2012. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bosch:2011:IAC

[BBS11] Jan Bosch and Petra M. Bosch-Sijtsema. Introducing agile customer-centered development in a legacy software product line. Software—Practice and Experience, 41(8):871–882, July 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bruneau:2013:ECP

[BC13] Julien Bruneau and Charles Consel. Extended conference papers: DiaSim: a simulator for pervasive computing applications. Software—Practice and Experience, 43(8):885– 909, August 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bispo:2017:MSC

[BC17] João Bispo and João M. P.

Cardoso. A MATLAB subset to C compiler targeting embedded systems. Software—Practice and Experience, 47(2):249–272, February 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:2021:ISI

[BC21] Judith Bishop and Kendra M. L. Cooper. Introduction to the Special Issue on Software Engineering in Practice. Software—Practice and Experience, 51(4):689– 690, April 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bacocco:2022:DIN

[BCC^{+}22] Duilio Luca Bacocco, Eugenio Carrani, Bruno Ciciani, Pierangelo Di Sanzo, Francesco Leotta, and Marina Torre. Design and implementation of the new Italian healthcare digital interoperable registry for implantable medical devices. Software—Practice and Experience, 52 (11):2368–2392, November 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bertino:2000:IAT

[BCF00] E. Bertino, B. Catania, and A. Filippone. An index allocation tool for objectoriented database systems. Software—Practice and Experience, 30(9):973– 1002, July 25, 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72503901/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72503901&PLACEBO=IE.
pdf.

Brandis:1995:OSF

[BCFT95] Marc M. Brandis, R. Crelier, Michael Franz, and Josef Templ. The Oberon system family. Software—Practice and Experience, 25(12):1331–1366, December 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Banatre:1981:EIA

[BCHR81] M. Banatre, A. Couvert, D. Herman, and M. Raynal. An experience in implementing abstract data types. Software—Practice and Experience, 11(3):315– 320, March 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Briggs:1998:PIC

[BCHS98] Preston Briggs, Keith D.

Cooper, Timothy J. Harvey, and L. Taylor Simpson. Practical improvements to the construction and destruction of static single assignment form. Software—Practice and Experience, 28(8):859– 881, July 10, 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1814;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1814&
PLACEBO=IE.pdf.

Bell:1994:UPW

[BCL^{+}94] Brigham Bell, Wayne Citrin, Clayton Lexis, John Rieman, Robert Weaver, Nick Wilder, and Benjamin G. Zorn. Using the programming walkthrough to aid in programming language design. Software —Practice and Experience, 24(1):1–25, January 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bruneton:2006:FCM

[BCL^{+}06] Eric Bruneton, Thierry Coupaye, Matthieu Leclercq, Vivien Quéma, and Jean-Bernard Stefani. The FRACTAL component model and its support in Java. Software—Practice and Experience, 36(11– 12):1257–1284, September/ October 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Berbecaru:2013:FSO

[BCL13] Diana Berbecaru, Matteo M. Casalino, and Antonio Lioy. FcgiOCSP: a scalable OCSP-based certificate validation system exploiting the FastCGI interface. Software—Practice and Experience, 43 (12):1489–1518, December 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brisaboa:2007:CPL

[BCLF^{+}07] N. R. Brisaboa, J. A. Cotelo-Lema, A. Fariña, M. R. Luaces, J. R. Parama, and J. R. R. Viqueira. Collecting and publishing large multiscale geographic datasets. Software—Practice and Experience, 37(12):1319–1348, October 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1971:SCP

[BCP71] F. D. Brown, Valerie J.

Calderbank, and M. D. Poole. Some comments on the portability of a large ALGOL program — the implementation of SID on KDF9. Software—Practice and Experience, 1(4):367– 371, October/December 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barringer:1979:PCS

[BCP79] H. Barringer, P. C. Capon, and R. Philips. The portable compiling systems of MUSS. Software—Practice and Experience, 9(8): 645–655, August 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bosch:2019:ISI

[BCP19] Jan Bosch, Kendra M. L.

Cooper, and Frances Paulisch. Introduction to the special issue on software engineering in practice. Software —Practice and Experience, 49(2):151–152, February 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baca:2013:ISS

[BCPL13] Dejan Baca, Bengt Carlsson, Kai Petersen, and Lars Lundberg. Improving software security with static automated code analysis in an industry setting. Software—Practice and Experience, 43(3):259–279, March 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brisaboa:2018:SPA

[BCPSC18] Nieves R. Brisaboa, Ricardo Cao, José R. Paramá, and Fernando Silva-Coira. Scalable processing and autocovariance computation of big functional data. Software—Practice and Experience, 48(1):123–140, January 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bufano:2022:PGD

[BCRF22] Roberto Bufano, Gennaro Costagliola, Mattia De Rosa, and Vittorio Fuccella. PolyRec Gesture Design Tool : a tool for fast prototyping of gesture-based mobile applications. Software—Practice and Experience, 52(2): 594–618, February 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Briggs:1997:VN

[BCS97] Preston Briggs, Keith D.

Cooper, and L. Taylor Simpson. Value numbering. Software—Practice and Experience, 27(6):701– 724, June 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7312;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7312&
PLACEBO=IE.pdf.

Boldi:2004:USF

[BCSV04] Paolo Boldi, Bruno Codenotti, Massimo Santini, and Sebastiano Vigna. UbiCrawler: a scalable fully distributed Web crawler. Software—Practice and Experience, 34(8): 711–726, July 10, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:2020:ISI

[BCSW20] Judith Bishop, Kendra M. L. Cooper, Helen Sharp, and Michael Whalen. Introduction to the special issue on software engineering in practice. Software—Practice and Experience, 50(3):187–188, March 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bettini:2006:DDC

[BCV06] Lorenzo Bettini, Sara Capecchi, and Betti Venneri. Double dispatch in C++. Software—Practice and Experience, 36(6):581– 613, May 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bron:1976:PCP

[BD76] C. Bron and W. De Vries. A PASCAL compiler for PDP 11 minicomputers. Software—Practice and Experience, 6(1):109– 116, January/March 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bruns:2014:TPB

[BD14] Ralf Bruns and Jürgen Dunkel. Towards patternbased architectures for event processing systems. Software—Practice and Experience, 44(11):1395– 1416, November 2014. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bhatt:2020:SAB

[BDA20] Ashutosh Bhatt, Priti Dimri, and Ambika Aggarwal. Self-adaptive brainstorming for jobshop scheduling in multicloud environment. Software— Practice and Experience, 50(8):1381–1398, August 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brinkmann:2009:RUC

[BDD09] Gunnar Brinkmann, Jan Degraer, and Karel De Loof. Rehabilitation of an unloved child: semisplaying. Software—Practice and Experience, 39(1): 33–45, January ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Biliris:1993:MCO

[BDG93] Alexandros Biliris, Shaul Dar, and Narain H. Gehani. Making C++ objects persistent: the hidden pointers. Software—Practice and Experience, 23 (12):1285–1303, December 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Berry:2000:LSG

[BDG^{+}00] J. Berry, Nathaniel Dean, Mark K. Goldberg, Gregory E. Shannon, and Steven Skiena. LINK: a system for graph computation. Software—Practice and Experience, 30 (11):1285–1302, September 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.

                                        com/cgi-bin/abstract/
                                        72516934/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=72516934&PLACEBO=IE.
                                        pdf.

Boom:1980:CCS

[BdJ80] H. J. Boom and E. de Jong. A critical comparison of several programming language implementations. Software—Practice and Experience, 10(6):435–473, June 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Berbecaru:2009:UFS

[BDL09] Diana Berbecaru, Amarkumar Desai, and Antonio Lioy. A unified and flexible solution for integrating CRL and OCSP into PKI applications. Software —Practice and Experience, 39(10):891–921, July ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Behrmann:2011:DUY

[BDL^{+}11] Gerd Behrmann, Alexandre David, Kim Guldstrand Larsen, Paul Pettersson, and Wang Yi. Developing Uppaal over 15 years. Software—Practice and Experience, 41(2):133– 142, February 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bernardeschi:2004:CSI

[BDLM04] C. Bernardeschi, N. De Francesco, G. Lettieri, and L. Martini. Checking secure information flow in Java bytecode by code transformation and standard bytecode verification. Software—Practice and Experience, 34(13): 1225–1255, November 10, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Buttazzo:2016:DAT

[BDM16] Giorgio Buttazzo, Carmelo Di Franco, and Mauro Marinoni. Design and analysis of target-sensitive real-time systems. Software—Practice and Experience, 46(9):1181–1200, September 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Briola:2017:AOO

[BDMP17] Daniela Briola, Vincenzo Deufemia, Viviana Mascardi, and Luca Paolino. Agent-oriented and ontologydriven digital libraries: the IndianaMAS experience. Software—Practice and Experience, 47 (11):1773–1799, November 2017. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Bettini:2002:KJP

[BDP02] Lorenzo Bettini, Rocco De Nicola, and Rosario Pugliese. Klava: a Java package for distributed and mobile applications. Software—Practice and Experience, 32(14):1365–1394, November 25, 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
98518911/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=98518911&PLACEBO=IE.
pdf.

Boyer:2014:FAR

[BdPGS14] Fabienne Boyer, Noel de Palma, Olivier Gruber, and Sylvain Sicard. Full autonomic repair for distributed applications. Software—Practice and Experience, 44(9):1027–1045, September 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baresi:2023:MFA

[BDQR23] Luciano Baresi, Tommaso Dolci, Giovanni Quattrocchi, and Nicholas Rasi. A multi-faceted analysis of the performance variability of virtual machines. Software—Practice and Experience, 53(11):2067–2091, November 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Buhr:1992:COO

[BDS^{+}92] Peter A. Buhr, Glen Ditchfield, Richard A. Stroobosscher, B. M. Younger, and C. Robert Zarnke. μC++: Concurrency in the object-oriented language C++. Software—Practice and Experience, 22(2):137– 172, February 1992. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barbuti:1999:LTR

[BDSV99] Roberto Barbuti, Nicoletta De Francesco, Antonella Santone, and Gigliola Vaglini. LORETO: a tool for reducing state explosion in verification of LOTOS programs. Software—Practice and Experience, 29(12):1123–1147, October 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
66000402/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=66000402&PLACEBO=IE.
pdf.

Bernstein:1981:MBL

[BE81] Arthur J. Bernstein and J. R. Ensor. A Modula based language supporting hierarchical development and verification. Software—Practice and Experience, 11(3):237–255, March 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Burrows:2002:JGE

[BE02] Anthony L. Burrows and David England. Java 3D, 3D graphical environments and behaviour. Software—Practice and Experience, 32(4):359–376, April 10, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        91016434/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=91016434&PLACEBO=IE.
                                        pdf.

Beaumont:1978:ISM

[Bea78] W. P. Beaumont. An implementation of structured multiprogramming. Software—Practice and Experience, 8(3):313–322, May 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Beckman:1991:SLL

[Bec91] Brian Beckman. A scheme for little languages in interactive graphics. Software —Practice and Experience, 21(2):187–207, February 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Beech:1982:MCL

[Bee82]  David Beech.  Modu-

larity of computer languages. Software—Practice and Experience, 12(10): 929–958, October 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bell:1974:RCS

[Bel74] James R. Bell. Rapid calculations of subscripted array addresses. Software—Practice and Experience, 4(3):275–277, July/ September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Belli:2014:HAM

[BELS14] Fevzi Belli, Andre Takeshi Endo, Michael Linschulte, and Adenilso Simao. A holistic approach to modelbased testing of Web service compositions. Software —Practice and Experience, 44(2):201–234, February 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Benediktsson:1977:SFP

[Ben77]  O. Benediktsson.  Se-

quential file processing in Fortran IV. Software— Practice and Experience, 7(5):655–659, September/ October 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bengtson:1989:MVM

[Ben89] Neal M. Bengtson. Microcomputers vs. mainframe simulations: a case study. Software—Practice and Experience, 19(10): 957–965, October 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bennett:1990:EDS

[Ben90] John K. Bennett. Experience with distributed Smalltalk. Software—Practice and Experience, 20(2): 157–180, February 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Berry:1978:EPP

[Ber78]  R. E. Berry.  Experi-

ence with the Pascal Pcompiler. Software—Practice and Experience, 8 (5):617–627, September/ October 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bersohn:1982:RTM

[Ber82]  Malcolm Bersohn.  On

the removal of two major obstacles to a linear flow of inspection in a program. Software—Practice and Experience, 12(1):96– 98, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bernstein:1985:PGC

[Ber85a] Robert L. Bernstein. Producing good code for the case statement. Software—Practice and Experience, 15(10):1021–1024, October 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See correction [KP94].

Bertsch:1985:OES

[Ber85b] Eberhard Bertsch. Optimization of expression syntax: an experimental approach. Software—Practice and Experience, 15(3):269– 276, March 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bernstein:1986:MIC

[Ber86] Robert L. Bernstein. Multiplication by integer constants. Software—Practice and Experience, 16(7):641– 652, July 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Berztiss:1988:PG

[Ber88] Alfs Berztiss. Programming with generators. Software—Practice and Experience, 18(1):73–81, January 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Berry:1999:SLS

[Ber99] Daniel M. Berry. Stretching letter and slantedbaseline formatting for Arabic, Hebrew, and Persian with ditroff/ffortid and dynamic PostScript fonts. Software—Practice and Experience, 29 (15):1417–1457, December 25, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
68501289/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=68501289&PLACEBO=IE.
pdf.

Berkani:2020:SSB

[Ber20] Lamia Berkani. A semantic and social-based collaborative recommendation of friends in social networks. Software—Practice and Experience, 50(8): 1498–1519, August 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bogott:1975:EMP

[BF75] Robert P. Bogott and Mark A. Franklin. Evaluation of Markov program models in virtual memory systems. Software—Practice and Experience, 5(4):337–346, October/December 1975. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barach:1980:NEI

[BF80] David R. Barach and David M. Fram. NPP: An easy to implement preprocessor for text formatting. Software—Practice and Experience, 10(5):335– 346, May 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bickmore:1997:MPL

[BF97] Timothy Bickmore and Robert E. Filman. MultiLex, a pipelined lexical analyzer. Software—Practice and Experience, 27(1): 25–32, January 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7255;   http://www3.
interscience.wiley.com/

cgi-bin/fulltext?ID=7255&
PLACEBO=IE.pdf.

Bellifemine:2011:SDS

[BFG^{+}11] Fabio Bellifemine, Giancarlo Fortino, Roberta Giannantonio, Raffaele Gravina, Antonio Guerrieri, and Marco Sgroi. SPINE: a domain-specific framework for rapid prototyping of WBSN applications. Software—Practice and Experience, 41(3):237–265, March 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brogi:2020:HPY

[BFGL20] Antonio Brogi, Stefano Forti, Carlos Guerrero, and Isaac Lera. How to place your apps in the fog: State of the art and open challenges. Software—Practice and Experience, 50(5):719– 740, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Becucci:2005:CBH

[BFGS05] Marcello Becucci, Alessandro Fantechi, Marco Giromini, and Emilio Spinicci. A comparison between handwritten and automatic generation of C code from SDL using static analysis. Software—Practice and Experience, 35(14):1317– 1347, November 25, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Birman:1999:MSD

[BFHR99] Kenneth P. Birman, Roy Friedman, Mark Hayden, and Injong Rhee. Middleware support for distributed multimedia and collaborative computing. Software—Practice and Experience, 29 (14):1285–1312, December 10, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
67500435/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=67500435&PLACEBO=IE.
pdf.

Bendisposto:2011:DCT

[BFJ^{+}11] Jens Bendisposto, Fabian Fritz, Michael Jastram, Michael Leuschel, and Ingo Weigelt. Developing Camille, a text editor for Rodin. Software—Practice and Experience, 41(2):189– 198, February 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brisaboa:2008:NAC

[BFNP08] N. R. Brisaboa, A. Fariña, G. Navarro, and J. R. Parama. New adaptive compressors for natural language text. Software —Practice and Experience, 38(13):1429–1450, November 10, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Blanco-Fernandez:2008:MFP

[BFPAGS^{+}08] Yolanda Blanco-Fernández, José J. Pazos-Arias, Alberto Gil-Solla, Manuel Ramos-Cabrer, Martín López-Nores, Jorge García-Duque, Ana Fernández-Vilas, Rebeca P. Díaz-Redondo, and Jesús Bermejo-Muñoz. An MHP framework to provide intelligent personalized recommendations about digital TV contents. Software—Practice and Experience, 38(9):925– 960, July 25, 2008. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bocanegra:2022:EST

[BFRFB22] Jose Bocanegra, Rafael Z. Frantz, Fabricia Roos-Frantz, and Fabio P. Basso. Evaluating the semantic transparency of Guaraná: a domain-specific language for enterprise application integration. Software— Practice and Experience, 52(4):967–983, April 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bell:1993:ESA

[BG93] Jim Bell and Gopal K.

Gupta. An evaluation of self-adjusting binary search tree techniques. Software —Practice and Experience, 23(4):369–382, April 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bartholdi:2001:VLA

[BG01] John J. Bartholdi III and Paul Goldsman. Vertexlabeling algorithms for the Hilbert spacefilling curve. Software—Practice and Experience, 31(5):395– 408, April 25, 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
77502753/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=77502753&PLACEBO=IE.
pdf.

Benuwa:2020:DLS

[BGA20] Ben-Bright Benuwa, Benjamin Ghansah, and Ernest K. Ansah. Deep localitysensitive discriminative dictionary learning for semantic video analysis. Software—Practice and Experience, 50(4):388–406, April 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Butler:2001:OOD

[BGG01] Greg Butler, Andrea Gantchev, and Peter Grogono. Objectoriented design of the subsumption architecture. Software—Practice and Experience, 31(9):911–923, July 25, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
80502221/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=80502221&PLACEBO=IE.
pdf.

B:1999:DIF

[BGM99] Praveen B., Deepak Gupta, and Rajat Moona. Design and implementation of a file system with onthe-fly data compression for GNU/ Linux. Software—Practice and Experience, 29(10):863–874, August 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=63001367;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=63001367&PLACEBO=IE.
pdf.

Bombonatti:2017:STS

[BGM17] Denise Bombonatti, Miguel Goulão, and Ana Moreira.

Synergies and tradeoffs in software reuse — a systematic mapping study. Software—Practice and Experience, 47(7):943–957, July 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Boyer:2017:RRP

[BGP17] Fabienne Boyer, Olivier Gruber, and Damien Pous. A robust reconfiguration protocol for the dynamic update of componentbased software systems. Software—Practice and Experience, 47(11):1729– 1753, November 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bedi:2013:ECP

[BGS^{+}13] Punam Bedi, Vandana Gandotra, Archana Singhal, Himanshi Narang, and Sumit Sharma. Extended conference paper: Threat-oriented security framework in risk management using multiagent system. Software—Practice and Experience, 43 (9):1013–1038, September 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ballesteros:2018:ZNF

[BGS18] Francisco J. Ballesteros, Gorka Guardiola, and Enrique Soriano. ZX: a network file system for high-latency networks. Software—Practice and Experience, 48(3):578– 599, March 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2550.

Bouzerzour:2020:SSI

[BGS20] Nour El Houda Bouzerzour, Souad Ghazouani, and Yahya Slimani. A survey on the service interoperability in cloud computing: Client-centric and provider-centric perspectives. Software—Practice and Experience, 50(7): 1025–1060, July 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bhardwaj:2020:HUP

[BGSG20] Amit Kumar Bhardwaj, Yuvraj Gajpal, Chirag Surti, and Sukhpal Singh Gill. HEART: Unrelated parallel machines problem with precedence constraints for task scheduling in cloud computing using heuristic and metaheuristic algorithms. Software—Practice and Experience, 50(12):2231–2251, December 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baxter:1982:SSV

[BH82] Anthony Q. Baxter and Johnson M. Hart. Software segmented virtual memory. Software—Practice and Experience, 12(2):185– 194, February 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Broom:1987:IUT

[BH87] M. A. Broom and T. R.

Hopkins. The implementation and use of Toolpack/1 on a graphics workstation. Software—Practice and Experience, 17(9): 561–573, September 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Busby:1992:PIM

[BH92] J. S. Busby and D. Hutchison. The practical integration of manufacturing applications. Software— Practice and Experience, 22(2):183–207, February 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Blostein:1994:LME

[BH94] D. Blostein and L. Haken.

The Lime Music Editor: a diagram editor involving complex transformations. Software—Practice and Experience, 24(3):289– 306, March 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baker:2001:DIG

[BH01] Scott Baker and John H. Hartman. The design and implementation of the Gecko NFS Web proxy. Software—Practice and Experience, 31(7):637– 665, June 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
78003147/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=78003147&PLACEBO=IE.
pdf.

Bhasker:1988:PGA

[Bha88] J. Bhasker. Process-graph analyser: a front-end tool for VHDL behavioural synthesis. Software—Practice and Experience, 18(5):469– 483, May 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Balalaie:2018:MMP

[BHJ^{+}18] Armin Balalaie, Abbas Heydarnoori, Pooyan Jamshidi, Damian A. Tamburri, and Theo Lynn. Microservices migration patterns. Software—Practice and Experience, 48(11):2019–2042, November 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bouchenak:2004:EIE

[BHK^{+}04] S. Bouchenak, D. Hagimont, S. Krakowiak, N. De Palma, and F. Boyer. Experiences implementing efficient Java thread serialization, mobility and persistence. Software—Practice and Experience, 34(4): 355–393, April 10, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bell:1973:UB

[BHL73] A. G. Bell, P. J. Hallowell, and D. H. Long. A universal benchmark? Software—Practice and Experience, 3(4):355–358, October/December 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Binder:2009:PIP

[BHMV09] Walter Binder, Jarle Hulaas, Philippe Moret, and Alex Villazón. Platformindependent profiling in a virtual execution environment. Software—Practice and Experience, 39(1):47– 79, January ??, 2009. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Benitez-Hidalgo:2023:NSO

[BHNDdMRG23] Antonio Benítez-Hidalgo, Ismael Navas-Delgado, and María del Mar Roldán-García. NORA: Scalable OWL reasoner based on NoSQL databases and Apache Spark. Special focus section: Benchmarking, Experimentation Tools, and Reproducible Practices for Data-Intensive Systems from Edge to Cloud: Software: Practice and Experience, 53(12):2377–2392, December 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baumann:2002:MMA

[BHR^{+}02] J. Baumann, F. Hohl, K. Rothermel, M. Strasser, and W. Theilmann. MOLE: a mobile agent system. Software—Practice and Experience, 32(6):575–603, May 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
93513618/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=93513618&PLACEBO=IE.
pdf.

Baude:2015:PDA

[BHR15] Françoise Baude, Ludovic Henrio, and Cristian Ruz.

Programming distributed and adaptable autonomous components — the GCM/ ProActive framework. Software—Practice and Experience, 45(9):1189–1227, September 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bian:2023:NPS

[BHT^{+}23] Haodong Bian, Jianqiang Huang, Jiahao Tang, Runting Dong, Li Wu, and Xiaoying Wang. PAS: a new powerful and simple quantum computing simulator. Software—Practice and Experience, 53(1):142– 159, January 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barr:2005:JEA

[BHvR05] Rimon Barr, Zygmunt J.

Haas, and Robbert van Renesse. JiST: an efficient approach to simulation using virtual machines. Software —Practice and Experience, 35(6):539–576, May 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Basney:2005:MOC

[BHW05] Jim Basney, Marty Humphrey, and Von Welch. The MyProxy online credential repository. Software —Practice and Experience, 35(9):801–816, July 25, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Budinsky:1985:SSR

[BHZ85] Frank J. Budinsky, Richard C. Holt, and Safwat G. Zaky. SRE: a syntax recognizing editor. Software —Practice and Experience, 15(5):489–497, May 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Buyya:2024:EES

[BIA24] Rajkumar Buyya, Shashikant Ilager, and Patricia Arroba. Energy-efficiency and sustainability in new generation cloud computing: a vision and directions for integrated management of data centre resources and workloads. Software —Practice and Experience, 54(1):24–38, January 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Binder:2006:PAS

[Bin06] Walter Binder. Portable and accurate sampling profiling for Java. Software —Practice and Experience, 36(6):615–650, May 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Butler:1994:DGP

[BIO94] Gregory Butler, Sridhar S.

Iyer, and E. A. O’Brien. A database of groups of prime-power order. Software—Practice and Experience, 24(10):911–951, October 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Birman:1999:RER

[Bir99] Kenneth P. Birman. A review of experiences with reliable multicast. Software—Practice and Experience, 29(9):741–774, July 25, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=62501862;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=62501862&PLACEBO=IE.
pdf.

Bishop:1979:BRBa

[Bis79a] J. M. Bishop. Book review:

A practical introduction to Pascal, I. R. Wilson and A. M. Addyman, Macmillan Press, 1978. No. of pages:

148. Price: £3.50. Software —Practice and Experience, 9(3):252–253, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1979:BRBb

[Bis79b] Judy M. Bishop. Book review: Programming in Pascal, Peter Grogono, Addison-Wesley Series in Computer Science, 1978. No. of pages: 359. Price: £7.50. Software—Practice and Experience, 9(5):425– 426, May 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1979:ISP

[Bis79c] Judy M. Bishop. Implementing strings in Pascal. Software—Practice and Experience, 9(9):779– 788, September 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1979:PP

[Bis79d] Judy M. Bishop. On Publication Pascal. Software —Practice and Experience, 9(9):711–717, September 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1980:LES

[Bis80] Judy Bishop. Letter to the Editor: A sense of proportion. Software—Practice and Experience, 10(2):160, February 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1981:BRBa

[Bis81a] J. M. Bishop. Book review:

Fortran for students, Roger Hutty, Macmillan, London, 1980. Software—Practice and Experience, 11(1):101, January 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1981:BRBb

[Bis81b] Judy M. Bishop. Book review: Programming via Pascal, J. S. Rohl and H. J. Barrett, Cambridge University Press, 1980. No. of pages: 327, E12.50 (cloth). Software—Practice and Experience, 11(8):880– 881, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1982:BRB

[Bis82] J. M. Bishop. Book review: Pascal programming, Laurence Atkinson, Wiley, Chichester, 1980. No. of pages: 428. Price £16.50 (Hardback); £6.95 (paperback). Software—Practice and Experience, 12(1):99, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1984:BRI

[Bis84] J. M. Bishop. Book review:

Information Representation and Manipulation Using Pascal, E. S. Page and L. B. Wilson. Cambridge University Press, 1983. No. of pages: 272. Price: hardback, £15.00; paperback, £6.95. Software—Practice and Experience, 14(2):195, February 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1986:BRB

[Bis86] J. M. Bishop. Book review: Fundamentals of programming languages, Ellis Horowitz, Springer-Verlag, 1983. no. of pages:

450. Software—Practice and Experience, 16(1):100, January 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1987:PUU

[Bis87] Matt Bishop. Profiling under UNIX by patching. Software—Practice and Experience, 17(10): 729–739, October 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1990:CUR

[Bis90] Matt Bishop. Collaboration using roles. Software —Practice and Experience, 20(5):485–497, May 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:2019:ROD

[Bis19] Judith Bishop. Recognising our distinguished referees. Software—Practice and Experience, 49(5): 697, May 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1972:EJC

[BJ72] D. W. Barron and I. R.

Jackson. The evolution of job control languages. Software—Practice and Experience, 2(2):143–164, April/ June 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Balasubramanian:2021:SFH

[BJ21] Venki Balasubramanian and Alireza Jolfaei. A scalable framework for healthcare monitoring application using the Internet of Medical Things. Software—Practice and Experience, 51(12):2457–2468, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Buchheim:2006:DRT

[BJL06] Christoph Buchheim, Michael Jünger, and Sebastian Leipert. Drawing rooted trees in linear time. Software—Practice and Experience, 36(6):651–665, May 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ballesteros:2000:UIC

[BJP^{+}00] F. J. Ballesteros, Ricardo Jimenez, Marta Patiño, Fabio Kon, Sergio Arevalo, and Roy Campbell. Using interpreted Composite-Calls to improve operating system services. Software—Practice and Experience, 30(6):589–615, May 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
71007582/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=71007582&PLACEBO=IE.
pdf.

Bernstein:1977:NGP

[BK77] N. Bernstein and B. Knobe. A non-general-purpose operating system. Software —Practice and Experience, 7(2):223–226, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bentley:1986:GLT

[BK86a] Jon Louis Bentley and Brian W. Kernighan. GRAP — a language for typesetting graphs. Communications of the Association for Computing Machinery, 29(8):782–792, August 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/
Abstracts/0001-0782/6429.
html.

Burns:1986:CIM

[BK86b] Alan Burns and J. A.

Kirkham. The construction of information management system prototypes in Ada. Software—Practice and Experience, 16(4):341– 350, April 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Butler:1987:SMS

[BK87] Gregory Butler and Matthew J.

Kendall. The suitability for master/slave concurrency of Concurrent Euclid, Ada and Modula. Software— Practice and Experience, 17(2):117–134, February 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bell:1993:LMS

[BK93] Timothy C. Bell and David Kulp. Longest-match string searching for Ziv– Lempel compression. Software—Practice and Experience, 23(7):757–771, July 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bartel:2023:GLD

[BK23] Silas Bartel and Matthias Korch. Generation of logic designs for efficiently solving ordinary differential equations on field programmable gate arrays. Software—Practice and Experience, 53(1):27– 52, January 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Battou:2002:CCA

[BKL^{+}02] Abdella Battou, Bilal Khan, Daniel C. Lee, Spencer Marsh, Sean Mountcastle, and David Talmage. CASiNO: component architecture for simulating network objects. Software—Practice and Experience, 32(11):1099–1128, September 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        97519257/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=97519257&PLACEBO=IE.
                                        pdf.

Bijlsma:2022:EDP

[BKP^{+}22] Lex A. Bijlsma, Arjan J. F. Kok, Harrie J. M.

Passier, Harold J. Pootjes, and Sylvia Stuurman. Evaluation of design pattern alternatives in Java. Software—Practice and Experience, 52(5): 1305–1315, May 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bowie:1978:STF

[BL78] W. S. Bowie and J. G. Linders. A software trace facility for OS/MVT. Software —Practice and Experience, 9(??):535–545, ?? 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bowie:1979:STF

[BL79] William S. Bowie and J. G.

Linders. A software trace facility for OS/MVT. Software—Practice and Experience, 9(7):535–545, July 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bull:1983:RTB

[BL83] Gordon Bull and Alan Lewis. Real-time BA-SIC. Software—Practice and Experience, 13 (11):1075–1092, November 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barak:1985:MMD

[BL85] Amnon Barak and Ami Litman. MOS: a multicomputer distributed operating system. Software—Practice and Experience, 15(8):725– 737, August 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barbosa:1990:DPS

[BL90a] Valmir C. Barbosa and Priscilla M. V. Lima. On the distributed parallel simulation of Hopfield’s neural networks. Software —Practice and Experience, 20(10):967–983, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1990:SEY

[BL90b] David Barron and Charles Lang. SPE — the early years. Software—Practice and Experience, 20(1):3–4, January 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Berbecaru:2015:EEU

[BL15] Diana Berbecaru and Antonio Lioy. Exploiting the European Union trusted service status list for certificate validation in STORK: design, implementation, and lessons learnt. Software—Practice and Experience, 45(11):1457–1477, November 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Blake:1992:AIT

[Bla92] Ben A. Blake. Assignment of independent tasks to minimize completion time. Software—Practice and Experience, 22(9):723– 734, September 1992. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Blake:2004:SLS

[Bla04]  M. Brian Blake.  A

specification language and service-oriented architecture to support distributed data management. Software—Practice and Experience, 34(11):1091–1117, September 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Berbecaru:2019:PDI

[BLC19] Diana Berbecaru, Antonio Lioy, and Cesare Cameroni. Providing digital identity and academic attributes through European eID infrastructures: Results achieved, limitations, and future steps. Software—Practice and Experience, 49(11):1643–1662, November 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Boloni:2008:CSP

[BLE^{+}08] Ladislau Bölöni, Linus J. Luotsinen, Joakim N. Ekblad, T. Ryan Fitz-Gibbon, Charles Houchin, Justin L. Key, Majid Ali Khan, Jin Lyu, Johann Nguyen, Rex Oleson, Gary Stein, Scott A. Vander Weide, and Viet Trinh. A comparison study of 12 paradigms for developing embodied agents. Software—Practice and Experience, 38(3): 259–305, March ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bershad:1988:PSO

[BLL88] Brian N. Bershad, Edward D. Lazowska, and Henry M. Levy. PRESTO: a system for objectoriented parallel programming. Software—Practice and Experience, 18(8):713– 732, August 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bernard:2004:GTS

[BLLP04] Eddy Bernard, Bruno Legeard, Xavier Luck, and Fabien Peureux. Generation of test sequences from formal specifications: GSM 11-11 standard case study. Software—Practice and Experience, 34 (10):915–948, August 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Berkovich:2000:BCA

[BLM00] Simon Berkovich, Gennadi M. Lapir, and Marilyn Mack. A bit-counting algorithm using the frequency division principle. Software—Practice and Experience, 30(14):1531–1540, November 25, 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
73001229/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=73001229&PLACEBO=IE.
pdf.

Buccafurri:2015:SES

[BLNU15] Francesco Buccafurri, Gianluca Lax, Antonino Nocera, and Domenico Ursino. A system for extracting structural information from Social Network accounts. Software—Practice and Experience, 45 (9):1251–1275, September 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bloesch:1993:ALG

[Blo93] Anthony Bloesch. Aesthetic layout of generalized trees. Software—Practice and Experience, 23(8):817– 827, August 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bellodi:2017:WSR

[BLR^{+}17] Elena Bellodi, Evelina Lamma, Fabrizio Riguzzi, Riccardo Zese, and Giuseppe Cota. A web system for reasoning with probabilistic OWL. Software—Practice and Experience, 47 (1):125–142, January 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Briand:2003:IUA

[BLS03] L. C. Briand, Y. Labiche, and H. Sun. Investigating the use of analysis contracts to improve the testability of objectoriented code. Software —Practice and Experience, 33(7):637–672, June 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Blum:1986:IDI

[Blu86] Bruce I. Blum. Iterative development of information systems: a case study. Software—Practice and Experience, 16(6):503– 515, June 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bratley:1972:CR

[BM72] Paul Bratley and Jean Millo. Computer recreations. Software—Practice and Experience, 2(4):397– 400, October 1972. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bentley:1993:ESF

[BM93] Jon Louis Bentley and M. Douglas McIlroy. Engineering a sort function. Software—Practice and Experience, 23 (11):1249–1265, November 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Borba:1997:STE

[BM97] Paulo Borba and Silvio Meira. A system for translating executable VDM specifications into Lazy ML. Software—Practice and Experience, 27(3):271– 289, March 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7291;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7291&
PLACEBO=IE.pdf.

Bederson:1998:IZU

[BM98] Ben Bederson and Jon Meyer. Implementing a zooming User Interface: experience building Pad++. Software —Practice and Experience, 28(10):1101–1135, August 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=10007356;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=10007356&PLACEBO=IE.
pdf.

Bosi:2001:ECB

[BM01] F. Bosi and M. Milano. Enhancing CLP branch and bound techniques for scheduling problems. Software—Practice and Experience, 31(1):17–42, January 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
76502347/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=76502347&PLACEBO=IE.
pdf.

Baldoni:2003:TTR

[BM03] Roberto Baldoni and Carlo Marchetti. Three-tier replication for FT-CORBA infrastructures. Software— Practice and Experience, 33(8):767–797, July 10, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barbosa:2006:DGS

[BM06] Denilson Barbosa and Alberto O. Mendelzon. Declarative generation of synthetic XML data. Software—Practice and Experience, 36(10):1051–1079, August 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bracher:1972:LDC

[BMA72] B. H. Bracher, J. F. Macewan, and A. G. Abbott. An on-line data collection system for film measurement. Software—Practice and Experience, 2(4):389– 396, October/December 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Buyya:2005:SPS

[BMAV05] Rajkumar Buyya, Manzur Murshed, David Abramson, and Srikumar Venugopal. Scheduling parameter sweep applications on global Grids: a deadline and budget constrained cost-time optimization algorithm. Software—Practice and Experience, 35(5): 491–512, April 25, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barbaglia:2017:SPD

[BMC17] Guido Barbaglia, Simone Murzilli, and Stefano Cudini. Survery paper: Definition of REST web services with JSON schema. Software—Practice and Experience, 47(6):907–920, June 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Batista:2024:MAW

[BMC^{+}24] Cesar Batista, Felipe Morais, Everton Cavalcante, Thais Batista, Bruno Proença, and William Breno Rodrigues Cavalcante. Managing asynchronous workloads in a multi-tenant microservice enterprise environment. Software—Practice and Experience, 54(2): 334–359, February 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Black:1998:EMM

[BMD^{+}98] David L. Black, Dejan S. Milojičić, Randall W. Dean, Michelle Dominijanni, Alan Langerman, and Steven J. Sears. Extended memory management (XMM): Lessons learned. Software—Practice and Experience, 28(9): 1011–1031, July 25, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www3.interscience.
wiley.com/cgi-bin/abstract?
ID=1819;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1819&
PLACEBO=IE.pdf.

Blair:1985:CU

[BMM85] Gordon S. Blair, Jon R.

Malone, and John A. Mariani. A critique of UNIX. Software—Practice and Experience, 15 (12):1125–1139, December 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Basso:2018:PIA

[BMM^{+}18] Tania Basso, Leonardo Montecchi, Regina Moraes, Mario Jino, and Andrea Bondavalli. PrivAPP: an integrated approach for the design of privacy-aware applications. Software—Practice and Experience, 48(3):499– 527, March 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2546.

Briola:2019:PPA

[BMM19] Daniela Briola, Daniela Micucci, and Leonardo Mariani. A platform for P2P agent-based collaborative applications. Software—Practice and Experience, 49(3):549–558, March 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brownbridge:1982:NCU

[BMR82] David R. Brownbridge, L. F. Marshall, and B. Randell. The Newcastle connection or UNIXes of the world unite! Software—Practice and Experience, 12(12):1147–1162, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bakic:2000:BPF

[BMR00] Aleksandar Bakić, Matt W. Mutka, and Diane T. Rover. BRISK: a portable and flexible distributed instrumentation system. Software—Practice and Experience, 30(12):1353– 1373, October 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        73000928/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=73000928&PLACEBO=IE.
                                        pdf.

Bakic:2003:LPV

[BMR03] Aleksandar M. Bakić, Matt W. Mutka, and Diane T. Rover. An online performance visualization technology. Software—Practice and Experience, 33(15):1447–1469, December 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Balland:2014:ESP

[BMR14] Emilie Balland, Pierre-Etienne Moreau, and Antoine Reilles. Effective strategic programming for Java developers. Software —Practice and Experience, 44(2):129–162, February 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Blair:1983:PEU

[BMS83] Gordon S. Blair, John A.

Mariani, and W. D. Shepherd. A practical extension to UNIX for interprocess communication. Software —Practice and Experience, 13(1):45–58, January 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bingmann:2021:EFS

[BMS21] Timo Bingmann, Jasper Marianczuk, and Peter Sanders. Engineering faster sorters for small sets of items. Software—Practice and Experience, 51(5):965– 1004, May 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bliudze:2017:ECC

[BMSZ17] Simon Bliudze, Anastasia Mavridou, Radoslaw Szymanek, and Alina Zolotukhina. Exogenous coordination of concurrent software components with JavaBIP. Software—Practice and Experience, 47 (11):1801–1836, November 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Binder:2016:PBI

[BMTA16] Walter Binder, Philippe Moret, Éric Tanter, and Danilo Ansaloni. Polymorphic bytecode instrumentation. Software—Practice and Experience, 46(10): 1351–1380, October 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bacon:2003:ACT

[BMY03] Jean Bacon, Ken Moody, and Walt Yao. Access control and trust in the use of widely distributed services. Software—Practice and Experience, 33(4):375– 394, April 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bachelet:2006:DGA

[BMY06] Bruno Bachelet, Antoine Mahul, and Loïc Yon. Designing generic algorithms for operations research. Software—Practice and Experience, 36(1):73– 93, January 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Buhr:1992:SAH

[BMZ92] Peter A. Buhr, Hamish I.

Macdonald, and C. Robert Zarnke. Synchronous and asynchronous handling of abnormal events in the muSystem. Software— Practice and Experience, 22(9):735–776, September 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Badros:2000:FPA

[BN00] Greg J. Badros and David Notkin. A framework for preprocessor-aware C source code analyses. Software—Practice and Experience, 30(8):907–924, July 10, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72502207/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72502207&PLACEBO=IE.
pdf.

Bishop:2013:EDT

[BN13] Judith Bishop and David Notkin. Editorial: Developing tools as plug-ins: TOPI 2011 special issue editorial. Software—Practice and Experience, 43(4):389– 390, April 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bustard:1992:EFD

[BNOW92] D. W. Bustard, M. T. Norris, R. A. Orr, and A. C. Winstanley. An exercise in formalizing the description of a concurrent system. Software—Practice and Experience, 22 (12):1069–1098, December 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Birrell:1995:NO

[BNOW95] Andrew Birrell, Greg Nelson, Susan Owicki, and Edward Wobber. Network objects. Software—Practice and Experience, 25(S4):87– 130, December 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brogi:2018:MBA

[BNS18] Antonio Brogi, Davide Neri, and Jacopo Soldani. A microservice-based architecture for (customisable) analyses of Docker images. Software—Practice and Experience, 48(8): 1461–1474, August 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2583.

Brown:1983:SMP

[BO83] P. J. Brown and J. A. Ogden. The SUPERMAC macro processor in Pascal. Software—Practice and Experience, 13(4):295– 304, April 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bouchebaba:2012:MFS

[BOPN12] Youcef Bouchebaba, Ali-Erdem Ozcan, Pierre Paulin, and Gabriela Nicolescu. MpAssign: a framework for solving the manycore platform mapping problem. Software—Practice and Experience, 42(7): 891–915, July 2012. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Boris:1983:CP

[Bor83] G. J. Boris. Correct programs? Software—Practice and Experience, 13(8):768– 769, August 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bornat:1986:PGO

[Bor86] Richard Bornat. A protocol for generalized occam. Software—Practice and Experience, 16(9):783– 799, September 1986. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Botting:1977:ESA

[Bot77]  R. J. Botting.  Effi-

cient storage for amorphous data. Software— Practice and Experience, 7 (2):201–203, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bourne:1971:DTE

[Bou71] S. R. Bourne. A design for a text editor. Software—Practice and Experience, 1(1):73–81, January/ March 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Boussinot:1991:RCE

[Bou91] Frédéric Boussinot. Reactive C: An extension of C to program reactive systems. Software—Practice and Experience, 21(4):401– 428, April 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bovey:1987:DGW

[Bov87] J. D. Bovey. A debugger for a graphical workstation. Software—Practice and Experience, 17(9):

647–662, September 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bowring:1973:WRR

[Bow73] C. A. Bowring. Writing a roll-out-roll-in facility for small machines. Software —Practice and Experience, 3(2):155–160, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bowen:1988:BRB

[Bow88] J. P. Bowen. Book review: Modula-2: Constructive program development, P. Messer and I. Marshall, Blackwell Scientific Publications, 1986. ISBN 0-632-01609-4. No. of pages: 544. Price: £35.00 (paperback £14.95). Software—Practice and Experience, 18(2): 167–168, February 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Boyland:2001:ABU

[Boy01]  John Boyland.  Alias

burying: Unique variables without destructive reads. Software—Practice and Experience, 31(6):533– 553, May 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/

                                        78003105/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=78003105&PLACEBO=IE.
                                        pdf.

Barcucci:1984:SDS

[BP84a] Elena Barcucci and Gianluca Pelacani. A software development system based on a macroprocessor. Software—Practice and Experience, 14(6):519– 531, June 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1984:MNC

[BP84b] Cynthia A. Brown and Paul Walton Purdom, Jr. A methodology and notation for compiler front end design. Software—Practice and Experience, 14(4):335– 346, April 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Binder:1990:FEL

[BP90] Leon H. Binder and J. H. Poore. Field experiments with local software quality metrics. Software—Practice and Experience, 20(7): 631–647, July 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brinkley:1997:SFS

[BP97] James F. Brinkley and Jeffrey S. Prothero. Slisp: a flexible software toolkit for hybrid, embedded and distributed applications. Software—Practice and Experience, 27(1):33–48, January 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7256;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7256&
PLACEBO=IE.pdf.

Bhamidipaty:1998:VFY

[BP98] A. Bhamidipaty and T. A.

Proebsting. Very fast YACC-compatible parsers (for very little effort). Software—Practice and Experience, 28(2):181–190, February 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1773;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1773&
PLACEBO=IE.pdf.

Bechini:2002:PSD

[BP02] Alessio Bechini and Cosimo Antonio Prete. Performancesteered design of software architectures for embedded multicore systems. Software—Practice and Experience, 32(12):1155–1173, October 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        98516187/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=98516187&PLACEBO=IE.
                                        pdf.

Bovet:2008:AAG

[BP08] Jean Bovet and Terence Parr. ANTLRWorks: an ANTLR grammar development environment. Software—Practice and Experience, 38(12):1305–1332, October ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bate:2009:CPP

[BP09] Iain Bate and Simon Poulding. Call for papers: Practical aspects of search-based software engineering. Software—Practice and Experience, 39(9): 867–868, June 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bate:2011:ESI

[BP11] Iain Bate and Simon Poulding. Editorial for the special issue on searchbased software engineering. Software—Practice and Experience, 41(5):467– 468, May 25, 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Buhr:2023:HPE

[BPDL23] Peter A. Buhr, Colby A.

Parsons, Thierry Delisle, and He Nan Li. Highperformance extended actors. Special focus section: Benchmarking, Experimentation Tools, and Reproducible Practices for Data-Intensive Systems from Edge to Cloud: Software: Practice and Experience, 53(12):2463–2500, December 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brandner:2013:ECP

[BPK13] Florian Brandner, Viktor Pavlu, and Andreas Krall. Extended conference papers: Automatic generation of compiler backends. Software—Practice and Experience, 43(2):207– 240, February 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Borie:1993:LKS

[BPM93] Richard B. Borie, Allen S.

Parrish, and Srinivas Mandyam. Lock-and-key strategies for handling undefined variables. Software—Practice and Experience, 23(7):693– 710, July 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brooke:2010:DCX

[BPP10] Phillip J. Brooke, Richard F. Paige, and Christopher Power. Documentcentric XML workflows with fragment digital signatures. Software—Practice and Experience, 40(8): 655–672, July 2010. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bellifemine:2001:DMA

[BPR01] Fabio Bellifemine, Agostino Poggi, and Giovanni Rimassa. Developing multiagent systems with a FIPAcompliant agent framework. Software—Practice and Experience, 31(2):103– 128, February 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        76504934/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=76504934&PLACEBO=IE.
                                        pdf.

Bush:2000:SAF

[BPS00] William R. Bush, Jonathan D.

Pincus, and David J. Sielaff. A static analyzer for finding dynamic programming errors. Software—Practice and Experience, 30(7):775–802, June 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72001834/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72001834&PLACEBO=IE.
pdf.

Bowman:1990:UAB

[BPY90] Mic Bowman, Larry L. Peterson, and Andrey Yeatts. Univers: An attributebased name server. Software—Practice and Experience, 20(4):403–424, April 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1988:CHS

[BR88] P. J. Brown and M. T. Russell. Converting help systems to hypertext. Software—Practice and Experience, 18(2):163–165, February 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baumgartner:1995:SLE

[BR95] Gerald Baumgartner and Vincent F. Russo. Signatures: a language extension for improving type abstraction and subtype polymorphism in C++. Software—Practice and Experience, 25(8):863–889, August 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Burgess:1997:DRA

[BR97] Mark Burgess and Ricky Ralston. Distributed resource administration using Cfengine. Software—Practice and Experience, 27(9):1083–1101, September 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7335;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7335&
PLACEBO=IE.pdf.

Begay:2001:RIF

[BR01a] D. Bégay and A. Rauzy. A realistic involvement of formal methods. Software—Practice and Experience, 31(2):191–208, February 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
76504937/START;   http:
//www3.interscience.wiley.

com/cgi-bin/fulltext?
ID=76504937&PLACEBO=IE.
pdf.

Bruneton:2001:AEM

[BR01b] Eric Bruneton and Michel Riveill. An architecture for extensible middleware platforms. Software—Practice and Experience, 31 (13):1237–1264, November 10, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
85011927/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=85011927&PLACEBO=IE.
pdf.

Braid:1975:BRB

[Bra75] I. C. Braid. Book review: Principles of Interactive Computer Graphics, W. M. Newman and R. F. Sproull, McGraw-Hill, New York, 1973. No. of pages:

607. Price: £14.95. Software—Practice and Experience, 5(3):315, July 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brailsford:1980:BRB

[Bra80] David F. Brailsford. Book review: Informal introduction to Algol 68 (revised edn), C. H. Lindsey and S. G. van der Meulen, North Holland, 1977. No. of pages: 361. Price: £14.00. Software —Practice and Experience, 10(2):161–162, February 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bradley:1999:EMD

[Bra99]  J. Bradley.  An effi-

cient modularized database structure for a highresolution column-gridded Mars global terrain database. Software—Practice and Experience, 29(5):437–456, April 25, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=55003856;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=55003856&PLACEBO=IE.
                                        pdf.

Brereton:1982:PFM

[Bre82] O. P. Brereton. Performance figures for messagepassing over a Cambridge ring. Software—Practice and Experience, 12(1):95– 96, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brereton:1986:MRF

[Bre86] O. P. Brereton. Management of replicated files in a UNIX environments. Software—Practice and Experience, 16(8):771–780, August 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brender:2002:BPL

[Bre02] Ronald F. Brender. The BLISS programming language: a history. Software—Practice and Experience, 32(10):955–981, August 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
96516179/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=96516179&PLACEBO=IE.
pdf.

Brignell:1982:BRU

[Bri82] J. E. Brignell. Book review: The use of microprocessors, M. Auminaux, Wiley, Chichester, 1980. No. of pages: viii + 198. Price: £13.20 ($31.35). Software —Practice and Experience, 12(5):496, May 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Briggs:1984:TIA

[Bri84] J. S. Briggs. Two implementations of the Ada program library. Software —Practice and Experience, 14(5):491–500, May 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Briggs:1987:GRP

[Bri87] J. S. Briggs. Generating reversible programs. Software—Practice and Experience, 17(7):439–453, July 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bissyande:2015:IEC

[BRL^{+}15] Tegawendé F. Bissyandé, Laurent Réveillère, Julia L. Lawall, Yérom-David Bromberg, and Gilles Muller. Implementing an embedded compiler using program transformation rules. Software—Practice and Experience, 45(2): 177–196, February 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bischof:1997:AEA

[BRMO97] C. H. Bischof, L. Roh, and A. J. Mauer-Oats. ADIC: an extensible automatic differentiation tool for ANSI-C. Software —Practice and Experience, 27(12):1427–1456, December 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.

com/cgi-bin/abstract?
ID=7279;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7279&
PLACEBO=IE.pdf.

Brown:1971:KLS

[Bro71] P. J. Brown. The Kent on-line system. Software—Practice and Experience, 1(3):269–277, July/ September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1972:RCS

[Bro72] P. J. Brown. Re-creation of source code from reverse Polish form. Software—Practice and Experience, 2(3):275–278, July/ September 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1974:WSA

[Bro74] P. J. Brown. Writing software in ALGOL. Software —Practice and Experience, 4(2):139–144, April/June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bron:1975:LE

[Bro75] C. Bron. Letter to the editor. Software—Practice and Experience, 5(3):312– 313, July 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1976:TC

[Bro76]  P. J. Brown.  Throw-

away compiling. Software—Practice and Experience, 6(3):423–434, July/ September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1977:MRC

[Bro77] P. J. Brown. More on the re-creation of source code from reverse polish. Software—Practice and Experience, 7(5):545– 551, September/October 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1978:LE

[Bro78] P. J. Brown. Letter to the editor. Software—Practice and Experience, 8(1):109, January 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1979:MT

[Bro79] P. J. Brown. Macros without tears. Software—Practice and Experience, 9(6): 433–437, June 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1980:SMF

[Bro80] P. J. Brown. SUPERMAC:

a macro facility that can be added to existing compilers. Software—Practice and Experience, 10(6):431– 434, June 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1981:DMI

[Bro81a] Frank M. Brown. Design of a MUMPS interpreter. Software—Practice and Experience, 11 (12):1293–1297, December 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1981:DPB

[Bro81b] P. J. Brown. Dynamic program building. Software—Practice and Experience, 11(8):831–843, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1982:MSG

[Bro82] P. J. Brown. “My system gives excellent error messages”–or does it? Software—Practice and Experience, 12(1):91–94, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1986:IDA

[Bro86a] Michael W. Brown. Integrating distributed array processing into EMAS 2900. Software—Practice and Experience, 16(6):517– 529, June 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1986:ID

[Bro86b] P. J. Brown. Interactive documentation. Software—Practice and Experience, 16(3):291–299, March 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brogi:2018:TSB

[BRS18] Antonio Brogi, Luca Rinaldi, and Jacopo Soldani. TosKer: a synergy between TOSCA and Docker for orchestrating multicomponent applications. Software—Practice and Experience, 48(11):2061–2079, November 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barioni:2009:SIS

[BRTT09] M. C. N. Barioni, H. L. Razente, A. J. M. Traina, and C. Traina, Jr. Seamlessly integrating similarity queries in SQL. Software —Practice and Experience, 39(4):355–384, March 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bruno:1984:UAD

[Bru84] Giorgio Bruno. Using Ada for discrete event simulation. Software—Practice and Experience, 14(7):685– 695, July 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bryant:1977:BRB

[Bry77] P. Bryant. Book review:

Benchmarking, edited by Nicholas Benwell, Halsted press, Hemisphere publishing corporation, Washington D.C. and London, 1975. no. of pages: 190. price: £11.80. Software— Practice and Experience, 7 (2):298–299, March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bauer:1974:DAR

[BS74] Alan M. Bauer and Harry J.

Saal. Does APL really need run-time checking? Software—Practice and Experience, 4(2):129– 138, April/June 1974. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barak:1980:USP

[BS80] Amnon B. Barak and Amos Shapir. UNIX with satellite processors. Software —Practice and Experience, 10(5):383–392, May 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1981:ESD

[BS81] J. D. Bishop and G. J. Smith. Experiences with a status database. Software—Practice and Experience, 11(12):1315–1329, December 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bailes:1984:SBF

[BS84] Paul A. Bailes and Antonio Salvadori. A semanticallybased formatting discipline for Pascal. Software—Practice and Experience, 14(3):235–251, March 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barak:1985:DLB

[BS85] Amnon Barak and Amnon Shiloh. A distributed load-balancing policy for a multicomputer. Software —Practice and Experience, 15(9):901–913, September 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bertran-Salvans:1988:FDA

[BS88] Miquel Bertran-Salvans. On a formal definition and application of dimensional design. Software—Practice and Experience, 18 (11):1029–1045, November 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Blaschek:1989:UAP

[BS89] G. Blaschek and J. Sametinger.

User-adaptable prettyprinting. Software—Practice and Experience, 19(7):687– 702, July 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bivens:1990:IRR

[BS90a] Mary P. Bivens and Mary Lou Soffa. Incremental register reallocation. Software—Practice and Experience, 20(10): 1015–1047, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bond:1990:IPC

[BS90b] Alan H. Bond and Basuki Soetarman. Integrating Prolog and CADAM to produce an intelligent CAD system. Software— Practice and Experience, 20(10):1049–1076, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Buhr:1990:SPL

[BS90c] Peter A. Buhr and Richard A. Stroobosscher. The μ-System: Providing lightweight concurrency on shared-memory multiprocessor computers running UNIX. Software—Practice and Experience, 20(9):929– 963, September 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Boysen:1993:ROS

[BS93] Pete Boysen and Pinaki Shah. Reducing object storage requirements in a multi-user environment. Software—Practice and Experience, 23(3):235– 241, March 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Boussinot:1998:STB

[BS98] Frédéric Boussinot and Jean-Ferdy Susini. The SugarCubes tool box: a reactive Java framework. Software—Practice and Experience, 28 (14):1531–1550, December 10, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=5001669;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?

ID=5001669&PLACEBO=IE.
pdf.

Blostein:1999:CGG

[BS99a] Dorothea Blostein and Andy Schürr. Computing with graphs and graph transformations. Software—Practice and Experience, 29(3):197–217, March 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55000310;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55000310&PLACEBO=IE.
pdf.

Brown:1999:PAM

[BS99b] Ian Brown and C. R.

Snow. A proxy approach to e-mail security. Software—Practice and Experience, 29(12):1049–1060, October 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
66000405/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=66000405&PLACEBO=IE.
pdf.

Boussinot:2000:JTS

[BS00] Frédéric Boussinot and Jean-Ferdy Susini. Java threads and SugarCubes. Software—Practice and Experience, 30(5):545–566, April 25, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
71004433/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=71004433&PLACEBO=IE.
pdf.

Babu:2019:SLA

[BS19] Kaippilly Raman Remesh Babu and Philip Samuel. Service-level agreementaware scheduling and load balancing of tasks in cloud. Software—Practice and Experience, 49(6):995– 1012, June 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brostoff:2005:RWD

[BSC^{+}05] Sacha Brostoff, M. Angela Sasse, David Chadwick, James Cunningham, Uche Mbanaso, and Sassa Otenko. R-What? development of a role-based access control policy-writing tool for e-scientists. Software —Practice and Experience, 35(9):835–856, July 25, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bonfim:2020:RTA

[BSDF20] Michel Bonfim, Marcelo Santos, Kelvin Dias, and Stênio Fernandes. A real-time attack defense framework for 5G network slicing. Software—Practice and Experience, 50(7): 1228–1257, July 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bracke:2021:DES

[BSM^{+}21] Vincent Bracke, Merlijn Sebrechts, Bart Moons, Jeroen Hoebeke, Filip De Turck, and Bruno Volckaert. Design and evaluation of a scalable Internet of Things backend for smart ports. Software—Practice and Experience, 51(7): 1557–1579, July 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Binder:2009:CPJ

[BSMV09] Walter Binder, Martin Schoeberl, Philippe Moret, and Alex Villazón. Crossprofiling for Java processors. Software—Practice and Experience, 39(18): 1439–1465, December 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bogo:2020:CAO

[BSNB20] Matteo Bogo, Jacopo Soldani, Davide Neri, and Antonio Brogi. Componentaware orchestration of cloud-based enterprise applications, from TOSCA to Docker and Kubernetes. Software—Practice and Experience, 50 (9):1793–1821, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bruestle:1985:ISD

[BSRS85] H. R. Bruestle, D. M. Snider, K. D. Russell, and H. D. Stewart. ISDMS: a scientific data management system. Software—Practice and Experience, 15(5):473– 487, May 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barsky:2010:SPA

[BST10] M. Barsky, U. Stege, and A. Thomo. A survey of practical algorithms for suffix tree construction in external memory. Software —Practice and Experience, 40(11):965–988, October 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bryant:1974:GUG

[BT74] P. E. Bryant and R. E. Thomas. Graphics under GEORGE 4. Software— Practice and Experience, 4(3):299, July/September 1974. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Basili:1975:TEC

[BT75] Victor R. Basili and Albert J. Turner. A transportable extendable compiler. Software—Practice and Experience, 5(3):269– 278, July/September 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brunt:1976:UOA

[BT76] R. F. Brunt and D. E.

Tuffs. A user-oriented approach to control languages. Software—Practice and Experience, 6(1):93– 108, January/March 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brain:1989:NPH

[BT89] Marshall D. Brain and Alan L. Tharp. Nearperfect hashing of large word sets. Software— Practice and Experience, 19(10):967–978, October 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bida:2007:ATS

[BT07] Eran Bida and Sivan Toledo. An automaticallytuned sorting library. Software—Practice and Experience, 37(11):1161–1192, September 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bodkhe:2021:SDD

[BT21a] Umesh Bodkhe and Sudeep Tanwar. Secure data dissemination techniques for IoT applications: Research challenges and opportunities. Software—Practice and Experience, 51 (12):2469–2491, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Borsotti:2021:EPS

[BT21b] Angelo Borsotti and Ulya Trofimovich. Efficient POSIX submatch extraction on nondeterministic finite automata. Software —Practice and Experience, 51(2):159–192, February 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Black:1981:CSF

[BTM81] James P. Black, David J. Taylor, and D. E. Morgan. A case study in fault tolerant software. Software —Practice and Experience, 11(2):145–157, February 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Blaich:2009:RVM

[BTS09] Andrew Blaich, Douglas Thain, and Aaron Striegel.

Reflections on the virtues of modularity: a case study in Linux security modules. Software—Practice and Experience, 39(15): 1235–1251, October ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Buhr:1994:ASM

[BTZ94] P. A. Buhr, David Till, and C. Robert Zarnke. Assignment as the sole means of updating objects. Software —Practice and Experience, 24(9):835–870, September 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Burrows:2007:EPD

[BTZ07] Steven Burrows, S. M. M.

Tahaghoghi, and Justin Zobel. Efficient plagiarism detection for large code repositories. Software—Practice and Experience, 37(2):151– 175, February 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Budgen:1985:CMM

[Bud85] David Budgen. Combining MASCOT with MODULA-2 to aid the engineering of real-time systems. Software—Practice and Experience, 15(8):767–793, August 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Budgen:1986:LE

[Bud86] D. Budgen. Letter to the Editor. Software—Practice and Experience, 16(6): 603, June 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Budd:1989:DOO

[Bud89] Timothy A. Budd. The design of an object-oriented command interpreter. Software—Practice and Experience, 19(1):35–51, January 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Buhr:1993:PPD

[Buh93] Raymond J. A. Buhr. Pictures that play: Design notations for real-time and distributed systems. Software—Practice and Experience, 23(8):895–931, August 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bull:1972:BRBa

[Bul72a] G. M. Bull. Book review: Basic programming, 2nd edn., John G. Kemeny and Thomas E. Kurtz, Wiley, London, 1971. No. of pages:

150. Price: £2.65. Software—Practice and Experience, 2(3):304, July 1972.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bull:1972:BRBb

[Bul72b] G. M. Bull. Book review:

Computing Terminals, Infotech State of the Art Report 4, F. D. Sherwood (Ed.), Infotech, Maidenhead, 1971. No. of pages:

358. Price: £40. Software —Practice and Experience, 2(3):304–305, July 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bull:1973:BRB

[Bul73] G. M. Bull. Book review: Programming timeshared computers in basic, Eugene H. Barnett, John Wiley, New York, 1972. No. of pages: 366. Price: £4.85. Software—Practice and Experience, 3(2):190, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bullers:1987:PAM

[Bul87] William I. Bullers, Jr. A processing algorithm for master-detail records in a relational database. Software—Practice and Experience, 17(10):701–717, October 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Burgess:1998:ASA

[Bur98] Mark Burgess. Automated system administration with feedback regulation. Software—Practice and Experience, 28(14):1519–1530, December 10, 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=5001668;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=5001668&PLACEBO=IE.
                                        pdf.

Burroughs:2016:RAS

[Bur16] Neil Burroughs. Register allocation and spilling using the expected distance heuristic. Software—Practice and Experience, 46 (11):1499–1523, November 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Burrows:2014:CTA

[BUT14] Steven Burrows, Alexandra L. Uitdenbogerd, and Andrew Turpin. Comparing techniques for authorship attribution of source code. Software—Practice and Experience, 44(1):1– 32, January 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Buxton:1978:BRB

[Bux78] J. N. Buxton. Book review: Software engineering. Edited by R. H. Perrott, Academic Press, London, 1977. No. of pages:

203. Price: £8.20. Software —Practice and Experience, 8(6):760, November 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Buyya:2000:PPS

[Buy00] Rajkumar Buyya. PAR-MON: a portable and scalable monitoring system for clusters. Software—Practice and Experience, 30(7): 723–739, June 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72001829/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72001829&PLACEBO=IE.
pdf.

Buyya:2021:GJY

[Buy21] R Buyya. Golden jubilee year of the software journal: Celebrating its vision, ground-breaking contributions and impact. Software —Practice and Experience, 51(1):3–4, January 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barford:1989:AGC

[BV89] Lee Alton Barford and Bradley T. Vander Zanden. Attribute grammars in constraint-based graphics systems. Software— Practice and Experience, 19(4):309–328, April 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Breuer:2006:RNO

[BV06] Peter T. Breuer and Marisol García Valls. Raiding the Noosphere: the open development of networked RAID support for the Linux kernel. Software —Practice and Experience, 36(4):365–395, April 10, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bainomugisha:2012:BSP

[BVB^{+}12] Engineer Bainomugisha, Jorge Vallejos, Elisa Gonzalez Boix, Pascal Costanza, Theo D’Hondt, and Wolfgang De Meuter. Bringing Scheme programming to the iPhone — experience. Software—Practice and Experience, 42(3):331– 356, March 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Basanta-Val:2011:NFI

[BVGVEA11] P. Basanta-Val, M. Garcia-Valls, and I. Estevez-Ayres.

Non-functional information transmission patterns for distributed real-time Java. Software—Practice and Experience, 41 (12):1409–1435, November/December 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Basanta-Val:2013:EOR

[BVGVEA13] P. Basanta-Val, M. García-Valls, and I. Estévez-Ayres. Enhancing OSGi with realtime Java support. Software—Practice and Experience, 43(1):33–65, January 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bell:1971:ALA

[BW71] D. A. Bell and B. A.

Wichmann. An ALGOLlike assembly language for a small computer. Software—Practice and Experience, 1(1):61–72, January/ March 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Boehm:1988:GCU

[BW88] Hans-Juergen Boehm and Mark Weiser. Garbage collection in an uncooperative environment. Software —Practice and Experience, 18(9):807–820, September 1988. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Burns:1995:EHR

[BW95] Alan Burns and A. J. Wellings. Engineering a hard real-time system: From theory to practice. Software—Practice and Experience, 25(7):705– 726, July 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Broughan:1996:FLT

[BW96] K. A. Broughan and D. M. K. Willcock. Fortran to Lisp translation using f2cl. Software—Practice and Experience, 26(10):1127–1139, October 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16765.

Brookes:1982:SAP

[BWA82] G. R. Brookes, I. R. Wilson, and A. M. Addyman. A static analysis of Pascal program structures. Software—Practice and Experience, 12(10):959–963, October 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baeza-Yates:1989:ISS

[BY89] Ricardo A. Baeza-Yates.

Improved string searching. Software—Practice and Experience, 19(3):257– 271, March 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Burton:1990:MMD

[BY90] F. W. Burton and Hsi-Kai Yang. Manipulating multilinked data structures in a pure functional language. Software—Practice and Experience, 20 (11):1167–1185, November 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bachelet:2017:DET

[BY17] Bruno Bachelet and Loïc Yon. Designing expression templates with concepts. Software—Practice and Experience, 47 (11):1521–1537, November 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Byrne:1991:SRE

[Byr91] Eric J. Byrne. Software reverse engineering: a case study. Software—Practice and Experience, 21 (12):1349–1364, December 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Briggs:2017:COI

[BZD17] Kim T. Briggs, Baoguo Zhou, and Gerhard W. Dueck. Cold object identification in the Java Virtual Machine. Software —Practice and Experience, 47(1):79–95, January 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brucker:2017:MVR

[BZM^{+}17] Achim D. Brucker, Bo Zhou, Francesco Malmignati, Qi Shi, and Madjid Merabti. Modelling, validating, and ranking of secure service compositions. Software—Practice and Experience, 47 (12):1923–1943, December 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cooke:1986:IFD

[CA86] D. J. Cooke and F. Abdollahzadeh. Insecurities in FORTRAN DOloops. Software—Practice and Experience, 16(3):201– 216 (or 201–215??), March 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cooke:2000:APC

[CA00] Daniel E. Cooke and Per Andersen. Automatic parallel control structures in SequenceL. Software— Practice and Experience, 30(14):1541–1570, November 25, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
73001716/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=73001716&PLACEBO=IE.
pdf.

Chatziantoniou:2008:SOA

[CA08a] Damianos Chatziantoniou and Achilleas Anagnostopoulos. A sessionoriented approach in modeling hierarchies of streams. Software—Practice and Experience, 38(1):95–115, January ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Crawford:2008:KAD

[CA08b] Heather Crawford and John Aycock. Kwyjibo: automatic domain name generation. Software— Practice and Experience, 38(14):1561–1567, November 25, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cuadrado:2014:SMM

[CA14] Jesús Sánchez Cuadrado and Jesús Perera Aracil. Scheduling model-to-model transformations with continuations. Software— Practice and Experience, 44(11):1351–1378, November 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cassol:2018:MIR

[CA18] Ignacio Cassol and Gabriela Arévalo. A methodology to infer and refactor an object-oriented model from C applications. Software—Practice and Experience, 48(3):550– 577, March 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2549.

Cockshott:1984:POM

[CAC^{+}84] W. P. Cockshott, Malcolm P. Atkinson, K. J. Chisholm, P. J. Bailey, and Ronald Morrison. Persistent object management system. Software—Practice and Experience, 14 (1):49–71, January 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Also published in/as: In ‘Readings in Object-Oriented Database Systems’, edited by S. Zdonik and D. Maier, Morgan Kaufman, 1990.

Chen:1994:ICT

[CAFH94] Shyh-Kwei Chen, Neal J.

Alewine, W. Kent Fuchs, and Wen-Mei W. Hwu. Incremental compiler transformations for multiple instruction retry. Software—Practice and Experience, 24(12):1179–1198, December 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cairns:1999:ETJ

[Cai99] Paul A. Cairns. Enumerated types in Java. Software—Practice and Experience, 29(3):291–297, March 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55000314;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55000314&PLACEBO=IE.
pdf.

Cucinotta:2018:IRT

[CALL18] Tommaso Cucinotta, Luca Abeni, Juri Lelli, and Giuseppe Lipari. Improving responsiveness of timesensitive applications by exploiting dynamic task dependencies. Software— Practice and Experience, 48(4):820–841, April 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2557.

Campbell:1985:BRA

[Cam85] J. A. Campbell. Book review: Artificial intelligence. P. H. Winston, second edition, Addison-Wesley, 1984. price: £15.95. Software—Practice and Experience, 15(6):623–624, June 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Carter:1979:FMU

[Car79] B. S. Carter. File maintenance using a line editor. Software—Practice and Experience, 9(6):457– 462, June 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cargill:1981:FSE

[Car81] T. A. Cargill. Full-screen editing in a hostile environment. Software—Practice and Experience, 11(9): 975–981, September 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cargill:1982:RDS

[Car82] T. A. Cargill. A robust distributed solution to the dining philosophers problem. Software—Practice and Experience, 12(10): 965–969, October 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cargill:1985:IBD

[Car85a] Thomas A. Cargill. Implementation of the Blit debugger. Software—Practice and Experience, 15(2):153– 168, February 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Carter:1985:NPM

[Car85b] B. S. Carter. NEWLONG:

a program for manipulation of data on populations of individuals subject to change. Software—Practice and Experience, 15(4):369– 377, April 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Carrington:1986:PUE

[Car86] David A. Carrington. Profiling under ELXSI UNIX. Software—Practice and Experience, 16(9):865–873, September 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Caron:1997:ASM

[Car97] Pascal Caron. AG: a set of Maple packages for manipulating automata and semigroups. Software—Practice and Experience, 27(8): 863–884, August 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7323;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7323&
PLACEBO=IE.pdf.

Carroll:1998:AOM

[Car98] Martin D. Carroll. Active objects made easy. Software—Practice and Experience, 28(1):1–21, January 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61001671;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=61001671&PLACEBO=IE.
pdf.

Carter:2022:CSB

[Car22] Lynn Robert Carter. Climbing out of a software black hole. Software—Practice and Experience, 52(2):319– 336, February 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Campanoni:2010:HFP

[CARB10] Simone Campanoni, Giovanni Agosta, Stefano Crespi Reghizzi, and Andrea Di Biagio. A highly flexible, parallel virtual machine: design and experience of ILDJIT. Software—Practice and Experience, 40 (2):177–207, February ??, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cashin:1992:ROS

[Cas92] Jerry Cashin. Road to open systems begins with first step. Software Magazine, 12(1):84–??, January 1992. CODEN SMWMEQ. ISSN 0897-8085.

Cavouras:1983:BRB

[Cav83a] John C. Cavouras. Book review: Coroutines, Christopher D. Marlin, Lecture notes in computer science, vol. 95, Springer-Verlag, Berlin, Heidelberg, 1980. ISBN 3-540-10256-6 and ISBN 0-387-10256-6. no. of pages: 246. Price: DM29.50. Software —Practice and Experience, 13(10):975–976, October 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cavouras:1983:IST

[Cav83b] John C. Cavouras. Implementing a simulation tool in a high-level language with no multitasking facilities. Software—Practice and Experience, 13(9):

809–815, September 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cronin:1972:HSC

[CB72] D. E. Cronin and J. P. Brandon. A high speed computer-to-computer data link. Software—Practice and Experience, 2(2):173– 186, April/June 1972. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chawla:2000:MMP

[CB00a] Rohit Chawla and Steve Baumel. Managing more physical with less virtual. Software—Practice and Experience, 30(6):639– 661, May 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        71007585/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=71007585&PLACEBO=IE.
                                        pdf.

Coulson:2000:EID

[CB00b] Geoff Coulson and Shakuntala Baichoo. Experiences in implementing a distributed object platform for multimedia applications. Software—Practice and Experience, 30(6): 663–683, May 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
71007584/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=71007584&PLACEBO=IE.
pdf.

Chauhan:2017:SPA

[CBB17] Muhammad Aufeef Chauhan, Muhammad Ali Babar, and Boualem Benatallah. Survery paper: Architecting cloud-enabled systems: a systematic survey of challenges and solutions. Software—Practice and Experience, 47(4):599–644, April 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Casini:2020:TII

[CBB20] Daniel Casini, Alessandro Biondi, and Giorgio Buttazzo. Timing isolation and improved scheduling of deep neural networks for real-time systems. Software—Practice and Experience, 50(9):1760–1777, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cortes:2000:SCR

[CBC00] Toni Cortes, Yolanda Becerra, and Raúl Cervera. Swap compression: resurrecting old ideas. Software—Practice and Experience, 30(5):567–587, April 25, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
71004434/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=71004434&PLACEBO=IE.
pdf.

Constantino:2023:DAH

[CBF23] Kattiana Constantino, Fabiano Belém, and Eduardo Figueiredo. Dual analysis for helping developers to find collaborators based on co-changed files: an empirical study. Software—Practice and Experience, 53(6): 1438–1464, June 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Calheiros:2010:BAS

[CBR10] Rodrigo N. Calheiros, Rajkumar Buyya, and César A. F. De Rose. Building an automated and self-configurable emulation testbed for grid applications. Software—Practice and Experience, 40(5):405– 429, April 25, 2010. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Calderbank:1973:PLS

[CC73] M. Calderbank and V. J.

Calderbank. A portable language for system development. Software— Practice and Experience, 3(4):309–321, October/ December 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cohen:1977:LIA

[CC77] Jacques Cohen and Neal Carpenter. A language for inquiring about the run-time behaviour of programs. Software—Practice and Experience, 7(4):445– 460, July/August 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Carpenter:1984:ERP

[CC84] B. E. Carpenter and Robert Cailliau. Experience with remote procedure calls in a real-time control system. Software—Practice and Experience, 14(9): 901–907, September 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chan:1987:ADD

[CC87] Fun Ting Chan and Tsong Yueh Chen. AIDA: a dynamic data flow anomaly detection system for Pascal programs. Software—Practice and Experience, 17(3):227– 239, March 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Campbell:1990:E

[CC90] John Campbell and Douglas Comer. Editorial. Software—Practice and Experience, 20(1):1, January 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:1997:RRE

[CC97] Ye-In Chang and Fwo-Long Chen. RBE: A ruleby-example active database system. Software—Practice and Experience, 27(4): 365–394, April 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7298;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7298&
PLACEBO=IE.pdf.

Chou:2000:PPC

[CC00] Shih-Chien Chou and Jen-Yen Jason Chen. Process program change control in a process environment. Software—Practice and Experience, 30(3):175– 197, March 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
71001773/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=71001773&PLACEBO=IE.
pdf.

Chang:2001:EEJ

[CC01] Da-Wei Chang and Ruei-Chuan Chang. EJVM: an economic Java runtime environment for embedded devices. Software—Practice and Experience, 31(2):129–146, February 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
76504935/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=76504935&PLACEBO=IE.
pdf.

Carey:2002:LLB

[CC02] James Carey and Brent Carlson. Lessons learned becoming a framework developer. Software—Practice and Experience, 32(8): 789–800, July 10, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www3.interscience.
wiley.com/cgi-bin/abstract/
94519520/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?

                                        ID=94519520&PLACEBO=IE.
                                        pdf.

Chan:2013:QSS

[CC13] W. K. Chan and Yan Cai. In quest of the science in statistical fault localization. Software—Practice and Experience, 43(8):971– 987, August 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:2018:PNC

[CC18] Hsung-Pin Chang and Cheng-Pang Chiang. PARC: a novel OS cache manager. Software—Practice and Experience, 48 (12):2193–2222, December 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chou:1996:BBC

[CCC96] Shih-Chien Chou, Jen-Yen Chen, and Chyan-Goei Chung. A behaviorbased classification and retrieval technique for object-oriented specification reuse. Software— Practice and Experience, 26(7):815–832, July 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        http://www3.interscience.
                                        wiley.com/cgi-bin/abstract?
                                        ID=16819.

Cardoso:2016:PDI

[CCC^{+}16] João M. P. Cardoso, José G. F. Coutinho, Tiago Carvalho, Pedro C. Diniz, Zlatko Petrov, Wayne Luk, and Fernando Gonçalves. Performance-driven instrumentation and mapping strategies using the LARA aspect-oriented programming approach. Software —Practice and Experience, 46(2):251–287, February 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chan:2005:VPS

[CCCZ05] Fan Chan, Jiannong Cao, Alvin T. S. Chan, and Kang Zhang. Visual programming support for graph-oriented parallel/ distributed processing. Software—Practice and Experience, 35(15):1409–1439, December 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:1999:MHP

[CCE99] Chi-Chao Chang, Grzegorz Czajkowski, and Thorsten Von Eicken. MRPC: a high performance RPC system for MPMD parallel computing. Software—Practice and Experience, 29(1):43– 66, January 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55002654;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55002654&PLACEBO=IE.
pdf.

Chong:2021:CLM

[CCE^{+}21] Nathan Chong, Byron Cook, Jonathan Eidelman, Konstantinos Kallas, Kareem Khazem, Felipe R. Monteiro, Daniel Schwartz-Narbonne, Serdar Tasiran, Michael Tautschnig, and Mark R. Tuttle. Code-level model checking in the software development workflow at Amazon Web Services. Software—Practice and Experience, 51(4):772– 797, April 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Castor:2009:MRE

[CCF^{+}09] Fernando Castor, Nélio Cacho, Eduardo Figueiredo, Alessandro Garcia, Cecília M. F. Rubira, Jefferson Silva de Amorim, and Hítalo Oliveira da Silva. On the modularization and reuse of exception handling with aspects. Software— Practice and Experience, 39(17):1377–1417, December 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chicote:2014:PJV

[CCG14] M. Chicote, D. Ciolek, and J. P. Galeotti. Practical JFSL verification using TACO. Software—Practice and Experience, 44(3):317– 334, March 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marimuthu:2021:HDO

[CCK21] Marimuthu C, Sridhar Chimalakonda, and Chandrasekaran K. How do open source app developers perceive API changes related to Android battery optimization? An empirical study. Software—Practice and Experience, 51(4):691– 710, April 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Camacho:2022:CSP

[CCLN22] Carlos Camacho, Pablo C.

Cañizares, Luis Llana, and Alberto Núñez. Chaos as a Software Product Line — a platform for improving open hybrid-cloud systems resiliency. Software—Practice and Experience, 52(7): 1581–1614, July 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Canfora:1996:IAI

[CCM96] G. Canfora, Aniello Cimitile, and M. Munro. An improved algorithm for identifying objects in code. Software—Practice and Experience, 26(1):25–48, January 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=61004284.

Chirouze:2005:SMA

[CCM05] Olivier Chirouze, David Cleary, and George G. Mitchell. A software methodology for applied research: eXtreme Researching. Software—Practice and Experience, 35 (15):1441–1454, December 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2006:MCM

[CCP06] Bo Chen, Harry H. Cheng, and Joe Palen. Mobile-C: a mobile agent platform for mobile C/C++ agents. Software—Practice and Experience, 36 (15):1711–1733, December 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cabodi:1991:TET

[CCPR91] G. Cabodi, P. Camurati, P. Prinetto, and M. Sonza Reorda. TPDL*: Extended temporal profile description language. Software —Practice and Experience, 21(4):355–374, April 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chan:2012:EES

[CCPY12] Eric Ying Kwong Chan, W. K. Chan, Pak-Lok Poon, and Yuen Tak Yu. An empirical evaluation of several test-a-few strategies for testing particular conditions. Software—Practice and Experience, 42(8):967– 994, August 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cabodi:2016:GLA

[CCQ16] Gianpiero Cabodi, Paolo Camurati, and Stefano Quer. A graph-labeling approach for efficient coneof-influence computation in model-checking problems with multiple properties. Software—Practice and Experience, 46(4):493– 511, April 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chawla:2019:FCB

[CCR19] Priyanka Chawla, Inderveer Chana, and Ajay Rana. Framework for cloud-based software test data generation service. Software—Practice and Experience, 49(8):1307– 1328, August 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Celentano:1980:CTU

[CCRD^{+}80] Augusto Celentano, Stefano Crespi-Reghizzi, Pierluigi Della Vigna, Carlo Ghezzi, G. Granata, and F. Savoretti. Compiler testing using a sentence generator. Software—Practice and Experience, 10(11): 897–918, November 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2001:JJB

[CCT01] Ding-Yi Chen, Tyng-Ruey Chuang, and Shi-Chun Tsai. JGAP: a Javabased graph algorithms platform. Software—Practice and Experience, 31(7): 615–635, June 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        78003148/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=78003148&PLACEBO=IE.
                                        pdf.

Citrin:1995:UFT

[CCvKH95] Wayne Citrin, Alistair Cockburn, Jürg von Känel, and Rainer Hauser. Using formalized temporal message-flow diagrams. Software—Practice and Experience, 25(12):1367– 1401, December 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Casey:1982:MVM

[CD82] Bernice E. Casey and B. Dasarathy. Modelling and validating the manmachine interface. Software —Practice and Experience, 12(6):557–569, June 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ciechanowicz:1984:CPT

[CD84] Z. J. Ciechanowicz and A. C. De Weever. The ‘completeness’ of the Pascal test suite. Software —Practice and Experience, 14(5):463–471, May 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cimitile:1991:REA

[Cd91] Aniello Cimitile and Ugo de Carlini. Reverse engineering: Algorithms for program graph production. Software—Practice and Experience, 21(5):519– 537, May 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cannon:1994:AFT

[CD94] Scott R. Cannon and David Dunn. Adding faulttolerant transaction processing to LINDA. Software —Practice and Experience, 24(5):449–466, May 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ciura:2001:HSL

[CD01] Marcin G. Ciura and Sebastian Deorowicz. How to squeeze a lexicon. Software—Practice and Experience, 31(11):1077–1090, September 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        85004330/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=85004330&PLACEBO=IE.
                                        pdf.

Cretella:2015:SEP

[CD15] Giuseppina Cretella and Beniamino Di Martino. A semantic engine for porting applications to the cloud and among clouds. Software—Practice and Experience, 45(12):1619–1637, December 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Claveirole:2012:MWF

[CdA12] Thomas Claveirole and Marcelo Dias de Amorim. Manipulating Wi-Fi packet traces with WiPal: design and experience. Software —Practice and Experience, 42(5):585–599, May 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Colarusso:2024:DTP

[CDF^{+}24] Carmine Colarusso, Assunta De Caro, Ida Falco, Lorenzo Goglia, and Eugenio Zimeo. A distributed tracing pipeline for improving locality awareness of microservices applications. Software—Practice and Experience, 54(6): 1118–1140, June 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ciancarini:2012:HQP

[CDFV12] Paolo Ciancarini, Angelo Di Iorio, Luca Furini, and Fabio Vitali. Highquality pagination for publishing. Software—Practice and Experience, 42(6):733– 751, June 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chien:1998:EHL

[CDG^{+}98] Andrew A. Chien, Julian Dolby, Bishwaroop Gangul, Vijay Karamcheti, and Xingbin Zhang. Evaluating high level parallel programming support for irregular applications in ICC++. Software—Practice and Experience, 28(11):1213–1243, September 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=10007358;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=10007358&PLACEBO=IE.
pdf.

Cowan:1993:CPG

[CDGP93] D. D. Cowan, C. M. Durance, E. Giguère, and G. M. Pianosi. CIRL/ PIWI: a GUI toolkit supporting retargetability. Software—Practice and Experience, 23(5):511– 527, May 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Coulouris:1976:DII

[CDH^{+}76] G. F. Coulouris, I. Durham, J. R. Hutchinson, M. H. Patel, T. Reeves, and D. G. Winderbank. The design and implementation of an interactive document editor. Software—Practice and Experience, 6(2):271– 279, April/June 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Carpenter:1977:NRR

[CDH77] B. E. Carpenter, R. W. Doran, and K. Hopper. Nonrecursive recursion. Software—Practice and Experience, 7(2):263–269, March/ April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chou:1985:DIW

[CDKK85] Hong-Tai Chou, David J.

DeWitt, Randy H. Katz, and Anthony C. Klug. Design and implementation of the Wisconsin Storage System. Software—Practice and Experience, 15(10): 943–962, October 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ciancarini:2016:BGB

[CDM^{+}16] Paolo Ciancarini, Angelo Di Iorio, Carlo Marchetti, Michele Schirinzi, and Fabio Vitali. Bridging the gap between tracking and detecting changes in XML. Software—Practice and Experience, 46(2):227– 250, February 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Canete:2013:WSN

[CDR13] Eduardo Cañete, Manuel Díaz, and Bartolomé Rubio. A wireless sensor network framework based on light databases. Software —Practice and Experience, 43(5):501–523, May 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Canonico:2003:IQA

[CDRV03] Roberto Canonico, Maurizio D’Arienzo, Simon Pietro Romano, and Giorgio Ventre. Introducing QoS awareness in distributed programming: QTcl. Software—Practice and Experience, 33(10):901–911, August 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ciminiera:1988:PSS

[CDV88] Luigi Ciminiera, C. Demartini, and Adriano Valenzano. A practical study of semi-automatic OSI protocol implementation. Software—Practice and Experience, 18(3):255–277, March 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Crookes:1984:ELD

[CE84] D. Crookes and J. W. G.

Elder. An experiment in language design for distributed systems. Software —Practice and Experience, 14(10):957–971, October 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chee:1997:IVC

[CE97] Chye-Lin L. Chee and Sevki S. Erdogan. An installable Version Control File System for UNIX. Software—Practice and Experience, 27(6):725–746, June 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7313;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7313&
PLACEBO=IE.pdf.

Constantinides:2002:EOM

[CEF02] Constantinos A. Constantinides, Tzilla Elrad, and Mohamed E. Fayad. Extending the object model to provide explicit support for crosscutting concerns. Software—Practice and Experience, 32(7):703– 734, June 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.

                                        com/cgi-bin/abstract/
                                        94515735/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=94515735&PLACEBO=IE.
                                        pdf.

Celko:1982:TTD

[Cel82] Joe Celko. Time token design methodology. Software —Practice and Experience, 12(10):889–895, October 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cederbladh:2024:ECD

[CEMS24] Johan Cederbladh, Romina Eramo, Vittoriano Muttillo, and Per Erik Strandberg. Experiences and challenges from developing cyber–physical systems in industry–academia collaboration. Software—Practice and Experience, 54(6): 1193–1212, June 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ceraolo:2018:ECS

[Cer18] Massimo Ceraolo. Experiences in creating a software tool to analyze and postprocess simulated and measured data. Software—Practice and Experience, 48(12):2380–2388, December 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Coutant:1980:DDD

[CF80] Cary A. Coutant and Christopher W. Fraser. A device driver for display terminals. Software—Practice and Experience, 10(3): 183–187, March 1980. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chivers:2005:ASD

[CF05] Howard Chivers and Martyn Fletcher. Applying Security Design Analysis to a service-based system. Software—Practice and Experience, 35(9):873– 897, July 25, 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Costantini:2015:SAD

[CFC15] Giulia Costantini, Pietro Ferrara, and Agostino Cortesi. A suite of abstract domains for static analysis of string values. Software —Practice and Experience, 45(2):245–287, February 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cuculovic:2022:SRD

[CFD^{+}22] Milos Cuculovic, Frederic Fondement, Maxime Devanne, Jonathan Weber, and Michel Hassenforder. Semantics to the rescue of document-based XML diff: a JATS case study. Software—Practice and Experience, 52(6): 1496–1516, June 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chhabra:2017:EOP

[CFKT17] Tamanna Chhabra, Simone Faro, M. Oğuzhan Külekci, and Jorma Tarhio. Engineering order-preserving pattern matching with SIMD parallelism. Software —Practice and Experience, 47(5):731–739, May 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Corsini:1984:DRA

[CFL84] P. Corsini, G. Frosini, and Lanfranco Lopriore. Distributing and revoking access authorizations on abstract objects: a capability approach. Software— Practice and Experience, 14(10):931–943, October 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Caprara:1998:ICL

[CFL^{+}98] A. Caprara, F. Focacci, E. Lamma, P. Mello, M. Milano, P. Toth, and D. Vigo. Integrating constraint logic programming and operations research techniques for the Crew Rostering Problem. Software—Practice and Experience, 28(1):49–76, January 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61001670;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=61001670&PLACEBO=IE.
pdf.

Cao:2014:CLP

[CFLC14] Jian Cao, Jiwen Fu, Minglu Li, and Jinjun Chen. CPU load prediction for cloud environment based on a dynamic ensemble model. Software—Practice and Experience, 44(7):793– 804, July 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Crookes:1983:BSG

[CFP83] D. Crookes, R. Fee, and V. Pickering. Building syntax graphs from syntax equations: a case study in modular programming. Software—Practice and Experience, 13 (12):1129–1139, December 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2022:BHI

[CFS^{+}22] Songling Chen, Xiangling Fu, Hongchao Si, Yi Wang, Songfeng Gao, and Chunyuan Wang. Blockchain for health IoT: a privacypreserving data sharing system. Software—Practice and Experience, 52 (9):2026–2044, September 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Collofello:1993:ACA

[CG93] James S. Collofello and Bakul P. Gosalla. An application of causal analysis to the software modification process. Software—Practice and Experience, 23(10):1095–1105, October 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chodrow:1995:ISS

[CG95a] Sarah E. Chodrow and Mohamed G. Gouda. Implementation of the sentry system. Software—Practice and Experience, 25(4):373– 387, April 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cifuentes:1995:DBP

[CG95b] Cristina Cifuentes and K. John Gough. Decompilation of binary programs. Software—Practice and Experience, 25(7):811– 829, July 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chan:1996:FVH

[CG96] Toong Shoon Chan and Ian Gorton. Formal validation of a high performance error control protocol using SPIN. Software—Practice and Experience, 26(1):105– 124, January 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61004283.

Chen:2004:WBD

[CGH^{+}04] Shu-Ching Chen, Sneh Gulati, Shahid Hamid, Xin Huang, Lin Luo, Nirva Morisseau-Leroy, Mark D. Powell, Chengjun Zhan, and Chengcui Zhang. A Web-based distributed system for hurricane occurrence projection. Software —Practice and Experience, 34(6):549–571, May 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2008:SED

[CGH08] Qi Chen, John Grundy, and John Hosking. SUM-LOW: early design-stage sketching of UML diagrams on an E-whiteboard. Software—Practice and Experience, 38(9):961–994, July 25, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2015:MCA

[CGH^{+}15] Zhe Chen, Yi Gu, Zhiqiu Huang, Jun Zheng, Chang Liu, and Ziyi Liu. Model checking aircraft controller software: a case study. Software—Practice and Experience, 45(7):989– 1017, July 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Coleman:1979:ACP

[CGHP79] Derek Coleman, Robin M. Gallimore, J. W. Hughes, and M. S. Powell. An assessment of Concurrent Pascal. Software—Practice and Experience, 9(10):827– 837, October 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Castiglione:2015:MPC

[CGIP15] Aniello Castiglione, Marco Gribaudo, Mauro Iacono, and Francesco Palmieri. Modeling performances of concurrent big data applications. Software—Practice and Experience, 45(8): 1127–1144, August 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:1989:DIV

[CGK89] Ellis E. Chang, David Gedye, and Randy H. Katz. The design and implementation of a version server for computer-aided design. Software—Practice and Experience, 19(3):199– 222, March 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Coleman:1976:VSS

[CGL76] M. J. Coleman, G. S.

Godliman, and G. L. Leonard. The Viscom — a system of ‘soft’ videoterminals. Software— Practice and Experience, 6(4):569–576, October/ December 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Calvert:2003:LEN

[CGM^{+}03] K. Calvert, J. Griffioen, B. Mullins, S. Natarajan, L. Poutievski, A. Sehgal, and S. Wen. Leveraging emerging network services to scale multimedia applications. Software— Practice and Experience, 33(14):1377–1397, November 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Canfora:2006:AFI

[CGP^{+}06] Gerardo Canfora, Félix García, Mario Piattini, Francisco Ruiz, and C. A. Visaggio. Applying a framework for the improvement of software process maturity. Software—Practice and Experience, 36(3): 283–304, March 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cohen:2000:DRN

[CGR00] W. E. Cohen, R. K. Gaede, and J. B. Rodgers. A demonstration of repeatable, non-intrusive measurement of program performance and compiler optimization in Linux using IN-Tune. Software —Practice and Experience, 30(8):895–906, July 10, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72502206/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72502206&PLACEBO=IE.
pdf.

Cowan:1980:DDA

[CGWL80] D. D. Cowan, J. W. Graham, J. W. Welch, and Carlos J. P. Lucena. A data-directed approach to program construction. Software—Practice and Experience, 10(5): 355–372, May 1980. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cheng:2020:ILT

[CGZ^{+}20] Wen Cheng, Tao Guo, Lingfang Zeng, Yang Wang, Lars Nagel, Tim Süß, and André Brinkmann. Improving LSM-trie performance by parallel search. Software—Practice and Experience, 50(10):1952– 1965, October 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Charlton:1973:NRS

[CH73] C. C. Charlton and P. G.

Hibbard. A note on recreating source code from the reverse Polish form. Software —Practice and Experience, 3(2):151–153, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:1988:IPP

[CH88] Pehong Chen and Michael A.

Harrison. Index preparation and processing. Software—Practice and Experience, 18(9):897–915, September 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). The L^{A}TEX text of this paper is included in the makeindex software distribution.

Cordy:1990:CGU

[CH90] James R. Cordy and Richard C. Holt. Code generation using an orthogonal model. Software—Practice and Experience, 20(3):301– 320, March 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:2006:SXD

[CH06] Tao-Ku Chang and Gwan-Hwan Hwang. To secure XML documents with the extension function of XSLT. Software—Practice and Experience, 36(5): 539–555, April 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chapin:1974:NFF

[Cha74] Ned Chapin. New format for flowcharts. Software—Practice and Experience, 4(4):341–357, October/December 1974. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Challab:1988:EMB

[Cha88] D. J. Challab. Elastic memory: Buddy systems interface. Software—Practice and Experience, 18 (12):1139–1155, December 1988. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2017:BBD

[CHC^{+}17] Dan Chen, Yangyang Hu, Chang Cai, Ke Zeng, and Xiaoli Li. Brain big data processing with massively parallel computing technology: challenges and opportunities. Software—Practice and Experience, 47(3): 405–420, March 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:2007:VMS

[CHCC07] Da-Wei Chang, Cheng-En Hsieh, Yan-Pai Chen, and Kwo-Cheng Chiu. Virtual machine support for zeroloss Internet service recovery and upgrade. Software —Practice and Experience, 37(13):1349–1376, November 10, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2004:BNS

[Che04] Jing-Chao Chen. Building a new sort function for a C library. Software—Practice and Experience, 34(8): 777–795, July 10, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2008:SPS

[Che08] Jing-Chao Chen. Symmetry Partition Sort. Software —Practice and Experience, 38(7):761–773, June ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2022:RRL

[Che22]  Zhaoyun Chen.  RI-

FLING: a reinforcement learning-based GPU scheduler for deep learning research and development platforms. Software—Practice and Experience, 52(6): 1319–1336, June 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chivers:2017:OUR

[Chi17] Howard Chivers. Optimising Unicode regular expression evaluation with previews. Software—Practice and Experience, 47(5):669– 688, May 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chou:1996:CAL

[Cho96] Seng-Cho Timothy Chou. Colony: An artificial life model for active autonomous objects. Software—Practice and Experience, 26(12):1373–1384, December 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16779.

Chou:1998:MMP

[Cho98] Seng-Cho T. Chou. Managing multi-party documents in a network computing environment. Software—Practice and Experience, 28(12):1315–1325, October 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1766;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1766&
PLACEBO=IE.pdf.

Christopher:1984:RCG

[Chr84] Thomas W. Christopher.

Reference count garbage collection. Software—Practice and Experience, 14(6): 503–507, June 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Czajkowski:2005:RMI

[CHS^{+}05] Grzegorz Czajkowski, Stephen Hahn, Glenn Skinner, Pete Soper, and Ciarán Bryce. A resource management interface for the Java^{T M} platform. Software—Practice and Experience, 35(2): 123–157, February 2005.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cooper:1991:EIS

[CHT91] Keith D. Cooper, Mary W. Hall, and Linda Torczon. An experiment with inline substitution. Software —Practice and Experience, 21(6):581–601, June 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cooper:1998:HBI

[CHT98] Keith D. Cooper, Timothy J. Harvey, and Linda Torczon. How to build an interference graph. Software—Practice and Experience, 28(4):425–444, April 10, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=1788;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=1788&
                                        PLACEBO=IE.pdf.

Chvalosky:1979:NTC

[Chv79] Václav Chvaloský. New techniques for COBOL programmers. Software— Practice and Experience, 9 (2):87–95, February 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chvalosky:1983:DT

[Chv83] Václav Chvaloský. Decision tables. Software —Practice and Experience, 13(5):423–429, May 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cai:2022:SCO

[CHY^{+}22] Weiwei Cai, Fazhi He, Shangxu Yang, Xiao Lv, and Yuan Cheng. Selfcompressing object sequence for consistency maintenance in co-editors. Software—Practice and Experience, 52(8):1802– 1825, August 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cheung:2003:DOO

[CI03] Kent K. T. Cheung and Horace H. S. Ip. Developing an object-oriented framework for contentbased image retrieval. Software—Practice and Experience, 33(6):523–565, May 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ciabrini:2007:SVS

[Cia07] Damien Ciabrini. Stack virtualization for source level debugging. Software —Practice and Experience, 37(7):693–725, June 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Correia:2022:EDM

[CIM22] Alexandre R. S. Correia, Juliano M. Iyoda, and Alexandre C. Mota. The effect of distance metrics in a general purpose synthesizer: an empirical study on integer domain imperative programs. Software—Practice and Experience, 52(5): 1294–1304, May 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Correia:2024:EDM

[CIM24] Alexandre R. S. Correia, Juliano M. Iyoda, and Alexandre C. Mota. The effect of distance metrics in a general purpose synthesizer of imperative programs: a second empirical study using enlarged search spaces. Software—Practice and Experience, 54(3):528– 540, March 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cutbill:1973:GEN

[CJ73] John L. Cutbill and Karen Sparck Jones. Guest editorials: The needs of the pseudo-pro. Software—Practice and Experience, 3(2):83–86, April/ June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cooper:1988:MOI

[CJ88] R. E. M. Cooper and G. Jones. A microprogrammed occam interpreter for the HLH Orion. Software—Practice and Experience, 18(1):63– 71, January 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chhikara:2022:DDR

[CJTK22] Prateek Chhikara, Nikhil Jain, Rajkumar Tekchandani, and Neeraj Kumar. Data dimensionality reduction techniques for Industry 4.0: Research results, challenges, and future research directions. Software—Practice and Experience, 52(3):658–688, March 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cunningham:1978:EPD

[CK78] R. J. Cunningham and J. Kramer. An exercise in program design using SIMULA class invariants. Software—Practice and Experience, 8(3):355– 369, May/June 1978. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cantoni:1986:TWB

[CK86] A. Cantoni and L. Kleeman. Three way branching self consistency checking of hardware and software. Software—Practice and Experience, 16 (11):1053–1054, November 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Carr:1994:SRP

[CK94] Steve Carr and Ken Kennedy. Scalar replacement in the presence of conditional control flow. Software—Practice and Experience, 24(1):51– 77, January 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Choi:1997:EMV

[CK97] Esther Jinee Choi and Yong Rae Kwon. An efficient method for version control of a tree data structure. Software—Practice and Experience, 27(7):797– 811, July 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7318;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7318&
PLACEBO=IE.pdf.

Carlberger:1999:IEP

[CK99] Johan Carlberger and Viggo Kann. Implementing an efficient part-of-speech tagger. Software—Practice and Experience, 29(9):815– 832, July 25, 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=62501863;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=62501863&PLACEBO=IE.
pdf.

Campbell-Kelly:2013:ODB

[CK13] M. Campbell-Kelly. Obituaries: David Barron: a life in software, 1935– 2012. Software—Practice and Experience, 43(6):733– 741, June 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Choi:2015:IMA

[CK15] Dongjin Choi and Pankoo Kim. Identifying the most appropriate expansion of acronyms used in Wikipedia text. Software —Practice and Experience, 45(8):1073–1086, August 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chae:2000:CMO

[CKB00] Heung Seok Chae, Yong Rae Kwon, and Doo Hwan Bae. A cohesion measure for object-oriented classes. Software—Practice and Experience, 30(12):1405– 1431, October 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
73000927/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=73000927&PLACEBO=IE.

pdf. See comments [XZ01, CKB01, XZ03, CKB03].

Chae:2001:RCC

[CKB01] Heung-Seok Chae, Yong-Rae Kwon, and Doo-Hwan Bae. Response to comments on: “A cohesion measure for object-oriented classes”. Software—Practice and Experience, 31 (14):1389–1392, November 25, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
85515671/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=85515671&PLACEBO=IE.

pdf. See comments [CKB00, XZ01, XZ03, CKB03].

Chae:2003:RMC

[CKB03] Heung-Seok Chae, Yong-Rae Kwon, and Doo-Hwan Bae. Response to More comments on: “A cohesion measure for object-oriented classes”. Software—Practice and Experience, 33(6): 589–594, May 2003. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [CKB00, CKB01, XZ01, XZ03].

Chakraborty:2023:JCT

[CKCG23] Ananya Chakraborty, Mohit Kumar, Nisha Chaurasia, and Sukhpal Singh Gill. Journey from cloud of things to fog of things: Survey, new trends, and research directions. Software —Practice and Experience, 53(2):496–551, February 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cha:2002:MME

[CKL^{+}02] Sang K. Cha, Kihong Kim, Byung S. Lee, Changbin Song, Sangyong Hwang, and Yongsik Kwon. MEADOW: a middleware for efficient access to multiple geographic databases through OpenGIS wrappers. Software—Practice and Experience, 32(4):377– 402, April 10, 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
91016435/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=91016435&PLACEBO=IE.
pdf.

Capilla:2020:UIS

[CKRC20] Rafael Capilla, Rick Kazman, Carlos Romera, and Carlos Carrillo. Usability implications in software architecture: the case study of a mobile app. Software—Practice and Experience, 50(12):2145–2168, December 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chuang:2002:NIO

[CKW02] Tyng-Ruey Chuang, Y. S. Kuo, and Chien-Min Wang. Non-intrusive object introspection in C++. Software—Practice and Experience, 32(2):191–207, February 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
88511128/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=88511128&PLACEBO=IE.
pdf.

Charlton:1981:ETP

[CL81] C. C. Charlton and P. H.

Leng. Editors: Two for the price of one. Software —Practice and Experience, 11(2):195–202, February 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cook:1982:CAP

[CL82] Robert P. Cook and Insup Lee. A contextual analysis of Pascal programs. Software—Practice and Experience, 12(2):195– 203, February 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Charlton:1983:APE

[CL83] C. C. Charlton and P. H.

Leng. Aids for pragmatic error detection. Software —Practice and Experience, 13(1):59–66, January 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Corsini:1995:ISM

[CL95] Paolo Corsini and Lanfranco Lopriore. An implementation of storage management in capability environments. Software—Practice and Experience, 25(5): 501–520, May 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Calefato:2009:UFD

[CL09] Fabio Calefato and Filippo Lanubile. Using frameworks to develop a distributed conferencing system: an experience report. Software—Practice and Experience, 39(15): 1293–1311, October ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ceraolo:2019:UML

[CL19] M. Ceraolo and G. Lutzemberger. Use of Modelica language to simulate electrified railway lines and trains. Software—Practice and Experience, 49(7): 1114–1130, July 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Clausecker:2023:TUC

[CL23] Robert Clausecker and Daniel Lemire. Transcoding Unicode characters with AVX-512 instructions. Special focus section: Benchmarking, Experimentation Tools, and Reproducible Practices for Data-Intensive Systems from Edge to Cloud: Software: Practice and Experience, 53(12):2430–2462, December 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Clayton:1986:CGS

[Cla86] Peter G. Clayton. A code generator synthesizer for the non-specialist. Software—Practice and Experience, 16(8):751–760, August 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Clarke:1989:OPC

[Cla89]  Keith Clarke.  One-

pass code generation using continuations. Software—Practice and Experience, 19(12):1175–1192, December 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Clark:1998:BRB

[Cla98]  John A. Clark.  Book

review: Software fault injection. Software— Practice and Experience, 28(15):1637, December 25, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=10050308;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=10050308&PLACEBO=IE.
pdf.

Chiang:1999:UDC

[CLC99] Mei-Ling Chiang, Paul C. H. Lee, and Ruei-Chuan Chang. Using data clustering to improve cleaning performance for flash memory. Software—Practice and Experience, 29(3): 267–290, March 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55000313;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55000313&PLACEBO=IE.
pdf.

Chang:2009:SSP

[CLC09] Hsung-Pin Chang, Yu-Chieh Lin, and Shiu-Hau Chen. System supports for protocol and application adaptation in vertical handoffs. Software—Practice and Experience, 39(8): 739–771, June 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:2015:PDC

[CLCC15] Hsung-Pin Chang, Syuan-You Liao, Da-Wei Chang, and Guo-Wei Chen. Profit data caching and hybrid disk-aware Completely Fair Queuing scheduling algorithms for hybrid disks. Software—Practice and Experience, 45(9):1229– 1249, September 2015. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2017:FHL

[CLD^{+}17] Yunliang Chen, Fangyuan Li, Ze Deng, Xiaodao Chen, and Jijun He. P M_{2.5} forecasting with hybrid LSE model-based approach. Software—Practice and Experience, 47(3):379– 390, March 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chambi:2016:SCB

[CLKG16] Samy Chambi, Daniel Lemire, Owen Kaser, and Robert Godin. Short communication: Better bitmap performance with Roaring bitmaps. Software—Practice and Experience, 46(5): 709–719, May 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Charlton:1991:VMN

[CLL91] C. C. Charlton, P. H. Leng, and J. Little. Vertical migration of numerical routines in software and microcode. Software—Practice and Experience, 21(3): 287–297, March 1991. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Charlton:1998:MCR

[CLLT98] C. C. Charlton, P. H. Leng, R. K. Lloyd, and R. Turnbull. Maintaining consistency in a replicated software archive. Software—Practice and Experience, 28(3):285–295, March 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1781;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1781&
PLACEBO=IE.pdf.

Clocksin:1985:ITP

[Clo85] W. F. Clocksin. Implementation techniques for Prolog databases. Software —Practice and Experience, 15(7):669–675, July 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Coenen:2009:IME

[CLP^{+}09] F. Coenen, P. Leng, A. Pagourtzis, W. Rytter, and D. Souliou. Improved methods for extracting frequent itemsets from interim-support trees. Software—Practice and Experience, 39(6):551– 571, April 25, 2009. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cornelius:1984:SSC

[CLR84] B. J. Cornelius, I. R.

Lowman, and D. J. Robson. Steady-state compilers. Software—Practice and Experience, 14(8):705– 709, August 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cerutti:2007:DEA

[CLS^{+}07] S. Cerutti, G. Lamperti, M. Scaroni, M. Zanella, and D. Zanni. A diagnostic environment for automaton networks. Software—Practice and Experience, 37(4): 365–415, April 10, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cheon:2005:MVC

[CLSE05] Yoonsik Cheon, Gary Leavens, Murali Sitaraman, and Stephen Edwards. Model variables: cleanly supporting abstraction in design by contract. Software—Practice and Experience, 35(6): 583–599, May 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Charlton:1990:PMA

[CLW90] C. C. Charlton, P. H.

Leng, and D. M. Wilkinson. Program monitoring and analysis: Software structures and architectural support. Software —Practice and Experience, 20(9):859–867, September 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Corradi:1998:SPH

[CLZ98] Antonio Corradi, Letizia Leonardi, and Franco Zambonelli. Strategies and protocols for highly parallel Linda servers. Software —Practice and Experience, 28(14):1493–1517, December 10, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=5001667;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=5001667&PLACEBO=IE.
pdf.

Cabri:1999:PBF

[CLZ99] Giacomo Cabri, Letizia Leonardi, and Franco Zambonelli. A proxy-based framework to support synchronous cooperation on the Web. Software— Practice and Experience, 29(14):1241–1263, December 10, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
67500433/START;   http:
//www3.interscience.wiley.

com/cgi-bin/fulltext?
ID=67500433&PLACEBO=IE.
pdf.

Cole:1982:TSI

[CM82] A. J. Cole and Ronald Morrison. Triplex: a system for interval arithmetic. Software—Practice and Experience, 12(4):341– 350, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Charlton:1983:TTT

[CM83] C. C. Charlton and K. C.

Mander. Tools and techniques for teaching microprocessor software development. Software—Practice and Experience, 13(10): 909–920, October 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cowling:1985:HSH

[CM85] A. J. Cowling and J. J.

McGregor. HANDIN: a system for helping with the teaching of programming. Software—Practice and Experience, 15(6):611– 622, June 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Clarke:1996:CGC

[CM96] Charles L. A. Clarke and David V. Mason. Compacting garbage collection can be fast and simple. Software—Practice and Experience, 26(2):177–194, February 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61004287.

Chapin:1998:GEG

[CM98a] Steve J. Chapin and Arthur B. ‘Barney’ Maccabe. Guest editorial: Guest editorial on multiprocessor operating systems. Software—Practice and Experience, 28(9):899, July 25, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chapin:1998:GEM

[CM98b] Steve J. Chapin and Arthur B. ‘Barney’ Maccabe. Guest editorial on multiprocessor operating systems. Software—Practice and Experience, 28(9): 899, July 25, 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1817;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1817&
PLACEBO=IE.pdf.

Chivers:2005:E

[CM05] Howard Chivers and Andrew Martin. Editorial. Software—Practice and Experience, 35(9):799– 800, July 25, 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cooper:2008:MDA

[CM08] Jonathan Cooper and Steve McKeever. A modeldriven approach to automatic conversion of physical units. Software—Practice and Experience, 38(4): 337–359, April 10, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:1992:PGA

[CMCH92] Pohua P. Chang, Scott A.

Mahlke, William Y. Chen, and Wen-Mei W. Hwu. Profile-guided automatic inline expansion for C programs. Software—Practice and Experience, 22(5):349– 369, May 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cao:2003:AID

[CMCL03] Jiannong Cao, Xiaoxing Ma, Alvin T. S. Chan, and Jian Lu. Architecting and implementing distributed Web applications using the graph-oriented approach. Software—Practice and Experience, 33(9): 799–820, July 25, 2003.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chung:2005:SS

[CMES05] Yoo C. Chung, Soo-Mook Moon, Kemal Ebcioğlu, and Dan Sahlin. Selective sweeping. Software —Practice and Experience, 35(1):15–26, January 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:1998:SDI

[CMF^{+}98] Kai H. Chang, Liam Murphy, Jonathan D. Fouss, Timothy D. Dollar II, Byong G. Lee, and Yifang Chang. Software development and integration in a computer supported cooperative work environment. Software—Practice and Experience, 28(6):657– 679, May 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=1800;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=1800&
                                        PLACEBO=IE.pdf.

Celesti:2017:EAD

[CMF^{+}17] Antonio Celesti, Davide Mulfari, Maria Fazio, Antonio Puliafito, and Massimo Villari. Evaluating alternative DaaS solutions in private and public OpenStack Clouds. Software—Practice and Experience, 47(9):1185–1200, September 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cooper:1985:EMP

[CMH85] R. E. M. Cooper, Bruce J.

McKenzie, and R. Harries. Extensions to a microcoded Pascal compiler. Software —Practice and Experience, 15(5):519–522, May 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:1991:UPI

[CMH91] Pohua P. Chang, Scott A.

Mahlke, and Wen-Mei W. Hwu. Using profile information to assist classic code optimizations. Software—Practice and Experience, 21(12):1301–1321, December 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2018:DUA

[CmJHL18] Huifeng Chen, Jian min Jiang, Zhong Hong, and Ling Lin. Decomposition of UML activity diagrams. Software—Practice and Experience, 48(1):105– 122, January 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Colin:1975:PSC

[CMM75] A. J. T. Colin, D. R. Mc-Gregor, and D. Mitchell. Paging stab-1 code. Software—Practice and Experience, 5(3):309–312, July 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Carna:2023:SSS

[CMPQ23] Stefano Carnà, Romolo Marotta, Alessandro Pellegrini, and Francesco Quaglia. Strategies and software support for the management of hardware performance counters. Software—Practice and Experience, 53(10):1928–1957, October 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cunto:1992:SMT

[CMR92] Walter Cunto, J. Ian Munro, and Manuel Rey. Selecting the median and two quartiles in a set of numbers. Software—Practice and Experience, 22(6): 439–454, June 1992. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cotroneo:2007:EBC

[CMR07] Domenico Cotroneo, Armando Migliaccio, and Stefano Russo. The Esperanto Broker: a communication platform for nomadic computing systems. Software—Practice and Experience, 37(10):1017–1046, August 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Collberg:2007:ESJ

[CMS07] Christian Collberg, Ginger Myles, and Michael Stepp. An empirical study of Java bytecode programs. Software—Practice and Experience, 37(6):581–641, May 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Castello:2002:ALS

[CMT02] Rodolfo Castello, Rym Mili, and Ioannis G. Tollis. Automatic layout of statecharts. Software—Practice and Experience, 32(1): 25–55, January 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
88013122/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=88013122&PLACEBO=IE.
pdf.

Caporuscio:2017:BDT

[CMT17] Mauro Caporuscio, Raffaela Mirandola, and Catia Trubiani. Building designtime and run-time knowledge for QoS-based component assembly. Software—Practice and Experience, 47(12):1905–1922, December 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Costa-Montenegro:2017:MDS

[CMTCC^{+}17] Enrique Costa-Montenegro, Alexander Tsybanev, Héctor Cerezo-Costas, Francisco Javier González-Castaño, Felipe Gil-Castiñeira, Belén Barragáns-Martínez, and Diego Almuiña-Troncoso. Inmemory distributed software solution to improve the performance of recommender systems. Software —Practice and Experience, 47(6):867–889, June 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cucurull:2010:ESA

[CNAM^{+}10] J. Cucurull, G. Navarro-Arribas, R. Martí, S. Robles, and J. Borrell. An efficient and secure agent code distribution service. Software—Practice and Experience, 40(4):363–386, April 10, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Collins:1983:CTA

[CNG^{+}83] W. Robert Collins, Robert E.

Noonan, Samuel T. Gregory, John C. Knight, and Roy W. Hamm. Comprehensive tools for assembler construction. Software— Practice and Experience, 13(5):447–451, May 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Calheiros:2013:EIE

[CNRB13] Rodrigo N. Calheiros, Marco A. S. Netto, César A. F. De Rose, and Rajkumar Buyya. EMUSIM: an integrated emulation and simulation environment for modeling, evaluation, and validation of performance of Cloud computing applications. Software—Practice and Experience, 43(5): 595–612, May 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cardell-Oliver:1988:BRB

[CO88] Rachel Cardell-Oliver. Book review: Distributed algorithms and protocols, Michel Raynal, Wiley, Chichester, 1988. ISBN 0-471-91754-0. No. of pages:

163. Price: £13.95. Software—Practice and Experience, 18(7):711, July 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cohn:1973:IED

[Coh73] Charles Erwin Cohn. Improving the efficiency of direct-memory-access output operations. Software —Practice and Experience, 3(2):179–183, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cohn:1974:FPB

[Coh74] Charles E. Cohn. FOR-TRAN programming for a batch-transmit CRT terminal. Software—Practice and Experience, 4(4):402– 404, October/December 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cohen:1975:ECP

[Coh75] Jacques Cohen. Experience with a conversational parser generating system. Software—Practice and Experience, 5(2): 169–180, April/June 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cohen:1998:GHS

[Coh98] Jonathan D. Cohen. An n-gram hash and skip algorithm for finding large numbers of keywords in continuous text streams. Software—Practice and Experience, 28(15):1605– 1635, December 25, 1998.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www3.interscience.
wiley.com/cgi-bin/abstract?
ID=10050312;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=10050312&PLACEBO=IE.
pdf.

Colin:1972:LE

[Col72a] Andrew Colin. Letters to the editor. Software— Practice and Experience, 2 (4):404–406, October 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Colin:1972:IS

[Col72b] Andrew J. T. Colin. The implementation of stab-1. Software—Practice and Experience, 2(2):137–142, April/June 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Coleman:1977:SDF

[Col77a] D. Coleman. The systematic design of fileprocessing programs. Software—Practice and Experience, 7(3):371–381, May/ June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Coleman:1977:BRB

[Col77b] M. J. Coleman. Book review: A practical approach to computing, W. Y. Arms, J. E. Baker and R. M. Pengelly, Wiley, London, 1976. No. of pages: 353. Price: £4.95. Software—Practice and Experience, 7(3):431– 432, June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Colquhoun:1977:FAS

[Col77c]  D. G. Colquhoun.  A

fast approximation to the sine function. Software— Practice and Experience, 7 (2):227–229, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Coleman:1979:DOS

[Col79] M. J. Coleman. Demonstrating operating systems algorithms. Software— Practice and Experience, 9(9):789–791, September 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Colijn:1981:NMC

[Col81] A. W. Colijn. A note on the Multics command language. Software—Practice and Experience, 11(7):741– 744, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Collinson:1982:CRU

[Col82] R. P. A. Collinson. The Cambridge ring and Unix. Software—Practice and Experience, 12(6):583–594, June 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cole:1983:NSF

[Col83] A. J. Cole. A note on space filling curves. Software—Practice and Experience, 13(12):1181–1189, December 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cole:1987:MIH

[Col87] Robert Cole. A method for interconnecting heterogeneous computer networks. Software—Practice and Experience, 17(6):387– 397, June 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Colomb:1988:ARE

[Col88] Robert M. Colomb. Assert, retract and external processes in Prolog. Software—Practice and Experience, 18(3):205–220, March 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Comer:1978:MII

[Com78] Douglas Comer. MOUSE4: An improved implementation of the RATFOR preprocessor. Software—Practice and Experience, 8(1): 35–40, January/February 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Comer:1979:MPM

[Com79] Douglas Comer. MAP: A Pascal macro preprocessor for large program development. Software—Practice and Experience, 9(3):203– 209, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Comer:1982:FFS

[Com82]  Douglas Comer.  The

flat file system FFG: a database system consisting of primitives. Software—Practice and Experience, 12(11):1069–1079, November 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Comer:1983:CBL

[Com83] Douglas Comer. The costs and benefits of lazy interactive input in Pascal. Software—Practice and Experience, 13(3):287–293, March 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Conway:1977:BRB

[Con77] D. E. Conway. Book review: Flowcharting, B. B. Bycer, Wiley, New York, 1975. no. of pages: 272. price: £8 cloth, £5.25 paper. Software—Practice and Experience, 7(1):152, January 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Contla:1984:CCS

[Con84] Jose Felipe Contla. Compact coding of syntax tables. Software—Practice and Experience, 14(10): 909–919, October 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Contla:1985:CCS

[Con85] Jose Felipe Contla. Compact coding of syntactically correct source programs. Software—Practice and Experience, 15(7):625– 636, July 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cook:1983:PCP

[Coo83] Steve Cook. Playing cards on the PERQ: An algorithm for overlapping rectangles. Software—Practice and Experience, 13 (11):1043–1053, November 1983. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Cooper:1985:SUQ

[Coo85] R. E. M. Cooper. On squeezing the UNIX quart into a Data General Eclipse pint pot. Software—Practice and Experience, 15(6): 571–581, June 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Coombs:1986:MSD

[Coo86] R. F. Coombs. A management system for dirty data. Software—Practice and Experience, 16(6):549– 558, June 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cooke:1996:ISL

[Coo96] Daniel E. Cooke. An introduction to SequenceL: a language to experiment with constructs for processing nonscalars. Software—Practice and Experience, 26(11):1205–1246, November 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16771.

Cook:2004:PCP

[Coo04] Jonathan J. Cook. P#: a concurrent Prolog for the .NET framework. Software —Practice and Experience, 34(9):815–845, July 25, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cook:2005:HCE

[Coo05] Robert P. Cook. Heuristic compression of an English word list. Software —Practice and Experience, 35(6):577–581, May 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cook:2008:INS

[Coo08] R. P. Cook. An introduction to the NASA Shuttle Ground Operations Simulator (SGOS). Software —Practice and Experience, 38(6):575–587, May ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cornelius:1982:BRB

[Cor82] B. J. Cornelius. Book review: Introductory Algol 68 programming, D. F. Brailsford and A. N. Walker, Ellis Horwood, Chichester. No. of pages: 281. Price: £14.00 (hardback), £5.95 (paperback). Software—Practice and Experience, 12(12):1166–1167, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cordy:1984:CTD

[Cor84] James R. Cordy. Compiletime detection of aliasing in Euclid programs. Software—Practice and Experience, 14(8):755–768, August 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cormack:1988:MKC

[Cor88a] Gordon V. Cormack. A micro-kernel for concurrency in C. Software— Practice and Experience, 18(5):485–491, May 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cornelius:1988:PLM

[Cor88b] B. J. Cornelius. Problems with the language Modula-2. Software—Practice and Experience, 18(6): 529–543, June 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cornwell:1999:BRB

[Cor99a] Pete Cornwell. Book review: UML Toolkit by Hans-Erik Eriksson and Magnus Penker, Wiley, 1997. ISBN: 0-471-19161-2, £39.95/$32.08. Software —Practice and Experience, 29(2):195, February 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cornwell:1999:UTH

[Cor99b]  Pete Cornwell.  ‘UML

Toolkit’ by Hans-Erik Eriksson and Magnus Penker, Wiley, 1997. ISBN: 0-471-19161-2, £39.95/ $32.08. Software—Practice and Experience, 29(2): 195, February 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55000308;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55000308&PLACEBO=IE.
pdf.

Cormack:2008:CPS

[Cor08] Gordon V. Cormack. Call for papers: Special issue on ‘Spam Detection’ Software: Practice and Experience (SPE). Software—Practice and Experience, 38(3): 335–336, March ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Counihan:1984:BRBa

[Cou84a] Martin Counihan. Book review: Beginning Basic, 2nd Edn., P. E. Gosling, Macmillan Press Ltd., 1982. Price: £2.95. Software—Practice and Experience, 14(4):400, April 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Counihan:1984:BRBb

[Cou84b] Martin Counihan. Book review: Pocket Guide to Basic, Roger Hunt, Pitman Programming Pocket Guides, 1982. Price: £2.25. Software—Practice and Experience, 14(4):400, April 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Counihan:1985:BRBa

[Cou85a] Martin Counihan. Book review: A structured approach to Fortran 77 programming, T. M. R. Ellis, Addison-Wesley International Computer Science Series, 1982. no. of pages:

350. price: £8.95. Software—Practice and Experience, 15(3):319, March 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Counihan:1985:BRBb

[Cou85b] Martin Counihan. Book review: Structured Fortran 77 for engineers and scientists, D. M. Etter, Benjamin/Cummings Publishing Company, 1983. No. of pages: 357. Price: £9.95. Software—Practice and Experience, 15(3): 319, March 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Courtney:1992:AEI

[Cou92] Phil Courtney. Automation education industry growing. Software Magazine, 12(1):58–??, January 1992. CODEN SMWMEQ. ISSN 0897-8085.

Cowie:1987:DAT

[Cow87] J. R. Cowie. A direct access technique for sequential files with variable length records. Software —Practice and Experience, 17(10):719–728, October 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cox:1976:FMA

[Cox76]  Lyle A. Cox, Jr.  A

fast method for approximate calculation of the Kolmogorov–Smironov statistic. Software—Practice and Experience, 6(2):281– 283, April 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cox:1985:TLM

[Cox85] M. G. Cox. Topic libraries for mathematical computation. Software—Practice and Experience, 15(4):397– 411, April 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cunningham:1976:LIS

[CP76] R. J. Cunningham and C. G. Pugh. A languageindependent system to aid the development of structured programs. Software—Practice and Experience, 6(4):487–503, October/December 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Coleman:1996:ABG

[CP96] Michael K. Coleman and D. Stott Parker. Aestheticsbased graph layout for human consumption. Software—Practice and Experience, 26(12):1415–1438, December 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16781.

Chimaris:2007:IGC

[CP07] Avraam N. Chimaris and George A. Papadopoulos. Implementing a generic component-based framework for telecontrol applications. Software— Practice and Experience, 37(10):1087–1132, August 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chahal:2022:EHB

[CP22] Prabhjot Kaur Chahal and Shreelekha Pandey. An efficient Hadoop-based brain tumor detection framework using big data analytic. Software—Practice and Experience, 52(3):805– 818, March 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chae:2010:ALB

[CPCL10] Heung Seok Chae, Jae Geol Park, Jian Feng Cui, and Joon Sang Lee. An adaptive load balancing management technique for RFID middleware systems. Software—Practice and Experience, 40(6):485– 506, May 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Carreton:2013:SAM

[CPD13] Andoni Lombide Carreton, Kevin Pinte, and Wolfgang De Meuter. Software abstractions for mobile RFID-enabled applications. Software—Practice and Experience, 43(10): 1219–1239, October 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Clint:1983:IHS

[CPHS83] M. Clint, R. Perrott, C. Holt, and A. Stewart. The influence of hardware and software considerations on the design of synchronous parallel algorithms. Software—Practice and Experience, 13(10): 961–974, October 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chirivella-Perez:2020:SME

[CPMAH^{+}20] Enrique Chirivella-Perez, Ricardo Marco-Alaez, Alba Hita, Ana Serrano, Jose M. Alcaraz Calero, Qi Wang, Pedro M. Neves, Giacomo Bernini, Konstantinos Koutsopoulos, Manuel Gil Pérez, Gregorio Martínez Pérez, Maria João Barros, and Anastasius Gavras. SELFNET 5G mobile edge computing infrastructure: Design and prototyping. Software—Practice and Experience, 50(5):741– 756, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Culpepper:2012:RBC

[CPP12] J. Shane Culpepper, Matthias Petri, and Simon J. Puglisi. Revisiting bounded context block-sorting transformations. Software—Practice and Experience, 42(8): 1037–1054, August 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Clowes:1973:ANI

[CPW73] J. S. Clowes, E. S. Page, and L. B. Wilson. Another note on internal merging. Software—Practice and Experience, 3(2):185, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Coleman:1974:MPS

[CPW74] S. S. Coleman, P. C. Poole, and W. M. Waite. The mobile programming system, Janus. Software—Practice and Experience, 4(1): 5–23, January/March 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Campo:2002:DOO

[CPZ02] Marcelo Campo, Andrés Díaz Pace, and Mario Zito. Developing object-oriented enterprise quality frameworks using proto-frameworks. Software—Practice and Experience, 32(8):837–843, July 10, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
94519517/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=94519517&PLACEBO=IE.
pdf.

Cabodi:1998:MOF

[CQC98] G. Cabodi, S. Quer, and P. Camurati. Memory optimization in function and set manipulation with BDDs. Software—Practice and Experience, 28(1):99– 120, January 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61001667;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=61001667&PLACEBO=IE.
pdf.

Chen:2013:ECP

[CQH^{+}13] Gengbiao Chen, Zhengwei Qi, Shiqiu Huang, Kangqi Ni, Yudi Zheng, Walter Binder, and Haibing Guan. Extended conference papers: A refined decompiler to generate C code with high readability. Software—Practice and Experience, 43(11):1337–1358, November 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chondros:2018:DIS

[CR18] Nikos Chondros and Mema Roussopoulos. Developing IntegrityCatalog, a software system for managing integrity-related metadata in digital repositories. Software—Practice and Experience, 48(1):45– 64, January 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Craddock:1976:DFB

[Cra76] J. M. Craddock. Designing a FORTRAN based system for meteorological statistics. Software—Practice and Experience, 6(3):301– 319, July/September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Craddock:1977:PSP

[Cra77] J. M. Craddock. Planning the software for a personal research contract. Software —Practice and Experience, 7(5):645–653, September/ October 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Calheiros:2011:CTM

[CRB^{+}11] Rodrigo N. Calheiros, Rajiv Ranjan, Anton Beloglazov, César A. F. De Rose, and Rajkumar Buyya. CloudSim: a toolkit for modeling and simulation of cloud computing environments and evaluation of resource provisioning algorithms. Software—Practice and Experience, 41(1):23– 50, January 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cappellari:2018:ODS

[CRC18] Paolo Cappellari, Mark Roantree, and Soon Ae Chun. Optimizing data stream processing for largescale applications. Software—Practice and Experience, 48(9):1607–1641, September 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Criado:2015:TAC

[CRGIP15] Javier Criado, Diego Rodríguez-Gracia, Luis Iribarne, and Nicolás Padilla. Toward the adaptation of component-based architectures by model transformation: behind smart user interfaces. Software—Practice and Experience, 45 (12):1677–1718, December 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Crowe:1987:DCU

[Cro87] M. K. Crowe. Dynamic compilation in the Unix environment. Software— Practice and Experience, 17(7):455–467, July 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Crowe:1991:NNA

[Cro91] David R. Crowe. NovAtel’s novel approach to CPU usage measurement. Software —Practice and Experience, 21(5):465–477, May 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:1994:MIA

[CRR94] Ta Chen, I. V. Ramakrishnan, and R. Ramesh. Multistage indexing algorithms for speeding Prolog execution. Software —Practice and Experience, 24(12):1097–1119, December 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Preliminary version: ICLP 1992, pp. 639– 653.

Cornelius:1980:MPP

[CRT80] B. J. Cornelius, D. J. Robson, and M. I. Thomas. Modification of the Pascal-P compiler for a singleaccumulator one-address minicomputer. Software —Practice and Experience, 10(3):241–246, March 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Comer:1982:HBS

[CS82] Douglas Comer and Vincent Yun Shen. Hashbucket search: a fast technique for searching an English spelling dictionary. Software—Practice and Experience, 12(7):669– 682, July 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Callison:1991:BRT

[CS91a] H. Rebecca Callison and Alan C. Shaw. Building a real-time kernel: First steps in validating a pure process/ADT model. Software—Practice and Experience, 21(4):337–354, April 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cheng:1991:DID

[CS91b] Hsiao-Chung Cheng and Jang-Ping Sheu. Design and implementation of a distributed file system. Software—Practice and Experience, 21(7):657– 675, July 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:1997:IIR

[CS97] Zhanliang Chen and Richard D. Schlichting. Interconnecting interactive and remote parallel components using Schooner. Software—Practice and Experience, 27(11):1349–1367, November 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7274;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7274&
PLACEBO=IE.pdf.

Chanchio:2002:DCR

[CS02] Kasidit Chanchio and Xian-He Sun. Data collection and restoration for heterogeneous process migration. Software—Practice and Experience, 32(9):845–871, July 25, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
94519537/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=94519537&PLACEBO=IE.
pdf.

Carr:2003:EES

[CS03] Steve Carr and Philip Sweany. An experimental evaluation of scalar replacement on scientific benchmarks. Software —Practice and Experience, 33(15):1419–1445, December 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Csallner:2004:JAR

[CS04] Christoph Csallner and Yannis Smaragdakis. JCrasher: an automatic robustness tester for Java. Software—Practice and Experience, 34(11):1025–1050, September 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Colnet:2015:EAM

[CS15] Dominique Colnet and Benoît Sonntag. Exploiting array manipulation habits to optimize garbage collection and type flow analysis. Software—Practice and Experience, 45(12):1639– 1657, December 2015. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Constantinou:2017:IEP

[CS17] Eleni Constantinou and Ioannis Stamelos. Identifying evolution patterns: a metrics-based approach for external library reuse. Software—Practice and Experience, 47(7):1027–1039, July 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chainbi:2018:MCS

[CS18] W. Chainbi and E. Sassi. A multiswarm for composite SaaS placement optimization based on PSO. Software—Practice and Experience, 48(10):1847–1864, October 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Correia:2022:IMF

[CS22] José Correia and António Rito Silva. Identification of monolith functionality refactorings for microservices migration. Software— Practice and Experience, 52(12):2664–2683, December 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cai:2023:ARC

[CSD23] Cheng-Hao Cai, Jing Sun, and Gillian Dobbie. Automatic refactoring of conditions and substitutions for B state transition models. Software—Practice and Experience, 53(11):2241– 2270, November 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cowan:1993:AIC

[CSIL93] D. D. Cowan, T. M.

Stepien, R. Ierusalimschy, and Carlos J. P. Lucena. Application integration: Constructing composite applications from interactive components. Software—Practice and Experience, 23(3):255–275, March 1993. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2016:NUN

[CSM^{+}16] Renhai Chen, Zhaoyan Shen, Chenlin Ma, Zili Shao, and Yong Guan. NVMRA: utilizing NVM to improve the random write operations for NANDflash-based mobile devices. Software—Practice and Experience, 46 (9):1263–1284, September 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Caymes-Scutari:2012:MTK

[CSMML12] P. Caymes-Scutari, A. Morajko, T. Margalef, and E. Luque. A methodology for transparent knowledge specification in a dynamic tuning environment. Software—Practice and Experience, 42(3):281–302, March 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chung:1993:PCE

[CSR93] Kehsiung Chung, Janche Sang, and Vernon Rego. A performance comparison of event calendar algorithms: an empirical approach. Software—Practice and Experience, 23(10): 1107–1138, October 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cook:2015:ERE

[CSS15] Charles T. Cook, Yu-Shan Sun, and Murali Sitaraman. Experience report: evolution of a webintegrated software development and verification environment. Software— Practice and Experience, 45(6):857–872, June 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Colin:1975:TIS

[CST75] Andrew J. T. Colin, K. Shorey, and W. Teasdale. The translation and interpretation of STAB-12. Software—Practice and Experience, 5(2):123–138, April/June 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chiang:2019:MAK

[CSTL19] Mei-Ling Chiang, Wei-Lun Su, Shu-Wei Tu, and Zhen-Wei Lin. Memoryaware kernel mechanism and policies for improving internode load balancing on NUMA systems. Software—Practice and Experience, 49(10):1485–1508, October 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chandrasekaran:2023:SIN

[CSZO23] Sunita Chandrasekaran, Min Si, Jidong Zhai, and Lena Oden. Special issue on new trends in high-performance computing: Software systems and applications. Software— Practice and Experience, 53(1):3–5, January 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cowell:1990:TAD

[CT90] Wayne R. Cowell and Christopher P. Thompson. Tools to aid in discovering parallelism and localizing arithmetic in Fortran programs. Software—Practice and Experience, 20(1):25– 47, January 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:1992:CDC

[CT92] Kuang-Yao Chang and Gregory Y. Tang. CCIGS: a data compression system for Chinese fonts and binary images using classification techniques. Software—Practice and Experience, 22(12):1027–1047, December 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:2007:KSZ

[CTLL07] Da-Wei Chang, Chuan-Ming Tsai, Wei-Kou Li, and Tzu-Rung Lee. Kernel support for zero-loss Internet service restart. Software—Practice and Experience, 37(8):833–855, July 10, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cukic:2016:CBT

[Čuk16] Ivan Čukić. A continuationbased task programming model for C++: design of the Causeway library. Software—Practice and Experience, 46(12):1617–1656, December 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cuming:1971:MOS

[Cum71]  C. B. Cuming.  Mon-

itoring the operation of system software. Software—Practice and Experience, 1(4):383–389, October/December 1971. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Clint:1984:UGV

[CV84] M. Clint and C. Vicent.

The use of ghost variables and virtual programming in the documentation and verification of programs. Software—Practice and Experience, 14(8):711– 737, August 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Costello:1998:RBT

[CV98] Adam M. Costello and George Varghese. Redesigning the BSD timer facilities. Software—Practice and Experience, 28(8): 883–896, July 10, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www3.interscience.
wiley.com/cgi-bin/abstract?
ID=1813;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1813&
PLACEBO=IE.pdf.

Caromel:2003:SFR

[CV03] Denis Caromel and Julien Vayssière. A security framework for reflective Java applications. Software —Practice and Experience, 33(9):821–846, July 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cunei:2008:EFT

[CV08] A. Cunei and J. Vitek. An efficient and flexible toolkit for composing customized method dispatchers. Software—Practice and Experience, 38(1):33–73, January ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Candal-Ventureira:2023:ERN

[CVGCGCFF23] David Candal-Ventureira, Francisco Javier González-Castaño, Felipe Gil-Castiñeira, and Pablo Fondo-Ferreiro. Is the edge really necessary for drone computing offloading? An experimental assessment in carriergrade 5G operator networks. Software—Practice and Experience, 53(3):579– 599, March 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chiueh:1997:DIS

[CVV97] Tzi-Cker Chiueh, Chitra Venkatramani, and Michael Vernick. Design and implementation of the Stony Brook Video Server. Software—Practice and Experience, 27(2):139– 154, February 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7283;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7283&
PLACEBO=IE.pdf.

Cox:1980:HSU

[CW80] G. B. Cox and B. C.

Walsh. A HELP system for the user community. Software—Practice and Experience, 10(3):219– 229, March 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Comer:1982:AYM

[CW82a] Douglas Comer and Keith Williamson. An alternative to Young’s module facility for Pascal. Software —Practice and Experience, 12(10):907–913, October 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Covington:1982:TDS

[CW82b] W. O. Covington, Jr. and Richard L. Wexelblat. A test data synthesizer for telephone company accounting office processing of call data. Software— Practice and Experience, 12(6):505–516, June 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cowderoy:1982:TBC

[CW82c] Richard I. Cowderoy and Peter J. L. Wallis. The transfer of a BCPL compiler to the Z80 microcomputer. Software—Practice and Experience, 12(3):235– 239, March 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:1991:LOP

[CW91] Chin-Chen Chang and Tzong-Chen Wu. A letteroriented perfect hashing scheme based upon Sparse table compression. Software—Practice and Experience, 21(1):35–49, January 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cohen:1992:STM

[CW92] Jacques Cohen and Aline Weitzman. Software tools for micro-analysis of programs. Software—Practice and Experience, 22(9): 777–808, September 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chao:1994:ITD

[CW94] Daniel Y. Chao and David T. Wang. An interactive tool for design, simulation, verification, and synthesis of protocols. Software—Practice and Experience, 24(8):747–783, August 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cooper:1997:AIC

[CW97] Tim Cooper and Michael Wise. Achieving incremental compilation through fine-grained builds. Software—Practice and Experience, 27(5):497–517, May 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7305;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7305&
PLACEBO=IE.pdf.

Cook:2001:IPL

[CW01] Phil Cook and Jim Welsh. Incremental parsing in language-based editors: user needs and how to meet them. Software —Practice and Experience, 31(15):1461–1486, December 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
85515685/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=85515685&PLACEBO=IE.
pdf.

Cutcutache:2008:FFB

[CW08] Ioana Cutcutache and Weng-Fai Wong. Fast, frequency-based, integrated register allocation and instruction scheduling. Software—Practice and Experience, 38(11):1105–1126, September ??, 2008. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cao:2021:ERE

[CWC^{+}21] Kun Cao, Tongquan Wei, Mingsong Chen, Keqin Li, Jian Weng, and Wuzheng Tan. Exploring reliable edge-cloud computing for service latency optimization in sustainable cyberphysical systems. Software—Practice and Experience, 51(11):2225–2237, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2008:CRA

[CWD08] Zhongqiang Chen, Peter Wei, and Alex Delis. Catching Remote Administration Trojans ( RATs ). Software—Practice and Experience, 38(7):667–703, June ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2007:NAE

[CWS07] Ningtao Chen, Nengchao Wang, and Baochang Shi. A new algorithm for encoding and decoding the Hilbert order. Software— Practice and Experience, 37(8):897–908, July 10, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2017:ESS

[CWZ17] Dan Chen, Lizhe Wang, and Suiping Zhou. Editorial: Software systems for data-centric smart city applications. Software— Practice and Experience, 47(8):1043–1044, August 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chiao:2001:ETS

[CY01a] Hsin-Ta Chiao and Shyan-Ming Yuan. An enhanced thread synchronization mechanism for Java. Software—Practice and Experience, 31(7):667– 695, June 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
78003134/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=78003134&PLACEBO=IE.

pdf. See correction [CY01b].

Chiao:2001:RIM

[CY01b] Hsin-Ta Chiao and Shyan-Ming Yuan. The real inventor of the monitor concept: a short note to correct the SP&E paper entitled “An enhanced thread synchronization mechanism for Java”. Software—Practice and Experience, 31(14):1393, November 25, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
85515675/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=85515675&PLACEBO=IE.

pdf. See [CY01a].

Chen:2022:RAG

[CYJ^{+}22] Hanhua Chen, Jie Yuan, Hai Jin, Yonghui Wang, Sijie Wu, and Zhihao Jiang. RGraph: Asynchronous graph processing based on asymmetry of remote direct memory access. Software —Practice and Experience, 52(2):374–393, February 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2015:EMF

[CYW^{+}15] Tseng-Yi Chen, Hsiu-Lien Yeh, Hsin-Wen Wei, Mei ju Sun, Tsan sheng Hsu, and Wei-Kuan Shih. An effective monitoring framework and user interface design. Software—Practice and Experience, 45(4):549– 570, April 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2004:EDC

[CZ04] Songqing Chen and Xiaodong Zhang. Enforcing direct communications between clients and Web servers to improve proxy performance and security. Software—Practice and Experience, 34(14): 1381–1395, November 25, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cunningham:1983:STF

[CZA83] R. James Cunningham and Silvana Zappacosta-Amboldi. Software tools for first-order logic. Software—Practice and Experience, 13(11):1019–1025, November 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2021:SGL

[CZL21] Fangwei Chen, Li Zhang, and Xiaoli Lian. A systematic gray literature review: the technologies and concerns of microservice application programming interfaces. Software—Practice and Experience, 51(7): 1483–1508, July 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2024:CAE

[CZL24] Fangwei Chen, Li Zhang, and Xiaoli Lian. CDTC: Automatically establishing the trace links between class diagrams in design phase and source code. Software—Practice and Experience, 54(2):281– 307, February 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DAgapeyeff:1973:EGE

[D’A73] A. D’Agapeyeff. Editorials: Guest editorial. Software—Practice and Experience, 3(1):3–4, January/ March 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dong:2006:AAD

[DAC06] Jing Dong, Paulo S. C.

Alencar, and Donald D. Cowan. Automating the analysis of design component contracts. Software —Practice and Experience, 36(1):27–71, January 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dewangan:2021:ERC

[DAC^{+}21] Bhupesh Kumar Dewangan, Amit Agarwal, Tanupriya Choudhury, Ashutosh Pasricha, and Suresh Chandra Satapathy. Extensive review of cloud resource management techniques in Industry 4.0: Issue and challenges. Software—Practice and Experience, 51(12):2373–2392, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deAlmeida:2018:ACR

[dAHCdAC18] Dayse Silveira de Almeida, Carmem Satie Hara, Ricardo Rodrigues Ciferri, and Cristina Dutra de Aguiar Ciferri. An asynchronous collaborative reconciliation model based on data provenance. Software—Practice and Experience, 48(1):197– 232, January 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dinh:2015:DCF

[DAJ^{+}15] Minh Ngoc Dinh, David Abramson, Chao Jin, Andrew Gontarek, Bob Moench, and Luiz DeRose. A data-centric framework for debugging highly parallel applications. Software —Practice and Experience, 45(4):501–526, April 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deAlbuquerque:2011:SMB

[dAKdGJ11] João Porto de Albuquerque, Heiko Krumm, Paulo Lício de Geus, and René Jeruschkat. Scalable model-based configuration management of security services in complex enterprise networks. Software—Practice and Experience, 41(3):307–338, March 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dannenberg:1982:LER

[Dan82] Roger B. Dannenberg. Letters to the Editor: Run time data structures. Software—Practice and Experience, 12(4):394, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dannenberg:1990:SEU

[Dan90] Roger B. Dannenberg. A structure for efficient update, incremental redisplay and undo in graphical editors. Software—Practice and Experience, 20(2):109– 132, February 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sanzo:2021:ARC

[DAP21] Pierangelo Di Sanzo, Dimiter R.

Avresky, and Alessandro Pellegrini. Autonomic rejuvenation of cloud applications as a countermeasure to software anomalies. Software—Practice and Experience, 51(1):46– 71, January 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

delAmo:2010:SMA

[dAPMV10] Ignacio J. G. del Amo, David A. Pelta, Antonio D. Masegosa, and Jose L. Verdegay. A software modeling approach for the design and analysis of cooperative optimization systems. Software—Practice and Experience, 40(9):811– 823, August 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Darmont:2000:DCD

[Dar00] Jérôme Darmont. DESP-C++: a discrete-event simulation package for C++. Software—Practice and Experience, 30(1):37– 60, January 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
69503096/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=69503096&PLACEBO=IE.
pdf.

Davies:1974:BRB

[Dav74] N. R. Davies. Book review: Computer simulation application, Julian Reitman, Wiley-Interscience, New York, 1971. No. of pages:

422. Price: £8.50. Software—Practice and Experience, 4(4):406–407, October/December 1974. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Davies:1978:BRB

[Dav78]  N. Ri. Davies.  Book

review: Computer methods in operational research, Arne Thesen, Academic Press, New York, 1978. No. of pages 268. Price: $19.50. Software—Practice and Experience, 8(6):756, November 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Davies:1982:SST

[Dav82] D. Julian M. Davies. String searching in text editors. Software—Practice and Experience, 12(8):709– 717, August 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dawes:1977:SNI

[Daw77] N. W. Dawes. A Simple Network Interacting Program’s Executive (SNIPE). Software—Practice and Experience, 7(3):341–345, May/June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Day:1983:TMM

[Day83] Robert A. Day. Typesetting mathematics on multiaccess systems. Software —Practice and Experience, 13(2):131–138, February 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Daynes:2000:IAF

[Day00] L. Daynès. Implementation of automated finegranularity locking in a persistent programming language. Software—Practice and Experience, 30(4):325–361, April 10, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        71004139/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=71004139&PLACEBO=IE.
                                        pdf.

Davis:1983:ZZD

[DB83] R. H. Davis and A. J. Bathgate. A zilog Z8000 disassembler. Software—Practice and Experience, 13 (11):1055–1073, November 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Davies:1986:APM

[DB86] G. Davies and S. Bowsher. Algorithms for pattern matching. Software —Practice and Experience, 16(6):575–601, June 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deBruin:2000:BBC

[dB00] Hans de Bruin. BCOOPL: Basic concurrent objectoriented programming language. Software—Practice and Experience, 30(8):849– 894, July 10, 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72502205/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72502205&PLACEBO=IE.
pdf.

Dangelmayr:2009:AOC

[DB09] C. Dangelmayr and W. Blochinger.

Aspect-oriented component assembly — a case study in parallel software design. Software—Practice and Experience, 39(9):807– 832, June 25, 2009. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Daghaghzadeh:2021:MDC

[DB21a] Mohammad Daghaghzadeh and Seyed Morteza Babamir. A model driven and clustering method for service identification directed by metrics. Software—Practice and Experience, 51(2): 449–484, February 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Delisle:2021:ACF

[DB21b] Thierry Delisle and Peter A. Buhr. Advanced control-flow and concurrency in C∀. Software—Practice and Experience, 51(5):1005–1042, May 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dunkel:2004:CJP

[DBH04] Jürgen Dunkel, Ralf Bruns, and Andreas Holitschke. Comparison of JavaServer Pages and XSLT: a software engineering perspective. Software—Practice and Experience, 34(1):1– 13, January 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Denz:2018:SMB

[DBO^{+}18] Robert Denz, Scott Brookes, Martin Osterloh, Stephen Kuhn, and Stephen Taylor. Symmetric multiprocessing from boot to virtualization. Software—Practice and Experience, 48 (3):681–718, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2505.

Daun:2023:ISA

[DBOT23] Marian Daun, Jennifer Brings, Patricia Aluko Obe, and Bastian Tenbergen. An industry survey on approaches, success factors, and barriers for technology transfer in software engineering. Software—Practice and Experience, 53(7): 1496–1524, July 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Denkers:2023:TCI

[DBvG^{+}23] Jasper Denkers, Marvin Brunner, Louis van Gool, Jurgen J. Vinju, Andy Zaidman, and Eelco Visser. Taming complexity of industrial printing systems using a constraint-based DSL: an industrial experience report. Software— Practice and Experience, 53(10):2026–2064, October 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Depradine:2003:CDC

[DC03] Colin Depradine and Pranay Chaudhuri. P^{3}: a code and design conventions preprocessor for Java. Software—Practice and Experience, 33(1):61–76, January 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
100019567/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?

                                        ID=100019567&PLACEBO=
                                        IE.pdf.

Di:2015:ECP

[DC15] Sheng Di and Franck Cappello. Extended conference papers: GloudSim: Google trace based cloud simulator with virtual machines. Software—Practice and Experience, 45 (11):1571–1590, November 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dawson:1982:HIC

[DCA82] A. F. Dawson, M. J. Coombs, and J. L. Alty. How to improve computer advisory services. Software —Practice and Experience, 12(9):857–878 (or 857– 877??), September 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Doyle:2004:DIM

[DCA04] Patrick Doyle, Carlos Cavanna, and Tarek S. Abdelrahman. The design and implementation of a modular and extensible Java Virtual Machine. Software—Practice and Experience, 34(3):287–313, March 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Castro:2020:ASA

[dCCCdAC20] João Pedro de Carvalho Castro, Anderson Chaves Carniel, and Cristina Dutra de Aguiar Ciferri. Analyzing spatial analytics systems based on Hadoop and Spark: a user perspective. Software—Practice and Experience, 50 (12):2121–2144, December 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Martino:2023:TSA

[DCEG23] Beniamino Di Martino, Luigi Colucci Cante, Antonio Esposito, and Mariangela Graziano. A tool for the semantic annotation, validation and optimization of business process models. Software—Practice and Experience, 53(5): 1174–1195, May 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deCaso:2013:IPV

[dCGG13] Guido de Caso, Diego Garbervetsky, and Daniel Gorín. Integrated program verification tools in education. Software—Practice and Experience, 43(4):403– 418, April 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Denninnart:2023:ESC

[DCS23] Chavit Denninnart, Thanawat Chanikaphon, and Mohsen Amini Salehi. Efficiency in the serverless cloud paradigm: a survey on the reusing and approximation aspects. Software—Practice and Experience, 53(10): 1853–1886, October 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deCarlini:1988:SAC

[dCV88] Ugo de Carlini and Umberto Villano. A simple algorithm for clock synchronization in transputer networks. Software—Practice and Experience, 18(4):331– 347, April 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Darragh:1993:BCR

[DCW93] John. J. Darragh, John G. Cleary, and Ian H. Witten. Bonsai: a compact representation of trees. Software—Practice and Experience, 23(3):277–291, March 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Droms:1990:PMX

[DD90] Ralph Droms and Wayne R. Dyksen. Performance measurements of the X Window System communication protocol. Software— Practice and Experience, 20(S2):S119–S136, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Desnoyers:2010:SFR

[DD10] Mathieu Desnoyers and Michel R. Dagenais. Synchronization for fast and reentrant operating system kernel tracing. Software—Practice and Experience, 40(12):1053–1072, November 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Daoud:2018:RDS

[DD18] Houssem Daoud and Michel R.

Dagenais. Recovering disk storage metrics from low-level trace events. Software—Practice and Experience, 48(5):1019– 1041, May 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2566.

Daoud:2021:PAD

[DD21] Houssem Daoud and Michel R.

Dagenais. Performance analysis of distributed storage clusters based on kernel and userspace traces. Software—Practice and Experience, 51(1):5–24, January 2021. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Delahaye:2015:SSE

[DdB15] Mickaël Delahaye and Lydie du Bousquet. Selecting a software engineering tool: lessons learnt from mutation analysis. Software —Practice and Experience, 45(7):875–891, July 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dautov:2018:CIV

[DDB^{+}18a] Rustem Dautov, Salvatore Distefano, Dario Bruneo, Francesco Longo, Giovanni Merlino, Antonio Puliafito, and Rajkumar Buyya. Cover image volume 48, issue 8. Software—Practice and Experience, 48(8):i, August 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2609.

Dautov:2018:MIS

[DDB^{+}18b] Rustem Dautov, Salvatore Distefano, Dario Bruneo, Francesco Longo, Giovanni Merlino, Antonio Puliafito, and Rajkumar Buyya. Metropolitan intelligent surveillance systems for urban areas by harnessing IoT and edge computing paradigms.

Software—Practice and Experience, 48(8):1475– 1492, August 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2586.

Desfossez:2016:RLD

[DDD16] Julien Desfossez, Mathieu Desnoyers, and Michel R. Dagenais. Runtime latency detection and analysis. Software—Practice and Experience, 46(10):1397– 1409, October 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DeBeukelaer:2017:ECP

[DDDF17] Herman De Beukelaer, Guy F. Davenport, Geert De Meyer, and Veerle Fack. Extended conference paper: JAMES: an objectoriented Java framework for discrete optimization using local search metaheuristics. Software—Practice and Experience, 47(6): 921–938, June 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DElia:2016:ECP

[DDF16] Daniele Cono D’Elia, Camil Demetrescu, and Irene Finocchi. Extended conference paper: Mining hot calling contexts in small space. Software—Practice and Experience, 46(8): 1131–1152, August 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DSouza:2018:EGW

[DDGP18] Chris D’Souza, Vincenzo Deufemia, Athula Ginige, and Giuseppe Polese. Enabling the generation of web applications from mockups. Software—Practice and Experience, 48 (4):945–973, April 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2559.

DeCremer:2020:SES

[DDMD20] Pieter De Cremer, Nathan Desmet, Matias Madou, and Bjorn De Sutter. Sensei: Enforcing secure coding guidelines in the integrated development environment. Software—Practice and Experience, 50 (9):1682–1718, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DalPalu:2007:CSD

[DDP07] Alessandro Dal Palù, Agostino Dovier, and Enrico Pontelli. A constraint solver for discrete lattices, its parallelization, and application to protein structure prediction. Software —Practice and Experience, 37(13):1405–1449, November 10, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DiBattista:2002:DDS

[DDPP02] Giuseppe Di Battista, Walter Didimo, Maurizio Patrignani, and Maurizio Pizzonia. Drawing database schemas. Software—Practice and Experience, 32(11):1065–1098, September 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
97519258/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=97519258&PLACEBO=IE.
pdf.

Detlefs:1994:MAC

[DDZ94] David Detlefs, Al Dosser, and Benjamin G. Zorn. Memory allocation costs in large C and C++ programs. Software—Practice and Experience, 24(6):527– 542, June 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deVSmit:1982:CTS

[de 82] G. de V.Smit. A comparison of three string matching algorithms. Software —Practice and Experience, 12(1):57–66, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DeBosschere:1996:OPP

[De 96] Koen De Bosschere. An operator precedence parser for standard Prolog text. Software—Practice and Experience, 26(7):763–779, July 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16820.

DiMartino:2016:ECP

[DE16] Beniamino Di Martino and Antonio Esposito. Extended conference papers: A rule-based procedure for automatic recognition of design patterns in UML diagrams. Software—Practice and Experience, 46(7): 983–1007, July 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Deasington:1986:BRB

[Dea86] R. J. Deasington. Book review: Operating Systems— Structures and Mechanisms, Philippe A. Janson, Academic Press Inc., 1985. ISBN 0-12-380230-X. No. of pages: 267. Price: £19.50. Software—Practice and Experience, 16 (11):1055–1056, November 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Debray:1988:PPP

[Deb88] Saumya K. Debray. Profiling Prolog programs. Software—Practice and Experience, 18(9):821–839, September 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Debray:1993:QJS

[Deb93] Saumya K. Debray. QD-Janus: a sequential implementation of Janus in Prolog. Software—Practice and Experience, 23 (12):1337–1360, December 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dellar:1982:FSN

[Del82] Carl N. R. Dellar. A file server for a network of low cost personal microcomputers. Software—Practice and Experience, 12 (11):1051–1068, November 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Deorowicz:2000:IBW

[Deo00] Sebastian Deorowicz. Improvements to Burrows– Wheeler compression algorithm. Software—Practice and Experience, 30 (13):1465–1483, November 10, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
73001202/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=73001202&PLACEBO=IE.
pdf.

Deorowicz:2002:SSA

[Deo02] Sebastian Deorowicz. Second step algorithms in the Burrows–Wheeler compression algorithm. Software—Practice and Experience, 32(2):99–111, February 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
88013462/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=88013462&PLACEBO=IE.
pdf.

Deorowicz:2010:SLC

[Deo10] Sebastian Deorowicz. Solving longest common subsequence and related problems on graphical processing units. Software—Practice and Experience, 40(8): 673–700, July 2010. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Desjardins:1974:DDS

[Des74] Pierre Desjardins. Dynamic data structure mapping. Software—Practice and Experience, 4(2):155– 162, April/June 1974. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Desmond:1992:MRI

[Des92] John Desmond. Mapping real to the ideal model. Software Magazine, 12(1): 46–??, January 1992. CO-DEN SMWMEQ. ISSN 0897-8085.

Deubler:1999:VSS

[Deu99] Hanns-Helmuth Deubler.

A viable system structure for large-scale software systems. Software —Practice and Experience, 29(12):1025–1047, October 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
66000406/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=66000406&PLACEBO=IE.
pdf.

Dewey:1984:QTG

[Dew84] M. E. Dewey. QS: a tool for general data input. Software—Practice and Experience, 14(12):1113–1121, December 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dewey:1986:TSP

[Dew86] Michael E. Dewey. TBFLP: a simple preprocessor for tables. Software—Practice and Experience, 16(8):731– 738, August 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dewhurst:1987:FST

[Dew87] Stephen C. Dewhurst. Flexible symbol table structures for compiling C++. Software—Practice and Experience, 17(8):503– 512, August 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dewan:1991:IMS

[Dew91] Prasun Dewan. An inheritance model for supporting flexible displays of data structures. Software —Practice and Experience, 21(7):719–738, July 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dewan:1993:DIM

[Dew93] Prasun Dewan. Designing and implementing multiuser applications: a case study. Software—Practice and Experience, 23(1):75– 93, January 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Davidson:1984:RAE

[DF84] Jack W. Davidson and Christopher W. Fraser. Register allocation and exhaustive peephole optimization. Software—Practice and Experience, 14(9): 857–865, September 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Davidson:1987:AIF

[DF87] Jack W. Davidson and Christopher W. Fraser. Automatic inference and fast interpretation of peephole optimization rules. Software—Practice and Experience, 17(11):801–812, November 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dunstan:1995:PSU

[DF95] Neil Dunstan and Ivan Fris.

Process scheduling and UNIX semaphores. Software—Practice and Experience, 25(10):1141–1153, October 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dvinsky:2015:ERC

[DF15] Alex Dvinsky and Roy Friedman. Experience reports: Chameleon — a group communication framework for smartphones. Software—Practice and Experience, 45(10):1429– 1455, October 2015. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Denys:2023:DCC

[DFD23] Pierre-Frédérick Denys, Quentin Fournier, and Michel R. Dagenais. Distributed computation of the critical path from execution traces. Software —Practice and Experience, 53(8):1722–1737, August 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DeLucia:2010:FGM

[DFOT10] Andrea De Lucia, Fausto Fasano, Rocco Oliveto, and Genoveffa Tortora. Finegrained management of software artefacts: the ADAMS system. Software —Practice and Experience, 40(11):1007–1034, October 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DeLucia:2008:MLV

[DFPT08] Andrea De Lucia, Rita Francese, Ignazio Passero, and Genoveffa Tortora. Migrating legacy video lectures to multimedia learning objects. Software— Practice and Experience, 38(14):1499–1530, November 25, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DeLucia:2009:DES

[DFPT09] Andrea De Lucia, Rita Francese, Ignazio Passero, and Genoveffa Tortora. Development and evaluation of a system enhancing Second Life to support synchronous role-based collaborative learning. Software—Practice and Experience, 39(12):1025–1054, August 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Deng:2022:FLB

[DFQ^{+}22] Fengyang Deng, Cai Fu, Yekui Qian, Jia Yang, Shuai He, and Hao Xu. Federated learning based multi-task feature fusion framework for code expressive semantic extraction. Software—Practice and Experience, 52(8): 1849–1866, August 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Duran-Faundez:2015:ERA

[DFRR15] Cristian Duran-Faundez, Mario Ramos, and Pedro Rodriguez. Experience report: Applying Gaia and AUML for the development of multiagent-based control software for flexible manufacturing systems: addressing methodological and implementation issues. Software—Practice and Experience, 45(12):1719–1737, December 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DeLucia:2008:DLS

[DFST08] Andrea De Lucia, Rita Francese, Giuseppe Scanniello, and Genoveffa Tortora. Developing legacy system migration methods and tools for technology transfer. Software— Practice and Experience, 38(13):1333–1364, November 10, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dai:2012:HBB

[DFW^{+}12] Shih-Yao Dai, Yarochkin Fyodor, Ming-Wei Wu, Yennun Huang, and Sy-Yen Kuo. Holography: a behavior-based profiler for malware analysis. Software—Practice and Experience, 42(9):1107–1136, September 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dedourek:1980:SD

[DGM80] John M. Dedourek, Uday G.

Gujar, and Marion E. McIntyre. Scanner design. Software—Practice and Experience, 10(12): 959–972, December 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Decaroli:2019:CIO

[DGM19] Gianni Decaroli, Travis Gagie, and Giovanni Manzini. A compact index for orderpreserving pattern matching. Software—Practice and Experience, 49(6): 1041–1051, June 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Das:2022:REG

[DGM^{+}22] Jaydeep Das, Shreya Ghosh, Anwesha Mukherjee, Soumya K. Ghosh, and Rajkumar Buyya. RESCUE: Enabling green healthcare services using integrated IoT-edge-fog-cloud computing environments. Software—Practice and Experience, 52(7):1615–1642, July 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Deufemia:2014:VLB

[DGPT14] Vincenzo Deufemia, Massimiliano Giordano, Giuseppe Polese, and Genoveffa Tortora. A visual languagebased system for extraction– transformation–loading development. Software— Practice and Experience, 44(12):1417–1440, December 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Diaz:2006:ECO

[DGR^{+}06] Manuel Díaz, Daniel Garrido, Sergio Romero, Bartolomé Rubio, Enrique Soler, and José M. Troya. Experiences with componentoriented technologies in nuclear power plant simulators. Software—Practice and Experience, 36(13): 1489–1512, November 10, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dastjerdi:2015:CFQ

[DGRB15] Amir Vahid Dastjerdi, Saurabh Kumar Garg, Omer F. Rana, and Rajkumar Buyya. Cloud-Pick: a framework for QoSaware and ontology-based service deployment across clouds. Software—Practice and Experience, 45(2):197– 231, February 2015. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dongarra:1979:ULF

[DH79] J. J. Dongarra and A. R.

Hinds. Unrolling loops in FORTRAN. Software— Practice and Experience, 9 (3):219–226, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Davidson:1988:SCF

[DH88] Jack W. Davidson and Anne M. Holler. A study of a C function inliner. Software—Practice and Experience, 18(8):775–790, August 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dearle:2000:OSS

[DH00] Alan Dearle and David Hulse. Operating system support for persistent systems: past, present and future. Software—Practice and Experience, 30(4):295– 324, April 10, 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
71004138/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?

ID=71004138&PLACEBO=IE.
pdf.

DiPenta:2011:USB

[DHA11] Massimiliano Di Penta, Mark Harman, and Giuliano Antoniol. The use of search-based optimization techniques to schedule and staff software projects: an approach and an empirical study. Software—Practice and Experience, 41(5):495– 519, May 25, 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Drave:2019:SMA

[DHG^{+}19] Imke Drave, Steffen Hillemacher, Timo Greifenberg, Stefan Kriebel, Evgeny Kusmenko, Matthias Markthaler, Philipp Orth, Karin Samira Salman, Johannes Richenhagen, Bernhard Rumpe, Christoph Schulze, Michael von Wenckstern, and Andreas Wortmann. SMArDT modeling for automotive software testing. Software —Practice and Experience, 49(2):301–328, February 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Diaz-Herrera:1992:AMK

[DHGR92] Jorge L. Díaz-Herrera, Ronald D. Graft, and Douglas B. Rupp. ARTK-M2: A kernel for Ada tasking requirements: An implementation and an automatic generator. Software—Practice and Experience, 22(4): 317–348, April 1992. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Diwan:2011:TSP

[DHMS11] Amer Diwan, Matthias Hauswirth, Todd Mytkowicz, and Peter F. Sweeney. TraceAnalyzer: a system for processing performance traces. Software—Practice and Experience, 41(3):267– 282, March 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Das:2001:SCR

[DHS01] Rajsekhar Das, Matti A.

Hiltunen, and Richard D. Schlichting. Supporting configurability and real time in RTD channels. Software—Practice and Experience, 31(12):1183– 1209, October 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
85011922/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=85011922&PLACEBO=IE.
pdf.

Daley:2002:FTD

[DHS02] Nigel Daley, Daniel Hoffman, and Paul Strooper. A framework for table driven testing of Java classes. Software—Practice and Experience, 32(5):465– 493, April 25, 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
93513336/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=93513336&PLACEBO=IE.
pdf.

Ding:2014:FBH

[DHWZ14] Weilong Ding, Yanbo Han, Jing Wang, and Zhuofeng Zhao. Feature-based highavailability mechanism for quantile tasks in realtime data stream processing. Software—Practice and Experience, 44(7):855– 871, July 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Diehl:1997:EAM

[Die97] Stephan Diehl. An experiment in abstract machine design. Software—Practice and Experience, 27(1):49– 62, January 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7257;   http://www3.

interscience.wiley.com/
cgi-bin/fulltext?ID=7257&
PLACEBO=IE.pdf.

Diehl:1998:FIC

[Die98] Stephan Diehl. A formal introduction to the compilation of Java. Software—Practice and Experience, 28(3):297–327, March 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1782;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1782&
PLACEBO=IE.pdf.

Demartini:1999:DDT

[DIS99] Claudio Demartini, Radu Iosif, and Riccardo Sisto. A deadlock detection tool for concurrent Java programs. Software—Practice and Experience, 29(7): 577–603, June 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=62003577;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=62003577&PLACEBO=IE.
pdf.

Dawson:1997:ESC

[DJM97] Scott Dawson, Farnam Jahanian, and Todd Mitton.

Experiments on six commercial TCP implementations using a software fault injection tool. Software—Practice and Experience, 27(12):1385–1410, December 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7277;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7277&
PLACEBO=IE.pdf.

deKretser:2004:SSE

[dKM04] Owen de Kretser and Alistair Moffat. SEFT: a search engine for text. Software—Practice and Experience, 34(10):1011–1023, August 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dubey:2011:CMH

[DKM11] Abhishek Dubey, Gabor Karsai, and Nagabhushan Mahadevan. A component model for hard realtime systems: CCM with ARINC-653. Software— Practice and Experience, 41(12):1517–1550, November/December 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dementiev:2008:SST

[DKS08] R. Dementiev, L. Kettner, and P. Sanders. Stxxl: standard template library for XXL data sets. Software —Practice and Experience, 38(6):589–637, May ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dusi:2023:QSC

[DL23] Michele Dusi and Gianfranco Lamperti. Quick subset construction. Software—Practice and Experience, 53(11):2092–2132, November 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dakin:1985:LSN

[DLP85] R. J. Dakin, B. R. Lederer, and K. R. Parker. A large scale network storage facility. Software—Practice and Experience, 15(9):889–899, September 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

laRosa:2023:EUB

[dlRNGCC23] Miguel Sánchez de la Rosa, Carlos Núñez-Gómez, M. Blanca Caminero, and Carmen Carrión. Exploring the use of blockchain in resourceconstrained fog computing environments. Software— Practice and Experience, 53(4):971–987, April 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dayarathna:2017:ECA

[DLWF17] Miyuru Dayarathna, Yuanlong Li, Yonggang Wen, and Rui Fan. Energy consumption analysis of data stream processing: a benchmarking approach. Software—Practice and Experience, 47(10): 1443–1462, October 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dai:2022:CCC

[DLX^{+}22] Fei Dai, Guozhi Liu, Xiaolong Xu, Qi Mo, Zhenping Qiang, and Zhihong Liang. Compatibility checking for cyber-physical systems based on microservices. Software—Practice and Experience, 52 (11):2393–2410, November 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dewar:1977:MSS

[DM77] Robert B. K. Dewar and A. P. McCann. Macro SPITBOL — a SNOBOL4 compiler. Software—Practice and Experience, 7(1): 95–113, January/March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Doyle:1984:PPS

[DM84] J. K. Doyle and K. I.

Mandelberg. A portable PDP-11 simulator. Software—Practice and Experience, 14(11):1047–1059, November 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Drechsler:2007:YSL

[DM07] R. L. Drechsler and J. M.

Mocenigo. The Yoix(R) scripting language: a different way of writing Java^{T M} applications. Software— Practice and Experience, 37(6):643–667, May 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ducournau:2011:PCH

[DM11] Roland Ducournau and Floréal Morandat. Perfect class hashing and numbering for object-oriented implementation. Software —Practice and Experience, 41(6):661–694, May 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dhillon:2015:EFC

[DM15] Sunny Dhillon and Qusay H.

Mahmoud. An evaluation framework for crossplatform mobile application development tools. Software—Practice and Experience, 45(10):1331– 1357, October 2015. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Do:2017:DYW

[DMC17] Quang Do, Ben Martini, and Kim-Kwang Raymond Choo. Is the data on your wearable device secure? An Android Wear smartwatch case study. Software—Practice and Experience, 47(3):391–403, March 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DeAntonellis:2006:LAF

[DMD^{+}06] Valeria De Antonellis, Michele Melchiori, Luca De Santis, Massimo Mecella, Enrico Mussi, Barbara Pernici, and Pierluigi Plebani. A layered architecture for flexible Web service invocation. Software—Practice and Experience, 36(2):191– 223, February 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deMoura:1999:SE

[dMdLvS99] Leonardo M. de Moura, Carlos José P. de Lucena, and Arndt von Staa. The Spider environment. Software—Practice and Experience, 29(2):99– 124, February 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55000304;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55000304&PLACEBO=IE.
pdf.

deMurillas:2017:PCR

[dMFÁE17] Eduardo González-López de Murillas, Javier Fabra, Pedro Álvarez, and Joaquín Ezpeleta. Parallel computation of the reachability graph of Petri net models with semantic information. Software—Practice and Experience, 47(5):647– 668, May 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dharminder:2022:PPS

[DMR^{+}22] Dharminder Dharminder, Dheerendra Mishra, Joel J. P. C. Rodrigues, Ricardo de A. L. Rabelo, and Kashif Saleem. PSSCC: Provably secure communication framework for crowdsourced industrial Internet of Things environments. Software—Practice and Experience, 52(3):744– 755, March 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dubnicki:1988:EEI

[DMW88] Cezary Dubnicki, Jan Madey, and Wojciech Wygladała. Edison-N: an Edison implementation for a network of microcomputers. Software—Practice and Experience, 18(4):349– 363, April 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dow:2020:CPR

[DNL^{+}20] Chyi-Ren Dow, Huu-Huy Ngo, Liang-Hsuan Lee, Po-Yu Lai, Kuan-Chieh Wang, and Van-Tung Bui. A crosswalk pedestrian recognition system by using deep learning and zebracrossing recognition techniques. Software—Practice and Experience, 50(5):630– 644, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dietrich:1989:TOO

[DNSG89] Walter C. Dietrich, Jr., Lee R. Nackman, Christine J. Sundaresan, and Franklin Gracer. TGMS: An object-oriented system for programming geometry. Software—Practice and Experience, 19(10): 979–1013, October 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Douglis:1991:TPM

[DO91] Fred Douglis and John K.

Ousterhout. Transparent process migration: Design alternatives and the Sprite implementation. Software—Practice and Experience, 21(8):757–785, August 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Denti:1999:ATB

[DO99] Enrico Denti and Andrea Omicini. An architecture for tuple-based coordination of multiagent systems. Software—Practice and Experience, 29(12):1103–1121, October 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
66000403/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=66000403&PLACEBO=IE.
pdf.

Distler:2007:CSD

[DO07] Thomas James Distler and Jinsong Ouyang. Clock synchronization for distributed media applications. Software—Practice and Experience, 37(14): 1489–1514, November 25, 2007. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Doddin:1978:POC

[Dod78] W. P. Doddin. In praise of optimizing compilers. Software—Practice and Experience, 8(6):755–756, November 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dodd:1982:ANT

[Dod82] W. P. Dodd. An assessment of a new technique for the representation of Boolean matrices. Software—Practice and Experience, 12(2): 115–139, February 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deOliveira:2016:TAP

[dOdO16] Daniel Bristot de Oliveira and Romulo Silva de Oliveira. Timing analysis of the PREEMPT RT Linux kernel. Software—Practice and Experience, 46(6):789– 819, June 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Araujo-de-Oliveira:2021:PFA

[dODP21] Patrícia Araújo de Oliveira, Francisco Durán, and Ernesto Pimentel. A procedural and flexible approach for specification, modeling, definition, and analysis for self-adaptive systems. Software—Practice and Experience, 51(6): 1387–1415, June 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deOliveira:2020:OCR

[dOED^{+}20] Edvard Martins de Oliveira, Júlio Cézar Estrella, Alexandre Claudio Botazzo Delbem, Mário Henrique Souza Pardo, Fausto Guzzo da Costa, Alexandre Defelicibus, and Stephan Reiff-Marganiec. Optimizing computational resource management for the scientific gateways ecosystems based on the service-oriented paradigm. Software—Practice and Experience, 50(6):899–924, June 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Doolan:1992:EFI

[Doo92] E. P. Doolan. Experience with Fagan’s inspection method. Software—Practice and Experience, 22(2): 173–182, February 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dubery:1985:SAP

[DP85] J. M. Dubery and A. J.

Pinches. Software for an air pollution measuring system: an application of Modula. Software —Practice and Experience, 15(4):413–422, April 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Degano:1995:CSE

[DP95] Pierpaolo Degano and Corrado Priami. Comparison of syntactic error handling in LR parsers. Software —Practice and Experience, 25(6):657–679, June 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Demko:2009:SOS

[DP09] Aleksander B. Demko and Nick J. Pizzi. Scopira: an open source C++ framework for biomedical data analysis applications. Software—Practice and Experience, 39(6):641–660, April 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Diaz:2011:APL

[DPAG11] Jessica Díaz, Jennifer Pérez, Pedro P. Alarcón, and Juan Garbajosa. Agile product line engineering — a systematic literature review. Software—Practice and Experience, 41(8):921– 941, July 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dias:2014:FFG

[DPDA14] Martín Dias, Mariano Martinez Peck, Stéphane Ducasse, and Gabriela Arévalo. Fuel: a fast general purpose object graph serializer. Software—Practice and Experience, 44(4):433–453, April 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Special issue on International Workshop on Smalltalk Technologies 2011.

Dagkakis:2016:MOS

[DPH16] Georgios Dagkakis, Ioannis Papagiannopoulos, and Cathal Heavey. ManPy: an open-source software tool for building discrete event simulation models of manufacturing systems. Software —Practice and Experience, 46(7):955–981, July 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Delange:2012:DIV

[DPK12] Julien Delange, Laurent Pautet, and Fabrice Kordon. Design, implementation and verification of MILS systems. Software —Practice and Experience, 42(7):799–816, July 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DeCapitanidiVimercati:2003:ACP

[DPS03] Sabrina De Capitani di Vimercati, Stefano Paraboschi, and Pierangela Samarati. Access control: principles and solutions. Software—Practice and Experience, 33(5):397– 421, April 25, 2003. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deRidder:1986:CCR

[dR86] Theo de Ridder. Coroutines for C reconsidered. Software—Practice and Experience, 16(3):301–302, March 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dershowitz:1990:CC

[DR90] Nachum Dershowitz and Edward M. Reingold. Calendrical calculations. Software—Practice and Experience, 20(9):899–928, September 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dodd:1992:MDD

[DR92] Paul S. Dodd and Chinya V. Ravishankar. Monitoring and debugging distributed real-time programs. Software—Practice and Experience, 22(10):863–877, October 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dembitz:2011:AOS

[DRG11] Šandor Dembitz, Mirko Randić, and Gordan Gledec. Advantages of online spellchecking: a Croatian example. Software—Practice and Experience, 41(11): 1203–1231, October 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Droge:2023:ECP

[DRH^{+}23] Bob Dröge, Victor Holanda Rusu, Kenneth Hoste, Caspar van Leeuwen, Alan O’Cais, and Thomas Röblitz. EESSI: a cross-platform ready-to-use optimised scientific software stack. Software—Practice and Experience, 53(1):176–210, January 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Drizis:1993:MFT

[Dri93] Leonidas Drizis. A method for fast tape backups and restores. Software—Practice and Experience, 23(7): 813–815, July 1993. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dodds:1982:DMS

[DRL82] Robert H. Dodds, Jr., Daniel R. Rehak, and Leonard A. Lopez. Development methodologies for scientific software. Software—Practice and Experience, 12(12):1085–1100, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dromey:1984:EPO

[Dro84] R. Geoff Dromey. Exploiting partial order with Quicksort. Software— Practice and Experience, 14(6):509–518, June 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dromey:1985:FTL

[Dro85a] R. Geoff Dromey. Forced termination of loops. Software—Practice and Experience, 15(1):29–39, January 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dromey:1985:PDI

[Dro85b] R. Geoff Dromey. Program development by inductive stepwise refinement. Software—Practice and Experience, 15(1):1–28, January 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dromey:1986:ASP

[Dro86] R. Geoff Dromey. An algorithm for the selection problem. Software— Practice and Experience, 16(11):981–986, November 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Righi:2015:ROR

[dRRGdC15] Rodrigo da Rosa Righi, Lucas Graebin, and Cristiano André da Costa. On the replacement of objects from round-based applications over heterogeneous environments. Software —Practice and Experience, 45(5):633–656, May 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DiIorio:2013:E

[DRZ13] Angelo Di Iorio, Davide Rossi, and Stefano Zacchiroli. Editorial. Software—Practice and Experience, 43(12):1393–1394, December 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dandamudi:1986:ABT

[DS86a] Sivarama P. Dandamudi and Paul G. Sorenson. Algorithms for BD trees. Software—Practice and Experience, 16(12):1077– 1096, December 1986. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dowsing:1986:WCA

[DS86b] R. D. Dowsing and M. T. Sanderson. Writing concurrent assemblers — a case study in Path Pascal. Software—Practice and Experience, 16(12):1117–1135, December 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dandamudi:1988:PAP

[DS88] Sivarama P. Dandamudi and Paul G. Sorenson. Performance analysis of partial-match search algorithms for BD trees. Software—Practice and Experience, 18(1):83–105, January 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dyreson:1994:ETI

[DS94] Curtis E. Dyreson and Richard T. Snodgrass. Efficient timestamp input and output. Software—Practice and Experience, 24(1):89– 109, January 1994. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dearnley:1999:DUE

[DS99] P. A. Dearnley and D. J. Smith. Discovering and using entity mappings in federated databases. Software—Practice and Experience, 29(1):17–42, January 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55002653;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55002653&PLACEBO=IE.
pdf.

DiGaspero:2003:EOO

[DS03] Luca Di Gaspero and Andrea Schaerf. EasyLocal++: an object-oriented framework for the flexible design of local-search algorithms. Software—Practice and Experience, 33(8): 733–765, July 10, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DiStefano:2009:AAA

[DS09] Antonella Di Stefano and Corrado Santoro. A^{3}M: an agent architecture for automated manufacturing. Software—Practice and Experience, 39(2):137– 162, February ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Drusinsky:2012:VQA

[DS12] Doron Drusinsky and Man-Tak Shing. Validating quality attribute requirements via execution-based model checking. Software—Practice and Experience, 42(7): 853–875, July 2012. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deSouza:2016:VLU

[dSC16] Alessandro J. de Souza and Anderson Luiz O. Cavalcanti. Visual language for use case description. Software—Practice and Experience, 46(9):1239–1261, September 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cimino:2019:MSI

[dSCdRS^{+}19] Leonardo de Souza Cimino, José Estevão Eugênio de Resende, Lucas Henrique Moreira Silva, Samuel Queiroz Souza Rocha, Matheus de Oliveira Correia, Guilherme Souza Monteiro, Gabriel Natã de Souza Fernandes, Renan da Silva Moreira, Junior Guilherme de Silva, Matheus Inácio Batista Santos, Andre Luiz Lins Aquino, André Luís Barroso Almeida, and Joubert de Castro Lima. A middleware solution for integrating and exploring IoT and HPC capabilities. Software —Practice and Experience, 49(4):584–616, April 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Davis:2005:UCH

[DSD^{+}05] Michael Davis, Randy Smith, Brandon Dixon, Allen Parrish, and David Cordes. Utilizing commodity hardware and software to distribute a realworld application: maximizing reuse while improving performance. Software —Practice and Experience, 35(7):621–641, June 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Duggan:2019:MSA

[DSD^{+}19] M. Duggan, R. Shaw, J. Duggan, E. Howley, and E. Barrett. A multitimesteps-ahead prediction approach for scheduling live migration in cloud data centers. Software—Practice and Experience, 49(4): 617–639, April 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

daSilva:2011:ASP

[dSdMSNO^{+}11] Ivonei Freitas da Silva, Paulo Anselmo da Mota Silveira Neto, Pádraig O’Leary, Eduardo Santana de Almeida, and Silvio Romero de Lemos Meira. Agile software product lines: a systematic mapping study. Software— Practice and Experience, 41(8):899–920, July 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Junior:2016:EIA

[dSJCM16] Jose Ricardo da Silva Junior, Esteban Clua, and Leonardo Murta. Efficient image-aware version control systems using GPU. Software—Practice and Experience, 46(8): 1011–1033, August 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Maciel:2013:IPF

[dSMH13] Luiz Alexandre Hiane da Silva Maciel and Celso Massaki Hirata. Invited paper: Fault-tolerant timestampbased two-phase commit protocol for RESTful services. Software—Practice and Experience, 43 (12):1459–1488, December 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Magalhaes:2022:AIP

[dSMLP22] José Wesley de Souza Magalhães, Chunhua Liao, and Fernando Magno Quintão Pereira. Automatic inspection of program state in an uncooperative environment. Software—Practice and Experience, 52 (12):2727–2758, December 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rocha:2021:ADC

[dSRdSS^{+}21] Élisson da Silva Rocha, Leylane G. F. da Silva, Guto L. Santos, Diego Bezerra, André Moreira, Glauco Gonçalves, Maria Valéria Marquezini, Amardeep Mehta, Mattias Wildeman, Judith Kelner, Djamel Sadok, and Patricia T. Endo. Aggregating data center measurements for availability analysis. Software—Practice and Experience, 51(5):868–892, May 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dunman:1982:MIC

[DSW82] B. R. Dunman, Stephen R.

Schach, and Peter T. Wood. A mainframe implementation of Concurrent Pascal. Software—Practice and Experience, 12(1):85– 89, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DeBosschere:1996:BBE

[DT96] Koenraad De Bosschere and Paul Tarau. Blackboardbased extensions in Prolog. Software—Practice and Experience, 26(1):49– 69, January 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=61004280.

Dastjerdi:2012:DAO

[DTB12] Amir Vahid Dastjerdi, Sayed Gholam Hassan Tabatabaei, and Rajkumar Buyya. A dependencyaware ontology-based approach for deploying service level agreement monitoring services in Cloud. Software—Practice and Experience, 42(4):501–518, April 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dikshit:1989:STB

[DTJ89] Piyush Dikshit, Satish K. Tripathi, and Pankaj Jalote. SAHAYOG: A test bed for evaluating dynamic load-sharing policies. Software—Practice and Experience, 19(5):411–435, May 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ducournau:2011:CVT

[Duc11] Roland Ducournau. Coloring, a versatile technique for implementing objectoriented languages. Software—Practice and Experience, 41(6):627–659, May 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dunn:1974:AVO

[Dun74] Richard C. Dunn. APL versus operator precedence. Software—Practice and Experience, 4(3):299– 300, July/September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dunn:1975:PPS

[Dun75] Raymond D. Dunn. POP-2 performance on a small 1900 machine. Software —Practice and Experience, 5(2):216–217, April 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dundas:1991:IDM

[Dun91] John A. Dundas III. Implementing dynamic minimalprefix tries. Software— Practice and Experience, 21(10):1027–1040, October 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dunlavey:1993:DEC

[Dun93] Michael R. Dunlavey. Differential evaluation: a cache-based technique for incremental update of graphical displays of structures. Software—Practice and Experience, 23(8):871– 893, August 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Derman:1984:SES

[DV84] Emanuel Derman and Christopher J. Van Wyk. A simple equation solver and its application to financial modelling. Software—Practice and Experience, 14(12):1169–1181, December 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DeBruin:1985:CVN

[DV85] R. De Bruin and C. G. Van Der Laan. The creation of a virtual NAG–Algol 68 program library. Software —Practice and Experience, 15(10):963–972, October 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deVet:1989:PAE

[dV89] John H. M. de Vet. A practical algorithm for evaluating database queries. Software—Practice and Experience, 19(5):491–504, May 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Davies:1973:PPU

[DW73] G. Davies and R. Welland. A pre-processor using rule mask techniques for extended entry decision tables. Software—Practice and Experience, 3(3):227– 234, July/September 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Davidson:1989:QCU

[DW89] Jack W. Davidson and David B. Whalley. Quick compilers using peephole optimization. Software— Practice and Experience, 19(1):79–97, January 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Du:1990:EIT

[DW90] Weichang Du and William W.

Wadge. The eductive implementation of a threedimensional spreadsheet. Software—Practice and Experience, 20(11):1097– 1114, November 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Davidson:1991:MSR

[DW91] Jack W. Davidson and David B. Whalley. Methods for saving and restoring register values across function calls. Software—Practice and Experience, 21(2): 149–165, February 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Debroy:2013:CBS

[DW13] Vidroha Debroy and W. Eric Wong. A consensus-based strategy to improve the quality of fault localization. Software—Practice and Experience, 43(8):989– 1011, August 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dubey:2015:OVM

[DWL^{+}15] Anshu Dubey, Klaus Weide, Dongwook Lee, John Bachan, Christopher Daley, Samuel Olofin, Noel Taylor, Paul M. Rich, and Lynn B. Reid. Ongoing verification of a multiphysics community code: FLASH. Software —Practice and Experience, 45(2):233–244, February 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ding:2017:ELT

[DWL^{+}17] Yue Ding, Dong Wang, Guoqiang Li, Daniel Sun, Xin Xin, and Shiyou Qian. Exploiting long-term and short-term preferences and RFID trajectories in shop recommendation. Software —Practice and Experience, 47(6):849–865, June 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DeIasio:2021:FMS

[DZ21] Antonio De Iasio and Eugenio Zimeo. A framework for microservices synchronization. Software—Practice and Experience, 51(1):25– 45, January 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dong:2009:XBE

[DZS09] Jing Dong, Yajing Zhao, and Yongtao Sun. XSLTbased evolutions and analyses of design patterns. Software—Practice and Experience, 39(8):773–805, June 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Evans:2003:PEG

[EAB^{+}03] Huw Evans, Malcolm Atkinson, Margaret Brown, Julie Cargill, Murray Crease, Steve Draper, Phil Gray, and Richard Thomas. The pervasiveness of evolution in GRUMPS software. Software—Practice and Experience, 33(2):99–120, February 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
102525915/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?

                                        ID=102525915&PLACEBO=
                                        IE.pdf.

Earnshaw:1976:GPA

[Ear76] R. A. Earnshaw. Graph plotting in ALGOL 68r. Software—Practice and Experience, 6(1):51– 60, January/March 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Earnshaw:1977:BRB

[Ear77] R. A. Earnshaw. Book review: A course in APL with applications, L. D. Grey, Addison-Wesley Co., Reading, Mass. 1976. No. of pages: 358. Price: £8.00. Software—Practice and Experience, 7(6):798, November 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ebad:2018:ICS

[Eba18] Shouki A. Ebad. The influencing causes of software unavailability: a case study from industry. Software—Practice and Experience, 48(5): 1056–1076, May 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2569.

Ebad:2020:HSD

[Eba20] Shouki A. Ebad. Healthcare software design and implementation — a project failure case. Software— Practice and Experience, 50(7):1258–1276, July 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Elshoff:1974:HAI

[EBD^{+}74] James L. Elshoff, Robert Beckermeyer, John Dill, Michael Marcotty, and John Murray. Handling asynchronous interrupts in a PL/1-like language. Software—Practice and Experience, 4(2):117–124, April/ June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Estrella-Balderrama:2010:GTS

[EBFK10] A. Estrella-Balderrama, J. J. Fowler, and S. G. Kobourov. GraphSET, a tool for simultaneous graph drawing. Software—Practice and Experience, 40 (10):849–863, September 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Escolar:2013:OFT

[EC13] Soledad Escolar and Jesús Carretero. An open framework for translating portable applications into operating system-specific wireless sensor networks applications. Software —Practice and Experience, 43(3):333–357, March 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Edmunds:1982:BRB

[Edm82] G. Edmunds. Book review: A practical introduction to computer graphics, Ian O. Angell, Macmillan, London, 1981. No. of pages:

160. Price: £5.95. Software —Practice and Experience, 12(9):884, September 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Edmunds:1986:BRB

[Edm86] G. Edmunds. Book review: Essentials of Fortran 77, John Shelley, Wiley, Chichester. 1984. no of pages: 218. price: £8.25. Software—Practice and Experience, 16(1):99– 100, January 1986. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Edwards:1977:BRB

[Edw77] R. Edwards. Book review: Algorithms + Data Structures = Programs, N. Wirth, Prentice-Hall Inc., Engelwood Cliffs, N.J., 1976. no. of pages: 366. price: £9.70. Software— Practice and Experience, 7 (2):303–304, March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Edwards:1998:BRB

[Edw98a] Alistair D. N. Edwards.

Book review: The Essence of Human-computer Interaction by Christine Faulkner, Prentice Hall, 1998. ISBN: 0-13-751975-3. Software—Practice and Experience, 28(8):897, July 10, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Edwards:1998:BRE

[Edw98b] Alistair D. N. Edwards.

Book review; ‘The Essence of Human-computer Interaction’ by Christine Faulkner, Prentice Hall, 1998. ISBN: 0-13-751975-3. Software—Practice and Experience, 28(8):897, July 10, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1812;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1812&
PLACEBO=IE.pdf.

Eldin:1990:VSF

[EE90] A. Sharaf Eldin and D. J.

Evans. A virtual stack facility for mini-computer and microcomputer implementation of the finite element method. Software—Practice and Experience, 20(12):1209–1226, December 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ehsan:2013:GCS

[EF13] Nava Ehsan and Heshaam Faili. Grammatical and context-sensitive error correction using a statistical machine translation framework. Software—Practice and Experience, 43(2):187– 206, February 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Einarsson:1984:MLP

[EG84] Bo Einarsson and W. Morven Gentleman. Mixed language programming. Software—Practice and Experience, 14(4):383–396 (or 383–395??), April 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

ElHajAhmed:2021:KDK

[EGCCM21] Ghofrane El Haj Ahmed, Felipe Gil-Castiñeira, and Enrique Costa-Montenegro. KubCG: a dynamic Kubernetes scheduler for heterogeneous clusters. Software —Practice and Experience, 51(2):213–234, February 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ertl:2002:VGE

[EGKP02] M. Anton Ertl, David Gregg, Andreas Krall, and Bernd Paysan. Vmgen

— a generator of efficient virtual machine interpreters. Software—Practice and Experience, 32(3):265– 294, March 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www.
complang.tuwien.ac.at/
anton/vmgen/;   http:
//www.complang.tuwien.
ac.at/papers/ertl+02.
ps.gz;   http://www3.
interscience.wiley.com/
cgi-bin/abstract/90010508/
START;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=90010508&
PLACEBO=IE.pdf.

Echevarria:2018:ESA

[EGL18] Juan Jose Echevarria, Pablo Garaizar, and Jon Legarda. An experimental study on the applicability of SYN cookies to networked constrained devices. Software—Practice and Experience, 48 (3):740–749, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/abs/10.
                                        1002/spe.2510.

Ehrman:1973:LE

[Ehr73] John R. Ehrman. Letter to the editor. Software— Practice and Experience, 3 (2):187, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Etalle:1999:DSP

[EHV99] Sandro Etalle, Pieter H. Hartel, and Willem G. Vree. Declarative solutions to partitionedgrid problems. Software—Practice and Experience, 29(13):1173–1200, November 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        66501673/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=66501673&PLACEBO=IE.
                                        pdf.

Einbu:1988:AAI

[Ein88] John M. Einbu. An architectural approach to improved program maintainability. Software—Practice and Experience, 18(1):51– 62, January 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Egan:1999:FTR

[EKM^{+}99] A. Egan, D. Kutz, D. Mikulin, R. Melhem, and D. Mossé. Fault-tolerant RT-Mach (FT-RT-Mach) and an application to real-time train control. Software—Practice and Experience, 29(4):379–395, April 10, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55001841;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55001841&PLACEBO=IE.
pdf.

Eick:1996:DTF

[EL96] Stephen G. Eick and Paul J. Lucas. Displaying trace files. Software—Practice and Experience, 26(4): 399–409, April 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16805.

Elfatatry:2005:NDL

[EL05] Ahmed Elfatatry and Paul Layzell. A negotiation description language. Software—Practice and Experience, 35(4):323–343, April 10, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ellman:1972:BRB

[Ell72]  D. Ellman.  Book re-

view: Fitting equations to data: Computer analysis of multifactor data for scientists and engineers, C. Daniel and F. S. Wood, Wiley, New York, 1971. No. of pages: 368. Price £8. Software—Practice and Experience, 2(3):306– 307, July 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ellis:1979:PPA

[Ell79a] T. M. R. Ellis. Parallel processing in an adaptable application oriented language processor. Software—Practice and Experience, 9(3): 183–190, March 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ellis:1979:UDS

[Ell79b] T. M. R. Ellis. A useful data-structuring technique. Software—Practice and Experience, 9(6):477– 484, June 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Elliott:1982:DSS

[Ell82a] Brig Elliott. Design of a simple screen editor. Software—Practice and Experience, 12(4):375–384, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Elliott:1982:HLD

[Ell82b] Brig Elliott. A high-level debugger for PL/I, FOR-TRAN and BASIC. Software—Practice and Experience, 12(4):331–340, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Englebert:1993:GAI

[ELRV93] Vincent Englebert, Baudouin Le Charlier, Didier Roland, and Pascal Van Hentenryck. Generic abstract interpretation algorithms for Prolog: Two optimization techniques and their experimental evaluation. Software—Practice and Experience, 23(4):419– 459, April 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Elshoff:1976:NPC

[Els76] James L. Elshoff. A numerical profile of commercial PL/I programs. Software—Practice and Experience, 6(4):505–525, October/December 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ebenstein:1990:OTP

[EM90] Samuel E. Ebenstein and Tracy L. McDermott. Optimization techniques for parallel processing. Software—Practice and Experience, 20(8):833–849, August 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

ElBoussaidi:2012:UDP

[EM12] Ghizlane El Boussaidi and Hafedh Mili. Understanding design patterns — what is the problem? Software—Practice and Experience, 42(12):1495–1529, December 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Eynard:2013:ECP

[EMD13] Davide Eynard, Luca Mazzola, and Antonina Dattolo. Extended conference paper: Exploiting tag similarities to discover synonyms and homonyms in folksonomies. Software—Practice and Experience, 43(12):1437–1457, December 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Emery:1984:BRB

[Eme84]  G. Emery.  Book re-

view: Computer Studies, a Practical Approach, G. M. Croft, Hodder & Stoughton. No of pages:

234. Price: £4.25. Software —Practice and Experience, 14(6):602–603, June 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Elahi:2020:TSC

[EMRK20] Bukhtawar Elahi, Asad Waqar Malik, Anis U. Rahman, and Muazzam A. Khan. Toward scalable cloud data center simulation using high-level architecture. Software—Practice and Experience, 50(6): 827–843, June 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ellis:1983:TET

[EMVW83] John R. Ellis, Nathaniel Mishkin, Mary-Claire Van Leunen, and Steven R. Wood. Tools: An environment for time-shared computing and programming. Software—Practice and Experience, 13(10): 873–892, October 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Engebretsen:2006:PIC

[Eng06] Lars Engebretsen. Platformindependent code conversion within the C++ locale framework. Software—Practice and Experience, 36(15):1643–1654, December 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Eramo:2024:MDE

[ENP^{+}24] Romina Eramo, Martina Nolletti, Luigi Pomante, Laura Pasquale, and Dario Pascucci. Model-driven engineering for simulation models interoperability: a case study in space industry. Software—Practice and Experience, 54(6): 1010–1033, June 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Eaglestone:1979:CNB

[EP79] B. M. Eaglestone and N. J. Partington. A campus network based on ICL 2900 series protocol. Software —Practice and Experience, 9(11):959–967, November 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Eggert:2005:PEN

[EP05] P. R. Eggert and D. S. Parker. Perturbing and evaluating numerical programs without recompilation — the wonglediff way. Software—Practice and Experience, 35(4):313– 322, April 10, 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Er:1983:OPC

[Er83] M. C. Er. Optimizing procedure calls and returns. Software—Practice and Experience, 13(10): 921–939, October 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Er:1985:PCG

[Er85] M. C. Er. Practical considerations of global and local variables. Software —Practice and Experience, 15(5):499–502, May 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Etchevers:2017:RSD

[ESB^{+}17] Xavier Etchevers, Gwen Salaün, Fabienne Boyer, Thierry Coupaye, and Noel De Palma. Reliable selfdeployment of distributed cloud applications. Software—Practice and Experience, 47(1):3–20, January 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Estebanez:2014:PMC

[ESRI14] César Estébanez, Yago Saez, Gustavo Recio, and Pedro Isasi. Performance of the most common noncryptographic hash functions. Software—Practice and Experience, 44(6):681– 698, June 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Eberhard:2007:MOC

[ET07] John Eberhard and Anand Tripathi. Mechanisms for object caching in distributed applications using Java RMI. Software —Practice and Experience, 37(8):799–831, July 10, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Engmann:1989:GFP

[EV89] Rolf Engmann and Frans Van Hoeve. Generation of file processing programs based on JSP. Software —Practice and Experience, 19(4):393–409, April 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Evans:1971:IAS

[Eva71] J. B. Evans. Installing an ASP system. Software—Practice and Experience, 1(1):51–60, January/ March 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Eve:1973:BRB

[Eve73]  J. Eve.  Book review:

An introduction to logical design of digital circuits, C. M. Reeves, Cambridge University Press, 1972. No. of pages: 192. Price: £1.60. Software— Practice and Experience, 3(2):191–192, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Eichelberger:2004:OOP

[EvG04] Holger Eichelberger and Jürgen Wolff von Gudenberg. Object-oriented processing of Java source code. Software—Practice and Experience, 34(12): 1157–1185, October 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

ElMotaki:2022:MCB

[EYG22] Saloua El Motaki, Ali Yahyaouy, and Hamid Gualous. Modeling the correlation between the workload and the power consumed by a server using stochastic and nonparametric approaches. Software—Practice and Experience, 52(10):2177– 2190, October 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Faustino:2022:DBS

[FAA^{+}22] João Faustino, Daniel Adriano, Ricardo Amaro, Rubén Pereira, and Miguel Mira da Silva. DevOps benefits: a systematic literature review. Software—Practice and Experience, 52 (9):1905–1926, September 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fairbairn:1987:MFF

[Fai87] Jon Fairbairn. Making form follow function: An exercise in functional programming style. Software —Practice and Experience, 17(6):379–386, June 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Farr:1974:VSI

[Far74] G. L. Farr. Virtual storage on ICL 1900’s. Software —Practice and Experience, 4(2):125–128, April/June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Farnum:1988:CSF

[Far88] Charles Farnum. Compiler support for floating-point computation. Software— Practice and Experience, 18(7):701–709, July 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Favero:2007:SPY

[Fav07] E. L. Favero. The simple and powerful yfx operator precedence parser. Software—Practice and Experience, 37(14):1451–1474, November 25, 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fallah:2022:AMD

[FB22] Somayyeh Fallah and Amir Jalaly Bidgoly. Android malware detection using network traffic based on sequential deep learning models. Software—Practice and Experience, 52 (9):1987–2004, September 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fazul:2023:PPR

[FB23] Rhauani Weber Aita Fazul and Patrícia Pitthan Barcelos. PRBP: a prioritized replica balancing policy for HDFS balancer. Software—Practice and Experience, 53(3):600–630, March 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Falleri:2014:ECP

[FBB^{+}14] Jean-Rémy Falleri, Xavier Blanc, Reda Bendraou, Marcos Aurélio Almeida da Silva, and Cédric Teyton. Extended conference paper: Incremental inconsistency detection with low memory overhead. Software—Practice and Experience, 44(5):621–641, May 2014. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Ford:1979:NLM

[FBDH79] B. Ford, J. Bentley, J. J. Du Croz, and S. J. Hague. The NAG library ‘machine’. Software—Practice and Experience, 9(1):65– 72, January 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fritzsch:2023:AMD

[FBH^{+}23] Jonas Fritzsch, Justus Bogner, Markus Haug, Ana Cristina Franco da Silva, Carolin Rubner, Matthias Saft, Horst Sauer, and Stefan Wagner. Adopting microservices and DevOps in the cyber-physical systems domain: a rapid review and case study. Software—Practice and Experience, 53(3):790–810, March 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Forward:2012:MDR

[FBLS12] Andrew Forward, Omar Badreddin, Timothy C. Lethbridge, and Julian Solano. Model-driven rapid prototyping with Umple. Software—Practice and Experience, 42(7):781– 797, July 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Firth:2005:CBA

[FBMA05] Andrew Firth, Tim Bell, Amar Mukherjee, and Don Adjeroh. A comparison of BWT approaches to string pattern matching. Software —Practice and Experience, 35(13):1217–1258, November 10, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ferreira:2022:AJF

[FBV22] Fabio Ferreira, Hudson Silva Borges, and Marco Tulio Valente. On the (un- )adoption of JavaScript front-end frameworks. Software—Practice and Experience, 52(4):947–966, April 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frances:1983:LE

[FC83] Nissim Frances and Thomas A.

Cargill. Letter to the editor. Software—Practice and Experience, 13(6): 549, June 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fleisch:1998:WMC

[FC98] Brett D. Fleisch and Mark Allan A. Co. Workplace microkernel and OS: a case study. Software—Practice and Experience, 28(6):569– 591, May 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1798;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1798&
PLACEBO=IE.pdf.

Ferrer:2012:EAM

[FCA12] Javier Ferrer, Francisco Chicano, and Enrique Alba. Evolutionary algorithms for the multiobjective test data generation problem. Software—Practice and Experience, 42(11):1331–1362, November 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frantz:2021:CBI

[FCBF^{+}21] Rafael Z. Frantz, Rafael Corchuelo, Vitor Basto-Fernandes, Fernando Rosa-Sequeira, Fabricia Roos-Frantz, and José L. Arjona. A cloud-based integration platform for enterprise application integration: a Model-Driven Engineering approach. Software —Practice and Experience, 51(4):824–847, April 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fellows:1983:UFR

[FCG83] David M. Fellows, Brian P.

Cassidy, and Uday G. Gujar. Update to ‘FOR-TRAN routines with optional arguments’. Software—Practice and Experience, 13(11):1093–1094, November 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fornaro:2019:OSE

[FCO^{+}19] Claudio Fornaro, Francesco Saverio Cafagna, Giuseppe Osteria, Valentina Scotti, Francesco Perfetto, and Livio Conti. The onboard software of the EUSO-SPB pathfinder experiment. Software—Practice and Experience, 49(3):524– 539, March 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fonseca:2009:IEI

[FCR^{+}09] Nuno A. Fonseca, Vítor Santos Costa, Ricardo Rocha, Rui Camacho, and Fernando Silva. Improving the efficiency of inductive logic programming systems. Software—Practice and Experience, 39(2):189– 219, February ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fan:2018:FMA

[FCYL18] Guisheng Fan, Liqiong Chen, Huiqun Yu, and Dongmei Liu. Formally modeling and analyzing cost-aware job scheduling for cloud data center. Software—Practice and Experience, 48(9):1536–1559, September 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fan:1992:ETB

[FD92] Changguan Fan and Suzanne Wagner Dietrich. Extension table built-ins for Prolog. Software—Practice and Experience, 22(7):573– 597, July 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Foket:2020:EEJ

[FDD20] Christophe Foket, Koen De Bosschere, and Bjorn De Sutter. Effective and efficient Java-type obfuscation. Software—Practice and Experience, 50(2):136– 160, February 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fox:2004:CFA

[FDHH04] Chris Fox, Sebastian Danicic, Mark Harman, and Robert M. Hierons. Con-SIT: a fully automated conditioned program slicer. Software—Practice and Experience, 34(1):15–46, January 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Faragardi:2018:EAR

[FDN^{+}18] Hamid Reza Faragardi, Saeid Dehnavi, Thomas Nolte, Mehdi Kargahi, and Thomas Fahringer. An energy-aware resource provisioning scheme for realtime applications in a cloud data center. Software— Practice and Experience, 48(10):1734–1757, October 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Feldman:1979:MPM

[Fel79] Stuart I. Feldman. Make — a program for maintaining computer programs. Software—Practice and Experience, 9(4):255–265, April 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Feldman:1981:DAS

[Fel81] Michael B. Feldman. Data abstraction, structured programming, and the practicing programmer. Software—Practice and Experience, 11(7):697–710, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fenwick:1994:CND

[Fen94a] Peter M. Fenwick. Correction: A new data structure for cumulative frequency tables by Peter M. Fenwick – Vol. 24, No. 3, 1994, pp. 327–336. Software—Practice and Experience, 24(7): 677, July 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [Fen94b].

Fenwick:1994:NDS

[Fen94b] Peter M. Fenwick. A new data structure for cumulative frequency tables. Software—Practice and Experience, 24(3):327–336, March 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See correction [Fen94a].

Fenwick:1996:NDS

[Fen96] Peter M. Fenwick. A new data structure for cumulative probability tables: An improved frequency-to-symbol algorithm. Software—Practice and Experience, 26(4):489–490, April 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16804.

Fenwick:1998:SRT

[Fen98] Peter Fenwick. Symbol ranking text compressors:

Review and implementation. Software—Practice and Experience, 28(5):547– 559, April 25, 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1795;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1795&
PLACEBO=IE.pdf.

Fenwick:2001:FSM

[Fen01a]  Peter Fenwick.  Fast

string matching for multiple searches. Software—Practice and Experience, 31(9):815–833, July 25, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
78505028/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=78505028&PLACEBO=IE.
pdf.

Fenwick:2001:SPP

[Fen01b] Peter Fenwick. Some perils of performance prediction: a case study on pattern matching. Software—Practice and Experience, 31(9):835–843, July 25, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        78505029/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=78505029&PLACEBO=IE.
                                        pdf.

Fenwick:2002:BWC

[Fen02] Peter Fenwick. Burrows– Wheeler compression with variable length integer codes. Software—Practice and Experience, 32 (13):1307–1316, November 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        98517581/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=98517581&PLACEBO=IE.
                                        pdf.

Fenwick:2012:PCE

[Fen12]  P. M. Fenwick.  PPM

compression without escapes. Software—Practice and Experience, 42(2):255– 260, February 2012. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ferrara:2013:GSA

[Fer13] P. Ferrara. A generic static analyzer for multithreaded Java programs. Software —Practice and Experience, 43(6):663–684, June 2013.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frankowski:1980:POS

[FF80] Elaine N. Frankowski and W. R. Franta. A process oriented simulation model specification and documentation language. Software —Practice and Experience, 10(9):721–742, September 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Faustle:1996:RRC

[FFD96] S. Fäustle, M. G. Fugini, and E. Damiani. Retrieval of reusable components using functional similarity. Software—Practice and Experience, 26(5):491– 530, May 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16807.

Floch:2013:PMB

[FFF^{+}13] J. Floch, C. Frà, R. Fricke, K. Geihs, M. Wagner, J. Lorenzo, E. Soladana, S. Mehlhase, N. Paspallis, H. Rahnama, P. A. Ruiz, and U. Scholz. Playing MUSIC — building context-aware and selfadaptive mobile applications. Software—Practice and Experience, 43(3):359– 388, March 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Freire:2019:REA

[FFRF19] Daniela L. Freire, Rafael Z. Frantz, and Fabricia Roos-Frantz. Ranking enterprise application integration platforms from a performance perspective: an experience report. Software —Practice and Experience, 49(5):921–941, May 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Freire:2019:SRT

[FFRFS19] Daniela L. Freire, Rafael Z. Frantz, Fabricia Roos-Frantz, and Sandro Sawicki. Survey on the runtime systems of enterprise application integration platforms focusing on performance. Software—Practice and Experience, 49(3):341–360, March 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Firth:1996:CA

[FFW96] J. R. Firth, C. H. Forsyth, and I. C. Wand. The compilation of Ada. Software—Practice and Experience, 26(8):863–909, August 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16826.

Ferragina:2008:PSE

[FG08] P. Ferragina and A. Gulli.

A personalized search engine based on Websnippet hierarchical clustering. Software—Practice and Experience, 38(2):189– 225, February ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Farahbod:2011:CMF

[FG11] Roozbeh Farahbod and Uwe Glässer. The Core-ASM modeling framework. Software—Practice and Experience, 41(2):167– 178, February 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fabry:2014:PMA

[FG14] Johan Fabry and Daniel Galdames. PHANtom: a modern aspect language for Pharo Smalltalk. Software —Practice and Experience, 44(4):393–412, April 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Special issue on International Workshop on Smalltalk Technologies 2011.

Friedman:1997:MEP

[FGIS97] Roy Friedman, Maxim Goldin, Ayal Itzkovitz, and Assaf Schuster. MIL-LIPEDE: Easy parallel programming in available distributed environments. Software—Practice and Experience, 27(8):929– 965, August 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7326;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7326&
PLACEBO=IE.pdf.

Fabri:2000:DCC

[FGK^{+}00] Andreas Fabri, Geert-Jan Giezeman, Lutz Kettner, Stefan Schirra, and Sven Schönherr. On the design of CGAL a computational geometry algorithms library. Software—Practice and Experience, 30(11):1167–1202, September 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72517074/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72517074&PLACEBO=IE.
pdf.

Fuggetta:1993:ESD

[FGMM93] Alfonso Fuggetta, Carlo Ghezzi, Dino Mandrioli, and Angelo Morzenti. Executable specifications with data-flow diagrams. Software—Practice and Experience, 23(6):629–653, June 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frick:2000:CRC

[FGNZ00] A. Frick, G. Goos, R. Neumann, and W. Zimmermann. Construction of robust class hierarchies. Software—Practice and Experience, 30(5):481–543, April 25, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
71004432/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=71004432&PLACEBO=IE.
pdf.

Franta:1974:LCB

[FH74] W. R. Franta and P. A.

Houle. On a loose communication between dissimilar CDC6000 operating systems. Software—Practice and Experience, 4(3):231– 236, July/September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fraser:1982:EMS

[FH82a] Christopher W. Fraser and David R. Hanson. Exploiting machine-specific pointer operations in abstract machines. Software —Practice and Experience, 12(4):367–373, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fraser:1982:MIL

[FH82b] Christopher W. Fraser and David R. Hanson. A machine-independent linker. Software—Practice and Experience, 12(4):351– 366, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fraser:1991:CGI

[FH91a] Christopher W. Fraser and David R. Hanson. A code generation interface for ANSI C. Software— Practice and Experience, 21(9):963–988, September 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

                                        URL http://storage.
                                        webhop.net/documents/
                                        interface.pdf.

Fraser:1991:HCB

[FH91b] Christopher W. Fraser and Robert R. Henry. Hardcoding bottom-up code generation tables to save time and space. Software —Practice and Experience, 21(1):1–12, January 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fernandez:1992:GCA

[FH92a] Mary F. Fernandez and David R. Hanson. Garbage collection alternatives for Icon. Software—Practice and Experience, 22(8):659– 672, August 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fraser:1992:SRS

[FH92b] Christopher W. Fraser and David R. Hanson. Simple register spilling in a retargetable compiler. Software —Practice and Experience, 22(1):85–99, January 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://storage.webhop.
net/documents/spills.
pdf.

Fayad:2002:EEF

[FHB02] Mohamed Fayad, David Hamu, and Davide Brugali. Editorial: Enterprise frameworks. Software—Practice and Experience, 32(8):735–736, July 10, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
94519519/START;   http:
//www3.interscience.wiley.

                                        com/cgi-bin/fulltext?
                                        ID=94519519&PLACEBO=IE.
                                        pdf.

Fleisch:1994:MKI

[FHJ94] Brett D. Fleisch, Randall L. Hyde, and Nils Christian Juul. Mirage+: a kernel implementation of distributed shared memory on a network of personal computers. Software —Practice and Experience, 24(10):887–909, October 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fu:2018:ERS

[FHL^{+}18] Sheng-Yu Fu, Ding-Yong Hong, Yu-Ping Liu, Jan-Jan Wu, and Wei-Chung Hsu. Efficient and retargetable SIMD translation in a dynamic binary translator. Software—Practice and Experience, 48 (6):1312–1330, June 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2573.

Fischer:1992:ASG

[FHS92] Bernd Fischer, Carsten Hammer, and Werner Struckmann. ALADIN: a scanner generator for incremental programming environments. Software—Practice and Experience, 22 (11):1011–1025, November 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fernandez-Iglesias:2005:GHQ

[FIÁSLSAR05] M. J. Fernández-Iglesias, L. M. Álvarez-Sabucedo, C. Lojo-Silva, and L. E. Anido-Rifón. Generating high quality printouts from content management systems: a cost-effective approach. Software—Practice and Experience, 35 (11):1007–1026, September 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fiddian:1982:MAF

[Fid82]  N. J. Fiddian.  Meta-

assembly: a fresh assessment. Software—Practice and Experience, 12 (12):1101–1117, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fidge:1988:LIM

[Fid88] C. J. Fidge. A LISP implementation of the model for ‘communicating sequential processes’. Software —Practice and Experience, 18(10):923–943, October 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fantechi:1986:UHL

[FIL86] A. Fantechi, P. Invarardi, and N. Lijtmaer. Using high level languages for local computer network communication: a case study in Ada. Software—Practice and Experience, 16(8):701– 717, August 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Filgueiras:1998:ISM

[Fil98] Miguel Filgueiras. Implementing a symbolic music processing system. Software—Practice and Experience, 28(5):493–512, April 25, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1793;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1793&
PLACEBO=IE.pdf.

Findlay:1977:BRB

[Fin77] W. Findlay. Book review: A primer on Pascal, R. Conway, D. Gries and E. C. Zimmerman, Winthrop Publishers Inc., Cambridge, Mass., 1976. No. of pages: 433. E7-95 (paperback). Software— Practice and Experience, 7 (3):432, June 1977. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Finnie:1988:UNS

[Fin88] G. R. Finnie. Unscrambling non-sequential programs. Software—Practice and Experience, 18(1):39– 50, January 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Finkel:1997:PET

[Fin97] Raphael A. Finkel. Pulsar:

An extensible tool for monitoring large Unix sites. Software—Practice and Experience, 27(10):1163– 1176, October 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7264;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7264&
PLACEBO=IE.pdf.

Fisher:1982:SUD

[Fis82] A. J. Fisher. The syntax of user-defined dyadic operators. Software—Practice and Experience, 12(7):623– 625, July 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fisher:1983:GVV

[Fis83]  D. L. Fisher.  Global

variables versus local variables. Software—Practice and Experience, 13(5):467– 469, May 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fischer:1984:GUC

[Fis84] A. J. Fischer. Guarded and unguarded coroutines: an implementation in BCPL. Software—Practice and Experience, 14(4): 369–376, April 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fisher:1986:MPI

[Fis86a] A. J. Fisher. A multiprocessor implementation of Occam. Software— Practice and Experience, 16(10):875–892, October 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fisher:1986:NAG

[Fis86b] A. J. Fisher. A new algorithm for generating Hilbert curves. Software —Practice and Experience, 16(1):5–12, January 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fitch:1977:PLP

[Fit77] John Fitch. Profiling a large program. Software— Practice and Experience, 7 (4):511–518, July/August 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fatoohi:2003:MDA

[FJ03] R. Fatoohi and D. Jensen.

Migration of DCE applications into CORBA and SOAP environments. Software—Practice and Experience, 33(1):1–18, January 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
100019573/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=100019573&PLACEBO=
IE.pdf.

Fjellheim:1979:MDT

[Fje79] Roar A. Fjellheim. A message distribution technique and its application to network control. Software —Practice and Experience, 9(6):499–505, June 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fraser:1990:LT

[FK90] Christopher W. Fraser and Balachander Krishnamurthy. Live text. Software—Practice and Experience, 20(8):851–858, August 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fredriksson:2016:PEA

[FK16] Kimmo Fredriksson and Pekka Kilpeläinen. Practically efficient array initialization. Software—Practice and Experience, 46(4):435– 467, April 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

French:2014:PVI

[FKD14] G. W. French, J. R. Kennaway, and A. M. Day. Programs as visual, interactive documents. Software—Practice and Experience, 44(8):911–930, August 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Felber:2013:CPP

[FKL^{+}13] Pascal Felber, Peter Kropf, Lorenzo Leonini, Toan Luu, Martin Rajman, Etienne Rivière, Valerio Schiavoni, and José Valerio. CoFeed: privacypreserving Web search recommendation based on collaborative aggregation of interest feedback. Software—Practice and Experience, 43(10):1165–1184, October 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fitzgerald:2000:MOC

[FKR^{+}00] Robert Fitzgerald, Todd B.

Knoblock, Erik Ruf, Bjarne Steensgaard, and David Tarditi. Marmot: an optimizing compiler for Java. Software—Practice and Experience, 30(3):199– 232, March 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
71001774/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=71001774&PLACEBO=IE.
pdf.

Feijs:1998:RAS

[FKV98] L. Feijs, R. Krikhaar, and R. Van Ommering. A relational approach to support software architecture analysis. Software—Practice and Experience, 28(4):371– 400, April 10, 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1786;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1786&
PLACEBO=IE.pdf.

Ferrari:1975:GPS

[FL75a] Domenico Ferrari and Mark Liu. A generalpurpose software measurement tool. Software—Practice and Experience, 5(2): 181–192, April/June 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fox:1975:DAS

[FL75b] D. A. Fox and T. Lang. Development of an advanced satellite system. Software —Practice and Experience, 5(2):203–210, April/June 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Feyock:1976:SDC

[FL76] Stefan Feyock and Paul Lazarus. Syntax-directed correction of syntax errors. Software—Practice and Experience, 6(2):207– 219, April/June 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fedele:1992:TTB

[FL92] Carine Fédèle and Olivier Lecarme. Towards a toolkit for building language implementations. Software —Practice and Experience, 22(11):911–936, November 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Freeman:1994:ERM

[FL94] M. J. Freeman and P. J. Layzell. Experience realising a meta-model for wide system understanding: The global system model. Software—Practice and Experience, 24(8):703– 728, August 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fecko:2002:LLA

[FL02] Mariusz A. Fecko and Christopher M. Lott. Lessons learned from automating tests for an operations support system. Software—Practice and Experience, 32(15):1485–1506, December 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
99019922/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=99019922&PLACEBO=IE.
pdf.

Fleck:1982:VAD

[Fle82] A. C. Fleck. Verifying abstract data types with SNOBOL4. Software— Practice and Experience, 12(7):627–640, July 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fleck:1990:CSC

[Fle90] A. C. Fleck. A case study comparison of four declarative programming languages. Software—Practice and Experience, 20(1):49– 65, January 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fan:2022:TBB

[FLL^{+}22] Yongkai Fan, Xiaodong Lin, Wei Liang, Jinghan Wang, Gang Tan, Xia Lei, and Lei Jing. TraceChain: a blockchain-based scheme to protect data confidentiality and traceability. Software—Practice and Experience, 52(1):115–129, January 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Flores:1972:ID

[Flo72]  Ivan Flores.  Intrapro-

gram documentation. Software—Practice and Experience, 2(4):353–358, October/December 1972. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Florentin:1973:BRB

[Flo73] J. J. Florentin. Book review: File organisation— selected papers from File 68, an I.A.G. conference, Swets and Zeitlinger, Amsterdam, 1968. No. of pages: 396. Price: F99.50 (guilders). Software— Practice and Experience, 3(3):304, July/September 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Florentin:1974:BRB

[Flo74] J. J. Florentin. Book review: Data base management systems, Carl Cagan, Melville Publishing Co., Los Angeles, California, 1973. No. of pages:

141. Price: £7.30. Software—Practice and Experience, 4(4):405, October/ December 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Florentin:1979:BRB

[Flo79] J. J. Florentin. Book review: The Codasyl approach to data base management, T. William Olle, Wiley, Chichester, 1978. No. of pages: 287. Price: £8.75. Software—Practice and Experience, 9(7): 597, July 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Farhadi:2020:SAT

[FLPM20] Mozhdeh Farhadi, Jean-Louis Lanet, Guillaume Pierre, and Daniele Miorandi. A systematic approach toward security in Fog computing: Assets, vulnerabilities, possible countermeasures. Software—Practice and Experience, 50(6):973–997, June 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fuentes-Lorenzo:2015:RSF

[FLSCC15] Damaris Fuentes-Lorenzo, Luis Sánchez, Antonio Cuadra, and Mar Cutanda. A RESTful and semantic framework for data integration. Software—Practice and Experience, 45 (9):1161–1188, September 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fisher:1977:MHL

[FM77] Richard N. Fisher and Gordon W. McQuarrie. MPL1700: a high(er)-level microprogramming language. Software—Practice and Experience, 7(6):747– 757, November/December 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Foxley:1978:MRT

[FM78] Eric Foxley and D. J. Morgan. Monitoring the run-time activity of Algol 68-R programs. Software—Practice and Experience, 8(1):29–34, January/ February 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Forsythe:1986:CDS

[FM86] Kenneth Forsythe and Stanislaw Matwin. Copying of dynamic structures in a Pascal environment. Software—Practice and Experience, 16(4):335– 340, April 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fang:2002:JJB

[FMA02] Xiang Fang, John A.

Miller, and Jonathan Arnold. J3DV: a Javabased 3D database visualization tool. Software—Practice and Experience, 32(5):443–463, April 25, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
93513335/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=93513335&PLACEBO=IE.
pdf.

Fernandes:2018:EMF

[FMC18] Thiago Soares Fernandes, Álvaro Freitas Moreira, and Érika Cota. EPE-Mobile — a framework for early performance estimation of mobile applications. Software—Practice and Experience, 48(1):85– 104, January 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fetterly:2004:LSS

[FMNW04] Dennis Fetterly, Mark Manasse, Marc Najork, and Janet L. Wiener. A large-scale study of the evolution of Web pages. Software—Practice and Experience, 34(2):213–237, February 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fuhrer:2002:MSF

[FMPR02] Patrik Fuhrer, Ghita Kouadri Mostéfaoui, and Jacques Pasquier-Rocha. MaDVi-World: a software framework for massively distributed virtual worlds. Software—Practice and Experience, 32(7):645–668, June 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        94515733/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=94515733&PLACEBO=IE.
                                        pdf.

Finkel:2004:CLT

[FMT04] Raphael Finkel, Victor W. Marek, and Mirosław Truszczyński. Constraint Lingo: towards highlevel constraint programming. Software—Practice and Experience, 34 (15):1481–1504, December 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fitch:1977:ILH

[FN77] J. P. Fitch and A. C.

Norman. Implementing LISP in a high-level language. Software—Practice and Experience, 7(6):713– 725, November/December 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fayyoumi:2010:SSD

[FO10] Ebaa Fayyoumi and B. John Oommen. A survey on statistical disclosure control and micro-aggregation techniques for secure statistical databases. Software—Practice and Experience, 40(12):1161–1188, November 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fong:1985:NCT

[Fon85] Kirby W. Fong. The NM-FECC Cray Time-Sharing System. Software—Practice and Experience, 15(1): 87–103, January 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Foolery:1972:CR

[Foo72] April Foolery. Computer recreations. Software— Practice and Experience, 2(2):187–189, April 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Forrest:1972:BRB

[For72] A. R. Forrest. Book review: Advanced computer graphics—economics techniques and applications, Eds.: R. D. Parslow and R. Elliot Green, Plenum Press, London and New York, 1971. No. of pages: 1,230 Price: £15.00. Software—Practice and Experience, 2(3):308, July 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Foster:1986:SCM

[Fos86] David G. Foster. Separate compilation in a Modula-2 compiler. Software—Practice and Experience, 16(2): 101–106, February 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Foster:1989:IDS

[Fos89] Ian T. Foster. Implementation of a declarative statetransition system. Software —Practice and Experience, 19(4):351–370, April 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fowler:1990:CM

[Fow90] Glenn Fowler. A case for make. Software—Practice and Experience, 20 (S1):S35–S46, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Foxley:1978:PAT

[Fox78] Eric Foxley. Programming aids for the teaching of programming. Software— Practice and Experience, 8 (2):213–225, March/April 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Foxley:1979:BRB

[Fox79] Eric Foxley. Book review:

Algol 68: A first and second course. Andrew D. McGettrick, C.U.P., Cambridge, 1978. No. of pages: 360. Price: £12.50 (hard cover), £5.50 (paperback). Software—Practice and Experience, 9(5):425, May 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Foxley:1987:MLT

[Fox87] Eric Foxley. Music — a language for typesetting music scores. Software—Practice and Experience, 17(8):485– 502, August 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frost:1982:FGN

[FP82] R. A. Frost and M. M. Peterson. A function for generating nearly balanced binary search trees from sets of non-random keys. Software—Practice and Experience, 12(2):163– 168, February 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fenwick:1997:IEI

[FP97] James B. Fenwick, Jr. and Lori L. Pollock. Issues and experiences in implementing a distributed tuplespace. Software— Practice and Experience, 27(10):1199–1232, October 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7266;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7266&
PLACEBO=IE.pdf.

Friedman:2015:GDT

[FP15] Roy Friedman and Amit Portnoy. A generic decentralized trust management framework. Software —Practice and Experience, 45(4):435–454, April 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frias:2018:PTT

[FPAF18] Bernardo Villalba Frías, Luigi Palopoli, Luca Abeni, and Daniele Fontanelli. The PROSIT tool: Toward the optimal design of probabilistic soft real-time systems. Software—Practice and Experience, 48 (11):1940–1967, November 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fuentes:2007:SDC

[FPT07] L. Fuentes, M. Pinto, and J. M. Troya. Supporting the development of CAM/ DAOP applications: an integrated development process. Software—Practice and Experience, 37(1):21– 64, January 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Feng:1978:SSC

[FR78] David Y. Feng and Richard F. Riesenfeld. A symbolic system for computeraided development of surface interpolants. Software—Practice and Experience, 8(4):461–481, July/ August 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fruchterman:1991:GDF

[FR91] Thomas M. J. Fruchterman and Edward M. Reingold. Graph drawing by forcedirected placement. Software—Practice and Experience, 21(11):1129–1164, November 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ficco:2009:HPS

[FR09] Massimo Ficco and Stefano Russo. A hybrid positioning system for technology-independent location-aware computing. Software—Practice and Experience, 39(13): 1095–1125, September 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frailey:1974:NDT

[Fra74]  Dennis J. Frailey.  A

note on deleting trailing blanks. Software—Practice and Experience, 4(2): 189–191, April/June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frailey:1975:DSR

[Fra75] Dennis J. Frailey. DSOS: a skeletal, real-time, minicomputer operating system. Software—Practice and Experience, 5(1):5–18, January/March 1975. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fraser:1979:CPC

[Fra79] Christopher W. Fraser. A compact, portable CRTbased text editor. Software —Practice and Experience, 9(2):121–125, February 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fraser:1980:MPV

[Fra80] Christopher W. Fraser.

Maintaining program variants by merging editor scripts. Software—Practice and Experience, 10(10): 817–821, October 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fraser:1982:PTE

[Fra82] Christopher W. Fraser. A programmable text editor. Software—Practice and Experience, 12(3):241– 250, March 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Franz:1993:EOS

[Fra93] Michael Franz. Emulating an operating system on top of another. Software —Practice and Experience, 23(6):677–692, June 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frank:1999:EBR

[Fra99]  Lars Frank.  Evalua-

tion of the basic remote backup and replication methods for high availability databases. Software— Practice and Experience, 29(15):1339–1353, December 25, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
68501288/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=68501288&PLACEBO=IE.
pdf.

Fraser:2006:IDI

[Fra06] Christopher W. Fraser. An instruction for direct interpretation of LZ77compressed programs. Software—Practice and Experience, 36(4):397–411, April 10, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frantz:2019:ROS

[FRBRF19] Rafael Z. Frantz, Matheus H. Rehbein, Rodolfo Berlezi, and Fabricia Roos-Frantz. Ranking open source application integration frameworks based on maintainability metrics: a review of five-year evolution. Software—Practice and Experience, 49(10):1531–1549, October 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Freeman:1978:SDRa

[Fre78a] Peter Freeman. Software design representation: a case study. Software— Practice and Experience, 8(5):501–511, September/ October 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Freeman:1978:SDRb

[Fre78b]  Peter Freeman.  Soft-

ware design representation: Analysis and improvement. Software— Practice and Experience, 8(5):513–528, September/ October 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Freak:1981:FPT

[Fre81] R. A. Freak. A Fortran to Pascal translator. Software —Practice and Experience, 11(7):717–732, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fdez-Riverola:2012:JAF

[FRGPLF^{+}12] F. Fdez-Riverola, D. Glez-Peña, H. López-Fernández, M. Reboiro-Jato, and J. R. Méndez. A Java application framework for scientific software development. Software—Practice and Experience, 42(8): 1015–1036, August 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frieder:1992:PDD

[Fri92]  O. Frieder.  A paral-

lel database-driven protocol verification system prototype. Software—Practice and Experience, 22(3):245– 264, March 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Froggatt:1981:ID

[Fro81] T. J. Froggatt. Integer division. Software—Practice and Experience, 11(9): 1000, September 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frost:1993:GAG

[Fro93] R. A. Frost. Guarded attribute grammars. Software—Practice and Experience, 23(10):1139–1156, October 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frailey:1973:LE

[FS73] Dennis J. Frailey and Paul A. Samet. Letters to the editor. Software—Practice and Experience, 3(4):403–404, October/December 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fisher:1981:SEE

[FS81] Paul S. Fisher and Jacob Slonim. Software engineering: An example of misuse. Software—Practice and Experience, 11(6):533– 539, June 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ferstl:1982:DSC

[FS82] O. K. Ferstl and E. J.

Sinz. Designing structured COBOL programs. Software—Practice and Experience, 12(5):455–474, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fayad:2011:CPP

[FS11] M. E. Fayad and Shivanshu K. Singh. Call for papers: Pattern Languages: Addressing the Challenges (PLAC). Software—Practice and Experience, 41 (1):129–130, January 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fayad:2013:EPL

[FS13] Mohamed E. Fayad and Shivanshu K. Singh. Editorial: Pattern languages: addressing the challenges. Software—Practice and Experience, 43(7):743–745, July 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fong:2008:RXD

[FSC08] Joseph Fong, Herbert Shiu, and Davy Cheung. A relational-XML data warehouse for data aggregation with SQL and XQuery. Software—Practice and Experience, 38(11):1183– 1213, September ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Freire:2021:MPM

[FSC^{+}21] Augusto Flávio A. A. Freire, Américo Falcone Sampaio, Luis Heustakio L. Carvalho, Otávio Medeiros, and Nabor C. Mendonça. Migrating production monolithic systems to microservices using aspect oriented programming. Software—Practice and Experience, 51(6):1280–1307, June 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Furuta:1991:YPB

[FSO91] Richard Furuta, P. David Stotts, and Jefferson Ogata. Ytracc: a parse browser for Yacc grammers. Software —Practice and Experience, 21(2):119–132, February 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ferreira:2011:UED

[FSR11] Jennifer Ferreira, Helen Sharp, and Hugh Robinson. User experience design and agile development: managing cooperation through articulation work. Software—Practice and Experience, 41(9):963– 974, August 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frantz:2022:AMP

[FSRF^{+}22] Rafael Z. Frantz, Sandro Sawicki, Fabricia Roos-Frantz, Fábio P. Basso, Benjamim Zucoloto, and Raquel M. Pillat. On the analysis of makespan and performance of the taskbased execution model for enterprise application integration platforms: an empirical study. Software—Practice and Experience, 52(7):1717–1735, July 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Finkel:1999:EUS

[FSS99] Raphael A. Finkel, Brian Sturgill, and Harlan Stenn. Experience with a Unix system-administration tool. Software—Practice and Experience, 29(11):953– 971, September 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=63501202;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=63501202&PLACEBO=IE.
                                        pdf.

Frank:1979:DMO

[FT79a] G. R. Frank and C. J. Theaker. The design of the MUSS operating system. Software—Practice and Experience, 9(8):599– 620, August 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frank:1979:MUI

[FT79b] G. R. Frank and C. J. Theaker. MUSS — the user interface. Software— Practice and Experience, 9 (8):621–631, August 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fuentes:2001:CDC

[FT01] Lidia Fuentes and José María Troya. Coordinating distributed components on the web: an integrated development environment. Software—Practice and Experience, 31(3):209–233, March 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
76507639/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=76507639&PLACEBO=IE.
pdf.

Faust:2003:SPL

[FV03] D. Faust and C. Verhoef. Software product line migration and deployment. Software—Practice and Experience, 33 (10):933–955, August 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Faragardi:2018:EPS

[FVF^{+}18] Hamid Reza Faragardi, Maryam Vahabi, Hossein Fotouhi, Thomas Nolte, and Thomas Fahringer. An efficient placement of sinks and SDN controller nodes for optimizing the design cost of industrial IoT systems. Software—Practice and Experience, 48(10): 1893–1919, October 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Friedman:1978:UCS

[FW78] Daniel P. Friedman and David S. Wise. Unbounded computational structures. Software— Practice and Experience, 8 (4):407–416, July/August 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Footit:1974:UWM

[FWS74] R. M. Footit and C. Whitby-Strevens. The University of Warwick Modular One operating system. Software—Practice and Experience, 4(4):309–339, October/December 1974. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Flater:1993:ECP

[FYP93] David W. Flater, Yelena Yesha, and E. K. Park. Extensions to the C programming language for enhanced fault detection. Software—Practice and Experience, 23(6):617– 628, June 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Frank:1998:SAP

[FZ98] Lars Frank and Torben U. Zahle. Semantic ACID properties in multidatabases using remote procedure calls and update propagations. Software—Practice and Experience, 28(1):77–98, January 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61001668;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=61001668&PLACEBO=IE.
pdf.

Ferrari:2012:ARD

[FZ12] Luca Ferrari and Haibin Zhu. Autonomous role discovery for collaborating agents. Software—Practice and Experience, 42(6):707– 731, June 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fu:2017:RRA

[FZS^{+}17] Min Fu, Liming Zhu, Daniel Sun, Anna Liu, Len Bass, and Qinghua Lu. Runtime recovery actions selection for sporadic operations on public cloud. Software—Practice and Experience, 47(2):223– 248, February 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fu:2019:CBP

[FZW19] Jiaojiao Fu, Yangfan Zhou, and Xin Wang. Componentbased permission management of Android applications. Software—Practice and Experience, 49 (9):1402–1418, September 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gedik:2012:MBF

[GA12] Buğra Gedik and Henrique Andrade. A modelbased framework for building extensible, high performance stream processing middleware and programming language for IBM InfoSphere Streams. Software—Practice and Experience, 42(11):1363–1391, November 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ghobaei-Arani:2022:DIS

[GAAA22] Mostafa Ghobaei-Arani, Shiva Asadianfam, and Ahad Abolfathi. Deploying IoT services on the fog infrastructure: a graph partitioning-based approach. Software—Practice and Experience, 52 (9):1967–1986, September 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gruber:2023:LLL

[GAB^{+}23] Bernhard Manfred Gruber, Guilherme Amadio, Jakob Blomer, Alexander Matthes, René Widera, and Michael Bussmann. LLAMA: the low-level abstraction for memory access. Software—Practice and Experience, 53(1):115– 141, January 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gedik:2009:TSD

[GAF^{+}09] Buğra Gedik, Henrique Andrade, Andy Frenkiel, Wim De Pauw, Michael Pfeifer, Paul Allen, Norman Cohen, and Kun-Lung Wu. Tools and strategies for debugging distributed stream processing applications. Software—Practice and Experience, 39(16): 1347–1376, November ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Goulas:2005:SGB

[GAH05] G. Goulas, P. Alefragis, and E. Housos. SchedSP: a Grid-based application service provider of scheduling solutions. Software— Practice and Experience, 35(12):1195–1216, October 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gait:1982:AEC

[Gai82a] Jason Gait. An algorithm for an efficient command interpreter. Software— Practice and Experience, 12(10):981–982, October 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gait:1982:ASP

[Gai82b] Jason Gait. Algorithms for a self-prompting keyboard monitor. Software —Practice and Experience, 12(4):393–394, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gait:1985:DCP

[Gai85] Jason Gait. A debugger for concurrent programs. Software—Practice and Experience, 15(6):539– 554, June 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gait:1986:PEC

[Gai86]  Jason Gait.  A probe

effect in concurrent programs. Software—Practice and Experience, 16(3):225– 233, March 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Galkowski:1979:LEC

[Gal79] J. T. Galkowski. Letter to the Editor: Computer programming and the human thought process. Software —Practice and Experience, 9(9):792, September 1979.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gangopadhyay:1982:FMG

[Gan82] Dipayan Gangopadhyay. A framework for modelling graphic interactions. Software—Practice and Experience, 12(2):141–151, February 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grutzmacher:2023:UGM

[GAQO23] Thomas Grützmacher, Hartwig Anzt, and Enrique S. Quintana-Ortí. Using Ginkgo’s memory accessor for improving the accuracy of memory-bound low precision BLAS. Software— Practice and Experience, 53(1):81–98, January 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Garratt:1986:BRB

[Gar86] P. W. Garratt. Book review: A Practical Handbook for Software Development, N. D. Birrell and M. A. Ould, Cambridge University Press, 1985. No. of pages: 259. Price: £20.00. Software—Practice and Experience, 16 (11):1056–1057, November 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gardiner:1996:AAW

[Gar96] Charles W. Gardiner. ASN EZE: An analgesic for writers of ASN.1 applications. Software— Practice and Experience, 26(10):1087–1096, October 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16763.

Ghobaei-Arani:2018:MFO

[GARSR18] Mostafa Ghobaei-Arani, Ali.A Rahmanian, Alireza Souri, and Amir Masoud Rahmani. A moth-flame optimization algorithm for web service composition in cloud computing: Simulation and verification. Software—Practice and Experience, 48(10):1865–1892, October 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gauthier:1995:APC

[Gau95]  Michel Gauthier.  The

avatars of a package for calendars in Ada. Software —Practice and Experience, 25(4):403–427, April 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gay:1980:BMA

[Gay80] A. R. Gay. Benchmarking a multi-access system. Software—Practice and Experience, 10(1):45–55, January 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gudes:1987:GTD

[GB87] Ehud Gudes and Gilad Bracha. GCI: a tool for developing interactive CAD user interfaces. Software —Practice and Experience, 17(11):783–799, November 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gimenes:2002:EFW

[GB02] Itana M. S. Gimenes and Leonor Barroca. Enterprise frameworks for workflow management systems. Software—Practice and Experience, 32(8):755–769, July 10, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
94519516/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=94519516&PLACEBO=IE.
pdf.

Garg:2013:ECP

[GB13] Saurabh Kumar Garg and Rajkumar Buyya. Extended conference papers: An environment for modeling and simulation of message-passing parallel applications for cloud computing. Software—Practice and Experience, 43 (11):1359–1375, November 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grozev:2014:ICA

[GB14] Nikolay Grozev and Rajkumar Buyya. Inter-Cloud architectures and application brokering: taxonomy and survey. Software—Practice and Experience, 44(3):369– 390, March 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ghedass:2023:ACI

[GB23] Fedia Ghedass and Faouzi Ben Charrada. Autonomic computing and incremental learning for the management of big services. Software—Practice and Experience, 53(7): 1594–1628, July 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Geihs:2009:CSA

[GBE^{+}09] K. Geihs, P. Barone, F. Eliassen, J. Floch, R. Fricke, E. Gjorven, S. Hallsteinsen, G. Horn, M. U. Khan, A. Mamelli, G. A. Papadopoulos, N. Paspallis, R. Reichle, and E. Stav. A comprehensive solution for applicationlevel adaptation. Software —Practice and Experience, 39(4):385–422, March 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Garcia:2014:ECP

[GBG^{+}14] Miguel Garcia, Alysson Bessani, Ilir Gashi, Nuno Neves, and Rafael Obelheiro. Extended conference paper: Analysis of operating system diversity for intrusion tolerance. Software —Practice and Experience, 44(6):735–770, June 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gehani:1984:CPA

[GC84] Narain H. Gehani and T. A. Cargill. Concurrent programming in the Ada(R) language: The polling bias. Software— Practice and Experience, 14(5):413–427, May 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gallego:2020:TAB

[GC20] Fernando O. Gallego and Rafael Corchuelo. Torii: an aspect-based sentiment analysis system that can mine conditions. Software —Practice and Experience, 50(1):47–64, January 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gonzalez-Castano:2001:JCV

[GCARPC^{+}01] F. J. González-Castaño, L. Anido-Rifón, J. M. Pousada-Carballo, P. S. Rodríguez-Hernández, and R. López-Gómez. A Java/ CORBA virtual machine architecture for remote execution of optimization solvers in heterogeneous networks. Software—Practice and Experience, 31(1): 1–16, January 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        76502341/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=76502341&PLACEBO=IE.
                                        pdf.

Goonasekera:2015:LAS

[GCF15] Nuwan Goonasekera, William Caelli, and Colin Fidge. LibVM: an architecture for shared library sandboxing. Software—Practice and Experience, 45 (12):1597–1617, December 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gray:2002:DAP

[GCK^{+}02] Robert S. Gray, George Cybenko, David Kotz, Ronald A. Peterson, and Daniela Rus. D’Agents: Applications and performance of a mobile-agent system. Software—Practice and Experience, 32(6): 543–573, May 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
93513614/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=93513614&PLACEBO=IE.
pdf.

Ghanam:2011:E

[GCM11] Yaser Ghanam, Kendra Cooper, and Frank Maurer. Editorial. Software —Practice and Experience, 41(8):867–869, July 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Georges:2004:JPR

[GCRD04] A. Georges, M. Christiaens, M. Ronsse, and K. De Bosschere. JaRec: a portable record/replay environment for multithreaded Java applications. Software—Practice and Experience, 34(6):523– 547, May 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guo:2024:CBT

[GCZ^{+}24] Yi Guo, Yu Chai, Lehuan Zhang, Hui Li, Mengzhi Luo, and Shikai Guo. Context-based transfer learning for low resource code summarization. Software —Practice and Experience, 54(3):465–482, March 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Garcia:2023:LAS

[GDBP23] Nicolás Pozas García, Francisco Durán, Katia Moreno Berrocal, and Ernesto Pimentel. Locationaware scalable service composition. Special focus section: Benchmarking, Experimentation Tools, and Reproducible Practices for Data-Intensive Systems from Edge to Cloud: Software: Practice and Experience, 53(12):2408–2429, December 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Garcia:2018:RPS

[GdCF^{+}18] Álvaro López García, Enol Fernández del Castillo, Pablo Orviz Fernández, Isabel Campos Plasencia, and Jesús Marco de Lucas. Resource provisioning in Science Clouds: Requirements and challenges. Software—Practice and Experience, 48 (3):486–498, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2544.

Gupta:2017:ITM

[GDGB17] Harshit Gupta, Amir Vahid Dastjerdi, Soumya K. Ghosh, and Rajkumar Buyya. iFogSim: a toolkit for modeling and simulation of resource management techniques in the Internet of Things, Edge and Fog computing environments. Software—Practice and Experience, 47 (9):1275–1296, September 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gui:2013:TAF

[GDH13] Ning Gui, Vincenzo De Florio, and Tom Holvoet. Transformer: an adaptation framework supporting contextual adaptation behavior composition. Software—Practice and Experience, 43(8):937–967, August 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Garcia:2004:AOO

[GdLC04] Alessandro F. Garcia, Carlos J. P. de Lucena, and Donald D. Cowan. Agents in object-oriented software engineering. Software— Practice and Experience, 34(5):489–521, April 25, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guedria:2020:RSD

[GDRV20] Soulaimane Guedria, Noël De Palma, Félix Renard, and Nicolas Vuillerme. R2D2: a scalable deep learning toolkit for medical imaging segmentation. Software—Practice and Experience, 50(10): 1966–1985, October 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guo:2020:PDD

[GDW^{+}20] Junqi Guo, Yazhu Dai, Chixiang Wang, Hao Wu, Tianyou Xu, and Kai Lin. A physiological data-driven model for learners’ cognitive load detection using HRV-PRV feature fusion and optimized XG-Boost classification. Software—Practice and Experience, 50(11):2046–2064, November 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Goenawan:2023:DSB

[GDW^{+}23] Ivan Hendy Goenawan, Zhihui Du, Chao Wu, Yankui Sun, Jianyan Wei, and David A. Bader. Dynamics signature based anomaly detection. Software—Practice and Experience, 53(1):160–175, January 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gedik:2014:GWS

[Ged14]  Buğra Gedik.  Generic

windowing support for extensible stream processing systems. Software—Practice and Experience, 44 (9):1105–1128, September 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guerraoui:2000:EOG

[GEF^{+}00] Rachid Guerraoui, Patrick Eugster, Pascal Felber, Benoît Garbinato, and Karim Mazouni. Experiences with object group systems. Software— Practice and Experience, 30(12):1375–1404, October 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
73000926/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=73000926&PLACEBO=IE.
pdf.

Gehani:1982:SFI

[Geh82] Narain H. Gehani. Specifications: Formal and informal — a case study. Software—Practice and Experience, 12(5):433–444, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gehani:1983:EFS

[Geh83] Narain H. Gehani. An electronic form system: an experience in prototyping. Software—Practice and Experience, 13(6):479– 486, June 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gehani:1985:ADT

[Geh85] Narain H. Gehani. Ada’s derived types and units of measure. Software—Practice and Experience, 15(6): 555–569, June 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gehani:1990:MPC

[Geh90] Narain H. Gehani. Message passing in Concurrent C: Synchronous versus asynchronous. Software —Practice and Experience, 20(6):571–592, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gehani:1992:ECC

[Geh92] Narain H. Gehani. Exceptional C or C with exceptions. Software—Practice and Experience, 22(10):

827–848, October 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gary:2011:AMO

[GEI^{+}11] Kevin Gary, Andinet Enquobahrie, Luis Ibanez, Patrick Cheng, Ziv Yaniv, Kevin Cleary, Shylaja Kokoori, Benjamin Muffih, and John Heidenreich. Agile methods for open source safety-critical software. Software—Practice and Experience, 41(9):945– 962, August 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Geller:1975:DOL

[Gel75] Dennis P. Geller. Debugging other languages in APL. Software—Practice and Experience, 5(2):139– 145, April/June 1975. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gentleman:1981:MPB

[Gen81] W. Morven Gentleman.

Message passing between sequential processes: the reply primitive and the administrator concept. Software—Practice and Experience, 11(5):435–466, May 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

George:1977:EFD

[Geo77] Richard George. Eliminate flowchart drawings. Software—Practice and Experience, 7(6):727–732, November/December 1977. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gerritse:1982:NEP

[Ger82] J. J. Gerritse. A note on extension of Pascal by coroutines. Software—Practice and Experience, 12 (12):1163–1164, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gujar:1978:ICJ

[GF78] Uday G. Gujar and David M. Fellows. Implementation of a computer job scheduling algorithm: A case history. Software—Practice and Experience, 8(5):529– 542, September/October 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gujar:1980:APE

[GF80] Uday G. Gujar and J. Anthony Fitzgerald. An algorithm for programming electrostatic plotter type devices. Software— Practice and Experience, 10(10):823–833, October 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gujar:1981:FRO

[GF81] Uday G. Gujar and David M.

Fellows. FORTRAN routines with optional arguments. Software—Practice and Experience, 11(2):187– 193, February 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ganapathi:1984:ALI

[GF84] Mahadevan Ganapathi and Charles N. Fischer. Attributed linear intermediate representations for retargetable code generators. Software—Practice and Experience, 14(4):347– 364, April 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gervais:2011:ETB

[GF11] Frédéric Gervais and Benoît Fraikin. Editorials: Tool building in formal methods. Software—Practice and Experience, 41(2):131– 132, February 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ghosh:2005:MTA

[GFS^{+}05] Sudipto Ghosh, Robert B.

France, Devon M. Simmonds, Abhijit Bare, Brahmila Kamalakar, Roopashree P. Shankar, Gagan Tandon, Peter Vile, and Shuxin Yin.

A middleware-transparent approach to developing distributed applications. Software—Practice and Experience, 35(12):1131–1154, October 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Goodwin:1996:OPC

[GG96] Jeremy P. Goodwin and Edward F. Gehringer. Optimizing procedure calls in block-structured languages. Software—Practice and Experience, 26 (12):1385–1414, December 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16780.

Guo:2008:SDP

[GG08] Hai-Feng Guo and Gopal Gupta. Simplifying dynamic programming via mode-directed tabling. Software—Practice and Experience, 38(1):75–94, January ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gary:1972:EFC

[GH72] John Gary and Richard Helgason. An extension of FORTRAN containing finite difference operators. Software—Practice and Experience, 2(4):321– 336, October/December 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Griss:1981:PLC

[GH81] Martin L. Griss and Anthony C. Hearn. A portable LISP compiler. Software —Practice and Experience, 11(6):541–605, June 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gunn:1984:PPI

[GH84] Hamish I. E. Gunn and David M. Harland. Polymorphic programming II. an orthogonal tagged high level architecture abstract machine. Software—Practice and Experience, 14 (11):1021–1046, November 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gupta:1993:EPW

[GH93] Aloke Gupta and Wen-Mei W. Hwu. An execution profiler for windoworiented applications. Software—Practice and Experience, 23(5):487–510, May 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grundy:2002:EPS

[GH02] J. Grundy and J. Hosking.

Engineering plug-in software components to support collaborative work. Software—Practice and Experience, 32(10):983– 1013, August 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
96516182/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=96516182&PLACEBO=IE.
pdf.

Gachet:2003:JBS

[GH03] Alexandre Gachet and Pius Haettenschwiler. A Jinibased software framework for developing distributed cooperative decision support systems. Software —Practice and Experience, 33(3):221–258, March 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Greer:2009:CPA

[GH09] Des Greer and Yann Hamon. Call for papers: ‘Agile Software Development’. Software—Practice and Experience, 39(9):865– 866, June 25, 2009. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Greer:2011:EAS

[GH11] Des Greer and Yann Hamon. Editorial: Agile Software Development. Software—Practice and Experience, 41(9):943–944, August 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gholamshahi:2019:SCI

[GH19]  Shabnam  Gholamshahi

and Seyed Mohammad Hossein Hasheminejad. Software component identification and selection: a research review. Software —Practice and Experience, 49(1):40–69, January 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gold:2005:UPS

[GHBH05] N. E. Gold, M. Harman, D. Binkley, and R. M. Hierons. Unifying program slicing and concept assignment for higher-level executable source code extraction. Software—Practice and Experience, 35(10): 977–1006, August 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grundy:2007:EDA

[GHC^{+}07] John Grundy, John Hosking, Shuping Cao, Denjin Zhao, Nianping Zhu, Ewan Tempero, and Hermann Stoeckle. Experiences developing architectures for realizing thin-client diagram editing tools. Software—Practice and Experience, 37(12):1245–1283, October 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grundy:1996:SFC

[GHM96] John C. Grundy, John G. Hosking, and Warwick B. Mugridge. Supporting flexible consistency management via discrete change description propagation. Software—Practice and Experience, 26(9):1053– 1083, September 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16835.

Gonzalez:2006:SEW

[GHM^{+}06] E. J. González, A. F. Hamilton, L. Moreno, R. L. Marichal, and V. Muñoz. Software experience when using ontologies in a multi-agent system for automated planning and scheduling. Software— Practice and Experience, 36(7):667–688, June 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Glass:2001:LHL

[GIF01] S. Glass, D. Ince, and E. Fergus. Llun — a high-level debugger for generated parsers. Software—Practice and Experience, 31(10):983–1001, August 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
82002142/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=82002142&PLACEBO=IE.
pdf.

Grune:1988:PFL

[GJ88] Dick Grune and Ceriel J. H.

Jacobs. A programmerfriendly LL(1) parser generator. Software—Practice and Experience, 18(1):29– 38, January 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gupta:1993:LSV

[GJ93] Deepak Gupta and Pankaj Jalote. On-line software version change using state transfer between processes. Software—Practice and Experience, 23(9):949– 964, September 1993. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gunter:2000:PDC

[GJ00] Carl A. Gunter and Trevor Jim. Policy-directed certificate retrieval. Software—Practice and Experience, 30(15):1609–1640, December 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        73501272/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=73501272&PLACEBO=IE.
                                        pdf.

Gu:2022:OMT

[GJ22] Xiaoqing Gu and Yizhang Jiang. Online multitarget tracking system for autonomous vehicles using discriminative dictionary learning with embedded auto-encoder algorithm. Software—Practice and Experience, 52(8): 1785–1801, August 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gaglianello:1986:CM

[GK86] Robert D. Gaglianello and Howard P. Katseff. Communications in Meglos. Software—Practice and Experience, 16(10): 945–963, October 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Goldschmidt:2008:CKS

[GK08] David E. Goldschmidt and Mukkai Krishnamoorthy.

Comparing keyword search to semantic search: a case study in solving crossword puzzles using the Google^{T M} API. Software —Practice and Experience, 38(4):417–445, April 10, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Garbervetsky:2014:EDT

[GK14] Diego Garbervetsky and Sunghun Kim. Editorial: Developing tools as plugins: TOPI 2012 special issue. Software—Practice and Experience, 44(3):259– 260, March 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grabowski:2021:AAP

[GK21] Szymon Grabowski and Tomasz M. Kowalski. Algorithms for all-pairs Hamming distance based similarity. Software—Practice and Experience, 51(7): 1580–1590, July 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Go:2016:ERP

[GKBK16] Kyungmin Go, Sungwon Kang, Jongmoon Baik, and Myungchul Kim. Experience report: Pairwise testing for systems with data derived from real-valued variable inputs. Software—Practice and Experience, 46(3):381–403, March 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gammage:1987:RR

[GKC87] N. D. Gammage, R. F. Kamel, and L. M. Casey. Remote rendezvous. Software—Practice and Experience, 17(10):741–755, October 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Graef:1979:HDI

[GKLM79] Norwin Graef, Horst Kretschmar, Klaus-Peter Löhr, and Bernt Morawetz. How to design and implement small time-sharing systems using Concurrent Pascal. Software—Practice and Experience, 9(1):17– 24, January 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Graham:1983:EPM

[GKM83] Susan L. Graham, Peter B. Kessler, and Marshall K. McKusick. An execution profiler for modular programs. Software—Practice and Experience, 13(8):671– 685, August 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gettys:1990:XWS

[GKM90] Jim Gettys, Philip L. Karlton, and Scott McGregor. The X Window System, Version 11. Software— Practice and Experience, 20(S2):S35–S67, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Giegerich:2003:EIL

[GKS03] R. Giegerich, S. Kurtz, and J. Stoye. Efficient implementation of lazy suffix trees. Software—Practice and Experience, 33 (11):1035–1049, September 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Groote:2011:EDM

[GKS^{+}11] J. F. Groote, J. J. A.

Keiren, F. P. M. Stappers, J. W. Wesselink, and T. A. C. Willemse. Experiences in developing the mCRL2 toolset. Software —Practice and Experience, 41(2):143–153, February 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gill:2022:QCT

[GKS^{+}22] Sukhpal Singh Gill, Adarsh Kumar, Harvinder Singh, Manmeet Singh, Kamalpreet Kaur, Muhammad Usman, and Rajkumar Buyya. Quantum computing: a taxonomy, systematic review and future directions. Software—Practice and Experience, 52(1): 66–114, January 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gao:2011:CSM

[GKWS11] Kehan Gao, Taghi M. Khoshgoftaar, Huanjing Wang, and Naeem Seliya. Choosing software metrics for defect prediction: an investigation on feature selection techniques. Software —Practice and Experience, 41(5):579–606, May 25, 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gillett:1978:ESL

[GL78] Will D. Gillett and Sandra Leach. Embedding semantics in LR parser tables. Software—Practice and Experience, 8(6):731– 753, November/December 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

George:1985:OCM

[GL85] Geoffrey George and James F. Leathrum. Orthogonality of concerns in module closure. Software—Practice and Experience, 15(2):119– 130, February 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gehani:1997:OTM

[GL97] Narain Gehani and Daniel F.

Lieuwen. Ode triggers: Monitoring the stock market. Software—Practice and Experience, 27(8):905– 927, August 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7325;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7325&
PLACEBO=IE.pdf.

Garcia:2005:MCL

[GL05] Ronald Garcia and Andrew Lumsdaine. Multi-Array: a C++ library for generic programming with arrays. Software—Practice and Experience, 35(2):159– 188, February 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Glass:1982:RTC

[Gla82] Robert L. Glass. Real-time checkout: the “source error first” approach. Software —Practice and Experience, 12(1):77–83, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gong:2021:KDW

[GLD^{+}21] Wenwen Gong, Chao Lv, Yucong Duan, Zengguang Liu, Mohammad R. Khosravi, Lianyong Qi, and Wanchun Dou. Keywordsdriven web APIs group recommendation for automatic app service creation process. Software—Practice and Experience, 51 (11):2337–2354, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guo:2021:TEF

[GLKZ21] Kehua Guo, Nan Li, Jian Kang, and Jian Zhang. Towards efficient federated learning-based scheme in medical cyber-physical systems for distributed data. Software—Practice and Experience, 51 (11):2274–2289, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Giang:2020:DAL

[GLL20] Nam Ky Giang, Rodger Lea, and Victor C. M. Leung. Developing applications in large scale, dynamic fog computing: a case study. Software— Practice and Experience, 50(5):519–532, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guo:2024:PLD

[GLL^{+}24] Juncai Guo, Jin Liu, Xiao Liu, Yao Wan, Yanjie Zhao, Li Li, Kui Liu, Jacques Klein, and Tegawendé F. Bissyandé. PyScribe — learning to describe Python code. Software—Practice and Experience, 54(3):501– 527, March 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ganino:2018:OPO

[GLMS18] Giulio Ganino, Domenico Lembo, Massimo Mecella, and Federico Scafoglieri. Ontology population for open-source intelligence: a GATE-based solution. Software—Practice and Experience, 48(12):2302– 2330, December 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grosse-Lindemann:1976:PPC

[GLN76] C. O. Grosse-Lindemann and H. H. Nagel. Postlude to a PASCAL-compiler bootstrap on a DECSystem-10. Software—Practice and Experience, 6(1):29– 42, January/March 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guo:2023:CSS

[GLS^{+}23] Shikai Guo, Zhongyan Liu, Zixuan Song, Hui Li, and Rong Chen. Code samples summarization for knowledge exchange in developer community. Software —Practice and Experience, 53(2):347–365, February 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gorton:2008:ELA

[GLT08] Ian Gorton, Yan Liu, and Nihar Trivedi. An extensible and lightweight architecture for adaptive server applications. Software— Practice and Experience, 38(8):853–883, July 10, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Glushkov:1974:EMP

[Glu74] V. M. Glushkov. Editorial: Mathematics and programming. Software—Practice and Experience, 4(3):195– 197, July/September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gluck:2012:SAO

[Glü12] Robert Glück. A selfapplicable online partial evaluator for recursive flowchart languages. Software—Practice and Experience, 42(6):649–673, June 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gomaa:1982:SEM

[GLW82] H. Gomaa, J. Lui, and P. Woo. The software engineering of a microcomputer application system. Software—Practice and Experience, 12(4):309– 321, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Good:1973:FPS

[GM73] J. Good and B. A. M.

Moon. FORTRAN — as provided by some major machine manufacturers in 1970. Software—Practice and Experience, 3(1):9–14, January/March 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gentleman:1977:DOS

[GM77] W. Morven Gentleman and J. Ian Munro. Designing overlay structures. Software—Practice and Experience, 7(4):493–500, July/ August 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gai:1985:DSA

[GM85a] Silvano Gai and Marco Mezzalama. Dynamic storage allocation: Experiments using the C language. Software—Practice and Experience, 15(7):693– 704, July 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Garratt:1985:ADS

[GM85b] P. W. Garratt and G. G. Michaelson. Analysis and design for a standard transport protocol. Software —Practice and Experience, 15(5):427–438, May 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grit:1985:PDC

[GM85c] Dale H. Grit and James R. McGraw. Programming divide and conquer for a MIMD machine. Software —Practice and Experience, 15(1):41–53, January 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gouda:2022:NEB

[GM22] Sunil Kumar Gouda and Ashok Kumar Mehta. A new evolutionism based self-adaptive multi-objective optimization method to predict software cost estimation. Software—Practice and Experience, 52(8): 1826–1848, August 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Goswami:2000:SSU

[GMC00] D. Goswami, R. Mall, and P. Chatterjee. Static slicing in Unix process environment. Software—Practice and Experience, 30(1): 17–36, January 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
69503097/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=69503097&PLACEBO=IE.
pdf.

Gomez:2021:SUF

[GMC^{+}21] Ángel Luis Perales Gómez, Lorenzo Fernández Maimó, Alberto Huertas Celdrán, Félix J. García Clemente, Manuel Gil Pérez, and Gregorio Martínez Pérez. Safe-Man: a unified framework to manage cybersecurity and safety in manufacturing industry. Software—Practice and Experience, 51(3):607–627, March 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Godboley:2017:ECP

[GMDM17] Sangharatna Godboley, Durga Prasad Mohapatra, Avijit Das, and Rajib Mall. Extended conference papers: An improved distributed concolic testing approach. Software—Practice and Experience, 47(2): 311–342, February 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ghosh:2022:SST

[GMGB22] Shreya Ghosh, Anwesha Mukherjee, Soumya K. Ghosh, and Rajkumar Buyya. STOPPAGE: Spatio-temporal data driven cloud-fog-edge computing framework for pandemic monitoring and management. Software—Practice and Experience, 52 (12):2700–2726, December 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gutierrez-Madronal:2019:EMT

[GMGDMB19] Lorena Gutiérrez-Madroñal, Antonio García-Domínguez, and Inmaculada Medina-Bulo. Evolutionary mutation testing for IoT with recorded and generated events. Software—Practice and Experience, 49(4):640– 672, April 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gil:2024:SOS

[GMGL24] Santiago Gil, Peter H. Mikkelsen, Cláudio Gomes, and Peter G. Larsen. Survey on open-source digital twin frameworks-A case study approach. Software —Practice and Experience, 54(6):929–960, June 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gajewska:1990:WXO

[GMM90] Hania Gajewska, Mark S.

Manasse, and Joel McCormack. Why X is not our ideal window system. Software—Practice and Experience, 20(S2):S137–S171, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ghazvini:2020:MMM

[GMNR20] Golnaz Aghaee Ghazvini, Mehran Mohsenzadeh, Ramin Nasiri, and Amir Masoud Rahmani. MMLT: a mutual multilevel trust framework based on trusted third parties in multicloud environments. Software—Practice and Experience, 50(7): 1203–1227, July 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gebala:2001:CIE

[GMO01] Robert M. Gebala, Carole M. McNamee, and Ronald A. Olsson. Compiler to interpreter: experiences with a distributed programming language. Software—Practice and Experience, 31(9):893– 909, July 25, 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        80502222/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=80502222&PLACEBO=IE.
                                        pdf.

Geist:2021:LML

[GMP^{+}21] Verena Geist, Michael Moser, Josef Pichler, Rodolfo Santos, and Volkmar Wieser. Leveraging machine learning for software redocumentation — a comprehensive comparison of methods in practice. Software —Practice and Experience, 51(4):798–823, April 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gallardo:2011:PUM

[GMPL11] M. M. Gallardo, P. Merino, L. Panizo, and A. Linares. A practical use of model checking for synthesis: generating a dam controller for flood management. Software—Practice and Experience, 41(11):1329–1347, October 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ghazouani:2020:BSM

[GMS20] Souad Ghazouani, Haithem Mezni, and Yahya Slimani. Bringing semantics to multicloud service compositions. Software—Practice and Experience, 50(4):447– 469, April 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gansner:2000:OGV

[GN00] Emden R. Gansner and Stephen C. North. An open graph visualization system and its applications to software engineering. Software—Practice and Experience, 30(11):1203–1233, September 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72517075/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72517075&PLACEBO=IE.
pdf.

Gervasi:2002:LVN

[GN02] Vincenzo Gervasi and Bashar Nuseibeh. Lightweight validation of natural language requirements. Software—Practice and Experience, 32(2):113–133, February 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
88013470/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?

                                        ID=88013470&PLACEBO=IE.
                                        pdf.

Gupta:2016:LSA

[GN16] Kartik Gupta and V. Krishna Nandivada. Lexical state analyzer for JavaCC grammars. Software— Practice and Experience, 46(6):751–765, June 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Girbea:2012:EAS

[GNSP12] Alina Girbea, Septimiu Nechifor, Francisc Sisak, and Liviu Perniu. Efficient address space generation for an OPC UA server. Software—Practice and Experience, 42(5):543– 557, May 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gansner:1988:DPD

[GNV88] E. R. Gansner, S. C. North, and K. P. Vo. DAG: a program that draws directed graphs. Software—Practice and Experience, 18 (11):1047–1062, November 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Goble:1971:SSM

[Gob71] P. N. Goble. Simulating a small machine for teaching purposes. Software—Practice and Experience, 1(3):209–221, July/ September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Goldschlager:1981:RSS

[Gol81a] Leslie M. Goldschlager.

Recursion in small storage. Software—Practice and Experience, 11(7):745– 751, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Goldschlager:1981:SAS

[Gol81b] Leslie M. Goldschlager.

Short algorithms for spacefilling curves. Software— Practice and Experience, 11(1):99, January 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gomaa:1974:ERA

[Gom74] H. Gomaa. An exercise in resource allocation. Software—Practice and Experience, 4(3):199–213, July/ September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gomaa:1978:CVH

[Gom78] H. Gomaa. The calibration and validation of a hybrid simulation/regression model of a batch computer system. Software—Practice and Experience, 8(1):

11–28, January/February 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gomaa:1982:DCS

[Gom82] H. Gomaa. The design and calibration of a simulation model of a star computer network. Software —Practice and Experience, 12(7):599–610, July 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gondzio:1987:MDT

[Gon87] Marek Gondzio. Microprocessor debugging techniques and their application in debugger design. Software—Practice and Experience, 17(3):215– 226, March 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Garcia:2016:DIE

[GOQ16] Miguel Garcia, Francisco Ortin, and Jose Quiroga. Design and implementation of an efficient hybrid dynamic and static typing language. Software—Practice and Experience, 46(2): 199–226, February 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gorlen:1987:OOC

[Gor87]  Keith E. Gorlen.  An

object-oriented class library for C++ programs. Software—Practice and Experience, 17(12):899– 922, December 1987. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gordon:1994:FSM

[Gor94] Steven A. Gordon. A faster Scrabble move generation algorithm. Software —Practice and Experience, 24(2):219–232, February 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gostick:1981:LE

[Gos81] R. W. Gostick. Letter to the editor. Software —Practice and Experience, 11(1):100, January 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gourlay:1986:LMP

[Gou86] John S. Gourlay. A language for music printing. Software—Practice and Experience, 29(5):388– 401, May 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gupta:2001:OSP

[GP01] Gopal Gupta and Enrico Pontelli. Optimization schemas for parallel implementation of nondeterministic languages and systems. Software —Practice and Experience, 31(12):1143–1181, October 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
85011732/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=85011732&PLACEBO=IE.
pdf.

Gog:2014:OSD

[GP14] Simon Gog and Matthias Petri. Optimized succinct data structures for massive data. Software—Practice and Experience, 44 (11):1287–1314, November 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ghormley:1998:GGL

[GPR^{+}98] Douglas P. Ghormley, David Petrou, Steven H. Rodrigues, Amin M. Vahdat, and Thomas E. Anderson. GLUnix: a Global Layer Unix for a network of workstations. Software—Practice and Experience, 28(9):929–961, July 25, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1822;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1822&
PLACEBO=IE.pdf.

Guo:2015:ECP

[GQ15] Hai-Feng Guo and Zongyan Qiu. Extended conference papers: A dynamic stochastic model for automatic grammar-based test generation. Software—Practice and Experience, 45(11):1519–1547, November 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gudes:1973:EBE

[GR73] Ehud Gudes and A. Reiter. On evaluating Boolean expressions. Software— Practice and Experience, 3(4):345–350, October/ December 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gannon:1979:IDA

[GR79] John D. Gannon and J. Rosenberg. Implementing data abstraction features in a stack-based language. Software—Practice and Experience, 9(7):547– 560, July 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gehani:1986:CC

[GR86] Narain H. Gehani and William D. Roome. Concurrent C. Software—Practice and Experience, 16(9): 821–844, September 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gehani:1988:CCC

[GR88] Narain H. Gehani and William D. Roome. Concurrent C++: Concurrent programming with class(es). Software—Practice and Experience, 18 (12):1157–1177, December 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gburzynski:1991:LPM

[GR91] Paweł Gburzyński and Piotr Rudnicki. LANSF: a protocol modelling environment and its implementation. Software—Practice and Experience, 21(1):51– 76, January 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gehani:1992:ICC

[GR92] Narain H. Gehani and William D. Roome. Implementing Concurrent C. Software—Practice and Experience, 22(3):265–285, March 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Galpin:1995:LSP

[GR95] V. C. Galpin and S. T.

Rock. A lift simulation prototype. Software—Practice and Experience, 25(3):251– 270, March 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grabowski:2017:SSA

[GR17] Szymon Grabowski and Marcin Raniszewski. Sampled suffix array with minimizers. Software—Practice and Experience, 47 (11):1755–1771, November 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grayson:1981:RKF

[Gra81] Martin Grayson. RNFREE — keyword free-format input. Software—Practice and Experience, 11(9):907– 911, September 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Graver:1992:EOO

[Gra92] Justin O. Graver. The evolution of an objectoriented compiler framework. Software—Practice and Experience, 22(7):519– 535, July 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Graefe:1996:ISD

[Gra96]  Goetz Graefe.  Iter-

ators, schedulers, and distributed-memory parallelism. Software—Practice and Experience, 26(4):427– 452, April 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16802.

Gawade:2014:CCS

[GRA14] Makarand Gawade, K. Ravikanth, and Sanjeev Aggarwal. Constantine: configurable static analysis tool in Eclipse. Software—Practice and Experience, 44(5): 537–563, May 2014. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Green:1980:ITN

[Gre80] T. R. G. Green. If’s and then’s: Is nesting just for the birds? Software —Practice and Experience, 10(5):373–381, May 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Garcia-Rois:2021:EMO

[GRFFGC^{+}21] Juan García-Rois, Pablo Fondo-Ferreiro, Felipe Gil-Castiñeira, Francisco J.

González-Castaño, and David Candal-Ventureira. Evaluating management and orchestration impact on closed-loop orchestration delay. Software—Practice and Experience, 51(2): 193–212, February 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Griswold:1975:PDF

[Gri75] Ralph E. Griswold. A portable diagnostic facility for SNOBOL4. Software—Practice and Experience, 5(1):93–104, January/March 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Griswold:1980:LEA

[Gri80] Ralph E. Griswold. Linguistic extension of abstract machine modelling to aid software development. Software—Practice and Experience, 10(1):1–9, January 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Griswold:1982:TAI

[Gri82] Ralph E. Griswold. A tool to aid in the installation of complex software systems. Software—Practice and Experience, 12(3):251– 267, March 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Griffiths:1986:ADC

[Gri86] J. G. Griffiths. An algorithm for displaying a class of space-filling curves. Software—Practice and Experience, 16(5):403–411, May 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grochow:1972:CMM

[Gro72a] Jerrold M. Grochow. Comment. Moo in multies. Software—Practice and Experience, 2(3):303–304, July 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grosch:1972:EGE

[Gro72b]  H. Grosch.  Editorial:

Guest editorial. Software— Practice and Experience, 2 (1):1–5, January 1972. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grogono:1973:MSE

[Gro73] Peter Grogono. MUSYS: Software for an electronic music studio. Software —Practice and Experience, 3(4):369–383, October/December 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grosch:1989:EGL

[Gro89] J. Grosch. Efficient generation of lexical analysers. Software—Practice and Experience, 19 (11):1089–1103, November 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grosch:1990:LGE

[Gro90]  J. Grosch.  Lalr —

a generator for efficient parsers. Software—Practice and Experience, 20 (11):1115–1135, November 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grothoff:2008:R

[Gro08] Christian Grothoff. The Runabout. Software— Practice and Experience, 38(14):1531–1560, November 25, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gomez:2006:STC

[GRR06] Juan Carlos Gomez, Jorge R.

Ramos, and Vernon Rego. Signals, timers, and continuations for multithreaded user-level protocols. Software—Practice and Experience, 36(5):449–471, April 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Girard:1974:IGT

[GRS74] P. M. Girard, I. H. Rose, and D. B. Scott. Implementation of a graphics and text output retrieval system. Software—Practice and Experience, 4(3):279– 288, July/September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grune:1979:CTL

[Gru79] Dick Grune. Choosing a tag-list algorithm for a compiler with special application to the ALEPH compiler. Software—Practice and Experience, 9(7):575– 593, July 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grundy:1983:BRB

[Gru83] Frances Grundy. Book review: Graded problems in computer science, A. D. McGettrick and P. D. Smith, Addison-Wesley, 1983. No. of pages: 314. Price: £6.95. ISBN 0-201-13787-9. Software— Practice and Experience, 13(10):980, October 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gomez-Rodriguez:2009:CPS

[GRVA09] C. Gómez-Rodríguez, J. Vilares, and M. A. Alonso.

A compiler for parsing schemata. Software—Practice and Experience, 39(5): 441–470, April 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gomez:2023:FSI

[GRVC23] Manuel J. Gomez, José A.

Ruipérez-Valiente, and Félix J. García Clemente. A framework to support interoperable Game-based Assessments as a Service (GBAaaS): Design, development, and use cases. Software—Practice and Experience, 53(11):2222– 2240, November 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gat:1976:MEP

[GS76] Israel Gat and Harry J.

Saal. Memoryless execution: a programmer’s viewpoint. Software—Practice and Experience, 6(4):463– 471, October/December 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gujar:1985:FSC

[GS85] Uday G. Gujar and Frank W. L. So. A flexible software character generator. Software—Practice and Experience, 15(2):191– 218, February 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gross:1990:SDA

[GS90] Thomas Gross and Peter Steenkiste. Structured dataflow analysis for arrays and its use in an optimizing compiler. Software—Practice and Experience, 20(2): 133–155, February 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gibbs:2006:FDC

[GS06a] Michael Gibbs and Bjarne Stroustrup. Fast dynamic casting. Software—Practice and Experience, 36(2): 139–156, February 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gregor:2006:SLS

[GS06b] Douglas Gregor and Sibylle Schupp. STLlint: lifting static checking from languages to libraries. Software—Practice and Experience, 36(3):225–254, March 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gomes:2008:VNC

[GS08] Daniel Gomes and Mário J. Silva. The Viúva Negra crawler: an experience report. Software—Practice and Experience, 38(2):161– 188, February ??, 2008.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gulcu:2014:FMS

[GSAE14] Koray Gülcü, Hasan Sözer, Barış Aktemur, and Ali Özer Ercan. Fault masking as a service. Software—Practice and Experience, 44(7):835– 854, July 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gonzalez-Sanchez:2011:PTS

[GSPA^{+}11] Alberto Gonzalez-Sanchez, Éric Piel, Rui Abreu, Hans-Gerhard Gross, and Arjan J. C. van Gemund. Prioritizing tests for software fault diagnosis. Software—Practice and Experience, 41(10):1105–1129, September 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grabowski:2017:BFB

[GSR17] Szymon Grabowski, Robert Susik, and Marcin Raniszewski. A Bloom filter based semiindex on q-grams. Software —Practice and Experience, 47(6):799–811, June 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guo:2020:FEO

[GSS^{+}20] Junqi Guo, Guicheng Shen, Yichen Sun, Jin Zhao, Hao Wu, and Zhilin Zhu. Field of experts optimizationbased noisy image retrieval. Software—Practice and Experience, 50 (11):2110–2118, November 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gentleman:1992:AMR

[GST92] W. Morven Gentleman, Terry Shepard, and Douglas V. P. Thoreson. Administrators and multiprocessor rendezvous mechanisms. Software—Practice and Experience, 22(1):1– 39, January 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grant:1995:EPF

[GSWZ95] P. W. Grant, J. A. Sharp, M. F. Webster, and X. Zhang. Experiences of parallelising finite-element problems in a functional style. Software—Practice and Experience, 25(9):947– 974, September 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Glaser:1987:LGC

[GT87] H. W. Glaser and P. Thompson. Lazy garbage collection. Software—Practice and Experience, 17(1):1– 4, January 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Graefe:1992:TPD

[GT92] Goetz Graefe and Shreekant S. Thakkar. Tuning a parallel database algorithm on a shared-memory multiprocessor. Software —Practice and Experience, 22(7):495–517, July 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Griswold:1993:DID

[GT93] William G. Griswold and Gregg M. Townsend. The design and implementation of dynamic hashing for sets and tables in Icon. Software —Practice and Experience, 23(4):351–367, April 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gaertner:2000:FFK

[GT00] N. Gaertner and B. Thirion.

A framework for fuzzy knowledge based control. Software—Practice and Experience, 30(1):1– 15, January 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
69503098/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?

                                        ID=69503098&PLACEBO=IE.
                                        pdf.

Gu:2005:LAO

[Gu05] Hung-Yan Gu. A largealphabet-oriented scheme for Chinese and English text compression. Software—Practice and Experience, 35(11):1027–1039, September 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guerraoui:2003:EM

[Gue03] R. Guerraoui. Editorial: Middleware. Software— Practice and Experience, 33(4):293, April 2003. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guttmann:1976:MDS

[Gut76] A. J. Guttmann. Multidimensional summations in FORTRAN. Software— Practice and Experience, 6(2):221–224, April/June 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gutknecht:1987:OPC

[Gut87] Jürg Gutknecht. One-pass compilation at its limits — a Modula-2 compiler for the Xerox Dragon computer. Software—Practice and Experience, 17(7):469– 484, July 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gill:2022:ISS

[GVBG22] Sukhpal Singh Gill, Ricardo Vinuesa, Venki Balasubramanian, and Soumya K. Ghosh. Innovative software systems for managing the impact of the COVID-19 pandemic. Software—Practice and Experience, 52(4): 821–823, April 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guimaraes:2018:EAB

[GVG^{+}18] E. Guimarães, S. Vidal, A. Garcia, J. A. Diaz Pace, and C. Marcos. Exploring architecture blueprints for prioritizing critical code anomalies: Experiences and tool support. Software—Practice and Experience, 48 (5):1077–1106, May 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2563.

Gonzalez-Velez:2010:SAS

[GVL10] Horacio González-Vélez and Mario Leyton. A survey of algorithmic skeleton frameworks: high-level structured parallel programming enablers. Software—Practice and Experience, 40(12):1135–1160, November 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gampe:2011:SMB

[GvRN^{+}11] Andreas Gampe, Jeffery von Ronne, David Niedzielski, Jonathan Vasek, and Kleanthis Psarris. Safe, multiphase bounds check elimination in Java. Software—Practice and Experience, 41(7):753–788, June 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gries:1984:LE

[GW84a] David Gries and P. H. Welch. Letters to the Editor. Software—Practice and Experience, 14(6): 601, June 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gutknecht:1984:ADP

[GW84b] Jürg Gutknecht and Werner Winiger. Andra: The document preparation system of the personal workstation Lilith. Software—Practice and Experience, 14(1):73– 100, January 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ganzinger:1985:FTL

[GW85] Harald Ganzinger and Walter Willmertinger. FOAM: a two-level approach to text formatting on a microcomputer system. Software —Practice and Experience, 15(4):327–341, April 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Goodwin:1996:ONO

[GW96] David W. Goodwin and Kent D. Wilken. Optimal and near-optimal global register allocation using 0-1 integer programming. Software—Practice and Experience, 26(8):929– 965, August 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16828.

Gorschek:2004:PSP

[GW04] Tony Gorschek and Claes Wohlin. Packaging software process improvement issues: a method and a case study. Software— Practice and Experience, 34(14):1311–1344, November 25, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gothe:1991:DAR

[GWA91] M. Carlsson Göthe, D. Wengelin, and L. Asplund. The Distributed Ada Run-time System DARTS. Software—Practice and Experience, 21(11):1249–1263, November 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Griffin:1988:DPP

[GWM88] James H. Griffin, Harvey J. Wasserman, and Lauren P. McGavran. A debugger for parallel processes. Software—Practice and Experience, 18(12):1179–1190, December 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guo:2023:FTL

[GWX^{+}23] Shikai Guo, Jiahui Wang, Zhihao Xu, Lin Huang, Hui Li, and Rong Chen. Feature transfer learning by reinforcement learning for detecting software defect. Software—Practice and Experience, 53(2):366– 389, February 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grunske:2011:EFI

[GWY^{+}11] Lars Grunske, Kirsten Winter, Nisansala Yatapanage, Saad Zafar, and Peter A. Lindsay. Experience with fault injection experiments for FMEA. Software—Practice and Experience, 41(11):1233–1258, October 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guo:2020:UAA

[GWZ^{+}20] Yan Guo, Shangguang Wang, Ao Zhou, Jinliang Xu, Jie Yuan, and Ching-Hsien Hsu. User allocationaware edge cloud placement in mobile edge computing. Software—Practice and Experience, 50(5):489– 502, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gao:2010:LRT

[GXN10] Lin Gao, Jingling Xue, and Tin-Fook Ngai. Loop recreation for thread-level speculation on multicore processors. Software—Practice and Experience, 40(1): 45–72, January ??, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guo:2016:TAN

[GYCL16] Yu-Chuan Guo, Wuu Yang, Jiunn-Yeu Chen, and Jenq-Kuen Lee. Translating the ARM Neon and VFP instructions in a binary translator. Software—Practice and Experience, 46(12):1591–1615, December 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grunwald:1993:CES

[GZ93] Dirk Grunwald and Benjamin G. Zorn. CustoMalloc: Efficient synthesized memory allocators. Software—Practice and Experience, 23(8):851– 869, August 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gu:2022:AAB

[GZW^{+}22] Zhaoquan Gu, Yan Zhu, Yuexuan Wang, Xiaojiang Du, Mohsen Guizani, and Zhihong Tian. Applying artificial bee colony algorithm to the multidepot vehicle routing problem. Software—Practice and Experience, 52(3):756– 771, March 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gu:2021:ADF

[GZX^{+}21] Rong Gu, Zhixiang Zhang, Zhihao Xu, Zhaokang Wang, Kai Zhang, Chunfeng Yuan, and Yihua Huang. Alchemy: Distributed financial quantitative analysis system with high-level programming model. Software— Practice and Experience, 51(8):1676–1699, August 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hahn:1972:DM

[HA72] K. W. Hahn and J. G.

Athey. Diagnostic messages. Software—Practice and Experience, 2(4):347– 352, October/December 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hernek:1990:EAP

[HA90] Diane Hernek and David P.

Anderson. Efficient automated protocol implementation using RTAG. Software—Practice and Experience, 20(9):869–885, September 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Haa:1982:CSS

[Haá82] Anna Haá. Computer system simulation in PAS-CAL. Software—Practice and Experience, 12(8):777– 784, August 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hac:1984:PDE

[Hać84] Anna Hać. PL/I as a discrete event simulation tool. Software—Practice and Experience, 14(7):697– 702, July 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hafiz:2013:PLD

[Haf13] Munawar Hafiz. A pattern language for developing privacy enhancing technologies. Software—Practice and Experience, 43(7): 769–787, July 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hall:1971:EGE

[Hal71]  Peter Hall.  Editorial:

Guest editorial. Software—Practice and Experience, 1(3):207–208, July/ September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hall:1982:MPC

[Hal82] J. A. Hall. A microprogrammed P-CODE interpreter for the Data General Eclipse S/130 minicomputer. Software—Practice and Experience, 12(8): 755–765, August 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Halang:1986:SMS

[Hal86] Wolfgang A. Halang. Suitable multiprocessor structures and virtual storage management when applying a feasible scheduling algorithm in hard realtime environment. Software—Practice and Experience, 16(8):761–769, August 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hamlet:1974:ULE

[Ham74] R. G. Hamlet. User-like executives. Software— Practice and Experience, 4 (1):41–49, January/March 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hammond:1977:BEP

[Ham77] John Hammond. BASIC — an evaluation of processing methods and a study of some programs. Software—Practice and Experience, 7(6):697–711, November/December 1977. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hamann:1979:SPM

[Ham79] Klaus Hamann. Structured programming in macro assembly languages comments on a paper by Jim Welch. Software—Practice and Experience, 9(2):171, February 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hampton:1981:URM

[Ham81] John S. Hampton. An unlimited register machine simulator. Software—Practice and Experience, 11(9): 999–1000, September 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hamilton:1984:DIG

[Ham84] J. E. M. Hamilton. The design of an interactive graphics system for the display of meteorological fields. Software—Practice and Experience, 14(6):587– 600, June 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hamlet:1995:IPT

[Ham95]  Dick Hamlet.  Imple-

menting prototype testing tools. Software—Practice and Experience, 25(4):347– 371, April 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heinze:2018:SAP

[HAM18] Thomas S. Heinze, Wolfram Amme, and Simon Moser. Static analysis and process model transformation for an advanced business process to Petri net mapping. Software—Practice and Experience, 48 (1):161–195, January 2018.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanford:1972:BRB

[Han72] K. V. Hanford. Book review: Compiler construction for digital computers, David Gries, Wiley, New York, 1971. No. of pages:

493. Price: £5.75. Software—Practice and Experience, 2(4):407–408, October 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1973:TMS

[Han73] Per Brinch Hansen. Testing a multiprogramming system. Software—Practice and Experience, 3(2):145– 150, April/June 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1976:DSC

[Han76a] Per Brinch Hansen. Disk scheduling at compile time. Software—Practice and Experience, 6(2):201–205, April/June 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1976:SOSa

[Han76b] Per Brinch Hansen. The Solo operating system: a Concurrent Pascal program. Software—Practice and Experience, 6(2):141– 149, April/June 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1976:SOSb

[Han76c] Per Brinch Hansen. The Solo operating system: Job interface. Software—Practice and Experience, 6(2): 151–164, April/June 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1976:SOSc

[Han76d] Per Brinch Hansen. The Solo operating system: Processes, monitors, and classes. Software—Practice and Experience, 6(2): 165–200, April/June 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1976:VAS

[Han76e] David R. Hanson. Variable association in SNOBOL4. Software—Practice and Experience, 6(2):245–254, April/June 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1977:BRB

[Han77a] Per Brinch Hansen. Book review: Studies in operating systems, R. M. McKeag and R. Wilson, edited by D. H. R. Huxtable, academic press, London, 1976. no. of pages: 263. price: £8.50. Software—Practice and Experience, 7(2):302– 303, March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1977:RES

[Han77b] David R. Hanson. RAT-SNO — an experiment in software adaptability. Software—Practice and Experience, 7(5):625–630, September/October 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1977:SMI

[Han77c] David R. Hanson. Storage management for an implementation of SNOBOL4. Software—Practice and Experience, 7(2):179–192, March/April 1977. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanneman:1978:BRBa

[Han78a] D. C. Hanneman. Book review: Computing systems hardware, M. Wells, Cambridge University Press, 1976. No. of pages: 245. Price: £4.00. Software— Practice and Experience, 8 (2):237, March 1978. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanneman:1978:BRBb

[Han78b] D. C. Hanneman. Book review: Macro processors, A. J. Cole, Cambridge University Press, 1976. No. of pages: 222. Price: £3.95. Software—Practice and Experience, 8(2):237– 238, March 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1978:RTM

[Han78c] Per Brinch Hansen. Reproducible testing of monitors. Software—Practice and Experience, 8(6):721– 729, November/December 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1978:EAS

[Han78d] David R. Hanson. Event associations in SNOBOL4 for program debugging. Software—Practice and Experience, 8(2):115–129, March/April 1978. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1979:AC

[Han79a] Per Brinch Hansen. Author’s correction. Software—Practice and Experience, 9(10):879–880, October 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1979:STC

[Han79b] David R. Hanson. A simple technique for controlled communication among separately compiled modules. Software—Practice and Experience, 9(11):921– 924, November 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1980:PFD

[Han80a]  David R. Hanson.  A

portable file directory system. Software—Practice and Experience, 10(8):623– 634, August 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1980:PSM

[Han80b]  David R. Hanson.  A

portable storage management system for the Icon programming language. Software—Practice and Experience, 10(6):489– 500, June 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1981:DE

[Han81a] Per Brinch Hansen. The design of Edison. Software —Practice and Experience, 11(4):363–396, April 1981.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1981:EML

[Han81b] Per Brinch Hansen. Edison— a multiprocessor language. Software—Practice and Experience, 11(4):325–361, April 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1981:EP

[Han81c] Per Brinch Hansen. Edison programs. Software —Practice and Experience, 11(4):397–414, April 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1981:GEG

[Han81d] Per Brinch Hansen. Guest editorial: Guest editorial. Software—Practice and Experience, 11(4):323, April 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1981:BSN

[Han81e] David R. Hanson. Is block structure necessary? Software—Practice and Experience, 11(8):853–866, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansche:1983:IOH

[Han83a] Brian Hansche. Improving overlay handlers. Software —Practice and Experience, 13(6):545–548, June 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1983:PIO

[Han83b] D. R. Hanson. A portable input/output system. Software—Practice and Experience, 13(1):95–100, January 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1983:SCO

[Han83c] David R. Hanson. Simple code optimizations. Software—Practice and Experience, 13(8):745–763, August 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1984:E

[Han84] David R. Hanson. Editorial. Software—Practice and Experience, 14 (12):1111, December 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1985:CRD

[Han85] David R. Hanson. Compact recursive-descent parsing of expressions. Software—Practice and Experience, 15(12):1205–1212, December 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1987:JPL

[Han87a] Per Brinch Hansen. Joyce — a programming language for distributed systems. Software—Practice and Experience, 17(1):29– 50, January 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1987:JI

[Han87b] Per Brinch Hansen. A Joyce implementation. Software—Practice and Experience, 17(4):267–276, April 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1987:LE

[Han87c] Wilfred J. Hansen. Letter to the editor. Software —Practice and Experience, 17(1):77–78, January 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1988:E

[Han88] David R. Hanson. Editorial. Software—Practice and Experience, 18(1):iii, January 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1989:JLR

[Han89a] Per Brinch Hansen. The Joyce language report. Software—Practice and Experience, 19(6):553–578, June 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1989:MIJ

[Han89b] Per Brinch Hansen. A multiprocessor implementation of Joyce. Software—Practice and Experience, 19(6): 579–592, June 1989. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1990:FAD

[Han90] David R. Hanson. Fast allocation and deallocation of memory based on object lifetimes. Software —Practice and Experience, 20(1):5–12, January 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1994:MLD

[Han94a] Per Brinch Hansen. Multiplelength division revisited: a tour of the minefield. Software—Practice and Experience, 24(6):579–601, June 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). This paper derives an algorithm for division of long integers, and implements it as a literate program, although without identifier cross-references. See comment about another division algorithm [Han95].

Hansen:1994:PLS

[Han94b] Per Brinch Hansen. The programming language SuperPascal. Software— Practice and Experience, 24(5):467–483, May 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1995:LDA

[Han95] Per Brinch Hansen. The long division algorithm of Linger, Mills and Witt. Software—Practice and Experience, 25(1):109, January 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [Han94a].

Hanson:1999:EEA

[Han99a] David R. Hanson. Early experience with ASDL in lcc. Software—Practice and Experience, 29(5):417– 435, April 25, 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

//storage.webhop.net/
documents/asdl.pdf;

http://www.cs.princeton.
edu/software/lcc/asdl.
html;   http://www3.interscience.
wiley.com/cgi-bin/abstract?
ID=55003854;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55003854&PLACEBO=IE.
pdf.

Hanson:1999:MID

[Han99b]  David R. Hanson.  A

machine-independent debugger — revisited. Software—Practice and Experience, 29(10):849–862, August 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=63001366;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=63001366&PLACEBO=IE.
pdf.

Hanson:2004:LNT

[Han04] David R. Hanson. lcc.NET:

targeting the .NET Common Intermediate Language from Standard C. Software—Practice and Experience, 34(3):265–286, March 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanssen:2011:ASP

[Han11] Geir K. Hanssen. Agile software product line engineering: enabling factors. Software—Practice and Experience, 41(8):883– 897, July 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harrison:1971:IST

[Har71a] M. C. Harrison. Implementation of the substring test by hashing. Communications of the Association for Computing Machinery, 14(12):777–779, December 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [TT82].

Harvey:1971:ADT

[Har71b] D. W. Harvey. Answering Datel telephone calls by PDP-8. Software—Practice and Experience, 1(3):223– 229, July/September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harris:1977:LE

[Har77] M. L. Harris. Letters to the editor. Software—Practice and Experience, 7(6):796, November 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harland:1980:HSD

[Har80a] David M. Harland. High speed data acquisition: Running a realtime process and a time-shared system (UNIX) concurrently. Software—Practice and Experience, 10(4):273–281, April 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harrington:1980:IPS

[Har80b] Steven Harrington. Infinite power series. Software —Practice and Experience, 10(10):835–848, October 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hart:1980:PAT

[Har80c] Richard Hart. Pattern analysis as a tool for inventing algorithms. Software —Practice and Experience, 10(5):405–417, May 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harris:1981:SUP

[Har81] James D. Harris. Sorting unsorted and partially sorted lists using the natural merge sort. Software—Practice and Experience, 11(12):1339–1340, December 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hardy:1982:SIC

[Har82] I. Trotter Hardy, Jr. The syntax of interactive command languages: a framework for design. Software —Practice and Experience, 12(1):67–75, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harandi:1983:ECR

[Har83] Mehdi T. Harandi. An experimental COBOL restructuring system. Software—Practice and Experience, 13(9):825–846, September 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harrison:1984:AMC

[Har84a] Warren A. Harrison. Applying Mccabe’s complexity measure to multipleexit programs. Software —Practice and Experience, 14(10):1004–1007, October 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hartson:1984:IPB

[Har84b]  H. Rex Hartson.  Im-

plementation of predicatebased protection in MUL-TISAFE. Software—Practice and Experience, 14(3): 207–234, March 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harland:1985:TLC

[Har85] David M. Harland. Towards a language for concurrent processes. Software —Practice and Experience, 15(9):839–888, September 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hartel:1991:PLC

[Har91] Pieter H. Hartel. Performance of lazy combinator graph reduction. Software—Practice and Experience, 21(3):299–329, March 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hartley:1992:OSF

[Har92] David H. Hartley. Optimizing stack frame accesses for processors with restricted addressing modes. Software —Practice and Experience, 22(2):101–110, February 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hart:1995:ELC

[Har95] Johnson M. Hart. Experience with logical code analysis in software maintenance. Software—Practice and Experience, 25 (11):1243–1262, November 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hartel:1999:LLE

[Har99] Pieter H. Hartel. LETOS - a lightweight execution tool for operational semantics. Software—Practice and Experience, 29 (15):1379–1416, December 25, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
68501286/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=68501286&PLACEBO=IE.
pdf.

Hastings:1977:FPH

[Has77] Jordan Towner Hastings. Floating point half-word packing for Control Data Corporation 6000/7000 series hardware. Software— Practice and Experience, 7 (1):146–147, January 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hatvany:1973:GES

[Hat73] J. Hatvany. Guest editorials: Software costs. Software—Practice and Experience, 3(4):307–308, October/December 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoogerbrugge:1999:CCS

[HATvdW99] Jan Hoogerbrugge, Lex Augusteijn, Jeroen Trum, and Rik van de Wiel. A code compression system based on pipelined interpreters. Software —Practice and Experience, 29(11):1005–1023, September 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=63501205;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=63501205&PLACEBO=IE.
pdf.

Hayden:1980:LER

[Hay80] Charles Hayden. Letters to the Editor: Remarks on ‘Implementation of concurrent Pascal on LSI-11’. Software—Practice and Experience, 10(10): 849–850, October 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [Mat80].

Hayashi:1983:PSP

[Hay83] Tsunetoshi Hayashi. A program structuring preprocessor for a macro assembly language. Software —Practice and Experience, 13(6):487–494, June 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hayashi:1987:TTH

[Hay87] Tsunetoshi Hayashi. Transporting TEX into the HITAC VOS3 environment. Software—Practice and Experience, 17(1):17– 28, January 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hazel:1971:BRB

[Haz71] P. Hazel. Book review: IBM 360 Assembler language programming, Gopal K. Kapur, Wiley, New York, 1971. No. of pages:

560. Price: £4.50. Software—Practice and Experience, 1(3):302–303, July/ September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hazel:1972:BRB

[Haz72] P. Hazel. Book review: System/360 Programming, Alex Thomas Jr., Holt, Rinehart and Winston, London, 1971. No. of pages:

250. Price: £4.40. Software—Practice and Experience, 2(3):307, July 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hazel:1974:GPT

[Haz74]  P. Hazel.  A general-

purpose text editor for OS/360. Software— Practice and Experience, 4(4):389–399, October/ December 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hazel:1980:DZT

[Haz80] Philipp Hazel. Development of the ZED text editor. Software—Practice and Experience, 10(1):57– 76, January 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hervas:2011:CCA

[HB11] Ramón Hervás and José Bravo. COIVA: contextaware and ontology-powered information visualization architecture. Software— Practice and Experience, 41(4):403–426, April 10, 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heidari:2018:CEN

[HB18] Safiollah Heidari and Rajkumar Buyya. Costefficient and network-aware dynamic repartitioningbased algorithms for scheduling large-scale graphs in cloud computing environments. Software—Practice and Experience, 48 (12):2174–2192, December 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hassan:2022:SSA

[HBB22] Sara Hassan, Rami Bahsoon, and Rajkumar Buyya.

Systematic scalability analysis for microservices granularity adaptation design decisions. Software—Practice and Experience, 52(6): 1378–1401, June 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hurault:2015:SLA

[HBC15] Aurélie Hurault, Kyungim Baek, and Henri Casanova. Selecting linear algebra kernel composition using response time prediction. Software—Practice and Experience, 45 (12):1659–1676, December 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hourcade:2004:BKA

[HBD04] Juan Pablo Hourcade, Benjamin B. Bederson, and Allison Druin. Building Kid-Pad: an application for children’s collaborative storytelling. Software—Practice and Experience, 34(9): 895–914, July 25, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Humphrey:2005:CUB

[HBJ05] Marty Humphrey, Jim Basney, and Jim Jokl. The case for using Bridge Certificate Authorities for Grid computing. Software—Practice and Experience, 35(9):

817–826, July 25, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hassan:2020:MTG

[HBK20] Sara Hassan, Rami Bahsoon, and Rick Kazman. Microservice transition and its granularity problem: a systematic mapping study. Software—Practice and Experience, 50(9):1651–1681, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hartman:2006:CSS

[HBM06] John H. Hartman, Scott Baker, and Ian Murdock. Customizing the Swarm storage system using agents. Software— Practice and Experience, 36(2):117–137, February 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hall:1979:IAC

[HC79] Chris Hall and Richard Clayton. Isolation of assembler coding ‘optimizations’. Software—Practice and Experience, 9(3): 248, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Horspool:1987:HCT

[HC87a] R. Nigel Horspool and Gordon V. Cormack. Hashing as a compaction technique for LR parser tables. Software—Practice and Experience, 17(6):413– 416, June 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hughes:1987:PIM

[HC87b] J. G. Hughes and Michelle Connolly. A portable implementation of a modular multiprocessing database programming language. Software—Practice and Experience, 17(8):533–546, August 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hung:1993:RRA

[HC93] Yung-Chen Hung and Gen-Huey Chen. Reverse reachability analysis: a new technique for deadlock detection on communicating finite state machines. Software—Practice and Experience, 23(9):965–979, September 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hsu:1997:CCJ

[HC97a] Shang-Te T. Hsu and Ruei-Chuan C. Chang. Continuous checkpointing: Joining the checkpointing with virtual memory paging. Software—Practice and Experience, 27(9):1103–1120, September 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7336;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7336&
PLACEBO=IE.pdf.

Hui:1997:FEL

[HC97b] Chi-Chung Hui and Samuel T.

Chanson. Flexible and extensible load balancing. Software—Practice and Experience, 27(11):1283– 1306, November 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7271;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7271&
PLACEBO=IE.pdf.

Horspool:1998:TCJ

[HC98] R. Nigel Horspool and Jason Corless. Tailored compression of Java class files. Software—Practice and Experience, 28(12):1253– 1268, October 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.

                                        com/cgi-bin/abstract?
                                        ID=1762;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=1762&
                                        PLACEBO=IE.pdf.

Hsu:1999:IUR

[HC99] Shang-Te Hsu and Ruei-Chuan Chang. An implementation of using remote memory to checkpoint processes. Software—Practice and Experience, 29(11):985–1004, September 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=63501204;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=63501204&PLACEBO=IE.
                                        pdf.

Hosking:2000:EPO

[HC00] Antony Hosking and Quintin Cutts. Editorial: Persistent object systems. Software—Practice and Experience, 30(4):293–294, April 10, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        71004137/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=71004137&PLACEBO=IE.
                                        pdf.

Hsiao:2010:EST

[HC10] Chun-Feng Hsiao and Chih-Ping Chu. Enhancing SCORM through creating a reusable Java class repository. Software—Practice and Experience, 40(10): 865–881, September 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:2012:VAJ

[HC12] Ting-Chang Huang and Da-Wei Chang. VM aware journaling: improving journaling file system performance in virtualization environments. Software—Practice and Experience, 42(3):303–330, March 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hasso:2013:SPC

[HC13] Sargon Hasso and R. C.

Carlson. Software patterns catalog based on a new classification scheme. Software—Practice and Experience, 43(7):759–768, July 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:2016:THF

[HC16] Ting-Chang Huang and Da-Wei Chang. TridentFS: a hybrid file system for non-volatile RAM, flash memory and magnetic disk. Software—Practice and Experience, 46(3):291– 318, March 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:2020:PSG

[HC20] Chung-Ming Huang and Pin-Jui Chen. Proximate sharing of geo data using the credit-considered mobile edge computing server switching control scheme. Software—Practice and Experience, 50 (9):1822–1841, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:1996:OLM

[HCC96] Chung-Ming Huang, Yao-Fang Chu, and Ye-In Chang. An ODA-like multimedia document system. Software—Practice and Experience, 26(10):1097– 1126, October 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16764.

Henin:1984:LCG

[HCD84] Bernard Henin, Daniel Couprie, and Pierre Douspis. A LIS compiler for GCOS-7. Software—Practice and Experience, 14(3): 253–261, March 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harper-Cyr:2019:FFT

[HCDB19] Christian Harper-Cyr, Michel R.

Dagenais, and Ahmad S. Bushehri. Fast and flexible tracepoints in x86. Software—Practice and Experience, 49(12):1712–1727, December 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hussain:2016:ECP

[HCG^{+}16] Ishtiaque Hussain, Christoph Csallner, Mark Grechanik, Qing Xie, Sangmin Park, Kunal Taneja, and B. M. Mainul Hossain. Extended conference paper: RUGRAT: Evaluating program analysis and testing tools and compilers with large generated random benchmark applications. Software—Practice and Experience, 46(3):405–431, March 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hull:1986:CCS

[HD86] M. Elizabeth C. Hull and G. Donnan. Contextually communicating sequential processes: a software engineering environment. Software—Practice and Experience, 16(9):845–864, September 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hennessy:1982:DIP

[HE82] John L. Hennessy and Hilding Elmquist. The design and implementation of parametric types in PAS-CAL. Software—Practice and Experience, 12(2):169– 184, February 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heal:1981:SC

[Hea81] J. G. B. Heal. On sorting and counting. Software —Practice and Experience, 11(8):879, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heffler:1982:DMC

[Hef82] Michael J. Heffler. Description of a menu creation and interpretation system. Software—Practice and Experience, 12(3):269– 281, March 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heher:1976:SFR

[Heh76] A. D. Heher. Some features of a real-time BASIC executive. Software—Practice and Experience, 6(3):387– 391, July/September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Helsgaun:1995:CST

[Hel95] Keld Helsgaun. CBack: a simple tool for backtrack programming in C. Software—Practice and Experience, 25(8):905–934, August 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Henrici:1979:TMF

[Hen79] A. Henrici. The titan and the micro — a fable. Software—Practice and Experience, 9(7):595– 596, July 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Herbert:1977:LEM

[Her77] Andrew Herbert. Letters to the Editor: Moo on the CAP computer. Software —Practice and Experience, 7(6):797–798, November 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Herbert:1984:BRB

[Her84] Andrew Herbert. Book review: Local Area Networks — Issues, Products and Developments, V. E. Cheong and R. A. Hirscheim, Wiley, 98. No. of pages 190. Software— Practice and Experience, 14(2):196–197, February 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hesketh:1991:PUB

[Hes91] Richard Hesketh. Perly — UNIX with buttons. Software—Practice and Experience, 21(11):1165–1187, November 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heuring:1986:AGF

[Heu86] V. P. Heuring. The automatic generation of fast lexical analysers. Software —Practice and Experience, 16(9):801–808, September 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heiser:1998:MSA

[HEV^{+}98] Gernot Heiser, Kevin Elphinstone, Jerry Vochteloo, Stephen Russell, and Jochen Liedtke. The Mungi singleaddress-space operating system. Software—Practice and Experience, 28(9):901– 928, July 25, 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1820;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1820&
PLACEBO=IE.pdf.

Hedrick:1973:HLP

[HF73] G. E. Hedrick and J. R.

Forrest. A high level plotting system. Software—Practice and Experience, 3(4):359–368, October/December 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hague:1976:PPC

[HF76] Stephen J. Hague and Brian Ford. Portability— prediction and correction. Software—Practice and Experience, 6(1):61– 69, January/March 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1980:TOS

[HF80] Per Brinch Hansen and Jon Fellows. The Trio operating system. Software— Practice and Experience, 10(11):943–948, November 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hsieh:1998:TSC

[HFPB98] Wilson C. Hsieh, Marc E.

Fiuczynski, Przemysław Pardyak, and Brian N. Bershad. Type-safe casting. Software—Practice and Experience, 28(11):1245– 1252, September 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=10007360;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=10007360&PLACEBO=IE.
pdf.

Henderson:1981:MLP

[HG81] Peter Henderson and Roger B. Gimson. Modularization of large programs. Software —Practice and Experience, 11(5):497–520, May 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harland:1984:PPA

[HG84] David M. Harland and Hamish I. E. Gunn. Polymorphic programming I. another language designed on semantic principles. Software—Practice and Experience, 14(10):973– 997, October 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Henry:1989:CMG

[HG89] Sallie M. Henry and Roger Goff. Complexity measurement of a graphical programming language. Software—Practice and Experience, 19(11):1065–1088, November 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heiden:2019:EPS

[HGK^{+}19] Simon Heiden, Lars Grunske, Timo Kehrer, Fabian Keller, Andre van Hoorn, Antonio Filieri, and David Lo. An evaluation of pure spectrum-based fault localization techniques for large-scale software systems. Software—Practice and Experience, 49(8): 1197–1224, August 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hartel:1994:CFL

[HGW94] Pieter H. Hartel, Hugh Glaser, and John. M. Wild. Compilation of functional languages using flow graph analysis. Software—Practice and Experience, 24(2): 127–173, February 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Herman-Giddens:1975:BBS

[HGWBS75] G. S. Herman-Giddens, R. B. Warren, R. C. Barr, and M. S. Spach. Biomac: Block structured programming using PDP-11 assembler language. Software—Practice and Experience, 5(4):359–374, October/December 1975. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1979:MC

[HH79] Per Brinch Hansen and Charles Hayden. Microcomputer comparison. Software—Practice and Experience, 9(3):211–217, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hamlet:1980:TPS

[HH80] R. G. Hamlet and R. M. Haralick. Transportable package software. Software—Practice and Experience, 10(12):1009–1027, December 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hunter:1982:NSE

[HH82] J. Alan Hunter and Nigel F. Hall. A network screen editor implementation. Software—Practice and Experience, 12(9):843–856, September 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Halme:1988:GED

[HH88] Heikki Halme and Juha Heinänen. GNU Emacs as a dynamically extensible programming environment. Software—Practice and Experience, 18(10): 999–1009, October 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hartson:1990:DHC

[HHK90] H. Rex Hartson, Deborah Hix, and Thomas M. Kraly. Developing human — computer interface models and representation techniques. Software—Practice and Experience, 20(5):425– 457, May 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:1984:SRT

[HHL84] J. C. Huang, Melody Ho, and Ted Law. A simulator for real-time software debugging and testing. Software—Practice and Experience, 14(9):845– 855, September 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harford:1992:NPM

[HHM92] A. Gayler Harford, Vincent P. Heuring, and Michael G. Main. A new parsing method for non-LR(1) grammars. Software —Practice and Experience, 22(5):419–437, May 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hillenbrand:2012:DEE

[HHMMG12] M. Hillenbrand, M. Heinz, J. Matheis, and K. D.

Müller-Glaser. Development of electric/electronic architectures for safetyrelated vehicle functions. Software—Practice and Experience, 42(7):817–851, July 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hylton:2019:TPC

[HHPSS19] Alan Hylton, Gregory Henselman-Petrusek, Janche Sang, and Robert Short. Tuning the performance of a computational persistent homology package. Software—Practice and Experience, 49(5):885–905, May 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1993:EDS

[HHR93] Eric N. Hanson, Tina M. Harvey, and Mark A. Roth. Experiences in database system implementation using a persistent programming language. Software—Practice and Experience, 23(12):1361–1377, December 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

He:2003:QCD

[HHRS03] Jun He, Matti A. Hiltunen, Mohan Rajagopalan, and Richard D. Schlichting. QoS customization in distributed object systems. Software—Practice and Experience, 33(4):295–320, April 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harper:1995:IID

[HHZ^{+}95] Mary P. Harper, Randall A. Helzerman, Carla B. Zoltowski, Boon-Dock Yeo, Yin Chan, Todd Stewart, and Bryan L. Pellom. Implementation issues in the development of the PARSEC parser. Software—Practice and Experience, 25(8):831–862, August 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hikita:1985:MPT

[HI85] Teruo Hikita and Kiyoshi Ishihata. A method of program transformation between variable sharing and message passing. Software —Practice and Experience, 15(7):677–692, July 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Himsolt:2000:GDI

[Him00] Michael Himsolt. Graphlet:

design and implementation of a graph editor. Software—Practice and Experience, 30(11):1303–1324, September 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        72517076/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=72517076&PLACEBO=IE.
                                        pdf.

Henricksen:2006:UCP

[HIR06] Karen Henricksen, Jadwiga Indulska, and Andry Rakotonirainy. Using context and preferences to implement self-adapting pervasive computing applications. Software—Practice and Experience, 36(11– 12):1307–1330, September/ October 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hewage:2024:CEN

[HIRB24] Tharindu B. Hewage, Shashikant Ilager, Maria A. Rodriguez, and Rajkumar Buyya. CloudSim express: a novel framework for rapid low code simulation of cloud computing environments. Software—Practice and Experience, 54(3):483– 500, March 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heines:1988:CPA

[HJ88a] T. S. Heines and P. J. Jalics. Compiler performance analysis tools. Software —Practice and Experience, 18(9):917–921, September 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hobley:1988:RBS

[HJ88b] Philip Hobley and Peter Jesty. A recovery block scheme for a VAX11/750 running Unix. Software —Practice and Experience, 18(6):561–581, June 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hafiz:2008:EMA

[HJ08] Munawar Hafiz and Ralph E.

Johnson. Evolution of the MTA architecture: the impact of security. Software —Practice and Experience, 38(15):1569–1599, December ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Han:2014:RPI

[HJ14] Sanghyun Han and Hyun-Wook Jin. Resource partitioning for Integrated Modular Avionics: comparative study of implementation alternatives. Software—Practice and Experience, 44(12):1441–1466, December 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:2000:CIC

[HJC00] C. M. Huang, M. Y. Jang, and Y. C. Chao. CTW: an integrated computer and telephone-accessed WWW system. Software—Practice and Experience, 30 (13):1485–1507, November 10, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
73001201/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=73001201&PLACEBO=IE.
pdf.

Hassoun:2005:ADP

[HJC05] Youssef Hassoun, Roger Johnson, and Steve Counsell. Applications of dynamic proxies in distributed environments. Software—Practice and Experience, 35(1):75–99, January 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hac:1989:PSD

[HJS89] Anna Hác, Xiaowei Jin, and Jo-Han Soo. A performance study of deadlock prevention algorithms in a distributed file system. Software—Practice and Experience, 19(5):461– 489, May 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:2020:DYC

[HJS^{+}20] Yuan Huang, Nan Jia, Junhuai Shu, Xinyu Hu, Xiangping Chen, and Qiang Zhou. Does your code need comment? Software—Practice and Experience, 50(3):227–245, March 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heine:1984:EIH

[HK84a] P. Heine and F. Kaiser.

An economical implementation of the high level real-time language PEARL on microcomputers: Intel RMX86-PEARL. Software —Practice and Experience, 14(4):377–382, April 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Henry:1984:ESS

[HK84b] Sallie M. Henry and Dennis G. Kafura. The evaluation of software systems’ structure using quantitative software metrics. Software—Practice and Experience, 14(6):561–573, June 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hautamaki:2006:FDS

[HK06a] Juha Hautamäki and Kai Koskimies. Finding and documenting the specialization interface of an application framework. Software—Practice and Experience, 36(13):1443–1465, November 10, 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hirschfeld:2006:DSA

[HK06b] Robert Hirschfeld and Katsuya Kawamura. Dynamic service adaptation. Software—Practice and Experience, 36(11–12):1115–1131, September/October 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hashemian:2012:WWG

[HKA12] Raoufehsadat Hashemian, Diwakar Krishnamurthy, and Martin Arlitt. Web workload generation challenges — an empirical investigation. Software— Practice and Experience, 42(5):629–647, May 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hambury:1972:DPC

[HKB72] J. N. Hambury, S. Karniel, and G. C. Barney. A display package for CAD of control systems. Software—Practice and Experience, 2(1):29–41, January/ March 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:2012:DIA

[HKC^{+}12] Chung-Ming Huang, Hao-Hsiang Ku, Yu-Chang Chao, Cheng-Wei Lin, and Yen-Wen Chen. Design and implementation of an adaptive Web2.0 QoS-based home appliance control service platform. Software —Practice and Experience, 42(1):57–87, January 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hong:2009:CAT

[HKM^{+}09] Sunghyun Hong, Jin-Chul Kim, Soo-Mook Moon, Jin Woo Shin, Jaemok Lee, Hyeong-Seok Oh, and Hyung-Kyu Choi. Client ahead-of-time compiler for embedded Java platforms. Software—Practice and Experience, 39(3):259–278, March 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huemer:1995:MOO

[HKV95] Christian Huemer, Gerti Kappel, and Stefan Vieweg. Migration in object-oriented database systems — a practical approach. Software—Practice and Experience, 25(10):1065–1096, October 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hunter:1977:ASO

[HKW77] G. Hunter, M. Kuriyan, and A. Wallis. The advantages of stream output procedures over formatdominated record transmissions. Software—Practice and Experience, 7(3): 347–356, May/June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Humphries:2000:IGS

[HKWZ00] Thorna O. Humphries, Artur Klauser, Alexander L. Wolf, and Benjamin G. Zorn. An infrastructure for generating and sharing experimental workloads for persistent object systems. Software—Practice and Experience, 30(4):387– 417, April 10, 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
71004141/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=71004141&PLACEBO=IE.
pdf.

Hammerl:1979:CRT

[HL79] Lothar Hammerl and Klaus-Peter Löhr. Comments on reproducible testing of monitors. Software—Practice and Experience, 9(7): 596, July 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Horspool:1987:MIS

[HL87] R. Nigel Horspool and Michael R. Levy. Mkscan — an interactive scanner generator. Software—Practice and Experience, 17(6): 369–378, June 1987. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

He:1991:MCP

[HL91] Xudong He and John A. N. Lee. A methodology for constructing predicate transition net specifications. Software—Practice and Experience, 21(8): 845–875, August 1991. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hashemi:1992:IPS

[HL92] Reza Hashemi and Ronald J.

Leach. Issues in porting software from C to C++. Software—Practice and Experience, 22(7):599– 602, July 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:1994:MME

[HL94] Chung-Ming Huang and Chung-Ming Lo. Multimedia E-mail: the evolution approach based on adapters. Software—Practice and Experience, 24(9): 785–800, September 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:1998:DCI

[HL98] Sun-Jen Huang and Richard Lai. Deriving complexity information from a formal communication protocol specification. Software—Practice and Experience, 28(14):1465–1491, December 10, 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=5001666;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=5001666&PLACEBO=IE.
pdf.

Ho:2002:ISC

[HL02a] K. S. Ho and H. V. Leong. Improving the scalability of the CORBA event service with a multi-agent load balancing algorithm. Software—Practice and Experience, 32(5):417–441, April 25, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
93513337/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=93513337&PLACEBO=IE.
pdf.

Huang:2002:PCP

[HL02b] Sun-Jen Huang and Richard Lai. PSAMS: a communication protocol specification assessment and measurement system. Software—Practice and Experience, 32(12):1129–1154, October 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
98516188/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=98516188&PLACEBO=IE.
pdf.

Huang:2003:SSD

[HL03] Y. Huang and Y. H.

Lee. System and software designs for transaction processing in asymmetric communication environments. Software— Practice and Experience, 33(14):1359–1376, November 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hsu:2020:IML

[HL20] I-Ching Hsu and Yin-Hung Lin. Integrated machine learning with Semantic Web for open government data recommendation based on cloud computing. Software—Practice and Experience, 50 (12):2293–2312, December 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hsiung:2005:SOO

[HLFS05] Pao-Ann Hsiung, Trong-Yen Lee, Jih-Ming Fu, and Win-Bin See. SESAG: an object-oriented application framework for real-time systems. Software—Practice and Experience, 35 (10):899–921, August 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoffman:2011:GBT

[HLGSW11] Daniel Malcolm Hoffman, David Ly-Gagnon, Paul Strooper, and Hong-Yi Wang. Grammarbased test generation with YouGen. Software—Practice and Experience, 41(4): 427–447, April 10, 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:2015:PCA

[HLH15] Chung-Ming Huang, Shih-Yang Lin, and Tsung-Han Hsieh. The personalized context-aware mobile advertisement system using a novel approaching detection method over cellular networks. Software—Practice and Experience, 45(1): 31–52, January 2015. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

He:2021:SCO

[HLLZ21] Zhenli He, Kenli Li, Keqin Li, and Wei Zhou. Server configuration optimization in mobile edge computing: a cost-performance tradeoff perspective. Software—Practice and Experience, 51(9):1868–1895, September 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Houston:2003:CAS

[HLR^{+}03] I. Houston, M. C. Little, I. Robinson, S. K. Shrivastava, and S. M. Wheater. The CORBA Activity Service Framework for supporting extended transactions. Software—Practice and Experience, 33(4):351– 373, April 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hwang:2020:AAA

[HLR20] Sungjae Hwang, Sungho Lee, and Sukyoung Ryu. All about activity injection: Threats, semantics, detection, and defense. Software—Practice and Experience, 50(7): 1061–1086, July 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heilig:2018:MCC

[HLRVB18] Leonard Heilig, Eduardo Lalla-Ruiz, Stefan Voß, and Rajkumar Buyya. Metaheuristics in cloud computing. Software—Practice and Experience, 48(10): 1729–1733, October 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hazel:1973:SCF

[HLS73] P. Hazel, J. Larmouth, and A. Stoneley. Some comments on FORTRAN systems. Software—Practice and Experience, 3(2):185– 187, April/June 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:2008:UIA

[HLW08] Chung-Ming Huang, Ming-Sian Lin, and Hon-Long Wong. A ubiquitous IAs access platform (UIAP) over UPnP. Software— Practice and Experience, 38(11):1127–1147, September ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hennessy:1982:CPC

[HM82] John L. Hennessy and Noah Mendelsohn. Compilation of the Pascal case statement. Software— Practice and Experience, 12(9):879–882, September 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hull:1984:GAQ

[HM84] M. Elizabeth C. Hull and R. M. McKeag. A general approach to queuing in C. S. P. Software—Practice and Experience, 14(8):769– 773, August 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hudson:1990:GUI

[HM90] Scott E. Hudson and Shamim P. Mohamed. A graphical user interface server for UNIX. Software—Practice and Experience, 20(12):1227–1239, December 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hartel:2012:SAD

[HM12] Pieter H. Hartel and Henk L. Muller. Simple algebraic data types for C. Software—Practice and Experience, 42(2):191–210, February 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:2018:KMM

[HM18] Chung-Ming Huang and Chia-Ming Mai. K-MING: a mobile proxy handoff control scheme for proximate group-based geodata sharing. Software—Practice and Experience, 48 (3):600–620, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2551.

Huedo:2004:FAE

[HML04] Eduardo Huedo, Ruben S. Montero, and Ignacio M. Llorente. A framework for adaptive execution in grids. Software—Practice and Experience, 34(7):631– 651, June 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hierons:2011:SBT

[HMN11] Robert M. Hierons, Mercedes G. Merayo, and Manuel Núñez. Scenariosbased testing of systems with distributed ports. Software—Practice and Experience, 41(10):999– 1026, September 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hutchinson:1989:TIN

[HMPT89] Norman C. Hutchinson, Shivakant Mishra, Larry L. Peterson, and Vicraj T. Thomas. Tools for implementing network protocols. Software—Practice and Experience, 19(9):895– 916, September 1989. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hirsch:2020:DTD

[HMRZ20] Matías Hirsch, Cristian Mateos, Juan Manuel Rodriguez, and Alejandro Zunino. DewSim: a tracedriven toolkit for simulating mobile device clusters in Dew computing environments. Software—Practice and Experience, 50(5):688– 718, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hayes:1988:SSC

[HMS88] Roger Hayes, Steve W.

Manweiler, and Richard D. Schlichting. A simple system for constructing distributed, mixed-language programs. Software—Practice and Experience, 18(7): 641–660, July 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hwang:1995:RLS

[HMS^{+}95] Yuan-Shin Hwang, Bongki Moon, Shamik D. Sharma, Ravi Ponnusamy, Raja Das, and Joel H. Saltz. Runtime and language support for compiling adaptive irregular programs on distributed-memory machines. Software—Practice and Experience, 25(6):597– 621, June 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Humayun:2023:TRM

[HNJ^{+}23] Mamoona Humayun, Mahmood Niazi, Noor Zaman Jhanjhi, Sajjad Mahmood, and Mohammad Alshayeb. Toward a readiness model for secure software coding. Software—Practice and Experience, 53 (4):1013–1035, April 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hosking:2001:PRE

[HNW^{+}01] Antony L. Hosking, Nathaniel Nystrom, David Whitlock, Quintin Cutts, and Amer Diwan. Partial redundancy elimination for access path expressions. Software—Practice and Experience, 31(6):577–600, May 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
78003106/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=78003106&PLACEBO=IE.
pdf.

Ho:1991:AGD

[HO91] W. Wilson Ho and Ronald A.

Olsson. An approach to genuine dynamic linking. Software—Practice and Experience, 21(4):375– 390, April 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoare:1972:GEQ

[Hoa72] C. A. R. Hoare. Guest editorial: The quality of software. Software—Practice and Experience, 2(2):103– 105, April 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoaglin:1973:ALO

[Hoa73] David C. Hoaglin. An analysis of the loop optimization scores in Knuth’s ‘empirical study of FOR-TRAN programs’. Software —Practice and Experience, 3(2):161–169, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoffman:1989:PIS

[Hof89] Daniel Hoffman. Practical interface specification. Software—Practice and Experience, 19(2):127– 148, February 1989. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hohn:2004:LLM

[Hoh04]  Michael H. Hohn.  A

little language for modularizing numerical PDE solvers. Software—Practice and Experience, 34(9): 797–813, July 25, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Holdsworth:1977:SIA

[Hol77] D. Holdsworth. System implementation in Algol 68-r. Software—Practice and Experience, 7(3):331– 339, June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Holdsworth:1983:SAA

[Hol83] David Holdsworth. A system for analysing Ada programs at run-time. Software—Practice and Experience, 13(5):407–421, May 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Holzmann:1988:IPR

[Hol88] Gerard J. Holzmann. An improved protocol reachability analysis technique. Software—Practice and Experience, 18(2):137–161, February 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Holsti:1989:SEI

[Hol89] Niklas Holsti. A session editor with incremental execution functions. Software —Practice and Experience, 19(4):329–350, April 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Holzmann:1993:SPI

[Hol93] Gerard J. Holzmann. Standardized protocol interfaces. Software—Practice and Experience, 23(7):711– 731, July 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hope:1971:PGD

[Hop71] Alexander K. Hope. A planar graph drawing program. Software—Practice and Experience, 1(1):83– 91, January/March 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hopgood:1973:BRB

[Hop73] F. R. A. Hopgood. Book review: Compiler techniques, B. W. Pollack (ed.), Auerbach, New York, 1972. No. of pages: 558. Price: £8.00. Software—Practice and Experience, 3(2):188, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hopgood:1974:BRB

[Hop74] F. R. A. Hopgood. Book review: ALGOL 60 compilation and assessment, B. A. Wichmann, Academic Press, London, 1973. No. of pages: 320. Price: £6.50. Software—Practice and Experience, 4(4):405– 406, October/December 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hopkins:1980:PVB

[Hop80a] T. R. Hopkins. PBA-SIC: a verifier for BA-SIC. Software—Practice and Experience, 10(3):175– 181, March 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoppe:1980:SNW

[Hop80b] Jiri Hoppe. A simple nucleus written in Modula-2: a case study. Software —Practice and Experience, 10(9):697–706, September 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoppe:1986:AAI

[Hop86]  Jiri Hoppe.  Another

approach to the implementation of synchronization primitives. Software—Practice and Experience, 16(12):1109–1116, December 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hopkins:1996:RSC

[Hop96] T. R. Hopkins. Restructuring software: A case study. Software—Practice and Experience, 26(8):967– 982, August 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16829.

Horn:1978:RAM

[Hor78] Berthold K. P. Horn. Rational arithmetic for minicomputers. Software— Practice and Experience, 8 (2):171–176, March/April 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Horspool:1980:PFS

[Hor80] R. Nigel Horspool. Practical fast searching in strings. Software—Practice and Experience, 10(6):501– 506, June 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Horn:1981:LE

[Hor81] R. J. Horn III. Letter to the editor. Software—Practice and Experience, 11(7): 767, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Horspool:2007:ODR

[Hor07a] Nigel Horspool. Obituaries: Douglas T. Ross (1929– 2007). Software—Practice and Experience, 37(7): 691, June 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Horspool:2007:OPBa

[Hor07b] Nigel Horspool. Obituaries: Peter Brown (1939– 2007). Software—Practice and Experience, 37(5):463, April 25, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Horspool:2007:OPBb

[Hor07c] Nigel Horspool. Obituary: Per Brinch Hansen (1938–2007). Software— Practice and Experience, 37(12):1243, October 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Horspool:2012:Ea

[Hor12a] R. Nigel Horspool. Editorial. Software—Practice and Experience, 42(1): 1, January 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Horspool:2012:Eb

[Hor12b] R. Nigel Horspool. Editorial. Software—Practice and Experience, 42(1): 1, January 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Horspool:2014:EWR

[Hor14] Nigel Horspool. Editorial: A welcome to Raj Buyya as the new coeditor. Software—Practice and Experience, 44(8): 887, August 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hora:2021:AMA

[Hor21] Andre Hora. APISonar:

Mining API usage examples. Software—Practice and Experience, 51(2):319– 352, February 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoffmann:1985:IIA

[HOS85] Christoph M. Hoffmann, Michael J. O’Donnell, and Robert I. Strandh. Implementation of an interpreter for abstract equations. Software—Practice and Experience, 15 (12):1185–1204, December 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoshen:1998:GTM

[Hos98] Joseph Hoshen. A graph theoretical method for the management and synchronization of large software updates. Software—Practice and Experience, 28(8):845–857, July 10, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1815;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1815&
PLACEBO=IE.pdf.

Howarth:1976:BRB

[How76] D. J. Howarth. Book review: Principles of Systems Programming, Robert M. Graham, Wiley, New York, 1975. No. of Pages: 422. Price: £8.25. Software— Practice and Experience, 6 (4):584, October 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Howden:1978:EES

[How78] William E. Howden. An evaluation of the effectiveness of symbolic testing. Software—Practice and Experience, 8(4):381– 397, July/August 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heo:2017:SCC

[HOY17] Kihong Heo, Hakjoo Oh, and Kwangkeun Yi. Selective conjunction of contextsensitivity and octagon domain toward scalable and precise global static analysis. Software—Practice and Experience, 47(11):1677– 1705, November 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hughes:1983:DLD

[HP83a] J. W. Hughes and M. S.

Powell. DTL: a language for the design and implementation of concurrent programs as structured networks. Software—Practice and Experience, 13(12):1099–1112, December 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hughes:1983:ID

[HP83b] J. W. Hughes and M. S. Powell. The implementation of DTL. Software—Practice and Experience, 13(12):1113–1128, December 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Honeyman:1987:PAA

[HP87] Peter Honeyman and Patricia E. Parseghian. Parsing ambiguous addresses for electronic services. Software—Practice and Experience, 17(1):51–60, January 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hentzel:1988:PSL

[HP88] I. R. Hentzel and D. J. Pokrass. A practical solution for a large sparse matrix. Software—Practice and Experience, 18(3):279– 283, March 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:2004:RCC

[HP04] David R. Hanson and Todd A. Proebsting. A research C# compiler. Software—Practice and Experience, 34(13):1211–1224, November 10, 2004. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL ftp://

ftp.research.microsoft.
com/pub/tr/tr-2003-32.
pdf;   http://research.
microsoft.com/research/
pubs/view.aspx?msr_tr_
id=MSR-TR-2003-32.

Hnetynka:2011:UMM

[HP11] Petr Hnětynka and František Plášil. Using metamodeling in design and implementation of componentbased systems: the SOFA case study. Software— Practice and Experience, 41(11):1185–1201, October 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hartman:2000:EBC

[HPB^{+}00] John Hartman, Larry Peterson, Andy Bavier, Peter Bigot, Patrick Bridges, Brady Montz, Rob Piltz, Todd Proebsting, and Oliver Spatscheck. Experiences building a communicationoriented JavaOS. Software—Practice and Experience, 30(10):1107–1126, August 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72510350/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72510350&PLACEBO=IE.
pdf.

Hellmann:1996:TVN

[HPC^{+}96] Douglas R. Hellmann, Mitchell L. Patterson, Robert L. Cannon, Philip D. Moore, and Christopher G. St. C. Kendall. Tools for visual and numerical data entry. Software—Practice and Experience, 26(5):531– 544, May 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16808.

Hong:2012:TEE

[HPK^{+}12] Kirak Hong, Jiin Park, Sungho Kim, Taekhoon Kim, Hwangho Kim, Bernd Burgstaller, and Bernhard Scholz. TinyVM: an energy-efficient execution infrastructure for sensor networks. Software— Practice and Experience, 42(10):1193–1209, October 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Park:2015:SVD

[hPmKgH15] Seong hun Park, Sung min Kim, and Young guk Ha. Scalable visualization for DBpedia ontology analysis using Hadoop. Software —Practice and Experience, 45(8):1103–1114, August 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hevia:2022:QQS

[HPP22] Jose Luis Hevia, Guido Peterssen, and Mario Piattini. QuantumPath: a quantum software development platform. Software—Practice and Experience, 52(6): 1517–1530, June 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hu:2020:USR

[HPZ^{+}20] Gang Hu, Min Peng, Yihan Zhang, Qianqian Xie, Wang Gao, and Mengting Yuan. Unsupervised software repositories mining and its application to code search. Software—Practice and Experience, 50(3):299– 322, March 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heaps:1977:CDM

[HR77] H. S. Heaps and T. Radhakrishnan. Compaction of diagnostic messages for compilers. Software—Practice and Experience, 7(1): 139–144, January/March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1990:LE

[HR90] David R. Hanson and Charles W. Reynolds. Letter to the Editor. Software —Practice and Experience, 20(9):965, September 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1996:MID

[HR96] David R. Hanson and Mukund Raghavachari. A machine-independent debugger. Software—Practice and Experience, 26 (11):1277–1299, November 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

URL http://storage.
webhop.net/documents/
cdb.pdf;   http://storage.
webhop.net/documents/
cdbtalk.pdf;   http:/
/www.cs.princeton.edu/
software/lcc/cdb;   http:
//www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16773.

Hague:2006:EUP

[HR06] Rob Hague and Peter Robinson. End-user programming of reconfigurable systems. Software—Practice and Experience, 36(11– 12):1285–1306, September/ October 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hennen:2000:OJL

[HRM00] Dennis S. Hennen, Suresh Ramachandran, and Sandra A. Mamrak. The Object-JavaScript language.

Software—Practice and Experience, 30(14):1571– 1585, November 25, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www3.interscience.
wiley.com/cgi-bin/abstract/
73001731/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=73001731&PLACEBO=IE.
pdf.

Hammouda:2009:DPT

[HRS^{+}09] Imed Hammouda, Anna Ruokonen, Mika Siikarla, André L. Santos, Kai Koskimies, and Tarja Systä. Design profiles: toward unified tool support for design patterns and UML profiles. Software—Practice and Experience, 39(4): 331–354, March 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Holdsworth:1973:MTB

[HRW73] D. Holdsworth, G. W.

Robinson, and M. Wells. A multi-terminal benchmark. Software—Practice and Experience, 3(1):43– 59, January/March 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hillman:1977:EIN

[HS77] A. L. Hillman and D. Schofield.

EDIT — an interactive network service; design and implementation. Software —Practice and Experience, 7(5):595–611, September/ October 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Haring:1983:REC

[HS83] G. Haring and O. Schechtner. On the realization of extended control structures in FORTRAN. Software —Practice and Experience, 13(5):431–445, May 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Horspool:1985:ASC

[HS85] R. Nigel Horspool and André Scheunemann. Automating the selection of code templates. Software —Practice and Experience, 15(5):503–514, May 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Homer:1989:ITC

[HS89] William Homer and Richard Schooler. Independent testing of compiler phases using a test case generator. Software—Practice and Experience, 19(1):53– 62, January 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hume:1991:FSS

[HS91] Andrew Hume and Daniel Sunday. Fast string searching. Software—Practice and Experience, 21 (11):1221–1248, November 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoffman:1997:CFA

[HS97] Daniel Hoffman and Paul Strooper. ClassBench: a framework for automated class testing. Software—Practice and Experience, 27(5):573–597, May 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7308;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7308&
PLACEBO=IE.pdf.

Hassija:2021:MDO

[HSC21] Vikas Hassija, Vikas Saxena, and Vinay Chamola. A mobile data offloading framework based on a combination of blockchain and virtual voting. Software—Practice and Experience, 51(12):2428–2445, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hauswirth:2010:TVP

[HSD10] Matthias Hauswirth, Peter F. Sweeney, and Amer Diwan. Temporal vertical profiling. Software—Practice and Experience, 40(8): 627–654, July 2010. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansson:1981:DPS

[HSM81] Hasse Hansson and Jørgen Steensgaard-Madsen. Document preparation systems. Software—Practice and Experience, 11(9):983– 997, September 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hashemi:2024:EAR

[HSRB24] Sayed Mohsen Hashemi, Amir Sahafi, Amir Masoud Rahmani, and Mahdi Bohlouli. Energy-aware resource management in fog computing for IoT applications: a review, taxonomy, and future directions. Software—Practice and Experience, 54(2):109– 148, February 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hsu:2012:EUM

[Hsu12] I-Ching Hsu. Extending UML to model Web 2.0based context-aware applications. Software—Practice and Experience, 42 (10):1211–1227, October 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hill:1975:WAA

[HSW75] I. D. Hill, R. S. Scowen, and B. A. Wichmann. Writing algorithms in ALGOL 60. Software—Practice and Experience, 5(3):229–244, July/September 1975. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hu:2020:IWM

[HSY^{+}20] Chengyu Hu, Xin Shu, Xuesong Yan, Deze Zeng, Wenyin Gong, and Lei Wang. Inline wireless mobile sensors and fog nodes placement for leakage detection in water distribution systems. Software—Practice and Experience, 50(7):1152–1167, July 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hinxman:1982:PEC

[HT82] A. I. Hinxman and Austin Tate. Parsing an extension to CODASYL FORTRAN DML. Software—Practice and Experience, 12(3):205– 209, March 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Horton:1986:UCP

[HT86] I. A. Horton and S. J. Turner. Using coroutines in Pascal. Software—Practice and Experience, 16(1):45– 61, January 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hasanagic:2019:RDS

[HTJNL19] Miran Hasanagić, Peter W. V. Tran-Jørgensen, René S. Nilsson, and Peter Gorm Larsen. Realization of distributed system models using code generation extensions. Software—Practice and Experience, 49(3):478–497, March 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hung:2015:COS

[HTWS15] Shih-Hao Hung, Tien-Tzong Tzeng, Gyun-De Wu, and Jeng-Peng Shieh. A code offloading scheme for big-data processing in Android applications. Software—Practice and Experience, 45(8):1087–1101, August 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:2022:DBT

[HTX^{+}22] Rubing Huang, Dave Towey, Yinyin Xu, Yunan Zhou, and Ning Yang. Dissimilaritybased test case prioritization through data fusion. Software—Practice and Experience, 52(6): 1352–1377, June 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:1987:DFT

[Hua87] J. C. Huang. A depth-first and tool-supportable programming strategy. Software—Practice and Experience, 17(2):159–163, February 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hudson:1972:IPS

[Hud72] R. J. W. Hudson. An implementation of phrase structures in ALGOL. Software—Practice and Experience, 2(3):231–240, July/ September 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hugo:1977:NDC

[Hug77] Ian Hugo. A note on documentation for computer users. Software—Practice and Experience, 7(3): 425, June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hughes:1979:FEM

[Hug79] J. W. Hughes. A formalization and explication of the Michael Jackson method of program design. Software —Practice and Experience, 9(3):191–202, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hughes:1982:SIG

[Hug82] R. J. M. Hughes. A semiincremental garbage collection algorithm. Software—Practice and Experience, 12(11):1081–1082, November 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hughes:1988:MIU

[Hug88] Larry Hughes. A multicast interface for UNIX 4.3. Software—Practice and Experience, 18(1):15– 27, January 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hughes:1993:OIL

[Hug93] Larry Hughes. Object identification in the Lego kernel. Software—Practice and Experience, 23(4):405– 418, April 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hughes:1997:EHL

[Hug97] Larry Hughes. Event handling in the Lego system. Software—Practice and Experience, 27(3):291– 310, March 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7292;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7292&
PLACEBO=IE.pdf.

Hummel:1976:LLO

[Hum76] Helmut Hummel. LEK-TOR: a list-oriented, machineindependent programming system for conversational applications. Software— Practice and Experience, 6(4):447–462, October/ December 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hume:1988:TTG

[Hum88] Andrew Hume. A tale of two greps. Software—Practice and Experience, 18 (11):1063–1072, November 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hunter:1972:BRB

[Hun72] R. B. Hunter. Book review: Programming in Pop-2, R. M. Burstall, J. S. Collins and R. J. Popplestone, Edinburgh University Press, 1971. No. of pages: 290. Price: £3. Software—Practice and Experience, 2(1): 99, January 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hunt:1980:I

[Hun80]  J. G. Hunt.  Inter-

rupts. Software—Practice and Experience, 10(7):523– 530, July 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hunter:1981:PPA

[Hun81] Geoffrey Hunter. Partitioned and prioritized access to timesharing services: the reserved availability method. Software —Practice and Experience, 11(7):689–696, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hung:1997:HSG

[Hun97] Yung-Chen Hung. A hypervideo system generator. Software—Practice and Experience, 27(11):1263– 1281, November 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7270;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7270&
PLACEBO=IE.pdf.

Hung:2000:LHS

[Hun00] Yung-Chen Hung. A layered hyperanimation system. Software—Practice and Experience, 30(10): 1067–1094, August 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www3.interscience.
wiley.com/cgi-bin/abstract/
72506516/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72506516&PLACEBO=IE.
pdf.

Hurst:1980:PPP

[Hur80] A. John Hurst. Pascal-P, program structure and program behaviour. Software—Practice and Experience, 10(12):1029–1036, December 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huskamp:1986:MOS

[Hus86] Jeffrey C. Huskamp. A modular operating system for the CRAY-1. Software—Practice and Experience, 16(12):1059–1076, December 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hardwick:1991:IUI

[HUS^{+}91] Martin Hardwick, Wayne Uejio, David L. Spooner, Joe Czechowski, Phil Lohr, and Brion Sarachan. Implementing a user interface management system for existing applications using an object-oriented database tool. Software—Practice and Experience, 21(2):167– 185, February 1991. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hutt:1974:BRB

[Hut74] A. T. F. Hutt. Book review: Management reporting systems, J. M. McKeever, John Wiley, New York, 1971. No. of pages:

94. Price: £3.75. Software—Practice and Experience, 4(3):300–301, July/ September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hutt:1976:BRB

[Hut76] A. T. F. Hutt. Book review: Management Information Systems, Infotech State of the Art Report 21, Edited by C. Boon, Infotech Information, Maidenhead, Berks., England, 1974. No. of pages: 661.

Price: £60. Software —Practice and Experience, 6(3):439–440, July 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hutt:1978:DMA

[Hut78] Andrew T. F. Hutt. Data mappings again. Software—Practice and Experience, 8(4):483–493, July/ August 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hutt:1979:CMR

[Hut79a] Andrew T. F. Hutt. A compiler model for a relational data base management system. Software—Practice and Experience, 9(2):157– 169, February 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hutt:1979:ODR

[Hut79b] Andrew T. F. Hutt. Organizing the description of a relational data base. Software—Practice and Experience, 9(5):361–368, May 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hartel:1988:SGR

[HV88] Pieter H. Hartel and Arthur H. Veen. Statistics on graph reduction of SASL programs. Software—Practice and Experience, 18(3):239–253, March 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hasselbring:2002:SRB

[HvdH02] Wilhelm Hasselbring and Ralph van den Houdt. Specifying a role-based guide for learning to work with an enterprise framework. Software—Practice and Experience, 32(8):833– 836, July 10, 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
94519523/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=94519523&PLACEBO=IE.
pdf.

Hopgood:1977:BRB

[HW77] F. R. A. Hopgood and R. W. Witty. Book review: An introduction to compiler writing, J. S. Rohl, Macdonald and Jane’s, London, 1975. No. of pages:

307. price £4.50. Software—Practice and Experience, 7(1):149–150, January 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Haddon:1978:EUI

[HW78] B. K. Haddon and W. M.

Waite. Experience with the universal intermediate language Janus. Software —Practice and Experience, 8(5):601–616, September/ October 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Holden:1980:AM

[HW80] J. Holden and I. C. Wand.

An assessment of Modula. Software—Practice and Experience, 10(8):593– 621, August 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Halewood:1988:NSD

[HW88] K. Halewood and M. R.

Woodward. NSEDIT: a syntax-directed editor and testing tool based on nassishneiderman charts. Software—Practice and Experience, 18(10):987–998, October 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Horspool:1990:EFL

[HW90] R. Nigel Horspool and Michael Whitney. Even faster LR parsing. Software —Practice and Experience, 20(6):515–535, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hwang:1994:UPP

[HW94] Wen-Ping Hwang and Ching-Lin Wang. Using procedural parameters and continuations in combinatorial searches. Software —Practice and Experience, 24(4):377–386, April 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hu:1998:CIV

[HW98] Chung-Hua Hu and Feng-Jian Wang. Constructing an integrated visual programming environment. Software—Practice and Experience, 28(7):773– 798, July 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1808;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1808&
PLACEBO=IE.pdf.

Horspool:2010:EFS

[HW10a] Nigel Horspool and Andy Wellings. Editorials: Focus Section Editorial. Software —Practice and Experience, 40(12):1105, November 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Horspool:2010:FSE

[HW10b] Nigel Horspool and Andy Wellings. Focus section editorial. Software—Practice and Experience, 40 (12):1105, ???? 2010. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hume:2015:SCS

[HW15] Tom Hume and Des Watson. Short communication: Is superoptimization viable for VM instruction sets? Software—Practice and Experience, 45(4):571– 579, April 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hatton:1988:SKS

[HWS^{+}88] Les Hatton, Andy Wright, Stuart Smith, Gregg Parkes, Paddy Bennett, and Robert Laws. The Seismic Kernel System — a large-scale exercise in Fortran 77 portability. Software—Practice and Experience, 18(4):301– 329, April 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Describes portability issues in a 500,000+ line system.

Huang:2020:POL

[HY20] Ming Hsiang Huang and Wuu Yang. PFACC: an OpenACC-like programming model for irregular nested parallelism. Software—Practice and Experience, 50(10):1877–1904, October 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:2019:DLS

[HYC19] Qing Huang, Yang Yang, and Ming Cheng. Deep learning the semantics of change sequences for query expansion. Software—Practice and Experience, 49(11):1600–1617, November 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huh:2015:ECP

[HYH15] Sungju Huh, Jonghun Yoo, and Seongsoo Hong. Extended conference papers: Cross-layer resource control and scheduling for improving interactivity in Android. Software—Practice and Experience, 45 (11):1549–1570, November 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:2013:TSD

[HYT13] Chung-Ming Huang, Chia-Ching Yang, and Chun-Yu Tseng. A telematics service discovery platform using the distributed peerto-peer networking system. Software—Practice and Experience, 43(6):641– 662, June 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Han:2023:SMS

[HYY23] Xue Han, Tingting Yu, and Gongjun Yan. A systematic mapping study of software performance research. Software—Practice and Experience, 53(5): 1249–1270, May 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Huang:2018:QEB

[HYZ^{+}18] Qing Huang, Yangrui Yang, Xue Zhan, Hongyan Wan, and Guoqing Wu. Query expansion based on statistical learning from code changes. Software—Practice and Experience, 48(7):1333–1351, July 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL https:

//onlinelibrary.wiley.
com/doi/abs/10.1002/spe.
2574.

Henderson:1994:COO

[HZ94] Robert Henderson and Benjamin G. Zorn. A comparison of object-oriented programming in four modern languages. Software—Practice and Experience, 24(11):1077–1095, November 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hsu:1995:ASC

[HZ95] William H. Hsu and Amy E. Zwarico. Automatic synthesis of compression techniques for heterogeneous files. Software— Practice and Experience, 25(10):1097–1116, October 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

He:2022:DDA

[HZWC22] Jing He, Jun Zhou, Haoyu Wang, and Li Cai. DACA: Distributed adaptive grid decision graph based clustering algorithm. Software—Practice and Experience, 52(5):1199–1215, May 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Iqbal:2021:DIT

[IAA^{+}21] Muhammad Shahid Iqbal, Iftikhar Ahmad, Muhammad Asif, Sun-Hee Kim, and Raja Majid Mehmood. Drug investigation tool: Identifying the effect of drug on cell image by using improved correlation. Software—Practice and Experience, 51(2):260– 270, February 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Iannello:1990:PAD

[Ian90] Giulio Iannello. Programming abstract data types, iterators and generic modules in C. Software—Practice and Experience, 20(3): 243–260, March 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Iribarne:2017:MBD

[IAPC17] Luis Iribarne, José-Andrés Asensio, Nicolás Padilla, and Javier Criado. Modeling Big data-based systems through ontological trading. Software—Practice and Experience, 47 (11):1561–1596, November 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ireland:2016:SDT

[IASC16] Iain Ireland, José Nelson Amaral, Raúl Silvera, and Shimin Cui. SafeType: detecting type violations for type-basedalias analysis of C. Software—Practice and Experience, 46(11):1571– 1588, November 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ismail:2013:IPE

[IB13] Leila Ismail and Rajeev Barua. Implementation and performance evaluation of a distributed conjugate gradient method in a cloud computing environment. Software—Practice and Experience, 43(3):281– 304, March 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ibrahim:2021:EEL

[IBA^{+}21] Muhammad Ibrahim, Naveed Khan Baloch, Sheraz Anjum, Yousaf Bin Zikria, and Sung Won Kim. An energy efficient and low overhead fault mitigation technique for Internet of Thing edge devices reliable onchip communication. Software—Practice and Experience, 51(12):2393–2410, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ibsen:1984:PVM

[Ibs84] Leif Ibsen. A portable virtual machine for Ada. Software—Practice and Experience, 14(1):17–29, January 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Iyengar:1985:EAC

[IC85] S. Sitharama Iyengar and Hsi Chang. Efficient algorithms to create and maintain balanced and threaded binary search trees. Software—Practice and Experience, 15(10):925–941, October 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ierusalimschy:1996:LEE

[IdFF96] Roberto Ierusalimschy, Luiz Henrique de Figueiredo, and Waldemar Celes Filho. Lua — an extensible extension language. Software—Practice and Experience, 26(6): 635–652, June 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16813.

Ierusalimschy:2009:TPM

[Ier09] Roberto Ierusalimschy. A text pattern-matching tool based on Parsing Expression Grammars. Software —Practice and Experience, 39(3):221–258, March 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Iyer:2001:JBR

[IH01] K. Viswanathan Iyer and M. Harinath. Java based replicated server objects and their remote method invocations in a distributed environment. Software—Practice and Experience, 31(4):357–370, April 10, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        77004438/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=77004438&PLACEBO=IE.
                                        pdf.

Inzinger:2014:GEB

[IHS^{+}14] Christian Inzinger, Waldemar Hummer, Benjamin Satzger, Philipp Leitner, and Schahram Dustdar. Generic event-based monitoring and adaptation methodology for heterogeneous distributed systems. Software—Practice and Experience, 44(7):805– 822, July 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ivanchykhin:2017:RAU

[IIL17] Dmytro Ivanchykhin, Sergey Ignatchenko, and Daniel Lemire. Regular and almost universal hashing: an efficient implementation. Software—Practice and Experience, 47(10): 1299–1323, October 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ismail:2015:IPE

[IK15] Leila Ismail and Latifur Khan. Implementation and performance evaluation of a scheduling algorithm for divisible load parallel applications in a cloud computing environment. Software —Practice and Experience, 45(6):765–781, June 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Inverardi:1993:EDL

[IM93] Paola Inverardi and Franco Mazzanti. Experimenting with dynamic linking with Ada. Software—Practice and Experience, 23(1):1– 14, January 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ingalls:2020:TDL

[IMBB20] Daniel Ingalls, Eliot Miranda, Clément Béra, and Elisa Gonzalez Boix. Two decades of live coding and debugging of virtual machines through simulation. Software—Practice and Experience, 50 (9):1629–1650, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Iwendi:2021:MOA

[IMG^{+}21] Celestine Iwendi, Praveen Kumar Reddy Maddikunta, Thippa Reddy Gadekallu, Kuruva Lakshmanna, Ali Kashif Bashir, and Md. Jalil Piran. A metaheuristic optimization approach for energy efficiency in the IoT networks. Software—Practice and Experience, 51 (12):2558–2571, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Inoue:2012:HPS

[IMKN12] Hiroshi Inoue, Takao Moriyama, Hideaki Komatsu, and Toshio Nakatani. A highperformance sorting algorithm for multicore singleinstruction multiple-data processors. Software— Practice and Experience, 42(6):753–777, June 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ince:1981:DTA

[Inc81] Darrel C. Ince. Decision tables in Algol 68. Software —Practice and Experience, 11(7):753–756, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ince:1983:STT

[Inc83] Darrel C. Ince. A software tool for top-down programming. Software—Practice and Experience, 13(8):687– 695, August 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ince:1984:SCC

[Inc84] D. C. Ince. A source code control system based on semantic nets. Software—Practice and Experience, 14(12):1159–1168, December 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ince:1985:PDL

[Inc85] D. C. Ince. A program design language based software maintenance tool. Software—Practice and Experience, 15(6):583–594, June 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ince:1986:BRB

[Inc86] D. C. Ince. Book review:

Program evolution, M. M. Lehman and L. A. Belady, APIC Studies in Data Processing, Academic Press, 1985. No. of pages: 538. Price: $35.00 (paperback £21.50). Software—Practice and Experience, 16 (11):1055, November 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Innes:1977:ELR

[Inn77] Donal R. Innes. Exploiting the least recently used page replacement algorithm. Software—Practice and Experience, 7(2):271– 273, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ince:1980:ABA

[IR80] Darrel C. Ince and Keith Robson. An Algol 68 based algebraic manipulation system. Software—Practice and Experience, 10(6):427– 430, June 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Iosevich:2005:SDS

[IS05] Vadim Iosevich and Assaf Schuster. Software Distributed Shared Memory: a VIA-based implementation and comparison of sequential consistency with homebased lazy release consistency. Software—Practice and Experience, 35(8):755– 786, July 10, 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Iseli:1990:MCW

[Ise90] Christian Iseli. A microcode compiler for the watch-oriented RISC processor. Software—Practice and Experience, 20(7):729– 747, July 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ilager:2021:SIE

[ISPB21] Shashikant Ilager, Vlado Stankovski, Shrideep Pallickarar, and Rajkumar Buyya. Special issue: Elastic computing from edge to the cloud environments. Software—Practice and Experience, 51 (9):1849–1851, September 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ishikawa:2006:EAM

[ISUG06] Norihiro Ishikawa, Hideharu Suzuki, Hidetoshi Ueno, and Tetsuya Gotoh. Experiment on and analysis of mobile content transformation using XSLT. Software—Practice and Experience, 36(7):761– 783, June 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Iwasaki:2002:DLB

[Iwa02] Hideya Iwasaki. Developing a Lisp-based preprocessor for TEX documents. Software—Practice and Experience, 32 (14):1345–1363, November 25, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
98518913/START;   http:
//www3.interscience.wiley.

                                        com/cgi-bin/fulltext?
                                        ID=98518913&PLACEBO=IE.
                                        pdf.

Izatt:1980:DAI

[Iza80] W. T. Izatt. Domain architecture and the ICL 2900 series. Software—Practice and Experience, 10(4):319– 328, April 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jaaksi:1995:IIA

[Jaa95a] Ari Jaaksi. Implementing interactive applications in C++. Software—Practice and Experience, 25(3):271– 289, March 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jaaksi:1995:OOS

[Jaa95b] Ari Jaaksi. Object-oriented specification of user interfaces. Software—Practice and Experience, 25 (11):1203–1221, November 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jha:2020:IES

[JAA^{+}20] Devki Nandan Jha, Khaled Alwasel, Areeb Alshoshan, Xianghua Huang, Ranesh Kumar Naha, Sudheer Kumar Battula, Saurabh Garg, Deepak Puthal, Philip James, Albert Zomaya, Schahram Dustdar, and Rajiv Ranjan. IoTSim-Edge: a simulation framework for modeling the behavior of Internet of Things and edge computing environments. Software—Practice and Experience, 50(6): 844–867, June 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jackson:1971:BRB

[Jac71] M. A. Jackson. Book review: Design of sequential file systems, Thomas R. Gildersleeve, Wiley– Interscience, 1971. No. of pages: 47. Price: £2.50. Software—Practice and Experience, 1(4):412, October/December 1971. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jackson:1984:BRB

[Jac84] Adrienne Jackson. Book review: The Unix System, S. R. Bourne, Addison Wesley, London, 1982. No. of pages: 351. Price: £11.95. Software—Practice and Experience, 14(7):703– 704, July 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jackson:1985:DAP

[Jac85] M. I. Jackson. Developing Ada programs using the Vienna Development Method (VDM). Software—Practice and Experience, 15(3): 305–318, March 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jain:2004:IME

[JAJB04] Rahul Jain, Sanjeev Kumar Aggarwal, Pankaj Jalote, and Shiladitya Biswas. An interactive method for extracting grammar from programs. Software—Practice and Experience, 34(5): 433–447, April 25, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jaksic:2004:MBS

[Jak04] Mirjana Jakšić. Mapping of bibliographical standards into XML. Software—Practice and Experience, 34(11):1051–1064, September 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jarrah:2021:GGB

[JAkM^{+}21] Moath Jarrah, Bahaa Al-khatieb, Naseem Mahasneh, Baghdad Alkhateeb, and Yaser Jararweh. GDBApex: a graph-based system to enable efficient transformation of enterprise infrastructures. Software—Practice and Experience, 51(3): 517–531, March 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jalics:1982:PCP

[Jal82] Paul J. Jalics. On the performance of COBOL programs in large vs. mini computers. Software— Practice and Experience, 12(2):103–114, February 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jalote:1987:SIA

[Jal87] Pankaj Jalote. Synthesizing implementations of abstract data types from axiomatic specifications. Software—Practice and Experience, 17(11):847–858, November 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jamieson:1980:LEI

[Jam80] M. J. Jamieson. Letter to the Editor: Integer division. Software—Practice and Experience, 10(4): 333, April 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jarvis:1975:TSW

[Jar75]  John F. Jarvis.  Two

simple windowing algorithms. Software—Practice and Experience, 5(2): 115–122, April/June 1975.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ju:1984:IFT

[JB84] J. B. Ju and K. D. Baker. Implementing a file transfer protocol between PDP11 and PRIME 550 computers. Software—Practice and Experience, 14 (11):1009–1019, November 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Joisha:2007:TSM

[JB07] Pramod G. Joisha and Prithviraj Banerjee. A translator system for the MATLAB language. Software—Practice and Experience, 37(5):535–578, April 25, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jones:1979:PHL

[JBCB79] Douglas Jones, A. B. Baskin, Thomas Chen, and Louis Bloomfield. Programs as higher level subroutines. Software—Practice and Experience, 9(2): 149–155, February 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Joseph:2019:SCR

[JC19] Christina Terese Joseph and K. Chandrasekaran.

Straddling the crevasse: a review of microservice software architecture foundations and recent advancements. Software—Practice and Experience, 49(10): 1448–1484, October 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jimenez:2022:DAB

[JC22] Patricia Jiménez and Rafael Corchuelo. On the design of an advanced business rule engine. Software—Practice and Experience, 52(10):2097–2126, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jesshope:1985:IPE

[JCL85] C. R. Jesshope, M. J.

Crawley, and G. L. Lovegrove. An intelligent Pascal editor for a graphical oriented workstation. Software—Practice and Experience, 15(11):1103–1119, November 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jann:2004:DRC

[JDBP04] Joefon Jann, Niteesh Dubey, R. S. Burugula, and Pratap Pattnaik. Dynamic reconfiguration of CPU and WebSphere on IBM pSeries servers. Software—Practice and Experience, 34(13): 1257–1272, November 10, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jimenez-Diaz:2012:RPV

[JDGCGA12] Guillermo Jimenez-Diaz, Pedro A. Gonzalez-Calero, and Mercedes Gomez-Albarran. Role-play virtual worlds for teaching objectoriented design: the ViR-Play development experience. Software—Practice and Experience, 42(2):235– 253, February 2012. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jordan:2006:SJA

[JDJ^{+}06] Mick Jordan, Laurent Daynès, Marcin Jarzab, Ciarán Bryce, and Grzegorz Czajkowski. Scaling J2EE^{T M} application servers with the multitasking virtual machine. Software—Practice and Experience, 36(6):557–580, May 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jann:2008:EMP

[JDPB08] Joefon Jann, Niteesh Dubey, Pratap Pattnaik, and R. Sarma Burugula. Effectiveness of multiple pageable page sizes for commercial applications. Software—Practice and Experience, 38(12):1261–1272, October ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jegado:1983:RAD

[Jeg83] M. Jegado. Recoverability aspects of a distributed file system. Software—Practice and Experience, 13(1):33– 44, January 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

James:1999:DAB

[JEG99] Perry R. James, Markus Endler, and Marie-Claude Gaudel. Development of an atomic-broadcast protocol using LOTOS. Software—Practice and Experience, 29(8):699–719, July 10, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=62004389;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=62004389&PLACEBO=IE.
pdf.

Jenkins:1989:QPI

[Jen89] Michael A. Jenkins. Q’Nial; a portable interpreter for the nested interactive array language, Nial. Software —Practice and Experience, 19(2):111–126, February 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jin:2021:DDA

[JFZ^{+}21] Yanrui Jin, Qiang Feng, Xiping Zhang, Peili Lu, Jiaqi Shen, Yihui Tu, and Zhiquan Wu. Deep domain adversarial residual neural network for sustainable wind turbine cyberphysical system fault diagnosis. Software—Practice and Experience, 51 (11):2128–2142, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jablonowski:1989:GTM

[JG89] David Jablonowski and Vincent A. Guarna, Jr. GMB: a tool for manipulating and animating graph data structures. Software—Practice and Experience, 19(3):283–301, March 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jeffery:1994:FEM

[JG94] Clinton L. Jeffery and Ralph E. Griswold. A framework for execution monitoring in Icon. Software—Practice and Experience, 24(11):1025–1049, November 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jara:2015:BDS

[JGB15] Antonio J. Jara, Dominique Genoud, and Yann Bocchi. Big data for smart cities with KNIME a real experience in the Smart-Santander testbed. Software—Practice and Experience, 45(8):1145–1160, August 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jindal:2021:FDN

[JGC^{+}21] Anshul Jindal, Michael Gerndt, Mohak Chadha, Vladimir Podolskiy, and Pengfei Chen. Function delivery network: Extending serverless computing for heterogeneous platforms. Software—Practice and Experience, 51 (9):1936–1963, September 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jones:1989:PPR

[JGR89] Leslie P. Jones, Edward W. Gassie, Jr., and Sridhar Radhakrishnan. PORTREP: a portable repeated string finder. Software—Practice and Experience, 19(1):63–77, January 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Janakiram:2008:OOW

[JGS^{+}08] D. Janakiram, Ashok Gunnam, N. Suneetha, Vineet Rajani, and K. Vinay Kumar Reddy. Objectoriented wrappers for the Linux kernel. Software— Practice and Experience, 38(13):1411–1427, November 10, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jung:2021:TPI

[JGSG^{+}21] Kumseok Jung, Julien Gascon-Samson, Shivanshu Goyal, Armin Rezaiean-Asel, and Karthik Pattabiraman. ThingsMigrate: Platform-independent migration of stateful JavaScript Internet of Things applications. Software—Practice and Experience, 51(1):117– 155, January 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jeffery:1995:AGH

[JGT95] Clinton L. Jeffery, Ralph E. Griswold, and Gregg M. Townsend. Adding graphics to a high-level programming language. Software —Practice and Experience, 25(6):637–655, June 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Johnson:2003:ENS

[JH03] C. W. Johnson and C. M.

Holloway. The ESA/NASA SOHO mission interruption: using the STAMP accident analysis technique for a software related mishap. Software— Practice and Experience, 33(12):1177–1198, October 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jammal:2019:GIT

[JHKS19] Manar Jammal, Hassan Hawilo, Ali Kanso, and Abdallah Shami. Generic input template for cloud simulators: a case study of CloudSim. Software —Practice and Experience, 49(5):720–747, May 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

James:1980:MPI

[JI80] E. B. James and D. Ireland.

Microcomputers as protective interfaces in computing networks. Software— Practice and Experience, 10(12):953–958, December 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jaber:2021:ESM

[JI21] Alaa Shawqi Jaber and Ali Kadhum Idrees. Energysaving multisensor data sampling and fusion with decision-making for monitoring health risk using WBSNs. Software—Practice and Experience, 51(2): 271–293, February 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jia:1997:IRM

[Jia97] Weijia Jia. Implementation of a reliable multicast protocol. Software—Practice and Experience, 27(7): 813–850, July 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7319;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7319&
PLACEBO=IE.pdf.

Jawaddi:2022:RMA

[JJI22] Siti Nuraishah Agos Jawaddi, Muhammad Hamizan Johari, and Azlan Ismail. A review of microservices autoscaling with formal verification perspective. Software—Practice and Experience, 52(11):2476–2495, November 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jan:2012:FEF

[JJK^{+}12] Mathieu Jan, Christophe Jouvray, Fabrice Kordon, Antonio Kung, Jimmy Lalande, Frédéric Loiret, Juan Navas, Laurent Pautet, Jacques Pulou, Ansgar Radermacher, and Lionel Seinturier. Flex-eWare: a flexible model driven solution for designing and implementing embedded distributed systems. Software—Practice and Experience, 42(12):1467–1494, December 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jayaraman:2017:CVJ

[JJL17] S. Jayaraman, B. Jayaraman, and D. Lessa. Compact visualization of Java program execution. Software—Practice and Experience, 47(2):163–191, February 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jones:1983:XSE

[JK83] Rondall E. Jones and David K. Kahaner. XER-ROR, the SLATEC errorhandling package. Software—Practice and Experience, 13(3):251–257, March 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jantz:2014:AAF

[JK14] Michael R. Jantz and Prasad A. Kulkarni. Analyzing and addressing false interactions during compiler optimization phase ordering. Software—Practice and Experience, 44(6):643– 679, June 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jaring:2004:RVF

[JKB04] M. Jaring, R. L. Krikhaar, and J. Bosch. Representing variability in a family of MRI scanners. Software—Practice and Experience, 34(1):69–100, January 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jayasekara:2022:OCB

[JKH22] Sachini Jayasekara, Shanika Karunasekera, and Aaron Harwood. Optimizing checkpoint-based faulttolerance in distributed stream processing systems: Theory to practice. Software—Practice and Experience, 52(1):296–315, January 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jankowitz:1985:PHL

[JKRS85] Hugo T. Jankowitz, Paul W. Kilfoil, Ilana Rabkin, and Stephen R. Schach. A portable high-level database debugger and software performance monitor. Software—Practice and Experience, 15(6):523–538, June 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Joyce:1974:RUA

[JKW74] J. D. Joyce, V. Kevin, and M. Whitney. Reuse of archival storage media. Software—Practice and Experience, 4(2):191, April/June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Johnston:1980:SSS

[JL80] D. B. Johnston and A. M.

Lister. Software science and student programs. Software—Practice and Experience, 10(2):159– 160, February 1980. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Johnston:1981:NSS

[JL81] D. B. Johnston and A. M.

Lister. A note on the software science length equation. Software—Practice and Experience, 11(8):875– 877, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jones:1991:MFL

[JL91] Simon L. Peyton Jones and David Lester. A modular fully-lazy lambda lifter in HASKELL. Software —Practice and Experience, 21(5):479–506, May 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Javed:2022:SAR

[JLBR22] Amir Javed, Mike Lakoju, Pete Burnap, and Omer Rana. Security analytics for real-time forecasting of cyberattacks. Software—Practice and Experience, 52(3):788–804, March 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jensen:1979:SUC

[JLR79] Jørn Jensen, Søren Lauesen, and A. P. Ravn. Synchronization under a commercial operating system. Software—Practice and Experience, 9(9):729– 740, September 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Johansen:2002:TR

[JLvR^{+}02] Dag Johansen, Kåre J. Lauvset, Robbert van Renesse, Fred B. Schneider, Nils P. Sudmann, and Kjetil Jacobsen. A TACOMA retrospective. Software—Practice and Experience, 32(6):605–619, May 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
93513616/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=93513616&PLACEBO=IE.
pdf.

Jiang:2009:MES

[JLZ09] Yuan Jiang, Ming Li, and Zhi-Hua Zhou. Mining extremely small data sets with application to software reuse. Software— Practice and Experience, 39(4):423–440, March 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jansen:2008:SVC

[JM08] Sam Jansen and Anthony McGregor. Static virtualization of C source code. Software—Practice and Experience, 38(4):397– 416, April 10, 2008. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jump:2010:DML

[JM10] Maria Jump and Kathryn S.

McKinley. Detecting memory leaks in managed languages with Cork. Software—Practice and Experience, 40(1):1–22, January ??, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Joshi:2003:FOJ

[JMM03] Rushikesh K. Joshi, Maureen Mascarenhas, and Yogesh Murarka. Filter objects for Java. Software —Practice and Experience, 33(6):509–522, May 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Johnsen:1978:SCT

[Joh78] Kari Johnsen. A simplification of code tables in a onepass PL/M compiler. Software—Practice and Experience, 8(4):399–405, July/ August 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Johnson:1979:TDS

[Joh79] Mark Scott Johnson. Translator design to support run-time debugging. Software—Practice and Experience, 9(12):1035–1041, December 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Johnson:1984:AEN

[Joh84] James L. Johnson. Anatomy of an educational network database system. Software—Practice and Experience, 14(8):739–754, August 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jokinen:1989:LIP

[Jok89]  Matti O. Jokinen.  A

language-independent prettyprinter. Software—Practice and Experience, 19(9): 839–856, September 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jones:1971:MFI

[Jon71] W. R. Jones. A MACRO facility for interactive display. Software—Practice and Experience, 1(2):159– 166, April/June 1971. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jones:1972:NIM

[Jon72] Bush Jones. A note on internal merging. Software—Practice and Experience, 2(3):241–243, July/ September 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jones:1974:BRB

[Jon74] Karen Sparck Jones. Book review: Advances in computers, Ed. F. L. Alt and M. Rubinoff, Volume 11 Guest Ed. M. C. Yovits, Academic Press, New York, 1971. No. of pages: 390. Price: $18.50 ( £8.65). Software—Practice and Experience, 4(3):303– 304, July/September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jones:1983:ALO

[Jon83] Douglas W. Jones. Assembly language as object code. Software—Practice and Experience, 13(8):715– 725, August 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jones:1985:YSE

[Jon85] Simon L. Peyton Jones. YACC in SASL — an exercise in functional programming. Software—Practice and Experience, 15(8):807– 820, August 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jordan:1978:SPP

[Jor78] M. J. Jordan. SLP: a paged processor for compact lists. Software—Practice and Experience, 8(3): 285–300, May/June 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jordan:1990:ESS

[Jor90] J. E. Jordan. Experiences structuring software in a periodic real-time environment. Software—Practice and Experience, 20(7):707– 718, July 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Joslin:1979:CAP

[Jos79] D. A. Joslin. A case for acquiring Pascal. Software— Practice and Experience, 9 (8):691–692, August 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See comment [Ano80b].

Joslin:1980:LES

[Jos80] D. A. Joslin. Letters to the Editor: Some defensive comments on ‘a case for acquiring Pascal’. Software—Practice and Experience, 10(7):590, July 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jones:1974:STE

[JP74] D. A. Jones and N. J. Partington. A software teletype exchange. Software—Practice and Experience, 4(2): 163–169, April/June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jones:1979:LEP

[JP79] Peter R. Jones and Arthur Pyster. Letters to the Editor: Programming language standards. Software —Practice and Experience, 9(3):249–250, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jammalamadaka:2022:EMI

[JP22]  Kiran  Jammalamadaka

and Nikhat Parveen. Equivalent mutant identification using hybrid wavelet convolutional rain optimization. Software—Practice and Experience, 52(2):576– 593, February 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jayaraman:2017:ASM

[JPG^{+}17] Prem Prakash Jayaraman, Charith Perera, Dimitrios Georgakopoulos, Schahram Dustdar, Dhavalkumar Thakker, and Rajiv Ranjan. Analytics-as-a-service in a multi-cloud environment through semanticallyenabled hierarchical data processing. Software— Practice and Experience, 47(8):1139–1156, August 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jarvi:2003:LLU

[JPL03] Jaakko Järvi, Gary Powell, and Andrew Lumsdaine. The Lambda Library: unnamed functions in C++. Software—Practice and Experience, 33(3): 259–291, March 2003. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jamshidi:2017:PBM

[JPM17] Pooyan Jamshidi, Claus Pahl, and Nabor C. Mendonça. Pattern-based multi-cloud architecture migration. Software— Practice and Experience, 47(9):1159–1184, September 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Joskowski:2023:SSC

[JPM23] Andrzej Joskowski, Adam Przybyłek, and Bartosz Marcinkowski. Scaling scrum with a customized nexus framework: a report from a joint industry– academia research project. Software—Practice and Experience, 53(7):1525– 1542, July 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Johnson:1992:RTF

[JR92] Paul A. Johnson and Ceri Rees. Reusability through fine-grain inheritance. Software—Practice and Experience, 22(12):1049–1068, December 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jimenez:2020:SMT

[JRGC20] Patricia Jiménez, Juan C. Roldán, Fernando O. Gallego, and Rafael Corchuelo. On the synthesis of metadata tags for HTML files. Software—Practice and Experience, 50(12):2169– 2192, December 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jeong:2010:KKA

[JSC^{+}10] Jinkyu Jeong, Euiseong Seo, Jeonghwan Choi, Hwanju Kim, Heeseung Jo, and Joonwon Lee. KAL: kernel-assisted noninvasive memory leak tolerance with a generalpurpose memory allocator. Software—Practice and Experience, 40(8):605– 625, July 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jacobson:2022:BST

[JSKB22] Ivar Jacobson, Jeff Sutherland, Brian Kerr, and Barbora Buhnova. Better Scrum through Essence. Software—Practice and Experience, 52(6):1531– 1540, June 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Javanmardi:2021:FJF

[JSPP21] Saeed Javanmardi, Mohammad Shojafar, Valerio Persico, and Antonio Pescapè. FPFTS: a joint fuzzy particle swarm optimization mobility-aware approach to fog task scheduling algorithm for Internet of Things devices. Software—Practice and Experience, 51 (12):2519–2539, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Junior:2018:CAA

[JSRM18] Ronaldo Gonçalves Junior, Americo Sampaio, Tiago Rolim, and Nabor C. Mendonça. Cloud application architecture appraiser (CA3): a multicriteria approach and tool for assessing cloud deployment options based on nonfunctional requirements. Software—Practice and Experience, 48(12):2278–2301, December 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Junger:2000:ASB

[JT00] Michael Jünger and Stefan Thienel. The ABA-CUS system for branchand-cut-and-price algorithms in integer programming and combinatorial optimization. Software—Practice and Experience, 30(11):1325–1352, September 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72517077/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72517077&PLACEBO=IE.
pdf.

Jiang:2011:AME

[JTG^{+}11] Bo Jiang, T. H. Tse, Wolfgang Grieskamp, Nicolas Kicillof, Yiming Cao, Xiang Li, and W. K. Chan. Assuring the model evolution of protocol software specifications by regression testing process improvement. Software—Practice and Experience, 41 (10):1073–1103, September 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jokinen:1996:CAS

[JTU96] Petteri Jokinen, Jorma Tarhio, and Esko Ukkonen. A comparison of approximate string matching algorithms. Software—Practice and Experience, 26(12):1439–1458, December 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.

com/cgi-bin/abstract?
ID=16782.

Joshi:1997:MFO

[JVR97] Rushikesh K. Joshi, N. Vivekananda, and D. Janaki Ram. Message filters for objectoriented systems. Software—Practice and Experience, 27(6):677–699, June 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7311;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7311&
PLACEBO=IE.pdf.

Jones:1975:TDS

[JW75] Anita K. Jones and William A.

Wulf. Towards the design of secure systems. Software—Practice and Experience, 5(4):321–336, October/December 1975. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jeffrey:2011:IBM

[JWTG11] Dennis Jeffrey, Yan Wang, Chen Tian, and Rajiv Gupta. Isolating bugs in multithreaded programs using execution suppression. Software—Practice and Experience, 41(11): 1259–1288, October 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jiang:2021:CRA

[JXG^{+}21] Wenjing Jiang, Zichen Xu, Cuiying Gao, Jingyun Gu, and Yuhao Wang. Cost risk analysis for instance recommendation in a sustainable cloud-cyber-physical system framework. Software—Practice and Experience, 51(11):2317–2336, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Johnson:1993:PHP

[JZ93] David B. Johnson and Willy Zwaenepoel. The Peregrine high-performance RPC system. Software— Practice and Experience, 23(2):201–221, February 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jiang:2002:TDS

[JZ02] Song Jiang and Xiaodong Zhang. TPF: a dynamic system thrashing protection facility. Software—Practice and Experience, 32(3):295–318, March 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        90010509/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?

ID=90010509&PLACEBO=IE.
pdf.

Janik:2010:AMA

[JZ10] Arkadiusz Janik and Krzysztof Zielinski. Adaptability mechanisms for autonomic system implementation with AAOP. Software —Practice and Experience, 40(3):209–223, March ??, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ji:2020:AAD

[JZLP20] Cun Ji, Xiunan Zou, Shijun Liu, and Li Pan. ADARC: an anomaly detection algorithm based on relative outlier distance and biseries correlation. Software—Practice and Experience, 50(11):2065–2081, November 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kulkarni:1987:IEF

[KA87] K. G. Kulkarni and Malcolm P. Atkinson. Implementing an extended functional data model using PS-Algol. Software—Practice and Experience, 17(3):171– 185, March 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Khwaja:2013:TDP

[KA13] Salman Khwaja and Mohammad Alshayeb. Towards design pattern definition language. Software —Practice and Experience, 43(7):747–757, July 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kaur:2022:SAT

[KA22] Amanjot Kaur and Nitin Auluck. Scheduling algorithms for truly heterogeneous hierarchical fog networks. Software—Practice and Experience, 52 (11):2411–2438, November 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kahrs:1995:HI

[Kah95] Mark Kahrs. The heart of IGES. Software—Practice and Experience, 25(8):935– 946, August 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kantorowitz:1997:AST

[Kan97] Eliezer Kantorowitz. Algorithm simplification through object orientation. Software—Practice and Experience, 27(2):173–183, February 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7286;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7286&
PLACEBO=IE.pdf.

Kang:2018:FCI

[Kan18]  Pilsung Kang.  Func-

tion call interception techniques. Software—Practice and Experience, 48 (3):385–401, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2501.

Kanda:2023:EFD

[KAO23] Shunsuke Kanda, Koichi Akabe, and Yusuke Oda. Engineering faster doublearray Aho–Corasick automata. Software—Practice and Experience, 53(6): 1332–1361, June 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kapitza:2013:EDA

[Kap13]  Rüdiger Kapitza.  Ed-

itorial: Distributed applications and interoperable systems (extended papers from DAIS’10). Software—Practice and Experience, 43(10):1143–1144, October 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Karn:1976:PCT

[Kar76] Ronald Karn. Parallel computing: Timing for Illiac IV. Software—Practice and Experience, 6(4):579–580, October 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Karus:2014:XDP

[Kar14] S. Karus. XML development with plug-ins as a service. Software—Practice and Experience, 44(3):335– 352, March 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Karthikeyan:2021:ELB

[Kar21] Periyasami Karthikeyan. An efficient load balancing using seven stone game optimization in cloud computing. Software—Practice and Experience, 51(6): 1242–1258, June 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kakkad:2014:CSP

[KAS^{+}14] Vasvi Kakkad, Saeed Attar, Andrew E. Santosa, Alan Fekete, and Bernhard Scholz. Curracurrong: a stream programming environment for wireless sensor networks. Software—Practice and Experience, 44(2): 175–199, February 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Karaa:2016:ABC

[KAS^{+}16] Wahiba Ben Abdessalem Karaa, Zeineb Ben Azzouz, Aarti Singh, Nilanjan Dey, Amira S. Ashour, and Henda Ben Ghazala. Automatic builder of class diagram (ABCD): an application of UML generation from functional requirements. Software—Practice and Experience, 46 (11):1443–1458, November 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Katzenelson:1971:DMS

[Kat71] Jacob Katzenelson. Documentation and the management of a software project — a case study. Software —Practice and Experience, 1(2):147–157, April/June 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Katzenelson:1983:HLP

[Kat83a] Jacob Katzenelson. Higher level programming and data abstractions — a case study using enhanced C. Software—Practice and Experience, 13(7):577– 595, July 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Katzenelson:1983:IEC

[Kat83b] Jacob Katzenelson. Introduction to enhanced C (EC). Software—Practice and Experience, 13(7):551– 576, July 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Katajainen:2017:AOI

[Kat17] Jyrki Katajainen. All-in-one implementation framework for binary heaps. Software—Practice and Experience, 47(4):523–558, April 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kawai:1979:LSS

[Kaw79]  Satoru Kawai.  Lat-

tice structure segmentation of ALGOL-like programs. Software—Practice and Experience, 9(6):485– 498, June 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kawai:1980:SSL

[Kaw80] Satoru Kawai. A semiblock structure for low-level languages. Software—Practice and Experience, 10(1):11– 19, January 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ko:1999:TCS

[KAYH^{+}99] Lo Ko, Nagham Al-Yaqoubi, Christopher Healy, Emily Ratliff, Robert Arnold, David Whalley, and Marion Harmon. Timing constraint specification and analysis. Software—Practice and Experience, 29(1):77–98, January 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55002656;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55002656&PLACEBO=IE.
pdf.

Kamal:2013:UPP

[KAZ13] Ahmad Waqas Kamal, Paris Avgeriou, and Uwe Zdun. The use of pattern participants relationships for integrating patterns: a controlled experiment. Software—Practice and Experience, 43(7):807– 833, July 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2006:AFB

[KB06] Jungyoon Kim and Doo Hwan Bae. An approach to feature-based software construction for enhancing maintainability. Software —Practice and Experience, 36(9):923–948, July 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2020:REM

[KBB^{+}20] Taewoo Kim, Alexander Behm, Michael Blow, Vinayak Borkar, Yingyi Bu, Michael J. Carey, Murtadha Hubail, Shiva Jahangiri, Jianfeng Jia, Chen Li, Chen Luo, Ian Maxon, and Pouria Pirzadeh. Robust and efficient memory management in Apache AsterixDB. Software—Practice and Experience, 50(7): 1114–1151, July 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Karasik:2005:GUI

[KBBS05] Dmitry Karasik, Anton Berezin, Vadim Belman, and Jon Sporring. A graphical user-interface and an image processing toolkit for Perl. Software—Practice and Experience, 35(5):467– 490, April 25, 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Koster:2003:TTI

[KBH^{+}03] Rainer Koster, Andrew P. Black, Jie Huang, Jonathan Walpole, and Calton Pu.

Thread transparency in information flow middleware. Software—Practice and Experience, 33(4):321– 349, April 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Krauter:2002:TSG

[KBM02] Klaus Krauter, Rajkumar Buyya, and Muthucumaru Maheswaran. A taxonomy and survey of grid resource management systems for distributed computing. Software—Practice and Experience, 32(2):135– 164, February 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
88511120/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=88511120&PLACEBO=IE.
pdf.

Koziolek:2020:AII

[KBPM^{+}20] Heiko Koziolek, Andreas Burger, Marie Platenius-Mohr, Julius Rückert, Francisco Mendoza, and Roland Braun. Automated industrial IoTdevice integration using the OpenPnP reference architecture. Software—Practice and Experience, 50(3): 246–274, March 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kounavis:2005:PDP

[KCCV05] Michael E. Kounavis, Andrew T. Campbell, Stephen T. Chou, and John Vicente. Programming the data path in network processor-based routers. Software—Practice and Experience, 35 (11):1041–1078, September 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kirschnick:2012:TAD

[KCG^{+}12] Johannes Kirschnick, Jose M. Alcaraz Calero, Patrick Goldsack, Andrew Farrell, Julio Guijarro, Steve Loughran, Nigel Edwards, and Lawrence Wilcock. Towards an architecture for deploying elastic services in the cloud. Software—Practice and Experience, 42(4): 395–408, April 2012. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2007:DRR

[KCH07] Hyoseung Kim, Hojung Cha, and Rhan Ha. Dynamic refresh-rate scaling via frame buffer monitoring for power-aware LCD management. Software—Practice and Experience, 37(2): 193–206, February 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kemball:2008:CBF

[KCH08] A. J. Kemball, R. M.

Crutcher, and R. Hasan. A component-based framework for radio-astronomical imaging software systems. Software—Practice and Experience, 38(5):493–507, April 25, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kaya:2023:DIB

[KCK^{+}23] M. Cagri Kaya, Erkin Cilden, Alper Karamanlioglu, Furkan Murat, Ali H. Dogru, and Halit Oguztuzun. Data interoperability between DDS and HLA through a dynamically reconfigurable gateway. Software—Practice and Experience, 53(11):2200–2221, November 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Krzywaniak:2022:DDE

[KCP22] Adam Krzywaniak, Paweł Czarnul, and Jerzy Proficz. DEPO: a dynamic energyperformance optimizer tool for automatic power capping for energy efficient high-performance computing. Software—Practice and Experience, 52 (12):2598–2634, December 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ke:2020:CAO

[KCS^{+}20] Hengjin Ke, Dan Chen, Tejal Shah, Xianzeng Liu, Xinhua Zhang, Lei Zhang, and Xiaoli Li. Cloud-aided online EEG classification system for brain healthcare: a case study of depression evaluation with a lightweight CNN. Software —Practice and Experience, 50(5):596–610, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kardas:2012:DIM

[KCYY12] Geylani Kardas, Moharram Challenger, Suleyman Yildirim, and Ali Yamuc. Design and implementation of a multiagent stock trading system. Software—Practice and Experience, 42(10):1247–1273, October 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Knight:1983:DSP

[KD83] John C. Knight and Douglas D. Dunlop. On the design of a special-purpose scientific programming language. Software—Practice and Experience, 13(10): 893–907, October 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kam:2013:LST

[KD13] Ben W. Y. Kam and Thomas R. Dean. Linguistic security testing for text communication protocols. Software—Practice and Experience, 43 (9):1103–1120, September 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Krishnan:2020:SNS

[KDA20] Prabhakar Krishnan, Subhasri Duttagupta, and Krishnashree Achuthan. SDN/NFV security framework for fog-to-things computing infrastructure. Software—Practice and Experience, 50(5):757–800, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Karshmer:1983:NMS

[KDP83] Arthur J. Karshmer, Dirk J.

Depree, and James Phelan. The New Mexico State University ringstar system: a distributed UNIX environment. Software—Practice and Experience, 13(12):1157–1168, December 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kemmerer:1985:UUB

[KE85] Richard A. Kemmerer and Steven T. Eckmann. UNI-SEX: A UNIX-based symbolic EXecutor for Pascal. Software—Practice and Experience, 15(5):439– 458, May 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kearns:1991:ERE

[Kea91a] Steven M. Kearns. Extending regular expressions with context operators and parse extraction. Software—Practice and Experience, 21(8):787– 804, August 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kearns:1991:T

[Kea91b] Steven M. Kearns. TLex. Software—Practice and Experience, 21(8):805–821, August 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Krishnaraj:2021:ERT

[KEL^{+}21] N. Krishnaraj, Mohamed Elhoseny, E. Laxmi Lydia, K. Shankar, and Omar Aldabbas. An efficient radix trie-based semantic visual indexing model for large-scale image retrieval in cloud environment. Software—Practice and Experience, 51(3):489–502, March 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kent:1977:BRB

[Ken77] P. Kent. Book review: Statistical package for the social sciences (second edition), N. H. Nie, C. H. Hull, J. G. Jenkins, K. Steinbrenner and D. H. Bent, McGraw-Hill, New York, 1975. No. of pages: 675. price: £5.90. Software— Practice and Experience, 7 (2):300–301, March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kent:1990:E

[Ken90] Christopher A. Kent. Editorial. Software—Practice and Experience, 20(S2):S1– S3, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kernighan:1975:RPR

[Ker75] Brian W. Kernighan. RAT-FOR — a preprocessor for a rational Fortran. Software—Practice and Experience, 5(4):395–406, October/December 1975. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kerr:1980:FEF

[Ker80] John D. Kerr. FORALL: An extensible Fortran system for conversationally accessing subroutine libraries. Software—Practice and Experience, 10(11): 889–896, November 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kernighan:1982:PLT

[Ker82a] Brian W. Kernighan. PIC: a language for typesetting graphics. Software— Practice and Experience, 12(1):1–21, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.4380120102.

Kerridge:1982:FIC

[Ker82b] Jon M. Kerridge. A FOR-TRAN implementation of Concurrent Pascal. Software—Practice and Experience, 12(1):45–55, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kermia:2017:SAE

[Ker17] Omar Kermia. Schedulability analysis and efficient scheduling of rate constrained messages in the TTEthernet protocol. Software—Practice and Experience, 47(11):1485–1499, November 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Keyes:1992:CBS

[Key92] Jessica Keyes. Changing the basics in software development. Software Magazine, 12(1):38–??, January 1992. CODEN SMWMEQ. ISSN 0897-8085.

Kendall:2002:DIE

[KF02] H. Richard Kendall and Vincent W. Freeh. The design and implementation of the exported procedure call. Software—Practice and Experience, 32(1):83– 98, January 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
88013194/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=88013194&PLACEBO=IE.
pdf.

Kaiser:1988:RDI

[KFJS88] Gail E. Kaiser, Peter H.

Feiler, Fahimeh Jalili, and Johann H. Schlichter. A retrospective on DOSE: An interpretive approach to structure editor generation. Software—Practice and Experience, 18(8):733– 748, August 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kanda:2018:PRM

[KFMF18] Shunsuke Kanda, Yuma Fujita, Kazuhiro Morita, and Masao Fuketa. Practical rearrangement methods for dynamic doublearray dictionaries. Software —Practice and Experience, 48(1):65–83, January 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Knight:1995:UMC

[KG95a] Jon Knight and Steve Guest. Using multicast communications to distribute code and data in Wide Area Networks. Software—Practice and Experience, 25(5):563–577, May 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Korfhage:1995:HLE

[KG95b] Willard Korfhage and Arthur P. Goldberg. Hermes language experiences. Software—Practice and Experience, 25(4):389–402, April 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kowalski:2018:FRM

[KG18] Tomasz M. Kowalski and Szymon Grabowski. Faster range minimum queries.

Software—Practice and Experience, 48(11):2043– 2060, November 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Khorsand:2018:FAA

[KGAR18] Reihaneh Khorsand, Mostafa Ghobaei-Arani, and Mohammadreza Ramezanpour. FAHP approach for autonomic resource provisioning of multitier applications in cloud computing environments. Software—Practice and Experience, 48(12):2147–2173, December 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Khorsand:2019:SLF

[KGAR19] Reihaneh Khorsand, Mostafa Ghobaei-Arani, and Mohammadreza Ramezanpour. A self-learning fuzzy approach for proactive resource provisioning in cloud environment. Software—Practice and Experience, 49(11):1618–1642, November 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kaagstrom:2006:AKA

[KGL06] Simon Kågström, Håkan Grahn, and Lars Lundberg. The application kernel approach — a novel approach for adding SMP support to uniprocessor operating systems. Software —Practice and Experience, 36(14):1563–1583, November 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:1996:PCS

[KGP96] Chinhyun Kim, Jean-Luc Gaudiot, and Włodek Proskurowski. Parallel computing with the Sisal applicative language: Programmability and performance issues. Software—Practice and Experience, 26(9):1025–1051, September 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16834.

Krintz:2001:ROD

[KGSC01] Chandra J. Krintz, David Grove, Vivek Sarkar, and Brad Calder. Reducing the overhead of dynamic compilation. Software—Practice and Experience, 31(8):717–738, July 10, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
78502910/START;   http:
//www3.interscience.wiley.

com/cgi-bin/fulltext?
ID=78502910&PLACEBO=IE.
pdf.

Kougiouris:1996:BMI

[KH96] Panos Kougiouris and Graham Hamilton. Buffer management issues in the implementation of fast interprocess communication. Software—Practice and Experience, 26(2):195– 211, February 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61004291.

Koehler:1997:CCC

[KH97] Brian Koehler and R. Nigel Horspool. CCC: a caching compiler for C. Software—Practice and Experience, 27(2):155–165, February 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7284;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7284&
PLACEBO=IE.pdf.

Krissinel:2004:CSI

[KH04] Evgeny B. Krissinel and Kim Henrick. Common subgraph isomorphism detection by backtracking search. Software—Practice and Experience, 34(6):591– 607, May 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kao:2007:SWA

[KH07] Shang-Juh Kao and I-Ching Hsu. Semantic Web approach to smart link generation for Web navigations. Software—Practice and Experience, 37(8):857– 879, July 10, 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kent:2012:ESI

[KH12] Kenneth B. Kent and Jérôme Hugues. Editorial to the Special Issue of Rapid System Prototyping’10. Software—Practice and Experience, 42(7): 779, July 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kan:2018:DSR

[KH18] Shuanglong Kan and Zhiqiu Huang. Detecting safetyrelated components in statecharts through traceability and model slicing. Software—Practice and Experience, 48(3):428– 448, March 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.

wiley.com/doi/abs/10.
1002/spe.2526.

Khayat:1986:PAT

[Kha86] M. G. Khayat. Printing Arabic text using dot matrix printers. Software— Practice and Experience, 16(2):165–172, February 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kumara:2019:SME

[KHC^{+}19] Indika Kumara, Jun Han, Alan Colman, Willem-Jan van den Heuvel, Damian A. Tamburri, and Malinda Kapuruge. SDSN@RT: a middleware environment for single-instance multitenant cloud applications. Software—Practice and Experience, 49(5):813– 839, May 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Khajeh-Hosseini:2012:CAT

[KHGSS12] Ali Khajeh-Hosseini, David Greenwood, James W. Smith, and Ian Sommerville. The Cloud Adoption Toolkit: supporting cloud adoption decisions in the enterprise. Software —Practice and Experience, 42(4):447–465, April 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2015:TTT

[KHH^{+}15] Taehong Kim, Myunggwon Hwang, Mi-Nyeong Hwang, Sa kwang Song, Do-Heon Jeong, and Hanmin Jung. Translation of technical terminologies between English and Korean based on textual big data. Software —Practice and Experience, 45(8):1115–1126, August 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kammer:2015:GFM

[KHHG15] D. Kammer, D. Henkens, C. Henzen, and R. Groh. Gesture formalization for multitouch. Software— Practice and Experience, 45(4):527–548, April 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kupsch:2017:ERB

[KHMB17] James A. Kupsch, Elisa Heymann, Barton Miller, and Vamshi Basupalli. Experience report: Bad and good news about using software assurance tools. Software—Practice and Experience, 47(1):143– 156, January 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2016:WTB

[KHOY16] Sol Kim, Kihong Heo, Hakjoo Oh, and Kwangkeun Yi. Widening with thresholds via binary search. Software—Practice and Experience, 46(10):1317–1328, October 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kausar:2020:SED

[KHS^{+}20] Samina Kausar, Muhammad Habib, Muhammad Yasir Shabir, Ata Ullah, Huahu Xu, Rashid Mehmood, Rongfang Bie, and Muhammad Shahid Iqbal. Secure and efficient data transfer using spreading and assimilation in MANET. Software—Practice and Experience, 50(11):2095–2109, November 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2024:DAF

[KHS^{+}24] Dasol Kim, Jiwoo Han, Siwoon Son, Myeong-Seon Gil, Yang-Sae Moon, and Heesun Won. Demeter: an automatic framework for data migration in open data lakes. Software— Practice and Experience, 54(5):721–743, May 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kurbalija:2009:CBC

[KIB09] Vladimir Kurbalija, Mirjana Ivanović, and Zoran Budimac. Case-based curve behaviour prediction. Software—Practice and Experience, 39(1):81–103, January ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kilgour:1971:EGS

[Kil71] A. C. Kilgour. The evolution of a graphics system for linked computers. Software—Practice and Experience, 1(3):259–268, July/ September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kilgour:1981:GNR

[Kil81] A. C. Kilgour. Generalized non-recursive traversal of binary trees. Software—Practice and Experience, 11(12):1299–1306, December 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kilpelainen:2012:UXP

[Kil12] Pekka Kilpeläinen. Using XQuery for problem solving. Software—Practice and Experience, 42 (12):1433–1465, December 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kilinc:2019:SBB

[Kil19] Deniz Kilinç. A Sparkbased Big Data analysis framework for realtime sentiment prediction on streaming data. Software—Practice and Experience, 49(9):1352–1364, September 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2002:HDR

[Kim02] Jeong Ah Kim. How to develop and to reuse the UniPDM framework. Software—Practice and Experience, 32(8):737–754, July 10, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
94519518/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=94519518&PLACEBO=IE.
pdf.

Kim:2015:DPB

[Kim15] Dae-Kyoo Kim. Design pattern based model transformation with tool support. Software—Practice and Experience, 45(4):473– 499, April 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kingslake:1971:TIS

[Kin71]  R. Kingslake.  TALK:

An interactive system for a small computer. Software—Practice and Experience, 1(4):391–401, October/December 1971. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kingston:1993:DIL

[Kin93] Jeffrey H. Kingston. The design and implementation of the Lout document formatting language. Software—Practice and Experience, 23(9):1001–1041, September 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kirkham:2007:RRS

[Kir07] C. C. Kirkham. Ranked reservoir sampling: an extension to the reservoir sampling algorithm. Software—Practice and Experience, 37(12):1285–1288, October 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kang:2011:CBH

[KJB11] Dongwon Kang, Jinhwan Jung, and Doo-Hwan Bae. Constraint-based human resource allocation in software projects. Software —Practice and Experience, 41(5):551–577, May 25, 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kvalnes:2010:SEC

[KJHG10] Åge Kvalnes, Dag Johansen, Pål Halvorsen, and Carsten Griwodz. Support for enterprise consolidation of I/O bound services. Software—Practice and Experience, 40(12):1035–1051, November 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kumar:2021:RMI

[KJVS21] Neeraj Kumar, Anish Jindal, Massimo Villari, and Satish Narayana Srirama. Resource management of IoT edge devices: Challenges, techniques, and solutions. Software—Practice and Experience, 51 (12):2357–2359, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Korn:1990:NDU

[KK90] David G. Korn and Eduardo Krell. A new dimension for the Unix(R) file system. Software—Practice and Experience, 20 (S1):S19–S34, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kruger:1997:ESW

[KK97] A. Kruger and W. F. Krajewski. Efficient storage of weather radar data. Software—Practice and Experience, 27(6):623–635, June 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7314;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7314&
PLACEBO=IE.pdf.

Kune:2016:SPA

[KKA^{+}16] Raghavendra Kune, Pramod Kumar Konugurthi, Arun Agarwal, Raghavendra Rao Chillarige, and Rajkumar Buyya. Survery papers: The anatomy of big data computing. Software—Practice and Experience, 46(1):79–105, January 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kune:2017:XEH

[KKA^{+}17] Raghavendra Kune, Pramod Kumar Konugurthi, Arun Agarwal, Raghavendra Rao Chillarige, and Rajkumar Buyya. XHAMI — extended HDFS and MapReduce interface for Big Data image processing applications in cloud computing environments. Software—Practice and Experience, 47(3):455–472, March 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kuzma:2023:FMM

[KKdC^{+}23] Braedy Kuzma, Ivan Korostelev, João P. L. de Carvalho, José E. Moreira, Christopher Barton, Guido Araujo, and José Nelson Amaral. Fast matrix multiplication via compiler-only layered data reorganization and intrinsic lowering. Software—Practice and Experience, 53(9):1793–1814, September 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kandalintsev:2017:FEP

[KKL17] Alexander Kandalintsev, Dzmitry Kliazovich, and Renato Lo Cigno. Freeze’nSense: estimation of performance isolation in cloud environments. Software—Practice and Experience, 47(6):831– 847, June 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kang:1999:FOE

[KKLL99] Kyo C. Kang, Sajoong Kim, Jaejoon Lee, and Kwanwoo Lee. Featureoriented engineering of PBX software for adaptability and reuseability.

Software—Practice and Experience, 29(10):875– 896, August 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=63001368;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=63001368&PLACEBO=IE.
                                        pdf.

Kornerup:1980:ICG

[KKM80] Peter Kornerup, Bent Bruun Kristensen, and Ole Lehrmann Madsen. Interpretation and code generation based on intermediate languages. Software—Practice and Experience, 10(8):635–658, August 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kawahito:2004:PRE

[KKN04] Motohiro Kawahito, Hideaki Komatsu, and Toshio Nakatani. Partial redundancy elimination for access expressions by speculative code motion. Software—Practice and Experience, 34(11):1065–1090, September 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2020:ARA

[KKPP20] Taeyoung Kim, Suntae Kim, Sooyong Park, and YoungBeom Park. Automatic recommendation to appropriate log levels. Software—Practice and Experience, 50(3):189–209, March 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kaaniche:2003:MLM

[KKR03] M. Kaâniche, K. Kanoun, and M. Rabah. Multi-level modeling approach for the availability assessment of ebusiness applications. Software—Practice and Experience, 33(14):1323–1341, November 25, 2003. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:1988:PGQ

[KKS88] Hyoung-Joo Kim, Henry F.

Korth, and Abraham Silberschatz. PICASSO: a graphical query language. Software—Practice and Experience, 18(3):169– 203, March 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Koskinen:2010:BPW

[KKS10] Johannes Koskinen, Markus Kettunen, and Tarja Systä. Behavioral profiles — a way to model and validate program behavior. Software—Practice and Experience, 40(8):701–733, July 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kantorowitz:1986:AGU

[KL86] E. Kantorowitz and H. Laor. Automatic generation of useful syntax error messages. Software—Practice and Experience, 16(7):627– 640, July 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kretz:2012:VCL

[KL12] Matthias Kretz and Volker Lindenstruth. Vc: A C++ library for explicit vectorization. Software—Practice and Experience, 42 (11):1409–1430, November 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kaser:2016:CBI

[KL16] Owen Kaser and Daniel Lemire. Compressed bitmap indexes: beyond unions and intersections. Software—Practice and Experience, 46(2):167–198, February 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Keiser:2021:VUL

[KL21] John Keiser and Daniel Lemire. Validating UTF-8 in less than one instruction per byte. Software —Practice and Experience, 51(5):950–964, May 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Keiser:2024:DJB

[KL24] John Keiser and Daniel Lemire. On-demand JSON: a better way to parse documents? Software—Practice and Experience, 54(6): 1074–1086, June 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Klint:1981:IT

[Kli81] Paul Klint. Interpretation techniques. Software —Practice and Experience, 11(9):963–973, September 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kang:1998:ASI

[KLLK98] Kyo C. Kang, Kwan W.

Lee, Ji Y. Lee, and Gerard J. Kim. ASADAL/ SIM: An incremental multilevel simulation and analysis tool for real-time software specifications. Software—Practice and Experience, 28(4):445–462, April 10, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1789;   http://www3.

                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=1789&
                                        PLACEBO=IE.pdf.

Kang:2020:PMT

[KLY20] JiHun Kang, JongBeom Lim, and HeonChang Yu. Partial migration technique for GPGPU tasks to prevent GPU memory starvation in RPC-based GPU virtualization. Software —Practice and Experience, 50(6):948–972, June 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kernighan:1979:UPE

[KM79] Brian W. Kernighan and John R. Mashey. The UNIX programming environment. Software—Practice and Experience, 9(1): 1–15, January 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Also in IEEE Computer, Vol. 14 (4), April 1981.

King:1983:DIC

[KM83] T. J. King and J. K. M. Moody. The design and implementation of CODD. Software—Practice and Experience, 13(1):67–78, January 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Klein:1989:PGS

[KM89] E. Klein and M. Martin. The parser generating system PGS. Software—Practice and Experience, 19(11):1015–1028, November 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Koskimies:1994:ASS

[KM94] Kai Koskimies and Erkki Mäkinen. Automatic synthesis of state machines from trace diagrams. Software—Practice and Experience, 24(7):643–658, July 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kiefer:2013:RDN

[KM13] Kurt E. Kiefer and Louise E.

Moser. Replay debugging of non-deterministic executions in the Kernel-based Virtual Machine. Software—Practice and Experience, 43(11):1261–1281, November 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kermarrec:1998:DIE

[KMB98] A.-M. Kermarrec, C. Morin, and M. Banâtre. Design, implementation and evaluation of ICARE: an efficient recoverable DSM. Software—Practice and Experience, 28(9):981– 1010, July 25, 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1821;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1821&
PLACEBO=IE.pdf.

Khan:2021:CBA

[KMB^{+}21] Osama Amir Khan, Saif U. R. Malik, Faizan M. Baig, Saif Ul Islam, Haris Pervaiz, Hassan Malik, and Syed Hassan Ahmed. A cache-based approach toward improved scheduling in fog computing. Software—Practice and Experience, 51(12):2360–2372, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kirby:1998:LRJ

[KMS98] Graham Kirby, Ron Morrison, and David Stemple. Linguistic reflection in Java. Software —Practice and Experience, 28(10):1045–1077, August 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=10007353;   http://
www3.interscience.wiley.

com/cgi-bin/fulltext?
ID=10007353&PLACEBO=IE.
pdf.

Kelly:1998:POO

[KMSS98] Andrew D. Kelly, Kim Marriott, Harald Søndergaard, and Peter J. Stuckey. A practical object-oriented analysis engine for CLP. Software—Practice and Experience, 28(2):199–224, February 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1776;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1776&
PLACEBO=IE.pdf.

Kon:2005:DIP

[KMY^{+}05] Fabio Kon, Jeferson Roberto Marques, Tomonori Yamane, Roy H. Campbell, and M. Dennis Mickunas. Design, implementation, and performance of an automatic configuration service for distributed component systems. Software —Practice and Experience, 35(7):667–703, June 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kramer:1988:ARS

[KN88] Jeff Kramer and Keng Ng. Animation of requirements specifications. Software—Practice and Experience, 18(8):749–774, August 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kuhl:1994:ORB

[KNC94] Frederick Kuhl, William Neal, and Howard Cohen. Object Request Broker: Foundation for distributed simulation. Software—Practice and Experience, 24(12):1145–1157, December 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Knott:1981:PME

[Kno81] Gary D. Knott. Procedures for managing extendible array files. Software—Practice and Experience, 11(1): 63–84, January 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Koskimies:1988:DLP

[KNPS88] Kai Koskimies, Otto Nurmi, Jukka Paakki, and Seppo Sippu. The design of a language processor generator. Software—Practice and Experience, 18(2):107– 135, February 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kakugawa:2001:GPF

[KNT^{+}01] H. Kakugawa, M. Nishikimi, N. Takahashi, S. Tomura, and K. Handa. A general purpose font module for multilingual application programs. Software—Practice and Experience, 31(15):1487–1508, December 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
86511587/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=86511587&PLACEBO=IE.
pdf.

Knuth:1971:ESF

[Knu71] Donald E. Knuth. An empirical study of FORTRAN programs. Software—Practice and Experience, 1(2): 105–133, April/June 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Knudsen:1984:EHS

[Knu84] Jørgen Lindskov Knudsen. Exception handling: a static approach. Software —Practice and Experience, 14(5):429–449, May 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Knuth:1988:ET

[Knu88] Donald E. Knuth. The errors of TEX. Technical Report STAN-CS-88-1223, Stanford University, Department of Computer Science, September 1988. ?? pp. See [Knu89].

Knuth:1989:ET

[Knu89] Donald E. Knuth. The errors of TEX. Software —Practice and Experience, 19(7):607–685, July 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). This is an updated version of [Knu88]. Reprinted with additions and corrections in [Knu92a, pp. 243–339].

Knuth:1992:LP

[Knu92a] Donald E. Knuth. Literate Programming. CSLI Lecture Notes Number 27. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1992. ISBN 0-937073-80-6 (paperback), 0-937073-81-4 (hardcover). xiii + 368 pp. LCCN QA76.6 .K644 1992. US$24.95 (paperback), US$59.95 (hardcover).

Knuutila:1992:EPP

[Knu92b] Timo Knuutila. Efficient Prolog programming. Software—Practice and Experience, 22(3):209–221, March 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Knudsen:2011:UIE

[Knu11] Svend E. Knudsen. Using independence to enable parallelism on multicore computers. Software —Practice and Experience, 41(4):393–402, April 10, 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Krogdahl:1986:ASS

[KO86] S. Krogdahl and K. A.

Olsen. Ada, as seen from Simula. Software—Practice and Experience, 16(8):689– 700, August 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

King:1991:FLS

[KO91] K. N. King and A. Jefferson Offutt. A Fortran language system for mutation-based software testing. Software —Practice and Experience, 21(7):685–718, July 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kobayashi:1977:SSI

[Kob77]  Makoto Kobayashi.  A

set of strategy-independent restructuring algorithms. Software—Practice and Experience, 7(5):585–594, September/October 1977.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Koopman:1987:IPF

[Koo87] Pieter W. M. Koopman. Interactive programs in a functional language: a functional implementation of an editor. Software— Practice and Experience, 17(9):609–622, September 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Koppler:1997:SAF

[Kop97] Rainer Koppler. A systematic approach to fuzzy parsing. Software—Practice and Experience, 27(6): 637–649, June 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7315;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7315&
PLACEBO=IE.pdf.

Korzeniowski:1992:MBF

[Kor92] Paul Korzeniowski. E-Mail becoming foundation for applications. Software Magazine, 12(1):95– ??, January 1992. CODEN SMWMEQ. ISSN 0897-8085.

Koskimies:1990:LRD

[Kos90]  Kai Koskimies.  Lazy

recursive descent parsing for modular language implementation. Software—Practice and Experience, 20(8):749–772, August 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kotula:1996:DPI

[Kot96]  Jeff Kotula.  Discover-

ing patterns: An industry report. Software—Practice and Experience, 26(11):1261–1276, November 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16774.

Kotula:2001:BIT

[Kot01] Jeffrey Kotula. Branded interface toolkits. Software—Practice and Experience, 31(12):1131–1142, October 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
85008993/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=85008993&PLACEBO=IE.
pdf.

Kourie:1987:DUP

[Kou87] D. G. Kourie. The design and use of a Prolog trace generator for CSP. Software—Practice and Experience, 17(7):423–438, July 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kowaltowski:1981:PPM

[Kow81] Tomasz Kowaltowski. Parameter passing mechanisms and run time data structures. Software— Practice and Experience, 11(7):757–765, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Knuth:1981:BPL

[KP81] Donald E. Knuth and Michael F. Plass. Breaking paragraphs into lines. Software—Practice and Experience, 11(11):1119–1184, November 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Koster:1990:RML

[KP90] Alexis Koster and D. Stott Parker. Regulation management and logic programming. Software —Practice and Experience, 20(1):79–107, January 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kannan:1994:CPG

[KP94] Sampath Kannan and Todd A. Proebsting. Correction to “Producing Good Code for the case Statement”, SPE 15: 1021–1024 (October 1985). Software—Practice and Experience, 24(2):233, February 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [Ber85a].

Koulopoulos:2002:PIB

[KPGH02] D. Koulopoulos, K. Papoutsis, G. Goulas, and E. Housos. PLEIADES: an Internet-based parallel/ distributed system. Software—Practice and Experience, 32(11):1035–1049, September 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
97519259/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=97519259&PLACEBO=IE.
pdf.

Kaubisch:1976:QP

[KPH76] W. H. Kaubisch, R. H.

Perrott, and C. A. R. Hoare. Quasiparallel programming. Software —Practice and Experience, 6(3):341–356, July/ September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2017:EEU

[KPJ^{+}17] Neunghoe Kim, Soojin Park, Dongwon Jeong, Mansoo Hwang, Sooyong Park, and Hoh Peter In. EURECA: Enduser requirements engineering with collaborative animation. Software—Practice and Experience, 47(7): 1001–1012, July 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kruger:2018:CAR

[KPK^{+}18] Jacob Krüger, Marcus Pinnecke, Andy Kenner, Christopher Kruczek, Fabian Benduhn, Thomas Leich, and Gunter Saake. Composing annotations without regret? Practical experiences using FeatureC. Software—Practice and Experience, 48 (3):402–427, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2525.

Katajainen:1986:SDC

[KPT86] Jyrki Katajainen, Martti Penttonen, and Jukka Teuhola. Syntax-directed compression of program files. Software—Practice and Experience, 16(3):269– 276, March 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2004:COV

[KPU04] Jinhwan Kim, Yunheung Paek, and Gangryung Uh. Code optimizations for a VLIW-style network processing unit. Software— Practice and Experience, 34(9):847–874, July 25, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2011:SRR

[KQZ^{+}11] Kwang-Hae (Kane) Kim, Jing Qian, Zhen Zhang, Qian Zhou, Kyung-Deok Moon, Jun-Hee Park, Kwang-Roh Park, and Doo-Hyun Kim. A scheme for reliable real-time messaging with bounded delays. Software—Practice and Experience, 41 (12):1387–1407, November/December 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Koskimies:1983:MSE

[KR83] Kai Koskimies and Kari-Jouko Räihä. Modelling of space-efficient one-pass translation using attribute grammars. Software— Practice and Experience, 13(2):119–129, February 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Knight:1985:SDT

[KR85] John C. Knight and James L. Russell. Software development tools for microcontrollers. Software —Practice and Experience, 15(5):459–472, May 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kragelund:1997:STP

[Kra97] Lars Vestergaard Kragelund. Solving a timetabling problem using hybrid genetic algorithms. Software—Practice and Experience, 27(10):1121–1134, October 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=7262;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=7262&
                                        PLACEBO=IE.pdf.

Kramer:2010:ADR

[Kra10] Thomas R. Kramer. Automatic detection and replacement of syntactic constructs causing shift/ reduce conflicts. Software —Practice and Experience, 40(5):387–404, April 25, 2010. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Kang:2021:PJT

[KRB21] Jonggu Kang, Duksan Ryu, and Jongmoon Baik. Predicting just-in-time software defects to reduce postrelease quality costs in the maritime industry. Software—Practice and Experience, 51(4):748–771, April 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Krishnamurthy:1990:E

[Kri90] Balachander Krishnamurthy.

Editorial. Software—Practice and Experience, 20 (S1):S1–S2, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Krishnamurthy:2004:E

[Kri04] Balachander Krishnamurthy.

Editorial. Software—Practice and Experience, 34(2): 101–102, February 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kallel:2021:IFC

[KRK21] Ameni Kallel, Molka Rekik, and Mahdi Khemakhem. IoT-fog-cloud based architecture for smart systems: Prototypes of autism and COVID-19 monitoring systems. Software—Practice and Experience, 51(1):91– 116, January 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kohlert:1993:IGM

[KRO93] Doug Kohlert, Ken Rodham, and Dan Olsen. Implementing a graphical multi-user interface toolkit. Software—Practice and Experience, 23(9): 981–999, September 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ko:2019:WSA

[KRR19] Yoonseok Ko, Xavier Rival, and Sukyoung Ryu. Weakly sensitive analysis for JavaScript objectmanipulating programs. Software—Practice and Experience, 49(5):840–884, May 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kronental:1981:LTP

[KRTW81] M. Kronental, J. W. Roberts, K. H. Timmesfeld, and I. C. Wand. The LTPL-E tasking proposals. Software—Practice and Experience, 11(1):85– 97, January 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kruijer:1982:MUO

[Kru82]  H. S. M. Kruijer.  A

multi-user operating system for transaction processing, written in Concurrent Pascal. Software —Practice and Experience, 12(5):445–454, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Klimek:2002:ERS

[KRZ02] Bartosz Klimek, Dominik Radziszowski, and Krzysztof Zieliński. An event routing service for distributed applications: computer telephony — a case study. Software—Practice and Experience, 32(15):1467–1483, December 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
99019928/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=99019928&PLACEBO=IE.
pdf.

Kriz:1980:EPC

[KS80] J. Kriz and H. Sandmayr. Extension of Pascal by coroutines and its application to quasi-parallel programming and simulation. Software—Practice and Experience, 10(10): 773–789, October 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kaiser:1982:DG

[KS82] Pavel Kaiser and Ivan Stetina. A dialogue generator. Software—Practice and Experience, 12(8):693– 707, August 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kerridge:1984:TSR

[KS84] Jon M. Kerridge and Dan Simpson. Three solutions for a robot arm controller using Pascal-Plus, occam and Edison. Software— Practice and Experience, 14(1):3–15, January 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kerridge:1986:CPP

[KS86] Jon M. Kerridge and Dan Simpson. Communicating parallel processes. Software —Practice and Experience, 16(1):63–86, January 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Katzenelson:1987:DPU

[KS87] Jacob Katzenelson and Avy Strominger. Debugging programs that use macro-oriented data abstractions. Software— Practice and Experience, 17(2):79–103, February 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Krishnamoorthy:1989:PTA

[KS89] Mukkai S. Krishnamoorthy and Ramesh Swaminathan. Program tools for algorithm animation. Software —Practice and Experience, 19(6):505–513, June 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Koenig:1995:FNC

[KS95] Andrew Koenig and Bjarne Stroustrup. Foundations for native C++ styles. Software—Practice and Experience, 25(S4):45–86, December 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kandogan:1998:EWD

[KS98] Eser Kandogan and Ben Shneiderman. Elastic windows: Design, implementation, and evaluation of multi-window operations. Software—Practice and Experience, 28(3): 225–248, March 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1779;   http://www3.
interscience.wiley.com/

cgi-bin/fulltext?ID=1779&
PLACEBO=IE.pdf.

Khanvilkar:2001:TIF

[KS01a] Shashank Khanvilkar and Sol M. Shatz. Tool integration for flexible simulation of distributed algorithms. Software—Practice and Experience, 31 (14):1363–1380, November 25, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
85515669/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=85515669&PLACEBO=IE.
pdf.

Klemm:2001:EJS

[KS01b] Reinhard Klemm and Navjot Singh. Enhancing Java server availability with JAS. Software—Practice and Experience, 31(7): 697–716, June 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
78003146/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=78003146&PLACEBO=IE.
pdf.

Karakoidas:2008:FJO

[KS08] Vassilios Karakoidas and Diomidis Spinellis. FIRE/ J — optimizing regular expression searches with generative programming. Software—Practice and Experience, 38(6):557–573, May ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kuperman:2010:ACH

[KS10] Benjamin A. Kuperman and Eugene H. Spafford. Audlib: a configurable, high-fidelity application audit mechanism. Software —Practice and Experience, 40(11):989–1005, October 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kourai:2020:FSC

[KS20] Kenichi Kourai and Kouta Sannomiya. Flexible service consolidation with nested virtualization and library operating systems. Software—Practice and Experience, 50(1):3– 21, January 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kumar:2024:FFI

[KS24] Aditya Kumar and Satish Narayana Srirama. FIDEL: Fog integrated federated learning framework to train neural networks. Software—Practice and Experience, 54(2):

186–207, February 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kaur:2022:DQL

[KSBL22] Avinash Kaur, Parminder Singh, Ranbir Singh Batth, and Chee Peng Lim. Deep-Q learning-based heterogeneous earliest finish time scheduling algorithm for scientific workflows in cloud. Software—Practice and Experience, 52(3):689– 709, March 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kobylinski:2018:HRB

[KSBW18] Pawel Kobyliński, Konrad Siek, Jan Baranowski, and Paweł T. Wojciechowski. Helenos: a realistic benchmark for distributed transactional memory. Software—Practice and Experience, 48 (3):528–549, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2548.

Kumar:2024:QAR

[KSDG24] Mohit Kumar, Jitendra Kumar Samriya, Kalka Dubey, and Sukhpal Singh Gill. QoS-aware resource scheduling using whale optimization algorithm for microservice applications. Efficient Management of Microservice-based Systems and Applications: Software: Practice and Experience, 54(4):546–565, April 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2011:ZAS

[KSH11] SeongKi Kim, HaYoon Song, and SangYong Han. ZipfAllocation: an algorithm for static allocation of movies in a cluster of video servers. Software— Practice and Experience, 41(6):695–716, May 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2015:HSP

[KSH^{+}15] Jung Woo Kim, Jungjoo Seo, Jin Hong, Kunsoo Park, and Sung-Ryul Kim. High-speed parallel implementations of the rainbow method based on perfect tables in a heterogeneous system. Software—Practice and Experience, 45(6):837– 855, June 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kanade:2009:VGO

[KSK09] Aditya Kanade, Amitabha Sanyal, and Uday P. Khedker. Validation of GCC optimizers through trace generation. Software—Practice and Experience, 39(6): 611–639, April 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kumar:2015:EET

[KSK15] Manoj Kumar, Arun Sharma, and Rajesh Kumar. An empirical evaluation of a three-tier conduit framework for multifaceted test case classification and selection using fuzzy-ant colony optimisation approach. Software—Practice and Experience, 45(7):949– 971, July 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kumar:2012:QAC

[KSKG12] Vijai Kumar, Arun Sharma, Rajesh Kumar, and P. S. Grover. Quality aspects for component-based systems: a metrics based approach. Software—Practice and Experience, 42 (12):1531–1548, December 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:2017:ESI

[KSRR17] Dae-Kyoo Kim, Eunjee Song, Jungwoo Ryoo, and Y. Raghu Reddy. Editorial: Special issue on software reuse. Software—Practice and Experience, 47(7):941– 942, July 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kalaki:2023:ADO

[KSSA23] Parisa Sadat Kalaki, Alireza Shameli-Sendi, and Behzad Khalaji Emamzadeh Abbasi. Anomaly detection on OpenStack logs based on an improved robust principal component analysis model and its projection onto column space. Software—Practice and Experience, 53(3):665–681, March 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Khanahmadi:2023:DMB

[KSSJ^{+}23] Mohammad Khanahmadi, Alireza Shameli-Sendi, Masoume Jabbarifar, Quentin Fournier, and Michel Dagenais. Detection of microservice-based software anomalies based on OpenTracing in cloud. Software—Practice and Experience, 53(8):1681– 1699, August 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kumari:2022:PQC

[KSST22] Swati Kumari, Maninder Singh, Raman Singh, and Hitesh Tewari. Postquantum cryptography techniques for secure communication in resourceconstrained Internet of Things devices: a comprehensive survey. Software—Practice and Experience, 52(10):2047–2076, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kim:1994:FSM

[KST94] Jong Yong Kim and John Shawe-Taylor. Fast string matching using an n-gram algorithm. Software— Practice and Experience, 24(1):79–88, January 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kersten:1984:AOC

[KT84] Martin L. Kersten and Hans Tebra. Application of an optimistic concurrency control method. Software —Practice and Experience, 14(2):153–168, February 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Karnik:2001:SAM

[KT01a] Neeran M. Karnik and Anand R. Tripathi. Security in the Ajanta mobile agent system. Software—Practice and Experience, 31(4):301–329, April 10, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
76508663/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=76508663&PLACEBO=IE.
pdf.

Kniesel:2001:JAR

[KT01b] Günter Kniesel and Dirk Theisen. JAC — Access right based encapsulation for Java. Software—Practice and Experience, 31(6): 555–576, May 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
78003102/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=78003102&PLACEBO=IE.
pdf.

Khiat:2020:MMS

[KTG20] Abdelhamid Khiat, Abdelkamel Tari, and Tom Guérout. MFHS: a modular scheduling framework for heterogeneous system. Software—Practice and Experience, 50(8): 1463–1497, August 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Khwaja:1997:VSD

[KU97] Amir A. Khwaja and Joseph E. Urban. A visual syntax-directed editor for the Descartes Specification Language. Software—Practice and Experience, 27(6):651–676, June 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=7310;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=7310&
                                        PLACEBO=IE.pdf.

Kuenning:1995:KPI

[Kue95] Geoffrey H. Kuenning. Kitrace: Precise interactive measurement of operating system kernels. Software —Practice and Experience, 25(1):1–21, January 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kuhl:1990:OOP

[Kuh90] Frederick S. Kuhl. Objectoriented programming applied to a prototype workstation. Software—Practice and Experience, 20(9): 887–898, September 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kulsrud:1974:SSR

[Kul74]  H. E. Kulsrud.  Some

statistics on the reasons for compiler use. Software—Practice and Experience, 4(3):241–249, July/ September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kurokawa:1978:IOF

[Kur78] Toshiaki Kurokawa. Input/ output facilities in LISP 1.9. Software—Practice and Experience, 8(3):277– 284, May/June 1978. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kurokawa:1981:NFS

[Kur81] Toshiaki Kurokawa. A new fast and safe marking algorithm. Software—Practice and Experience, 11(7):671– 682, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kurtz:1999:RSR

[Kur99]  Stefan Kurtz.  Reduc-

ing the space requirement of suffix trees. Software—Practice and Experience, 29(13):1149–1171, November 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
66501674/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?

ID=66501674&PLACEBO=IE.
pdf.

Kernighan:1998:TTT

[KV98] Brian W. Kernighan and Christopher J. Van Wyk. Timing trials, or the trials of timing: Experiments with scripting and userinterface languages. Software—Practice and Experience, 28(8):819–843, July 10, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1811;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1811&
PLACEBO=IE.pdf.

Kabanov:2014:TYP

[KV14] Jevgeni Kabanov and Varmo Vene. A thousand years of productivity: the JRebel story. Software—Practice and Experience, 44(1):105–127, January 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kulkarni:2017:POR

[KV17] Naveen Kulkarni and Vasudeva Varma. Perils of opportunistically reusing software module. Software —Practice and Experience, 47(7):971–984, July 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Koopman:1995:OMS

[KvEP95] Pieter W. M. Koopman, M. C. J. D. van Eekelen, and M. J. Plasmeijer. Operational machine specification in a functional programming language. Software—Practice and Experience, 25(5):463– 499, May 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Karna:2019:ADM

[KVG19] Hrvoje Karna, Linda Vicković, and Sven Gotovac. Application of data mining methods for effort estimation of software projects. Software—Practice and Experience, 49(2): 171–191, February 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kempton:1990:RTD

[KW90] Willett Kempton and Brian A. Wichmann. Runtime detection of undefined variables considered essential. Software—Practice and Experience, 20(4):391– 402, April 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kiong:1992:ISE

[KW92] Derek Kiong and Jim Welsh. Incremental semantic evaluation in languagebased editors. Software— Practice and Experience, 22(2):111–135, February 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kastens:2009:RSM

[KW09] Uwe Kastens and William M. Waite. Reusable specification modules for type analysis. Software—Practice and Experience, 39(9):833– 864, June 25, 2009. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kastens:2017:NAM

[KW17] Uwe Kastens and William Waite. Name analysis for modern languages: a general solution. Software—Practice and Experience, 47(11):1597–1631, November 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kreahling:2005:BEC

[KWB^{+}05] William C. Kreahling, David Whalley, Mark W. Bailey, Xin Yuan, Gang-Ryung Uh, and Robert van Engelen. Branch elimination by condition merging. Software—Practice and Experience, 35(1):51– 74, January 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kersten:1981:APD

[KWW81] Martin L. Kersten, Anthony I. Wasserman, and Anthony I. Wasserman. The architecture of the PLAIN data base handler. Software—Practice and Experience, 11(2):175– 186, February 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Knobe:1977:CMC

[KY77] Bruce Knobe and Gideon Yuval. The cost of making a compiler indent. Software—Practice and Experience, 7(1):145, January 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kao:2005:AAM

[KY05] Tzu-Han Kao and Shyan-Ming Yuan. Automatic adaptation of mobile applications to different user devices using modular mobile agents. Software— Practice and Experience, 35(14):1349–1391, November 25, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Khan:2022:VRA

[KYF^{+}22] Javed Ali Khan, Affan Yasin, Rubia Fatima, Danish Vasan, Arif Ali Khan, and Abdul Wahid Khan. Valuating requirements arguments in the online user’s forum for requirements decisionmaking: The CrowdRE-VArg framework. Software—Practice and Experience, 52(12):2537–2573, December 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Legge:1990:UFS

[LA90] Gaynor Legge and Mahir Ali. UNIX file system behaviour and machine architecture dependency. Software—Practice and Experience, 20(11):1077–1096, November 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lanna:2011:SD

[LA11] Marconi Lanna and Daniel Amyot. Spotting the difference. Software—Practice and Experience, 41(6):607– 626, May 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Linos:1994:VPD

[LAD^{+}94] Panagiotis K. Linos, Philippe Aubet, Laurent Dumas, Yann Helleboid, Patricia Lejeune, and Philippe Tulula. Visualizing program dependencies: An experimental study. Software —Practice and Experience, 24(4):387–403, April 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lieuwen:2000:LTG

[LAG00] Daniel Lieuwen, Robert Arlein, and Narain Gehani. The LTAP trigger gateway for LDAP directories. Software—Practice and Experience, 30(9):1003–1024, July 25, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72504004/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72504004&PLACEBO=IE.
pdf.

Lai:1995:UPV

[Lai95] Richard Lai. Using PRO-TEAN for verifying a complex protocol — a case study. Software—Practice and Experience, 25 (9):1021–1044, September 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lakos:1980:IBB

[Lak80] C. A. Lakos. Implementing BCPL on the Burroughs B6700. Software—Practice and Experience, 10(8):673– 683, August 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lampson:1972:EGE

[Lam72] Butler W. Lampson. Editorial: Guest editorial. Software—Practice and Experience, 2(3):195–196, July 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lamb:1981:CPO

[Lam81] David Alex Lamb. Construction of a peephole optimizer. Software—Practice and Experience, 11(6): 639–647, June 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lamperti:2020:TDM

[Lam20] Gianfranco Lamperti. Temporal determinization of mutating finite automata: Reconstructing or restructuring. Software—Practice and Experience, 50(4):335– 367, April 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Landy:1971:DSS

[Lan71] B. Landy. Development of scheduling strategies in the TITAN operating system. Software—Practice and Experience, 1(3):279– 295, July/September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lang:1974:BRB

[Lan74a] C. A. Lang. Book review: Graphic languages, proceedings of the IFIP working conference on graphic languages, Ed. F. Nake and A. Rosenfeld, North-Holland, Amsterdam, 1972. No. of pages:

442. Price: $19.30. Software—Practice and Experience, 4(2):192, April/June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lang:1974:EPL

[Lan74b] C. A. Lang. Editorial:

Programming languages for writing system programs. Software—Practice and Experience, 4(1): 1–4, January/March 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lang:1975:BRB

[Lan75] T. Lang. Book review:

Macro Processors and Techniques for Portable Software, P. J. Brown, Wiley, New York, 1974. No. of pages: 244. Price: £5.25. Software—Practice and Experience, 5(1):110– 111, January 1975. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lang:1976:ECT

[Lan76]  C. A. Lang.  Edito-

rial: Computer typesetting. Software—Practice and Experience, 6(3): 291, July 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lang:1982:APD

[Lan82] T. Lang. Another program for drawing diagrams. Software—Practice and Experience, 12(5):397–404, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Langston:1990:UMT

[Lan90] Peter S. Langston. Unix music tools at Bellcore. Software—Practice and Experience, 20(S1):S47– S61, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Larmouth:1971:BRB

[Lar71] J. Larmouth. Book review: OS/360 Job control language, Harry W. Cadow, Prentice–Hall, New Jersey, 1970. No. of pages: 301. Price: £6. Software—Practice and Experience, 1(3):

301, July/September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Larmouth:1973:SF

[Lar73a] J. Larmouth. Serious FOR-TRAN. Software—Practice and Experience, 3(2): 87–107, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Larmouth:1973:SFP

[Lar73b] J. Larmouth. Serious FOR-TRAN — part 2. Software—Practice and Experience, 3(3):197–225, July/ September 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Larmouth:1975:BRB

[Lar75a] J. Larmouth. Book review: The Macro Implementation of Snobol 4, Ralph E. Griswold, W. H. Freeman & Co., 1972. No. of pages: 310. Price £7.00. Software—Practice and Experience, 5(1):111– 112, January 1975. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Larmouth:1975:SSM

[Lar75b] J. Larmouth. Scheduling for a share of the machine. Software—Practice and Experience, 5(1):29– 49, January/March 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Larmouth:1978:SIT

[Lar78] J. Larmouth. Scheduling for immediate turnround. Software—Practice and Experience, 8(5):559– 578, September/October 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Larmouth:1981:FP

[Lar81] J. Larmouth. Fortran 77 portability. Software— Practice and Experience, 11(10):1071–1117, October 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Larus:1990:AET

[Lar90]  James R. Larus.  Ab-

stract execution: a technique for efficiently tracing programs. Software—Practice and Experience, 20 (12):1241–1258, December 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Laramee:2008:CEC

[Lar08] Robert S. Laramee. Comparing and evaluating computer graphics and visualization software. Software —Practice and Experience, 38(7):735–760, June ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Laracy:2009:RVG

[Lar09] Joseph R. Laracy. Random variate generator. Software—Practice and Experience, 39(1):105–110, January ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lauesen:1979:DT

[Lau79] Søren Lauesen. Debugging techniques. Software —Practice and Experience, 9(1):51–63, January 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Laver:1977:BRB

[Lav77] Murray Laver. Book review: Data processing in 1980–1985, A Study Of Potential Limitations To Progress, T. A. Dolotta, M. I. Bernstein, R. S. Dickson Jr., N. A. France, B. A. Rosenblatt, D. M. Smith and T. B. Steel Jr., Wiley, New York, 1976. No. of pages: 191. Price: £9.80/$16.60 (cloth), £4.30 / $7.95 (paper). Software —Practice and Experience, 7(3):428–429, June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Laver:1978:BRB

[Lav78] Murray Laver. Book review: Using computers, Brian L. Meek and Simon Fairthorne, Ellis Horwood, Chichester, 1977. No. of pages: 208. Price: £7.50 (cloth), £2-75 (paper). Software—Practice and Experience, 8(4):496– 497, July 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lawrence:1978:SSC

[Law78] A. R. Lawrence. SCRUBsystematically clean and renumber users BASIC. Software—Practice and Experience, 8(2):227–232, March/April 1978. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lunn:1981:MTC

[LB81] K. Lunn and K. H. Bennett. Message transport on the Cambridge Ring — a simulation study. Software —Practice and Experience, 11(7):711–716, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Larus:1994:REF

[LB94] James R. Larus and Thomas Ball. Rewriting executable files to measure program behavior. Software—Practice and Experience, 24(2):197–218, February 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:2002:ERM

[LB02] Joon-Sang Lee and Doo-Hwan Bae. An enhanced role model for alleviating the role-binding anomaly. Software—Practice and Experience, 32 (14):1317–1344, November 25, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
98517569/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=98517569&PLACEBO=IE.
pdf.

Lemire:2015:DBI

[LB15] D. Lemire and L. Boytsov.

Decoding billions of integers per second through vectorization. Software— Practice and Experience, 45(1):1–29, January 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:2011:ASC

[LBC^{+}11] Sukhee Lee, Gigon Bae, Heung Seok Chae, Doo-Hwan Bae, and Yong Rae Kwon. Automated scheduling for clone-based refactoring using a competent GA. Software—Practice and Experience, 41(5):521– 550, May 25, 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Laso:2021:HMF

[LBGA^{+}21] Sergio Laso, Javier Berrocal, José García-Alonso, Carlos Canal, and Juan Manuel Murillo. Human microservices: a framework for turning humans into service providers. Software—Practice and Experience, 51(9):1910–1935, September 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lemire:2016:SCI

[LBK16] Daniel Lemire, Leonid Boytsov, and Nathan Kurz. SIMD compression and the intersection of sorted integers. Software—Practice and Experience, 46(6):723– 749, June 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lawall:2013:WEF

[LBP^{+}13] Julia L. Lawall, Julien Brunel, Nicolas Palix, René Rydhof Hansen, Henrik Stuart, and Gilles Muller. WYSIWIB: exploiting fine-grained program structure in a scriptable API-usage protocolfinding process. Software —Practice and Experience, 43(1):67–92, January 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lawson:1978:FDH

[LBS78] Harold W. Lawson, Jr., Miquel Berthan, and Javier Sanagustin. The formal definition of human/machine communications. Software—Practice and Experience, 8(1): 51–58, January/February 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lapalme:1986:LIP

[LC86] Guy Lapalme and Suzanne Chapleau. Logicon: An integration of Prolog into Icon. Software—Practice and Experience, 16(10): 925–944, October 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lhee:2003:BOF

[LC03] Kyung-Suk Lhee and Steve J.

Chapin. Buffer overflow and format string overflow vulnerabilities. Software —Practice and Experience, 33(5):423–460, April 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:2005:DDR

[LC05] Yueh-Feng Lee and Ruei-Chuan Chang. Developing dynamic-reconfigurable communication protocol stacks using Java. Software —Practice and Experience, 35(6):601–620, May 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liao:2007:TPS

[LC07] Hsien-Chou Liao and Wen-Feng Chen. Two-phase scale-based reduction method for fulfilling monitoring service on mobile devices. Software—Practice and Experience, 37(1):1– 19, January 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lin:2012:ATE

[LC12] Kai-Wei Lin and Peng-Sheng Chen. An assistance tool employing a systematic methodology for GCC retargeting. Software —Practice and Experience, 42(1):19–36, January 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lhotak:2009:UXZ

[LCA09] Ondřej Lhoták, Stephen Curial, and José Nelson Amaral. Using XBDDs and ZBDDs in points-to analysis. Software—Practice and Experience, 39(2):163– 188, February ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:1997:HSA

[LCC97] Paul C. H. Lee, Ruei-Chuan Chang, and Meng Chang Chen. Hipec: a system for application-customized virtual-memory caching management. Software— Practice and Experience, 27(5):547–571, May 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www3.interscience.
wiley.com/cgi-bin/abstract?
ID=7307;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7307&
PLACEBO=IE.pdf.

Liao:2014:NMM

[LCC14] Chih-Shan Liao, Chen-Po Chuang, and Ruay-Shiung Chang. A novel monitoring mechanism by event trigger for Hadoop system performance analysis. Software —Practice and Experience, 44(7):823–834, July 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2017:LCR

[LCGS17] Leyuan Liu, Jingying Chen, Changxin Gao, and Nong Sang. A lowcost real-time face tracking system for ITSs and SDASs. Software—Practice and Experience, 47(8): 1111–1126, August 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lai:2021:DFA

[LCT^{+}21] Ying Hsun Lai, Yao Chung Chang, Chia Wei Tsai, Chih Hsun Lin, and Mu Yen Chen. Data fusion analysis for attentiondeficit hyperactivity disorder emotion recognition with thermal image and Internet of Things devices. Software—Practice and Experience, 51(3):595– 606, March 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liao:1998:DIS

[LCW98] Hsien-Chou Liao, Ming-Feng Chen, and Feng-Jian Wang. A domainindependent software reuse framework based on a hierarchical thesaurus. Software—Practice and Experience, 28(8):799–818, July 10, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1810;   http://www3.
interscience.wiley.com/

cgi-bin/fulltext?ID=1810&
PLACEBO=IE.pdf.

Liu:2007:ESC

[LCW07] Ho-Han Liu, Mei-Ling Chiang, and Men-Chao Wu. Efficient support for content-aware request distribution and persistent connection in Web clusters. Software—Practice and Experience, 37 (11):1215–1241, September 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Luo:2007:TPI

[LCY07] Gang Luo, Tong Chen, and Hao Yu. Toward a progress indicator for program compilation. Software—Practice and Experience, 37(9): 909–933, July 25, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lim:2008:DIP

[LCZ08] Sang Seok Lim, Jong Hyuk Choi, and Kurt D. Zeilenga. Design, implementation, and performance analysis of PKI certificate repository using LDAP Component Matching. Software —Practice and Experience, 38(8):827–851, July 10, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lirov:1987:IDD

[LD87] Yuval Lirov and Nissim Daunov. An integrated data dictionary to facilitate automatic report generation in a network database. Software—Practice and Experience, 17(3): 187–195, March 1987. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lin:1995:FGU

[LD95] Yi-Bing B. Lin and Dan Daly. A flexible graphical user interface for performance modeling. Software —Practice and Experience, 25(2):193–216, February 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:1999:OPB

[LD99] Wenyin Liu and Dov Dori. Object-process based graphics recognition class library: principles and applications. Software— Practice and Experience, 29(15):1355–1378, December 25, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
68501287/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=68501287&PLACEBO=IE.
pdf.

Laval:2014:RCD

[LD14] Jannik Laval and Stéphane Ducasse. Resolving cyclic dependencies between packages with enriched dependency structural matrix. Software—Practice and Experience, 44(2):235– 257, February 2014. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Levy:1996:ILP

[LDG^{+}96] C. H. Levy, L. H. De Figueiredo, M. Gatass, C. J. P. Lucena, and D. D. Cowan. IUP/LED: a portable user interface development tool. Software—Practice and Experience, 26(7):737–762, July 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16818.

Leavitt:1992:LE

[LDH92] Thomas C. Leavitt and Jorge L. Diaz-Herrera. Letter to the editor. Software —Practice and Experience, 22(10):909–910, October 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:1998:IAS

[LDI98] Z. Liu, X. Du, and N. Ishii.

An improved adaptive string searching algorithm. Software—Practice and Experience, 28(2):191–198, February 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1775;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1775&
PLACEBO=IE.pdf.

LeRiche:1988:KPM

[Le 88] P. J. Le Riche. K6 — a protected mode kernel for the iAPX286. Software—Practice and Experience, 18(12):1109–1124, December 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Leavenworth:1977:SDU

[Lea77] B. M. Leavenworth. Structured debugging using a domain specific language. Software—Practice and Experience, 7(4):475– 482, July/August 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Leavitt:1981:LE

[Lea81] Randal Leavitt. Letter to the editor. Software —Practice and Experience, 11(3):321, March 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Leathrum:1982:DMS

[Lea82] J. F. Leathrum. A design medium for software. Software—Practice and Experience, 12(6):497–503, June 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lecroq:1995:ERS

[Lec95] Thierry Lecroq. Experimental results on string matching algorithms. Software—Practice and Experience, 25(7):727–765, July 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lecroq:1998:ESM

[Lec98] Thierry Lecroq. Experiments on string matching in memory structures. Software—Practice and Experience, 28(5):561–568, April 25, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1796;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1796&
PLACEBO=IE.pdf.

Lee:1980:BM

[Lee80]  Martin P. Lee.  Bib-

liographic manipulation. Software—Practice and Experience, 10(5):347–354, May 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:1983:EHC

[Lee83] P. A. Lee. Exception handling in C programs. Software—Practice and Experience, 13(5):389–405, May 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Leith:1984:TDW

[Lei84] Philip Leith. Top-down design within a functional environment. Software— Practice and Experience, 14(10):921–930, October 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Leith:1985:II

[Lei85] Philip Leith. An IKBS implementation. Software —Practice and Experience, 15(1):65–86, January 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lemire:2021:NPG

[Lem21] Daniel Lemire. Number parsing at a gigabyte per second. Software—Practice and Experience, 51(8): 1700–1727, August 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See improvement [ML23].

Lengauer:1990:CGS

[Len90] Christian Lengauer. Code generation for a systolic computer. Software—Practice and Experience, 20(3): 261–282, March 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Leroy:2002:BVJ

[Ler02] Xavier Leroy. Bytecode verification on Java smart cards. Software—Practice and Experience, 32(4):319– 340, April 10, 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
91016433/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=91016433&PLACEBO=IE.
pdf.

Letrache:2017:ACO

[LER17] Khadija Letrache, Omar El Beggar, and Mohammed Ramdani. The automatic creation of OLAP cube using an MDA approach. Software—Practice and Experience, 47 (12):1887–1903, December 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lesk:1972:GID

[Les72] Arthur M. Lesk. Generation of interactive displays from FORTRAN using the PDP-10/LSD-1 computer graphics system. Software—Practice and Experience, 2(3):259–273, July/ September 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Luders:1995:AAD

[LES95] Peter Lüders, Rolf Ernst, and Stefan Stille. An approach to automatic display layout using combinatorial optimization algorithms. Software—Practice and Experience, 25 (11):1183–1202, November 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Levison:1980:SHL

[Lev80] Michael Levison. Simulating high-level control structures in assembly language. Software—Practice and Experience, 10(11): 949–950, November 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Levine:1982:OYG

[Lev82a] John Levine. An overview of the Yale gem system. Software—Practice and Experience, 12 (12):1133–1145, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Levison:1982:PTE

[Lev82b] Michael Levison. A programmable text-editing system. Software—Practice and Experience, 12(7):611– 621, July 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Levison:1983:EMF

[Lev83] Michael Levison. Editing mathematical formulae. Software—Practice and Experience, 13(2):189– 195, February 1983. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Levy:1995:IOF

[Lev95] G. F. Levy. Improving the output of the FOR-TRAN to C translator, f2c. Software—Practice and Experience, 25(2):217– 227, February 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Levy:1997:USH

[Lev97]  G. F. Levy.  Using

<stdio.h> in the output of the FORTRAN to C translator, f2c. Software—Practice and Experience, 27(12):1369–1384, December 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7276;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7276&
PLACEBO=IE.pdf.

Levy:1998:WPG

[Lev98] Michael R. Levy. Web programming in Guide. Software—Practice and Experience, 28(15):1581–1603, December 25, 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=10050311;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=10050311&PLACEBO=IE.
pdf.

Levy:2001:NAC

[Lev01] G. F. Levy. Numeric ActiveX components. Software—Practice and Experience, 31(2):147–189, February 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
76504936/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?

ID=76504936&PLACEBO=IE.
pdf.

Lew:1983:DTG

[Lew83] Art Lew. Decision tables for general-purpose scientific programming. Software—Practice and Experience, 13(2):181–188, February 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lang:1974:GPT

[LF74] T. Lang and D. A. Fox.

A general purpose task queueing mechanism for small machines. Software—Practice and Experience, 4(1):63–70, January/ March 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Leblanc:1982:CSR

[LF82] Richard J. Leblanc and Charles N. Fischer. A case study of run-time errors in Pascal programs. Software —Practice and Experience, 12(9):825–834, September 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lyttle:1990:SDR

[LF90] Dan Lyttle and Ray Ford.

A symbolic debugger for real-time embedded Ada software. Software—Practice and Experience, 20(5):

499–514, May 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lago-Fernandez:2014:NAA

[LFGCGCRP14] J. Lago-Fernández, F. Gil-Castiñeira, F. J. González-Castaño, and A. Román-Portabales. A new approach to authenticating and encrypting Voice over Internet Protocol communications. Software—Practice and Experience, 44(5): 593–619, May 2014. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liang:2022:NAM

[LFH^{+}22] Jianguo Liang, You Fu, Rong Hua, Hao Zhang, Wenqiang Zhu, and Yuxi Ye. A novel acceleration method for molecular dynamics of crystal silicon on GPUs using OpenACC. Software—Practice and Experience, 52(10): 2226–2240, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Li:2022:SRS

[LFHL22] Zongwei Li, Dan Feng, Yuchong Hu, and Mengting Lu. SPOPB: Reducing solid state drive write traffic for flash-based key– value caching. Software —Practice and Experience, 52(2):353–373, February 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:1996:DIC

[LFW96] Tsing-Gen Lee, Wen-Jong Fang, and Allen C.-H. Wu. The design and implementation of a cooperative design-view environment for interactive partitioning applications. Software—Practice and Experience, 26(10):1141–1160, October 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16766.

Leach:1973:BXX

[LG73] Geoffrey Leach and Hellmut Golde. Bootstrapping XPL to an XDS Sigma 5 computer. Software—Practice and Experience, 3(3):235–244, July/ September 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Loeser:1976:SLD

[LG76] R. Loeser and E. M.

Gaposchkin. The second law of debugging. Software—Practice and Experience, 6(4):577–578, October/December 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lieuwen:1999:VOI

[LG99] Daniel Lieuwen and Narain Gehani. Versions in Ode: implementation and experiences. Software—Practice and Experience, 29(5):397– 416, April 25, 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=55003858;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=55003858&PLACEBO=IE.
                                        pdf.

Li:2019:IAH

[LG19] Changqing Li and Jianhua Gu. An integration approach of hybrid databases based on SQL in cloud computing environment. Software—Practice and Experience, 49(3):401– 422, March 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

LeBlanc:1984:SDP

[LGC84] Thomas J. LeBlanc, Robert H. Gerber, and Robert P. Cook. The StarMod distributed programming kernel. Software—Practice and Experience, 14 (12):1123–1139, December 1984. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Loyall:2011:DPD

[LGP^{+}11] Joseph P. Loyall, Matthew Gillen, Aaron Paulos, Larry Bunch, Marco Carvalho, James Edmondson, Douglas C. Schmidt, Andrew Martignoni III, and Asher Sinclair. Dynamic policy-driven quality of service in serviceoriented information management systems. Software —Practice and Experience, 41(12):1459–1489, November/December 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lobato:2008:AOS

[LGRL08] Cidiane Lobato, Alessandro Garcia, Alexander Romanovsky, and Carlos Lucena. An aspect-oriented software architecture for code mobility. Software —Practice and Experience, 38(13):1365–1392, November 10, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lin:2008:MRT

[LGZ^{+}08] Yuehua Lin, Jeff Gray, Jing Zhang, Steve Nordstrom, Aniruddha Gokhale, Sandeep Neema, and Swapna Gokhale. Model replication: transformations to address model scalability. Software—Practice and Experience, 38(14):1475– 1497, November 25, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Leverett:1982:ASD

[LH82] Bruce W. Leverett and Peter G. Hibbard. An adaptive system for dynamic storage allocation. Software—Practice and Experience, 12(6):543–555, June 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Li:1986:NLC

[LH86] Kai Li and Paul Hudak. A new list compaction method. Software—Practice and Experience, 16(2): 145–163, February 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lehrig:2018:ATM

[LHB18] Sebastian Lehrig, Marcus Hilbrich, and Steffen Becker. The architectural template method: templating architectural knowledge to efficiently conduct quality-of-service analyses. Software—Practice and Experience, 48(2): 268–299, February 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2517.

Lo:1997:FRT

[LHC97] Siu Ling Ann Lo, Norman C. Hutchinson, and Samuel T. Chanson. A flexible real-time scheduling abstraction: Design and implementation. Software—Practice and Experience, 27(9):1055–1066, September 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7333;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7333&
PLACEBO=IE.pdf.

Li:2015:BGG

[LHC15] Chen Li, Linpeng Huang, and Luxi Chen. Breeze graph grammar: a graph grammar approach for modeling the software architecture of big dataoriented software systems. Software—Practice and Experience, 45(8):1023– 1050, August 2015. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lin:2007:SIM

[LHFL07] Jim-Min Lin, Zeng-Wei Hong, Guo-Ming Fang, and Che-Tai Lee. A style for integrating MS-Windows software applications to client-server systems using Java technology. Software —Practice and Experience, 37(4):417–440, April 10, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:2015:DME

[LHGM15] Byeongcheol Lee, Martin Hirzel, Robert Grimm, and Kathryn S. McKinley. Debugging mixedenvironment programs with Blink. Software—Practice and Experience, 45 (9):1277–1306, September 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Luecke:1991:EFV

[LHH^{+}91] Glenn Luecke, Waqar Haque, James Hoekstra, Howard Jespersen, and James Coyle. Evaluation of Fortran vector compilers and preprocessors. Software—Practice and Experience, 21(9):891–905, September 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Loh:1999:VPC

[LHK99] Peter K. K. Loh, Gurdeep S. Hura, and Chia Cheng Khoon. Virtual prototyping of cellular phones. Software—Practice and Experience, 29(10):897–929, August 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=63001369;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=63001369&PLACEBO=IE.
pdf.

Lim:1995:MDI

[LHS^{+}95] Ee-Peng Lim, San-Yih Hwang, Jaideep Srivastava, Dave Clements, and M. Ganesh. Myriad: Design and implementation of a federated database prototype. Software—Practice and Experience, 25(5):533– 562, May 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Li:2018:OLD

[Li18] Keqin Li. Optimal load distribution for multiple classes of applications on heterogeneous servers with variable speeds. Software—Practice and Experience, 48(10):1805–1819, October 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Li:2022:WMS

[Li22]  Keqin Li.  Workload

management and server speed setting for costperformance ratio optimization. Software—Practice and Experience, 52 (12):2635–2663, December 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Libes:1993:KCM

[Lib93] Don Libes. Kibitz — connecting multiple interactive programs together. Software—Practice and Experience, 23(5):465–475, May 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Libes:1997:ATC

[Lib97a] Don Libes. Automation and testing of charactergraphic programs. Software—Practice and Experience, 27(2):123–137, February 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=7282;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=7282&
                                        PLACEBO=IE.pdf.

Libes:1997:TTB

[Lib97b]  Don Libes.  Tcl/Tk-

based agents for mail and news notification. Software—Practice and Experience, 27(4):481–493, April 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7302;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7302&
PLACEBO=IE.pdf.

Lichtman:1977:ICU

[Lic77] Zavdi L. Lichtman. Interpretation control using a linear tree representation. Software—Practice and Experience, 7(2):275– 278, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lichtman:1986:FNL

[Lic86] Zavdi L. Lichtman. The function of T and NIL in LISP. Software—Practice and Experience, 16(1):1–3, January 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lindgaard:1979:PTO

[Lin79] Anders Lindgård. P: a timesharing operating system for laboratory automation. Software—Practice and Experience, 9(12):971– 986, December 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Linnainmaa:1986:ITS

[Lin86] Seppo Linnainmaa. Icecream, transportable software for creating fiendly human interfaces. Software—Practice and Experience, 16(8):739–750, August 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lins:1987:ECC

[Lin87] Rafael D. Lins. On the efficiency of categorical combinators as a rewriting system. Software—Practice and Experience, 17(8):547– 559, August 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Linderman:1998:LEA

[Lin98a] Jp Linderman. Letter to the Editor. Algorithms for the Hilbert order. Software—Practice and Experience, 28(1):121–123, January 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=61001669;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=61001669&PLACEBO=IE.
                                        pdf.

Lindvall:1998:LCC

[Lin98b]  Mikael Lindvall.  Are

large C++ classes changeprone? An empirical investigation. Software— Practice and Experience, 28(15):1551–1558, December 25, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=10050309;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=10050309&PLACEBO=IE.
pdf.

Lions:1979:EUT

[Lio79]  John Lions.  Experi-

ences with the UNIX timesharing system. Software —Practice and Experience, 9(9):701–709, September 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Litman:1993:IPH

[Lit93] Andy Litman. An implementation of precompiled headers. Software—Practice and Experience, 23(3): 341–350, March 1993. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:1986:SPM

[Liu86] Ken-Chih Liu. A string pattern matching extension to Pascal and some comparisons with SNOBOL4. Software—Practice and Experience, 16(6):541–548, June 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2001:RSP

[Liu01] Mengchi Liu. The Relationlog system prototype. Software—Practice and Experience, 31(5):409– 443, April 25, 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
77502755/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=77502755&PLACEBO=IE.
pdf.

Liu:2003:DUD

[Liu03] Mengchi Liu. DrawCAD: using deductive objectrelational databases in CAD. Software—Practice and Experience, 33(2):143– 172, February 2003. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
102525917/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=102525917&PLACEBO=
IE.pdf.

Livesley:1975:BRB

[Liv75] R. K. Livesley. Book review: Structural Mechanics Computer Programs, Surveys, Assessments, and Availability, Edited by W. Pilkey, K. Saczalski and H. Schaeffer, University Press of Virginia, Charlottesville, 1975. No. of pages: 1100. Price: £20. Software— Practice and Experience, 5(3):315–316, July 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Luck:1999:SLS

[LJ99] Michael Luck and Mike Joy. A secure on-line submission system. Software—Practice and Experience, 29(8):721–740, July 10, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=62004391;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=62004391&PLACEBO=IE.
pdf.

Lee:2010:TPT

[LJL^{+}10] Hyunjin Lee, Lei Jin, Kiyeon Lee, Socrates Demetriades, Michael Moeng, and Sangyeun Cho. Twophase trace-driven simulation (TPTS): a fast multicore processor architecture simulation approach. Software—Practice and Experience, 40(3):239–258, March ??, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2020:DSV

[LJS20] Richen Liu, Genlin Ji, and Mingjun Su. Domainspecific visualization system based on automatic multiseed recommendations: Extracting stratigraphic structures. Software—Practice and Experience, 50(2):98–115, February 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lucchesi:1993:AFA

[LK93] Cláudio L. Lucchesi and Tomasz Kowaltowski. Applications of finite automata representing large vocabularies. Software— Practice and Experience, 23(1):15–30, January 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:1999:OVO

[LK99] Sang-Won Lee and Hyoung-Joo Kim. Object versioning in an ODMGcompliant object database system. Software—Practice and Experience, 29(5):479– 500, April 25, 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55003855;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55003855&PLACEBO=IE.
pdf.

Levelt:1992:CTP

[LKBT92] Willem G. Levelt, M. Frans Kaashoek, Henri E. Bal, and Andrew S. Tanenbaum. A comparison of two paradigms for distributed shared memory. Software—Practice and Experience, 22(11):985–1010, November 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2012:CAR

[LKC12] Huai Liu, Fei-Ching Kuo, and Tsong Yueh Chen. Comparison of adaptive random testing and random testing under various testing and debugging scenarios. Software—Practice and Experience, 42(8): 1055–1074, August 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:2000:FBA

[LKCC00] Kwanwoo Lee, Kyo C.

Kang, Wonsuk Chae, and Byoung Wook Choi.

Feature-based approach to object-oriented engineering of applications for reuse. Software—Practice and Experience, 30(9):1025– 1046, July 25, 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72504243/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72504243&PLACEBO=IE.
pdf.

Lee:2013:DVO

[LKCW13] Yun-Jung Lee, Eun-Kyung Kim, Hwan-Gue Cho, and Gyun Woo. Detecting and visualizing online dispute dynamics in replying comments. Software—Practice and Experience, 43 (12):1395–1413, December 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lemire:2018:RBI

[LKK^{+}18] Daniel Lemire, Owen Kaser, Nathan Kurz, Luca Deri, Chris O’Hara, François Saint-Jacques, and Gregory Ssi-Yan-Kai. Roaring bitmaps: Implementation of an optimized software library. Software— Practice and Experience, 48(4):867–895, April 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2560.

Lemire:2019:FRD

[LKK19] Daniel Lemire, Owen Kaser, and Nathan Kurz. Faster remainder by direct computation: Applications to compilers and software libraries. Software—Practice and Experience, 49(6): 953–970, June 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lyon:1995:STP

[LKL95] Gordon Lyon, Raghu Kacker, and Arnaud Linz. A scalability test for parallel code. Software—Practice and Experience, 25 (12):1299–1314, December 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Levary:1991:MSD

[LL91] Reuven R. Levary and Chi Y. Lin. Modelling the software development process using an expert simulation system having fuzzy logic. Software—Practice and Experience, 21(2):133– 148, February 1991. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lai:1996:ESE

[LL96] Richard Lai and A. Lo. EASE: a software environment for automatic implementation of application protocol. Software—Practice and Experience, 26(1): 83–103, January 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61004285.

LeBlanc:1989:EMO

[LLCG^{+}89] Thomas J. LeBlanc, John M. Lellor-Crummey, Neal M. Gafter, Lawrence A. Crowl, and Peter C. Dibble. The elmwood multiprocessor operating system. Software—Practice and Experience, 19(11):1029–1055, November 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liang:2014:RBA

[LLH14] Qianhui Liang, Bu-Sung Lee, and Patrick C. K. Hung. A rule-based approach for availability of service by automated service substitution. Software —Practice and Experience, 44(1):47–76, January 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lafi:2012:AHR

[LLJ12] Walid Lafi, Didier Lattard, and Ahmed Jerrya. An asynchronous hierarchical router for networks-on-chip-based three-dimensional multi-processor system-on-chip. Software—Practice and Experience, 42(7):877– 890, July 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2023:BBP

[LLJ^{+}23] Yang Liu, Liangjie Lin, Lin Jiang, Weizhe Zhang, Xuan Wang, Mehdi Gheisari, Tao Gong, Chao Gao, and Hamid Esmaeili Najafabadi. A blockchainbased privacy-preserving advertising attribution architecture: Requirements, design, and a prototype implementation. Software— Practice and Experience, 53(8):1700–1721, August 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:2004:ITS

[LLK04] Min-Gu Lee, Sunggu Lee, and K. H. (Kane) Kim. Implementation of a TMOstructured real-time airplanelanding simulator on a distributed computing environment. Software—Practice and Experience, 34 (15):1441–1462, December 2004. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Luo:2018:PTV

[LLLY18] Haoyu Luo, Jin Liu, Xiao Liu, and Yun Yang. Predicting temporal violations for parallel business cloud workflows. Software— Practice and Experience, 48(4):775–795, April 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2554.

Liu:2019:ASM

[LLLY19] Yuzhou Liu, Lei Liu, Huaxiao Liu, and Xinglong Yin. App store mining for iterative domain analysis: Combine app descriptions with user reviews. Software—Practice and Experience, 49(6): 1013–1040, June 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Latorre:2005:SPD

[LLM05] Roberto Latorre, Francisco López, and Antonio E. Martínez. Sharing of precompiled database statements in J2EE applications. Software—Practice and Experience, 35(3):301– 311, March 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2016:EOC

[LLN16] Fei Liu, Bixin Li, and Rupesh Nasre. Efficient online cycle detection technique combining with Steensgaard points-to information. Software—Practice and Experience, 46(5):601– 623, May 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lloyd:1982:BRB

[Llo82] E. Keith Lloyd. Book review: The art of computer programming, vol. 2, Seminumerical algorithms (2nd edition), Donald E. Knuth, Addison-Wesley, Reading, Mass, 1981. No. of pages: xiv + 688. Price: £17.95. ISBN 0-201-03822-6. Software—Practice and Experience, 12(9):883–884, September 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2006:AWD

[LLS06] Yen-Tso Liu, Tyng-Yeu Liang, and Ce-Kuen Shieh. Adapting workload distribution on software DSM clusters. Software—Practice and Experience, 36(11– 12):1133–1155, September/ October 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Leung:1998:DGD

[LLW98] Kwong-Sak Leung, Kin-Hong Lee, and Yuk-Yin Wong. DJM: a global distributed virtual machine on the Internet. Software—Practice and Experience, 28(12):1269–1297, October 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=1763;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=1763&
                                        PLACEBO=IE.pdf.

Liu:2023:CFL

[LLW^{+}23] Hengyuan Liu, Zheng Li, Haifeng Wang, Yong Liu, and Xiang Chen. CRMF: a fault localization approach based on class reduction and method call frequency. Software—Practice and Experience, 53 (4):1061–1090, April 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lin:2014:BAD

[LLWB14] Weiwei Lin, Chen Liang, James Z. Wang, and Rajkumar Buyya. Bandwidthaware divisible task scheduling for cloud computing. Software—Practice and Experience, 44(2):163– 174, February 2014. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Li:2020:TEB

[LLZ20] Jingyao Li, Lei Liu, and Peng Zhang. Tabularexpression-based method for constructing metamorphic relations. Software —Practice and Experience, 50(8):1345–1380, August 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lister:1976:IM

[LM76] A. M. Lister and K. J.

Maynard. An implementation of monitors. Software—Practice and Experience, 6(3):377–385, July/ September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lemoine:1981:STP

[LM81a] M. Lemoine and J. Mullor. Software transferability: a practical approach. Software—Practice and Experience, 11(5):425– 433, May 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Logothetis:1981:CSC

[LM81b] George Logothetis and Prateek Mishra. Compiling short-circuit Boolean expressions in one pass. Software—Practice and Experience, 11(11):1197–1214, November 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lange:2002:EMA

[LM02] Danny B. Lange and Dejan S. Milojicic. Editorial: Mobile agent systems. Software—Practice and Experience, 32(6):495– 496, May 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
93513615/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=93513615&PLACEBO=IE.
pdf.

Liddell:2006:DPC

[LM06] Mike Liddell and Alistair Moffat. Decoding prefix codes. Software—Practice and Experience, 36 (15):1687–1710, December 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:2007:WUS

[LM07] Eric K. Lee and Charles U. Martel. When to use splay trees. Software— Practice and Experience, 37(15):1559–1575, December ??, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Leotta:2015:PML

[LM15] Francesco Leotta and Massimo Mecella. PLaTHEA: a marker-less people localization and tracking system for home automation. Software—Practice and Experience, 45(6):801–835, June 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lemire:2022:TBU

[LM22] Daniel Lemire and Wojciech Muła. Transcoding billions of Unicode characters per second with SIMD instructions. Software— Practice and Experience, 52(2):555–575, February 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:2016:ECP

[LMK16] Seong-Won Lee, Soo-Mook Moon, and Seong-Moo Kim. Extended conference papers: Flow-sensitive runtime estimation: an enhanced hot spot detection heuristics for embedded Java just-in-time compilers. Software—Practice and Experience, 46(6):841– 864, June 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lamma:1991:RMC

[LMN91] Evelina Lamma, Paola Mello, and Antonio Natali. Reflection mechanisms for combining Prolog databases. Software —Practice and Experience, 21(6):603–624, June 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lukovic:2007:ADC

[LMPR07] Ivan Luković, Pavle Mogin, Jelena Pavićević, and Sonja Ristić. An approach to developing complex database schemas using form types. Software —Practice and Experience, 37(15):1621–1656, December ??, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lander:1992:EOP

[LMSP92] Leslie C. Lander, Sandeep Mitra, Nitin Singhvi, and Thomas F. Piatkowski. The elaboration order problem of Ada. Software —Practice and Experience, 22(5):391–417, May 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Landy:1971:SET

[LN71] B. Landy and Roger M.

Needham. Software engineering techniques used in the development of the Cambridge Multiple-Access System. Software —Practice and Experience, 1(2):167–173, April/June 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2016:ILL

[LNhCW16] Wei Jun Liu, Ge Nong, Wai hong Chan, and Yi Wu. Improving a lightweight LZ77 computation algorithm for running faster. Software—Practice and Experience, 46 (9):1201–1217, September 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lang:1982:EBS

[LNW82] T. Lang, J. P. Newbury, and D. E. P. Watkins. Experiences in benchmarking for the selection of an interactive system. Software —Practice and Experience, 12(6):531–542, June 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lobelle:1985:IDW

[Lob85] Marc C. Lobelle. Integration of diskless workstations in UNIX United. Software—Practice and Experience, 15(10):997–1010, October 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Logrippo:1988:ILS

[LOBF88] L. Logrippo, A. Obaid, J. P. Briand, and M. C. Fehri. An interpreter for LOTOS, a specification language for distributed systems. Software—Practice and Experience, 18(4): 365–385, April 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Loewe:2007:EOP

[Loe07] Laurence Loewe. Evolution@home: observations on participant choice, work unit variation and low-effort global computing. Software—Practice and Experience, 37(12): 1289–1318, October 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Long:1988:BRB

[Lon88] D. Long. Book review: A practical introduction to denotational semantics, Lloyd Allison, Cambridge University Press, 1987. ISBN 0-521-31423-2 (paper-back), 0-521-30689-2 (hard cover). No. of pages: xii + 132. Price:

£22.50 (hard cover), £7.95 (paperback). Software— Practice and Experience, 18(5):493, May 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lopriore:1989:UIS

[Lop89] Lanfranco Lopriore. A user interface specification for a program debugging and measuring environment. Software—Practice and Experience, 19(5):437– 460, May 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lor:1991:ODS

[Lor91] Kar-Wing Edward Lor.

Operational definitions for system requirements as the basis of design automation. Software—Practice and Experience, 21(10): 1103–1124, October 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lindgaard:1983:HML

[LOS83] Anders Lindgård, Jørgen Oxenbøll, and Ejvind Sørensen. Hierarchical multi-level computer network for laboratory automation. Software—Practice and Experience, 13(3): 227–240, March 1983. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Leece:1978:UMS

[LP78] John Leece and Frank Parker. Use and misuse of SPSS. Software—Practice and Experience, 8(3): 301–311, May/June 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lalonde:1983:STC

[LP83] Wilf R. Lalonde and John R. Pugh. A simple technique for converting from a PASCAL shop to a C shop. Software—Practice and Experience, 13(9): 771–775, September 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Linton:1986:CSS

[LP86] A. Linton and F. Panzieri. A communication system supporting large datagrams on a local area network. Software—Practice and Experience, 16(3):277– 289, March 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Loureiro:2013:EDS

[LPA13] André Loureiro, João Paulo Porto, and Guido Araujo. Extending decoupled software pipeline to parallelize Java programs. Software —Practice and Experience, 43(5):525–541, May 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lokuciejewski:2011:APO

[LPF^{+}11] Paul Lokuciejewski, Sascha Plazar, Heiko Falk, Peter Marwedel, and Lothar Thiele. Approximating Pareto optimal compiler optimization sequences — a trade-off between WCET, ACET and code size. Software—Practice and Experience, 41(12):1437–1458, November/December 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lopez-Pintado:2019:CBP

[LPGBD^{+}19] Orlenys López-Pintado, Luciano García-Bañuelos, Marlon Dumas, Ingo Weber, and Alexander Ponomarev. Caterpillar: a business process execution engine on the Ethereum blockchain. Software— Practice and Experience, 49(7):1162–1193, July 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lin:2009:IRC

[LPP09] L. Lin, S. J. Prowell, and J. H. Poore. The impact of requirements changes on specifications and state machines. Software—Practice and Experience, 39(6): 573–610, April 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lecarme:1978:SCC

[LPT78] Olivier Lecarme and Marie-Claude Peyrolle-Thomas. Self-compiling compilers: An appraisal of their implementation and portability. Software—Practice and Experience, 8(2):149–170, March/April 1978. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lecarme:1982:CAP

[LPT82] Olivier Lecarme, Mireille Pellissier, and Marie-Claude Thomas. Computeraided production of language implementation systems: a review and classification. Software—Practice and Experience, 12(9): 785–824, September 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Loia:1993:HLM

[LQ93] Vincenzo Loia and Michel Quaggetto. High-level management of computation history for the design and implementation of a Prolog system. Software —Practice and Experience, 23(2):119–150, February 1993. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Loia:1996:OPC

[LQ96] Vincenzo Loia and Michel Quaggetto. Optimizing procedure calls for delayed non-local execution protocol. Software —Practice and Experience, 26(12):1303–1334, December 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16776.

Loia:1999:EFD

[LQ99] Vincenzo Loia and Michel Quaggetto. Embed finite domain constraint programming into Java and some Web-based applications. Software—Practice and Experience, 29(4):311– 339, April 10, 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55001842;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55001842&PLACEBO=IE.
pdf.

Li:2004:LLC

[LQ04] M. Li and M. Qi. Leveraging legacy codes to distributed problem-solving environments: a Web services approach. Software —Practice and Experience, 34(13):1297–1309, November 10, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Li:2022:SAT

[LQL^{+}22] Hui Li, Yang Qu, Yong Liu, Rong Chen, Jun Ai, and Shikai Guo. Self-admitted technical debt detection by learning its comprehensive semantics via graph neural networks. Software— Practice and Experience, 52(10):2152–2176, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2024:MDP

[LQS^{+}24] Hanwen Liu, Lianyong Qi, Shigen Shen, Arif Ali Khan, Shunmei Meng, and Qianmu Li. Microservicedriven privacy-aware crossplatform social relationship prediction based on sequential information. Software—Practice and Experience, 54(1):85–105, January 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lashkari:1993:VDM

[LRMM93] Yezdezard Lashkari, Viswanath Ramachandran, Sanjay Malpani, and S. L. Mehndiratta. Vartalaap: a distributed multicast communication system. Software —Practice and Experience, 23(7):799–811, July 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lyon:1975:STI

[LS75] Gordon Lyon and Rona B.

Stillman. Simple transforms for instrumenting FORTRAN decks. Software—Practice and Experience, 5(4):347–358, October/December 1975. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lancaster:1976:QCC

[LS76] Ronald L. Lancaster and Victor B. Schneider. Quick compiler construction using uniform code generators. Software—Practice and Experience, 6(1):83– 91, January/March 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lister:1977:HM

[LS77] A. M. Lister and P. J.

Sayer. Hierarchical monitors. Software—Practice and Experience, 7(5):613– 623, September/October 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lindstrom:1981:HMY

[LS81] Hans Lindstrom and Jan Skansholm. How to make your own simulation system. Software—Practice and Experience, 11(6):629– 637, June 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Li:1984:ISC

[LS84] Kuo-Cheng Li and Herbert D. Schwetman. Implementing a scaler C compiler on the Cyber 205. Software —Practice and Experience, 14(9):867–888, September 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lindvall:1996:PIT

[LS96a] Mikael Lindvall and Kristian Sandahl. Practical implications of traceability. Software—Practice and Experience, 26(10):1161– 1180, October 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16767.

Liu:1996:EDH

[LS96b] Xian Liu and Günther Schrack. Encoding and decoding the Hilbert order. Software—Practice and Experience, 26(12):1335– 1346, December 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16777.

Lopes:1997:TPB

[LS97] Luís M. B. Lopes and Fernando M. A. Silva. Threadand process-based implementations of the pSystem parallel programming environment. Software—Practice and Experience, 27(3): 329–351, March 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7294;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7294&
PLACEBO=IE.pdf.

Leal:2003:MWB

[LS03] José Paulo Leal and Fernando Silva. Mooshak: a Web-based multi-site programming contest system. Software—Practice and Experience, 33(6):567– 581, May 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lilis:2015:IIF

[LS15] Yannis Lilis and Anthony Savidis. An integrated implementation framework for compile-time metaprogramming. Software— Practice and Experience, 45(6):727–763, June 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2016:PPT

[LS16] Yang Liu and Andrew Simpson. Privacy-preserving targeted mobile advertising: requirements, design and a prototype implementation. Software—Practice and Experience, 46 (12):1657–1684, December 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Little:2022:EAT

[LS22] Mark C. Little and Santosh K. Shrivastava. The evolution of the Arjuna transaction processing system. Software—Practice and Experience, 52 (4):1035–1064, April 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lelli:2016:DSL

[LSAF16] Juri Lelli, Claudio Scordino, Luca Abeni, and Dario Faggioli. Deadline scheduling in the Linux kernel. Software—Practice and Experience, 46(6):821– 839, June 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Li:1994:CAF

[LSF94] Chung-Chi Jim Li, Elliot M. Stewart, and W. Kent Fuchs. Compilerassisted full checkpointing. Software—Practice and Experience, 24(10): 871–886, October 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lekidis:2018:MBD

[LSK^{+}18] Alexios Lekidis, Emmanouela Stachtiari, Panagiotis Katsaros, Marius Bozga, and Christos K. Georgiadis. Model-based design of IoT systems with the BIP component framework. Software—Practice and Experience, 48 (6):1167–1194, June 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2568.

Lemire:2016:CFS

[LSYKK16] Daniel Lemire, Gregory Ssi-Yan-Kai, and Owen Kaser. Consistently faster and smaller compressed bitmaps with Roaring. Software—Practice and Experience, 46(11):1547– 1569, November 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lamperti:2016:ECP

[LSZ16] Gianfranco Lamperti, Michele Scandale, and Marina Zanella. Extended conference papers: Determinization and minimization of finite acyclic automata by incremental techniques. Software— Practice and Experience, 46(4):513–549, April 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Layzell:1983:SDC

[LT83] P. J. Layzell and J. M. Triance. Syntax driven COBOL compilers: an assessment. Software—Practice and Experience, 13 (12):1141–1155, December 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lester:1985:SPA

[LT85] L. N. Lester and H. A. Thelander. A SIMULA program to automate wirewrap schedule generation. Software—Practice and Experience, 15(3):259– 268, March 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lins:1990:ISU

[LT90] Rafael D. Lins and Simon J. Thompson. Implementing SASL using categorical multi-combinators.

Software—Practice and Experience, 20(11):1137– 1165, November 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Levy:1991:MOD

[LT91] Henry M. Levy and Ewan D.

Tempero. Modules, objects and distributed programming: Issues in RPC and remote object invocation. Software—Practice and Experience, 21(1):77– 90, January 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Larsson:2020:IED

[LTK^{+}20] Lars Larsson, William Tärneberg, Cristian Klein, Erik Elmroth, and Maria Kihl. Impact of etcd deployment on Kubernetes, Istio, and application performance. Software—Practice and Experience, 50 (10):1986–2007, October 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Luk:2003:BLD

[LTL^{+}03] R. W. P. Luk, B. K. Y.

T’sou, T. B. Y. Lai, O. O. Y. Kwong, F. C. Y. Chik, and L. Y. L. Cheung. Bilingual legal document retrieval and management using XML. Software—Practice and Experience, 33(1):41–59, January 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
100019569/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=100019569&PLACEBO=
IE.pdf.

Linden:1996:AGP

[LTV96] Greger Lindén, Henry Tirri, and A. Inkeri Verkamo. ALCHEMIST: a general purpose transformation generator. Software—Practice and Experience, 26(6): 653–675, June 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16814.

Lu:2021:FC

[LTW^{+}21a] Qinghua Lu, An Binh Tran, Ingo Weber, Hugo O’Connor, Paul Rimba, Xiwei Xu, Mark Staples, Liming Zhu, and Ross Jeffery. Featured cover. Software —Practice and Experience, 51(5):i, May 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lu:2021:IMD

[LTW^{+}21b] Qinghua Lu, An Binh Tran, Ingo Weber, Hugo O’Connor, Paul Rimba, Xiwei Xu, Mark Staples, Liming Zhu, and Ross Jeffery. Integrated model-driven engineering of blockchain applications for business processes and asset management. Software—Practice and Experience, 51(5): 1059–1079, May 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lunbeck:1986:FUR

[Lun86] R. J. Lunbeck. File updating revisited. Software—Practice and Experience, 16(3):235–241, March 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lundberg:1989:PAS

[Lun89] Lars Lundberg. A parallel Ada system on an experimental multiprocessor. Software—Practice and Experience, 19(8):787– 800, August 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lurie:1973:SFI

[LV73] D. Lurié and C. Vandoni.

Statistics for FORTRAN identifiers and scatter storage techniques. Software —Practice and Experience, 3(2):171–177, April/June 1973. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Lammel:2001:SAG

[LV01] R. Lämmel and C. Verhoef. Semi-automatic grammar recovery. Software—Practice and Experience, 31(15):1395–1438, December 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
85515683/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=85515683&PLACEBO=IE.
pdf.

Llorens:2020:ESD

[LV20] David Llorens and Juan Miguel Vilar. Easily solving dynamic programming problems in Haskell by memoization of hylomorphisms. Software—Practice and Experience, 50(12):2193– 2211, December 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lv:2021:SIT

[Lv21] Zhihan Lv. Security of Internet of Things edge devices. Software—Practice and Experience, 51 (12):2446–2456, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:2006:UBC

[LvDDM06] Han Lee, Daniel von Dincklage, Amer Diwan, and J. Eliot B. Moss. Understanding the behavior of compiler optimizations. Software—Practice and Experience, 36(8):835– 844, July 10, 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Luckham:1984:AAB

[LvLS84] David C. Luckham, Friedrich W.

von Henke, H. J. Larsen, and D. R. Stevenson. Adam: An Ada-based language for multiprocessing. Software—Practice and Experience, 14(7):605– 642, July 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lu:2004:RXE

[LW04] Eric Jui-Lin Lu and Chang-Chuan Wu. A ReScUE XML/ EDI model. Software—Practice and Experience, 34(3):315–338, March 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liaw:2014:OOO

[LW14] Heh-Tyan Liaw and Shih-Chieh Wei. Obfuscation for object-oriented programs: dismantling instance methods. Software—Practice and Experience, 44(9):1077–1104, September 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Leszczyna:2019:TIP

[LW19] Rafał Leszczyna and Michał R. Wróbel. Threat intelligence platform for the energy sector. Software—Practice and Experience, 49(8): 1225–1254, August 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2021:MCI

[LWJ^{+}21] Richen Liu, Xiaodong Wen, Meng Jiang, Guang Yang, Chuyu Zhang, and Xiaojian Chen. Multiuser collaborative illustration and visualization for volumetric scientific data. Software—Practice and Experience, 51(5):1080–1096, May 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lu:2023:RPC

[LWW^{+}23] Shuaibing Lu, Jie Wu, Ning Wang, Yubin Duan, Haiming Liu, Jiayue Zhang, and Juan Fang. Resource provisioning in collaborative fog computing for multiple delay-sensitive users. Software—Practice and Experience, 53(2):243– 262, February 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2019:SSD

[LWZ^{+}19] Jiahao Liu, Fang Wang, Lingfang Zeng, Dan Feng, and Tingwei Zhu. SDFS: a software-defined file system for multitenant cloud storage. Software—Practice and Experience, 49(3):361– 379, March 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lu:2021:NMP

[LWZ^{+}21] Ning Lu, Bin Wang, Yongxin Zhang, Wenbo Shi, and Christian Esposito. NeuCheck: a more practical Ethereum smart contract security analysis tool. Software—Practice and Experience, 51(10): 2065–2084, October 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2004:SMS

[LX04] Xian Liu and Feilong Xu.

A softerware monitor for shared-memory multiprocessor computers. Software —Practice and Experience, 34(8):757–776, July 10, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2011:ART

[LXY^{+}11] Huai Liu, Xiaodong Xie, Jing Yang, Yansheng Lu, and Tsong Yueh Chen. Adaptive random testing through test profiles. Software—Practice and Experience, 41(10):1131–1154, September 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lien:1992:SQA

[LY92] Chih-Cheng Lien and Chien-Chiao Yang. Specification and quality assurance of timing constraints in real-time systems development. Software— Practice and Experience, 22(11):963–984, November 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Luk:2003:ACS

[LYL^{+}03] R. W. P. Luk, D. S. Yeung, Q. Lu, H. L. Leung, S. Y. Li, and F. Leung. ASAB: a Chinese screen reader. Software—Practice and Experience, 33(3):201– 219, March 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2020:RSF

[LYLY20] Chun Liu, Wei Yang, Zheng Li, and Yijun Yu. Recommending software features to designers: From the perspective of users. Software—Practice and Experience, 50 (9):1778–1792, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:2004:EJE

[LYM04] SeungIl Lee, Byung-Sun Yang, and Soo-Mook Moon. Efficient Java exception handling in just-in-time compilation. Software —Practice and Experience, 34(15):1463–1480, December 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lyon:1985:SDS

[Lyo85] Gordon Lyon. Structural dimensions of small programming environments. Software—Practice and Experience, 15(1):105–117, January 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2017:CBT

[LYX^{+}17] Jin Liu, Xiao Yu, Zheng Xu, Kim-Kwang Raymond Choo, Liang Hong, and Xiaohui Cui. A cloud-based taxi trace mining framework for smart city. Software—Practice and Experience, 47(8):1081–1094, August 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2010:SRP

[LZ10] Yongpeng Liu and Hong Zhu. A survey of the research on power management techniques for highperformance systems. Software—Practice and Experience, 40(11):943–964, October 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Longo:2020:OCI

[LZD20] Antonella Longo, Marco Zappatore, and Andrea De Matteis. An osmotic computing infrastructure for urban pollution monitoring. Software—Practice and Experience, 50(5):533– 557, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Li:2023:RVI

[LZK^{+}23] Fuyang Li, Kuan Zou, Jacky Wai Keung, Xiao Yu, Shuo Feng, and Yan Xiao. On the relative value of imbalanced learning for code smell detection. Software—Practice and Experience, 53(10):1902–1927, October 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2017:RPO

[LZL^{+}17] Zhixin Liu, Peng Zhang, Hak-Keung Lam, Kit Yan Chan, and Kai Ma. Robust power optimization scheme for cooperative wireless relay system in smart city. Software— Practice and Experience, 47(8):1045–1059, August 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2022:MAC

[LZLL22] Huaxiao Liu, Mengxi Zhang, Lei Liu, and Zhou Liu. A method to acquire cross-domain requirements based on Syntax Direct Technique. Software—Practice and Experience, 52(1):236–253, January 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2022:MPD

[LZWX22] Linfeng Liu, Houqian Zhang, Jiagao Wu, and Jia Xu. Message piece dissemination approach for opportunistic underwater sensor network invaded by underwater spy-robots. Software—Practice and Experience, 52(5):1242–1261, May 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2018:GGB

[LZZZ18] Yufeng Liu, Xiaoqin Zeng, Yang Zou, and Kang Zhang. A graph grammarbased approach for graph layout. Software—Practice and Experience, 48 (9):1523–1535, September 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Makofske:2000:RTM

[MA00] David B. Makofske and Kevin C. Almeroth. Realtime multicast tree visualization and monitoring. Software—Practice and Experience, 30(9):1047– 1065, July 25, 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        72506499/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=72506499&PLACEBO=IE.
                                        pdf.

Makofske:2001:BTI

[MA01] David B. Makofske and Kevin C. Almeroth. From broadcast television to Internet audio/video: techniques and tools for VCRstyle interactivity. Software—Practice and Experience, 31(8):781–801, July 10, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
78502909/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=78502909&PLACEBO=IE.
pdf.

Magableh:2020:DRQ

[MA20a] Basel Magableh and Muder Almiani. A deep recurrent Q network towards self-adapting distributed microservice architecture. Software— Practice and Experience, 50(2):116–135, February 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mohamed:2020:MOM

[MA20b] Amany M. Mohamed and Hisham M. Abdelsalam. A multicriteria optimization model for cloud service provider selection in multicloud environments. Software—Practice and Experience, 50(6):925–947, June 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Maass:2006:MSE

[Maa06] Moritz G. Maaß. Matching statistics: efficient computation and a new practical algorithm for the multiple common substring problem. Software—Practice and Experience, 36(3):305– 331, March 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mohan:2022:AFI

[MAA^{+}22] Senthilkumar Mohan, John A, Ahed Abugabah, Adimoolam M, Shubham Kumar Singh, Ali kashif Bashir, and Louis Sanzogni. An approach to forecast impact of Covid-19 using supervised machine learning model. Software —Practice and Experience, 52(4):824–840, April 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Macewen:1977:SSA

[Mac77a] Glenn H. Macewen. A simulation sequencing algorithm. Software—Practice and Experience, 7(2):165– 172, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Macleod:1977:DID

[Mac77b] Ian A. Macleod. Design and implementation of a display oriented text editor. Software—Practice and Experience, 7(6):771– 778, November/December 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Machura:1979:ISP

[Mac79] Marek Machura. Implementation of a specialpurpose language using Pascal implementation methodology. Software—Practice and Experience, 9(11):931– 945, November 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

MacCallum:1996:RPA

[Mac96a] Iain MacCallum. Roundpound arithmetic. Software—Practice and Experience, 26(3):293–307, March 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16796.

Machura:1996:MIC

[Mac96b] Marek Machura. Managing information in a cooperative object database system. Software—Practice and Experience, 26(5): 545–579, May 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16809.

Madsen:1979:CHL

[Mad79] Johannes Madsen. CCL — a high-level command language. Software—Practice and Experience, 9(1):25– 30, January 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Madhavji:1982:BRB

[Mad82] N. H. Madhavji. Book review: Information representation and manipulation in a computer, E. S. Page and L. B. Wilson, Cambridge University Press, Second Edition, 1978. No. of pages: 271. Price: £17.50 (hardback), £5.75 (paperback). Software—Practice and Experience, 12(7):691, July 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Madsen:1995:OIO

[Mad95] Ole Lehrmann Madsen. Open issues in objectoriented programming — a Scandinavian perspective. Software—Practice and Experience, 25(S4):3– 43, December 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mills:1991:CIS

[MAF91] Christopher Mills, Stanley C. Ahalt, and Jim Fowler. Compiled instruction set simulation. Software—Practice and Experience, 21(8):877–889, August 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Malik:2021:EEE

[MAG^{+}21] Saif U. R. Malik, Hina Akram, Sukhpal Singh Gill, Haris Pervaiz, and Hassan Malik. EFFORT: Energy efficient framework for offload communication in mobile cloud computing. Software—Practice and Experience, 51 (9):1896–1909, September 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mohamed:2015:MAB

[MAJ15] Nader Mohamed and Jameela Al-Jaroodi. MidCloud: an agent-based middleware for effective utilization of replicated Cloud services. Software—Practice and Experience, 45(3):343–363, March 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Malcolm:1980:LEM

[Mal80] James A. Malcolm. Letters to the Editor: More on publishing programs. Software—Practice and Experience, 10(7):591, July 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Malone:1983:IRT

[Mal83] J. R. Malone. Implementation of a retrospective tracing facility. Software —Practice and Experience, 13(9):791–796, September 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Malik:2011:RC

[Mal11] Petra Malik. A retrospective on CZT. Software— Practice and Experience, 41(2):179–188, February 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Malakuti:2017:MCM

[Mal17] Somayeh Malakuti. Modular composition of multiple applications with architectural event modules. Software—Practice and Experience, 47(7):1013–1025, July 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mancini:1988:TSI

[Man88]  Luigi V. Mancini.  A

technique for subclassing and its implementation exploiting polymorphic procedures. Software—Practice and Experience, 18(4): 287–300, April 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Manduchi:2001:DJA

[Man01] Gabriele Manduchi. Developing Java applications for a nuclear fusion experiment: a test case for Java applicability in a demanding environment. Software—Practice and Experience, 31(11):1025–1042, September 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
83001949/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=83001949&PLACEBO=IE.
pdf.

Mann:2018:CSI

[Man18] Zoltán Ádám Mann. Cloud simulators in the implementation and evaluation of virtual machine placement algorithms. Software—Practice and Experience, 48 (7):1368–1389, July 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2579.

Marlin:1979:HBI

[Mar79] Chris D. Marlin. A heapbased implementation of the programming language Pascal. Software—Practice and Experience, 9(2):101– 119, February 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marti:1983:LMT

[Mar83]  Jed Marti.  The little

META translator writing system. Software—Practice and Experience, 13(10): 941–959, October 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marks:1984:TPS

[Mar84a] B. L. Marks. Taming the PL/I syntax. Software—Practice and Experience, 14(8):775–789, August 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marsden:1984:SPE

[Mar84b] B. W. Marsden. A STandard Pascal event simulation package. Software —Practice and Experience, 14(7):659–684, July 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marsland:1985:MBS

[Mar85] T. A. Marsland. Management of block-structured programs. Software—Practice and Experience, 15(7): 715–723, July 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marinescu:1986:IPC

[Mar86] Dan C. Marinescu. Interprocess communication in MVS/XA and applications for scientific and engineering information processing. Software—Practice and Experience, 16(5):489– 501, May 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marshall:1988:BRB

[Mar88] L. Marshall. Book review:

C for professional programmers, Keith Tizzard, Ellis Horwood, 1986. ISBN: 0-7458-0092-0. No. of Pages:

211. Price: £18.50. Software—Practice and Experience, 18(8):805–806, August 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mahmud:2016:MQE

[MAR^{+}16] Md. Redowan Mahmud, Mahbuba Afrin, Md. Abdur Razzaque, Mohammad Mehedi Hassan, Abdulhameed Alelaiwi, and Majed Alrubaian. Maximizing quality of experience through contextaware mobile application scheduling in cloudlet infrastructure. Software —Practice and Experience, 46(11):1525–1545, November 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marowka:2022:RPP

[Mar22] Ami Marowka. Reformulation of the performance portability metric. Software—Practice and Experience, 52(1):154–171, January 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mattsson:1980:ICP

[Mat80] Sven Erik Mattsson. Implementation of concurrent Pascal on LSI-11. Software—Practice and Experience, 10(3):205–217, March 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See remarks [Hay80].

Mateti:1983:SSI

[Mat83a] Prabhaker Mateti. A specification schema for indenting programs. Software— Practice and Experience, 13(2):163–179, February 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mathewson:1983:UAD

[Mat83b] S. C. Mathewson. User acceptance: Design considerations for a program generator. Software—Practice and Experience, 13(2):101– 117, February 1983. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Maeda:1994:SCB

[MAT94a] Alfredo M. Maeda, Jun-Ichi Aoe, and Hideto Tomabechi. Signaturecheck based unification filter. Software—Practice and Experience, 24(7):603– 622, July 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Matos:1994:MMF

[Mat94b] Armando B. Matos. A matrix model for the flow of control in Prolog programs with applications to profiling. Software—Practice and Experience, 24(8):729– 746, August 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Maude:1982:RSE

[Mau82] T. Maude. Reporting syntax errors. Software—Practice and Experience, 12(5): 493, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Maurer:1992:DIG

[Mau92] Peter M. Maurer. The design and implementation of a grammar-based data generator. Software—Practice and Experience, 22(3):223– 244, March 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Maurer:2005:CCL

[Mau05] Peter M. Maurer. Converting command-line applications into binary components. Software—Practice and Experience, 35(8): 787–797, July 10, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mohebi:2016:SPI

[MAW^{+}16] Amin Mohebi, Saeed Aghabozorgi, Teh Ying Wah, Tutut Herawan, and Ramin Yahyapour. Survery papers: Iterative big data clustering algorithms: a review. Software—Practice and Experience, 46(1):107– 129, January 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mulani:1996:GNV

[MB96] Janak Mulani and Arun Bahulkar. A graphical navigator for viewing databases. Software—Practice and Experience, 26(4):411–426, April 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16801.

Mulvaney:1997:RBE

[MB97] David Mulvaney and Chris Bristow. A rule-based extension to the C++ language. Software—Practice and Experience, 27(7):747– 761, July 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7321;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7321&
PLACEBO=IE.pdf.

Morrison:1986:PGF

[MBB^{+}86] Ronald Morrison, Alfred L. Brown, P. J. Bailey, A. J. T. Davie, and Alan Dearle. A persistent graphics facility for the ICL PERQ. Software—Practice and Experience, 16(4):351– 367, April 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McKeever:2021:UML

[MBBS21] Steve McKeever, Oscar Bennich-Björkman, and Omar-Alfred Salah. Unit of measurement libraries, their popularity and suitability. Software—Practice and Experience, 51(4):711– 734, April 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Morrison:2000:CPA

[MBG^{+}00] Ron Morrison, Dharini Balasubramaniam, Mark Greenwood, Graham Kirby, Ken Mayes, Dave Munro, and Brian Warboys. A compliant persistent architecture. Software —Practice and Experience, 30(4):363–386, April 10, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        71004140/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=71004140&PLACEBO=IE.
                                        pdf.

Martens:2019:CIV

[MBG19a] Andreas Martens, Matthias Book, and Volker Gruhn. Cover image volume 49, issue 2. Software—Practice and Experience, 49(2): i, February 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Martens:2019:DDM

[MBG19b] Andreas Martens, Matthias Book, and Volker Gruhn. A data decomposition method for stepwise migration of complex legacy data. Software—Practice and Experience, 49(2):214– 232, February 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See image caption correction [Ano19a].

Mendoza:2021:AVD

[MBGC21] Camilo Mendoza, José Bocanegra, Kelly Garcés, and Rubby Casallas. Architecture violations detection and visualization in the continuous integration pipeline. Software—Practice and Experience, 51(8): 1822–1845, August 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mamrak:1997:BIS

[MBO97] Sandra A. Mamrak, John Boyd, and Iván Ordóñez. Building an information system for collaborative researchers: a case study from the brain-tumor research domain. Software—Practice and Experience, 27(3):353–363, March 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7295;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7295&
PLACEBO=IE.pdf.

Moret:2010:VEP

[MBV^{+}10] Philippe Moret, Walter Binder, Alex Villazón, Danilo Ansaloni, and Abbas Heydarnoori. Visualizing and exploring profiles with calling context ring charts. Software—Practice and Experience, 40(9):825– 847, August 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Magee:1991:PAD

[MC91] J. N. Magee and Shing-Chi C. Cheung. Parallel algorithm design for workstation clusters. Software—Practice and Experience, 21(3):235–250, March 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Moe:2002:UET

[MC02] Johan Moe and David A. Carr. Using execution trace data to improve distributed systems. Software—Practice and Experience, 32(9):889–906, July 25, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        94519526/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=94519526&PLACEBO=IE.
                                        pdf.

Mehboob:2023:MDP

[MC23] Bilal Mehboob and Chun Yong Chong. A metadata driven process for assessing stability and reusability based on risk of change of software systems. Software—Practice and Experience, 53(5): 1218–1248, May 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McCaig:1983:FPP

[McC83] J. M. McCaig. FUMBLR:

a Pascal program for initializing memory. Software—Practice and Experience, 13(12):1191–1200, December 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McCormack:1990:WFX

[McC90] Joel McCormack. Writing fast X servers for dumb color frame buffers. Software—Practice and Experience, 20(S2):S83–S108, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mi:2023:WMR

[MCCJ23] Qing Mi, Mingjie Chen, Zhi Cai, and Xibin Jia. What makes a readable code? A causal analysis method. Software—Practice and Experience, 53(6):

1391–1409, June 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McDowell:1971:BRB

[McD71] B. A. J. McDowell. Book review: A guide to Cobol programming, 2nd edn., D. D. McCracken and U. Garbassi, Wiley, New York, 1970. No. of pages: 209. Price: £3.15. Software— Practice and Experience, 1(3):301, July/September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McDonald:1987:FFU

[McD87] Chris S. McDonald. fsh — a functional UNIX command interpreter. Software —Practice and Experience, 17(10):685–700, October 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McGregor:1982:BSA

[McG82] James J. McGregor. Backtrack search algorithms and the maximal common subgraph problem. Software —Practice and Experience, 12(1):23–34, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marques:1988:DOS

[MCG^{+}88] J. Alves Marques, J. P. Cunha, P. Guedes, N. Guimarães, and A. Cunha. The distributed operating system of the SMD project. Software—Practice and Experience, 18(9):859–877, September 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McGlinn:1989:PVC

[McG89] Robert J. McGlinn. A parallel version of Cook and Kim’s algorithm for presorted lists. Software— Practice and Experience, 19(10):917–930, October 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Martinez-Carreras:2008:TIW

[MCGS08] M. Antonia Martínez-Carreras and Antonio F. Gómez-Skarmeta. Towards interoperability: a wrapper model for integrating remote laboratories in a collaborative discovery learning environment. Software —Practice and Experience, 38(15):1601–1620, December ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Myles:2005:ETS

[MCHN05] Ginger Myles, Christian Collberg, Zachary Heidepriem, and Armand Navabi. The evaluation of two software watermarking algorithms. Software—Practice and Experience, 35 (10):923–938, August 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McIlroy:1990:SPS

[McI90] M. Douglas McIlroy. Squinting at power series. Software—Practice and Experience, 20(7):661–683, July 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McIlroy:1999:KAQ

[McI99] M. D. McIlroy. A killer adversary for quicksort. Software—Practice and Experience, 29(4):341–344, April 10, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=55001843;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=55001843&PLACEBO=IE.
                                        pdf.

McKenzie:1989:FPO

[McK89] Bruce J. McKenzie. Fast peephole optimization techniques. Software—Practice and Experience, 19 (12):1151–1162, December 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McKenzie:1990:LPC

[McK90] Bruce J. McKenzie. LR parsing of CFGs with restrictions. Software—Practice and Experience, 20(8): 823–832, August 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McKenney:1999:DP

[McK99] Paul E. McKenney. Differential profiling. Software—Practice and Experience, 29(3):219–234, March 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55000311;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55000311&PLACEBO=IE.
pdf.

Mehboob:2021:RAF

[MCLL21] Bilal Mehboob, Chun Yong Chong, Sai Peck Lee, and Joanne Mun Yee Lim. Reusability affecting factors and software metrics for reusability: a systematic literature review. Software—Practice and Experience, 51(6): 1416–1458, June 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McNab:2005:GWG

[McN05] Andrew McNab. The Grid-Site Web/ Grid security system. Software—Practice and Experience, 35(9): 827–834, July 25, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McDonald:1988:SGP

[MD88] Chris McDonald and Trevor I. Dix. Support for graphs of processes in a command interpreter. Software—Practice and Experience, 18(10):1011–1016, October 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mukherjee:2019:RFE

[MDB19] Anwesha Mukherjee, Debashis De, and Rajkumar Buyya. E^{2}R-F^{2}N: Energy-efficient retailing using a femtolet-based fog network. Software—Practice and Experience, 49(3): 498–523, March 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mirachi:2017:AAM

[MdCGdC^{+}17] Samoel Mirachi, Valdir da Costa Guerra, Adilson Marques da Cunha, Luiz Alberto Vieira Dias, and Emilia Villani. Applying agile methods to aircraft embedded software: an experimental analysis. Software—Practice and Experience, 47(11):1465– 1484, November 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marback:2013:ECP

[MDH^{+}13] Aaron Marback, Hyunsook Do, Ke He, Samuel Kondamarri, and Dianxiang Xu. Extended conference papers: a threat modelbased approach to security testing. Software—Practice and Experience, 43(2):241– 258, February 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mosberger:1996:IAS

[MDP96] David Mosberger, Peter Druschel, and Larry L. Peterson. Implementing atomic sequences on uniprocessors using rollforward. Software—Practice and Experience, 26(1):1– 23, January 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61004282.

Muth:2001:ALT

[MDWD01] Robert Muth, Saumya K.

Debray, Scott Watterson, and Koen De Bosschere. alto: a link-time optimizer for the Compaq Alpha. Software—Practice and Experience, 31(1):67– 101, January 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        76502429/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=76502429&PLACEBO=IE.
                                        pdf.

Meek:1987:BRB

[Mee87] Brian Meek. Book review: A Guide to Modula-2, Kaare Christian, Springer-Verlag, 1986. ISBN 0-387-9642–5. No. of pages: 436 + xix. Price: £33.00 (hardback). Software—Practice and Experience, 17(2):167– 169, February 1987. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mei:1980:LLC

[Mei80] Tung Yun Mei. LCCD, a language for Chinese character design. Report STAN-CS-80-824, Stanford University, Department of Computer Science, 1980. ?? pp. See also [Mei81].

Mei:1981:LLC

[Mei81] Tung Yun Mei. LCCD, a language for Chinese character design. Software—Practice and Experience, 11(12):1273–1292, December 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See also [Mei80].

Mejuev:2003:DEU

[Mej03] I. Mejuev. Developing enduser tailorable Web applications using a compositional framework. Software—Practice and Experience, 33(10):913–931, August 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Menzies:1997:OOP

[Men97]  Tim Menzies.  Object-

oriented patterns: Lessons from expert systems. Software—Practice and Experience, 27(12):1457–1478, December 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7280;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7280&
PLACEBO=IE.pdf.

Moffat:1996:SFV

[MEP96] Alistair Moffat, Gary Eddy, and Ola Petersson. Splaysort: Fast, versatile, practical. Software—Practice and Experience, 26(7): 781–797, July 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16822.

Mercer:1973:BSH

[Mer73] C. A. Mercer. Buffering for sustained, high speed transfers. Software—Practice and Experience, 3(4):351–354, October/December 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mercer:1974:BRB

[Mer74] C. A. Mercer. Book review: Mini computers, hardware, software and applications, James D. Schoeffler and Ronald H. Temple, I.E.E.E. Press, 1973. No. of pages: 366. Price: £5.75. Software—Practice and Experience, 4(4):408, October/December 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McKeag:1984:DSC

[MER84] R. M. McKeag, Ann Eves, and E. J. Rooney. Designing software for children. Software—Practice and Experience, 14(5):451– 461, May 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Merrill:1993:PNL

[Mer93] Gary H. Merrill. Parsing non-LR(k) grammars with yacc. Software—Practice and Experience, 23(8):829– 850, August 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Messerschmidt:1980:PPC

[Mes80] H. J. Messerschmidt. Parallel programming for a chess endgame database. Software—Practice and Experience, 10(6):475–487, June 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Messerschmidt:1996:LIC

[Mes96] H. J. Messerschmidt. List iterators in C++. Software—Practice and Experience, 26(11):1197–1203, November 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16770.

Meyer:1978:NCM

[Mey78] Bertrand Meyer. A note on computing multiple sums. Software—Practice and Experience, 8(1):1– 9, January/February 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Monteiro:2008:IER

[MF08] Miguel P. Monteiro and João M. Fernandes. An illustrative example of refactoring object-oriented source code with aspectoriented mechanisms. Software—Practice and Experience, 38(4):361–396, April 10, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mezni:2018:TAS

[MF18] Haithem Mezni and Mayssa Fayala. Time-aware service recommendation: Taxonomy, review, and challenges. Software—Practice and Experience, 48 (11):2080–2108, November 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mili:2002:EFI

[MFB^{+}02] Hafedh Mili, Mohamed Fayad, Davide Brugali, David Hamu, and Dov Dori. Enterprise frameworks: issues and research directions. Software—Practice and Experience, 32(8):801–831, July 10, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
94519521/START;   http:
//www3.interscience.wiley.

com/cgi-bin/fulltext?
ID=94519521&PLACEBO=IE.
pdf.

Monzon:2012:ADR

[MFdlP12] Antonio Monzón, José L.

Fernández, and Juan A. de la Puente. Application of deadlock risk evaluation of architectural models. Software—Practice and Experience, 42 (9):1137–1163, September 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Meister:2010:SCI

[MFH10] Jeffrey A. Meister, Jeffrey S. Foster, and Michael Hicks. Serializing C intermediate representations for efficient and portable parsing. Software—Practice and Experience, 40(3): 225–238, March ??, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Morita:2001:FIM

[MFYiA01] Kazuhiro Morita, Masao Fuketa, Yoshihiro Yamakawa, and Jun ichi Aoe. Fast insertion methods of a double-array structure. Software—Practice and Experience, 31(1):43– 65, January 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.

                                        com/cgi-bin/abstract/
                                        76502428/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=76502428&PLACEBO=IE.
                                        pdf.

McClure:1976:MCC

[MG76] Robert M. McClure and Richard L. Gauthier. Minivital: a compiler case study. Software—Practice and Experience, 6(2):225– 237, April/June 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

MacGregor:1994:OSP

[MG94] M. H. MacGregor and W. D. Grover. Optimized κ-shortest-paths algorithm for facility restoration. Software—Practice and Experience, 24(9):823– 834, September 1994. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Milano:2009:BEC

[MG09] Michela Milano and Alessio Guerri. Bid evaluation in combinatorial auctions: optimization and learning. Software—Practice and Experience, 39(13): 1127–1155, September 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Monteiro:2013:IDP

[MG13] Miguel P. Monteiro and João Gomes. Implementing design patterns in Object Teams. Software—Practice and Experience, 43 (12):1519–1551, December 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marimuthu:2022:IAF

[MGBO22] Satheesh Kumar Marimuthu, Srinivasagan Kalampatti Gopalasamy, and Jalel Ben-Othman. Intelligent antiphishing framework to detect phishing scam: a hybrid classification approach. Software—Practice and Experience, 52(2):459– 481, February 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Milanovic:2009:BCA

[MGG^{+}09] Milan Milanović, Dragan Gašević, Adrian Giurca, Gerd Wagner, and Vladan Devedžić. Bridging concrete and abstract syntaxes in model-driven engineering: a case of rule languages. Software—Practice and Experience, 39(16): 1313–1346, November ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Milazzo:2018:KDM

[MGGS18] Fabrizio Milazzo, Vito Gentile, Antonio Gentile, and Salvatore Sorce. KIND-DAMA: a modular middleware for Kinectlike device data management. Software—Practice and Experience, 48(1):141– 160, January 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Munshi:2024:QML

[MGJ^{+}24] Manushi Munshi, Rajesh Gupta, Nilesh Kumar Jadav, Zdzislaw Polkowski, Sudeep Tanwar, Fayez Alqahtani, and Wael Said. Quantum machine learningbased framework to detect heart failures in Healthcare 4.0. Software—Practice and Experience, 54(2): 168–185, February 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marijan:2019:LAO

[MGL19] Dusica Marijan, Arnaud Gotlieb, and Marius Liaaen. A learning algorithm for optimizing continuous integration development and testing practice. Software—Practice and Experience, 49(2):192– 213, February 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Malloy:2003:DTF

[MGP03] Brian A. Malloy, Tanton H. Gibbs, and James F. Power. Decorating tokens to facilitate recognition of ambiguous language constructs. Software—Practice and Experience, 33(1): 19–39, January 2003. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
100019568/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=100019568&PLACEBO=
IE.pdf.

Moreno:2020:ASR

[MGS^{+}20] Julio Moreno, Javier Gómez, Manuel A. Serrano, Eduardo B. Fernandez, and Eduardo Fernández-Medina. Application of security reference architecture to Big Data ecosystems in an industrial scenario. Software —Practice and Experience, 50(8):1520–1538, August 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mazidi:2020:ARP

[MGT20] Arash Mazidi, Mehdi Golsorkhtabaramiri, and Meisam Yadollahzadeh Tabari. Autonomic resource provisioning for multilayer cloud applications with K-nearest neighbor resource scaling and priority-based resource allocation. Software—Practice and Experience, 50(8): 1600–1625, August 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McMullin:1982:ICB

[MGW82] Paul R. McMullin, John D. Gannon, and Mark D. Weiser. Implementing a compiler-based test tool. Software—Practice and Experience, 12(10):971– 979, October 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Maccari:2005:MIV

[MH05] Alessandro Maccari and Anders Heie. Managing infinite variability in mobile terminal software. Software —Practice and Experience, 35(6):513–537, May 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mhashi:2005:EMR

[Mha05] Mahmoud Moh’d Mhashi. The effect of multiple reference characters on detecting matches in stringsearching algorithms. Software—Practice and Experience, 35(13):1299–1315, November 10, 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McKenzie:1990:SHA

[MHB90] Bruce J. McKenzie, R. Harries, and Timothy C. Bell. Selecting a hashing algorithm. Software—Practice and Experience, 20(2):209– 224, February 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marshall:2001:OOD

[MHM01] M. S. Marshall, I. Herman, and G. Melançon. An object-oriented design for graph visualization. Software—Practice and Experience, 31(8):739–756, July 10, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
78502908/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=78502908&PLACEBO=IE.
pdf.

Malekhosseini:2018:EUP

[MHN18] Razieh Malekhosseini, Mehdi Hosseinzadeh, and Keyvan Navi. Evaluation of users’ privacy concerns by checking of their WhatsApp status. Software—Practice and Experience, 48 (5):1143–1164, May 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2565.

Morimoto:1994:MCT

[MIA94] Katsushi Morimoto, Hirokazu Iriguchi, and Jun-Ichi Aoe. A method of compressing trie structures. Software—Practice and Experience, 24(3):265– 288, March 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Middleton:1979:AGA

[Mid79] Tony Middleton. Automatic generation of aggregate operators in ALGOL 68. Software—Practice and Experience, 9(8):685–690, August 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Middleton:1986:RAA

[Mid86] A. G. Middleton. Routines: An argument against the conventional approach to functions and procedures. Software—Practice and Experience, 16(2):119– 130, February 1986. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mills:1972:BRB

[Mil72] J. R. Mills. Book review: A guide to packaged systems, Robert V. Head, Wiley, New York, 1971. No. of pages: 116. Price: £3.75. Software—Practice and Experience, 2(4):406, October 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Milne:1974:SIG

[Mil74] W. J. Milne. A simple interactive graphics facility. Software—Practice and Experience, 4(1):71– 78, January/March 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Milkowski:2010:DOS

[Mił10] Marcin Miłkowski. Developing an open-source, rule-based proofreading tool. Software—Practice and Experience, 40(7):543– 566, June 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Muxworthy:1978:LE

[MIR78] David T. Muxworthy, J. Inglis, and Lawrence A. Rowe. Letter to the editor. Software—Practice and Experience, 8(2):236– 237, March 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mitchell:1973:LE

[Mit73] Edward E. L. Mitchell. Letter to the editor. Software —Practice and Experience, 3(3):303, July/September 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mateti:1983:CPI

[MJ83] Prabhaker Mateti and Joxan Jaffar. A correctness proof of an indenting program. Software—Practice and Experience, 13(3):199– 226, March 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Miles:1998:IGO

[MJ98] William S. Miles and Leroy F. Johnson. Implementing generalized operator overloading. Software—Practice and Experience, 28(6):593–610, May 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=1799;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=1799&
                                        PLACEBO=IE.pdf.

Meehan:1999:CLF

[MJ99] Gary Meehan and Mike Joy. Compiling lazy functional programs to Java bytecode. Software—Practice and Experience, 29(7): 617–645, June 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=62003575;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=62003575&PLACEBO=IE.
pdf.

Ma:1990:TTI

[MK90] Kwan-Liu Ma and Robert R.

Kessler. TICL — a type inference system for Common Lisp. Software—Practice and Experience, 20(6):593– 623, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mossenbock:1996:ATS

[MK96] Hanspeter Mössenböck and Kai Koskimies. Active text for structuring and understanding source code. Software—Practice and Experience, 26(7):833– 850, July 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16821.

Manolache:2001:STU

[MK01] L. I. Manolache and D. G. Kourie. Software testing using model programs. Software—Practice and Experience, 31 (13):1211–1236, November 10, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        85011926/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=85011926&PLACEBO=IE.
                                        pdf.

Muller:2003:GPA

[MK03] Jens Müller and Jan G. Korvink. A general purpose adaptivity driver for FE software. Software— Practice and Experience, 33(12):1097–1116, October 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mann:2004:TSL

[MK04] Zoltán Ádám Mann and Károly Kondorosi. Tracing system-level communication in distributed systems. Software—Practice and Experience, 34(8):727– 755, July 10, 2004. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Momeni:2018:LAR

[MK18] Behnam Momeni and Mehdi Kharrazi. LDMBL: an architecture for reducing code duplication in heavyweight binary instrumentations. Software—Practice and Experience, 48(9):1642–1659, September 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mahmood:2022:SEE

[MKA^{+}22] Yasir Mahmood, Nazri Kama, Azri Azmi, Ahmad Salman Khan, and Mazlan Ali. Software effort estimation accuracy prediction of machine learning techniques: a systematic performance evaluation. Software—Practice and Experience, 52(1):39– 65, January 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Malik:2011:SMP

[MKC11] Najmus Saqib Malik, David Ko, and Harry H. Cheng. A secure migration process for mobile agents. Software—Practice and Experience, 41(1):87–101, January 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Martin:2020:CCR

[MKC20] John Paul Martin, A. Kandasamy, and K. Chandrasekaran. CREW: Cost and Reliability aware Eagle–Whale optimiser for service placement in Fog. Software—Practice and Experience, 50(12):2337– 2360, December 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mikkilineni:1998:ERO

[MKD98] Rani Mikkilineni, M. A. Ketabchi, and S. Dasananda. External relationships of objects. Software—Practice and Experience, 28(7): 737–747, July 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1806;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1806&
PLACEBO=IE.pdf.

McCarthy:2022:BLG

[MKD^{+}22] Harry McCarthy, Abigail Koay, Michael Dawson, Kenneth B. Kent, and Panos Patros. Benchmarking and learning garbage collection delays for resource-restricted graphical user interfaces. Software—Practice and Experience, 52(1):172–193, January 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mashkoor:2018:ESS

[MKE18] Atif Mashkoor, Felix Kossak, and Alexander Egyed. Evaluating the suitability of state-based formal methods for industrial deployment. Software—Practice and Experience, 48 (12):2350–2379, December 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mohammed:2017:FSF

[MKM^{+}17] Bashir Mohammed, Mariam Kiran, Kabiru M. Maiyama, Mumtaz M. Kamala, and Irfan-Ullah Awan. Failover strategy for fault tolerance in cloud computing environment. Software—Practice and Experience, 47 (9):1243–1274, September 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Moeyersons:2022:TCB

[MKW^{+}22] Jerico Moeyersons, Sarah Kerkhove, Tim Wauters, Filip De Turck, and Bruno Volckaert. Towards cloudbased unobtrusive monitoring in remote multi-vendor environments. Software— Practice and Experience, 52(2):427–442, February 2022. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Mahmood:2008:CMU

[ML08] Sajjad Mahmood and Richard Lai. A complexity measure for UML componentbased system specification. Software—Practice and Experience, 38(2):117– 134, February ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mula:2020:BED

[ML20] Wojciech Muła and Daniel Lemire. Base64 encoding and decoding at almost the speed of a memory copy. Software—Practice and Experience, 50(2): 89–97, February 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mushtak:2023:FNP

[ML23] Noble Mushtak and Daniel Lemire. Fast number parsing without fallback. Software—Practice and Experience, 53(7):1467–1471, July 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See original work [Lem21].

Markiewicz:2002:TAC

[MLC02] Marcus E. Markiewicz, Carlos J. P. Lucena, and Donald D. Cowan. Taming access control security: extending capabilities using the views relationship. Software—Practice and Experience, 32(4):341– 358, April 10, 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
91016432/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=91016432&PLACEBO=IE.
pdf.

Maffione:2019:CAD

[MLR19] Vincenzo Maffione, Giuseppe Lettieri, and Luigi Rizzo. Cache-aware design of general-purpose Single-Producer-Single-Consumer queues. Software—Practice and Experience, 49(5):748– 779, May 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mendes:2018:TCH

[MLV18] Joaquim Mendes, Nuno Laranjeiro, and Marco Vieira. Toward characterizing HTML defects on the Web. Software—Practice and Experience, 48 (3):750–757, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/abs/10.
                                        1002/spe.2545.

McGregor:1980:SDI

[MM80a] D. R. McGregor and J. R. Malone. Stabdump — a dump interpreter program to assist debugging. Software—Practice and Experience, 10(4):329–332, April 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McKeag:1980:EPP

[MM80b] R. M. McKeag and P. Milligan. An experiment in parallel program design. Software—Practice and Experience, 10(9):687– 696, September 1980. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McGregor:1981:DRS

[MM81] D. R. McGregor and J. R. Malone. Design for a robust, simple and highly reliable filestore. Software —Practice and Experience, 11(9):943–947, September 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McGregor:1982:FTF

[MM82] D. R. McGregor and J. A. Mariani. ‘Fingerprinting’ — a technique for file identification and maintenance. Software—Practice and Experience, 12 (12):1165–1166, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Miller:1985:FCP

[MM85] Webb Miller and Eugene W. Myers. A file comparison program. Software—Practice and Experience, 15(11):1025–1040, November 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Miller:1986:SEA

[MM86] Webb Miller and Eugene W. Myers. Side-effects in automatic file updating. Software—Practice and Experience, 16(9):809– 820, September 1986. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Miller:1988:SRR

[MM88] Webb Miller and Eugene W. Myers. A simple row-replacement method. Software—Practice and Experience, 18(7): 597–611, July 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mullin:1990:TTS

[MM90] James K. Mullin and Daniel J. Margoliash. A tale of three spelling checkers. Software—Practice and Experience, 20(6):625– 630, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Montague:1997:SRP

[MM97] Bruce R. Montague and Charles E. Mcdowell. Synchronous/reactive programming of concurrent system software. Software—Practice and Experience, 27(3): 207–243, March 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7289;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7289&
PLACEBO=IE.pdf.

Morin:2001:GTT

[MM01] Johanne Morin and Stan Matwin. GenEx: a tool for testing in ILP. Software—Practice and Experience, 31(10):1003–1023, August 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
82002143/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=82002143&PLACEBO=IE.
pdf.

Manzini:2002:OOI

[MM02] G. Manzini and S. Mazet.

An object-oriented interface for the dynamic memory management of sparse discrete mathematical operators in numerical scientific applications. Software—Practice and Experience, 32(7):621–644, June 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
94515734/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=94515734&PLACEBO=IE.
pdf.

McIlwain:2006:TCL

[MM06] Sean McIlwain and Barton P. Miller. A tool for converting Linux device drivers into Solaris compatible binaries. Software —Practice and Experience, 36(7):689–710, June 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Matera:2003:MDD

[MMCF03] M. Matera, A. Maurino, S. Ceri, and P. Fraternali. Model-driven design of collaborative Web applications. Software—Practice and Experience, 33(8): 701–732, July 10, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Meling:2008:JAD

[MMHB08] Hein Meling, Alberto Montresor, Bjarne E. Helvik, and Ozalp Babaoglu. Jgroup/ ARM: a distributed object group platform with autonomous replication management. Software—Practice and Experience, 38(9): 885–923, July 25, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Milenkovic:2004:MDB

[MMK04] Milena Milenkovic, Aleksandar Milenkovic, and Jeffrey Kulick. Microbenchmarks for determining branch predictor organization. Software—Practice and Experience, 34(5):465– 487, April 25, 2004. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Martin:2018:ABP

[MMM18] Alexis Martin and Vania Marangozova-Martin. Automatic benchmark profiling through advanced workflow-based trace analysis. Software—Practice and Experience, 48 (6):1195–1217, June 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.

wiley.com/doi/abs/10.
1002/spe.2570.

Moitra:1979:DAH

[MMN79] Abha Moitra, S. P. Mudur, and A. W. Narwekar. Design and analysis of a hyphenation procedure. Software—Practice and Experience, 9(4):325–337, April 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Maenhaut:2016:ECP

[MMOD16] Pieter-Jan Maenhaut, Hendrik Moens, Veerle Ongenae, and Filip De Turck. Extended conference papers: Migrating legacy software to the cloud: approach and verification by means of two medical software use cases. Software —Practice and Experience, 46(1):31–54, January 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Miller:1986:DPM

[MMS86] Barton P. Miller, Cathryn Macrander, and Stuart Sechrest. A distributed programs monitor for Berkeley UNIX. Software —Practice and Experience, 16(2):183–200, February 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Moser:1990:FVS

[MMS90] Louise E. Moser and P. M. Melliar-Smith. Formal verification of safety-critical systems. Software—Practice and Experience, 20(8): 799–821, August 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Muramatsu:1980:PRA

[MN80] Hiroshi Muramatsu and Hiroaki Negishi. Page replacement algorithm for large-array manipulation. Software—Practice and Experience, 10(7):575–587, July 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mertz:2018:AAL

[MN18] Jhonny Mertz and Ingrid Nunes. Automation of application-level caching in a seamless way. Software—Practice and Experience, 48 (6):1218–1237, June 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2571.

Mozaffari:2021:CCO

[MNEM21] Maryam Mozaffari, Eslam Nazemi, and Amir-Masoud Eftekhari-Moghadam. CONST:

Continuous online NoSQL schema tuning. Software—Practice and Experience, 51(5):1147–1169, May 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Makaroff:2004:PEV

[MNH04] D. J. Makaroff, G. W.

Neufeld, and N. C. Hutchinson. Performance evaluation for VBR Continuous Media File Server admission control. Software —Practice and Experience, 34(12):1187–1210, October 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mudur:1979:DST

[MNM79] S. P. Mudur, A. W. Narwekar, and Abha Moitra. Design of software for text composition. Software— Practice and Experience, 9(4):313–323, April 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mistry:2014:ERA

[MNW14] James Mistry, Matthew Naylor, and Jim Woodcock. Experience report: Adapting FreeRTOS for multicores: an experience report. Software—Practice and Experience, 44 (9):1129–1154, September 2014. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Moffat:1989:WBT

[Mof89] Alistair Moffat. Wordbased text compression. Software—Practice and Experience, 19(2):185–198, February 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Moffat:1999:IDS

[Mof99] Alistair Moffat. An improved data structure for cumulative probability tables. Software—Practice and Experience, 29(7):647– 659, June 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=62003576;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=62003576&PLACEBO=IE.
                                        pdf.

Mogul:2004:CFH

[Mog04] Jeffrey C. Mogul. Clarifying the fundamentals of HTTP. Software—Practice and Experience, 34(2):103– 134, February 2004. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mohilner:1977:UPF

[Moh77] Patricia R. Mohilner. Using Pascal in a Fortran environment. Software— Practice and Experience, 7(3):357–362, May/June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mohanty:1981:SCE

[Moh81] Siba N. Mohanty. Software cost estimation: Present and future. Software— Practice and Experience, 11(2):103–121, February 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Monfroglio:1996:HGA

[Mon96a] A. Monfroglio. Hybrid genetic algorithms for a rostering problem. Software—Practice and Experience, 26(7):851–862, July 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16824.

Monfroglio:1996:TTC

[Mon96b] Angelo Monfroglio. Timetabling through constrained heuristic search and genetic algorithms. Software—Practice and Experience, 26(3): 251–279, March 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16797.

Morrison:1977:MIP

[Mor77] B. Morrison. A method of implementing procedure entry and exit in block structured high level language. Software—Practice and Experience, 7(4):537– 539, July 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Morris:1980:PSR

[Mor80] James B. Morris. Programming by successive refinement of data abstractions. Software—Practice and Experience, 10(4):249– 263, April 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Morrison:1982:LCC

[Mor82] Ronald Morrison. Low cost computer graphics for micro computers. Software—Practice and Experience, 12(8):767–776, August 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mosedale:1973:PCS

[Mos73] T. W. Mosedale. PEDANT: A computerized support to program modularity under limited memory conditions. Software—Practice and Experience, 3(2): 121–143, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mossenbock:1988:CWI

[Mös88] Hanspeter Mössenböck. A convenient way to incorporate semantic actions in two-pass compiling schemes. Software—Practice and Experience, 18(7): 691–700, July 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mostefaoui:2006:MAF

[Mos06] A. Mostefaoui. A modular and adaptive framework for large scale video indexing and content-based retrieval: the SIRSALE system. Software—Practice and Experience, 36(8):871– 890, July 10, 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Motzkin:1981:SQ

[Mot81] Dalia Motzkin. A stable Quicksort. Software —Practice and Experience, 11(6):607–611, June 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mora:2018:DMS

[MOTG18] M. Mora, R. V. O’Connor, F. Tsui, and J. Marx Gómez. Design methods for software architectures in the serviceoriented computing and cloud paradigms. Software—Practice and Experience, 48(2):263–267, February 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2547.

Moudry:1972:NDC

[Mou72] J. Moudrý. A notation describing corrections in files. Software—Practice and Experience, 2(3):279– 285, July/September 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Meeson:1979:OFP

[MP79] Reginald Meeson and Arthur Pyster. Overhead in FORTRAN preprocessors. Software—Practice and Experience, 9(12):987– 999, December 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

MacLean:1981:CVP

[MP81] M. A. MacLean and J. E. L. Peck. CHEF: a versatile portable text editor. Software—Practice and Experience, 11(5):467– 477, May 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mezzalama:1982:MIA

[MP82] M. Mezzalama and P. Prinetto. A machine-independent approach to microprogram synthesis. Software—Practice and Experience, 12 (11):985–1010, November 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Minsky:2000:PMV

[MP00] Naftaly H. Minsky and Partha Pratim Pal. Providing multiple views for objects. Software—Practice and Experience, 30(7): 803–823, June 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72500203/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72500203&PLACEBO=IE.
pdf.

Mamrak:2002:AFG

[MP02] Sandra A. Mamrak and Sheetal Pole. Automatic form generation. Software—Practice and Experience, 32(11):1051–1063, September 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        97519256/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=97519256&PLACEBO=IE.
                                        pdf.

Mariani:2013:MTI

[MP13] L. Mariani and F. Pastore. MASH: tool integration made easy. Software —Practice and Experience, 43(4):419–433, April 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marton:2018:SFC

[MP18] Gábor Márton and Zoltán Porkoláb. Selective friends in C++. Software—Practice and Experience, 48(8): 1493–1519, August 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2587.

Malloy:2019:GEM

[MP19] Brian A. Malloy and James F. Power. Grammar engineering for multiple front-ends for Python. Software—Practice and Experience, 49(3):380–400, March 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Malohlava:2013:IDS

[MPBH13] Michal Malohlava, Frantisek Plasil, Tomas Bures, and Petr Hnetynka. Interoperable domain-specific languages families for code generation. Software— Practice and Experience, 43(5):479–499, May 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mandal:2019:SAA

[MPC^{+}19] Amit Kr Mandal, Federica Panarotto, Agostino Cortesi, Pietro Ferrara, and Fausto Spoto. Static analysis of Android Auto infotainment and on-board diagnostics II apps. Software—Practice and Experience, 49(7):1131–1161, July 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mugarza:2020:CFZ

[MPJ20] Imanol Mugarza, Jorge Parra, and Eduardo Jacob. Cetratus: a framework for zero downtime secure software updates in safetycritical systems. Software —Practice and Experience, 50(8):1399–1424, August 2020. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

McCluskey:1995:RCM

[MPN^{+}95] T. L. McCluskey, J. M. Porteous, Y. Naik, C. N. Taylor, and S. Jones. A requirements capture method and its use in an air traffic control application. Software—Practice and Experience, 25(1):47–71, January 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Miller:1987:DMD

[MPP87] Barton P. Miller, Michael L. Powell, and David L. Presotto. DEMOS/MP: The development of a distributed operating system. Software—Practice and Experience, 17(4):277– 290, April 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marcen:2022:EBE

[MPPC22] Ana C. Marcén, Francisca Pérez, Óscar Pastor, and Carlos Cetina. Evaluating the benefits of empowering model-driven development with a machine learning classifier. Software—Practice and Experience, 52(11):2439–2455, November 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mishra:1993:EMC

[MPS93] Shivakant Mishra, Larry L.

Peterson, and Richard D. Schlichting. Experience with modularity in Consul. Software—Practice and Experience, 23(10): 1059–1075, October 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Moody:1980:CMB

[MR80] Ken Moody and Martin Richards. A coroutine mechanism for BCPL. Software—Practice and Experience, 10(10):765–771, October 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McIlroy:1992:MSU

[MR92] M. D. McIlroy and J. A.

Reeds. Multilevel security in the UNIX tradition. Software—Practice and Experience, 22(8):673– 694, August 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mascarenhas:1996:AAP

[MR96] Edward Mascarenhas and Vernon Rego. Ariadne: Architecture of a portable threads system supporting thread migration. Software—Practice and Experience, 26(3):327–356, March 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16793.

Manzini:2004:SFD

[MR04] Giovanni Manzini and Marcella Rastero. A simple and fast DNA compressor. Software—Practice and Experience, 34(14): 1397–1411, November 25, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Milne:2005:ICC

[MR05] Iain Milne and Glenn Rowe. Interpreting computer code in a computerbased learning system for novice programmers. Software—Practice and Experience, 35(15):1477–1493, December 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ma:2007:DAI

[MR07] Haoxue Ma and Tore Risch. A database approach for information communication in a peer-to-peer collaborative CAD environment. Software—Practice and Experience, 37 (11):1193–1213, September 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marchezan:2019:PFR

[MRBB19] Luciano Marchezan, Elder Macedo Rodrigues, Maicon Bernardino, and Fábio Paulo Basso. PAxSPL: a feature retrieval process for software product line reengineering. Software— Practice and Experience, 49(8):1278–1306, August 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Moreira:2019:DSS

[MRG^{+}19] Andre Moreira, Daniel Rosendo, Demis Gomes, Guto Leoni Santos, Leylane Silva, Carolina Cani, Judith Kelner, Djamel Sadok, Glauco Gonçalves, Amardeep Mehta, Mattias Wildeman, and Patricia Takako Endo. DCAV: a software system to evaluate next-generation cloud data center availability through a friendly graphical interface. Software—Practice and Experience, 49 (11):1573–1599, November 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mark:1992:IMN

[MRNL92] Leo Mark, Nick Roussopoulos, Tina Newsome, and Piriya Laohapipattana. Incrementally maintained network–relational database mapping. Software—Practice and Experience, 22(12):1099–1131, December 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Masmano:2008:ICT

[MRR^{+}08] M. Masmano, I. Ripoll, J. Real, A. Crespo, and A. J. Wellings. Implementation of a constanttime dynamic storage allocator. Software—Practice and Experience, 38(10): 995–1026, August ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mateos:2015:TIC

[MRZ15] Cristian Mateos, Juan Manuel Rodriguez, and Alejandro Zunino. A tool to improve code-first Web services discoverability through text mining techniques. Software—Practice and Experience, 45(7):925–948, July 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Maccallum:1974:MLS

[MS74a] K. J. Maccallum and L. T. Shafe. A mixed language system POP-2 and FOR-TRAN. Software—Practice and Experience, 4(2): 145–154, April/June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marcotty:1974:SPL

[MS74b] Michael Marcotty and Henrik Schutz. The systems programming language, Malus. Software— Practice and Experience, 4 (1):79–90, January/March 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marsland:1980:HDP

[MS80a] T. A. Marsland and S. F.

Sutphen. A heterogeneous dual processor. Software —Practice and Experience, 10(1):21–28, January 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Munn:1980:RPW

[MS80b] R. J. Munn and J. M. Stewart. RATMAC: a preprocessor for writing portable scientific software. Software—Practice and Experience, 10(9):743–749, September 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Murali:1983:SGC

[MS83] V. Murali and R. K. Shyamasundar. A sentence generator for a compiler for PT, a PASCAL subset. Software—Practice and Experience, 13(9):857– 869, September 1983. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Malloy:1990:CSP

[MS90] Brian Malloy and Mary Lou Soffa. Conversion of simulation processes to Pascal constructs. Software —Practice and Experience, 20(2):191–207, February 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Musser:1994:AOG

[MS94] David R. Musser and Alexander A. Stepanov. Algorithm-oriented generic libraries. Software—Practice and Experience, 24(7): 623–642, July 1994. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mummert:1996:LTD

[MS96] L. Mummert and M. Satyanarayanan. Long term distributed file reference tracing: Implementation and experience. Software—Practice and Experience, 26(6):705–736, June 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16815.

McAllister:1998:ADA

[MS98] Andrew J. McAllister and David Sharpe. An approach for decomposing N-ary data relationships. Software—Practice and Experience, 28(2):125–154, February 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1774;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1774&
PLACEBO=IE.pdf.

Mamrak:1999:CSP

[MS99] Sandra A. Mamrak and Saurabh Sinha. A case study: productivity and quality gains using an object-oriented framework. Software—Practice and Experience, 29(6):501–518, May 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61000494;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=61000494&PLACEBO=IE.
pdf.

Mezni:2018:NBS

[MS18] Haithem Mezni and Mokhtar Sellami. A negotiationbased service selection approach using swarm intelligence and kernel density estimation. Software—Practice and Experience, 48 (6):1285–1311, June 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2575.

Malhotra:2022:SNN

[MS22] Sheenam Malhotra and Williamjeet Singh. A secure neural networkbased ranking approach for document searching in cloud data center. Software—Practice and Experience, 52(9):1927–1946, September 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marx:2013:RRR

[MSB^{+}13] Edgard Marx, Percy Salas, Karin Breitman, José Viterbo, and Marco Antonio Casanova. RDB2RDF: a relational to RDF plugin for Eclipse. Software —Practice and Experience, 43(4):435–447, April 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Moss:2018:CAM

[MSB18] Aaron Moss, Robert Schluntz, and Peter A. Buhr. C: Adding modern programming language features to C. Software—Practice and Experience, 48(12):2111– 2146, December 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Maciel:2020:MTS

[MSB20] Bruno I. F. Maciel, Silas G.

T. C. Santos, and Roberto S. M. Barros. MOA-Manager: a tool to support data stream experiments. Software—Practice and Experience, 50(4):325– 334, April 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mao:2023:LAR

[MSC23] Chengying Mao, Yue Song, and Jifu Chen. A lightweight adaptive random testing method for deep learning systems. Software—Practice and Experience, 53(11):2271– 2295, November 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McKenney:2001:EEP

[MSK01] Paul E. McKenney, Jack Slingwine, and Phil Krueger. Experience with an efficient parallel kernel memory allocator. Software—Practice and Experience, 31(3): 235–257, March 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/

                                        76507638/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=76507638&PLACEBO=IE.
                                        pdf.

Muller:2024:GBF

[MSK^{+}24] Sebastian Müller, Jan Arne Sparka, Martin Kuban, Claudia Draxl, and Lars Grunske. Grammar-based fuzzing of data integration parsers in computational materials science. Software —Practice and Experience, 54(2):208–224, February 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mishra:2007:PSE

[MSR^{+}07] Dhananjay Mishra, Clifford A. Shaffer, Naren Ramakrishnan, Layne T. Watson, Kyung K. Bae, Jian He, Alex A. Verstak, and William H. Tranter. S^{4}W: a problem-solving environment for wireless system design. Software— Practice and Experience, 37(14):1539–1558, November 25, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Menzel:2013:MCR

[MST13] Michael Menzel, Marten Schönherr, and Stefan Tai. (MC^{2})^{2}: criteria, requirements and a software prototype for Cloud infrastructure decisions. Software—Practice and Experience, 43(11):1283–1297, November 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mirampalli:2023:HFB

[MSWC23]  Sreenivasu  Mirampalli,

Satish Narayana Srirama, Rajeev Wankar, and Raghavendra Rao Chillarige. Hierarchical fuzzy-based Quality of Experience (QoE)-aware application placement in fog nodes. Software—Practice and Experience, 53(2): 263–282, February 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Muhammad:2024:ALM

[MSZ^{+}24] Ahmad Ahmad Muhammad, Abdelrahman Soliman, Hala Zayed, Ahmed H. Yousef, and Sahar Selim. Automated library mapping approach based on cross-platform for mobile development programming languages. Software —Practice and Experience, 54(5):683–703, May 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mucke:1978:IRL

[MT78] F. U. Mucke and W. Trier.

Improvement of relocatable linkage editor (loader) input. Software—Practice and Experience, 8(4):417– 419, July/August 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mincy:1984:PVL

[MT84a] Jeffrey W. Mincy and Alan L. Tharp. Processing variable-length abbreviations. Software—Practice and Experience, 14(4): 397–398, April 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mullender:1984:IF

[MT84b] Sape J. Mullender and Andrew S. Tanenbaum. Immediate files. Software —Practice and Experience, 14(4):365–368, April 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Maskit:1994:MDP

[MT94] Daniel Maskit and Stephen Taylor. A message-driven programming system for fine-grain multicomputers. Software—Practice and Experience, 24(10): 953–980, October 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Maccarone:1993:PPD

[MTdT93] Maria Concetta Maccarone, Mario Tripiciano, Vito di Gesù, and Domenico Tegolo. PDB: a pictorial database oriented to data analysis. Software—Practice and Experience, 23(1): 55–73, January 1993. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mboli:2022:ITE

[MTM22] Julius Sechang Mboli, Dhavalkumar Thakker, and Jyoti L. Mishra. An Internet of Things-enabled decision support system for circular economy business model. Software—Practice and Experience, 52(3):772– 787, March 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Moore:2014:PDS

[MTPC14] Thomas Moore, Sameer Tilak, Phillip Papadouplous, and Luca Clementi. Programmatically defining the software footprint of sensor networks using the Android platform. Software—Practice and Experience, 44(10):1277–1286, October 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mzali:1983:LE

[MTRC83] Jalel Mzali, Jean-Jacques Thiel, Jeff Reeves, and Jim Conrad. Letters to the editor. Software—Practice and Experience, 13 (11):1095–1097, November 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Magnenat-Thalmann:1981:GPE

[MTT81] Nadia Magnenat-Thalmann and Daniel Thalmann. A graphical Pascal extension based on graphical types. Software—Practice and Experience, 11(1):53– 62, January 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Magnenat-Thalmann:1983:MTD

[MTT83] Nadia Magnenat-Thalmann and Daniel Thalmann. MIRA-3D: a three-dimensional graphical extension of Pascal. Software—Practice and Experience, 13(9):797– 808, September 1983. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Madhura:2023:ELB

[MUE23] Rajarethinam Madhura, Vaidyanathan Rhymend Uthariaraj, and Benjamin Lydia Elizabeth. An efficient list-based task scheduling algorithm for heterogeneous distributed computing environment. Software—Practice and Experience, 53(2):390–412, February 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mullins:1976:BRB

[Mul76] J. M. Mullins. Book review: Lang-Pak — An Interactive Language Design System, Lee E. Heindel and Jerry T. Roberto, Elsevier Publishing Company, Amsterdam, 1975. No. of pages: 184. Price: Dfl 23.00 ($9.75). Software—Practice and Experience, 6(3): 438–439, July 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Musstopf:1979:CPO

[Mus79] Guenter Musstopf. Catalogue on problem-oriented languages. Software— Practice and Experience, 9(10):879, October 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Musser:1997:ISS

[Mus97] David R. Musser. Introspective sorting and selection algorithms. Software—Practice and Experience, 27(8):983–993, August 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7328;   http://www3.

                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=7328&
                                        PLACEBO=IE.pdf.

Mustacoglu:2017:NMD

[Mus17] Ahmet F. Mustacoglu. A novel model-driven approach for seamless integration. Software—Practice and Experience, 47(5): 689–708, May 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Merrett:1986:FPR

[MV86] T. H. Merrett and Ted Van Rossum. A functional processor for the relational algebra on a microcomputer. Software—Practice and Experience, 16(11): 987–1002, November 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mazzeo:1995:PFC

[MV95] Antonino Mazzeo and Umberto Villano. Parallel 1D-FFT computation on constant-valence multicomputers. Software—Practice and Experience, 25(6):681– 704, June 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Min:2016:NMS

[MV16] Byungho Min and Vijay Varadharajan. A novel malware for subversion of self-protection in antivirus. Software—Practice and Experience, 46(3):361– 379, March 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marchini:2020:CFT

[MV20] Stefano Marchini and Sebastiano Vigna. Compact Fenwick trees for dynamic ranking and selection. Software—Practice and Experience, 50(7): 1184–1202, July 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Maiti:2022:ADI

[MVM^{+}22] Moinak Maiti, Darko Vuković, Amrit Mukherjee, Pavan D. Paikarao, and Janardan Krishna Yadav. Advanced data integration in banking, financial, and insurance software in the age of COVID-19. Software —Practice and Experience, 52(4):887–903, April 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Maenhaut:2019:ERM

[MVOD19] Pieter-Jan Maenhaut, Bruno Volckaert, Veerle Ongenae, and Filip De Turck. Efficient resource management in the cloud: From simulation to experimental validation using a low-cost Raspberry Pi testbed. Software—Practice and Experience, 49(3):449–477, March 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mili:2018:OBM

[MVS^{+}18] Hafedh Mili, Petko Valtchev, Laszlo Szathmary, Anis Boubaker, Abderrahmane Leshob, Yasmine Charif, and Louis Martin. Ontologybased model-driven development of a destination management portal: Experience and lessons learned. Software—Practice and Experience, 48(8): 1438–1460, August 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2581.

Magalhaes:2009:ERO

[MvSdL09] João Magalhães, Arndt von Staa, and Carlos José Pereira de Lucena. Evaluating the recoveryoriented approach through the systematic development of real complex applications. Software—Practice and Experience, 39(3): 315–330, March 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mostinckx:2009:MBR

[MVT^{+}09] Stijn Mostinckx, Tom Van Cutsem, Stijn Timbermont, Elisa Gonzalez Boix, Éric Tanter, and Wolfgang De Meuter. Mirrorbased reflection in AmbientTalk. Software—Practice and Experience, 39(7): 661–699, May ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Min:2014:ASN

[MVTH14] Byungho Min, Vijay Varadharajan, Udaya Tupakula, and Michael Hitchens. Antivirus security: naked during updates. Software— Practice and Experience, 44(10):1201–1222, October 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mannaert:2012:TES

[MVV12] Herwig Mannaert, Jan Verelst, and Kris Ven. Towards evolvable software architectures based on systems theoretic stability. Software—Practice and Experience, 42(1):89–116, January 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Madhavji:1981:DSD

[MW81] N. H. Madhavji and I. R.

Wilson. Dynamically structured data. Software—Practice and Experience, 11(12):1235–1260, December 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Montuelle:1982:LE

[MW82] Jean Montuelle and Ian Willers. Letter to the editor. Software—Practice and Experience, 12(9):883, September 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Moynihan:1991:DIH

[MW91] Vincent D. Moynihan and Peter J. L. Wallis. The design and implementation of a high-level language converter. Software—Practice and Experience, 21(4):391– 400, April 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mayer:1993:IAA

[MW93] Herbert G. Mayer and Michael Wolfe. Interprocedural alias analysis: Implementation and empirical results. Software—Practice and Experience, 23 (11):1201–1233, November 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Makady:2013:VPR

[MW13] Soha Makady and Robert J.

Walker. Validating pragmatic reuse tasks by leveraging existing test suites. Software—Practice and Experience, 43(9):1039– 1070, September 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Magavi:1995:DIH

[MWB95] Sunil Magavi, Johnny Wong, and Prakash Bodla. Design and implementation of heterogeneous distributed multimedia system using Mosaic GSQL. Software—Practice and Experience, 25(11):1223– 1241, November 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mili:1986:SMI

[MXYQ86] Ali Mili, Wang Xiao-Yang, and Yu Qing. Specification methodology: An integrated relational approach. Software—Practice and Experience, 16 (11):1003–1030, November 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Munakata:1987:MSP

[MY87] Toshinori Munakata and D. Yun Yeh. Multiple stack performance near saturation. Software—Practice and Experience, 17(6):399– 411, June 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mencnarowski:2000:ISE

[MZ00] Daniel Mencnarowski and Krzysztof Zieliński. Improving scalability of eventdriven distributed objects architectures. Software— Practice and Experience, 30(13):1509–1529, November 10, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
73001206/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=73001206&PLACEBO=IE.
pdf.

Marquez:2000:FPO

[MZB00] Alonso Marquez, John N. Zigman, and Stephen M. Blackburn. Fast portable orthogonally persistent Java^{T M}. Software—Practice and Experience, 30(4): 449–479, April 10, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www3.interscience.
wiley.com/cgi-bin/abstract/
71004143/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?

ID=71004143&PLACEBO=IE.
pdf.

Mateos:2008:SAG

[MZC08] Cristian Mateos, Alejandro Zunino, and Marcelo Campo. A survey on approaches to gridification. Software—Practice and Experience, 38(5):523– 556, April 25, 2008. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mateos:2010:MJN

[MZC10] Cristian Mateos, Alejandro Zunino, and Marcelo Campo. m-JGRIM: a novel middleware for Gridifying Java applications into mobile Grid services. Software —Practice and Experience, 40(4):331–362, April 10, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Neira-Ayuso:2010:CBK

[NAGL10] Pablo Neira-Ayuso, Rafael M.

Gasca, and Laurent Lefevre. Communicating between the kernel and user-space in Linux using Netlink sockets. Software—Practice and Experience, 40(9):797– 810, August 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Narayanan:1994:DSS

[Nar94] Ajith K. Narayanan. Design of a safe string library for C. Software—Practice and Experience, 24(6):565– 578, June 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Naz:2021:TBI

[NAU^{+}21] Ayesha Naz, Hafiz M. Asif, Tariq Umer, Shahid Ayub, and Fadi Al-Turjman. Trilaterationbased indoor localization engineering technique for visible light communication system. Software—Practice and Experience, 51(3):503– 516, March 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Navarro:2001:NGF

[Nav01] Gonzalo Navarro. NR-grep: a fast and flexible patternmatching tool. Software —Practice and Experience, 31(13):1265–1312, November 10, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
85512195/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=85512195&PLACEBO=IE.
pdf.

Nabavi:2019:APO

[NB19] Seyyed Yahya Nabavi and Omid Bushehrian. An adaptive plan-oriented and continuous software migration to cloud in dynamic enterprises. Software—Practice and Experience, 49(9):1365–1378, September 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nanda:1999:ACT

[NBOS99] Mangala Gowri Nanda, Purandar Bhaduri, Sundeep Oberoi, and Amitabha Sanyal. An application of compiler technology to the year 2000 problem. Software—Practice and Experience, 29(4):359–377, April 10, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55001840;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55001840&PLACEBO=IE.
pdf.

Nelson:1975:PPF

[NC75] James M. Nelson and Charles E. Cohn. Parallel processing in FORTRAN with floating-point hardware. Software—Practice and Experience, 5(1):65– 68, January/March 1975.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Navarro:2012:AMV

[NCFCFV12] Antonio Navarro, Jesús Cristóbal, Carmen Fernández-Chamizo, and Alfredo Fernández-Valmayor. Architecture of a multiplatform virtual campus. Software—Practice and Experience, 42(10):1229–1246, October 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nguyen-Duc:2024:WHI

[NDKL^{+}24] Anh Nguyen-Duc, Dron Khanna, Giang Huong Le, Des Greer, Xiaofeng Wang, Luciana Martinez Zaina, Gerardo Matturro, Jorge Melegati, Eduardo Guerra, Petri Kettunen, Sami Hyrynsalmi, Henry Edison, Afonso Sales, Rafael Chanin, Didzis Rutitis, Kai-Kristian Kemell, Abdullah Aldaeej, Tommi Mikkonen, Juan Garbajosa, and Pekka Abrahamsson. Work-from-home impacts on software project: a global study on software development practices and stakeholder perceptions. Software—Practice and Experience, 54(5):896– 926, May 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Neely:1975:EAF

[Nee75]  Peter M. Neely.  Edi-

torial: After FORTRAN, what? Software—Practice and Experience, 5(1):1–4, January 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Neely:1976:NPD

[Nee76]  Peter M. Neely.  The

new programming discipline. Software—Practice and Experience, 6(1):1–27, January/March 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Neely:1977:BRB

[Nee77a] P. M. Neely. Book review:

Logical construction of programs, J.-D. Warnier, H. E. Stenfert Kroese B.V., Leiden, 1974. No. of pages:

221. Price: 40.00 guilders ( £10.00). Software—Practice and Experience, 7(3): 429–431, June 1977. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Neely:1977:IIA

[Nee77b] Peter M. Neely. Implementation independent arithmetic: Speculation for discussion. Software—Practice and Experience, 7 (4):461–468, July/August 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Neely:1977:UIA

[Nee77c] Peter M. Neely. On the use of integer arithmetic to achieve confirmably correct computation. Software— Practice and Experience, 7 (2):159–163, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nentwich:2000:BBO

[NEFZ00] Christian Nentwich, Wolfgang Emmerich, Anthony Finkelstein, and Andrea Zisman. BOX: Browsing objects in XML. Software—Practice and Experience, 30(15):1661–1676, December 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
73501273/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=73501273&PLACEBO=IE.
pdf.

Nehmer:1979:ICP

[Neh79] J. Nehmer. The implementation of concurrency for a PL/I-like language. Software—Practice and Experience, 9(12):1043–1057, December 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nunes:2017:MCI

[NEP^{+}17] Luiz Henrique Nunes, Julio Cezar Estrella, Charith Perera, Stephan Reiff-Marganiec, and Alexandre Cláudio Botazzo Delbem. Multi-criteria IoT resource discovery: a comparative analysis. Software—Practice and Experience, 47 (10):1325–1341, October 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Newbury:1982:ITE

[New82] J. P. Newbury. Immediate turnround: an elusive goal. Software—Practice and Experience, 12(10): 897–906 (or 897–895??), October 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Newman:1986:PVL

[New86] I. A. Newman. Processing variable length abbreviations: Some observations. Software—Practice and Experience, 16(5):413– 423, May 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nunes:2014:HEF

[NGLL14] Camila Nunes, Alessandro Garcia, Carlos Lucena, and Jaejoon Lee. Heuristic expansion of feature mappings in evolving program families. Software—Practice and Experience, 44(11):1315–1349, November 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nidd:2003:CPC

[NH03] M. Nidd and D. Husemann. CA-PK: conditional access for broadcast networks. Software—Practice and Experience, 33(5):481– 496, April 25, 2003. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Naghdipour:2023:SDP

[NHB23] Ameneh Naghdipour, Seyed Mohammad Hossein Hasheminejad, and Roghayeh Leila Barmaki. Software design pattern selection approaches: a systematic literature review. Software—Practice and Experience, 53(4):1091–1122, April 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Negus:1981:DSQ

[NHP81] B. Negus, M. J. Hunt, and J. A. Prentice. DIALOG: a scheme for the quick and effective production of interactive applications software. Software—Practice and Experience, 11(3):205– 224, March 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nishikimi:2008:WFD

[NHTT08] M. Nishikimi, K. Handa, N. Takahashi, and S. Tomura. When fonts do not know everything: an extensible system for multilingual text rendering. Software—Practice and Experience, 38(1):1–31, January ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nicholls:1972:BRB

[Nic72] J. E. Nicholls. Book review: PL/1 Programming in technological applications. Gabriel F. Groner, Wiley-Interscience, New York, 1971. No. of pages: 182+11. Price: Cloth £4.75, Paper £3.50. Software—Practice and Experience, 2(3):305–306, July 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nicholson:1998:BRG

[Nic98] Mark Nicholson. Book review: Genetic Algorithms and grouping problems.

Software—Practice and Experience, 28(10):1137– 1138, August 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=10007352;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=10007352&PLACEBO=IE.
                                        pdf.

Nicart:2008:TSV

[Nic08]  Florent Nicart.  To-

wards scalable virtuality in C++. Software—Practice and Experience, 38(14): 1451–1473, November 25, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nielsen:1979:LEM

[Nie79] Gunnar Nielsen. Letter to the Editor: Macro assembly languages. Software —Practice and Experience, 9(2):172, February 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nakamura:1985:NML

[NIEN85] Tomoo Nakamura, Katsuo Ikeda, Yoshihiko Ebihara, and Masafumi Nishikawa. Network management in a local computer network. Software—Practice and Experience, 15(4):343– 358, April 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nilsen:1988:GCS

[Nil88] Kevin D. Nilsen. Garbage collection of strings and linked data structures in real time. Software—Practice and Experience, 18(7): 613–640, July 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nilsen:1990:HLG

[Nil90] Kelvin Nilsen. High-level goal-directed concurrent processing in Icon. Software—Practice and Experience, 20(12):1273–1290, December 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nadella:2011:MFH

[NJ11] Suneetha Nadella and Dharanipragada Janakiram. Message filters for hardening the Linux kernel. Software—Practice and Experience, 41(1):51–62, January 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nagarajan:2014:ESD

[NJG14] Vijay Nagarajan, Dennis Jeffrey, and Rajiv Gupta. Erratum: A system for debugging via online tracing and dynamic slicing. Software—Practice and Experience, 44 (9):1155, September 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [NJGG12b].

Nagarajan:2012:ESD

[NJGG12a] Vijay Nagarajan, Dennis Jeffrey, Rajiv Gupta, and Neelam Gupta. Erratum: A system for debugging via online tracing and dynamic slicing. Software—Practice and Experience, 42 (11):1431, November 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [NJGG12b].

Nagarajan:2012:SDO

[NJGG12b] Vijay Nagarajan, Dennis Jeffrey, Rajiv Gupta, and Neelam Gupta. A system for debugging via online tracing and dynamic slicing. Software—Practice and Experience, 42(8):995– 1014, August 2012. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See errata [NJGG12a, NJG14].

Nieminen:2007:EIA

[NK07] Janne Nieminen and Pekka Kilpeläinen. Efficient implementation of Aho– Corasick pattern matching automata using Unicode. Software—Practice and Experience, 37(6):669– 690, May 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ngassam:2006:PHF

[NKW06] E. Ketcha Ngassam, Derrick G. Kourie, and Bruce W. Watson. Performance of hardcoded finite automata. Software—Practice and Experience, 36(5): 525–538, April 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Neely:1975:LE

[NL75] Peter M. Neely and John Larmouth. Letter to the editor. Software—Practice and Experience, 5(2):213– 215, April 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Newman:1976:DCU

[NL76] N. Newman and T. Lang.

Documentation for computer users. Software —Practice and Experience, 6(3):321–326, July/ September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Noble:2001:EAO

[NL01] James Noble and Doug Lea. Editorial: Aliasing in object-oriented systems. Software—Practice and Experience, 31(6): 505, May 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
78003103/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=78003103&PLACEBO=IE.
pdf.

Nizipli:2024:PMU

[NL24] Yagiz Nizipli and Daniel Lemire. Parsing millions of URLs per second. Software —Practice and Experience, 54(5):744–758, May 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nesmachnow:2015:ETA

[NLA15] Sergio Nesmachnow, Francisco Luna, and Enrique Alba. An empirical time analysis of evolutionary algorithms as C programs. Software—Practice and Experience, 45(1):111– 142, January 2015. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nehmer:1977:LE

[NM77] Jürgen Nehmer and Anthony Mark. Letter to the editor. Software—Practice and Experience, 7(2):297– 298, March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ng:1978:IGC

[NM78] Nam Ng and T. A. Marsland. Introducing graphics capabilities to several highlevel languages. Software —Practice and Experience, 8(5):629–639, September/ October 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nikora:2006:BHQ

[NM06] Allen P. Nikora and John C. Munson. Building high-quality software fault predictors. Software —Practice and Experience, 36(9):949–969, July 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Narayanan:2019:HSS

[NM19] Munisamy Eswara Narayanan and Balasundaram Muthukumar. A highly secured and streamlined cloud collaborative editing scheme along with an efficient user revocation in cloud computing. Software—Practice and Experience, 49 (12):1728–1747, December 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nelson:2011:SEU

[NMG11] Adam Nelson, Tim Menzies, and Gregory Gay. Sharing experiments using open-source software. Software—Practice and Experience, 41(3):283–305, March 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nakano:2024:DSS

[NMH^{+}24] Bunta Nakano, Takaki Matsumoto, Naoto Hoshikawa, Tomoyoshi Ito, and Atsushi Shiraki. Development of a second-screen system for sharing virtual reality information. Software— Practice and Experience, 54(5):796–812, May 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Narasimhan:2002:ECB

[NMMS02] P. Narasimhan, L. E. Moser, and P. M. Melliar-Smith. Eternal — a component-based framework for transparent faulttolerant CORBA. Software—Practice and Experience, 32(8):771–788, July 10, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
94519522/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?

ID=94519522&PLACEBO=IE.
pdf.

Nevill-Manning:1998:ETP

[NMRW98] Craig G. Nevill-Manning, Todd Reed, and Ian H. Witten. Extracting text from PostScript. Software—Practice and Experience, 28(5):481–491, May 25, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1792;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1792&
PLACEBO=IE.pdf.

Niyogi:2024:FSV

[NN24] Rajdeep Niyogi and Amar Nath. Formal specification and verification of a team formation protocol using TLA^{+}. Software—Practice and Experience, 54(6):961– 984, June 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Najafizadegan:2021:AMS

[NNK21] Negin Najafizadegan, Eslam Nazemi, and Vahid Khajehvand. An autonomous model for selfoptimizing virtual machine selection by learning automata in cloud environment. Software—Practice and Experience, 51(6):

1352–1386, June 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Naujokat:2014:SFM

[NNL^{+}14] Stefan Naujokat, Johannes Neubauer, Anna-Lena Lamprecht, Bernhard Steffen, Sven Jörges, and Tiziana Margaria. Simplicity-first model-based plug-in development. Software—Practice and Experience, 44(3): 277–297, March 2014. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Neelofar:2017:ISB

[NNLR17] Neelofar Neelofar, Lee Naish, Jason Lee, and Kotagiri Ramamohanarao. Improving spectral-based fault localization using static analysis. Software—Practice and Experience, 47(11):1633–1655, November 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Neelofar:2018:SBF

[NNR18] N. Neelofar, Lee Naish, and Kotagiri Ramamohanarao. Spectral-based fault localization using hyperbolic function. Software—Practice and Experience, 48 (3):641–664, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2527.

Nica:2013:UMT

[NNW13] Mihai Nica, Simona Nica, and Franz Wotawa. On the use of mutations and testing for debugging. Software—Practice and Experience, 43(9):1121–1142, September 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Noot:1983:STF

[Noo83]  Han Noot.  Structured

text formatting. Software —Practice and Experience, 13(1):79–94, January 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Norvig:1991:CWE

[Nør91] Peter Nørvig. Correcting a widespread error in unification algorithms. Software —Practice and Experience, 21(2):231–233, February 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Notkin:1990:PSS

[Not90] David Notkin. Proxies: a software structure for accommodating heterogeneity. Software—Practice and Experience, 20(4):357– 364, April 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Norris:1998:DIR

[NP98] Cindy Norris and Lori L. Pollock. The design and implementation of RAP: a PDG-based register allocator. Software—Practice and Experience, 28(4):401– 424, April 10, 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1787;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1787&
PLACEBO=IE.pdf.

Nejedly:2018:CSL

[NPHJ18] Petr Nejedly, Filip Plesinger, Josef Halamek, and Pavel Jurak. CudaFilters: a SignalPlant library for GPUaccelerated FFT and FIR filtering. Software—Practice and Experience, 48(1): 3–9, January 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Narayana:1979:SAC

[NPJ79] K. T. Narayana, V. R. Prasad, and M. Joseph. Some aspects of concurrent programming in CCN-PASCAL. Software—Practice and Experience, 9(9): 749–770, September 1979.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Newey:1972:AMM

[NPW72] M. C. Newey, P. C. Poole, and W. M. Waite. Abstract machine modelling to produce portable software — a review and evaluation. Software—Practice and Experience, 2(2): 107–136, April/June 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nystrom:2004:EII

[NR04] Mattias Nyström and Tore Risch. Engineering information integration using object-oriented mediator technology. Software—Practice and Experience, 34(10):949–975, August 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Noureddine:2013:SPR

[NRS13] Adel Noureddine, Romain Rouvoy, and Lionel Seinturier. Survey paper: A review of middleware approaches for energy management in distributed environments. Software— Practice and Experience, 43(9):1071–1100, September 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nasar:2021:DDA

[NRUP21] Nasreen Nasar, Sumati Ray, Saiyed Umer, and Hari Mohan Pandey. Design and data analytics of electronic human resource management activities through Internet of Things in an organization. Software—Practice and Experience, 51 (12):2411–2427, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nolan:1974:WCT

[NS74] Lawrence E. Nolan and Jon C. Strauss. Workload characterization for timesharing system selection. Software—Practice and Experience, 4(1):25– 39, January/March 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Natarajan:1979:LII

[NS79] N. Natarajan and Mukul K. Sinha. Language issues in the implementation of a kernel. Software—Practice and Experience, 9(9): 771–778, September 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nishanov:2001:MCC

[NS01a] Gor V. Nishanov and Sibylle Schupp. A mostlycopying collector component for class templates. Software—Practice and Experience, 31(5):445–470, April 25, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
77502756/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=77502756&PLACEBO=IE.
pdf.

Niv:2001:TAS

[NS01b] Nitzan Niv and Assaf Schuster. Transparent adaptation of sharing granularity in MultiView-based DSM systems. Software—Practice and Experience, 31(15):1439–1459, December 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
85515684/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=85515684&PLACEBO=IE.
pdf.

Nagpal:2008:PIS

[NS08] Rahul Nagpal and Y. N.

Srikant. Pragmatic integrated scheduling for clustered VLIW architectures. Software—Practice and Experience, 38(3):227– 257, March ??, 2008. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Neri:2023:EAF

[NS23] Cassio Neri and Lorenz Schneider. Euclidean affine functions and their application to calendar algorithms. Software—Practice and Experience, 53(4):937– 970, April 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nadrchal:1983:IAS

[NSKK83] Jaroslav Nadrchal, Ivan Sklenář, Václav Kříž, and Jaroslav Král. An implementation of Algol 68 for a small computer. Software —Practice and Experience, 13(7):597–660, July 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Norris-Sherborn:1986:PAD

[NSM86] A. Norris-Sherborn and W. J. Milne. A practical approach to data modelling in spatial applications. Software—Practice and Experience, 16(10): 893–913, October 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Naish:2016:ALE

[NSM16] Lee Naish, Peter Schachte, and Aleck M. MacNally. Adtpp: lightweight efficient safe polymorphic algebraic data types for C. Software—Practice and Experience, 46(12):1685– 1703, December 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

N:2022:ESS

[NSM^{+}22] Balaji N, Lakshmi S, Anand M, Anbarasan M, and Mathiyalagan P. An efficient scheme for secure feature location using data fusion and data mining in Internet of Things environment. Software—Practice and Experience, 52(3):642– 657, March 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nutt:1977:GCH

[NSW77] Gary J. Nutt, Waldean A.

Schulz, and Karl H. Williamson. Generating code for a hypothetical computer using a production assembler. Software— Practice and Experience, 7 (1):147–148, January 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nackman:1984:HEH

[NT84] Lee R. Nackman and Russell H. Taylor. A hierarchical exception handler binding mechanism. Software —Practice and Experience, 14(10):999–1003, October 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Navarro:2005:LBM

[NT05] Gonzalo Navarro and Jorma Tarhio. LZgrep: a Boyer– Moore string matching tool for Ziv–Lempel compressed text. Software—Practice and Experience, 35(12): 1107–1130, October 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        http://www.dcc.uchile.
                                        cl/~gnavarro/pubcode/.

Niemi:2020:BWP

[NT20] Arto Niemi and Jukka Teuhola. Burrows–Wheeler post-transformation with effective clustering and interpolative coding. Software—Practice and Experience, 50(9):1858–1874, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nitu:2017:SEV

[NTF^{+}17] Vlad Nitu, Boris Teabe, Leon Fopa, Alain Tchana, and Daniel Hagimont.

StopGap: elastic VMs to enhance server consolidation. Software—Practice and Experience, 47 (11):1501–1519, November 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nguyen:2024:ITM

[NUB24] Hoa T. Nguyen, Muhammad Usman, and Rajkumar Buyya. iQuantum: a toolkit for modeling and simulation of quantum computing environments. Software—Practice and Experience, 54(6): 1141–1171, June 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nutt:1976:CSR

[Nut76]  Gary J. Nutt.  Com-

puter system resource requirements of novice programming students. Software—Practice and Experience, 6(1):43–50, January/ March 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Neal:1978:EPC

[NW78] David Neal and Virgil Wallentine. Experiences with the portability of concurrent PASCAL. Software —Practice and Experience, 8(3):341–353, May/June 1978. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Nordstrom:1984:DIP

[NW84] Bengt Nordström and Åke Wikström. The design of an interactive program development system for Pascal. Software—Practice and Experience, 14(2):177– 190, February 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Newey:1985:RIS

[NW85] M. C. Newey and W. M. Waite. The robust implementation of sequencecontrolled iteration. Software—Practice and Experience, 15(7):655–668, July 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nedjah:1999:EAD

[NWE99] Nadia Nedjah, Colin D. Walter, and Stephen E. Eldridge. Efficient automatadriven pattern-matching for equational programs. Software—Practice and Experience, 29(9):793–813, July 25, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=62501864;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?

ID=62501864&PLACEBO=IE.
pdf.

Ng:1978:FPM

[NY78] P. H. Ng and G. Young.

A 1900 FORTRAN post mortem dump system. Software—Practice and Experience, 8(4):421–427, July/August 1978. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Naeen:2020:AMB

[NZH20] Hossein Monshizadeh Naeen, Esmaeil Zeinali, and Abolfazl Toroghi Haghighat. Adaptive Markov-based approach for dynamic virtual machine consolidation in cloud data centers with quality-of-service constraints. Software—Practice and Experience, 50(2): 161–183, February 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nirumand:2019:VFV

[NZL19] Atefeh Nirumand, Bahman Zamani, and Behrouz Tork Ladani. VAnDroid: a framework for vulnerability analysis of Android applications using a modeldriven reverse engineering technique. Software—Practice and Experience, 49(1): 70–99, January 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nourisa:2022:COS

[NZPWR22] Jalil Nourisa, Berit Zeller-Plumhoff, and Regine Willumeit-Römer. CppyABM: an open-source agent-based modeling library to integrate C++ and Python. Software—Practice and Experience, 52(6):1337– 1351, June 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nirumand:2023:MBF

[NZTL^{+}23] Atefeh Nirumand, Bahman Zamani, Behrouz Tork-Ladani, Jacques Klein, and Tegawendé F. Bissyandé. A model-based framework for inter-app vulnerability analysis of Android applications. Software—Practice and Experience, 53(4): 895–936, April 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Oono:2003:FCE

[OAF^{+}03] Masaki Oono, El-Sayed Atlam, Masao Fuketa, Kazuhiro Morita, and Jun ichi Aoe. A fast and compact elimination method of empty elements from a double-array structure. Software—Practice and Experience, 33(13): 1229–1249, November 10, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Owusu:2019:FDB

[OAZ19] Ebenezer Owusu, Jamal-Deen Abdulai, and Yongzhao Zhan. Face detection based on multilayer feed-forward neural network and Haar features. Software—Practice and Experience, 49 (1):120–129, January 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Obermaisser:2011:ECS

[Obe11] Roman Obermaisser. Editorials: Component and service-oriented distributed embedded real-time systems (extended papers from ISORC 2010). Software—Practice and Experience, 41(12):1385–1386, November/December 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Olsson:1991:DAE

[OCH91] Ronald A. Olsson, Richard H. Crawford, and W. Wilson Ho. A dataflow approach to event-based debugging. Software— Practice and Experience, 21(2):209–230, February 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ottenstein:1992:ECF

[OE92] Kar L. J. Ottenstein and Steven J. Ellcey. Experience compiling Fortran to program dependence graphs. Software—Practice and Experience, 22(1):41– 62, January 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

OGorman:2005:MQO

[OEA05] K. O’Gorman, A. El Abbadi, and D. Agrawal. Multiple query optimization in middleware using query teamwork. Software —Practice and Experience, 35(4):361–391, April 10, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Oestreicher:1971:DIS

[Oes71] M. D. Oestreicher. The design of the internal structure of the ICL GEORGE 3 operating system. Software —Practice and Experience, 1(2):189–200, April/June 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Osterweil:1976:DVE

[OF76] Leon J. Osterweil and Lloyd D. Fosdick. DAVE: a validation error detection and documentation system for Fortran programs. Software—Practice and Experience, 6(4):473–486, October/December 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Omoronyia:2010:RAD

[OFRW10] Inah Omoronyia, John Ferguson, Marc Roper, and Murray Wood. A review of awareness in distributed collaborative software engineering. Software—Practice and Experience, 40(12):1107–1133, November 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Osorio:2016:FAC

[OG16] Roberto R. Osorio and Patricia González. A fast algorithm for constructing nearly optimal prefix codes. Software—Practice and Experience, 46(10): 1299–1316, October 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Odoom:2022:CFP

[OHD22] Justice Odoom, Xiaofang Huang, and Samuel Akwasi Danso. COVID-19 and future pandemics: a blockchain-based privacyaware secure borderless travel solution from electronic health records. Software—Practice and Experience, 52(10):2263–2287, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Offutt:1999:DDR

[OJP99] A. Jefferson Offutt, Zhenyi Jin, and Jie Pan. The dynamic domain reduction procedure for test data generation. Software—Practice and Experience, 29(2):167–193, February 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=55000306;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=55000306&PLACEBO=IE.
                                        pdf.

Ogasawara:2004:OPO

[OKN04] Takeshi Ogasawara, Hideaki Komatsu, and Toshio Nakatani. Optimizing precision overhead for x86 processors. Software—Practice and Experience, 34(9):875– 893, July 25, 2004. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Oliver:1983:NAC

[Oli83] J. Oliver. The necessity for accurate compiler-provided routines when evaluating special functions. Software—Practice and Experience, 13(12):1169–1174, December 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

ONeal:1989:SFA

[OLS89] Michael Ralph O’Neal, William Lively, and Sallie Sheppard. Software function allocation methodology. Software—Practice and Experience, 19(8):775– 786, August 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Olsson:1990:USD

[Ols90] Ronald A. Olsson. Using SR for discrete event simulation: a study in concurrent programming. Software—Practice and Experience, 20(12):1187–1208, December 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Owolabi:1988:FAS

[OM88] O. Owolabi and D. R. Mc-Gregor. Fast approximate string matching. Software —Practice and Experience, 18(4):387–393, April 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Olsson:1996:EUC

[OM96] Ronald A. Olsson and Carole M. McNamee. Experience using the C preprocessor to implement CCR, monitor, and CSP preprocessors for SR. Software—Practice and Experience, 26(2):125–134, February 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61004288.

Oliver:2016:ERD

[OM16] Helen Oliver and Patrick McSharry. Experience reports: A design protopattern for continuously evaluated forecasting in IBM(R) InfoSphere(R) Streams. Software—Practice and Experience, 46(5): 645–684, May 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Okuno:1996:TFF

[OMA96] Hirotomo Okuno, Hideki Matsumoto, and Hironori Asai. TableSpec: Free format specification table and source code generation. Software—Practice and Experience, 26(2):213– 235, February 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61004286.

Ortin:2014:SDL

[OMGDG14] Francisco Ortin, Sheila Mendez, Vicente García-Díaz, and Miguel Garcia. On the suitability of dynamic languages for hotreprogramming a robotics framework: a Python case study. Software—Practice and Experience, 44(1):77– 104, January 2014. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ortac:2015:AML

[OMM15] Alper Ortac, Martin Monperrus, and Mira Mezini. Abmash: mashing up legacy Web applications by automated imitation of human actions. Software— Practice and Experience, 45(5):581–612, May 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

ONeill:1988:GIS

[O’N88] M. A. O’Neill. GPROC:

An integrated system for the processing of numerical scientific data. Software —Practice and Experience, 18(9):841–857, September 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Onibere:1985:WPF

[Oni85]  E. A. Onibere.  Writ-

ing portable Fortran programs for microcomputers. Software—Practice and Experience, 15(4):321– 326, April 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Onodera:1993:GCC

[Ono93a] Tamiya Onodera. A generational and conservative copying collector for hybrid object-oriented languages. Software—Practice and Experience, 23(10): 1077–1093, October 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Onodera:1993:RCT

[Ono93b] Tamiya Onodera. Reducing compilation time by a compilation server. Software—Practice and Experience, 23(5):477–485, May 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

OShea:2019:VTA

[OOG19] Donna O’Shea, Francisco Ortin, and Kevin Geary. A virtualized test automation framework: a DellEMC case study of test automation practice. Software— Practice and Experience, 49(2):329–337, February 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Offutt:1996:EED

[OPTZ96] A. Jefferson Offutt, Jie Pan, Kanupriya Tewary, and Tong Zhang. An experimental evaluation of data flow and mutation testing. Software—Practice and Experience, 26(2):165– 176, February 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61004289.

Orgass:1981:FIE

[Org81] Richard J. Orgass. Files in an interactive environment. Software—Practice and Experience, 11 (12):1265–1271, December 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ormicki:1977:RTB

[Orm77] Andrzej Ormicki. Realtime BASIC for laboratory use. Software—Practice and Experience, 7(4):435– 443, July/August 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Oldehoeft:1981:IMP

[ORT81] R. R. Oldehoeft, W. D.

Ralph, and M. H. Tindall. An interactive manager for Pascal software. Software—Practice and Experience, 11(8):867–873, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ozcan:1996:ISI

[ÖS96] Mehmet Bülent Özcan and Jawed Siddiqi. Interchanging specifications and implementations in evolutionary prototyping. Software—Practice and Experience, 26(9):999–1023, September 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16833.

Ogle:1992:PEI

[OSW92] David M. Ogle, Neil G. Sullivan, and E. Hollins Williams, Jr. Practical experience with OS/2 installable file systems. Software —Practice and Experience, 22(7):537–551, July 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ono:2002:SSA

[OT02] Kouichi Ono and Hideki Tai. A security scheme for Aglets. Software—Practice and Experience, 32(6): 497–514, May 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.

com/cgi-bin/abstract/
93513617/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=93513617&PLACEBO=IE.
pdf.

Ottmann:1982:DSV

[Ott82] Josef S. Ottmann. DASD sharing in DOS/VSE. Software—Practice and Experience, 12(9):835–844 (or 835–842??), September 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Oliveira:1983:AMI

[OW83] J. N. Oliveira and I. R.

Wilson. An analysis of microcomputer implementation of Pascal. Software —Practice and Experience, 13(4):373–384, April 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Olsson:1989:STA

[OW89] Ronald A. Olsson and Gregory R. Whitehead. A simple technique for automatic recompilation in modular programming languages. Software—Practice and Experience, 19(8):757– 773, August 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Olsson:2016:ERR

[OW16] Ronald A. Olsson and Todd Williamson. Experience reports: RJ: a Java package providing JRlike concurrent programming. Software—Practice and Experience, 46(5):685– 708, May 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Oh:2010:AML

[OY10] Hakjoo Oh and Kwangkeun Yi. An algorithmic mitigation of large spurious interprocedural cycles in static analysis. Software—Practice and Experience, 40(8): 585–603, July 2010. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ozcan:1998:UEF

[Özc98] Mehmet Bülent Özcan. Use of executable formal specifications in user validation. Software—Practice and Experience, 28(13):1359– 1385, November 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1768;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1768&
PLACEBO=IE.pdf.

Ozkaya:2018:AAL

[Ozk18] Mert Ozkaya. The analysis of architectural languages for the needs of practitioners. Software— Practice and Experience, 48(5):985–1018, May 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2561.

Purtilo:1991:MRI

[PA91] James M. Purtilo and Joanne M. Atlee. Module reuse by interface adaptation. Software—Practice and Experience, 21(6):539– 556, June 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Printezis:2001:EOP

[PA01] T. Printezis and M. P.

Atkinson. An efficient object promotion algorithm for persistent object systems. Software—Practice and Experience, 31(10): 941–981, August 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
82002141/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=82002141&PLACEBO=IE.
pdf.

Paek:2007:EEC

[PACK07] Yunheung Paek, Minwook Ahn, Doosan Cho, and Taehwan Kim. Efficient embedded code generation with multiple load/ store instructions. Software—Practice and Experience, 37(11):1133–1159, September 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pagan:1979:HSI

[Pag79] Frank G. Pagan. A highlystructured interpreter for a SNOBOL4 subset. Software—Practice and Experience, 9(4):281–312, April 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pagan:1984:TCP

[Pag84]  Frank G. Pagan.  To-

ward complete programming language descriptions that are both formal and understandable. Software—Practice and Experience, 14(3):199–206, March 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pagan:1988:CIC

[Pag88] Frank G. Pagan. Converting interpreters into compilers. Software—Practice and Experience, 18(6):509– 527, June 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Palme:1974:LSS

[Pal74] Jacob Palme. List structures in SIMULA and PL/1 — a comparison. Software—Practice and Experience, 4(4):379–388, October/December 1974. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Palme:1976:ESS

[Pal76]  Jacob Palme.  Experi-

ence from the standardization of the SIMULA programming language. Software—Practice and Experience, 6(3):405–409, July/ September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Palme:1978:HFH

[Pal78a] Jacob Palme. How I fought with hardware and software and succeeded. Software—Practice and Experience, 8(1):77–83, January/ February 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Palme:1978:PMD

[Pal78b] Jacob Palme. Performance measurements in the development of a Simula compiler. Software—Practice and Experience, 8(2):234– 235, March 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Palme:1979:HCI

[Pal79] Jacob Palme. A humancomputer interface for noncomputer specialists. Software—Practice and Experience, 9(9):741–747, September 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Palme:1980:VIA

[Pal80] Jacob Palme. A viewdata interface for arbitrary computer programs. Software—Practice and Experience, 10(12):987–991, December 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Palme:1982:USP

[Pal82] Jacob Palme. Uses of the SIMULA process concept. Software—Practice and Experience, 12(2):153– 161, February 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Palmer:1986:FPD

[Pal86] J. A. B. Palmer. A Fortran procedure for drawing some space-filling curves. Software—Practice and Experience, 16(6):559–574, June 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pazos-Arias:2006:AFP

[PALNGD^{+}06] José J. Pazos-Arias, Martín López-Nores, Jorge García-Duque, Alberto Gil-Solla, Manuel Ramos-Cabrer, Yolanda Blanco-Fernández, Rebeca P. Díaz-Redondo, and Ana Fernández-Vilas. ATLAS: a framework to provide multiuser and distributed t-learning services over MHP. Software— Practice and Experience, 36(8):845–869, July 10, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pankhurst:1972:SSL

[Pan72] R. J. Pankhurst. SOLO; system for on-line optimization. Software—Practice and Experience, 2(3): 245–258, July/September 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Papakonstantinou:1979:PMR

[Pap79] G. Papakonstantinou. A poor man’s realization of attribute grammars. Software—Practice and Experience, 9(9):719–728, September 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Parsons:1975:HLJ

[Par75] I. T. Parsons. A high-level job control language. Software—Practice and Experience, 5(1):69–82, January/ March 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Parker:1978:MDM

[Par78] R. A. Parker. More on detecting misspelt identifiers in Fortran. Software— Practice and Experience, 8 (1):109, January 1978. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Parsons:1979:SSI

[Par79] I. T. Parsons. A support system for interactive languages. Software—Practice and Experience, 9(1):73– 86, January 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Parker:1985:GCI

[Par85a] J. R. Parker. A general character to integer conversion method. Software—Practice and Experience, 15(8):761–766, August 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Partridge:1985:SIT

[Par85b] Derek Partridge. Specifications and an implementation of the type-ambiguity problem in Pascal. Software—Practice and Experience, 15(12):1141–1158, December 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pashtan:1987:PII

[Pas87] Ariel Pashtan. A Prolog implementation of an instruction-level processor simulator. Software—Practice and Experience, 17(5): 309–318, May 1987. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Patterson:1983:LE

[Pat83] E. G. Patterson. Letter to the editor. Software —Practice and Experience, 13(2):197, February 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Paton:1994:DES

[Pat94]  Keith Paton.  Descrip-

tion of essential system behaviour using messagestate tables. Software— Practice and Experience, 24(4):405–420, April 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Purdom:1987:TMS

[PB87] Paul Walton Purdom, Jr. and Cynthia A. Brown. Tree matching and simplification. Software—Practice and Experience, 17(2):105– 115, February 1987. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Probets:2003:SOF

[PB03] S. G. Probets and D. F. Brailsford. Substituting outline fonts for bitmap fonts in archived PDF files. Software—Practice and Experience, 33(9):885– 899, July 25, 2003. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://brain.mhri.edu.
au/text/texpdf/texpdf-
ffont.html;   http://
www.cs.nott.ac.uk/~dfb/
Publications/Download/
2003/Probets03.pdf;
http://www.eprg.org/research/

.

Perez-Berenguer:2018:SBA

[PBGM18] Daniel Pérez-Berenguer and Jesús García-Molina. A standard-based architecture to support learning interoperability: a practical experience in gamification. Software—Practice and Experience, 48 (6):1238–1268, June 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2572.

Phillips:1978:TCL

[PBW78] J. B. Phillips, M. F.

Burke, and G. S. Wilson. Threaded code for laboratory computers. Software —Practice and Experience, 8(3):257–263, May/June 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Parrish:1996:ICI

[PCBE96] A. Parrish, D. Cordes, R. Borie, and S. Edara. Illustrating client and implementation readability tradeoffs in Ada and C++. Software—Practice and Experience, 26(7):799– 814, July 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16823.

Pinto:2018:WPC

[PCBR18] Gustavo Pinto, Fernando Castor, Rodrigo Bonifacio, and Marcel Rebouças. Work practices and challenges in continuous integration: a survey with Travis CI users. Software—Practice and Experience, 48(12):2223–2236, December 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Plebani:2012:MEO

[PCC^{+}12] Pierluigi Plebani, Cinzia Cappiello, Marco Comuzzi, Barbara Pernici, and Sandeep Yadav. Micro-MAIS: executing and orchestrating Web services on constrained mobile devices. Software—Practice and Experience, 42 (9):1075–1094, September 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Perez-Castillo:2012:CSB

[PCdGPP12] Ricardo Pérez-Castillo, Ignacio García-Rodríguez de Guzmán, Mario Piattini, and Ángeles S. Places. A case study on business process recovery using an e-government system. Software—Practice and Experience, 42(2):159– 189, February 2012. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Plank:1999:MEO

[PCL^{+}99] James S. Plank, Yuqun Chen, Kai Li, Micah Beck, and Gerry Kingsley. Memory exclusion: optimizing the performance of checkpointing systems. Software—Practice and Experience, 29(2):125–142, February 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55000305;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55000305&PLACEBO=IE.
pdf.

Perrott:1983:PLA

[PCM83] R. H. Perrott, D. Crookes, and P. Milligan. The programming language AC-TUS. Software—Practice and Experience, 13(4):305– 322, April 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Palopoli:2009:AAQ

[PCML09] L. Palopoli, T. Cucinotta, L. Marzario, and G. Lipari. AQuoSA — adaptive quality of service architecture. Software—Practice and Experience, 39(1): 1–31, January ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pyster:1978:ECC

[PD78] Arthur Pyster and Amitava Dutta. Error-checking compilers and portability. Software—Practice and Experience, 8(1):99–108, January/February 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Perrott:1981:EFP

[PD81] R. H. Perrott and P. S. Dhillon. An experiment with Fortran and Pascal. Software—Practice and Experience, 11(5):491– 496, May 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pohle:2000:FEU

[PD00] Hans-Peter Pohle and Bernd Drescher. A flexible and easy to use molecular biology workbench efficiently developed in Tcl/Tk. Software—Practice and Experience, 30(12):1433–1445, October 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        73000929/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=73000929&PLACEBO=IE.
                                        pdf.

Plank:2005:NCT

[PD05] James S. Plank and Ying Ding. Note: Correction to the 1997 tutorial on Reed–Solomon coding. Software—Practice and Experience, 35(2):189– 194, February 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [Pla97].

Philippe:2010:SAS

[PDBG10] Jeremy Philippe, Noel De Palma, Fabienne Boyer, and Olivier Gruber. Selfadaptation of service level in distributed systems. Software—Practice and Experience, 40(3):259–283, March ??, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Parrish:1998:IPD

[PDC^{+}98] Allen Parrish, Brandon Dixon, David Cordes, Susan Vrbsky, and John Lusth. Implementing persistent data structures using C++. Software— Practice and Experience, 28(15):1559–1579, December 25, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=10050310;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=10050310&PLACEBO=IE.
pdf.

Piraghaj:2017:CEM

[PDCB17] Sareh Fotuhi Piraghaj, Amir Vahid Dastjerdi, Rodrigo N. Calheiros, and Rajkumar Buyya. Container-CloudSim: an environment for modeling and simulation of containers in cloud data centers. Software— Practice and Experience, 47(4):505–521, April 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Portillo-Dominguez:2016:ECP

[PDPM^{+}16] A. Omar Portillo-Dominguez, Philip Perry, Damien Magoni, Miao Wang, and John Murphy. Extended conference paper: TRINI: an adaptive load balancing strategy based on garbage collection for clustered Java systems. Software—Practice and Experience, 46(12):1705–1733, December 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Portillo-Dominguez:2017:PAF

[PDPMM17] A. Omar Portillo-Dominguez, Philip Perry, Damien Magoni, and John Murphy. PHOEBE: an automation framework for the effective usage of diagnosis tools in the performance testing of clustered systems. Software—Practice and Experience, 47(11):1837–1874, November 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pellegrini:2022:DIF

[PDPQ22] Alessandro Pellegrini, Pierangelo Di Sanzo, Andrea Piccione, and Francesco Quaglia. Design and implementation of a fully transparent partial abort support for software transactional memory. Software—Practice and Experience, 52 (11):2456–2475, November 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Perez-Diaz:2013:WNF

[PDROFRM13] N. Pérez-Díaz, D. Ruano-Ordas, F. Fdez-Riverola, and J. R. Méndez. Wirebrush4SPAM: a novel framework for improving efficiency on spam filtering services. Software—Practice and Experience, 43 (11):1299–1318, November 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Prutchi:2022:HAF

[PdSCJM22] Eduardo S. Prutchi, Heleno de S. Campos Junior, and Leonardo G. P. Murta. How the adoption of feature toggles correlates with branch merges and defects in open-source projects? Software—Practice and Experience, 52(2): 506–536, February 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pedersen:1986:PAH

[Ped86] T. Pedersen. Process administration in a high level language. Software—Practice and Experience, 16(4): 303–333, April 1986. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Peine:2002:APE

[Pei02] Holger Peine. Application and programming experience with the Ara mobile agent system. Software—Practice and Experience, 32(6):515–541, May 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        93513613/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=93513613&PLACEBO=IE.
                                        pdf.

Pemberton:1980:CER

[Pem80] Steven Pemberton. Comments on an error-recovery scheme by Hartmann. Software—Practice and Experience, 10(3):231–240, March 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Perko:1985:IDS

[Per85] A. Perko. On implicit data structures for priority queues. Software—Practice and Experience, 15(1):55– 57, January 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Prasser:2020:FDA

[PES^{+}20] Fabian Prasser, Johanna Eicher, Helmut Spengler, Raffael Bild, and Klaus A. Kuhn. Flexible data anonymization using ARX — current status and challenges ahead. Software—Practice and Experience, 50(7):1277–1304, July 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Peterson:1976:CGS

[Pet76] Norman D. Peterson. Cobol generation of source programs and reports. Software—Practice and Experience, 6(1):117–131, January/March 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Petyt:1977:BRB

[Pet77] M. Petyt. Book review:

Shock and vibration computer programs: Reviews and summaries, edited by W. and B. Pilkey, The Shock and Vibration Information Center, Washington, DC, 1975. No. of pages: 663, price:

$25.00 in USA, $31-25 elsewhere. Software—Practice and Experience, 7(2):301– 302, March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Peterson:1988:DED

[Pet88] Larry L. Peterson. Dragonmail: An exercise in distributed computing. Software—Practice and Experience, 18(8):791–803, August 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Petty:2001:ACC

[Pet01] Grant W. Petty. Automated computation and consistency checking of physical dimensions and units in scientific programs. Software—Practice and Experience, 31(11):1067– 1076, September 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        84502037/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=84502037&PLACEBO=IE.
                                        pdf.

Ponder:1988:IPP

[PF88] Carl Ponder and Richard J. Fateman. Inaccuracies in program profilers. Software —Practice and Experience, 18(5):459–467, May 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Patil:1997:LCC

[PF97] Harish Patil and Charles Fischer. Low-cost, concurrent checking of pointer and array accesses in C programs. Software—Practice and Experience, 27(1):87– 110, January 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7259;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7259&
PLACEBO=IE.pdf.

Parrend:2009:SBO

[PF09] P. Parrend and S. Frenot.

Security benchmarks of OSGi platforms: toward Hardened OSGi. Software —Practice and Experience, 39(5):471–499, April 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pfeiffer:1984:FCG

[Pfe84] G. Pfeiffer. A flexible command generation technique for application in dialogue systems. Software —Practice and Experience, 14(5):483–489, May 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Prakash:1981:ERR

[PG81] Naveen Prakash and P. N. Gejji. Error reporting and recovery in Sangrah. Software—Practice and Experience, 11(6):613–621, June 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Page:1998:POR

[PGH^{+}98] T. W. Page, Jr., R. G. Guy, J. S. Heidemann, D. H. Ratner, P. L. Reiher, A. Goel, G. H. Kuenning, and G. J. Popek. Perspectives on optimistically replicated, peer-to-peer filing. Software—Practice and Experience, 28(2):155– 180, February 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=1777;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=1777&
                                        PLACEBO=IE.pdf.

Pervez:2010:FMA

[PGK^{+}10] Salman Pervez, Ganesh Gopalakrishnan, Robert M. Kirby, Rajeev Thakur, and William Gropp. Formal methods applied to high-performance computing software design: a case study of MPI one-sided communication-based locking. Software—Practice and Experience, 40(1):23– 43, January ??, 2010. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Posch:1984:ACR

[PH84] R. Posch and G. Haring. Allocating computer resources to satisfy user and management demands. Software—Practice and Experience, 14(4):311– 322 (or 311–321??), April 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Patel:1986:IAP

[PH86] R. R. Patel and R. B.

Hunter. Implementing AL-GOL 68 predicates. Software—Practice and Experience, 16(1):87–97, January 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pati:2014:SPS

[PH14] Tanumoy Pati and James H.

Hill. Survey paper: A survey report of enhancements to the visitor software design pattern. Software— Practice and Experience, 44(6):699–733, June 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pfandzelter:2021:ZFE

[PHB21] Tobias Pfandzelter, Jonathan Hasenburg, and David Bermbach. From zero to fog: Efficient engineering of fog-based Internet of Things applications. Software—Practice and Experience, 51(8):1798–1821, August 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Phipps:1999:COB

[Phi99] Geoffrey Phipps. Comparing observed bug and productivity rates for Java and C++. Software—Practice and Experience, 29(4):345– 358, April 10, 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55001844;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55001844&PLACEBO=IE.
pdf.

Purser:1984:PP

[PHS84] Michael Purser, Chris Horn, and John Sheehan. The PHS protocol. Software—Practice and Experience, 14(4):299–310, April 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pike:1987:TES

[Pik87] Rob Pike. The text editor sam. Software—Practice and Experience, 17(11): 813–845, November 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pike:1990:IN

[Pik90] Rob Pike. The implementation of Newsqueak. Software—Practice and Experience, 20(7):649–659, July 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pilkey:1975:SMC

[Pil75] Walter D. Pilkey. Structural member computer programs. Software—Practice and Experience, 5(1): 109–110, January 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pitkin:1982:BRB

[Pit82] W. J. Pitkin. Book review: Entity relationship approach to systems analysis and design, Peter Pin-Shan Chen (ed.), North-Holland 1980, No. of pages:

663. Price $68.25. Software—Practice and Experience, 12(5):495, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Purser:1975:DRT

[PJ75] W. F. C. Purser and D. M. Jennings. The design of a real-time operating system for a minicomputer. part 1. Software—Practice and Experience, 5(2): 147–167, April/June 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Partridge:1976:CTE

[PJ76] D. Partridge and E. B. James. Compiling techniques to exploit the pattern of language use. Software—Practice and Experience, 6(4):527–539, October/December 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

P:2021:FSM

[PJJM21] Jevitha K. P., Swaminathan Jayaraman, Bharat Jayaraman, and Sethumadhavan M. Finite-state model extraction and visualization from Java program execution. Software —Practice and Experience, 51(2):409–437, February 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pfandzelter:2023:MDR

[PJS^{+}23] Tobias Pfandzelter, Nils Japke, Trever Schirmer, Jonathan Hasenburg, and David Bermbach. Managing data replication and distribution in the fog with FReD. Software—Practice and Experience, 53(10): 1958–1981, October 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Powroz:1982:LE

[PK82] N. E. Powroz and A. C.

Kilgour. Letters to the editor. Software—Practice and Experience, 12(5):493– 494, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Park:1989:SAP

[PK89] Jong Soo Park and Myunghwan Kim. A selection algorithm with a practical upper bound on expected number of comparisons. Software—Practice and Experience, 19 (11):1105–1110, November 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Papaspyrou:2004:GEC

[PK04] Nikolaos S. Papaspyrou and Ioannis T. Kassios. GLU^{♮} embedded in C++: a marriage between multidimensional and object-oriented programming. Software— Practice and Experience, 34(7):609–630, June 2004.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pardo:2011:SLS

[PK11] Abelardo Pardo and Carlos Delgado Kloos. Sub-Collaboration: large-scale group management in collaborative learning. Software—Practice and Experience, 41(4):449–465, April 10, 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Park:2023:DCC

[PKB23] Hyukwoo Park, Seonghyun Kim, and Boram Bae. Dynamic code compression for JavaScript engine. Software—Practice and Experience, 53(5): 1196–1217, May 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pukall:2013:JFR

[PKC^{+}13] Mario Pukall, Christian Kästner, Walter Cazzola, Sebastian Götz, Alexander Grebhahn, Reimar Schröter, and Gunter Saake. JavAdaptor — flexible runtime updates of Java applications. Software —Practice and Experience, 43(2):153–185, February 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pryanishnikov:2007:COP

[PKH07] Ivan Pryanishnikov, Andreas Krall, and Nigel Horspool. Compiler optimizations for processors with SIMD instructions. Software—Practice and Experience, 37(1):93–113, January 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Park:2012:TBR

[PKK12] Yu-Sik Park, Hyung-Min Koo, and In-Young Ko. A task-based and resourceaware approach to dynamically generate optimal software architecture for intelligent service robots. Software—Practice and Experience, 42(5):519–541, May 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Park:2012:EHV

[PKN^{+}12] Yoonho Park, Richard King, Senthil Nathan, Wesley Most, and Henrique Andrade. Evaluation of a high-volume, low-latency market data processing system implemented with IBM middleware. Software —Practice and Experience, 42(1):37–56, January 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pruijt:2017:ECP

[PKvdWB17] Leo Pruijt, Christian Köppe, Jan Martijn van der Werf, and Sjaak Brinkkemper. Extended conference papers: The accuracy of dependency analysis in static architecture compliance checking. Software —Practice and Experience, 47(2):273–309, February 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Perrott:1991:SDI

[PL91] R. H. Perrott and T. F. Lunney. A syntax-directed integrated programming environment for developing SIMD supercomputer software. Software—Practice and Experience, 21(3):269– 286, March 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Paulino:2008:PLS

[PL08] Hervé Paulino and Luís Lopes. A programming language for service-oriented computing with mobile agents. Software—Practice and Experience, 38(7):705– 734, June ??, 2008. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Plantec:2014:EIW

[PL14] A. Plantec and L. Lagadec. Editorial: International Workshop on Smalltalk Technologies 2011, special issue. Software—Practice and Experience, 44(4):391– 392, April 2014. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Special issue on International Workshop on Smalltalk Technologies 2011.

Plank:1997:TRS

[Pla97] James S. Plank. A tutorial on Reed–Solomon coding for fault-tolerance in RAID-like systems. Software—Practice and Experience, 27(9):995–1012, September 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7330;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7330&
PLACEBO=IE.pdf. See cor-

rection [PD05].

Plestenjak:1999:ADP

[Ple99] Bor Plestenjak. An algorithm for drawing planar graphs. Software— Practice and Experience, 29(11):973–984, September 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=63501203;   http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=63501203&PLACEBO=IE.
                                        pdf.

Pacheco-Lorenzo:2023:ISC

[PLFIVRAR23] Moisés Pacheco-Lorenzo, Manuel J. Fernández-Iglesias, Sonia Valladares-Rodriguez, and Luis E. Anido-Rifón. Implementing scripted conversations by means of smart assistants. Software—Practice and Experience, 53(5): 1271–1283, May 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Palopoli:2002:OOT

[PLL^{+}02] Luigi Palopoli, Giuseppe Lipari, Gerardo Lamastra, Luca Abeni, Gabriele Bolognini, and Paolo Ancilotti. An object-oriented tool for simulating distributed real-time control systems. Software —Practice and Experience, 32(9):907–932, July 25, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        94519527/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=94519527&PLACEBO=IE.
                                        pdf.

Passini:2022:DFS

[PLPA22] William Filisbino Passini, Cristiane Aparecida Lana, Vinicius Pfeifer, and Frank José Affonso. Design of frameworks for self-adaptive service-oriented applications: a systematic analysis. Software—Practice and Experience, 52(1):5– 38, January 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pike:1985:HST

[PLR85] Rob Pike, Bart Locanthi, and John Reiser. Hardware/software trade-offs for bitmap graphics on the Blit. Software—Practice and Experience, 15(2):131– 151, February 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Park:2013:UDP

[PLR13] Jaeyong Park, Seok-Won Lee, and David C. Rine. UML design pattern metamodellevel constraints for the maintenance of software evolution. Software—Practice and Experience, 43(7): 835–866, July 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Park:2018:SAJ

[PLR18] Changhee Park, Hongki Lee, and Sukyoung Ryu.

Static analysis of JavaScript libraries in a scalable and precise way using loop sensitivity. Software—Practice and Experience, 48 (4):911–944, April 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2552.

Plum:1974:RSQ

[Plu74] Thomas W.-S. Plum. Random search on the 8-queens problem or a simple solution for a simple problem. Software—Practice and Experience, 4(3):251– 253, July/September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Plum:1977:FUP

[Plu77] Thomas Plum. Fooling the user of a programming language. Software—Practice and Experience, 7(2):215– 221, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Peck:1981:CPE

[PM81] J. E. L. Peck and M. A. Maclean. The construction of a portable editor. Software—Practice and Experience, 11(5):479– 489, May 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Parkin:2012:TRS

[PM12] Simon E. Parkin and Graham Morgan. Toward reusable SLA monitoring capabilities. Software—Practice and Experience, 42(3):261–280, March 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Panda:2017:RTS

[PM17] S. Panda and D. P. Mohapatra. Regression test suite minimization using integer linear programming model. Software—Practice and Experience, 47 (11):1539–1560, November 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Petri:2018:CII

[PM18] Matthias Petri and Alistair Moffat. Compact inverted index storage using general-purpose compression libraries. Software —Practice and Experience, 48(4):974–982, April 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2556.

Parson:2005:OOD

[PMC05] Dale Parson, David J. Murray, and Yu Chen. Objectoriented design patterns for debugging heterogeneous languages and virtual machines. Software—Practice and Experience, 35(3):255– 279, March 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Poulos:2022:PSN

[PMC22] Alexandra Poulos, Sally A. McKee, and Jon C. Calhoun. Posits and the state of numerical representations in the age of exascale and edge computing. Software—Practice and Experience, 52(2):619– 635, February 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pyle:1971:SOB

[PMG71] I. C. Pyle, R. C. F. McLatchie, and B. Grandage. A second-order bug with delayed effect. Software —Practice and Experience, 1(3):231–233, July/ September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pawlak:2016:SLI

[PMP^{+}16] Renaud Pawlak, Martin Monperrus, Nicolas Petitprez, Carlos Noguera, and Lionel Seinturier. SPOON: a library for implementing analyses and transformations of Java source code. Software—Practice and Experience, 46 (9):1155–1179, September 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pentakalos:1997:PPW

[PMY97] Odysseas I. Pentakalos, Daniel A. Menascé, and Yelena Yesha. Pythia and Pythia/WK: Tools for the performance analysis of mass storage systems. Software—Practice and Experience, 27(9):1035–1054, September 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7332;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7332&
PLACEBO=IE.pdf.

Pong:1983:PSP

[PN83] M. C. Pong and N. Ng.

PIGS: a system for programming with interactive graphical support. Software—Practice and Experience, 13(9):847–855, September 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Plesinger:2023:DOS

[PNK^{+}23] Filip Plesinger, Petr Nejedly, Zuzana Koscova, Maurice Rohr, Ivo Viscor, Radovan Smisek, Adam Ivora, Pavel Leinveber, Karol Curila, and Christoph Hoog Antink. DeepPlayer: an opensource SignalPlant plugin for deep learning inference. Software—Practice and Experience, 53(2):455– 464, February 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Perez:2020:OPN

[PNM^{+}20] Taciano D. Perez, Marcelo V. Neves, Diego Medaglia, Pedro H. G. Monteiro, and César A. F. De Rose. Orthogonal persistence in nonvolatile memory architectures: a persistent heap design and its implementation for a Java Virtual Machine. Software—Practice and Experience, 50(4):368– 387, April 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pattipati:2020:OEF

[PNP20] Dileep Kumar Pattipati, Rupesh Nasre, and Sreenivasa Kumar Puligundla. OPAL: an extensible framework for ontology-based program analysis. Software —Practice and Experience, 50(8):1425–1462, August 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pohjanpalo:1981:MMR

[Poh81] Hannu Pohjanpalo. MROS-68K, a memory resident operating system, for MC68000. Software—Practice and Experience, 11(8): 845–852, August 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Polack:2001:CSU

[Pol01] Fiona Polack. A case study using lightweight formalism to review an information system specification. Software—Practice and Experience, 31(8):757– 780, July 10, 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
78502911/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=78502911&PLACEBO=IE.
pdf.

Pichler:2003:ACM

[POM03] Roman Pichler, Klaus Ostermann, and Mira Mezini. On aspectualizing component models. Software— Practice and Experience, 33(10):957–974, August 2003. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Poole:1971:IEA

[Poo71a] M. D. Poole. Implementation of an editing algorithm allowing repeating corrections. Software—Practice and Experience, 1(4):373– 381, October/December 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Poole:1971:DMA

[Poo71b] P. C. Poole. Developing a multi-access system online. Software—Practice and Experience, 1(1):39– 50, January/March 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Poore:1988:DLS

[Poo88] J. H. Poore. Derivation of local software quality metrics (software quality circles). Software—Practice and Experience, 18 (11):1017–1027, November 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Powell:1979:ETU

[Pow79] M. S. Powell. Experience of transporting and using the SOLO operating system. Software—Practice and Experience, 9(7):561– 569, July 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Powell:1987:STU

[Pow87] M. S. Powell. Strongly typed user interfaces in an abstract data store. Software—Practice and Experience, 17(4):241–266, April 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Powell:1995:APO

[Pow95] Michael L. Powell. Alternative perspectives on object technology. Software—Practice and Experience, 25(S4):131–141, December 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Petkovich:2016:ECP

[POZ^{+}16] J. C. Petkovich, A. Oliveira, Y. Zhang, T. Reidemeister, and S. Fischmeister. Extended conference paper: DataMill: a distributed heterogeneous infrastructure for robust experimentation. Software—Practice and Experience, 46(10): 1411–1440, October 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Patel:1980:SPD

[PP80] Ahmed Patel and Michael Purser. Systems programming for data communications on minicomputers. Software—Practice and Experience, 10(4):283– 305, April 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Parkyn:1984:DME

[PP84] D. G. Parkyn and N. J. Pilbeam. The direct mapping of the entity-relationship model. Software—Practice and Experience, 14 (11):1109–1110, November 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Prowell:1998:SBS

[PP98] Stacy J. Prowell and Jesse H. Poore. Sequencebased software specification of deterministic systems. Software—Practice and Experience, 28(3):329– 345, March 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1783;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1783&
PLACEBO=IE.pdf.

Purdila:2016:SSF

[PP16] Vasile Purdilă and Ştefan-Gheorghe Pentiuc. Singlescan: a fast star-join query processing algorithm. Software—Practice and Experience, 46(3):319–339, March 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Prastowo:2020:TRT

[PPA20] Tadeus Prastowo, Luigi Palopoli, and Luca Abeni. Tice: a real-time language compilable using C++ compilers. Software—Practice and Experience, 50(12):2252–2292, December 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Parama:2006:DVL

[PPBP06] José R. Paramá, Ángeles S.

Places, Nieves R. Brisaboa, and Miguel R. Penabad. The design of a Virtual Library of Emblem Books. Software—Practice and Experience, 36(5): 473–494, April 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Polyzoidou:2023:SDP

[PPN^{+}23] Eleni Polyzoidou, Evangelia Papagiannaki, Nikolaos Nikolaidis, Apostolos Ampatzoglou, Nikolaos Mittas, Elvira Maria Arvanitou, Alexander Chatzigeorgiou, George Manolis, and Evdoxia Manganopoulou. SmartCLIDE design pattern assistant: a decisiontree based approach. Software—Practice and Experience, 53(6):1304–1331, June 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pati:2023:IIB

[PPPP23] Abhilash Pati, Manoranjan Parhi, Binod Kumar Pattanayak, and Bibudhendu Pati. IFCnCov: an IoT-based smart diagnostic architecture for COVID-19. Software—Practice and Experience, 53(11):2133– 2162, November 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Polo:2002:UQR

[PPR02] Macario Polo, Mario Piattini, and Francisco Ruiz. Using a qualitative research method for building a software maintenance methodology. Software—Practice and Experience, 32(13):1239–1260, November 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
98517572/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=98517572&PLACEBO=IE.
pdf.

Parimala:2021:SBS

[PPR^{+}21] M. Parimala, R. M. Swarna Priya, M. Praveen Kumar Reddy, Chiranji Lal Chowdhary, Ravi Kumar Poluru, and Suleman Khan. Spatiotemporalbased sentiment analysis on tweets for risk assessment of event using deep learning approach. Software—Practice and Experience, 51(3):550–570, March 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pereira:2017:MAD

[PPSO17] Rui Humberto R. Pereira, J. Baltasar García Perez-Schofield, and Francisco Ortin. Modularizing application and database evolution — an aspect-oriented framework for orthogonal persistence. Software— Practice and Experience, 47(2):193–221, February 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Power:2005:TSG

[PPSS05] D. J. Power, E. A. Politou, M. A. Slaymaker, and A. C. Simpson. Towards secure Grid-enabled healthcare. Software—Practice and Experience, 35(9):857– 871, July 25, 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Parr:1995:APL

[PQ95] Terence J. Parr and Russell W. Quong. ANTLR: A predicated- LL ( k ) parser generator. Software—Practice and Experience, 25(7):789–810, July 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL ftp://ftp.

                                        parr-research.com/pub/
                                        pccts/papers/antlr.ps;
                                         ftp://ftp.parr-research.
                                        com/pub/pccts/papers/
                                        needlook.ps;   ftp://
                                        ftp.parr-research.com/
                                        pub/pccts/papers/predicates.
                                        ps.Z;   ftp://ftp.parr-
                                        research.com/pub/pccts/
                                        papers/sorcerer.ps.Z;
                                         http://java.magelang.
                                        com/antlr/entry.html;
                                         http://www.parr-research.
                                        com/antlr/parr.phd.thesis.
                                        ps.zip.

Perrott:1977:QT

[PR77] R. H. Perrott and A. K. Raja. Quasiparallel tracing. Software—Practice and Experience, 7(4):483– 492, July/August 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Presotto:1990:ICN

[PR90] David L. Presotto and Dennis M. Ritchie. Interprocess communication in the Ninth Edition Unix system. Software—Practice and Experience, 20(S1):S3– S17, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Poggi:1998:EFC

[PR98] Agostino Poggi and Giovanni Rimassa. An efficient and flexible C++ library for concurrent programming. Software— Practice and Experience, 28(13):1437–1463, November 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1771;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1771&
PLACEBO=IE.pdf.

Pizzonia:2016:NNE

[PR16] Maurizio Pizzonia and Massimo Rimondini. Netkit: network emulation for education. Software—Practice and Experience, 46(2):133– 165, February 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Prasad:1980:VNP

[Pra80] V. R. Prasad. Variable number of parameters in typed languages. Software —Practice and Experience, 10(7):507–517, July 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Prasad:1996:BRBa

[Pra96a] Divya Prasad. Book review: Software Quality Assurance and Measurement. A Worldwide Perspective, Edited by Norman Fenton, Robin Whitty and Yoshinori Lizuka, International Thomson Computer Press, 1995. Price: £35. No. of pages: 315. Software—Practice and Experience, 26(8):983, August 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16830.

Prasad:1996:BRBb

[Pra96b] Divya Prasad. Book review: Software Quality Assurance And Measurement. A Worldwide Perspective, Edited by Norman Fenton, Robin Whitty and Yoshinori Lizuka, International Thomson Computer Press, 1995. Price: £35. No. of pages: 315. Software —Practice and Experience, 26(8):983, August 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pal:2006:AAI

[PRA^{+}06] Partha Pal, Paul Rubel, Michael Atighetchi, Franklin Webber, William H. Sanders, Mouna Seri, HariGovind Ramasamy, James Lyons, Tod Courtney, Adnan Agbaria, Michel Cukier, Jeanna Gossett, and Idit Keidar. An architecture for adaptive intrusiontolerant applications. Software—Practice and Experience, 36(11–12):1331–1354, September/October 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Presotto:1985:USA

[Pre85] David L. Presotto. Upas — a simpler approach to network mail. In USENIX Association [USE85], pages 533–538. LCCN QA76.8.U65 U8 1985.

Pronk:1992:STC

[Pro92] C. Pronk. Stress testing of compilers for Modula-2. Software—Practice and Experience, 22(10):885– 897, October 1992. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Piantadosi:2023:DFS

[PRP^{+}23] Valentina Piantadosi, Giovanni Rosa, Davide Placella, Simone Scalabrino, and Rocco Oliveto. Detecting functional and securityrelated issues in smart contracts: a systematic literature review. Software —Practice and Experience, 53(2):465–495, February 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Parsons:2006:APD

[PRTS06] David Parsons, Awais Rashid, Alexandru Telea, and Andreas Speck. An architectural pattern for designing component-based application frameworks. Software—Practice and Experience, 36(2):157–190, February 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pryce:1985:EWL

[Pry85] J. D. Pryce. Experiences with writing library software for an attached processor. Software—Practice and Experience, 15(7):705– 714, July 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pauli:1980:CBI

[PS80] W. Pauli and Mary Lou Soffa. Coroutine behaviour and implementation. Software—Practice and Experience, 10(3):189–204, March 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Perrott:1981:CDA

[PS81] R. H. Perrott and D. K.

Stevenson. Considerations for the design of array processing languages. Software —Practice and Experience, 11(7):683–688, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Purdin:1987:FRF

[PSA87] Titus D. M. Purdin, Richard D. Schlichting, and Gregory R. Andrews. A file replication facility for Berkeley Unix. Software —Practice and Experience, 17(12):923–940, December 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pawlak:2004:JAB

[PSD^{+}04] Renaud Pawlak, Lionel Seinturier, Laurence Duchien, Gérard Florin, Fabrice Legond-Aubry, and Laurent Martelli. JAC: an aspect-based distributed dynamic framework. Software—Practice and Experience, 34(12):1119–1148, October 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Patnaik:1983:APQ

[PSR83] L. M. Patnaik, Phule Shailendra, and K. Venkateswara Rao. Access path query language for relational database systems. Software—Practice and Experience, 13(8):661–670, August 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Perez-Schofield:2002:MSE

[PSRCC02] J. Baltasar García Perez-Schofield, Emilio García Roselló, Tim B. Cooper, and Manuel Pérez Cota. Managing schema evolution in a container-based persistent system. Software—Practice and Experience, 32(14):1395–1410, November 25, 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
98518883/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=98518883&PLACEBO=IE.
pdf.

Paolino:2010:TNA

[PSTV10] Luca Paolino, Monica Sebillo, Genoveffa Tortora, and Giuliana Vitiello. Towards a new approach to query search engines: the Search Tree visual language. Software—Practice and Experience, 40(8):735– 750, July 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Panchapakesan:1985:IAL

[PSV85] S. Panchapakesan, S. Subramanian, and H. Venkateswaran. An interactive assembly level debugging system. Software—Practice and Experience, 15(1):59–64, January 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Paulisch:1990:EEG

[PT90] Frances Newbery Paulisch and Walter F. Tichy. EDGE: An extendible graph editor. Software— Practice and Experience, 20(S1):S63–S88, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Petrakis:2000:SSC

[PT00a] Euripides G. M. Petrakis and Kostas Tzeras. Similarity searching in the CORDIS text database. Software—Practice and Experience, 30(13):1447– 1464, November 10, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www3.interscience.
wiley.com/cgi-bin/abstract/
73001208/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?

ID=73001208&PLACEBO=IE.
pdf.

Proebsting:2000:NII

[PT00b] Todd A. Proebsting and Gregg M. Townsend. A new implementation of the Icon language. Software—Practice and Experience, 30(8):925–972, July 10, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72502208/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72502208&PLACEBO=IE.
pdf.

Padhye:2014:ECP

[PT14] Vinit Padhye and Anand Tripathi. Extended conference paper: Mechanisms for building autonomically scalable services on cooperatively shared computing platforms. Software —Practice and Experience, 44(10):1251–1276, October 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Paliyawan:2017:UUK

[PT17] Pujana Paliyawan and Ruck Thawonmas. UKI: universal Kinect-type controller by ICE Lab. Software—Practice and Experience, 47(10):1343–1363, October 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Palopoli:2003:DSS

[PTU03] Luigi Palopoli, Giorgio Terracina, and Domenico Ursino. DIKE: a system supporting the semiautomatic construction of cooperative information systems from heterogeneous databases. Software —Practice and Experience, 33(9):847–884, July 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pashtan:1984:RMD

[PU84] Ariel Pashtan and Elizabeth A. Unger. Resource monitors: a design methodology for operating systems. Software—Practice and Experience, 14(8):791– 806, August 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Purser:1976:DRT

[Pur76] W. F. C. Purser. The design of a real-time operating system for a minicomputer. part II. Software —Practice and Experience, 6(3):327–340, July 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Perkins:1984:UPV

[PV84] Daniel R. Perkins and Dennis Volper. UCSD Pascal on the VAX, portability and performance. Software —Practice and Experience, 14(5):473–482, May 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pibiri:2021:PTO

[PV21] Giulio Ermanno Pibiri and Rossano Venturini. Practical trade-offs for the prefixsum problem. Software —Practice and Experience, 51(5):921–949, May 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pozzan:2022:RMV

[PV22] Gabriele Pozzan and Tullio Vardanega. Rafting multiplayer video games. Software—Practice and Experience, 52(4):1065–1091, April 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Paredes-Valverde:2015:SRT

[PVAHRG^{+}15] Mario Andrés Paredes-Valverde, Giner Alor-Hernández, Alejandro Rodríguez-González, Rafael Valencia-García, and Enrique Jiménez-Domingo. A systematic review of tools, languages, and methodologies for mashup development. Software—Practice and Experience, 45(3):365– 397, March 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pereira:2006:AFO

[PVBB06] F. M. Q. Pereira, M. T. O.

Valente, R. S. Bigonha, and M. A. S. Bigonha. Arcademis: a framework for object-oriented communication middleware development. Software—Practice and Experience, 36(5): 495–512, April 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Potok:1999:PAO

[PVR99] Thomas E. Potok, Mladen Vouk, and Andy Rindos. Productivity analysis of object-oriented software developed in a commercial environment. Software—Practice and Experience, 29(10):833–847, August 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=63001370;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=63001370&PLACEBO=IE.
pdf.

Panchapakesan:1979:AM

[PVS79] S. Panchapakesan, H. Venkateswaran, and S. Subramanian. Assemblers for microcomputers. Software—Practice and Experience, 9(10):843– 852, October 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Plaice:1993:UTM

[PW93] John Plaice and William W.

Wadge. A Unix tool for managing reusable software components. Software— Practice and Experience, 23(9):933–948, September 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pedrycz:1997:FCS

[PW97] Witold Pedrycz and James Waletzky. Fuzzy clustering in software reusability. Software—Practice and Experience, 27(3):245– 270, March 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=7290;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=7290&
                                        PLACEBO=IE.pdf.

Petersen:2011:MFL

[PW11] K. Petersen and C. Wohlin.

Measuring the flow in lean software development. Software—Practice and Experience, 41(9):975– 996, August 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pearce:2007:PA

[PWBK07] David J. Pearce, Matthew Webster, Robert Berry, and Paul H. J. Kelly. Profiling with AspectJ. Software —Practice and Experience, 37(7):747–777, June 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pham:2021:ASA

[PWI^{+}21] Truong An Pham, Junjue Wang, Roger Iyengar, Yu Xiao, Padmanabhan Pillai, Roberta Klatzky, and Mahadev Satyanarayanan. Ajalon: Simplifying the authoring of wearable cognitive assistants. Software—Practice and Experience, 51(8): 1773–1797, August 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pyle:1972:STM

[Pyl72] I. C. Pyle. Some techniques in multi-computer system software design. Software—Practice and Experience, 2(1):43–54, January/ March 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pyle:1979:IOH

[Pyl79] I. C. Pyle. Input/output in high level programming languages. Software —Practice and Experience, 9(11):907–914, November 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pyle:1980:AUD

[Pyl80] I. C. Pyle. Axioms for userdefined operators. Software —Practice and Experience, 10(4):307–318, April 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pyle:1984:TTC

[Pyl84] I. C. Pyle. Transducers are type-converters. Software —Practice and Experience, 14(9):889–899, September 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pong:1992:CCT

[PZ92] Man-Chi Pong and Yongguang Zhang. cxterm: A Chinese terminal emulator for the X Window System. Software—Practice and Experience, 22(10): 809–826, October 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Parson:2000:JNI

[PZ00] Dale Parson and Zhenyu Zhu. Java Native Interface idioms for C++ class hierarchies. Software—Practice and Experience, 30(15):1641–1660, December 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
73501271/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=73501271&PLACEBO=IE.
pdf.

Perrott:1987:SPD

[PZA87] R. H. Perrott and Adib Zarea-Aliabadi. A supercomputer program development system. Software —Practice and Experience, 17(10):663–683, October 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Psiuk:2013:DOB

[PŻZ13] Marek Psiuk, Daniel Żmuda, and Krzysztof Zieliński. Distributed OSGi built over message-oriented middleware. Software—Practice and Experience, 43(1): 1–31, January 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Qian:1983:AXP

[QC83] Lu Ru Qian and Wei Zi Chu. An assessment of the XR project for compiler development and transportation. Software—Practice and Experience, 13(11): 993–1004, November 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Quick:2017:BFD

[QC17] Darren Quick and Kim-Kwang Raymond Choo. Big forensic data management in heterogeneous distributed systems: quick analysis of multimedia forensic data. Software— Practice and Experience, 47(8):1095–1109, August 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Qadeer:2021:PES

[QH21] Abdul Qadeer and John Heidemann. Plumb: Efficient stream processing of multi-user pipelines. Software—Practice and Experience, 51(2):385–408, February 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Quittner:1978:CDD

[QK78] P. Quittner and D. Kotsis. Comparison of different disk searching methods. Software—Practice and Experience, 8(6):673– 679, November/December 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Queiros:2013:ECP

[QL13] Ricardo Queirós and José Paulo Leal. Extended conference papers: crimsonHex: a learning objects repository for programming exercises. Software—Practice and Experience, 43(8):911– 935, August 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Qasem:2013:ECP

[QM13] Apan Qasem and Josh Magee. Extended conference paper: Improving TLB performance on current chip multiprocessor architectures through demand-driven superpaging. Software—Practice and Experience, 43(6):705– 729, June 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Quinton:2016:ECP

[QRD16] Clément Quinton, Daniel Romero, and Laurence Duchien. Extended conference papers: SALOON: a platform for selecting and configuring cloud environments. Software—Practice and Experience, 46(1):55– 78, January 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Qin:1988:RSS

[QSA88] Bin Qin, Howard A. Sholl, and Reda A. Ammar. RTS: a system to simulate the real time cost behaviour of parallel computations. Software—Practice and Experience, 18(10): 967–985, October 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Qin:1990:TMT

[QSA90] Bin Qin, Howard A. Sholl, and Reda A. Ammar. A tool to minimize the time costs of parallel computations through optimal processing power allocation. Software—Practice and Experience, 20(3):283– 300, March 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Quattrocchi:2023:MSC

[QTV23] Giovanni Quattrocchi, Damian Andrew Tamburri, and Willem-Jan Van Den Heuvel. Making service continuity smarter with artificial intelligence: an approach and its evaluation. Software—Practice and Experience, 53(10):1887–1901, October 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Quittner:1983:ECI

[Qui83] Pál Quittner. Efficient combination of index tables and hashing. Software —Practice and Experience, 13(6):471–478, June 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Quinlan:1991:CWF

[Qui91] Sean Quinlan. A cached WORM file system. Software—Practice and Experience, 21(12):1289–1299, December 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rasmussen:1987:RTI

[RA87] Jørgen Born Rasmussen and Bill Appelbe. Realtime interrupt handling in Ada. Software—Practice and Experience, 17(3):197– 213, March 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robinson:1995:DPC

[RA95] Patrick G. Robinson and James D. Arthur. Distributed process creation within a shared data space framework. Software— Practice and Experience, 25(2):175–191, February 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Richards:1979:TPO

[RAB^{+}79] M. Richards, A. R. Aylward, P. Bond, R. D. Evans, and B. J. Knight. TRIPOS: a portable operating system for minicomputers. Software— Practice and Experience, 9(7):513–526, July 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Radford:1980:CCP

[Rad80] P. Radford. Checking cyclic pointers. Software—Practice and Experience, 10(7): 519–521, July 1980. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reyes-Alvarez:2019:TMO

[RÁdMRGAM19] Liudmila Reyes-Álvarez, María del Mar Roldán-García, and José F. Aldana-Montes. Tool for materializing OWL ontologies in a column-oriented database. Software—Practice and Experience, 49 (1):100–119, January 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ragan:1986:CLD

[Rag86] Richard R. Ragan. CON-TEXT: An on-line documentation system. Software—Practice and Experience, 16(3):217–224, March 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rain:1972:SCC

[Rai72] Mark Rain. Short communications and comment: Letters to the editor. Software—Practice and Experience, 2(1):97, January 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rain:1973:TUM

[Rai73]  Mark Rain.  Two un-

usual methods for debugging system software. Software—Practice and Experience, 3(1):61–63, January/ March 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rain:1981:SMC

[Rai81] Mark Rain. The structure of the MARY/2 compiler. Software—Practice and Experience, 11(3):225– 235, March 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rain:1984:ATR

[Rai84]  Mark Rain.  Avoiding

trickle-down recompilation in the Mary2 implementation. Software—Practice and Experience, 14 (12):1149–1157, December 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Raita:1992:TBM

[Rai92] Timo Raita. Tuning the Boyer–Moore–Horspool string searching algorithm. Software—Practice and Experience, 22(10):879–884, October 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Raita:1999:GSD

[Rai99] Timo Raita. On guards and symbol dependencies in substring search. Software—Practice and Experience, 29(11):931–941, September 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=63501200;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=63501200&PLACEBO=IE.
pdf.

Ramsay:1983:DPA

[Ram83] Allan Ramsay. A distributed programming assistant. Software—Practice and Experience, 13(11): 983–992, November 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ramsey:1996:SSL

[Ram96] Norman Ramsey. A simple solver for linear equations containing nonlinear operators. Software—Practice and Experience, 26(4): 467–487, April 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16800.

Ramsey:1998:UEP

[Ram98] Norman Ramsey. Unparsing expressions with prefix and postfix operators. Software—Practice and Experience, 28(12):1327–1356, October 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1761;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1761&
PLACEBO=IE.pdf.

Rak-amnouykit:2024:PPS

[RaMB^{+}24] Ingkarat Rak-amnouykit, Ana Milanova, Guillaume Baudart, Martin Hirzel, and Julian Dolby. Principled and practical static analysis for Python: Weakest precondition inference of hyperparameter constraints. Software—Practice and Experience, 54(3): 363–393, March 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ruddle:2003:ALW

[RAN03] Alan Ruddle, Colin Allison, and R. Nicoll. Analysing the latency of World Wide Web applications. Software —Practice and Experience, 33(14):1301–1322, November 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rimsa:2021:PDR

[RAP21] Andrei Rimsa, José Nelson Amaral, and Fernando M. Q. Pereira. Practical dynamic reconstruction of control flow graphs. Software—Practice and Experience, 51(2):353–384, February 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ravn:1982:PVC

[Rav82] A. P. Ravn. Pointer variables in Concurrent Pascal. Software—Practice and Experience, 12(3):211– 222, March 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rayner:1975:RDM

[Ray75] D. Rayner. Recent developments in machineindependent job control languages. Software— Practice and Experience, 5(4):375–393, October/ December 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rin:1975:OSA

[RB75] N. Adam Rin and Maxine Brown. An overview of a system for automatic generation of file conversion programs. Software—Practice and Experience, 5(2): 193–202, April/June 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Roper:1981:CSP

[RB81] T. J. Roper and Chris J.

Barter. A communicating sequential process language and implementation. Software—Practice and Experience, 11(11):1215–1234, November 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rees:1982:BRBb

[RB82] Michael J. Rees and D. W.

Barron. Book review: Programming in standard FORTRAN 77, A. Balfour and D. H. Marwick, Heinemann Educational Books, London, 1979. No. of pages:

388. Price: £4.50 (softback). Software—Practice and Experience, 12(10): 983–984, October 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rosenberg:1989:MMS

[RB89] Jonathan B. Rosenberg and Jonathan D. Becker. Mapping massive SIMD parallelism onto vector architectures for simulation. Software—Practice and Experience, 19(8):739– 756, August 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rees:1991:EID

[RB91] Stephen A. Rees and James P. Black. An experimental investigation of distributed matrix multiplication techniques. Software—Practice and Experience, 21(10):1041–1063, October 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rodriguez:2019:CBC

[RB19] Maria A. Rodriguez and Rajkumar Buyya. Containerbased cluster orchestration systems: a taxonomy and future directions. Software —Practice and Experience, 49(5):698–719, May 2019.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ranjan:2012:ESS

[RBB12] Rajiv Ranjan, Rajkumar Buyya, and Boualem Benatallah. Editorial: Special section: Software architectures and application development environments for Cloud computing. Software—Practice and Experience, 42(4):391–394, April 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Roy:2024:SSW

[RBD^{+}24] Mandira Roy, Raunak Bag, Novarun Deb, Agostino Cortesi, Rituparna Chaki, and Nabendu Chaki. SCARS: Suturing wounds due to conflicts between nonfunctional requirements in autonomous and robotic systems. Software—Practice and Experience, 54(5): 759–795, May 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ranjan:2014:ENS

[RBL^{+}14a] Rajiv Ranjan, Rajkumar Buyya, Philipp Leitner, Armin Haller, and Stefan Tai. Editorial: A note on software tools and techniques for monitoring and prediction of cloud services. Software—Practice and Experience, 44(7):771– 775, July 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reiss:2014:PCB

[RBL14b] Steven P. Reiss, Jared N. Bott, and Joseph J. La Viola, Jr. Plugging in and into code bubbles: the code bubbles architecture. Software—Practice and Experience, 44(3):261– 276, March 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Riguzzi:2016:PLP

[RBL^{+}16] Fabrizio Riguzzi, Elena Bellodi, Evelina Lamma, Riccardo Zese, and Giuseppe Cota. Probabilistic logic programming on the web. Software—Practice and Experience, 46(10):1381– 1396, October 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rahad:2021:HMD

[RBR21] Khandoker Rahad, Omar Badreddin, and Sayed Mohsin Reza. The human in model-driven engineering loop: a case study on integrating handwritten code in model-driven engineering repositories. Software—Practice and Experience, 51(6):1308–1321, June 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rosales:2023:LSC

[RBRB23] Eduardo Rosales, Matteo Basso, Andrea Rosà, and Walter Binder. Large-scale characterization of Java streams. Software—Practice and Experience, 53 (9):1763–1792, September 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Richardson:2014:GSR

[RBS14] Thomas Richardson, Liz Burd, and Shamus Smith. Guidelines for supporting real-time multi-touch applications. Software —Practice and Experience, 44(8):931–949, August 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Richardson:1989:PLI

[RC89] Joel E. Richardson and Michael J. Carey. Persistence in the E language: Issues and implementation. Software—Practice and Experience, 19 (12):1115–1150, December 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rising:1992:PDP

[RC92] Linda Rising and Frank W. Calliss. Problems with determining package cohesion and coupling. Software —Practice and Experience, 22(7):553–571, July 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robinson:2010:PPS

[RC10] Peter J. Robinson and Keith L. Clark. Pedro: a publish/subscribe server using Prolog technology. Software—Practice and Experience, 40(4): 313–329, April 10, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rodrigues:2019:CBP

[RCA^{+}19] Leonardo R. Rodrigues, Euclides Cardoso Jr., Omir C. Alves Jr., Fernando F. Redígolo, Maurício A. Pillon, Charles C. Miers, and Guilherme P. Koslovski. Cloud broker proposal based on multicriteria decision-making and virtual infrastructure migration. Software—Practice and Experience, 49 (9):1331–1351, September 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robillard:1991:PST

[RCC91] Pierre N. Robillard, Daniel Coupal, and François Coallier. Profiling software through the use of metrics. Software—Practice and Experience, 21(5):507– 518, May 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ramadasan:2017:LGE

[RCC17] Datta Ramadasan, Marc Chevaldonné, and Thierry Chateau. LMA: a generic and efficient implementation of the Levenberg– Marquardt Algorithm. Software—Practice and Experience, 47(11):1707–1727, November 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Richmond:2023:FGF

[RCH^{+}23] Paul Richmond, Robert Chisholm, Peter Heywood, Mozhgan Kabiri Chimeh, and Matthew Leach. FLAME GPU 2: a framework for flexible and performant agent based simulation on GPUs. Software—Practice and Experience, 53(8):1659–1680, August 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rodriguez:2013:BPD

[RCMZ13] Juan Manuel Rodriguez, Marco Crasso, Cristian Mateos, and Alejandro Zunino. Best practices for describing, consuming, and discovering Web services: a comprehensive toolset. Software—Practice and Experience, 43(6):613– 639, June 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ryckbosch:2014:APT

[RD14] Frederick Ryckbosch and Amer Diwan. Analyzing performance traces using temporal formulas. Software—Practice and Experience, 44(7):777–792, July 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rowe:1989:VSI

[RDC89] Lawrence A. Rowe, Peter Danzig, and Wilson Choi. A visual shell interface to a database. Software—Practice and Experience, 19(6): 515–528, June 1989. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reingold:1993:CCI

[RDC93] Edward M. Reingold, Nachum Dershowitz, and Stewart M. Clamen. Calendrical calculations, II:

Three historical calendars. Software—Practice and Experience, 23(4):383– 404, April 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rubira:2005:EHD

[RdLFF05] C. M. F. Rubira, R. de Lemos, G. R. M. Ferreira, and F. Castor Filho. Exception handling in the development of dependable component-based systems. Software—Practice and Experience, 35(3):195– 236, March 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rajlich:1990:VTS

[RDLK90] Vaclav Rajlich, Nicholas Damaskinos, Panagiotis Linos, and Wafa Khorshid. VIFOR: a tool for software maintenance. Software— Practice and Experience, 20(1):67–77, January 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rowe:1987:BDG

[RDM^{+}87] Lawrence A. Rowe, Michael Davis, Eli Messinger, Carl Meyer, Charles Spirakis, and Allen Tuan. A browser for directed graphs. Software—Practice and Experience, 17(1):61–76, January 1987. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Rehman:2014:UMM

[RdOTF14] Ahmad Saif Ur Rehman, Augusto Born de Oliveira, Mahesh Tripunitara, and Sebastian Fischmeister. The use of mTags for mandatory security: a case study. Software—Practice and Experience, 44 (12):1511–1523, December 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reavley:1973:TEV

[Rea73] P. A. Reavley. Two examples of virtual records. Software—Practice and Experience, 3(4):333–343, October/December 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rohl:1975:CBS

[REC75] J. S. Rohl, H. D. Ellison, and R. J. Collins. An in-core batching standard FORTRAN compiler for large ICL 1900 machine. Software—Practice and Experience, 5(1):19– 28, January/March 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rechenberg:1979:LES

[Rec79] P. Rechenberg. Letter to the Editor: Syntax analysis. Software—Practice and Experience, 9(12):1059, December 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rees:1971:SIM

[Ree71] M. J. Rees. Some improvements to the MINIMOP multi-access operating system. Software—Practice and Experience, 1(2):175– 188, April/June 1971. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reeves:1973:BRB

[Ree73] C. M. Reeves. Book review: Numerical methods with FORTRAN IV case studies, W. S. Dorn and D. D. McCracken, Wiley, London, 1972. No. of pages:

447. Price: £6.00. Software —Practice and Experience, 3(2):190–191, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reeves:1975:BRB

[Ree75]  C. M. Reeves.  Book

review: Digital Computing and Numerical Methods (With Fortran IV, Watfor, and Watfiv Programming) Brice Carnahan and James O. Wilkes, John Wiley, 1973. No. of pages: 477. Price: £7.00. Software— Practice and Experience, 5 (1):111, January 1975. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reeves:1976:BRB

[Ree76] C. M. Reeves. Book review: Electronic Data Processing (second edition), G. Emery, Pitman, London, 1975. No. of pages: 422. Price: £4.95. Software— Practice and Experience, 6 (3):440, July 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rees:1978:BRB

[Ree78] M. J. Rees. Book review: Illustrating basic, D. Alcock, Cambridge University Press, Cambridge, 1977. No. of pages: 134. Price: £4.95 (hard cover), £1.90 (limp cover). Software—Practice and Experience, 8(6):756–759, November 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rees:1982:BRBa

[Ree82] M. J. Rees. Book review: Writing interactive compilers and interpreters, P. J. Brown, Wiley, Chichester, 1979. Number of pages:

265. Price: £10.75. Software—Practice and Experience, 12(1):100, January 1982. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Rees:1984:BRP

[Ree84a] M. J. Rees. Book review:

Pascal Compiler Validation, B. A. Wichmann and Z. J. Ciechanowicz (eds.), Wiley, Chichester, 1983. No. of pages: 176. Price £9.75 Hardback. Software —Practice and Experience, 14(6):601–602, June 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rees:1984:BRB

[Ree84b] M. J. Rees. Book review:

The Unix Operating System, Kaare Christian, Wiley, Chichester, 1983. ISBN 0-471-89052-9 (paperback). No. of pages: 318. Price: £15.50. Software—Practice and Experience, 14(4): 399, April 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reiter:1972:ROT

[Rei72] Allen Reiter. A resourceoriented time-sharing monitor. Software—Practice and Experience, 2(1):55– 71, January/March 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reiser:1982:LEP

[Rei82] John F. Reiser. Letters to the Editor: A portable LISP complier. Software —Practice and Experience, 12(4):394–396, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reimer:1984:IDP

[Rei84] Manuel Reimer. Implementation of the database programming language Modula/R on the personal computer Lilith. Software— Practice and Experience, 14(10):945–956, October 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reiss:1990:IFE

[Rei90] Steven P. Reiss. Interacting with the FIELD environment. Software—Practice and Experience, 20 (S1):S89–S115, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reich:1999:DIC

[Rei99] Jacqueline R. Reich. Design and implementation of a client-server architecture for taxonomy manager. Software—Practice and Experience, 29(2):143– 166, February 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55000307;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55000307&PLACEBO=IE.
pdf.

Roberts:1981:TMA

[REMC81] Eric S. Roberts, Arthur Evans, Jr., C. Robert Morgan, and Edmund M. Clarke. Task management in Ada — a critical evaluation for real-time multiprocessors. Software— Practice and Experience, 11(10):1019–1051, October 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ren:2023:PSA

[Ren23]  Xiquan Ren.  Python

shared atomic data types. Special focus section: Benchmarking, Experimentation Tools, and Reproducible Practices for Data-Intensive Systems from Edge to Cloud: Software: Practice and Experience, 53(12):2393–2407, December 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Revesz:1985:NMG

[Rév85] György E. Révész. A note on macro generation. Software—Practice and Experience, 15(5):423–426, May 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reynolds:1987:UCL

[Rey87] C. F. Reynolds. The use of colour in language syntax analysis. Software—Practice and Experience, 17(8): 513–519, August 1987. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reynolds:1990:SRM

[Rey90] Charles W. Reynolds. Signalling regions: Multiprocessing in a shared memory reconsidered. Software —Practice and Experience, 20(4):325–356, April 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Richardson:1989:IOE

[RG89] Stephen Richardson and Mahadevan Ganapathi. Interprocedural optimization: Experimental results. Software—Practice and Experience, 19(2):149– 169, February 1989. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rodriguez:2021:MBA

[RGC^{+}21] Cristo Rodríguez, Kelly Garcés, Jordi Cabot, Rubby Casallas, Fabián Melo, Daniel Escobar, and Alejandro Salamanca. Modelbased assisted migration of oracle forms applications: the overall process in an industrial setting. Software —Practice and Experience, 51(8):1641–1675, August 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rodgers:1999:TSN

[RGK99] Jeremy B. Rodgers, Rhonda Kay Gaede, and Jeffrey H. Kulick. IN-Tune: an In-Situ non-invasive performance tuning tool for multi-threaded Linux on symmetric multiprocessing Pentium workstations. Software—Practice and Experience, 29(9):775–792, July 25, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=62501865;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=62501865&PLACEBO=IE.
pdf.

Ressia:2014:TED

[RGN^{+}14] Jorge Ressia, Tudor Gîrba, Oscar Nierstrasz, Fabrizio Perin, and Lukas Renggli. Talents: an environment for dynamically composing units of reuse. Software —Practice and Experience, 44(4):413–432, April 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Special issue on International Workshop on Smalltalk Technologies 2011.

Robert:2020:PHM

[RGS^{+}20a] Andrew Robert, Apaar Gupta, Vinayak Shenoy, Dinkar Sitaram, and Subramaniam Kalambur. Predicting Hadoop misconfigurations using machine learning. Software—Practice and Experience, 50(7): 1168–1183, July 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rosendo:2020:AAD

[RGS^{+}20b] Daniel Rosendo, Demis Gomes, Guto Leoni Santos, Leylane Silva, Andre Moreira, Judith Kelner, Djamel Sadok, Glauco Gonçalves, Amardeep Mehta, Mattias Wildeman, and Patricia Takako Endo. Availability analysis of design configurations to compose virtual performanceoptimized data center systems in next-generation cloud data centers. Software—Practice and Experience, 50(6):805–826, June 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rivero-Garcia:2021:UBF

[RGSGHGCG21] Alexandra Rivero-García, Iván Santos-González, Candelaria Hernández-Goya, and Pino Caballero-Gil. Using blockchain in the follow-up of emergency situations related to events. Software—Practice and Experience, 51(10):2000– 2014, October 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Romero:2014:SCR

[RGV14] Julio Cano Romero and Marisol García-Valls. Scheduling component replacement for timely execution in dynamic systems. Software—Practice and Experience, 44(8):889–910, August 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Richardson:1977:DIN

[RH77] M. G. Richardson and S. J.

Hague. The design and implementation of the NAG master library file system. Software—Practice and Experience, 7(1):127– 137, January/March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reghbati:1978:NSM

[RH78] H. K. Reghbati and V. C.

Hamacher. A note on the suitability of Modula for process control applications. Software—Practice and Experience, 8(2):233– 234, March 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Romero:2013:DSO

[RHT^{+}13] Daniel Romero, Gabriel Hermosillo, Amirhosein Taherkordi, Russel Nzekwa, Romain Rouvoy, and Frank Eliassen. The DigiHome Service-Oriented Platform. Software—Practice and Experience, 43(10):1205– 1218, October 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Richards:1971:PBC

[Ric71] M. Richards. The portability of the BCPL compiler. Software—Practice and Experience, 1(2):135– 146, April/June 1971. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Richards:1976:JDP

[Ric76] M. Richards. JUMBO: a demonstration program to illustrate the use of BCPL in a real-time graphics application requiring scaled arithmetic. Software— Practice and Experience, 6(2):255–259, April/June 1976. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Richards:1979:CFR

[Ric79] Martin Richards. A compact function for regular expression pattern matching. Software—Practice and Experience, 9(7):527– 534, July 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Richter:2000:IYA

[Ric00] Mathias W. Richter. Iava:

yet another interpreter for scripting within the Java platform. Software—Practice and Experience, 30(2):81–106, February 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
69503461/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=69503461&PLACEBO=IE.
pdf.

Ringland:1984:SED

[Rin84] Gill Ringland. Software engineering in a development group. Software—Practice and Experience, 14(6):533– 559, June 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rinaldi:1992:BCB

[Rin92] Damian Rinaldi. Balancing the cost/benefit equation. Software Magazine, 12(1):79–??, January 1992. CODEN SMWMEQ. ISSN 0897-8085.

Rintala:2007:ERP

[Rin07]  Matti Rintala.  Excep-

tions in remote procedure calls using C++ template metaprogramming. Software—Practice and Experience, 37(3):231–246, March 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ristov:2005:LTD

[Ris05] Strahil Ristov. LZ trie and dictionary compression. Software—Practice and Experience, 35(5):445– 465, April 25, 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rajapakse:2009:TGR

[RJ09] Damith C. Rajapakse and Stan Jarzabek. Towards generic representation of web applications: solutions and trade-offs. Software —Practice and Experience, 39(5):501–530, April 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rozman:2006:QQA

[RJGH06] Ivan Rozman, Matjaz B.

Juric, Izidor Golob, and Marjan Hericko. Qualitative and quantitative analysis and comparison of Java distributed architectures. Software—Practice and Experience, 36(14): 1543–1562, November 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rong:2020:DEC

[RJZ^{+}20] Guoping Rong, Zefeng Jin, He Zhang, Youwen Zhang, Wenhua Ye, and Dong Shao. DevDocOps: Enabling continuous documentation in alignment with DevOps. Software —Practice and Experience, 50(3):210–226, March 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ramachandran:1989:MBS

[RK89] Umakishore Ramachandran and M. Yousef Amin Khalidi. A measurementbased study of hardware support for object invocation. Software—Practice and Experience, 19(9):809– 828, September 1989. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ramachandran:1991:IDS

[RK91] Umakishore Ramachandran and Yousef Y. A. Khalidi. An implementation of distributed shared memory. Software—Practice and Experience, 21(5):443– 464, May 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rama:2015:SSM

[RK15a] Girish Maskeri Rama and Avinash Kak. Some structural measures of API usability. Software—Practice and Experience, 45(1):75– 110, January 2015. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ristov:2015:FCS

[RK15b] Strahil Ristov and Damir Korenčić. Fast construction of space-optimized recursive automaton. Software—Practice and Experience, 45(6):783–799, June 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Roantree:2014:HAS

[RL14] Mark Roantree and Jun Liu. A heuristic approach to selecting views for materialization. Software—Practice and Experience, 44(10):1157–1179, October 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rouhi:2023:TPB

[RL23] Alireza Rouhi and Kevin Lano. Towards a patternbased model transformation framework. Software—Practice and Experience, 53(9):1815–1849, September 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rosik:2011:AAD

[RLB^{+}11] Jacek Rosik, Andrew Le Gear, Jim Buckley, Muhammad Ali Babar, and Dave Connolly. Assessing architectural drift in commercial software development: a case study. Software—Practice and Experience, 41(1):63–86, January 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reshetova:2018:TLK

[RLPA18] Elena Reshetova, Hans Liljestrand, Andrew Paverd, and N. Asokan. Toward Linux kernel memory safety. Software—Practice and Experience, 48 (12):2237–2256, December 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Radue:1975:SSP

[RM75] J. E. Radue and J. M. Mullins. Solving synchronization problems using semaphores. Software—Practice and Experience, 5(1):51–64, January/ March 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ramsey:1991:LPT

[RM91] Norman Ramsey and Carla Marceau. Literate programming on a team project. Software—Practice and Experience, 21(7): 677–683, July 1991. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rutten:1997:ERD

[RMC97] E. P. Rutten, E. Marchand, and F. Chaumette. An experiment with reactive data-flow tasking in active robot vision. Software—Practice and Experience, 27(5):599–621, May 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7304;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7304&
PLACEBO=IE.pdf.

Russo:2012:RSO

[RMdL12] Alessandro Russo, Massimo Mecella, and Massimiliano de Leoni. ROME4EU — a service-oriented processaware information system for mobile devices. Software—Practice and Experience, 42(10):1275–1314, October 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Riganelli:2019:SCT

[RMM19] Oliviero Riganelli, Daniela Micucci, and Leonardo Mariani. From source code to test cases: a comprehensive benchmark for resource leak detection in Android apps. Software—Practice and Experience, 49(3):540–548, March 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Risco-Martin:2023:XTI

[RMMH^{+}23] José L. Risco-Martín, Saurabh Mittal, Kevin Henares, Román Cardenas, and Patricia Arroba. xDEVS: a toolkit for interoperable modeling and simulation of formal discrete event systems. Software—Practice and Experience, 53(3):748–789, March 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ruiz-Martinez:2014:SLM

[RMMLSME14] A. Ruiz-Martínez, C. Inmaculada Marín-López, D. Sánchez-Martínez, and I. Castell Egea. SIPmsign: a lightweight mobile signature service based on the Session Initiation Protocol. Software—Practice and Experience, 44(5):511– 535, May 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ruiz-Martinez:2011:ACV

[RMSMML^{+}11] Antonio Ruiz-Martínez, Daniel Sánchez-Martínez, C. Inmaculada Marín-López, Manuel Gil-Pérez, and Antonio F. Gómez-Skarmeta. An advanced certificate validation service and architecture based on XKMS. Software —Practice and Experience, 41(3):209–236, March 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rodriguez:2017:ERI

[RMZ17] Ana Rodriguez, Cristian Mateos, and Alejandro Zunino. Experience reports: Improving scientific application execution on Android mobile devices via code refactorings. Software —Practice and Experience, 47(5):763–796, May 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rine:2000:TES

[RN00] David C. Rine and Nader Nada. Three empirical studies of a software reuse reference model. Software—Practice and Experience, 30(6):685–722, May 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
71007586/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=71007586&PLACEBO=IE.
pdf.

Richard:2016:IJS

[RNS^{+}16] Adam Richard, Lai Nguyen, Peter Shipton, Kenneth B. Kent, Azden Bierbrauer, Konstantin Nasartschuk, and Marcel Dombrowski. Inter-JVM sharing. Software—Practice and Experience, 46(9):1285–1296, September 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rees:1977:SIB

[RO77] M. J. Rees and A. W.

Oppenheimer. SOBS — an incremental BASIC system. Software—Practice and Experience, 7(5):631– 643, September/October 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Roberts:1972:BRB

[Rob72] K. V. Roberts. Book review: Software engineering, proceedings of the third symposium on computer and information sciences held in Miami Beach, Florida, December 1969, Julius T. Tou (Ed.), Academic Press, New York and London, 1971. No. of pages: Vol. 1, 272; Vol. 2, 321. Price: Vol. 1, $12.00; Vol. 2, $13.00. Software—Practice and Experience, 2(1): 97–98, January 1972. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robertson:1979:VPA

[Rob79] D. A. Robertson. Variable precision arithmetic on CDC 6000/7000 machines. Software—Practice and Experience, 9(3):247– 248, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robson:1981:BRB

[Rob81] D. J. Robson. Book review: Understanding and writing compilers. Richard Bornat, Macmillan, London. No of pages: 396. Price: £12.00 (hard back), £5.95 (paper back). Software—Practice and Experience, 11(8): 880, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robson:1982:BRBb

[Rob82a] D. J. Robson. Book review: Pascal implementation: The P4 compiler, S. Pemberton and M. C. Danies. Ellis Horwood, Chichester. No. of pages: 254. Price: £22.50. Software—Practice and Experience, 12(12):1167, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robson:1982:BRBa

[Rob82b] D. J. Robson. Book review:

The design and construction of compilers, Robin Hunter, Wiley, Chichester, 1981. No. of pages: 272. Price: £10.50. Software —Practice and Experience, 12(5):494–495, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robson:1983:ETC

[Rob83a] D. J. Robson. An evaluation of throw-away compiling. Software—Practice and Experience, 13(3):241– 249, March 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robson:1983:TCL

[Rob83b] D. J. Robson. Towards a conversational languagesensitive system for Pascal. Software—Practice and Experience, 13 (11):1013–1017, November 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robinson:1984:SPC

[Rob84] John T. Robinson. Separating policy from correctness in concurrency control design. Software—Practice and Experience, 14(9): 827–844, September 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ruano-Ordas:2016:RTS

[ROFGFR^{+}16] David Ruano-Ordás, Jorge Fdez-Glez, Florentino Fdez-Riverola, Vitor Basto Fernandes, and José Ramón Méndez. RuleSIM: a toolkit for simulating the operation and improving throughput of rule-based spam filters. Software— Practice and Experience, 46(8):1091–1108, August 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ruano-Ordas:2016:UNS

[ROFGFRM16] David Ruano-Ordás, Jorge Fdez-Glez, Florentino Fdez-Riverola, and José Ramón Méndez. Using new scheduling heuristics based on resource consumption information for increasing throughput on rulebased spam filtering systems. Software—Practice and Experience, 46(8): 1035–1051, August 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rogers:1971:BRB

[Rog71] M. H. Rogers. Book review: Time sharing system design concepts, Richard W. Watson, McGraw-Hill, New York, 1970. No. of pages: 270. Price: £4.80. Software—Practice and Experience, 1(4):411– 412, October/December 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rogers:1973:BRB

[Rog73] M. H. Rogers. Book review: Computer networks, Infotech State of the art report no. 6, Infotech, Maidenhead, 1971. No. of pages, 623: Price: £40. Software —Practice and Experience, 3(2):188–189, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rogers:1974:BRB

[Rog74] M. H. Rogers. Book review: Interactive computing in Basic, P. C. Sanderson, Butterworths, London, 1973. No. of pages:

161. Price: £4.00. Software—Practice and Experience, 4(4):407, October/ December 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rogers:2023:TPW

[Rog23] David M. Rogers. Three practical workflow schedulers for easy maximum parallelism. Software —Practice and Experience, 53(1):99–114, January 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rohl:1977:BRB

[Roh77a] J. S. Rohl. Book review: Annual review in automatic programming, volume 7. edited by L. Bolliet, A. P. Ershov, M. I. Halpern, J. G. Laski and W. C. McGee, Pergamon, Oxford, 1975. no. of pages: 307. price: £15.00. Software— Practice and Experience, 7 (1):151–152, January 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rohl:1977:CCR

[Roh77b] J. S. Rohl. Converting a class of recursive procedures into non-recursive ones. Software—Practice and Experience, 7(2):231– 238, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rohl:1981:ERC

[Roh81] J. S. Rohl. Eliminating recursion from combinatoric procedures. Software—Practice and Experience, 11(8):803–817, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rondogiannis:1999:AMP

[Ron99] P. Rondogiannis. Adding multidimensionality to procedural programming languages. Software—Practice and Experience, 29 (13):1201–1221, November 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
66501672/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=66501672&PLACEBO=IE.
pdf.

Ronnblom:2007:HEA

[Rön07] Johan Rönnblom. Higherror approximate dictionary search using estimate hash comparisons. Software—Practice and Experience, 37(10):1047–1059, August 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Roper:1988:BRBb

[Rop88a] Marc Roper. Book review: Software — requirements, specification and testing, Tom Anderson (ed.), Blackwell Scientific Publications, 1985. ISBN 0-632-01309-5. No. of pages: 149. Price: £13.50. Software—Practice and Experience, 18(3):285– 286, March 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Roper:1988:BRBa

[Rop88b] Marc Roper. Book review: What is software engineering? G. L. Simons, NCC Publications, 1987. ISBN 0-85012-612-6. No. of pages: 65. Price: £4.50. Software—Practice and Experience, 18(2):166, February 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ross:1971:EGE

[Ros71] Douglas T. Ross. Editorial: Guest editorial. It’s time to ask ‘why?’. Software —Practice and Experience, 1(2):103–104, April/June 1971. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Ross:1974:BRB

[Ros74] D. T. Ross. Book review: Systematic programming: An introduction, N. Wirth, Prentice-Hall, Englewood Cliffs, N.J., 1973. No. of pages: 167. Price: £5.25. Software—Practice and Experience, 4(1):112, January/March 1974. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rosin:1975:MR

[Ros75] Robert F. Rosin. Mythology revisited. Software —Practice and Experience, 5(2):215–216, April 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rosin:1977:GND

[Ros77] Robert F. Rosin. A graphical notation for describing system implementation. Software—Practice and Experience, 7(2):239– 250, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ramasamy:2008:EBI

[RPCS08] HariGovind V. Ramasamy, Prashant Pandey, Michel Cukier, and William H. Sanders. Experiences with building an intrusiontolerant group communication system. Software —Practice and Experience, 38(6):639–666, May ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rossi:2007:JJL

[RPP07] G. Rossi, E. Panegai, and E. Poleo. JSetL: a Java library for supporting declarative programming in Java. Software—Practice and Experience, 37(2):115– 149, February 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rui:2020:SEB

[RQL^{+}20] Lanlan Rui, Yabin Qin, Biyao Li, Ying Wang, and Haoqiu Huang. SEWMS: an edge-based smart wearable maintenance system in communication network. Software—Practice and Experience, 50(5):611– 629, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rakity:1982:LE

[RR82] Philip Rakity and Peter Radford. Letter to the editor. Software—Practice and Experience, 12(3): 293, March 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rees:1985:VSP

[RR85] Michael J. Rees and David J. Robson. Video-Slide: a presentation aid for the BBC microcomputer. Software—Practice and Experience, 15(9):915– 923, September 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ramos:2005:EIM

[RR05] Jorge R. Ramos and Vernon Rego. Efficient implementation of multiprocessor scheduling algorithms on a simulation testbed. Software—Practice and Experience, 35(1): 27–50, January 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Raj:2022:SGB

[RR22] Vinay Raj and Sadam Ravichandra. A service graph based extraction of microservices from monolith services of serviceoriented architecture. Software—Practice and Experience, 52(7):1661–1678, July 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Raith:2023:FST

[RRFD23] Philipp Raith, Thomas Rausch, Alireza Furutanpey, and Schahram Dustdar. faas-sim: a tracedriven simulation framework for serverless edge computing platforms. Special focus section: Benchmarking, Experimentation Tools, and Reproducible Practices for Data-Intensive Systems from Edge to Cloud: Software: Practice and Experience, 53(12):2327–2361, December 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rahman:2018:MRA

[RRK^{+}18] Md. Masudur Rahman, Rashed Rubby Riyadh, Shah Mostafa Khaled, Abdus Satter, and Md. Rayhanur Rahman. MM-RUC3: a recommendation approach of move method refactoring using coupling, cohesion, and contextual similarity to enhance software design. Software—Practice and Experience, 48 (9):1560–1587, September 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Russo:1995:OSI

[RRP95] Guido Russo, Stefano Russo, and Benôit Pirenne.

An operating system independent WORM archival system. Software—Practice and Experience, 25(5):521– 532, May 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rodriguez-Rivera:1997:NCG

[RRR97] Gustavo Rodriguez-Rivera and Vincent F. Russo. Nonintrusive cloning garbage collection with stock operating system support. Software—Practice and Experience, 27(8):885–904, August 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7324;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7324&
PLACEBO=IE.pdf.

Russell:1976:IPC

[RS76] David L. Russell and Jeffrey V. Sue. Implementation of a Pascal compiler for the IBM 360. Software—Practice and Experience, 6(3):371–376, July/ September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rees:1982:KEO

[RS82] D. J. Rees and P. D.

Stephens. The kernel of the EMAS 2900 operating system. Software—Practice and Experience, 12(7):655– 667, July 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ramanathan:1986:TDF

[RS86] Jay Ramanathan and Charley Shubra. Template design for file processing applications. Software— Practice and Experience, 16(2):173–182, February 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Roper:1987:STM

[RS87] M. Roper and P. Smith. A structural testing method for JSP designed programs. Software—Practice and Experience, 17(2):135– 157, February 1987. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reyes:1990:IPS

[RS90] Tom C. Reyes and James L. Sills. Implementation of a portable screen manager. Software—Practice and Experience, 20(10): 1007–1014, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Read:1991:LWU

[RS91] Robert L. Read and Martin L. Smith. A lightweight UIMS. Software —Practice and Experience, 21(1):13–33, January 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rafea:1993:LAI

[RS93a] Ahmed A. Rafea and Khaled F. Shaalan. Lexical analysis of inflected Arabic words using exhaustive search of an augmented transition network. Software—Practice and Experience, 23(6):567–588, June 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robillard:1993:ICG

[RS93b] Pierre N. Robillard and Mario Simoneau. Iconic control graph representation. Software—Practice and Experience, 23(2):223– 234, February 1993. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Raju:1994:PES

[RS94] Sitaram C. V. Raju and Alan C. Shaw. A prototyping environment for specifying, executing and checking communicating realtime state machines. Software—Practice and Experience, 24(2):175–195, February 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Romig:1995:DDE

[RS95] Phillip R. Romig III and Ashok Samal. DeViouS: a distributed environment for computer vision. Software—Practice and Experience, 25(1):23–45, January 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Roy:2021:BBC

[RS21] Deepsubhra Guha Roy and Satish Narayana Srirama. A blockchainbased cyber attack detection scheme for decentralized Internet of Things using software-defined network. Software—Practice and Experience, 51(7): 1540–1556, July 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rodriguez-Silva:2016:IVR

[RSLAGCLB16] Daniel A. Rodríguez-Silva, Jaime Loureiro-Acuña, Francisco J. González-Castaño, and Cristina López-Bravo. Improving the virtualization of rich applications by combining VNC and streaming protocols at the hypervisor layer. Software—Practice and Experience, 46(5):625– 644, May 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rodriguez-Silva:2015:SAV

[RSRCGC15] Daniel A. Rodríguez-Silva, Iria Rodríguez-Cruz, and Francisco J. González-Castaño. A software architecture for virtualized educational multimedia games in IPTV environments. Software—Practice and Experience, 45(2):143– 160, February 2015. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ritschel:2023:TIE

[RSW^{+}23] Nico Ritschel, Anand Ashok Sawant, David Weintrop, Reid Holmes, Alberto Bacchelli, Ronald Garcia, Chandrika K R, Avijit Mandal, Patrick Francis, and David C. Shepherd. Training industrial enduser programmers with interactive tutorials. Software—Practice and Experience, 53(3):729–747, March 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robinson:1977:DAP

[RT77] S. K. Robinson and I. S.

Torsun. Dynamic analysis of program performance (DAP) in a Fortran batch environment. Software— Practice and Experience, 7(3):307–315, May/June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ringland:1978:PIR

[RT78] Gill Ringland and A. R. Trice. Pilot implementations of reliable systems. Software—Practice and Experience, 8(3):323– 339, May/June 1978. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rozin:1991:HIP

[RT91] Roman Rozin and Siegfried Treu. A hybrid implementation of a processoriented programming language for system simulation. Software—Practice and Experience, 21(6):557– 579, June 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Romei:2010:XDM

[RT10] Andrea Romei and Franco Turini. XML data mining. Software—Practice and Experience, 40(2):101– 130, February ??, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Raj:1991:EGP

[RTL^{+}91] Rajendra K. Raj, Ewan Tempero, Henry M. Levy, Andrew P. Black, Norman C. Hutchinson, and Eric Jul. Emerald: a general-purpose programming language. Software—Practice and Experience, 21(1):91–118, January 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rueher:1993:FEP

[Rue93] Michel Rueher. A first exploration of Prolog III’s capabilities. Software—Practice and Experience, 23(2): 177–200, February 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Russo:1995:EOO

[Rus95] Vincent F. Russo. Editorial: Object-orientation: Looking back, looking forward. Software—Practice and Experience, 25(S4):1– 2, December 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ranjan:2020:STT

[RVS^{+}20] Rajiv Ranjan, Massimo Villari, Haiying Shen, Omer Rana, and Rajkumar Buyya. Software tools and techniques for fog and edge computing. Software —Practice and Experience, 50(5):473–475, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rose:1981:FPL

[RW81] G. A. Rose and J. Welsh. Formatted programming languages. Software— Practice and Experience, 11(7):649–670, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rousskov:2004:HPB

[RW04] Alex Rousskov and Duane Wessels. High-performance benchmarking with Web Polygraph. Software— Practice and Experience, 34(2):187–211, February 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ryder:2012:LDA

[RW12] Barbara G. Ryder and Ben Wiedermann. Language design and analyzability: a retrospective. Software —Practice and Experience, 42(1):3–18, January 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rudafshani:2017:LDD

[RW17] Masoomeh Rudafshani and Paul A. S. Ward. LeakSpot: detection and diagnosis of memory leaks in JavaScript applications. Software —Practice and Experience, 47(1):97–123, January 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ranjan:2017:ESI

[RWJ^{+}17] Rajiv Ranjan, Lizhe Wang, Prem Prakash Jayaraman, Karan Mitra, and Dimitrios Georgakopoulos. Editorial: Special issue on Big Data and Cloud of Things (CoT). Software—Practice and Experience, 47(3):345– 347, March 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ryan:1980:LEJ

[Rya80] Kevin T. Ryan. Letters to the Editor: Jackson and sequence checking. Software —Practice and Experience, 10(10):849, October 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ryder:1974:PV

[Ryd74] Barbara G. Ryder. The PFORT verifier. Software—Practice and Experience, 4(4):359–377, October/December 1974. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ryu:2016:SFP

[Ryu16]  Sukyoung Ryu.  Scal-

able framework for parsing: from Fortress to JavaScript. Software— Practice and Experience, 46(9):1219–1238, September 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rouhi:2017:MBF

[RZ17] Alireza Rouhi and Bahman Zamani. A modelbased framework for automatic generation of a pattern language verifier. Software—Practice and Experience, 47(12):1945–1980, December 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sosic:1997:GRD

[SA97] Rok Sosič and David Abramson. Guard: a relative debugger. Software—Practice and Experience, 27(2):185–206, February 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7287;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7287&
PLACEBO=IE.pdf.

Stoecklin:2002:CRG

[SA02] Sara Stoecklin and Clement Allen. Creating a reusable GUI component. Software—Practice and Experience, 32(5):403–416, April 25, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
93513334/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=93513334&PLACEBO=IE.
pdf.

Singh:2020:LBA

[SA20] Anil Singh and Nitin Auluck. Load balancing aware scheduling algorithms for fog networks. Software—Practice and Experience, 50 (11):2012–2030, November 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Srivastava:2020:DCP

[SAA^{+}20] Aishwarya Srivastava, Siddhant Aggarwal, Amy Apon, Edward Duffy, Ken Kennedy, Andre Luckow, Brandon Posey, and Marcin Ziolkowski. Deployment of a cloud pipeline for real-time visual inspection using fast streaming high-definition images. Software—Practice and Experience, 50(6):868– 898, June 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sabin:1976:PSE

[Sab76] M. A. Sabin. Portability— some experiences with FORTRAN. Software —Practice and Experience, 6(3):393–396, July/ September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stevenson:1992:VCF

[SAC^{+}92] D. E. Stevenson, L. K. Ammons, W. G. Crosmun, A. Jackson, and G. L. Raj. A vector C and Fortran compiler for the FPS Tseries: Experiences with compiling to occam I. Software—Practice and Experience, 22(5):371–390, May 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Santhanakrishnan:2006:STC

[SAC06]  Ganesh  Santhanakrish-

nan, Ahmed Amer, and Panos K. Chrysanthis. Self-tuning caching: the Universal Caching algorithm. Software—Practice and Experience, 36(11– 12):1179–1188, September/ October 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Such:2011:GOS

[SAEGF11] Jose M. Such, Juan M. Alberola, Agustin Espinosa, and Ana Garcia-Fornes. A group-oriented secure multiagent platform. Software—Practice and Experience, 41(11):1289–1302, October 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shoyari:2021:AMR

[SAEMM21] Mahsa Faraji Shoyari, Ehsan Ataie, Reza Entezari-Maleki, and Ali Movaghar. Availability modeling in redundant OpenStack private clouds. Software— Practice and Experience, 51(6):1218–1241, June 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sale:1979:SSA

[Sal79a] A. H. J. Sale. Strings and the sequence abstraction in Pascal. Software—Practice and Experience, 9(8):671– 683, August 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sale:1979:ISP

[Sal79b]  Arthur Sale.  Imple-

menting strings in Pascal-Again. Software—Practice and Experience, 9(10):839– 841, October 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sale:1979:MM

[Sal79c] Arthur Sale. Miniscules and majuscules. Software —Practice and Experience, 9(11):915–919, November 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sale:1979:PSR

[Sal79d] Arthur Sale. Pascal stylistics and reserved words. Software—Practice and Experience, 9(10):821–825, October 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sale:1981:FDP

[Sal81a] A. H. J. Sale. Forwarddeclared procedures, parameterlists and scope. Software —Practice and Experience, 11(2):123–130, February 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sale:1981:ICS

[Sal81b] Arthur Sale. The implementation of case statements in Pascal. Software —Practice and Experience, 11(9):929–942, September 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sherwani:2004:LCE

[SAL^{+}04] Jahanzeb Sherwani, Nosheen Ali, Nausheen Lotia, Zahra Hayat, and Rajkumar Buyya. Libra: a computational economy-based job scheduling system for clusters. Software—Practice and Experience, 34(6):573– 590, May 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sebastio:2016:ERA

[SAL16] Stefano Sebastio, Michele Amoretti, and Alberto Lluch Lafuente. Experience report: AVOCLOUDY: a simulator of volunteer clouds. Software—Practice and Experience, 46(1):3– 30, January 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Samet:1971:NRC

[Sam71a] P. A. Samet. A note on radix conversion for integers. Software—Practice and Experience, 1(1):93– 96, January/March 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Samet:1971:SCC

[Sam71b] Paul A. Samet. Short communications and comment: Letters to the editor. Software—Practice and Experience, 1(4):411, October/ December 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Samet:1975:AVO

[Sam75] P. A. Samet. APL versus operator precedence. Software—Practice and Experience, 5(3):313, July 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Samet:1981:ESC

[Sam81] Hanan Samet. Experience with software conversion. Software—Practice and Experience, 11(10): 1053–1069, October 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shaw:1981:CPL

[SAN^{+}81] Mary Shaw, Guy T. Almes, Joseph M. Newcomer, Brian K. Reid, and Wm. A. Wulf. A comparison of programming languages for software engineering. Software—Practice and Experience, 11(1):1–52, January 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sandberg:1988:EOO

[San88] David W. Sandberg. Experience with an objectoriented virtual machine. Software—Practice and Experience, 18(5):415–425, May 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Santos:2017:VMP

[San17] André L. Santos. Variability management of pluginbased systems using feature models. Software—Practice and Experience, 47(7): 959–970, July 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sarwate:1977:EBF

[Sar77]  D. V. Sarwate.  Eval-

uation of boolean functions. Software—Practice and Experience, 7(3):426– 427, June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sassa:1979:PMM

[Sas79] Masataka Sassa. A pattern matching macro processor. Software—Practice and Experience, 9(6):439– 456, June 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Satterthwaite:1972:DTH

[Sat72] E. Satterthwaite. Debugging tools for high level languages. Software—Practice and Experience, 2(3):197– 217, July/September 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schuts:2022:LSS

[SATV22] Mathijs T. W. Schuts, Rodin T. A. Aarssen, Paul M. Tielemans, and Jurgen J. Vinju. Largescale semi-automated migration of legacy C/C++ test code. Software—Practice and Experience, 52(7): 1543–1580, July 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Saunders:1988:AGB

[Sau88] D. Saunders. An addendum to G. S. Blair, J. A. Mariani and W. D. Shepherd: ‘Practical Extension to UNIX for Interprocess Communication’. Software—Practice and Experience, 18(12):1191–1192, December 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Savidis:2004:IGS

[Sav04]  Anthony Savidis.  The

implementation of generic smart pointers for advanced defensive programming. Software—Practice and Experience, 34(10): 977–1009, August 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Savidis:2006:AID

[Sav06] Anthony Savidis. Application invariants: Design by Contract augmented with deployment correctness logic. Software—Practice and Experience, 36(3): 255–282, March 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Savidis:2007:RIL

[Sav07] Anthony Savidis. Rapidly implementing languages to compile as C++ without crafting a compiler. Software—Practice and Experience, 37(15):1577–1620, December ??, 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Savidis:2011:IID

[Sav11] Anthony Savidis. Integrated implementation of dynamic untyped objectbased operator overloading. Software—Practice and Experience, 41(11): 1155–1184, October 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sherwood:2016:FES

[SAY16] Thomas Sherwood, Ezak Ahmad, and Moi Hoon Yap. Formulating efficient software solution for digital image processing system. Software—Practice and Experience, 46(7):931– 954, July 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sartaj:2024:MBD

[SAYM24] Hassan Sartaj, Shaukat Ali, Tao Yue, and Kjetil Moberg. Model-based digital twins of medicine dispensers for healthcare IoT applications. Software—Practice and Experience, 54(6):1172–1192, June 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Smith:1982:MSM

[SB82] A. Smith and D. H. Bell. A multiprogramming system for a microcomputer. Software—Practice and Experience, 12(5):405–410, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sachs:1983:SIP

[SB83] J. M. Sachs and S. K. Burns. STOIC, an interactive programming system for dedicated computing. Software—Practice and Experience, 13(1):1– 16, January 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schaefer:1993:SAE

[SB93] Carl F. Schaefer and Gary N. Bundy. Static analysis of exception handling in Ada. Software —Practice and Experience, 23(10):1157–1174, October 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sweeney:2003:QES

[SB03] Peter F. Sweeney and Michael Burke. Quantifying and evaluating the space overhead for alternative C++ memory layouts. Software—Practice and Experience, 33(7):595– 636, June 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Simpson:2013:MES

[SB13] Matthew S. Simpson and Rajeev K. Barua. Mem-Safe: ensuring the spatial and temporal memory safety of C at runtime. Software—Practice and Experience, 43(1):93– 128, January 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sandhu:2021:SRA

[SB21] Amandeep Kaur Sandhu and Ranbir Singh Batth. Software reuse analytics using integrated random forest and gradient boosting machine learning algorithm. Software—Practice and Experience, 51(4):735– 747, April 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Srirama:2022:PGJ

[SB22] Satish Narayana Srirama and Rajkumar Buyya. Post golden jubilee year of the software journal: New research trends and strengthening advisory editorial team. Software—Practice and Experience, 52(1):3– 4, January 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shahpar:2022:EEA

[SBB22] Zahra Shahpar, Vahid Khatibi Bardsiri, and Amid Khatibi Bardsiri. An evolutionary ensemble analogybased software effort estimation. Software—Practice and Experience, 52(4): 929–946, April 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sharma:2007:OMI

[SBcC07] Srikant Sharma, Inho Baek, and Tzi cker Chiueh. OmniCon: a Mobile IP-based vertical handoff system for wireless LAN and GPRS links. Software —Practice and Experience, 37(7):779–798, June 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schneider:2015:SPS

[SBD15] Chris Schneider, Adam Barker, and Simon Dobson. Survey paper: A survey of self-healing systems frameworks. Software—Practice and Experience, 45(10):1375–1398, October 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sousa:2019:ESC

[SBF19] Bruno L. Sousa, Mariza A. S. Bigonha, and Kecia A. M. Ferreira. An exploratory study on cooccurrence of design patterns and bad smells using software metrics. Software—Practice and Experience, 49(7):1079–1113, July 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sousa:2024:EID

[SBFF24] Bruno L. Sousa, Mariza A. S. Bigonha, Kecia A. M. Ferreira, and Glaura C. Franco. Evolution of internal dimensions in object-oriented software — a time series based approach. Software—Practice and Experience, 54(6): 1034–1073, June 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shalaby:2005:BER

[SBG^{+}05] Nadia Shalaby, Andy Bavier, Yitzchak Gottlieb, Scott Karlin, Larry Peterson, Xiaohu Qie, Tammo Spalink, and Mike Wawrzoniak. Building extensible routers using network processors. Software—Practice and Experience, 35(12): 1155–1194, October 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Soares:2006:DPA

[SBL06] Sérgio Soares, Paulo Borba, and Eduardo Laureano. Distribution and persistence as aspects. Software —Practice and Experience, 36(7):711–759, June 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stanimirovic:2013:MIL

[SBS13] Aleksandar Stanimirović, Miloš Bogdanović, and Leonid Stoimenov. Methodology and intermediate layer for the automatic creation of ontology instances stored in relational databases. Software— Practice and Experience, 43(2):129–152, February 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sajedi-Badashian:2020:VTB

[SBS20] Ali Sajedi-Badashian and Eleni Stroulia. Vocabulary and time based bugassignment: a recommender system for opensource projects. Software —Practice and Experience, 50(8):1539–1564, August 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sudama:1990:EDS

[SC90] Ram Sudama and Dah-Ming Chiu. Experiences of designing a sophisticated network monitor. Software —Practice and Experience, 20(6):555–570, June 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sherrell:1994:ETZ

[SC94] Linda B. Sherrell and Doris L. Carver. Experiences in translating Z designs to Haskell implementations. Software—Practice and Experience, 24 (12):1159–1178, December 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sonntag:2014:ECS

[SC14] Benoît Sonntag and Dominique Colnet. Efficient compilation strategy for object-oriented languages under the closed-world assumption. Software—Practice and Experience, 44(5): 565–592, May 2014. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Somasekaram:2022:PLM

[SC22] Premathas Somasekaram and Radu Calinescu. Predicting locally manageable resource failures of high availability clusters. Software—Practice and Experience, 52(10):2191–2225, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sas:2023:GRG

[SCD^{+}23] Cezar Sas, Andrea Capiluppi, Claudio Di Sipio, Juri Di Rocco, and Davide Di Ruscio. GitRanking: a ranking of GitHub topics for software classification using active sampling. Software—Practice and Experience, 53(10):1982–2006, October 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Salehi:2017:RSR

[SCF^{+}17] Mohsen Amini Salehi, Thomas Caldwell, Alejandro Fernandez, Emmanuel Mickiewicz, Eric W. D. Rozier, Saman Zonouz, and David Redberg. RESeED: a secure regular-expression search tool for storage clouds. Software—Practice and Experience, 47 (9):1221–1241, September 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Song:2009:BNB

[ScG09] Yang Song, Aleksander Koł cz, and C. Lee Giles. Better Naive Bayes classification for high-precision spam detection. Software—Practice and Experience, 39(11):1003–1024, August 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schwabe:1992:HDU

[SCGP92] Daniel Schwabe, Andrea Caloini, Franca Garzotto, and Paolo Paolini. Hypertext development using a model-based approach. Software—Practice and Experience, 22(11): 937–962, November 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schumann:1972:MDB

[Sch72] U. Schumann. MAPLIB:

a data bank of FORTRAN functions describing material properties. Software—Practice and Experience, 2(1):21–28, January/ March 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schneck:1974:MM

[Sch74] Paul B. Schneck. The myth of multiprogramming. Software—Practice and Experience, 4(1):59–62, January/ March 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schneiderman:1976:RDT

[Sch76a]  Ben Schneiderman.  A

review of design techniques for programs and data. Software—Practice and Experience, 6(4):555– 567, October/December 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schonfelder:1976:PSF

[Sch76b] J. L. Schonfelder. The production of special function routines for a multimachine library. Software—Practice and Experience, 6(1):71–82, January/ March 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schwetman:1978:JSM

[Sch78] Herbert D. Schwetman. Job scheduling in multiprogrammed computer systems. Software—Practice and Experience, 8(3):241– 255, May/June 1978. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schach:1980:PTP

[Sch80] Stephen R. Schach. A portable trace for the Pascal heap. Software—Practice and Experience, 10(6): 421–426, June 1980. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schach:1982:UTS

[Sch82] Stephen R. Schach. A unified theory for software production. Software—Practice and Experience, 12(7): 683–689, July 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schneck:1983:MVM

[Sch83a] Paul B. Schneck. The myth of virtual memory. Software—Practice and Experience, 13(6):537–543, June 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schoppers:1983:SCL

[Sch83b] Marcel J. Schoppers. Short communication. logic-programming production systems with metalog. Software—Practice and Experience, 13(9): 871–872, September 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schneck:1986:SSH

[Sch86] Paul B. Schneck. Superlinear speed-up and the halting problem. Software—Practice and Experience, 16(8):781–782, August 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schneider:1989:CPP

[Sch89a] Victor Schneider. Converting a portable Pascal P-Code interpreter to a code generator. Software—Practice and Experience, 19(11):1111–1113, November 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schonfelder:1989:SEP

[Sch89b] J. L. Schonfelder. Semantic extension possibilities in the proposed new Fortran. Software—Practice and Experience, 19(6):529– 551, June 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schaerf:2000:LCF

[SCL00] Andrea Schaerf, Marco Cadoli, and Maurizio Lenzerini. LOCAL++: A C++ framework for local search algorithms. Software—Practice and Experience, 30(3):233–257, March 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
71001775/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=71001775&PLACEBO=IE.
pdf.

Sun:2021:CSO

[SCLD21] Lei Sun, Yanfeng Chen, Wenlong Li, and Qingxu Deng. Charging stationsoriented electric vehicle charging strategy based on battery characteristics. Software—Practice and Experience, 51 (11):2168–2184, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Scowen:1973:BSA

[Sco73] R. S. Scowen. Babel and SOAP, applications for extensible compilers. Software—Practice and Experience, 3(1):15–27, January/ March 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Scowen:1977:DMI

[Sco77a] R. S. Scowen. On detecting misspelt identifiers in Fortran. Software—Practice and Experience, 7(4): 536, July 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Scowen:1977:SAP

[Sco77b] R. S. Scowen. Some aids for program documentation. Software—Practice and Experience, 7(6):779– 792, November/December 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Scowen:1981:SST

[Sco81] R. S. Scowen. A survey of some text editors. Software —Practice and Experience, 11(9):883–906, September 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Silva-Coira:2023:MAR

[SCPL23] Fernando Silva-Coira, José R. Paramá, and Susana Ladra. Map algebra on raster datasets represented by compact data structures. Software—Practice and Experience, 53(6):1362– 1390, June 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sang:1994:STB

[SCR94] Janche Sang, Ke-Hsiung Chung, and Vernon Rego. A simulation testbed based on lightweight processes. Software—Practice and Experience, 24(5):485–505, May 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sorzano:2002:CLI

[SCT02] C. O. S. Sorzano, J. M.

Carazo, and O. Trelles. Command-line interfaces can be efficiently brought to graphics: COLIMATE (the COmmand LIne MATE). Software—Practice and Experience, 32(9):873–887, July 25, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
94519538/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=94519538&PLACEBO=IE.
pdf.

Samet:1975:DAP

[SD75] P. A. Samet and Raymond D. Dunn. Detailed analysis of a program— an instructive horror story. Software—Practice and Experience, 5(2):211–213, April 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sharma:2018:HTR

[SD18] Suchakrapani Datt Sharma and Michel R. Dagenais. Hardware trace reconstruction of runtime compiled code. Software—Practice and Experience, 48 (5):1042–1055, May 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2567.

Safaei:2022:GOD

[SDB^{+}22] Mahmood Safaei, Maha Driss, Wadii Boulila, Elankovan A. Sundararajan, and Mitra Safaei. Global outliers detection in wireless sensor networks: a novel approach integrating time-series analysis, entropy, and random forestbased classification. Software—Practice and Experience, 52(1):277–295, January 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sunderland:2004:FXB

[SDC04] R. M. Sunderland, R. I. Damper, and R. M. Crowder. Flexible XML-based configuration of physical simulations. Software— Practice and Experience, 34(12):1149–1155, October 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Scanniello:2010:ALR

[SDDD10] Giuseppe Scanniello, Anna D’Amico, Carmela D’Amico, and Teodora D’Amico. Architectural layer recovery for software system understanding and evolution. Software—Practice and Experience, 40(10): 897–916, September 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sotgiu:2021:CDA

[SDF^{+}21] Alessandro Sotgiu, Cinzia De Donato, Claudio Fornaro, Sandro Tassa, Marco Scannavini, Dario Iannaccio, Giovanni Ambrosi, Simona Bartocci, Laurent Basara, Roberto Battiston, William J. Burger, Donatella Campana, Luca Carfora, Guido Castellini, Piero Cipollone, Livio Conti, Andrea Contin, Fulvio De Persio, Cristian De Santis, Francesco M. Follega, Cristina Guandalini, Maria Ionica, Roberto Iuppa, Giuliano Laurenti, Ignazio Lazzizzera, Mauro Lolli, Christian Manea, Matteo Martucci, Giuseppe Masciantonio, Matteo Mergé, Giuseppe Osteria, Lorenzo Pacini, Francesco Palma, Federico Palmonari, Beatrice Panico, Alexandra Parmentier, Francesco Perfetto, Piergiorgio Picozza, Mirko Piersanti, Michele Pozzato, Matteo Puel, Irina Rashevskaya, Ester Ricci, Marco Ricci, Sergio Bruno Ricciarini, Valentina Scotti, Roberta Sparvoli, Bruno Spataro, Vincenzo Vitale, Simona Zoffoli, and Paolo Zuccon. Control and data acquisition software of the highenergy particle detector on board the China Seismo-Electromagnetic Satellite space mission. Software—Practice and Experience, 51(6):1459–1480, June 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sauvanaud:2020:BDD

[SDG^{+}20] Carla Sauvanaud, Ajay Dholakia, Jordi Guitart, Chulho Kim, and Peter Mayes. Big data deployment in containerized infrastructures through the interconnection of network namespaces. Software—Practice and Experience, 50(7):1087–1113, July 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sahay:2023:ABP

[SDIP23] Apurvanand Sahay, Davide Di Ruscio, Ludovico Iovino, and Alfonso Pierantonio. Analyzing business process management capabilities of low-code development platforms. Software—Practice and Experience, 53(4):1036–1060, April 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Saar:2016:BCB

[SDKS16] Tõnis Saar, Marlon Dumas, Marti Kaljuve, and Nataliia Semenenko. Browserbite: cross-browser testing via image processing. Software—Practice and Experience, 46(11):1459–1477, November 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Soares:2021:WFI

[SdLJMP21] Daricélio Moreira Soares, Manoel Limeira de Lima Júnior, Leonardo Murta, and Alexandre Plastino. What factors influence the lifetime of pull requests? Software—Practice and Experience, 51(6):1173–1193, June 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Selcuk:2011:RMI

[SE11] Yunus Emre Selçuk and Nadia Erdoğan. Role models — implementationbased approaches to using roles. Software—Practice and Experience, 41(1):1– 22, January 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Seffah:1997:OOF

[Sef97] Ahmed Seffah. An objectoriented framework for modeling and designing intelligent training systems. Software—Practice and Experience, 27(10):1233– 1243, October 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=7267;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=7267&
                                        PLACEBO=IE.pdf.

Self:1975:END

[Sel75] John A. Self. Embedding non-determinism. Software—Practice and Experience, 5(3):221–227, July/ September 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Santos:2023:SAA

[SESK23] Guto Leoni Santos, Patricia Takako Endo, Djamel Sadok, and Judith Kelner. SPIDER: an availabilityaware framework for the service function chain placement in distributed scenarios. Software—Practice and Experience, 53(2): 413–437, February 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Setzer:1979:NRT

[Set79]  V. W. Setzer.  Non-

recursive top-down syntax analysis. Software—Practice and Experience, 9(3): 237–245, March 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sethi:1981:UST

[Set81] Ravi Sethi. Uniform syntax for type expressions and declarators. Software —Practice and Experience, 11(6):623–628, June 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sethi:1984:PEA

[Set84] Ravi Sethi. Preprocessing embedded actions. Software—Practice and Experience, 14(3):291–297, March 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sewell:1982:RLT

[Sew82] B. T. Sewell. A rapid lookup table method for trigonometric functions. Software—Practice and Experience, 12(10):885–888 (or 885–887??), October 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Saxena:1985:PRT

[SF85] S. Saxena and J. A. Field. Portable real-time software for 8-bit microprocessors. Software—Practice and Experience, 15(3):277– 303, March 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shepherd:1988:LEC

[SF88] Roger Shepherd and Charles Farnum. Letter to the Editor: Compiler support for floating-point computation. Software—Practice and Experience, 18 (12):1193–1194, December 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sens:1998:SFM

[SF98] Pierre Sens and Bertil Folliot. The STAR fault manager for distributed operating environments: Design, implementation and performance. Software—Practice and Experience, 28(10):1079–1099, August 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=10007355;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=10007355&PLACEBO=IE.
pdf.

Simons:2013:EGE

[SFB13] Anthony J. H. Simons, Gordon Fraser, and Leonardo Bottaci. Editorial: Guest editor’s introduction to the special section on TAIC-PART 2010 — Testing: Academic and Industrial Conference — Practice and Research Techniques. Software—Practice and Experience, 43(9):1101–1102, September 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Santana:2021:AMU

[SFC^{+}21] Alexandre Santana, Vinicius Freitas, Márcio Castro, Laércio L. Pilla, and Jean-François Méhaut. ARTful: a model for user-defined schedulers targeting multiple high-performance computing runtime systems. Software—Practice and Experience, 51(7):1622– 1638, July 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shimasaki:1980:APP

[SFIK80] Masaaki Shimasaki, Shigeru Fukaya, Katsuo Ikeda, and Takeshi Kiyono. An analysis of Pascal programs in compiler writing. Software —Practice and Experience, 10(2):149–157, February 1980. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Sang:2001:MLS

[SFK^{+}01] Janche Sang, Gregory Follen, Chan Kim, Isaac Lopez, and Scott Townsend. Migrating legacy scientific applications towards CORBA-based client-server architectures. Software— Practice and Experience, 31(14):1313–1330, November 25, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        85512521/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=85512521&PLACEBO=IE.
                                        pdf.

Sreerama:1997:EGO

[SFS97a] Sethu Sreerama, David Fleming, and Murali Sitaraman. Errata: Graceful object-based performance evolution by sethu sreerama, David Fleming and murali sitaraman— vol. 27, no. 1, 1997, pp. 111–122. Software—Practice and Experience, 27(4): 495, April 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sreerama:1997:GOBa

[SFS97b] Sethu Sreerama, David Fleming, and Murali Sitaraman. Graceful objectbased performance evolution. Software—Practice and Experience, 27(1):111– 122, January 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7260;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7260&
PLACEBO=IE.pdf. See com-

ments [SFS97c].

Sreerama:1997:GOBb

[SFS97c] Sethu Sreerama, David Fleming, and Murali Sitaraman. Graceful objectbased performance evolution by Sethu Sreerama, David Fleming and Murali Sitaraman — Vol. 27, No. 1, 1997, pp. 111– 122. Software—Practice and Experience, 27(4):495, April 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7297;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7297&
PLACEBO=IE.pdf.  See

[SFS97b].

Sanders:1979:DM

[SG79] Alton F. Sanders and Herbert L. Gelernter. Data metastructures. Software —Practice and Experience, 9(10):795–811, October 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Scheifler:1990:XWS

[SG90] Robert W. Scheifler and Jim Gettys. The X Window System. Software—Practice and Experience, 20(S2):S5–S34, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Reprinted from TOG 5(2): 79-109 (1986).

Sayre:1993:BGD

[SG93] Kirk Sayre and Michael A. Gray. Backtalk: a generalized dynamic communication system for DAI. Software—Practice and Experience, 23(9):1043–1057, September 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shanbhag:1997:CSG

[SG97] Vivek K. Shanbhag and K. Gopinath. A C++ simulator generator from graphical specifications. Software—Practice and Experience, 27(4):395–423, April 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.

com/cgi-bin/abstract?
ID=7299;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7299&
PLACEBO=IE.pdf.

Shahidinejad:2020:JCO

[SGA20] Ali Shahidinejad and Mostafa Ghobaei-Arani. Joint computation offloading and resource provisioning for edge–cloud computing environment: a machine learning-based approach. Software—Practice and Experience, 50(12):2212– 2230, December 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Said:2021:UAV

[SGA21] Nidal Al Said, Yuri Gorbachev, and Aleksei Avdeenko. An unmanned aerial vehicles navigation system on the basis of pattern recognition applications-Review of implementation options and prospects for development. Software—Practice and Experience, 51(7): 1509–1517, July 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See retraction notice [Ano23m].

Salimian:2021:TAA

[SGAS21] Mahboubeh Salimian, Mostafa Ghobaei-Arani, and Ali Shahidinejad. Toward an autonomic approach for Internet of Things service placement using gray wolf optimization in the fog computing environment. Software—Practice and Experience, 51(8): 1745–1772, August 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

SantAnna:2013:MCA

[SGBR13] Claudio Sant’Anna, Alessandro Garcia, Thais Batista, and Awais Rashid. Mastering crosscutting architectural decisions with aspects. Software—Practice and Experience, 43(3):305– 332, March 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sztajnberg:2011:IES

[SGCM11] Alexandre Sztajnberg, Rodrigo Souza Granja, Jeane Cesário, and André Felipe Almeida Monteiro. An integration experience of a software architecture and a monitoring infrastructure to deploy applications with non-functional requirements in computing grids. Software—Practice and Experience, 41(1):103– 127, January 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Skibinski:2005:RDB

[SGD05] Przemysław Skibiński, Szymon Grabowski, and Sebastian Deorowicz. Revisiting dictionary-based compression. Software—Practice and Experience, 35 (15):1455–1476, December 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Snyder:2018:RNS

[SGDA18] Brett Snyder, Robert C.

Green II, Vijay Devabhaktuni, and Mansoor Alam. ReliaCloud-NS: a scalable web-based simulation platform for evaluating the reliability of cloud computing systems. Software—Practice and Experience, 48 (3):665–680, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2541.

Schnorf:1993:CTC

[SGH93] Peter Schnorf, Mahadevan Ganapathi, and John L. Hennessy. Compile-time copy elimination. Software—Practice and Experience, 23(11):1175–1200, November 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Skibinski:2008:EAX

[SGS08] Przemysław Skibiński, Szymon Grabowski, and Jakub Swacha. Effective asymmetric XML compression. Software—Practice and Experience, 38(10):1027– 1047, August ??, 2008. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sheikhalishahi:2015:ARC

[SGWVP15] Mehdi Sheikhalishahi, Lucio Grandinetti, Richard M. Wallace, and Jose Luis Vazquez-Poletti. Autonomic resource contentionaware scheduling. Software —Practice and Experience, 45(2):161–175, February 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sleep:1982:SNC

[SH82] M. R. Sleep and S. Holmström. A short note concerning lazy reduction rules for append. Software—Practice and Experience, 12(11):1082–1084, November 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Smith:1998:HPM

[SH98] Peter Smith and Norman C. Hutchinson. Heterogeneous process migration: The Tui system. Software—Practice and Experience, 28(6):611– 639, May 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1801;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1801&
PLACEBO=IE.pdf.

Saiedian:2003:CEG

[SH03] Hossein Saiedian and Steve Hill. A comparative evaluation of generic programming in Java and C++. Software—Practice and Experience, 33(2):121– 142, February 2003. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
102525916/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=102525916&PLACEBO=
IE.pdf.

Sirkia:2017:IOL

[SH17] Teemu Sirkiä and Lassi Haaranen. Improving online learning activity interoperability with Acos server. Software—Practice and Experience, 47 (11):1657–1676, November 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sun:2022:CGF

[SH22] Wen-Lin Sun and Yu-Lun Huang. Cross: a generic framework for system integration and its adaption in hospitals. Software—Practice and Experience, 52(7): 1643–1660, July 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shave:1972:BRB

[Sha72] M. J. R. Shave. Book review: Data structure: Theory and practice, A. T. Berztiss, Academic Press, New York, 1971, No. of pages: 442. Price: £6.75. Software—Practice and Experience, 2(2):191– 192, April 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shave:1977:SUT

[Sha77] M. J. R. Shave. The simulation and use of a TABLE data type in Algol W. Software—Practice and Experience, 7(6):733–745, November/December 1977. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shave:1978:PSR

[Sha78] M. J. R. Shave. The programming of structural relationships in dynamic environments. Software— Practice and Experience, 8 (2):199–211, March/April 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shave:1980:PID

[Sha80] M. J. R. Shave. Problems of integrity and distributed databases. Software—Practice and Experience, 10(2):135–147, February 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shave:1983:BRB

[Sha83] M. J. R. Shave. Book review: Distributed data bases, I. W. Draffan and F. Poole (Eds), Cambridge University Press, 1980. no. of pages: 389. Price: £18.50. Software—Practice and Experience, 13(2):197, February 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shaalan:2005:AGG

[Sha05] Khaled F. Shaalan. Arabic GramCheck: a grammar checker for Arabic. Software—Practice and Experience, 35(7):643–665, June 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Son:2019:CNM

[SHB19] Jungmin Son, TianZhang He, and Rajkumar Buyya. CloudSimSDN-NFV: Modeling and simulation of network function virtualization and service function chaining in edge computing environments. Software—Practice and Experience, 49(12):1748–1764, December 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Spier:1974:SMT

[SHC74] Michael J. Spier, Thomas N. Hastings, and David N. Cutler. A storage mapping technique for the implementation of protective domains. Software—Practice and Experience, 4(3):215– 230, July/September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shearn:1975:DES

[She75] D. C. S. Shearn. Discrete event simulation in AL-GOL 68. Software—Practice and Experience, 5(3): 279–293, July/September 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shen:1981:LE

[She81a] M. K. Shen. Letter to the editor. Software—Practice and Experience, 11(6):

649, June 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shepherd:1981:ASC

[She81b] W. D. Shepherd. Ancilla: a server for the Cambridge Model Distributed System. Software—Practice and Experience, 11 (11):1185–1195, November 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shelton:1992:WOM

[She92]  Jim Shelton.  A well-

oiled methodology. Software Magazine, 12(1):49– ??, January 1992. CODEN SMWMEQ. ISSN 0897-8085.

Sheridan:2007:PTC

[She07] Flash Sheridan. Practical testing of a C99 compiler using output comparison. Software—Practice and Experience, 37(14): 1475–1488, November 25, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Saur:2016:CST

[SHF16] Karla Saur, Michael Hicks, and Jeffrey S. Foster. Cstrider: type-aware heap traversal for C. Software —Practice and Experience, 46(6):767–788, June 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Soule:2016:RIL

[SHGG16] Robert Soulé, Martin Hirzel, Buğra Gedik, and Robert Grimm. River: an intermediate language for stream processing. Software—Practice and Experience, 46(7):891–929, July 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Scheiman:1999:PTC

[SHIS99] Chris J. Scheiman, Bjoern Haake, Maximilian Ibel, and Klaus E. Schauser. Profiling techniques for communication in finegrained parallel languages. Software—Practice and Experience, 29(6):519–550, May 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61000493;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=61000493&PLACEBO=IE.
pdf.

Shneiderman:1973:PS

[Shn73] Ben Shneiderman. Polynomial search. Software —Practice and Experience, 3(1):5–8, January/March 1973. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Shrivastava:1976:SPS

[Shr76] S. K. Shrivastava. Systematic programming of scheduling algorithms. Software—Practice and Experience, 6(3):357–370, July/ September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shrivastava:1978:SPR

[Shr78] S. K. Shrivastava. Sequential Pascal with recovery blocks. Software—Practice and Experience, 8(2):177– 185, March/April 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shrivastava:1979:CPBb

[Shr79a] S. K. Shrivastava. Concurrent Pascal with backward error recovery: Implementation. Software— Practice and Experience, 9 (12):1021–1033, December 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shrivastava:1979:CPBa

[Shr79b] S. K. Shrivastava. Concurrent Pascal with backward error recovery: Language features and examples. Software—Practice and Experience, 9 (12):1001–1020, December 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schofield:1980:MMM

[SHR80] D. Schofield, A. L. Hillman, and J. L. Rodgers. MM/1, a man-machine interface. Software—Practice and Experience, 10(9): 751–763, September 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Seshadri:2010:PDS

[SI10] Karthick Seshadri and K. Viswanathan Iyer. Parallelization of a dynamic SVD clustering algorithm and its application in information retrieval. Software —Practice and Experience, 40(10):883–896, September 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Safaei:2020:SNA

[SIC^{+}20] Mahmood Safaei, Abul Samad Ismail, Hassan Chizari, Maha Driss, Wadii Boulila, Shahla Asadi, and Mitra Safaei. Standalone noise and anomaly detection in wireless sensor networks: a novel time-series and adaptive Bayesiannetwork-based approach. Software—Practice and Experience, 50(4):428–446, April 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Solaimani:2016:OAD

[SIK^{+}16] Mohiuddin Solaimani, Mohammed Iftekhar, Latifur Khan, Bhavani Thuraisingham, Joe Ingram, and Sadi Evren Seker. Online anomaly detection for multi-source VMware using a distributed streaming framework. Software—Practice and Experience, 46(11):1479–1497, November 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Silberschatz:1981:ACM

[Sil81] Abraham Silberschatz. On the access-control mechanism of the program component manager. Software —Practice and Experience, 11(2):159–166, February 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Silberberg:1992:ISP

[Sil92]  David Silberberg.  In-

gest — a simple program for performing distributed relational database operations. Software—Practice and Experience, 22(6):455– 466, June 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Simpson:1983:BRB

[Sim83] D. Simpson. Book review: Theoretical foundations of programming methodology, Manfred Broy and Gunther Schmidt (eds.), D. Reidel Publishing Co, 1982. No. of pages: 658. Price $78.50 (hardback), $39.50 (paperback). Software— Practice and Experience, 13(10):977–979, October 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Singer:1981:SEM

[Sin81] D. W. Singer. SCENAR-IOS: An event management package. Software—Practice and Experience, 11(5): 521–529, May 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sassa:1995:RCG

[SIN95] Masataka Sassa, Harushi Ishizuka, and Ikuo Nakata. Rie, a compiler generator based on a one-pass-type attribute grammar. Software—Practice and Experience, 25(3):229–250, March 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sites:1979:RAI

[Sit79] Richard L. Sites. Resolving ambiguous identifiers. Software—Practice and Experience, 9(7):571– 574, July 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sawyer:1979:GMR

[SJ79] Dana C. Sawyer and Gerald C. Johns. ‘Guardians’— A method for running old programs under new operating systems. Software —Practice and Experience, 9(11):969–970, November 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sanghi:2004:TPE

[SJA^{+}04] Dheeraj Sanghi, Pankaj Jalote, Puneet Agarwal, Nitin Jain, and Supriyo Bose. A testbed for performance evaluation of loadbalancing strategies for Web server systems. Software—Practice and Experience, 34(4):339–353, April 10, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shihab:2011:PCU

[SJA^{+}11] Emad Shihab, Zhen Ming Jiang, Bram Adams, Ahmed E. Hassan, and Robert Bowerman. Prioritizing the creation of unit tests in legacy software systems. Software—Practice and Experience, 41(10):1027–1048, September 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sarwar:2021:PRP

[SJK^{+}21] Muhammad Usman Sarwar, Abdul Rehman Javed, Farzana Kulsoom, Suleman Khan, Usman Tariq, and Ali Kashif Bashir. PAR-CIV: Recognizing physical activities having complex interclass variations using semantic data of smartphone. Software—Practice and Experience, 51(3):532– 549, March 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Snelick:1994:SPT

[SJKL94] Robert Snelick, Joseph JáJá, Raghu Kacker, and Gordon Lyon. Syntheticperturbation techniques for screening shared memory programs. Software—Practice and Experience, 24(8): 679–701, August 1994. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Seo:2009:CTR

[SJP^{+}09] Euiseong Seo, Jinkyu Jeong, Seonyeong Park, Jinsoo Kim, and Joonwoon Lee. Catching two rabbits: adaptive realtime support for embedded Linux. Software—Practice and Experience, 39(5):531– 550, April 10, 2009. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stone:1996:TTS

[SK96] Roger G. Stone and Gada F.

Kadoda. Token-by-token syntax-directed editing (using an LR parser). Software—Practice and Experience, 26(11):1247–1260, November 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16772.

Schmidt:2003:IVL

[SK03] Carsten Schmidt and Uwe Kastens. Implementation of visual languages using pattern-based specifications. Software—Practice and Experience, 33 (15):1471–1505, December 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Singh:2022:SSP

[SKD^{+}22] Ashima Singh, Amrita Kaur, Arwinder Dhillon, Sahil Ahuja, and Harpreet Vohra. Software system to predict the infection in COVID-19 patients using deep learning and web of things. Software—Practice and Experience, 52(4):868– 886, April 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sugiki:2008:TMT

[SKI08] Akiyoshi Sugiki, Kenji Kono, and Hideya Iwasaki. Tuning mechanisms for two major parameters of Apache Web servers. Software—Practice and Experience, 38(12):1215–1240, October ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Systa:2001:SER

[SKM01] Tarja Systä, Kai Koskimies, and Hausi Müller. Shimba — an environment for reverse engineering Java software systems. Software—Practice and Experience, 31(4):371–394, April 10, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        77004439/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=77004439&PLACEBO=IE.
                                        pdf.

Schwabe:1978:DID

[SL78] Daniel Schwabe and Carlos J. P. Lucena. Design and implementation of data abstraction definition facility. Software—Practice and Experience, 8(6):709–719, November/December 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stepney:1987:FSA

[SL87] Susan Stepney and Stephen P.

Lord. Formal specification of an access control system. Software—Practice and Experience, 17(9): 575–593, September 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schwanke:2004:EAD

[SL04] Robert W. Schwanke and Robyn R. Lutz. Experience with the architectural design of a modest product family. Software —Practice and Experience, 34(13):1273–1296, November 10, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Slater:1986:SBM

[Sla86] Mel Slater. Segments on bit-mapped graphics displays. Software—Practice and Experience, 16(11): 965–980, November 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Savary-Leblanc:2023:SAS

[SLBC^{+}23] Maxime Savary-Leblanc, Lola Burgueño, Jordi Cabot, Xavier Le Pallec, and Sébastien Gérard. Software assistants in software engineering: a systematic mapping study. Software—Practice and Experience, 53(3):856–892, March 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Suk:2018:UCO

[SLJ^{+}18] Jae Hyuk Suk, Jae-Yung Lee, Hongjoo Jin, In Seok Kim, and Dong Hoon Lee. UnThemida: Commercial obfuscation technique analysis with a fully obfuscated program. Software—Practice and Experience, 48 (12):2331–2349, December 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Slonneger:1993:ECS

[Slo93] Ken Slonneger. Executing continuation semantics: a comparison. Software—Practice and Experience, 23(12):1379–1397, December 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schantz:2006:CQS

[SLRS06] Richard E. Schantz, Joseph P. Loyall, Craig Rodrigues, and Douglas C. Schmidt. Controlling quality-of-service in distributed real-time and embedded systems via adaptive middleware. Software—Practice and Experience, 36(11–12):1189–1208, September/October 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sauer:1979:QNS

[SM79] Charles H. Sauer and E. A.

MacNair. Queueing network software for systems modelling. Software— Practice and Experience, 9(5):369–380, May 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Steensgaard-Madsen:1981:MPP

[SM81] Jørgen Steensgaard-Madsen.

Modular programming with Pascal. Software—Practice and Experience, 11 (12):1331–1337, December 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sears:1985:SCR

[SM85] Ken H. Sears and Alan E.

Middleditch. Software concurrency in real-time control systems: a software nucleus. Software—Practice and Experience, 15(8):739– 759, August 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schonfelder:1990:DSF

[SM90] J. L. Schonfelder and J. S. Morgan. Dynamic strings in Fortran 90. Software—Practice and Experience, 20(12):1259–1271, December 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Steensgaard-Madsen:1999:HHE

[SM99] Jørgen Steensgaard-Madsen. HTEL: a HyperText Expression Language. Software—Practice and Experience, 29(8):661–675, July 10, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=62004392;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=62004392&PLACEBO=IE.
pdf.

Shah:2001:DIE

[SM01] Hemal V. Shah and Rajesh S. Madukkarumukumana. Design and implementation of efficient communication abstractions on the Virtual Interface Architecture: Stream sockets and RPC experience. Software—Practice and Experience, 31(11):1043–1065, September 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
83501987/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=83501987&PLACEBO=IE.
pdf.

Smaragdakis:2002:FFT

[SM02] Yannis Smaragdakis and Brian McNamara. FC++: Functional tools for objectoriented tasks. Software—Practice and Experience, 32(10):1015–1033, August 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
96516180/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=96516180&PLACEBO=IE.
pdf.

SilvaSouza:2015:DAC

[SM15] Vítor E. Silva Souza and John Mylopoulos. Designing an adaptive computeraided ambulance dispatch system with Zanshin: an experience report. Software —Practice and Experience, 45(5):689–725, May 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Singh:2018:DSC

[SM18] Jagannath Singh and Durga Prasad Mohapatra. Dynamic slicing of concurrent AspectJ programs: an explicit context-sensitive approach. Software—Practice and Experience, 48(1):233– 260, January 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sen:2020:ACM

[SM20] Amartya Sen and Sanjay Madria. Analysis of a cloud migration framework for offline risk assessment of cloud service providers. Software—Practice and Experience, 50(6): 998–1021, June 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sankranti:2024:DPE

[SM24] Sivaiah Sankranti and S. Roji Marjorie. Designing power-efficient SRAM cells with SGFinFETs using LECTOR technique. Software—Practice and Experience, 54(5):704–720, May 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sannella:1993:MWV

[SMFBB93] Michael Sannella, John Maloney, Bjørn N. Freeman-Benson, and Alan Borning.

Multi-way versus one-way constraints in user interfaces: Experience with the DeltaBlue algorithm. Software—Practice and Experience, 23(5):529–566, May 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sierra:2007:EHM

[SMGMOFM07a] José Luis Sierra, Pablo Moreno-Ger, Iván Martínez-Ortiz, and Baltasar Fernández-Manjón. Erratum: A highly modular and extensible architecture for an integrated IMS-based authoring system: the < e-Aula> experience. Software—Practice and Experience, 37(5):579, April 25, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [SMGMOFM07b].

Sierra:2007:HME

[SMGMOFM07b] José Luis Sierra, Pablo Moreno-Ger, Iván Martínez-Ortiz, and Baltasar Fernández-Manjón. A highly modular and extensible architecture for an integrated IMS-based authoring system: the < e-Aula> experience. Software—Practice and Experience, 37(4): 441–461, April 10, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See erratum [SMGMOFM07a].

Smith:1980:MMC

[Smi80] Alan Jay Smith. Multiprogramming and memory contention. Software —Practice and Experience, 10(7):531–552, July 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Smith:1985:DMB

[Smi85] Edward T. Smith. A debugger for message-based processes. Software—Practice and Experience, 15 (11):1073–1086, November 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Smith:1989:RLM

[Smi89] Jonathan M. Smith. Rapid location of mount points. Software—Practice and Experience, 19(9):829–838, September 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Smith:1991:EVF

[Smi91] P. D. Smith. Experiments with a very fast substring search algorithm. Software—Practice and Experience, 21(10):1065–1074, October 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Smith:1994:TBM

[Smi94] P. D. Smith. On tuning the Boyer–Moore– Horspool string searching algorithm. Software —Practice and Experience, 24(4):435–436, April 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sadjadi:2006:MDO

[SMKZ06] S. M. Sadjadi, P. K.

McKinley, E. P. Kasten, and Z. Zhou. MetaSockets: design and operation of runtime reconfigurable communication services. Software—Practice and Experience, 36(11– 12):1157–1178, September/ October 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Silva:2023:JSF

[SMLS23] Joaquim Silva, Eduardo R. B. Marques, Luís M. B. Lopes, and Fernando M. A. Silva. Jay: a software framework for prototyping and evaluating offloading applications in hybrid edge clouds. Software—Practice and Experience, 53(10): 2007–2025, October 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Spenke:1984:LIE

[SMM^{+}84] Michael Spenke, Heinz Mühlenbein, Monika Mevenkamp, Friedemann Mattern, and Christian Beilken. A language independent error recovery method for LL(1) parsers. Software—Practice and Experience, 14 (11):1095–1107, November 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Soares:2013:GNA

[SMM13] Luiz Fernando G. Soares, Marcio F. Moreno, and Rafael S. Marinho. Ginga-NCL architecture for plugins. Software—Practice and Experience, 43(4):449– 463, April 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Soldani:2021:TMA

[SMNB21] Jacopo Soldani, Giuseppe Muntoni, Davide Neri, and Antonio Brogi. The μTOSCA toolchain: Mining, analyzing, and refactoring microservice-based architectures. Software— Practice and Experience, 51(7):1591–1621, July 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Straw:1989:OMP

[SMR89] Andrew Straw, Fred Mellender, and Steve Riegel.

Object management in a persistent Smalltalk system. Software—Practice and Experience, 19(8):719– 737, August 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shrivastava:1993:DFT

[SMR93] Santosh K. Shrivastava, Luigi V. Mancini, and Brian Randell. The duality of fault-tolerant system structures. Software —Practice and Experience, 23(7):773–798, July 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Seinturier:2012:CBM

[SMR^{+}12] Lionel Seinturier, Philippe Merle, Romain Rouvoy, Daniel Romero, Valerio Schiavoni, and Jean-Bernard Stefani. A component-based middleware platform for reconfigurable service-oriented architectures. Software— Practice and Experience, 42(5):559–583, May 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Staron:2018:IEE

[SMT^{+}18] Mirosław Staron, Wilhelm Meding, Matthias Tichy, Jonas Bjurhede, Holger Giese, and Ola Söder. Industrial experiences from evolving measurement systems into self-healing systems for improved availability. Software—Practice and Experience, 48 (3):719–739, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2522.

Szafron:1990:LII

[SN90] Duane Szafron and Randy Ng. LexAGen: An interactive incremental scanner generator. Software —Practice and Experience, 20(5):459–483, May 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Santelices:2001:FDV

[SN01] Raúl A. Santelices and Miguel Nussbaum. A framework for the development of videogames. Software—Practice and Experience, 31(11):1091–1107, September 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shah:2007:SMP

[SN07] Sam Shah and Brian D. Noble. A study of email patterns. Software —Practice and Experience, 37(14):1515–1538, November 25, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sneeringer:1978:UID

[Sne78] James Sneeringer. Userinterface design for text editing: A case study. Software—Practice and Experience, 8(5):543–557, September/October 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sreenivas:2021:MDB

[SNK21] Velagapudi Sreenivas, Varsha Namdeo, and Eda Vijay Kumar. Modified deep belief network based human emotion recognition with multiscale features from video sequences. Software—Practice and Experience, 51(6):1259–1279, June 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Salli:2015:EDA

[SNL15] Olli Salli, Olli Nevalainen, and Ville Leppänen. Eventdriven asynchronous method calls with the D-Bus message system. Software— Practice and Experience, 45(1):53–74, January 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sreenivasan:1980:ESR

[SNM80] K. Sreenivasan, G. A. Nelson, and J. A. Maksin, Jr. An experimental study of relative throughput in a multiprocessor computer system. Software—Practice and Experience, 10(12): 973–986, December 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Snow:1978:ETO

[Sno78a] C. R. Snow. An exercise in the transportation of an operating system. Software—Practice and Experience, 8(1):41–50, January/ February 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Snow:1978:STP

[Sno78b] C. R. Snow. The software tools project. Software —Practice and Experience, 8(5):585–599, September/ October 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Snow:1991:MPC

[Sno91] C. R. Snow. A multiprotocol campus time server. Software—Practice and Experience, 21(9):907– 922, September 1991. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Snyder:2008:BPA

[Sny08] W. Van Snyder. The benefits of posing application software as a language interpreter. Software—Practice and Experience, 38(5): 509–522, April 25, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Symes:1977:SFA

[SO77] L. R. Symes and R. R. Oldehoeft. Success and failure analysis of a problem solving system implementation. Software—Practice and Experience, 7(3): 383–389, May/June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sterling:2007:ABI

[SO07] Chad D. Sterling and Ronald A. Olsson. Automated bug isolation via program chipping. Software—Practice and Experience, 37(10):1061–1086, August 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Saifan:2021:FLE

[SO21] Ahmad A. Saifan and Lana Obeidat. Feature location enhancement based on source code augmentation with synonyms of terms. Software—Practice and Experience, 51(2):235– 259, February 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

St-Onge:2023:NMR

[SOKE23] Cédric St-Onge, Nadjia Kara, and Claes Edstrom. NFVLearn: a multiresource, long short-term memory-based virtual network function resource usage prediction architecture. Software—Practice and Experience, 53(3):555– 578, March 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sommerville:1982:PMS

[Som82] Ian Sommerville. A pattern matching system. Software —Practice and Experience, 12(6):517–530, June 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sosic:1995:PIP

[Sos95] Rok Sosič. A procedural interface for program directing. Software—Practice and Experience, 25(7):767– 787, July 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sozer:2015:ISC

[Söz15] H. Sözer. Integrated static code analysis and runtime verification. Software— Practice and Experience, 45(10):1359–1373, October 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sethi:1979:MSI

[SP79] H. R. Sethi and D. H. Paul.

MPLOT3 — a systemindependent plotting package. Software—Practice and Experience, 9(11):891– 905, November 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sajaniemi:1988:EAS

[SP88] Jorma Sajaniemi and Jari Pekkanen. An empirical analysis of spreadsheet calculation. Software—Practice and Experience, 18(6): 583–596, June 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Spafford:1990:EMT

[Spa90] Eugene H. Spafford. Extending mutation testing to find environmental bugs. Software—Practice and Experience, 20(2):181– 189, February 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Silva:2011:DTE

[SPHB11] Renato Silva, Carine Pascal, Thai Son Hoang, and Michael Butler. Decomposition tool for Event-B. Software—Practice and Experience, 41(2):199–208, February 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Spier:1976:SMD

[Spi76] Michael J. Spier. Software malpractice — a distasteful experience. Software—Practice and Experience, 6(3):293–299, July/ September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Spinellis:2002:UTV

[Spi02] Diomidis Spinellis. Unix tools as visual programming components in a GUIbuilder environment. Software—Practice and Experience, 32(1):57–71, January 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
88013123/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=88013123&PLACEBO=IE.
pdf.

Spivey:2004:FAC

[Spi04] J. M. Spivey. Fast, accurate call graph profiling. Software—Practice and Experience, 34(3):249– 264, March 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Spinellis:2009:ULO

[Spi09] Diomidis Spinellis. Userlevel operating system transactions. Software— Practice and Experience, 39(14):1215–1233, September 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Spooner:1971:SAP

[Spo71] C. R. Spooner. A software architecture for the 70’s: Part I — the general approach. Software—Practice and Experience, 1(1): 5–37, January/March 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schoeberl:2010:WCE

[SPPH10] Martin Schoeberl, Wolfgang Puffitsch, Rasmus Ulslev Pedersen, and Benedikt Huber. Worst-case execution time analysis for a Java processor. Software —Practice and Experience, 40(6):507–542, May 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sabir:2019:SLR

[SPR^{+}19] Fatima Sabir, Francis Palma, Ghulam Rasool, Yann-Gaël Guéhéneuc, and Naouel Moha. A systematic literature review on the detection of smells and their evolution in object-oriented and service-oriented systems. Software—Practice and Experience, 49(1):3– 39, January 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sharma:1988:MDS

[SR88] Ravi Sharma and Lawrence L. Rose. Modular design for simulation. Software —Practice and Experience, 18(10):945–966, October 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sunderam:1991:ESH

[SR91] V. S. Sunderam and Vernon J. Rego. EcliPSe: a system for high performance concurrent simulation. Software—Practice and Experience, 21 (11):1189–1219, November 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Simon:2002:RSD

[SR02] Robert Simon and Paul Rabinovich. Rapid simulator development for multicast protocol analysis. Software—Practice and Experience, 32(1):1– 23, January 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
88013121/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=88013121&PLACEBO=IE.
pdf.

Simmhan:2018:CI

[SRC^{+}18a] Yogesh Simmhan, Pushkara Ravindra, Shilpa Chaturvedi, Malati Hegde, and Rashmi Ballamajalu. Cover image. Software—Practice and Experience, 48 (7):i, July 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2589.

Simmhan:2018:TDD

[SRC^{+}18b] Yogesh Simmhan, Pushkara Ravindra, Shilpa Chaturvedi, Malati Hegde, and Rashmi Ballamajalu. Towards a data-driven IoT software architecture for smart city utilities. Software—Practice and Experience, 48 (7):1390–1416, July 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2580.

Sicari:2019:HEI

[SRCP19] Sabrina Sicari, Alessandra Rizzardi, and Alberto Coen-Porisini. How to evaluate an Internet of Things system: Models, case studies, and real developments. Software—Practice and Experience, 49 (11):1663–1685, November 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sreenivasan:1976:AAD

[Sre76] K. Sreenivasan. Application of accounting data in evaluating computer system performance. Software —Practice and Experience, 6(2):239–244, April/June 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sendin-Rana:2009:IPF

[SRGCPB^{+}09] Pablo Sendín-Raña, Francisco J. González-Castaño, Enrique Pérez-Barros, Pedro S. Rodríguez-Hernández, Felipe Gil-Castiñeira, and José M. Pousada-Carballo. Improving the performance and functionality of Mondrian open-source OLAP systems. Software—Practice and Experience, 39(3): 279–298, March 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Srirama:2024:DRF

[Sri24] Satish Narayana Srirama.

A decade of research in fog computing: Relevance, challenges, and future directions. Software—Practice and Experience, 54(1): 3–23, January 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sendin-Rana:2008:EAG

[SROAdM^{+}08] Pablo Sendín-Raña, Nicolás Otero-Alonso, Vicente Goyanes de Miguel, Francisco J. González-Castaño, Pedro S. Rodríguez-Hernández, Felipe Gil-Castiñeira, and Enrique Costa-Montenegro. Emulating Access Grid features at Web endpoints: a developer’s view. Software —Practice and Experience, 38(13):1393–1410, November 10, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stoermer:2006:MCS

[SROV06] Christoph Stoermer, Anthony Rowe, Liam O’Brien, and Chris Verhoef. Modelcentric software architecture reconstruction. Software—Practice and Experience, 36(4):333–363, April 10, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sendin-Rana:2010:WOB

[SRRFGC^{+}10] P. Sendín-Raña, E. Rodríguez-Fernández, F. J. González-Castaño, E. Costa-Montenegro, P. S. Rodríguez-Hernández, J. M. Pousada-Carballo, and J. C. Burguillo-Rial. Web-oriented business intelligence solution based on Associative Query Logic. Software—Practice and Experience, 40(9):779–796, August 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Saghi:1998:MSH

[SRS98] Gene Saghi, Kirk Reinholtz, and Paul A. Savory. A multithreaded scheduler for a high-speed spacecraft simulator. Software—Practice and Experience, 28(6): 641–656, May 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=1802;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=1802&
                                        PLACEBO=IE.pdf.

Shirvani:2018:IMD

[SRS18] Mirsaeid Hosseini Shirvani, Amir Masoud Rahmani, and Amir Sahafi. An iterative mathematical decision model for cloud migration: a cost and security risk approach. Software—Practice and Experience, 48 (3):449–485, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2528.

Sommerville:1984:ES

[SS84] Ian Sommerville and D. J.

Smith. An electronic secretary. Software—Practice and Experience, 14(9): 817–825, September 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Silverman:1989:DBS

[SS89] Robert D. Silverman and Sidney J. Stuart. A distributed batching system for parallel processing. Software—Practice and Experience, 19 (12):1163–1174, December 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Santucci:1993:QDV

[SS93] Giuseppe Santucci and Pier Angelo Sottile. Query by diagram: a visual environment for querying databases. Software —Practice and Experience, 23(3):317–340, March 1993. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Side:1994:DDP

[SS94] R. S. Side and G. C. Shoja. A debugger for distributed programs. Software—Practice and Experience, 24(5): 507–525, May 1994. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sametinger:1995:DIA

[SS95] J. Sametinger and S. Schiffer. Design and implementation aspects of an experimental C++ programming environment. Software— Practice and Experience, 25(2):111–128, February 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sagonas:2003:EEI

[SS03] Konstantinos Sagonas and Erik Stenman. Experimental evaluation and improvements to linear scan register allocation. Software—Practice and Experience, 33(11):1003–1034, September 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schurmann:2007:IAF

[SS07] Klaus-Bernd Schürmann and Jens Stoye. An incomplex algorithm for fast suffix array construction. Software—Practice and Experience, 37(3):309–329, March 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Skrbic:2008:BRE

[ŠS08] Srd̄an Škrbić and Dušan Surla. Bibliographic records editor in XML native environment. Software —Practice and Experience, 38(5):471–491, April 25, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stefanov:2009:IBC

[SS09] Ekaterina Stefanov and Anthony M. Sloane. On the implementation of bytecode compression for interpreted languages. Software—Practice and Experience, 39(2):111–135, February ??, 2009. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sharma:2019:USO

[SS19] Shipra Sharma and Balwinder Sodhi. Using Stack Overflow content to assist in code review. Software —Practice and Experience, 49(8):1255–1277, August 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Singh:2021:GLY

[SS21] Robbert Gurdeep Singh and Christophe Scholliers. GraphRedex: Look at your research. Software—Practice and Experience, 51(6): 1322–1351, June 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Srinivasarao:2023:MOM

[SS23] Popuri Srinivasarao and Aravapalli Rama Satish. Multi-objective materialized view selection using flamingo search optimization algorithm. Software—Practice and Experience, 53(4):988–1012, April 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sarimbekov:2016:WCJ

[SSB^{+}16] Aibek Sarimbekov, Lukas Stadler, Lubomír Bulej, Andreas Sewe, Andrej Podzimek, Yudi Zheng, and Walter Binder. Workload characterization of JVM languages. Software —Practice and Experience, 46(8):1053–1089, August 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sanchez-Segura:2003:ATS

[SSCdA^{+}03] María I. Sánchez-Segura, Juan J. Cuadrado, Angelica de Antonio, Antonio de Amescua, and Luis García. Adapting traditional software processes to virtual environments development. Software—Practice and Experience, 33(11):1051–1080, September 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Saidis:2011:DVH

[SSD11] Kostas Saidis, Yannis Smaragdakis, and Alex Delis. DOLAR: virtualizing heterogeneous information spaces to support their expansion. Software —Practice and Experience, 41(11):1349–1383, October 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shakarami:2020:RCO

[SSGA20] Ali Shakarami, Ali Shahidinejad, and Mostafa Ghobaei-Arani. A review on the computation offloading approaches in mobile edge computing: a gametheoretic perspective. Software—Practice and Experience, 50(9):1719–1759, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Suh:2017:EET

[SSK^{+}17] Young-Kyoon Suh, Richard T. Snodgrass, John D. Kececioglu, Peter J. Downey, Robert S. Maier, and Cheng Yi. EMP: execution time measurement protocol for compute-bound programs. Software—Practice and Experience, 47(4):559– 597, April 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Szydlo:2022:IOR

[SSK^{+}22] Tomasz Szydlo, Amadeusz Szabala, Nazar Kordiumov, Konrad Siuzdak, Lukasz Wolski, Khaled Alwasel, Fawzy Habeeb, and Rajiv Ranjan. IoTSim-Osmosis-RES: Towards autonomic renewable energyaware osmotic computing. Software—Practice and Experience, 52(7): 1698–1716, July 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sinha:2022:CBH

[SSKG22] Akash Sinha, Gulshan Shrivastava, Prabhat Kumar, and Deepak Gupta. A community-based hierarchical user authentication scheme for Industry 4.0. Software—Practice and Experience, 52(3):729– 743, March 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schoofs:2011:PMP

[SSM11] Anthony Schoofs and Phillip Stanley-Marbell. Portability in MAC protocol and transceiver software implementations for LR-WPAN platforms. Software—Practice and Experience, 41(4): 339–361, April 10, 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stal:2013:ETA

[SSO13] Michael Stal, Douglas C.

Schmidt, and William R. Otte. Efficiently and transparently automating scalable on-demand activation and deactivation of services with the activator pattern. Software—Practice and Experience, 43(7):789– 806, July 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Scheler:2011:RTS

[SSP11] Fabian Scheler and Wolfgang Schröder-Preikschat. The Real-Time Systems Compiler: migrating eventtriggered systems to timetriggered systems. Software —Practice and Experience, 41(12):1491–1515, November/December 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shobaki:2015:EEV

[SSRAH15] Ghassan Shobaki, Laith Sakka, Najm Eldeen Abu Rmaileh, and Hasan Al-Hamash. Experimental evaluation of various register-pressure-reduction heuristics. Software—Practice and Experience, 45 (11):1497–1517, November 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Slottow:2002:ITD

[SSS^{+}02] Joan Slottow, Ali Shahriari, Michael Stein, Xiao Chen, Chris Thomas, and Philip B. Ender. Instrumenting and tuning dataView — a networked application for navigating through large scientific datasets. Software—Practice and Experience, 32(2):165–190, February 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
88511127/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=88511127&PLACEBO=IE.
pdf.

Sharma:2022:UOS

[SSS22] Pankajeshwara Nand Sharma, Bastin Tony Roy Savarimuthu, and Nigel Stanger. Unearthing open source decisionmaking processes: a case study of Python enhancement proposals. Software —Practice and Experience, 52(10):2312–2346, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sor:2015:MLD

[SSST15] Vladimir Sor, Satish Narayana Srirama, and Nikita Salnikov-Tarnovski. Memory leak detection in Plumbr. Software—Practice and Experience, 45(10):1307–1330, October 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sankar:2020:ISA

[SSV^{+}20] H. Sankar, V. Subramaniyaswamy, V. Vijayakumar, Sangaiah Arun Kumar, R. Logesh, and A. Umamakeswari. Intelligent sentiment analysis approach using edge computing-based deep learning technique. Software —Practice and Experience, 50(5):645–657, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Storey:1977:PAL

[ST77] Tony Storey and Stephen Todd. Performance analysis of large systems. Software—Practice and Experience, 7(3):363–369, May/ June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schonfelder:1979:APA

[ST79] J. L. Schonfelder and J. T. Thomason. Arbitrary precision arithmetic in Algol 68. Software—Practice and Experience, 9(3):173– 182, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Suzuki:2001:DCS

[ST01] Tetsuya Suzuki and Takehiro Tokuda. The DeltaUp constraint solver: minimizing the number of method selections in DeltaBlue. Software—Practice and Experience, 31(14):1351– 1361, November 25, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        http://www3.interscience.
                                        wiley.com/cgi-bin/abstract/
                                        85512985/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=85512985&PLACEBO=IE.
                                        pdf.

Schrefl:2004:URJ

[ST04] Michael Schrefl and Thomas Thalhammer. Using roles in Java. Software—Practice and Experience, 34(5): 449–464, April 25, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Salehie:2012:TGD

[ST12] Mazeiar Salehie and Ladan Tahvildari. Towards a goaldriven approach to action selection in self-adaptive software. Software—Practice and Experience, 42(2): 211–233, February 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shtern:2014:MSI

[ST14] Mark Shtern and Vassilios Tzerpos. Methods for selecting and improving software clustering algorithms. Software—Practice and Experience, 44(1): 33–46, January 2014. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Steinhauser:2019:DAE

[ST19] Antonín Steinhauser and Petr Tuma. DjangoChecker: Applying extended taint tracking and server side parsing for detection of context-sensitive XSS flaws. Software—Practice and Experience, 49(1):130–148, January 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Staunstrup:1982:MPC

[Sta82] Jørgen Staunstrup. Message passing communication versus procedure call communication. Software—Practice and Experience, 12(3):223–234, March 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Staelin:2005:LEM

[Sta05] Carl Staelin. lmbench: an extensible micro-benchmark suite. Software—Practice and Experience, 35 (11):1079–1105, September 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Starosolski:2007:SFA

[Sta07] Roman Starosolski. Simple fast and adaptive lossless image compression algorithm. Software—Practice and Experience, 37(1): 65–91, January 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sozer:2009:FFD

[STA09] Hasan Sözer, Bedir Tekinerdoğan, and Mehmet Akşit. FLORA: a framework for decomposing software architecture to introduce local recovery. Software—Practice and Experience, 39(10):869–889, July ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Salehi:2014:CMF

[STB14] Mohsen Amini Salehi, Adel Nadjaran Toosi, and Rajkumar Buyya. Contention management in federated virtualized distributed systems: implementation and evaluation. Software—Practice and Experience, 44(3):353– 368, March 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stewart:1979:LEM

[Ste79] H. D. Stewart. Letter to the Editor: Mini-computer requirements. Software— Practice and Experience, 9(6):510–511, June 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stevenson:1980:CGR

[Ste80] Gordon Stevenson. Code generation with a recursive optimizer. Software —Practice and Experience, 10(5):393–403, May 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Steffen:1984:EPD

[Ste84] Joseph L. Steffen. Experience with a portable debugging tool. Software —Practice and Experience, 14(4):323–334, April 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Steffen:1992:ART

[Ste92] Joseph L. Steffen. Adding run-time checking to the portable C compiler. Software—Practice and Experience, 22(4):305–316, April 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Steenkiste:1998:DIE

[Ste98]  Peter Steenkiste.  De-

sign, implementation, and evaluation of a singlecopy protocol stack. Software—Practice and Experience, 28(7):749–772, July 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1807;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1807&
PLACEBO=IE.pdf.

Steckler:2002:CSP

[Ste02] Paul A. Steckler. Component support in PLT scheme. Software—Practice and Experience, 32 (10):933–954, August 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www3.interscience.
wiley.com/cgi-bin/abstract/
96516181/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=96516181&PLACEBO=IE.
pdf.

Salampasis:1997:HSA

[STH97] Michail Salampasis, John Tait, and Colin Hardy. HyperTree: a structural approach to Web authoring. Software—Practice and Experience, 27(12):1411– 1426, December 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7278;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7278&
PLACEBO=IE.pdf.

Shiang:2018:ORM

[STH^{+}18] Cheah Wai Shiang, Fu Swee Tee, Alfian Abdul Halin, Ng Keng Yap, and Puah Chin Hong. Ontology reuse for multiagent system development through pattern classification. Software—Practice and Experience, 48 (11):1923–1939, November 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stirling:1978:UPO

[Sti78] W. Douglas Stirling. The use of a procedure-oriented language for process-oriented simulation. Software— Practice and Experience, 8 (2):137–148, March/April 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stiegler:1979:SAC

[Sti79] Helmut G. Stiegler. A structure for access control lists. Software—Practice and Experience, 9(10):813– 819, October 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stirling:1985:FSE

[Sti85]  Colin Stirling.  Follow

set error recovery. Software—Practice and Experience, 15(3):239–257, March 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stone:1988:BRB

[Sto88] Roger G. Stone. Book review: Concurrency and programming languages, D. M. Harland, Ellis Horwood, Ltd., 1986. ISBN 0-85312-846-4. No. of pages: 182. Price: £19.50. Software —Practice and Experience, 18(2):167, February 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stoyenko:1994:SRS

[Sto94] Alexander D. Stoyenko.

SUPRA-RPC: SUbprogram PaRAmeters in Remote Procedure Calls. Software—Practice and Experience, 24(1):27–49, January 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stone:2005:VDW

[Sto05] Roger G. Stone. Validation of dynamic Web pages generated by an embedded scripting language. Software—Practice and Experience, 35(13):1259–1274, November 10, 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Strawn:1977:DAR

[Str77] George O. Strawn. Does APL really need run-time parsing? Software— Practice and Experience, 7 (2):193–200, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stroustrup:1981:LRT

[Str81] Bjarne Stroustrup. ‘Long return’: a technique for improving the efficiency of inter-module communication. Software—Practice and Experience, 11(2):131– 143, February 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stroustrup:1982:EIP

[Str82] Bjarne Stroustrup. An experiment with the interchangeability of processes and monitors. Software—Practice and Experience, 12(11):1011–1025, November 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stroustrup:1983:ACC

[Str83a] Bjarne Stroustrup. Adding classes to the C language: An exercise in language evolution. Software—Practice and Experience, 13(2): 139–161, February 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Strubbe:1983:KRG

[Str83b] Hugo J. Strubbe. Kernel for a responsive and graphical user interface. Software—Practice and Experience, 13(11):1033–1042, November 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Strumpen:1995:CHW

[Str95] Volker Strumpen. Coupling hundreds of workstations for parallel molecular sequence analysis. Software—Practice and Experience, 25(3):291–304, March 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sykes:1983:ESD

[STS83] Floyd Sykes, Raymond T.

Tillman, and Ben Shneiderman. The effect of scope delimiters on program comprehension. Software— Practice and Experience, 13(9):817–824, September 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Soylemez:2024:MRA

[STT24] Mehmet Söylemez, Bedir Tekinerdogan, and Ayça Kolukisa Tarhan. Microservice reference architecture design: a multi-case study. Software—Practice and Experience, 54(1):58– 84, January 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Surla:2013:ERD

[Sur13] Bojana Dimić Surla. Experience report: Developing an Eclipse editor for MARC records using Xtext. Software—Practice and Experience, 43 (11):1377–1392, November 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Steele:2022:CES

[SV22] Guy L. Steele, Jr. and Sebastiano Vigna. Computationally easy, spectrally good multipliers for congruential pseudorandom number generators. Software—Practice and Experience, 52(2):443–458, February 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Svahnberg:2005:TVR

[SvGB05] Mikael Svahnberg, Jilles van Gurp, and Jan Bosch. A taxonomy of variability realization techniques. Software—Practice and Experience, 35(8):705–754, July 10, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Scowen:1974:DCP

[SW74] R. S. Scowen and B. A. Wichmann. The definition of comments in programming languages. Software —Practice and Experience, 4(2):181–188, April/June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Spector:1982:LE

[SW82] David Spector and Martin Wakeman. Letters to the editor. Software—Practice and Experience, 12(1):98– 99, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schach:1986:APF

[SW86a] Stephen R. Schach and Peter T. Wood. An almost path-free very highlevel interactive data manipulation language for a microcomputer-based database system. Software—Practice and Experience, 16(3):243–268, March 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Snow:1986:ENC

[SW86b] C. R. Snow and H. Whitfield. An experiment with the Newcastle Connection Protocol. Software—Practice and Experience, 16 (11):1031–1043, November 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Smith:1987:FTA

[SW87] John B. Smith and Stephen F.

Weiss. Formatting texts accessed randomly. Software —Practice and Experience, 17(1):5–16, January 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schleiermacher:1990:IPP

[SW90] Adolf Schleiermacher and Jürgen F. H. Winkler. The implementation of ProTest: a Prolog debugger for a refined box model. Software —Practice and Experience, 20(10):985–1006, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Seshadri:1991:ICS

[SW91] V. Seshadri and David B. Wortman. An investigation into concurrent semantic analysis. Software—Practice and Experience, 21(12):1323–1348, December 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Snow:1994:SA

[SW94] C. R. Snow and H. Whitfield. Simple authentication. Software—Practice and Experience, 24(5):437– 447, May 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stanier:2012:SIC

[SW12] James Stanier and Des Watson. A study of irreducibility in C programs. Software—Practice and Experience, 42(1):117– 130, January 2012. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shaikh:2014:FTU

[SW14] Asadullah Shaikh and Uffe Kock Wiil. A feedback technique for unsatisfiable UML/ OCL class diagrams. Software—Practice and Experience, 44 (11):1379–1393, November 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shaw:1975:MND

[SWA^{+}75] Alan Shaw, Nelson Weiderman, Gregory R. Andrews, Mary-Beth Felcyn, John Rieber, and Gain Wong. A multiprogramming nucleus with dynamic resource facilities. Software—Practice and Experience, 5(3):245–267, July/ September 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sjoberg:1997:EPB

[SWA^{+}97] Dag I. K. Sjøberg, Ray Welland, Malcolm P. Atkinson, Paul Philbrow, and Cathy Waite. Exploiting persistence in build management. Software—Practice and Experience, 27(4): 447–480, April 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7301;   http://www3.
interscience.wiley.com/

                                        cgi-bin/fulltext?ID=7301&
                                        PLACEBO=IE.pdf.

Sun:2017:CCB

[SWBS17] Hao Sun, Xiaofeng Wang, Rajkumar Buyya, and Jinshu Su. CloudEyes: Cloudbased malware detection with reversible sketch for resource-constrained Internet of Things (IoT) devices. Software—Practice and Experience, 47(3):421– 441, March 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sommerville:1986:SDM

[SWBT86] Ian Sommerville, R. Welland, I. Bennett, and R. Thomson. SOFTLIB: a documentation management system. Software—Practice and Experience, 16(2):131– 143, February 1986. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schmitz:1994:DIN

[SWN94] Daniel C. Schmitz, Johnny S. Wong, and Ron M. Nelson. Design and implementation of National Engineering Education Delivery System (NEEDS). Software—Practice and Experience, 24(11):1051–1076, November 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sommerville:1989:EUI

[SWPS89] Ian Sommerville, Ray Welland, Stuart Potter, and John Smart. The ECLIPSE user interface. Software—Practice and Experience, 19(4):371–391, April 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shi:2021:BBT

[SWY^{+}21] Peichang Shi, Huaimin Wang, Shangzhi Yang, Chang Chen, and Wentao Yang. Blockchainbased trusted data sharing among trusted stakeholders in IoT. Software —Practice and Experience, 51(10):2051–2064, October 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sun:2017:ATB

[SXWL17] Hailong Sun, Bang Xiao, Xu Wang, and Xudong Liu. Adaptive trade-off between consistency and performance in data replication. Software—Practice and Experience, 47(6):891– 906, June 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Spitz:1979:POP

[SY79] Eli Spitz and Asher Yuval. Public online pool — a method for managing online direct access space. Software—Practice and Experience, 9(2):139– 147, February 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stephens:1986:PMU

[SY86] P. D. Stephens and J. K. Yarwood. Providing multiuser access to distributed array processors. Software —Practice and Experience, 16(6):531–539, June 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sulistio:2004:TCB

[SYB04] Anthony Sulistio, Chee Shin Yeo, and Rajkumar Buyya. A taxonomy of computerbased simulations and its mapping to parallel and distributed systems simulation tools. Software—Practice and Experience, 34(7): 653–673, June 2004. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shi:2018:PPA

[SYG^{+}18] Kai Shi, Huiqun Yu, Jianmei Guo, Guisheng Fan, and Xingguang Yang. A parallel portfolio approach to configuration optimization for large software product lines. Software—Practice and Experience, 48(9):1588–1606, September 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stephens:1980:EOS

[SYRS80] P. D. Stephens, J. K.

Yarwood, D. J. Rees, and N. H. Shelness. The evolution of the operating system EMAS 2900. Software—Practice and Experience, 10(12):993–1008, December 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sui:2014:MCS

[SYXZ14] Yulei Sui, Sen Ye, Jingling Xue, and Jie Zhang. Making context-sensitive inclusion-based pointer analysis practical for compilers using parameterised summarisation. Software—Practice and Experience, 44(12):1485–1510, December 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Santana:1988:LBS

[SZ88] M. J. Santana and E. J. Zaluska. Load balancing in a session-based distributed file-store architecture. Software—Practice and Experience, 18(11):1091–1107, November 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stankovic:2000:EJI

[SZ00] Nenad Stankovic and Kang Zhang. An evaluation of Java implementations of message-passing. Software—Practice and Experience, 30(7):741–763, June 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72001830/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72001830&PLACEBO=IE.
pdf.

Seidl:2001:IHO

[SZ01] Matthew L. Seidl and Benjamin G. Zorn. Implementing heap-object behavior prediction efficiently and effectively. Software—Practice and Experience, 31(9):869–892, July 25, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
79503393/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=79503393&PLACEBO=IE.
pdf.

Strembeck:2009:ASD

[SZ09] Mark Strembeck and Uwe Zdun. An approach for the systematic development of domain-specific languages. Software— Practice and Experience, 39(15):1253–1292, October ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sharbaf:2020:CTW

[SZ20] Mohammadreza Sharbaf and Bahman Zamani. Configurable three-way model merging. Software—Practice and Experience, 50(8): 1565–1599, August 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Somani:2019:ICI

[SZSB19] Gaurav Somani, Xinghui Zhao, Satish Narayana Srirama, and Rajkumar Buyya. Integration of Cloud, Internet of Things, and Big Data Analytics. Software—Practice and Experience, 49(4):561– 564, April 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tzou:1991:PMP

[TA91] Shin-Yuan Tzou and David P.

Anderson. The performance of message-passing using restricted virtual memory remapping. Software—Practice and Experience, 21(3):251–267, March 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tsujino:1984:CCP

[TAAT84] Y. Tsujino, M. Ando, T. Araki, and N. Tokura. Concurrent C: a programming language for distributed multiprocessor systems. Software—Practice and Experience, 14 (11):1061–1078, November 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tonella:2000:REM

[TAFC00] P. Tonella, G. Antoniol, R. Fiutem, and F. Calzolari. Reverse engineering 4.7 million lines of code. Software—Practice and Experience, 30(2):129– 150, February 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
69503459/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=69503459&PLACEBO=IE.
pdf.

Tagg:1988:LLP

[Tag88] A. G. Tagg. Low-level programming in Modula-2: an experience with local area network access. Software —Practice and Experience, 18(6):495–507, June 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Turaga:2010:DPD

[TAG^{+}10] Deepak Turaga, Henrique Andrade, Buğra Gedik, Chitra Venkatramani, Olivier Verscheure, Jeffrey David Harris, John Cox, William Szewczyk, and Paul Jones. Design principles for developing stream processing applications. Software—Practice and Experience, 40 (12):1073–1104, November 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Torsun:1981:DAC

[TAJ81] I. S. Torsun and M. M. Al-Jarrah. Dynamic analysis of COBOL programs. Software—Practice and Experience, 11(9):949–961, September 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Taliaferro:1971:MKS

[Tal71] W. M. Taliaferro. Modularity. the key to system growth potential. Software—Practice and Experience, 1(3):245–257, July/ September 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tanenbaum:1973:CR

[Tan73] Andrew S. Tanenbaum. Computer recreations. Software—Practice and Experience, 3(4):397–399, October/December 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Taylor:1983:IVT

[Tay83] Richard N. Taylor. An integrated verification and testing environment. Software—Practice and Experience, 13(8):697–713, August 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thomas:1972:WWW

[TB72] R. E. Thomas and J. C. Baldwin. What was, what is and what should have been. A critical evaluation of the chilton multi-access system. Software—Practice and Experience, 2(4):313– 320, October/December 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tanenbaum:1973:PTS

[TB73] Andrew S. Tanenbaum and William H. Benson. The people’s time sharing system. Software—Practice and Experience, 3(2):109– 119, April/June 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Taylor:1986:EDS

[TB86] David J. Taylor and James P. Black. Experimenting with data structures. Software—Practice and Experience, 16(5):443– 456, May 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tripathi:1989:IOO

[TBA89] Anand Tripathi, Eric Berge, and Mehmet Aksit. An implementation of the object-oriented concurrent programming language SINA. Software —Practice and Experience, 19(3):235–256, March 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tomei:2022:CAP

[TBF^{+}22] Matteo Tomei, Lorenzo Baraldi, Giuseppe Fiameni, Simone Bronzin, and Rita Cucchiara. A computational approach for progressive architecture shrinkage in action recognition. Software—Practice and Experience, 52(2):537– 554, February 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tamburri:2020:CCM

[TBPK20] Damian Andrew Tamburri, Kelly Blincoe, Fabio Palomba, and Rick Kazman. “The Canary in the Coal Mine … ”: a cautionary tale from the decline of SourceForge. Software—Practice and Experience, 50(10):1930–1951, October 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tiwari:2018:OME

[TBSI18] Nidhi Tiwari, Umesh Bellur, Santonu Sarkar, and Maria Indrawan. Optimizing MapReduce for energy efficiency. Software—Practice and Experience, 48 (9):1660–1687, September 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thamsen:2023:SIB

[TBT23] Lauritz Thamsen, David Bermbach, and Demetris Trihinas. Special issue on benchmarking, experimentation tools, and reproducible practices for dataintensive systems from edge to cloud. Special focus section: Benchmarking, Experimentation Tools, and Reproducible Practices for Data-Intensive Systems from Edge to Cloud: Software: Practice and Experience, 53(12):2325–2326, December 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tambag:2003:MBL

[TC03] Y. Tambag and A. Cosar.

Managing business lifecycle using ebIOP. Software —Practice and Experience, 33(13):1217–1227, November 10, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tremblay:2007:MMD

[TC07] G. Tremblay and F. Champagne. Marking musical dictations using the edit distance algorithm. Software—Practice and Experience, 37(2):207–230, February 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Trotman:2019:MMO

[TC19] Andrew Trotman and Matt Crane. Micro- and macrooptimizations of SaaT search. Software—Practice and Experience, 49(5):942– 950, May 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tse:1994:APS

[TCC^{+}94] T. H. Tse, T. Y. Chen, F. T. Chan, H. Y. Chen, and H. L. Xie. The application of Prolog to structured design. Software—Practice and Experience, 24(7):659– 676, July 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Torrey:2007:CIL

[TCM07] Lisa A. Torrey, Joyce Coleman, and Barton P. Miller. A comparison of interactivity in the Linux 2.6 scheduler and an MLFQ scheduler. Software—Practice and Experience, 37(4):347– 364, April 10, 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tracey:2000:ATD

[TCMM00] N. Tracey, J. Clark, K. Mander, and J. Mc-Dermid. Automated testdata generation for exception conditions. Software—Practice and Experience, 30(1):61–79, January 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
69503095/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?

ID=69503095&PLACEBO=IE.
pdf.

Toyn:1994:EBT

[TD94] Ian Toyn and Alan J. Dix.

Efficient binary transfer of pointer structures. Software—Practice and Experience, 24(11):1001–1023, November 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tchana:2015:SSL

[TDDE15] Alain Tchana, Noel De Palma, Bruno Dillenseger, and Xavier Etchevers. A self-scalable load injection service. Software—Practice and Experience, 45(5):613– 632, May 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thrampoulidis:1997:ROL

[TDH97] K. Thrampoulidis, N. Diamantopoulos, and E. Housos. REDOM: An OO language to define and on-line manipulate regulations in the resource (re)Scheduling problem. Software— Practice and Experience, 27(10):1135–1161, October 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7263;   http://www3.

                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=7263&
                                        PLACEBO=IE.pdf.

Tripathi:1990:SNA

[TE90] Anand Tripathi and Zine Attia El Hili. Synchronization of nested atomic actions. Software—Practice and Experience, 20(8):773– 798, August 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Taschini:1999:SEC

[TEBK99] Stefano Taschini, Markus Emmenegger, Henry Baltes, and Jan G. Korvink. Smart enumeration in C++: virtual construction, message dispatching and tables. Software—Practice and Experience, 29(1):67– 76, January 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=55002655;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=55002655&PLACEBO=IE.
                                        pdf.

Terrasa:2008:LPT

[TEGF08] Andrés Terrasa, Agustín Espinosa, and Ana García-Fornes. Lightweight POSIX tracing. Software—Practice and Experience, 38(5): 447–469, April 25, 2008.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tennent:1978:ALT

[Ten78] Robert D. Tennent. Another look at type compatibility in Pascal. Software—Practice and Experience, 8(4):429–437, July/ August 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tennent:1982:TEB

[Ten82] Robert D. Tennent. Two examples of block structuring. Software—Practice and Experience, 12(4):385– 392, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tennent:1985:CAI

[Ten85] Robert D. Tennent. A comparison of the ANSI and ISO Pascal Standards. Software—Practice and Experience, 15(8):821– 822, August 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Terry:1986:MKS

[Ter86] P. D. Terry. A Modula-2 kernel for supporting monitors. Software—Practice and Experience, 16(5):457– 472, May 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Theaker:1979:AMO

[TF79a] C. J. Theaker and G. R. Frank. An assessment of the MUSS operating system. Software—Practice and Experience, 9(8):657– 670, August 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Theaker:1979:MPO

[TF79b] C. J. Theaker and G. R. Frank. MUSS: a portable operating system. Software —Practice and Experience, 9(8):633–643, August 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Taghizadeh:2022:MBD

[TGAS22] Jaber Taghizadeh, Mostafa Ghobaei-Arani, and Ali Shahidinejad. A metaheuristicbased data replica placement approach for dataintensive IoT applications in the fog computing environment. Software—Practice and Experience, 52(2): 482–505, February 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tse:2015:EFS

[TGC15] T. H. Tse, Arnaud Gotlieb, and Zhenyu Chen. Editorial: Focus section on quality software. Software —Practice and Experience, 45(7):873–874, July 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thomas:2008:DHF

[TGCF08] Gaël Thomas, Nicolas Geoffray, Charles Clément, and Bertil Folliot. Designing highly flexible virtual machines: the Jn-JVM experience. Software —Practice and Experience, 38(15):1643–1675, December ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tremblay:2008:OGE

[TGPS08] G. Tremblay, F. Guérin, A. Pons, and A. Salah. Oto, a generic and extensible tool for marking programming assignments. Software—Practice and Experience, 38(3):307– 333, March ??, 2008. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tuynman:1986:DRT

[TH86] F. Tuÿnman and L. O.

Hertzberger. A distributed real-time operating system. Software—Practice and Experience, 16(5):425– 441, May 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tavanapong:2001:DIV

[TH01] Wallapak Tavanapong and Kien A. Hua. Design and implementation of a video browsing system for the Internet. Software—Practice and Experience, 31(5):471– 503, April 25, 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
77502754/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=77502754&PLACEBO=IE.
pdf.

Thalmann:1984:IDV

[Tha84] Daniel Thalmann. An interactive data visualization system. Software—Practice and Experience, 14(3):277– 290, March 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thesen:1977:END

[The77] Abne Thesen. The evolution of a new discrete event simulation language for inexperienced users (WIDES). Software—Practice and Experience, 7(4): 519–533, July 1977. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tarhio:2017:TBA

[THG17] Jorma Tarhio, Jan Holub, and Emanuele Giaquinta. Technology beats algorithms (in exact string matching). Software— Practice and Experience, 47(12):1877–1885, December 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thimbleby:1980:LRP

[Thi80] Harold Thimbleby. Leave and recall: Primitives for procedural programming. Software—Practice and Experience, 10(2):127– 134, February 1980. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thimbleby:1987:DTI

[Thi87] Harold Thimbleby. The design of a terminal independent package. Software —Practice and Experience, 17(5):351–367, May 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thimbleby:1989:USI

[Thi89] Harold Thimbleby. Using sentinels in insert sort. Software—Practice and Experience, 19(3):303– 307, March 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thirion:1993:CIP

[Thi93] Bernard Thirion. Construction of an interactive programming environment for control of theatrical-set effects via object-oriented methods. Software—Practice and Experience, 23(6): 655–675, June 1993. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thimbleby:1996:ECA

[Thi96] Harold Thimbleby. An equivalence class algorithm for drawing autostereograms. Software—Practice and Experience, 26(3): 309–325, March 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16792.

Thiemann:1997:DSD

[Thi97] Peter J. Thiemann. Drawing syntax diagrams in Haskell. Software—Practice and Experience, 27(8): 967–982, August 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7327;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7327&
PLACEBO=IE.pdf.

Thimbleby:1999:CJ

[Thi99]  Harold Thimbleby.  A

critique of Java. Software—Practice and Experience, 29(5):457–478, April 25, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55003857;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55003857&PLACEBO=IE.
pdf.

Thimbleby:2003:DCP

[Thi03a] Harold Thimbleby. The directed Chinese Postman Problem. Software—Practice and Experience, 33 (11):1081–1096, September 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thimbleby:2003:ECP

[Thi03b] Harold Thimbleby. Explaining code for publication. Software—Practice and Experience, 33(10): 975–1001, August 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thimbleby:2012:HPI

[Thi12] Harold Thimbleby. Heedless programming: ignoring detectable error is a widespread hazard. Software—Practice and Experience, 42(11):1393–1407, November 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thomson:1974:BRB

[Tho74] N. D. Thomson. Book review: APL for management, Theodore J. Mock and Miklos A. Vasarhelyi, Wiley—Becker and Hayes, Los Angeles, 1972. No. of pages: 265. Price: $6.95. Software—Practice and Experience, 4(1):111– 112, January/March 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thomas:1977:BRB

[Tho77] M. Thomas. Book review: An introduction to the uses of computers, Murray Laver, Cambridge University Press, 1976. No. of pages: 194. Price £2.95. Software—Practice and Experience, 7(4):541– 542, July 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thorelli:1978:MSC

[Tho78]  Lars-Erik Thorelli.  A

monitor for small computers. Software—Practice and Experience, 8(4):439– 450, July/August 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Todter:1995:PPC

[THS95] Kai Tödter, Carsten Hammer, and Werner Struckmann. PARC++: a parallel C++. Software—Practice and Experience, 25(6): 623–636, June 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tichy:1985:RSV

[Tic85] Walter F. Tichy. RCS:

a system for version control. Software—Practice and Experience, 15(7):637– 654, July 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Taneja:2019:SMM

[TJB^{+}19] Mohit Taneja, Nikita Jalodia, John Byabazaire, Alan Davy, and Cristian Olariu. SmartHerd management: a microservices-based fog computing-assisted IoT platform towards datadriven smart dairy farming. Software—Practice and Experience, 49(7): 1055–1078, July 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Teperman:1972:FE

[TK72a] A. Teperman and Jacob Katzenelson. A format editor. Software—Practice and Experience, 2(3):219– 230, July/September 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thomas:1972:CQP

[TK72b] R. E. Thomas and P. Kent. Control of queues in a permissive society. Software—Practice and Experience, 2(1):79–91, January/ March 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thomas:2009:EEW

[TK09] Richard C. Thomas and Amela Karahasanović. Experience with an extensible workspace for analysis of low-level usage data. Software—Practice and Experience, 39(14):1185–1213, September 25, 2009. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tanenbaum:1978:GSP

[TKB78] Andrew S. Tanenbaum, Paul Klint, and Wim Bohm. Guidelines for software portability. Software —Practice and Experience, 8(6):681–698, November/ December 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Taiani:2009:CDR

[TKF09] François Taïani, Marc-Olivier Killijian, and Jean-Charles Fabre. CosmOpen: dynamic reverse engineering on a budget. How cheap observation techniques can be used to reconstruct complex multilevel behaviour. Software —Practice and Experience, 39(18):1467–1514, December 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tripathi:2007:ACR

[TKT^{+}07] Anand R. Tripathi, Devdatta Kulkarni, Harsha Talkad, Muralidhar Koka, Sandeep Karanth, Tanvir Ahmed, and Ivan Osipkov. Autonomic configuration and recovery in a mobile agent-based distributed event monitoring system. Software—Practice and Experience, 37(5): 493–522, April 25, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tomlinson:1985:PDF

[TKWW85] G. M. Tomlinson, D. Keeffe, I. C. Wand, and A. J. Wellings. The PULSE distributed file system. Software—Practice and Experience, 15(11):1087–1101, November 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tsang:1998:SVM

[TL98] Tony Tsang and Richard Lai. Specification and verification of multimedia synchronization scenarios using Time-Estelle. Software—Practice and Experience, 28(11):1185–1211, September 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=10007359;   http://
                                        www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=10007359&PLACEBO=IE.
                                        pdf.

Teodorov:2014:MDP

[TL14] Ciprian Teodorov and Loïc Lagadec. Model-driven physical-design automation for FPGAs: fast prototyping and legacy reuse. Software—Practice and Experience, 44(4):455–482, April 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Special issue on International Workshop on Smalltalk Technologies 2011.

Tsang:1998:EBB

[TLB98] Edward P. K. Tsang, Jin Li, and James M. Butler. EDDIE beats the bookies. Software—Practice and Experience, 28(10):1033– 1043, August 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=10007354;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=10007354&PLACEBO=IE.
pdf.

Tiberkak:2018:NAG

[TLB^{+}18] Allal Tiberkak, Tayeb Lemlouma, Abdelkader Belkhir, Ahmed Bouabdallah, and Abdelfetah Hentout. A novel approach for generic home emergency management and remote monitoring. Software —Practice and Experience, 48(4):761–774, April 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2553.

Tsai:2018:PPE

[TLC^{+}18] Meng-Han Tsai, Kuan-Chou Lai, Hsi-Ya Chang, Kuan Fu Chen, and Kuo-Chan Huang. Pewss: a platform of extensible workflow simulation service for workflow scheduling research. Software— Practice and Experience, 48(4):796–819, April 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2555.

Tzou:1993:DDE

[TLMP93] Shin-Yuan Tzou, Jyh-Jang Lim, Jai Menon, and David Palmer. A distributed development environment for embedded software. Software—Practice and Experience, 23(11):1235–1248, November 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Toyn:1995:CAZ

[TM95] Ian Toyn and John A. Mc-Dermid. CADiZ: An architecture for Z tools and its implementation. Software—Practice and Experience, 25(3):305–330, March 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tajalli:2014:IRA

[TM14] Hossein Tajalli and Nenad Medvidović. iDARE — a reference architecture for integrated software environments. Software—Practice and Experience, 44(3): 299–316, March 2014. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tziakouris:2022:MIF

[TMGR^{+}22] Giannis Tziakouris, Carlos Mera-Gómez, Francisco Ramírez, Rami Bahsoon, and Rajkumar Buyya. Market-inspired framework for securing assets in cloud computing environments. Software—Practice and Experience, 52 (9):2005–2025, September 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tiburski:2021:LVM

[TMJ^{+}21] Ramão T. Tiburski, Carlos R. Moratelli, Sérgio F. Johann, Everton de Matos, and Fabiano Hessel. A lightweight virtualization model to enable edge computing in deeply embedded systems. Software—Practice and Experience, 51 (9):1964–1981, September 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Takaoka:1982:IHV

[TMM82] Tadao Takaoka, Michael Maclean, and Bruce McKenzie. Introduction of history to variables. Software —Practice and Experience, 12(6):595–597, June 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tomar:2018:PQU

[TMS18] Pradeep Tomar, Rajesh Mishra, and Kavita Sheoran. Prediction of quality using ANN based on Teaching–Learning Optimization in componentbased software systems. Software—Practice and Experience, 48(4):896– 910, April 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2562.

Tripathi:1998:DRP

[TN98] A. R. Tripathi and T. Noonan. Design of a Remote Procedure Call system for object-oriented distributed programming. Software —Practice and Experience, 28(1):23–48, January 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61001666;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=61001666&PLACEBO=IE.
pdf.

Tian:2009:ADD

[TNGT09] Chen Tian, Vijay Nagarajan, Rajiv Gupta, and Sriraman Tallam. Automated dynamic detection of busy-wait synchronizations. Software—Practice and Experience, 39(11): 947–972, August 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Toroslu:2023:FWA

[Tor23] Ismail H. Toroslu. The Floyd–Warshall all-pairs shortest paths algorithm for disconnected and very sparse graphs. Software—Practice and Experience, 53(6):1287–1303, June 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Trammell:1992:GPD

[TP92] Carmen J. Trammell and J. H. Poore. A group process for defining local software quality: Field applications and validation experiments. Software—Practice and Experience, 22(8):603– 636, August 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tarhio:1997:SMD

[TP97] Jorma Tarhio and Hannu Peltola. String matching in the DNA alphabet. Software—Practice and Experience, 27(7):851– 861, July 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=7320;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=7320&
                                        PLACEBO=IE.pdf.

Tak:2003:ASS

[TP03] Sung Woo Tak and Eun Kyo Park. Adaptive secure software architecture for electronic commerce. Software —Practice and Experience, 33(14):1343–1357, November 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Torsun:1977:NIF

[TR77] I. S. Torsun and S. K. Robinson. Non-‘Interpretive’ FORTRAN input/output. Software—Practice and Experience, 7(2):205–213, March/April 1977. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tracz:1979:CPH

[Tra79a] William J. Tracz. Computer programming and the human thought process. Software—Practice and Experience, 9(2):127– 137, February 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tratner:1979:FAD

[Tra79b] M. Tratner. A fundamental approach to debugging. Software—Practice and Experience, 9(2):97– 99, February 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tofighy:2018:ECL

[TRGA18] Sajjad Tofighy, Ali A.

Rahmanian, and Mostafa Ghobaei-Arani. An ensemble CPU load prediction algorithm using a Bayesian information criterion and smooth filters in a cloud computing environment. Software—Practice and Experience, 48 (12):2257–2277, December 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tijero:2017:MPP

[TRO17] Héctor Pérez Tijero, Mario Aldea Rivas, and Daniel Medina Ortega. Multiprocessor platform for partitioned real-time systems. Software —Practice and Experience, 47(1):61–78, January 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Teles:2021:CSS

[TRRK21] Germanno Teles, Joel J.

P. C. Rodrigues, Ricardo A. L. Rabêlo, and Sergei A.

Kozlov. Comparative study of support vector machines and random forests machine learning algorithms on credit operation. Software—Practice and Experience, 51(12):2492–2500, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thomas:1981:DLD

[TS81] P. G. Thomas and B. E. Staff. A design language for the definition of a retrieval system interface for casual users of a relational database. Software—Practice and Experience, 11(5): 417–424, May 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Touati:1991:RMC

[TS91] Hervé Touati and Alan Jay Smith. Reducing and manipulating complex trace data. Software—Practice and Experience, 21(6):639– 655, June 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thalhammer:2002:RAD

[TS02] Thomas Thalhammer and Michael Schrefl. Realizing active data warehouses with off-the-shelf database technology. Software—Practice and Experience, 32(12):1193–1222, October 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
98516185/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=98516185&PLACEBO=IE.
pdf.

Tseng:1997:PPF

[Tse97] Chau-Wen Tseng. Performance of the prototype Fortran D compiler. Software—Practice and Experience, 27(7):763– 796, July 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7317;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7317&
PLACEBO=IE.pdf.

Tse:2013:EFS

[Tse13] T. H. Tse. Editorial: Focus section on program debugging. Software—Practice and Experience, 43(8):969– 970, August 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tsin:1982:EPP

[Tsi82] Yung H. Tsin. Extending the power of PASCAL’s external procedure mechanism. Software—Practice and Experience, 12(3):283– 292, March 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tolosa:2011:TSM

[TSMGD^{+}11] José Barranquero Tolosa, Oscar Sanjuán-Martínez, Vicente García-Díaz, B. Cristina Pelayo G-Bustelo, and Juan Manuel Cueva Lovelle. Towards the systematic measurement of ATL transformation models. Software—Practice and Experience, 41(7):789– 815, June 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Toffalini:2019:PSA

[TSO19] Flavio Toffalini, Jun Sun, and Martín Ochoa. Practical static analysis of context leaks in Android applications. Software—Practice and Experience, 49(2): 233–251, February 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thomas:2014:BFB

[TSZ14] Stephen W. Thomas, Richard T. Snodgrass, and Rui Zhang. Benchmark frameworks and τ Bench. Software—Practice and Experience, 44(9):1047–1075, September 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thomas:1974:IMG

[TT74] R. E. Thomas and D. C.

Toll. Implementation of a management game. Software—Practice and Experience, 4(2):171–179, April/ June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tharp:1982:PTS

[TT82] Alan L. Tharp and Kuo-Chung Tai. The practicality of text signatures for accelerating string searching. Software—Practice and Experience, 12(1):35– 44, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Expands Harrison’s work [Har71a].

Tamm:1996:LBV

[TT96] Boris Tamm and Kuldar Taveter. A list-based virtual machine for COBOL. Software—Practice and Experience, 26(12):1347– 1371, December 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16778.

Tan:2013:HFI

[TTC^{+}13] Yu Shyang Tan, Jiaqi Tan, Eng Siong Chng, Bu-Sung Lee, Jiaming Li, Susumu Date, Hui Ping Chak, Xiong Xiao, and Atsushi Narishige. Hadoop framework: impact of data organization on performance. Software—Practice and Experience, 43 (11):1241–1260, November 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tsai:1997:IDO

[TTH97] Lian-Jou Tsai, Shang-Rong Tsai, and Cheng-Liang Hou. An implementation of disk objects in the logical machine monitor. Software—Practice and Experience, 27(11):1335–1348, November 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7273;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7273&
PLACEBO=IE.pdf.

Tun:2009:SFE

[TTJ^{+}09] Thein Than Tun, Tim Trew, Michael Jackson, Robin Laney, and Bashar Nuseibeh. Specifying features of an evolving software system. Software —Practice and Experience, 39(11):973–1002, August 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Turner:1979:NIT

[Tur79] D. A. Turner. A new implementation technique for applicative languages. Software—Practice and Experience, 9(1):31–49, January 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Turner:2006:VFB

[Tur06] Kenneth J. Turner. Validating feature-based specifications. Software— Practice and Experience, 36(10):999–1027, August 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Turk:2022:SMS

[Tur22] Tomaz Turk. SDFunc:

Modular spreadsheet design with sheet-defined functions in Microsoft Excel. Software—Practice and Experience, 52(2):415–426, February 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Triantafyllos:1996:SRM

[TV96] George Triantafyllos and Stamatis Vassiliadis. Software reliability models for computer implementations — an empirical study. Software—Practice and Experience, 26(2):135–164, February 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61004292.

Terra:2009:DCL

[TV09] Ricardo Terra and Marco Tulio Valente. A dependency constraint language to manage object-oriented software architectures. Software—Practice and Experience, 39(12):1073–1094, August 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Terra:2015:RSR

[TVCB15] Ricardo Terra, Marco Tulio Valente, Krzysztof Czarnecki, and Roberto S. Bigonha. A recommendation system for repairing violations detected by static architecture conformance checking. Software—Practice and Experience, 45(3): 315–342, March 2015. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thorgeirsson:2021:ESA

[TVSG21] Adam Thor Thorgeirsson, Moritz Vaillant, Stefan Scheubner, and Frank Gauterin. Evaluating system architectures for driving range estimation and charge planning for electric vehicles. Software—Practice and Experience, 51(1): 72–90, January 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Terblanche:2016:OBE

[TW16] Chamonix Terblanche and Pornpit Wongthongtham. Ontology-based employer demand management. Software—Practice and Experience, 46(4):469–492, April 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tsuji:1988:SFP

[TWI88] Tatsuo Tsuji, Katsumasa Watanabe, and Atsushi Ikehata. Structured Fortran preprocessors generating optimized output. Software—Practice and Experience, 18(5):427–442, May 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tsai:2013:OPS

[TWJ^{+}13] Wei-Tek Tsai, Budan Wu, Zhi Jin, Yu Huang, and Wu Li. Ontology patterns for service-oriented software development. Software—Practice and Experience, 43(7):867–883, July 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thekkath:1994:TFS

[TWL94] Chandramohan A. Thekkath, John Wilkes, and Edward D. Lazowska. Techniques for file system simulation. Software—Practice and Experience, 24(11): 981–999, November 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Toegl:2012:SSJ

[TWNH12] Ronald Toegl, Thomas Winkler, Mohammad Nauman, and Theodore W. Hong. Specification and standardization of a Java Trusted Computing API. Software—Practice and Experience, 42(8):945–965, August 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tao:2022:FPH

[TWS^{+}22] Yudong Tao, Tianyi Wang, Anchen Sun, Shahid S. Hamid, Shu-Ching Chen, and Mei-Ling Shyu. Florida public hurricane loss model: Software system for insurance loss projection. Software—Practice and Experience, 52(7):1736–1755, July 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tian:2018:PCA

[TXHL18] Donghai Tian, Xi Xiong, Changzhen Hu, and Peng Liu. A policy-centric approach to protecting OS kernel from vulnerable LKMs. Software—Practice and Experience, 48 (6):1269–1284, June 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2576.

Triance:1980:ESL

[TY80] J. M. Triance and J. F. S.

Yow. Experiences with a schematic logic preprocessor. Software—Practice and Experience, 10(10): 791–800, October 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tien:2014:EOS

[TY14] Tsan-Rong Tien and Yi-Ping You. Enabling OpenCL support for GPGPU in kernel-based virtual machine. Software—Practice and Experience, 44(5):483– 510, May 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Urra:2019:HSF

[UCCPM19] Enrique Urra, Claudio Cubillos, Daniel Cabrera-Paniagua, and Rafael Mellado. hMod: a software framework for assembling highly detailed heuristics algorithms. Software— Practice and Experience, 49(6):971–994, June 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Urgun:2007:IMC

[UDS^{+}07] Bedirhan Urgun, Curtis E. Dyreson, Richard T. Snodgrass, Jessica K. Miller, Nick Kline, Michael D. Soo, and Christian S. Jensen. Integrating multiple calendars using τ Zaman. Software—Practice and Experience, 37(3):267–308, March 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Urbieta:2018:AIV

[UFR18] M. Urbieta, D. Frajberg, and G. Rossi. Assessing the impact of volatile functionality removal in web applications: Modeldriven vs code-based approaches. Software—Practice and Experience, 48(1): 11–43, January 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Urban:1999:IEU

[UFS99] Susan D. Urban, Ling Fu, and Jami J. Shah. The implementation and evaluation of the use of CORBA in an engineering design application. Software— Practice and Experience, 29(14):1313–1338, December 10, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
67500436/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=67500436&PLACEBO=IE.
pdf.

Ukelson:1991:CSU

[UGBW91] Jacob P. Ukelson, John D.

Gould, Stephen J. Boies, and Charles Wiecha. Case study: Using ITS style tools to implement IBM’s CUA-2 user interface style. Software—Practice and Experience, 21(12):1265– 1288, December 1991. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Uzun:2014:EEW

[UGK^{+}14] Erdinç Uzun, Edip Serdar Güner, Yılmaz Kılıçaslan, Tarık Yerlikaya, and Hayri Volkan Agun. An effective and efficient Web content extractor for optimizing the crawling process. Software—Practice and Experience, 44(10):1181–1199, October 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Umemura:1991:FPN

[Ume91] Kyoji Umemura. Floatingpoint number LISP. Software—Practice and Experience, 21(10):1015–1026, October 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Utture:2019:ELS

[UN19] Akshay Utture and V Krishna Nandivada. Efficient lock-step synchronization in task-parallel languages. Software—Practice and Experience, 49(9):1379– 1401, September 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

USENIX:1985:SCP

[USE85] USENIX Association, editor. Summer conference proceedings, Portland 1985: June 11–14, 1985, Portland, Oregon USA. USENIX Association, P.O. Box 7, El Cerrito 94530, CA, USA, 1985. LCCN QA76.8.U65 U8 1985.

Ushijima:1977:SEP

[Ush77] Kazuo Ushijima. Steps to an efficient program for floating-point summation. Software—Practice and Experience, 7(6):759– 769, November/December 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Uzun:2019:ACA

[UT19] Burak Uzun and Bedir Tekinerdogan. Architecture conformance analysis using model-based testing: a case study approach. Software—Practice and Experience, 49(3):423– 448, March 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Uh:1999:EEI

[UW99] Gang-Ryung Uh and David B.

Whalley. Effectively exploiting indirect jumps. Software—Practice and Experience, 29(12):1061– 1101, October 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
66000404/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=66000404&PLACEBO=IE.
pdf.

Uh:2005:CTE

[UWW^{+}05] Gang-Ryung Uh, Yuhong Wang, David Whalley, Sanjay Jinturkar, Yunheung Paek, Vincent Cao, and Chris Burns. Compiler transformations for effectively exploiting a zero overhead loop buffer. Software—Practice and Experience, 35(4):393–412, April 10, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Unlu:2022:TWP

[ÜY22] Hüseyin Ünlü and Yeliz Yesilada. Transcoding web pages via stylesheets and scripts for saving energy on the client. Software—Practice and Experience, 52(4): 984–1003, April 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Valentine:1976:BRBb

[Val76a] S. H. Valentine. Book review: Cobol Support Packages: Programming and Productivity Aids, Stanley M. Naftaly, Bruce G. Johnson and Michael C. Cohen, Wiley-Inter-science, New York, 1972. No. of pages:

182. Price: £9.45. Software—Practice and Experience, 6(2):285, April 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Valentine:1976:BRBa

[Val76b] S. H. Valentine. Book review: PL/I In easy stages, a programmed learning textbook, J. S. Roper, Elek Science, 1973. No. of pages:

240. Price: £2. Software— Practice and Experience, 6 (1):135–136, January 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Valentine:1977:BRBa

[Val77a] S. H. Valentine. Book review: Machine oriented higher level languages: Proceedings of IFIP working conference held in Trondheim, Norway, August 1973, edited by W. L. van der Poel and L. A. Maarssen, North Holland, 1974. no. of pages:

535. price: £12.30. ISBN: 0-7204-2802-5. Software— Practice and Experience, 7 (1):150–151, January 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Valentine:1977:BRBb

[Val77b] S. H. Valentine. Book review: Minicomputer forum. Online conferences, 1975. no of pages: 600. Price: £25. Software— Practice and Experience, 7 (2):300, March 1977. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Valentine:1978:BRB

[Val78] S. H. Valentine. Book review: Minicomputer software, Proceedings of IFIP Working Conference held at Lake Balaton, Hungary, September 1975. Edited by J. R. Bell and C. G. Bell, North Holland Publishing Co., Amsterdam, 1976. No. of pages: 333. Price: £17.85. ISBN: 0-7204-0437-1. Software— Practice and Experience, 8 (3):377, May 1978. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Valentine:1979:BRB

[Val79] S. H. Valentine. Book review: Computer programming methodology, W. M. Turski, Heyden, 1978. No. of pages: 208. Price: £14. Software—Practice and Experience, 9(3):253– 254, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Valentine:1980:BRB

[Val80] S. H. Valentine. Book review: Anatomy of LISP, John R. Allen, McGraw Hill, 1978. No. of pages:

446. Price: £11.20. Software—Practice and Experience, 10(11):951, November 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Valdorf:1984:DDP

[Val84] G. Valdorf. Dedicated, distributed and portable operating systems: a structuring concept. Software—Practice and Experience, 14(11):1079–1093, November 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Valois:2000:ISS

[Val00] John D. Valois. Introspective sorting and selection revisited. Software—Practice and Experience, 30(6): 617–638, May 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
71007583/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=71007583&PLACEBO=IE.
pdf.

VanTilborg:1982:ELG

[Van82] André M. Van Tilborg. Executing large graphics programs with a small computer. Software—Practice and Experience, 12(10): 915–927, October 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

VanWyk:1986:AGP

[Van86] Christopher J. Van Wyk. AWK as glue for programs. Software—Practice and Experience, 16(4):369– 388, April 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

VanWyk:1992:AEC

[Van92] Christopher J. Van Wyk. Arithmetic equality constraints as C++ statements. Software—Practice and Experience, 22(6):467– 494, June 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vasic:2017:ASC

[VAP^{+}17] Valter Vasić, Aleksandar Antonić, Krešimir Pripužić, Miljenko Mikuc, and Ivana Podnar Žarko. Adaptable secure communication for the Cloud of Things. Software—Practice and Experience, 47(3):489– 501, March 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Varley:1993:PEL

[Var93] Dominic A. Varley. Practical experience of the limitations of gprof. Software —Practice and Experience, 23(4):461–463, April 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vaucher:1979:SER

[Vau79] Jean G. Vaucher. Sequence error recovery considered misleading. Software— Practice and Experience, 9(11):925–929, November 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vaucher:1980:PPT

[Vau80] Jean G. Vaucher. Prettyprinting of trees. Software —Practice and Experience, 10(7):553–561, July 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vaucher:1989:RMP

[Vau89] Jean G. Vaucher. Record manipulation in Prolog. Software—Practice and Experience, 19(8):801–807, August 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vitek:2001:CTJ

[VB01] Jan Vitek and Boris Bokowski. Confined types in Java. Software—Practice and Experience, 31(6): 507–532, May 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
78003101/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=78003101&PLACEBO=IE.
pdf.

Vouillon:2014:BJJ

[VB14] Jérôme Vouillon and Vincent Balat. From bytecode to JavaScript: the Js of ocaml compiler. Software—Practice and Experience, 44(8):951–972, August 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

VanRenesse:1998:BAS

[VBH^{+}98] Robbert Van Renesse, Ken Birman, Mark Hayden, Alexey Vaysburd, and David Karr. Building adaptive systems using Ensemble. Software—Practice and Experience, 28(9):963– 979, July 25, 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1818;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1818&
PLACEBO=IE.pdf.

Venkateswaran:2022:ATS

[VBS22] Sreekrishnan Venkateswaran, Adwait Bauskar, and Santonu Sarkar. Architecture of a time-sensitive provisioning system for cloudnative software. Software—Practice and Experience, 52(5):1170–1198, May 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vongsathorn:1990:SAD

[VC90] Paul Vongsathorn and Scott D. Carson. A system for adaptive disk rearrangement. Software—Practice and Experience, 20(3):225– 242, March 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vardanega:2002:ESR

[VC02] Tullio Vardanega and Gert Caspersen. Engineering software reuse for on-board embedded real-time systems. Software—Practice and Experience, 32(3):233– 264, March 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
89011426/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=89011426&PLACEBO=IE.
pdf.

Venigalla:2021:CAP

[VC21] Akhila Sri Manasa Venigalla and Sridhar Chimalakonda. On the comprehension of application programming interface usability in game engines. Software—Practice and Experience, 51(8): 1728–1744, August 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vanDelft:1999:JES

[vD99]  André van Delft.  A

Java extension with support for dimensions. Software—Practice and Experience, 29(7):605–616, June 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=62003574;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=62003574&PLACEBO=IE.
pdf.

vandenBrand:2000:EAT

[vdBdJKO00] M. G. J. van den Brand, H. A. de Jong, P. Klint, and P. A. Olivier. Efficient annotated terms. Software—Practice and Experience, 30(3):259–291, March 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/

71001776/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=71001776&PLACEBO=IE.
pdf.

vandenBos:1977:CFT

[vdBT77] Jan van den Bos and Hendrik-Jan Thomassen. COMLNK — a file transport and job entry utility for a communication link. Software—Practice and Experience, 7(2):173– 177, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vonDincklage:2011:IPA

[vDD11] Daniel von Dincklage and Amer Diwan. Integrating program analyses with programmer productivity tools. Software—Practice and Experience, 41(7):817– 840, June 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vismara:2000:ESG

[VDG^{+}00] Luca Vismara, Giuseppe Di Battista, Ashim Garg, Giuseppe Liotta, Roberto Tamassia, and Francesco Vargiu. Experimental studies on graph drawing algorithms. Software—Practice and Experience, 30(11):1235–1284, September 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72516933/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72516933&PLACEBO=IE.
pdf.

vanderHoek:2003:SRM

[vdHW03] André van der Hoek and Alexander L. Wolf. Software release management for component-based software. Software—Practice and Experience, 33(1):77– 98, January 2003. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
101519987/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=101519987&PLACEBO=
IE.pdf.

vanderMeer:2013:ARI

[vdMF13] Jeroen van der Meer and Flavius Frasincar. Automatic review identification on the Web using pattern recognition. Software—Practice and Experience, 43(12):1415–1436, December 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vincenzi:2006:EST

[VDMW06] A. M. R. Vincenzi, M. E.

Delamaro, J. C. Maldonado, and W. E. Wong. Establishing structural testing criteria for Java bytecode. Software—Practice and Experience, 36(14): 1513–1541, November 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vanderPloeg:2014:DNL

[vdP14] Atze van der Ploeg. Drawing non-layered tidy trees in linear time. Software—Practice and Experience, 44(12):1467–1484, December 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vandeRiet:1979:PEB

[vdRW79] Reind P. van de Riet and R. Wiggers. Practice and experience with BASIS: an interactive programming system for introductory courses in informatics. Software—Practice and Experience, 9(6):463– 476, June 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vanDeursen:2004:SMA

[vDV04] Arie van Deursen and Joost Visser. Source model analysis using the JJTraveler visitor combinator framework. Software—Practice and Experience, 34(14): 1345–1379, November 25, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vanderWalt:2017:ERF

[vdWCB17] Paul van der Walt, Charles Consel, and Emilie Balland. Experience reports: Frameworks compiled from declarations: a language-independent approach. Software—Practice and Experience, 47(5):741– 762, May 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Venstermans:2006:BVB

[VED06] Kris Venstermans, Lieven Eeckhout, and Koen De Bosschere. 64-bit versus 32-bit Virtual Machines for Java. Software—Practice and Experience, 36(1):1– 26, January 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vella:1985:CSG

[Vel85]  A. D. Vella.  Compil-

ers — some good, some bad! Software—Practice and Experience, 15(10): 991–996, October 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vella:1988:BRB

[Vel88] C. A. Vella. Book review: X.25 explained — protocols for packet switching networks, R. J. Deasington, Ellis Honvood, 1986. ISBN 0-7458-0110-2. No. of pages: 125. Price: £16.50. Software—Practice and Experience, 18(5): 494, May 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

VanVliet:1985:ET

[VG85] J. C. Van Vliet and H. M.

Gladney. An evaluation of tagging. Software—Practice and Experience, 15(9): 823–837, September 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vanGurp:2001:DIE

[vGB01] J. van Gurp and J. Bosch.

Design, implementation and evolution of object oriented frameworks: concepts and guidelines. Software—Practice and Experience, 31(3):277–300, March 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
76507640/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?

                                        ID=76507640&PLACEBO=IE.
                                        pdf.

Vogel:2021:PHL

[VGF21] Adriano Vogel, Dalvan Griebler, and Luiz Gustavo Fernandes. Providing high-level self-adaptive abstractions for stream parallelism on multicores. Software—Practice and Experience, 51(6):1194–1217, June 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vanGurp:2010:CPR

[vGPB10] Jilles van Gurp, Christian Prehofer, and Jan Bosch. Comparing practices for reuse in integration-oriented software product lines and large open source software projects. Software—Practice and Experience, 40(4): 285–312, April 10, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

VanDrunen:2004:ABP

[VH04] Thomas VanDrunen and Antony L. Hosking. Anticipationbased partial redundancy elimination for static single assignment form. Software—Practice and Experience, 34(15):1413–1439, December 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vanHoeve:1987:OOA

[vHE87] Frans van Hoeve and Rolf Engmann. An objectoriented approach to application generation. Software —Practice and Experience, 17(9):623–645, September 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vanHorebeek:1988:EHM

[vHLB^{+}88] Ivo van Horebeek, Johan Lewi, Eddy Bevers, Luc Duponcheel, and Willy van Puymbroeck. An exception handling method for constructive algebraic specifications. Software—Practice and Experience, 18(5): 443–458, May 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

VanderZanden:2005:LLP

[VHM^{+}05] Bradley T. Vander Zanden, Richard Halterman, Brad A. Myers, Rob Miller, Pedro Szekely, Dario A. Giuse, David Kosbie, and Rich McDaniel. Lessons learned from programmers’ experiences with one-way constraints. Software— Practice and Experience, 35(13):1275–1298, November 10, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vildosola:1980:LE

[Vil80] Fernando Vildosola. Letter to the editor. Software—Practice and Experience, 10(5):419, May 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Visvalingam:1976:ICD

[Vis76] M. Visvalingam. Indexing with coded deltas— a data compaction technique. Software—Practice and Experience, 6(3):397– 403, July/September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vanKatwijk:1987:ATO

[vK87] J. van Katwijk. Addressing types and objects in Ada. Software—Practice and Experience, 17(5):319– 343, May 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vlietstra:1973:ASS

[VL73] J. Vlietstra and J. B. M. Lucassen. AEDBAR: a software system for designing and analysing planar bar mechanisms. Software—Practice and Experience, 3(1):29–42, January/ March 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vanMeurs:1977:IU

[vMC77] John van Meurs and Ernst Lopes Cardazo. Interfacing the user. Software—Practice and Experience, 7(1):85–93, January/March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vilela:1997:PGV

[VMJ97] Plínio R. S. Vilela, José C.

Maldonado, and Mario Jino. Program graph visualization. Software —Practice and Experience, 27(11):1245–1262, November 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7269;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7269&
PLACEBO=IE.pdf.

Verma:2024:SEE

[VMY24] Prateek Verma, Ashish Kumar Maurya, and Rama Shankar Yadav. A survey on energyefficient workflow scheduling algorithms in cloud computing. Software— Practice and Experience, 54(5):637–682, May 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Venugopal:2008:DRB

[VNGB08] Srikumar Venugopal, Krishna Nadiminti, Hussein Gibbins, and Rajkumar Buyya. Designing a resource broker for heterogeneous grids. Software —Practice and Experience, 38(8):793–825, July 10, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Varghese:2020:NGC

[VNLB20] Blesson Varghese, Marco Netto, Ignacio M. Llorente, and Rajkumar Buyya. New generation cloud computing. Software—Practice and Experience, 50(6):803– 804, June 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vo:1996:VGE

[Vo96] Kiem-Phong Vo. Vmalloc: A general and efficient memory allocator. Software—Practice and Experience, 26(3):357–374, March 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16795.

Vo:1997:CCD

[Vo97] Kiem-Phong P. Vo. CDT: a container data type library. Software—Practice and Experience, 27(10):1177– 1197, October 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7265;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7265&
PLACEBO=IE.pdf.

Vo:2000:DMA

[Vo00] Kiem-Phong Vo. The discipline and method architecture for reusable libraries. Software—Practice and Experience, 30(2):107– 128, February 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
69503460/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=69503460&PLACEBO=IE.
pdf.

vanOmmering:2003:HCS

[vO03] Rob van Ommering. Horizontal communication: a style to compose control software. Software—Practice and Experience, 33 (12):1117–1150, October 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Voros:1984:CCO

[Vör84]  G. Vörös.  Calculation

of co-ordinates for optimal hatching. Software—Practice and Experience, 14(2): 169–176, February 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vivanco:2005:SCJ

[VP05] Rodrigo A. Vivanco and Nicolino J. Pizzi. Scientific computing with Java and C++: a case study using functional magnetic resonance neuroimages. Software—Practice and Experience, 35(3):237–254, March 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vanReeuwijk:1992:TCG

[vR92] C. van Reeuwijk. Tm: a code generator for recursive data structures. Software —Practice and Experience, 22(10):899–908, October 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vidakovic:2006:GCD

[VR06] Jovana Vidaković and Miloš Racković. Generating content and display of library catalogue cards using XML technology. Software—Practice and Experience, 36(5):513– 524, April 25, 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vilas:2006:MOC

[VRC^{+}06] Ana Fernández Vilas, Rebeca P. Díaz Redondo, Manuel Ramos Cabrer, José J. Pazos Arias, Alberto Gil Solla, Jorge García Duque, Martín López Nores, and Yolanda Blanco Fernández. MHP-OSGi convergence: a new model for open residential gateways. Software—Practice and Experience, 36(13): 1421–1442, November 10, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vanRenesse:1989:PAD

[vRvST89] Robert van Renesse, Hans van Staveren, and Andrew S. Tanenbaum. The performance of the Amoeba distributed operating system. Software—Practice and Experience, 19(3):223– 234, March 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Verma:1980:MPF

[VS80] S. B. Verma and Maithili Sharan. Multiple precision floating-point computation in FORTRAN. Software—Practice and Experience, 10(3):163–173, March 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vernon:1988:VVI

[VS88] David Vernon and Giulio Sandini. VIS: a virtual image system for image-understanding research. Software—Practice and Experience, 18(5):395– 414, May 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Venkateswaran:2018:APD

[VS18] Sreekrishnan Venkateswaran and Santonu Sarkar. Architectural partitioning and deployment modeling on hybrid clouds. Software—Practice and Experience, 48(2):345–365, February 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2496.

Varshney:2020:CAS

[VS20] Prateeksha Varshney and Yogesh Simmhan. Characterizing application scheduling on edge, fog, and cloud computing resources. Software—Practice and Experience, 50(5):558–595, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vassiliades:1986:MTN

[VSB86] Synnove Vassiliades, Michael D. Sayers, and Jean M. Bacon. A monitor tool for a network based on the Cambridge Ring. Software —Practice and Experience, 16(7):671–687, July 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Valenzano:1993:RPP

[VSC93] Adriano Valenzano, R. Sisto, and Luigi Ciminiera. Rapid prototyping of protocols from LOTOS specifications. Software—Practice and Experience, 23(1):31– 54, January 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vogler:2017:ACB

[VSID17] Michael Vögler, Johannes M.

Schleicher, Christian Inzinger, and Schahram Dustdar. Ahab: a cloud-based distributed big data analytics framework for the Internet of Things. Software—Practice and Experience, 47(3): 443–454, March 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

VanBiljon:1987:RAP

[VSM87] W. R. Van Biljon, D. A.

Sewry, and M. A. Mulders. Register allocation in a pattern matching code generator. Software—Practice and Experience, 17(8):521– 531, August 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Verhelst:1984:PIP

[VV84] P. W. E. Verhelst and N. F. Verster. PEP: an interactive programming system with an ALGOL-like programming language. Software—Practice and Experience, 14(2):119–133, February 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Veerman:2006:CMD

[VV06] Niels Veerman and Ernst-Jan Verhoeven. Cobol minefield detection. Software—Practice and Experience, 36(14):1605–1642, November 25, 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Villarrubia:2024:DAE

[VVG^{+}24] Carlos Villarrubia, Juan Manuel Vara, David Granada, Cristian Gómez-Macías, and Francisco Javier Pérez-Blanco. DesignScrum — an agility educational resource powered by creativity. Software—Practice and Experience, 54(6):985–1009, June 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vardanega:1999:SPC

[VvK99] T. Vardanega and J. van Katwijk. A software process for the construction of predictable on-board embedded real-time systems. Software—Practice and Experience, 29(3):235– 266, March 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55000312;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55000312&PLACEBO=IE.
pdf.

Vaughan:1991:PID

[VWB91] Jon Vaughan, Duncan Whyatt, and Graham Brookes. A parallel implementation of the Douglas– Peucker line simplification algorithm. Software—Practice and Experience, 21(3):331–336, March 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Vines:1998:CTC

[VZ98] Phil Vines and Justin Zobel. Compression techniques for Chinese text. Software—Practice and Experience, 28(12):1299– 1314, October 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=1764;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=1764&
                                        PLACEBO=IE.pdf.

White:1977:SSD

[WA77] John R. White and Richard K. Anderson. Supporting the structured development of complex PL/I software systems. Software— Practice and Experience, 7 (2):279–293, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wadler:1985:SLF

[Wad85] Philip Wadler. A simple language is also a functional language. Software —Practice and Experience, 15(2):219, February 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wadler:1987:FSS

[Wad87] Philip Wadler. Fixing some space leaks with a garbage collector. Software—Practice and Experience, 17(9): 595–608, September 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wagstaff:1978:LE

[Wag78] S. P. Wagstaff. Letter to the editor. Software— Practice and Experience, 8 (4):496, July 1978. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wu:2012:DSM

[WAH^{+}12] Yali Wu, Andrew A. Allen, Frank Hernandez, Robert France, and Peter J. Clarke. A domain-specific modeling approach to realizing user-centric communication. Software—Practice and Experience, 42(3): 357–390, March 2012. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Waite:1973:SMA

[Wai73a] W. M. Waite. A sampling monitor for applications programs. Software—Practice and Experience, 3(1):75–79, January/ March 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Waite:1973:GEG

[Wai73b] William M. Waite. Guest editorials: Guest editorial. Software—Practice and Experience, 3(3):195– 196, July/September 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Waite:1975:HDP

[Wai75] W. M. Waite. Hints on distributing portable software. Software—Practice and Experience, 5(3):295– 308, July/September 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Waite:1985:TTC

[Wai85]  W. M. Waite.  Treat-

ment of tab characters by a compiler. Software—Practice and Experience, 15(11):1121–1123, November 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Waite:1986:CLA

[Wai86] W. M. Waite. The cost of lexical analysis. Software —Practice and Experience, 16(5):473–488, May 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wainer:2002:CTD

[Wai02] Gabriel Wainer. CD++: a toolkit to develop DEVS models. Software—Practice and Experience, 32 (13):1261–1306, November 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/

98517571/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=98517571&PLACEBO=IE.
pdf.

Wainer:2007:DST

[Wai07] Gabriel Wainer. Developing a software toolkit for urban traffic modeling. Software—Practice and Experience, 37(13): 1377–1404, November 10, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Walker:1980:PNG

[Wal80] John Q. Walker II. Positioning nodes for general trees. Software—Practice and Experience, 10(7):553– 561, July 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wallis:1981:BRB

[Wal81a] Peter Wallis. Book review: Practice in software adaption and maintenance. Edited by R. Ebert, J. Luegger and R. Goecke, North-Holland Publishing Company, 1980. No. of pages: 455. Price: US $58.50. Software—Practice and Experience, 11(5): 532, May 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wallis:1981:DSM

[Wal81b] Peter J. L. Wallis. Designing storage management schemes for blockstructured languages. Software—Practice and Experience, 11(8):789–801, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wallis:1981:HTI

[Wal81c] Peter J. L. Wallis. Handling type information when compiling a language with user-defined types. Software—Practice and Experience, 11(2):167– 173, February 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wallis:1982:BRB

[Wal82] Peter Wallis. Book review: Conversational languages, I. Kupka and N. Wilsing, Wiley, Chichester, 1980. No. of pages: 117. Price: £9.45. Software—Practice and Experience, 12(1):101, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wallace:1983:DCR

[Wal83a] John J. Wallace. DMERT crash resistant file systems. Software—Practice and Experience, 13(4):385– 387, April 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wallace:1983:BRB

[Wal83b] M. G. Wallace. Book review: Language as a cognitive process, vol I: Syntax, Terry Winograd, Addison-Wesley, 1983. No. of pages: 640. ISBN 0-201-08571-2. Software— Practice and Experience, 13(10):976–977, October 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wallis:1983:BRB

[Wal83c] Peter Wallis. Book review: Problem solving with ADA, Brian Mayoh, Wiley, Chichester, 1982. no. of pages: viii + 233. Price: £10.75. Software—Practice and Experience, 13(2):198, February 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Walden:1984:AGM

[Wal84a] Kim Walden. Automatic generation of make dependencies. Software—Practice and Experience, 14(6): 575–585, June 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wallis:1984:BRB

[Wal84b] Peter Wallis. Book review: Software Engineering with Ada, Grady Booch, Benjamin/Cummings, 1983. No. of pages: xix + 502. Price: £14.20. Software —Practice and Experience, 14(6):603, June 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Walker:1986:IPP

[Wal86a] A. N. Walker. The Ilem Project: Pascal to Algol in easy stages I: Outward bound. Software—Practice and Experience, 16 (12):1097–1107, December 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wallis:1986:BRB

[Wal86b] Peter Wallis. Book review: Polymorphic programming language: Design and implementation, David M. Harland, Ellis Horwood Ltd., 1984. no. of pages: 251, price: £19.50. Software—Practice and Experience, 16(1): 99, January 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Walker:1990:NPA

[Wal90] John Q. Walker II. A node-positioning algorithm for general trees. Software —Practice and Experience, 20(7):685–705, July 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Welponer:2012:MRI

[WAML12] Michele Welponer, Luca Abeni, Guido Marchetto, and Renato Lo Cigno. Measuring and reducing the impact of the operating system kernel on endto-end latencies in synchronous packet switched networks. Software—Practice and Experience, 42 (11):1315–1330, November 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wand:1979:SIL

[Wan79]  I. C. Wand.  Systems

implementation languages and IRONMAN. Software—Practice and Experience, 9(10):853–878, October 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wand:1982:BRI

[Wan82] I. C. Wand. Book review: Introduction to realtime software design, S. T. Allworth, Macmillan, 1981. No. of pages: 140. Price £5.50. Software—Practice and Experience, 12(3):295– 296, March 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Warren:1980:LPC

[War80] David H. D. Warren. Logic programming and compiler writing. Software—Practice and Experience, 10 (2):97–125, February 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wassenberg:2012:LAS

[Was12] Jan Wassenberg. Lossless asymmetric single instruction multiple data codec. Software—Practice and Experience, 42 (9):1095–1106, September 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Watt:1986:ESD

[Wat86]  David A. Watt.  Exe-

cutable semantic descriptions. Software—Practice and Experience, 16(1):13– 43, January 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Waters:1989:ASM

[Wat89] Richard C. Waters. Automated software management based on structural models. Software— Practice and Experience, 19(10):931–955, October 1989. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Watson:2004:RMR

[Wat04] Bruce W. Watson. Reducing memory requirements in reachability-based finite automata operations. Software—Practice and Experience, 34(3):239–248, March 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Williams:1977:AHC

[WB77] M. Howard Williams and A. R. Bulmer. Array handling in Cobol compilers. Software—Practice and Experience, 7(4):469– 474, July/August 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Williams:1978:UFN

[WB78] M. Howard Williams and A. R. Bulmer. Use of a formal notation for static semantics in compiler design. Software—Practice and Experience, 8(5):579– 584, September/October 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Welsh:1979:PPA

[WB79] J. Welsh and D. W. Bustard. Pascal-Plus — another language for modular multiprogramming. Software—Practice and Experience, 9(11):947–957, November 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

White:1985:PPR

[WB85a] N. H. White and K. H. Bennett. PRTDS — A Pascal run-time diagnostics system. Software—Practice and Experience, 15 (11):1041–1056, November 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

White:1985:RTD

[WB85b] N. H. White and K. H. Bennett. Run-time diagnostics in Pascal. Software —Practice and Experience, 15(4):359–367, April 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wonsil:2023:RS

[WBA^{+}23] Joseph Wonsil, Nichole Boufford, Prakhar Agrawal, Christopher Chen, Tianhang Cui, Akash Sivaram, and Margo Seltzer. Reproducibility as a service. Software—Practice and Experience, 53(7): 1543–1571, July 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Woodall:2007:ISO

[WBB07] Philip Woodall, Pearl Brereton, and David Budgen. Investigating serviceoriented system performance: a systematic study. Software—Practice and Experience, 37(2):177–191, February 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Willkomm:2015:ERR

[WBB15] Johannes Willkomm, Christian Bischof, and H. Martin Bücker. Experience reports: RIOS: efficient I/O in reverse direction. Software—Practice and Experience, 45(10):1399–1427, October 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wassenberg:2022:VPP

[WBGS22] Jan Wassenberg, Mark Blacher, Joachim Giesen, and Peter Sanders. Vectorized and performanceportable quicksort. Software—Practice and Experience, 52(12):2684–2699, December 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Welsh:1991:DRL

[WBK91] Jim Welsh, Brad Broom, and Derek Kiong. A design rationale for a languagebased editor. Software— Practice and Experience, 21(9):923–948, September 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Weir:2020:ILT

[WBN^{+}20] Charles Weir, Ingolf Becker, James Noble, Lynne Blair, M. Angela Sasse, and Awais Rashid. Interventions for long-term software security: Creating a lightweight program of assurance techniques for developers. Software—Practice and Experience, 50(3): 275–298, March 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wu:2020:SII

[WBPR20] Hao Wu, Rongfang Bie, Charith Pereira, and Omer Rana. Special issue: Identification, information, and knowledge in the Internet of Things. Software—Practice and Experience, 50 (11):2011, November 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Witten:1982:PTD

[WBS82] Ian H. Witten, Mike Bonham, and Evelyne Strong. On the power of traps and diversions in a document preparation language. Software—Practice and Experience, 12(12):1119–1131, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wolf:1996:ECS

[WBV96] Lars C. Wolf, Wolfgang Burke, and Carsten Vogt. Evaluation of a CPU scheduling mechanism for multimedia systems. Software—Practice and Experience, 26(4):375–398, April 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16799.

Waite:1981:ASI

[WC81] W. M. Waite and Lynn Robert Carter. An analysis/ synthesis interface for Pascal compilers. Software —Practice and Experience, 11(8):769–787, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Waite:1985:CGP

[WC85] W. M. Waite and Lynn Robert Carter. The cost of a generated parser. Software—Practice and Experience, 15(3):221–237, March 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Weaver:1987:RTM

[WC87] Alfred C. Weaver and M. Alex Colvin. A realtime messaging system for token ring networks. Software—Practice and Experience, 17(12):885–897, December 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Watson:2004:SPC

[WC04] Bruce W. Watson and Loek Cleophas. SPARE Parts: a C++ toolkit for string pattern recognition. Software —Practice and Experience, 34(7):697–710, June 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Whaley:2008:AAC

[WC08] R. Clint Whaley and Anthony M. Castaldo. Achieving accurate and contextsensitive timing for code optimization. Software— Practice and Experience, 38(15):1621–1642, December ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wiseman:1972:RMP

[WCE^{+}72] N. E. Wiseman, C. J. Cheney, M. Etherton, J. O. Hiles, and H. U. Lemke. RAINBOW: a multi-purpose CAD system. Software—Practice and Experience, 2(4):359– 375, October/December 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2011:EIS

[WCK11] Ji Wang, Wing Kwong Chan, and Fei-Ching Kuo. Editorial: Introduction to the special issue for the 10th International Conference on Quality Software (QSIC 2010). Software— Practice and Experience, 41(10):997–998, September 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wu:2017:FLU

[WCS^{+}17] Eric Hsiaokuang Wu, Shumin Chuang, Chen-Yen Shih, Hao-Che Hsueh, Shih-Syuan Huang, and Hsiao-Ping Huang. A flexible and lightweight user-demand DRM system for multimedia contents over multiple portable device platforms. Software—Practice and Experience, 47(10): 1417–1441, October 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wei:2016:BAT

[WCsH16] Hsin-Wen Wei, Tseng-Yi Chen, and Tsan sheng Hsu. BASE: an assistant tool to precisely simulate energy consumption and reliability of energy-efficient storage systems. Software —Practice and Experience, 46(5):581–599, May 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2019:TAV

[WCT19] Jing V. Wang, Chi-Tsun Cheng, and Chi K. Tse. A thermal-aware VM consolidation mechanism with outage avoidance. Software —Practice and Experience, 49(5):906–920, May 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Weber:1987:OSE

[Web87] Manfred Weber. Operatingsystem enhancements for a fault-tolerant dual-processor structure for the control of an industrial process. Software—Practice and Experience, 17(5):345–350, May 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Weinberg:1972:PCS

[Wei72] Gerald M. Weinberg. Programming and compiling strategies for paging systems. Software—Practice and Experience, 2(2):165– 171, April/June 1972. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Weiser:1985:CWS

[Wei85]  Mark Weiser.  CWSH:

The windowing shell of the Maryland Window System. Software—Practice and Experience, 15(5):515– 519, May 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wells:1972:BRB

[Wel72] M. Wells. Book review:

Introduction to operating systems, A. J. T. Colin, Macdonald, London and American Elsevier, 1971. No. of pages: 120. Price: £1.75. Software—Practice and Experience, 2(4):408, October 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Welch:1978:SPM

[Wel78a] Jim Welch. Structured programming in macro assembly languages. Software —Practice and Experience, 8(3):371–376, May/June 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Welsh:1978:ERC

[Wel78b] J. Welsh. Economic range checks in Pascal. Software—Practice and Experience, 8(1):85–97, January/ February 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Welch:1983:PAR

[Wel83] P. H. Welch. Parallel assignment revisited. Software—Practice and Experience, 13(12):1175–1180, December 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wendt:1980:MPN

[Wen80] Siegfried Wendt. Modified Petri nets as flowcharts for recursive programs. Software—Practice and Experience, 10(11):935–942, November 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wentworth:1990:PCG

[Wen90] E. P. Wentworth. Pitfalls of conservative garbage collection. Software—Practice and Experience, 20(7):719– 727, July 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

West:1983:ORD

[Wes83] Vincent West. An optimizer for a relational database command language. Software—Practice and Experience, 13 (11):1005–1012, November 1983. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Wettstein:1977:ISO

[Wet77] H. Wettstein. The implementation of synchronizing operations in various environments. Software—Practice and Experience, 7(1):115–126, January/March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wetherell:1980:DCA

[Wet80] Charles Wetherell. Design considerations for array processing languages. Software—Practice and Experience, 10(4):265–271, April 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wexelblat:1975:LE

[Wex75] R. L. Wexelblat. Letters to the editor. Software— Practice and Experience, 5 (4):407, October 1975. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wexelblat:1978:LE

[Wex78] R. L. Wexelblat. Letter to the editor. Software —Practice and Experience, 8(5):642, September 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wexelblat:1981:LE

[Wex81a] R. L. Wexelblat. Letter to the editor. Software—Practice and Experience, 11(5):531, May 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wexelblat:1981:COF

[Wex81b] Richard L. Wexelblat. The consequences of one’s first programming language. Software—Practice and Experience, 11(7):733– 740, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wampler:1983:IGG

[WG83] Stephen B. Wampler and Ralph E. Griswold. The implementation of generators and goal-directed evaluation in Icon. Software —Practice and Experience, 13(6):495–518, June 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1989:OS

[WG89] Niklaus Wirth and Jürg Gutknecht. The Oberon system. Software—Practice and Experience, 19(9): 857–893, September 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Walker:1992:OCI

[WG92a] Kenneth Walker and Ralph E.

Griswold. An optimizing compiler for the Icon programming language. Software—Practice and Experience, 22(8):637–657, August 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wood:1992:UIA

[WG92b] Catherine A. Wood and Philip D. Gray. User interface-application communication in the Chimera User Interface Management System. Software— Practice and Experience, 22(1):63–84, January 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Whitehead:2004:WPD

[WG04] E. James Whitehead, Jr.

and Yaron Y. Goland. The WebDAV property design. Software—Practice and Experience, 34(2):135– 161, February 2004. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wu:2008:GDG

[WGM08] Hui Wu, Jeff Gray, and Marjan Mernik. Grammardriven generation of domainspecific language debuggers. Software—Practice and Experience, 38 (10):1073–1103, August ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Williamson:1984:CCS

[WH84] Ronald Williamson and Ellis Horowitz. Concurrent communication and synchronization mechanisms. Software—Practice and Experience, 14(2):135– 151, February 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wu:1997:CSC

[WH97] Pei-Chi C. Wu and Kuo-Chan C. Huang. Case studies on cache performance and optimization of programs with unit strides. Software—Practice and Experience, 27(2):167– 172, February 1997. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7285;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7285&
PLACEBO=IE.pdf.

Wilkie:1998:MCC

[WH98] F. G. Wilkie and B. Hylands. Measuring complexity in C++ application software. Software—Practice and Experience, 28(5):513–546, April 25, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1794;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1794&
PLACEBO=IE.pdf.

Watkins:2006:ETD

[WH06] Alison Watkins and Ellen M.

Hufnagel. Evolutionary test data generation: a comparison of fitness functions. Software—Practice and Experience, 36(1):95– 116, January 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Whaley:1972:FTF

[Wha72] Alfred D. Whaley. A failure tolerant filing system. Software—Practice and Experience, 2(3):287–291, July/ September 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Whalley:1993:TFI

[Wha93] David B. Whalley. Techniques for fast instruction cache performance evaluation. Software—Practice and Experience, 23(1):95– 118, January 1993. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2021:DGR

[WHH21] Zhijin Wang, Yaohui Huang, and Bingyan He. Dualgrained representation for hand, foot, and mouth disease prediction within public health cyber-physical systems. Software—Practice and Experience, 51 (11):2290–2305, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Whiddett:1983:DDS

[Whi83] R. J. Whiddett. Dynamic distributed systems. Software—Practice and Experience, 13(4):355–371, April 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Whiddett:1987:BRB

[Whi87] Dick Whiddett. Book review: Synchronization Of Parallel Programs, F. Andre, D. Herman and J.- P Verjus, North Oxford Academic, Oxford, 1985. Price: £18.00. No. of pages:

103. Software—Practice and Experience, 17(2):165– 166, February 1987. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wong:1998:ETS

[WHLM98] W. Eric Wong, Joseph R. Horgan, Saul London, and Aditya P. Mathur. Effect of test set minimization on fault detection effectiveness. Software—Practice and Experience, 28(4):347– 369, April 10, 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1785;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1785&
PLACEBO=IE.pdf.

Wong:2000:ADM

[WHS^{+}00] W. Eric Wong, Joseph R.

Horgan, Michael Syring, Wayne Zage, and Dolores Zage. Applying design metrics to predict faultproneness: a case study on a large-scale software system. Software—Practice and Experience, 30 (14):1587–1608, November 25, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
73500504/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=73500504&PLACEBO=IE.
pdf.

Wang:2023:VGE

[WHYC23] Sixuan Wang, Chen Huang, Dongjin Yu, and Xin Chen. VulGraB: Graphembedding-based code vulnerability detection with bi-directional gated graph neural network. Software —Practice and Experience, 53(8):1631–1658, August 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Woodman:1985:STC

[WI85] M. Woodman and D. C. Ince. A software tool for the construction and maintenance of structured analysis notation. Software—Practice and Experience, 15(11):1057–1072, November 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wichmann:1972:BRB

[Wic72a] B. A. Wichmann. Book review: The Alpha programming system, A. P. Yershov (Ed.), APIC Studies in Data Processing No. 7, Academic Press, London, 1971. No. of pages:

247. Price: £5.00. Software—Practice and Experience, 2(3):305, July 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wichmann:1972:NII

[Wic72b]  B. A. Wichmann.  A

note on the IFIP inputoutput procedures for Algol 60. Software—Practice and Experience, 2(4):403–404, October 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wichmann:1977:HCP

[Wic77] B. A. Wichmann. How to call procedures, or second thoughts on Ackermann’s function. Software—Practice and Experience, 7(3): 317–329, May/June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wichmann:1979:ID

[Wic79] B. A. Wichmann. Integer division. Software—Practice and Experience, 9(6): 507–508, June 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wichmann:1981:PBA

[Wic81] B. A. Wichmann. Has the program been altered? Software—Practice and Experience, 11(8):877– 879, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Widener:1990:XIC

[Wid90]  Glenn Widener.  The

X11 Inter-Client Communication Conventions Manual. Software—Practice and Experience, 20(S2): S109–S118, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wieczerzycki:1996:SRT

[Wie96] Waldemar Wieczerzycki. Software reusability through versions. Software—Practice and Experience, 26(8): 911–927, August 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16827.

Wijnstra:2005:CPF

[Wij05] Jan Gerben Wijnstra. Classifying product families using platform coverage and variation mechanisms. Software—Practice and Experience, 35(5):413– 444, April 25, 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilkes:1972:BRB

[Wil72] M. V. Wilkes. Book review: Advances in computers, Vol. 10, Eds.: F. L. Alt and M. Rubinoff, Guest Ed.: W. Freiberger, Academic Press, New York, 1971. No. of pages: 313. Price: £6.75. Software —Practice and Experience, 2(2):191–192, April 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilkes:1973:CMA

[Wil73] M. V. Wilkes. The Cambridge Multiple-Access System in retrospect. Software—Practice and Experience, 3(4):323–332, October/December 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilkes:1974:BRB

[Wil74a] Maurice V. Wilkes. Book review: Data base management, Infotech state of the art report 15, Ed. C. Boon and C. J. Bunyan, Infotech Information, Maidenhead, Berks, England, 1973. No. of pages:

665. Price: £40. Software—Practice and Experience, 4(3):301–302, July/ September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Williams:1974:HPI

[Wil74b] M. R. Williams. Heuristic procedures (if they work— leave them alone). Software—Practice and Experience, 4(3):237–240, July/ September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilkes:1976:BRB

[Wil76]  M. V. Wilkes.  Book

review: Data base systems, Infotech State of the art Report, Infotech Information Ltd., Maidenhead, 1975. No. of pages: 722. Price: £60. Software— Practice and Experience, 6 (4):582, October 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Williams:1979:LSA

[Wil79] M. Howard Williams. Long/short address optimization in assemblers. Software—Practice and Experience, 9(3):227–235, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilson:1980:PSH

[Wil80] I. R. Wilson. Pascal for school and hobby use. Software—Practice and Experience, 10(8):659–671, August 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilkes:1982:HCE

[Wil82a] John Wilkes. HADES: a command environment that supports structure. Software—Practice and Experience, 12(7):641–654, July 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Williams:1982:SAS

[Wil82b] M. Howard Williams. A system to assist survey analysts. Software—Practice and Experience, 12(5):487– 491, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Williamson:1983:IBP

[Wil83] J. H. Williamson. Incremental back-up procedures. Software—Practice and Experience, 13 (11):1027–1031, November 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilson:1984:PPT

[Wil84a] Anne D. Wilson. Programs to process trees, representing program structures and data structures. Software —Practice and Experience, 14(9):807–816, September 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilson:1984:BRB

[Wil84b] L. B. Wilson. Book review: Algorithms, Robert Sedgewick, Addison-Wesley Publishing Company, 1983. No. of pages: 552. Price: £13.95. Software—Practice and Experience, 14(5):501– 502, May 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilson:1987:BRB

[Wil87] J. D. Wilson. Book review: Fortran Optimization (revised edition), Michael Metcalf, Academic Press, 1985. No. of pages: 253. Price: £20.00. Software —Practice and Experience, 17(2):166–167, February 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilson:1989:GPO

[Wil89] G. A. Wilson. Generic parameters in Occam. Software—Practice and Experience, 19(11):1057–1064, November 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Winkler:2002:SVU

[Win02] Jürgen F. H. Winkler. A safe variant of the unsafe integer arithmetic of Java^{T M}. Software—Practice and Experience, 32(7): 669–701, June 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        94515736/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=94515736&PLACEBO=IE.
                                        pdf.

Wirth:1971:DPC

[Wir71] Niklaus Wirth. The design of a PASCAL compiler. Software—Practice and Experience, 1(4):309– 333, October/December 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1972:GES

[Wir72] Niklaus Wirth. Guest editorial: Software and its portability. Software— Practice and Experience, 2 (4):311–312, October 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1977:GEG

[Wir77a] N. Wirth. Guest editorial:

Guest editorial extract from Professor Cleverbyte’s visit to heaven. Software —Practice and Experience, 7(2):155–158, March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1977:DIM

[Wir77b] Niklaus Wirth. Design and implementation of Modula. Software—Practice and Experience, 7(1):67– 84, January/March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1977:MLM

[Wir77c] Niklaus Wirth. Modula: a language for modular multiprogramming. Software—Practice and Experience, 7(1):3–35, January/ March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1977:UM

[Wir77d] Niklaus Wirth. The use of Modula. Software— Practice and Experience, 7 (1):37–65, January/March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1988:MO

[Wir88a] Niklaus Wirth. From Modula to Oberon. Software —Practice and Experience, 18(7):661–670, July 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1988:PLO

[Wir88b]  Niklaus Wirth.  The

programming language Oberon. Software—Practice and Experience, 18(7): 671–690, July 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1990:CNL

[Wir90] Niklaus Wirth. Ceres-Net: a low-cost computer network. Software—Practice and Experience, 20(1):13– 24, January 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wiseman:1974:BRB

[Wis74] N. E. Wiseman. Book review: Computer composition using Page-1, John Pierson, Wiley Interscience, New York, 1972. No. of pages: 135. Price: £3.50. Software—Practice and Experience, 4(3):302, July/September 1974. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wise:1993:EPP

[Wis93] Michael J. Wise. Experience with PMS-Prolog: a distributed coarse-grain-parallel Prolog with processes, modules and streams. Software—Practice and Experience, 23(2):151–175, February 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wijayarathna:1997:ABC

[WIS^{+}97] P. G. Wijayarathna, C. Ito, K. Sakamoto, S. Hamba, and T. Kushiya. The analysis of the bus chaining phenomenon in traffic congestion. Software—Practice and Experience, 27(9):1067–1081, September 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=7334;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=7334&
                                        PLACEBO=IE.pdf.

Witt:1977:PDB

[Wit77a] J. Witt. ‘Pun-Dora’s’ box or how to produce enough papers to wrap up structured programming. Software—Practice and Experience, 7(2):296, March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Witty:1977:DF

[Wit77b] Robert W. Witty. Dimensional flowcharting. Software—Practice and Experience, 7(5):553–584, September/October 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Witsel:1982:MLS

[Wit82] Willem Maarten Witsel. Meta: a language system for bitslice applications. Software—Practice and Experience, 12(4):297– 308, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Witschorik:1983:RTD

[Wit83] C. A. Witschorik. The realtime debugging monitor for the Bell System 1A processor. Software—Practice and Experience, 13(8):727– 743, August 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2020:VRO

[WIYC20] Yifan Wang, Kiran Ijaz, Dong Yuan, and Rafael A. Calvo. VR-Rides: an object-oriented application framework for immersive virtual reality exergames. Software—Practice and Experience, 50(7): 1305–1324, July 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wichmann:1976:TAC

[WJ76] B. A. Wichmann and B. Jones. Testing ALGOL 60 compilers. Software— Practice and Experience, 6(2):261–270, April/June 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wright:1993:SRI

[WJ93] William E. Wright and Sakthivel Jeyaratnam. Success rate of interpolation in subsegment prediction. Software—Practice and Experience, 23(3):243– 253, March 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2014:ECP

[WJC^{+}14] Yi Wang, Zhiping Jia, Renhai Chen, Meng Wang, Duo Liu, and Zili Shao. Extended conference papers: Loop scheduling with memory access reduction subject to register constraints for DSP applications. Software—Practice and Experience, 44(8):999– 1026, August 2014. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Westermann:2006:PSA

[WK06a] U. Westermann and W. Klas. PTDOM: a schema-aware XML database system for MPEG-7 media descriptions. Software—Practice and Experience, 36(8):785– 834, July 10, 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wong:2006:E

[WK06b] W. Eric Wong and Karama Kanoun. Editorial. Software—Practice and Experience, 36(9):891, July 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wong:1996:SSB

[WKD96] Johnny Wong, Sriram Kini, and Kishore Doobagunta. Synchronization in specification-based multimedia presentations. Software—Practice and Experience, 26(1):71–81, January 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61004281.

Wibisono:2013:OOO

[WKG^{+}13] Adianto Wibisono, Ralph Koning, Paola Grosso, Adam Belloum, Marian Bubak, and Cees de Laat. OIntEd: online ontology instance editor enabling a new approach to ontology development. Software—Practice and Experience, 43(11):1319–1335, November 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2015:IJV

[WKJ15] Yang Wang, Kenneth B.

Kent, and Graeme Johnson. Improving J9 virtual machine with LTTng for efficient and effective tracing. Software—Practice and Experience, 45(7):973– 987, July 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kwon:2018:LMW

[wKJM18] Jin woo Kwon, InChang Jeong, and Soo-Mook Moon. Lightweight migration for web applications with framework separation. Software—Practice and Experience, 48 (3):621–640, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/abs/10.
                                        1002/spe.2542.

Wurz:2024:MMA

[WKKK24] Hendrik M. Würz, Michel Krämer, Marvin Kaster, and Arjan Kuijper. Migrating monolithic applications to function as a service. Software—Practice and Experience, 54(2):149– 167, February 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wortman:1976:SPC

[WKL76] David B. Wortman, Philip J. Khaiat, and David M. Laskar. Six PL/I compilers. Software—Practice and Experience, 6(3):411– 422, July/September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wijayarathna:1998:GRS

[WKS^{+}98] P. G. Wijayarathna, Y. Kawata, A. Santosa, K. Isogai, and M. Maekawa. GSL: a requirements specification language for enduser intelligibility. Software—Practice and Experience, 28(13):1387–1414, November 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1769;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1769&
PLACEBO=IE.pdf.

Willis:1972:NPU

[WL72] Chris Willis and Lawrence Liddiard. A note on packing and unpacking of bytes. Software—Practice and Experience, 2(4):401– 402, October 1972. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Welsh:1981:CST

[WL81a] Jim Welsh and Andrew Lister. A comparative study of task communication in Ada. Software—Practice and Experience, 11(3):257– 290, March 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wood:1981:DVT

[WL81b] G. K. Wood and J. Larmouth. Distributing viewdata and teletext services to a user community. Software—Practice and Experience, 11(10):1009–1017, October 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wong:2003:EI

[WL03] W. Eric Wong and J. Jenny Li. Editorial: ICCCN 2001. Software—Practice and Experience, 33(14): 1299–1300, November 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wong:1998:CCT

[WLL98] Kam-Fai Wong, Vincent Y. Lum, and Wai-Ip Lam. Chicon — a Chinese text manipulation language. Software—Practice and Experience, 28(7):681– 701, July 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=1804;   http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/fulltext?ID=1804&
                                        PLACEBO=IE.pdf.

Wu:2021:GBC

[WLS^{+}21] Jimmy Ming-Tai Wu, Zhongcui Li, Gautam Srivastava, Meng-Hsiun Tasi, and Jerry Chun-Wei Lin. A graph-based convolutional neural network stock price prediction with leading indicators. Software—Practice and Experience, 51(3): 628–644, March 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Walraven:2013:PDC

[WLTJ13] Stefan Walraven, Bert Lagaisse, Eddy Truyen, and Wouter Joosen. Policydriven customization of cross-organizational features in distributed service systems. Software— Practice and Experience, 43(10):1145–1163, October 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2022:MNS

[WLZ^{+}22] Yunzhuo Wang, Jianfeng Li, Liangying Zhou, Jingwei Sun, and Guangzhong Sun. Multi-net strategy: Accelerating physicsinformed neural networks for solving partial differential equations. Software—Practice and Experience, 52(12):2513–2536, December 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wainer:2020:DRT

[WM20] Gabriel Wainer and Mohammad Moallemi. Designing real-time systems using imprecise discreteevent system specifications. Software—Practice and Experience, 50(8): 1327–1344, August 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wong:1994:RDA

[WMG94] Johnny Wong, William Marshall, and Robert Goodman. Remote database access in the distributed computing environment. Software—Practice and Experience, 24(4):421–434, April 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wu:2021:FLB

[WMHL21] Kaidong Wu, Yun Ma, Gang Huang, and Xuanzhe Liu. A first look at blockchain-based decentralized applications. Software—Practice and Experience, 51(10):2033–2050, October 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Watson:2004:EPA

[WMJ04] David Watson, G. Robert Malan, and Farnam Jahanian. An extensible probe architecture for network protocol performance measurement. Software —Practice and Experience, 34(1):47–67, January 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2012:PPP

[WMSY12] Peishun Wang, Yi Mu, Willy Susilo, and Jun Yan. Privacy preserving protocol for service aggregation in cloud computing. Software —Practice and Experience, 42(4):467–483, April 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Winder:1988:JAI

[WN88] Russel Winder and Joe Nicolson. JDB: an adaptable interface for debugging. Software—Practice and Experience, 18(3):221– 238, March 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wong:2006:SPP

[WN06] William Hak-Lim Wong and Joseph Kee-Yin Ng. Scalable peer-to-peer networking architecture: DI-VINE. Software—Practice and Experience, 36(13): 1467–1487, November 10, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Watanabe:1981:MMI

[WOKT81] Tan Watanabe, Tsuneharu Ohsawa, Hisao Kuma, and Wakunaga Tsukada. Micro MUMPS: An interactive database language for micro-computers. Software—Practice and Experience, 11(12):1307–1314, December 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wolberg:1982:CMS

[Wol82] John R. Wolberg. A costing model for software conversions. Software—Practice and Experience, 12 (11):1043–1049, November 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wolfsthal:1991:SCQ

[Wol91] Yaron Wolfsthal. Style control in the Quill Document Editing System. Software —Practice and Experience, 21(6):625–638, June 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wolf:1992:OOI

[Wol92]  Wayne Wolf.  Object-

oriented implementation issues in an experimental CAD system. Software —Practice and Experience, 22(4):287–304, April 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Woodsford:1971:DIG

[Woo71] P. A. Woodsford. The design and implementation of the GINO 3D graphics software package. Software—Practice and Experience, 1(4):335–365, October/December 1971. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Woodward:1972:PEA

[Woo72] P. M. Woodward. Practical experience with ALGOL 68. Software—Practice and Experience, 2(1):7–19, January/March 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Woodward:1974:BRB

[Woo74] P. M. Woodward. Book review: ALGOL 68 implementation, Ed. J. E. L. Peck, proceedings of the IPIP working conference on ALGOL 68 implementation, Munich, 20–24 July 1970, North-Holland, Amsterdam, 1971. No. of pages:

375. Price: £7.35. Software —Practice and Experience, 4(2):192, April/June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Woodward:1984:AHS

[Woo84] M. R. Woodward. The application of Halstead’s software science theory to Algol 68 programs. Software—Practice and Experience, 14(3):263–276, March 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Woodman:1986:FSM

[Woo86] Mark Woodman. Formatted syntaxes and Modula-2. Software—Practice and Experience, 16(7):605– 626, July 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Workman:1983:GSD

[Wor83] David A. Workman. GRASP: a software development system using D-Charts. Software—Practice and Experience, 13(1):17–32, January 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wong:1996:RAM

[WP96] Johnny S. Wong and Deepa S. Parthasarathy. Remote access to multimedia databases: An object oriented approach. Software—Practice and Experience, 26(6):677–704, June 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=16816.

Walton:2000:GTP

[WP00] Gwendolyn H. Walton and J. H. Poore. Generating transition probabilities to support model-based software testing. Software—Practice and Experience, 30(10):1095–1106, August 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72510349/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72510349&PLACEBO=IE.
pdf.

Whaley:2005:MDM

[WP05] R. Clint Whaley and Antoine Petitet. Minimizing development and maintenance costs in supporting persistently optimized BLAS. Software—Practice and Experience, 35(2):101– 121, February 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2021:MGP

[WPL^{+}21] Tingting Wang, Beibei Pang, Xinxi Li, Yan Wang, Chaoqiang Huang, Jian Gong, Guangai Sun, and Dong Liu. MyNeutron-DAQ: a general program for data acquisition of neutron scattering spectrometers at China Mianyang Research Reactor and online data-analysis. Software —Practice and Experience, 51(2):438–448, February 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Williams:1986:RSD

[WPN86] M. Howard Williams, I. M. Pattison, and J. C. Neves. Reorganization in a simple database system. Software—Practice and Experience, 16(8):719–729, August 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Walton:1995:STS

[WPT95] Gwendolyn H. Walton, J. H. Poore, and Carmen J. Trammell. Statistical testing of software based on a usage model. Software—Practice and Experience, 25(1):97– 108, January 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Welsh:1972:PCI

[WQ72] J. Welsh and C. Quinn. A PASCAL compiler for ICL 1900 series computers. Software—Practice and Experience, 2(1):73– 77, January/March 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wiseman:1977:OSI

[WR77] N. E. Wiseman and P. Robinson. An operating system for interactive terminals. Software—Practice and Experience, 7(4):501– 510, July/August 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wolberg:1978:CLP

[WR78] John R. Wolberg and Marshall Rafal. CONVERT: a language for program and data file conversions. Software—Practice and Experience, 8(2):187–198, March/ April 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wolberg:1979:UCT

[WR79] John R. Wolberg and Marshall Rafal. Using convert to transform source code. Software—Practice and Experience, 9(11):881– 890, November 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Winner:1984:OSS

[WR84] Robert I. Winner and L. B. Reed. Operating system support for sharing writable control store. Software—Practice and Experience, 14(12):1183–1196, December 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wallace:1995:EFP

[WR95] Malcolm Wallace and Colin Runciman. Extending a functional programming system for embedded applications. Software—Practice and Experience, 25(1): 73–96, January 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilson:2022:RTF

[WR22] Allan Jenose Wilson and Aravindhan Sarojini Radhamani. Real time flood disaster monitoring based on energy efficient ensemble clustering mechanism in wireless sensor network. Software—Practice and Experience, 52(1):254– 276, January 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

White:1999:SVL

[WRD99] Elizabeth L. White, Jeffrey Ruby, and Laura Denise Deddens. Software visualization of LR parsing and synthesized attribute evaluation. Software—Practice and Experience, 29(1): 1–16, January 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55002652;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55002652&PLACEBO=IE.
pdf.

Wrench:1988:CIC

[Wre88] K. L. Wrench. CSP-i:

An implementation of communicating sequential processes. Software—Practice and Experience, 18(6):545– 560, June 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wright:1994:ASM

[Wri94] Alden H. Wright. Approximate string matching using within-word parallelism. Software—Practice and Experience, 24(4):337– 362, April 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wright:1998:BRB

[Wri98] Peter Wright. Book review:

Human-computer interaction, by A. Dix, J. Finlay, G. Abowd, R. Beale, Prentice Hall, 1998. Software—Practice and Experience, 28(12):1357–1358, October 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=1765;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=1765&
PLACEBO=IE.pdf.

Wong:1997:MPT

[WRR97] Johnny Wong, Satish Rao, and Naveen Ramaiah. A multimedia presentation toolkit for the World Wide Web. Software—Practice and Experience, 27(4):425– 446, April 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7300;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7300&
PLACEBO=IE.pdf.

Wong:1974:USM

[WS74] K. Wong and Jon C. Strauss. Use of a software monitor in the validation of an analytic computer system model. Software—Practice and Experience, 4(3):255–263, July/ September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilk:1983:OPP

[WS83] A. Wilk and W. Silverman. OPTIMA: a portable PCODE optimizer. Software—Practice and Experience, 13(4):323–354, April 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Whitfield:1994:DIG

[WS94a] Deborah Whitfield and Mary Lou Soffa. The design and implementation of Genesis. Software—Practice and Experience, 24(3): 307–325, March 1994. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Woodside:1994:CPM

[WS94b] C. Murray Woodside and Cheryl Schramm. Complex performance measurements with NICE (Notation for Interval Combinations and Events). Software—Practice and Experience, 24(12):1121–1144, December 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Whyman:1999:EMT

[WS99] Edward K. Whyman and Harold L. Somers. Evaluation metrics for a translation memory system. Software—Practice and Experience, 29(14):1265–1284, December 10, 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
67500434/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=67500434&PLACEBO=IE.
pdf.

Weber:1996:VDP

[WSB96] D. Weber, M. Spezialetti, and H. Barada. Vid-Net: Distributed processing environment for computer generated animation. Software—Practice and Experience, 26(2):237– 250, February 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=61004290.

Welsh:1977:AIP

[WSH77] J. Welsh, W. J. Sneeringer, and C. A. R. Hoare. Ambiguities and insecurities in Pascal. Software— Practice and Experience, 7(6):685–696, November/ December 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wu:2003:OHP

[WSL03] Min-You Wu, Wei Shu, and Chow-Sing Lin. Odyssey: a high-performance clustered video server. Software —Practice and Experience, 33(7):673–700, June 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2020:PPC

[WSL^{+}20] Ruoyu Wang, Daniel Sun, Guoqiang Li, Raymond Wong, and Shiping Chen. Pipeline provenance for cloud-based big data analytics. Software—Practice and Experience, 50(5):658– 674, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wiesner:2023:SLS

[WSN^{+}23] Philipp Wiesner, Marvin Steinke, Henrik Nickel, Yazan Kitana, and Odej Kao. Software-in-the-loop simulation for developing and testing carbon-aware applications. Special focus section: Benchmarking, Experimentation Tools, and Reproducible Practices for Data-Intensive Systems from Edge to Cloud: Software: Practice and Experience, 53(12):2362–2376, December 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2023:HRD

[WSWR23] Ruoyu Wang, Daniel Sun, Raymond Wong, and Rajiv Ranjan. Horn rule discovery with batched caching and rule identifier for proficient compressor of knowledge data. Software—Practice and Experience, 53(3): 682–703, March 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wada:2011:EDO

[WSYO11] Hiroshi Wada, Junichi Suzuki, Yuji Yamano, and Katsuya Oba. Evolutionary deployment optimization for service-oriented clouds. Software—Practice and Experience, 41(5):469– 493, May 25, 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wu:2022:CLA

[WTF^{+}22] Song Wu, Zhiheng Tao, Hao Fan, Zhuo Huang, Xinmin Zhang, Hai Jin, Chen Yu, and Chun Cao. Container lifecycle-aware scheduling for serverless computing. Software— Practice and Experience, 52(2):337–352, February 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wu:1999:BWN

[Wu99] Pei-Chi Wu. A byte-wise normalization method in arithmetic coding. Software—Practice and Experience, 29(4):299–309, April 10, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=55001839;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=55001839&PLACEBO=IE.
pdf.

Wu:2000:TNS

[Wu00] Pei-Chi Wu. Transition from national standards to Unicode: multilingual support in operating systems and programming languages. Software—Practice and Experience, 30(7): 765–774, June 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
72001831/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=72001831&PLACEBO=IE.
pdf.

Wu:2001:BTF

[Wu01] Pei-Chi Wu. A base62 transformation format of ISO 10646 for multilingual identifiers. Software—Practice and Experience, 31(12):1125–1130, October 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        85008992/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=85008992&PLACEBO=IE.
                                        pdf.

Wu:2002:PST

[Wu02] Pei-Chi Wu. A pageshift transformation format of ISO 10646. Software—Practice and Experience, 32(1):73–82, January 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        88013124/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=88013124&PLACEBO=IE.
                                        pdf.

Wulf:1975:LE

[Wul75] William A. Wulf. Letters to the editor. Software— Practice and Experience, 5 (1):109–110, January 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Witten:1983:GUS

[WW83] Ian H. Witten and Brian Wyvill. On the generation and use of space-filling curves. Software—Practice and Experience, 13(6):519– 525, June 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilson:1989:CTT

[WW89] G. A. Wilson and G. A.

Wilson. Comparative timings of three different set implementations in occam. Software—Practice and Experience, 19(3):273– 281, March 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

White:1991:PTC

[WW91] Lee J. White and Bogdan Wiszniewski. Path testing of computer programs with loops using a tool for simple loop patterns. Software—Practice and Experience, 21(10):1075–1102, October 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:1995:IHA

[WW95] H. K. Wang and Jean-Lien C. Wu. Interactive hypermedia applications: a model and its implementation. Software—Practice and Experience, 25 (9):1045–1063, September 1995. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Wu:1996:EOC

[WW96] Pei-Chi Wu and Feng-Jian Wang. On efficiency and optimization of C++ programs. Software—Practice and Experience, 26(4): 453–465, April 1996. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16803.

Wagner:2000:EDA

[WW00] Dorothea Wagner and Karsten Weihe. Editorial: Discrete algorithm engineering. Software —Practice and Experience, 30(11):1165, September 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        72517073/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=72517073&PLACEBO=IE.
                                        pdf.

Wang:2009:AHC

[WW09] Yi-Hsien Wang and I-Chen Wu. Achieving high and consistent rendering performance of Java AWT/ Swing on multiple platforms. Software—Practice and Experience, 39(7):701– 736, May ??, 2009. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2003:DIA

[WWB03] Rong Wang, Feiyi Wang, and Gregory T. Byrd. Design and implementation of Acceptance Monitor for building intrusion tolerant systems. Software— Practice and Experience, 33(14):1399–1417, November 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2019:FEO

[WWCW19] Pei Wang, Dinghao Wu, Zhaofeng Chen, and Tao Wei. Field experience with obfuscating millionuser iOS apps in large enterprise mobile development. Software—Practice and Experience, 49(2):252– 273, February 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wolforth:2010:GSA

[WWGP10] Ian Wolforth, Martin Walker, Lars Grunske, and Yiannis Papadopoulos. Generalizable safety annotations for specification of failure patterns. Software —Practice and Experience, 40(5):453–483, April 25, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2007:PAS

[WWJ07] Yi-Hsien Wang, I-Chen Wu, and Jyh-Yaw Jiang. A portable AWT/Swing architecture for Java game development. Software— Practice and Experience, 37(7):727–745, June 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wei:2016:ESD

[WXR16] Shiyi Wei, Franceska Xhakaj, and Barbara G. Ryder. Empirical study of the dynamic behavior of JavaScript objects. Software—Practice and Experience, 46(7):867–889, July 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2018:CIV

[WY18a] Rui Wang and Shi Ying. Cover image volume 48, issue 11. Software—Practice and Experience, 48(11):i, November 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2018:SSP

[WY18b] Rui Wang and Shi Ying. SaaS software performance issue identification using HMRF–MAP framework. Software—Practice and Experience, 48 (11):2000–2018, November 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wyatt:1984:SPI

[Wya84] M. G. Wyatt. A shorthand preprocessor for IDMS COBOL. Software—Practice and Experience, 14(2): 191–194 (or 191–193??), February 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wu:2015:MMF

[WYAZ15] Ji Wu, Tao Yue, Shaukat Ali, and Huihui Zhang. A modeling methodology to facilitate safetyoriented architecture design of industrial avionics software. Software—Practice and Experience, 45(7): 893–924, July 2015. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wyvill:1977:PM

[Wyv77] B. L. M. Wyvill. PICTURES-68 MK1. Software—Practice and Experience, 7 (2):251–261, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wortman:1994:ADC

[WZF94] David B. Wortman, S. Zhou, and S. Fink. Automating data conversion for heterogeneous distributed shared memory. Software—Practice and Experience, 24 (1):111–125, January 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Williams:2001:SAT

[WZH01] Hugh E. Williams, Justin Zobel, and Steffen Heinz. Self-adjusting trees in practice for large text collections. Software—Practice and Experience, 31(10): 925–939, August 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
82002140/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=82002140&PLACEBO=IE.
pdf.

Ward:2008:CSS

[WZLN08] Martin Ward, Hussein Zedan, Matthias Ladkau, and Stefan Natelberg. Conditioned semantic slicing for abstraction; industrial experiment. Software— Practice and Experience, 38(12):1273–1304, October ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xiao:2007:HIB

[XAN07] Gaoyu Xiao, Aamer Aziz, and Wieslaw L. Nowinski. Hybrid image-based collision detection in Java 3D. Software—Practice and Experience, 37(9):963– 982, July 25, 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xue:2006:PDC

[XCG06] Jingling Xue, Qiong Cai, and Lin Gao. Partial dead code elimination on predicated code regions. Software—Practice and Experience, 36(15):1655–1685, December 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xie:2018:JSD

[XCL^{+}18] Guoqi Xie, Yuekun Chen, Yan Liu, Chunnian Fan, Renfa Li, and Keqin Li. JDAS: a software development framework for multidatabases. Software—Practice and Experience, 48(2):366–382, February 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2472.

Xu:2021:BBC

[XCZ^{+}21] Xiaolong Xu, Yi Chen, Xuyun Zhang, Qingxiang Liu, Xihua Liu, and Lianyong Qi. A blockchainbased computation offloading method for edge computing in 5G networks. Software—Practice and Experience, 51(10): 2015–2032, October 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xu:2024:SIE

[XDVB24] Minxian Xu, Schahram Dustdar, Massimo Villari, and Rajkumar Buyya. Special issue on efficient management of microservicebased systems and applications. Efficient Management of Microservicebased Systems and Applications: Software: Practice and Experience, 54(4):543– 545, April 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xu:2017:THD

[XDZ^{+}17] Xiaolong Xu, Wanchun Dou, Xuyun Zhang, Chunhua Hu, and Jinjun Chen. A traffic hotline discovery method over Cloud of Things using big taxi GPS data. Software—Practice and Experience, 47(3):361– 377, March 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xiang:2021:JDB

[XHP^{+}21] Fu Xiang, Wang Huaimin, Shi Peichang, Ouyang Xue, and Zhang Xunhui. Jointgraph: a DAGbased efficient consensus algorithm for consortium blockchains. Software— Practice and Experience, 51(10):1987–1999, October 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xue:2019:PPT

[XLLY19] Gang Xue, Di Liu, Junsong Liu, and Shaowen Yao. A process partitioning technique for constructing decentralized web service compositions. Software—Practice and Experience, 49(10):1550–1570, October 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xu:2020:AVE

[XLZ^{+}20] Xianghua Xu, LiQiming Liu, Lingjun Zhang, Ping Li, and Jinjun Chen. Abnormal visual event detection based on multi-instance learning and autoregressive integrated moving average model in edge-based Smart City surveillance. Software— Practice and Experience, 50(5):476–488, May 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xiao:2021:EAA

[XMTL21] Ling Xiao, Yu Meng, Xiaobing Tian, and Haibo Luo. Energy allocation for activity recognition in wearable devices with kinetic energy harvesting. Software—Practice and Experience, 51 (11):2185–2202, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xi:2023:LSE

[XP23] Rui Xi and Karthik Pattabiraman. A large-scale empirical study of low-level function use in Ethereum smart contracts and automated replacement. Software—Practice and Experience, 53(3):631–664, March 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xiao:2017:SOO

[XWC^{+}17] Jing Xiao, Zhongyuan Wang, Yu Chen, Liang Liao, Jun Xiao, Gen Zhan, and Ruimin Hu. A sensitive object-oriented approach to big surveillance data compression for social security applications in smart cities. Software —Practice and Experience, 47(8):1061–1080, August 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xie:2021:AVC

[XWHX21] Pingfu Xie, Yunlong Wang, Bo He, and Guoqi Xie. AutoToolVCC: Variable consistency check tool for automotive software projects. Software—Practice and Experience, 51(11):2306– 2316, November 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xie:2018:IBI

[XXJS18] Wuping Xie, Jinyun Xue, Dongming Jiang, and Lan Song. An iteration-based interactive analysis method to design dynamic serviceoriented systems. Software—Practice and Experience, 48(2):300–326, February 2018. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2543.

Xie:2013:AAE

[XXZ13] Xinwei Xie, Jingling Xue, and Jie Zhang. Acculock: accurate and efficient detection of data races. Software—Practice and Experience, 43(5):543– 576, May 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xu:2024:PAC

[XYW^{+}24] Minxian Xu, Lei Yang, Yang Wang, Chengxi Gao, Linfeng Wen, Guoyao Xu, Liping Zhang, Kejiang Ye, and Chengzhong Xu. Practice of Alibaba cloud on elastic resource provisioning for large-scale microservices cluster. Software —Practice and Experience, 54(1):39–57, January 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xu:2001:CCM

[XZ01] Baowen Xu and Yuming Zhou. Comments on “A cohesion measure for objectoriented classes” by Heung Seok Chae, Yong Rae Kwon and Doo Hwan Bae (Softw. Pract. Exper. 2000; 30:1405–1431). Software —Practice and Experience, 31(14):1381–1388, November 25, 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
85515670/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=85515670&PLACEBO=IE.

pdf. See [CKB00, CKB01, XZ03, CKB03].

Xu:2003:MCC

[XZ03] Baowen Xu and Yuming Zhou. More comments on “A cohesion measure for object-oriented classes” by Heung-Seok Chae, Yong-Rae Kwon and Doo-Hwan Bae (Softw. Pract. Exper. 2000; 30:1405–1431). Software—Practice and Experience, 33(6):583–588, May 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [CKB00, CKB01, XZ01, CKB03].

Xu:2022:PPS

[XZW^{+}22] Minxian Xu, Qiheng Zhou, Huaming Wu, Weiwei Lin, Kejiang Ye, and Chengzhong Xu. PDMA: Probabilistic service migration approach for delayaware and mobility-aware mobile edge computing. Software—Practice and Experience, 52(2):394–414, February 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xie:2021:CPV

[XZX^{+}21] Yong Xie, Yu Zhou, Jing Xu, Jian Zhou, Xiaobai Chen, and Fu Xiao. Cybersecurity protection on in-vehicle networks for distributed automotive cyberphysical systems: Stateof-the-art and future challenges. Software—Practice and Experience, 51 (11):2108–2127, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yaseen:2019:CBV

[YAFA19] Muhammad Usman Yaseen, Ashiq Anjum, Mohsen Farid, and Nick Antonopoulos. Cloud-based video analytics using convolutional neural networks. Software —Practice and Experience, 49(4):565–583, April 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yang:1991:ISD

[Yan91] Wuu Yang. Identifying syntactic differences between two programs. Software —Practice and Experience, 21(7):739–755, July 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yasrebi:1994:IPT

[Yas94]  Mehrad Yasrebi.  Im-

plementations and performances of two RPCbased LAN gateway applications. Software—Practice and Experience, 24(3): 235–264, March 1994. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yousefi-Azar:2021:MIF

[YAVHC21] Mahmood Yousefi-Azar, Vijay Varadharajan, Len Hamey, and Shiping Chen. Mutual information and feature importance gradient boosting: Automatic byte n-gram feature reranking for Android malware detection. Software—Practice and Experience, 51(7):1518–1539, July 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yeo:2006:TMB

[YB06] Chee Shin Yeo and Rajkumar Buyya. A taxonomy of market-based resource management systems for utility-driven cluster computing. Software —Practice and Experience, 36(13):1381–1419, November 10, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

You:2016:SRB

[YC16] Yi-Ping You and Jian-Ru Chen. A static region-based compiler for the Dalvik virtual machine. Software—Practice and Experience, 46(8): 1109–1130, August 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yang:2003:ICS

[YCY03] S. H. Yang, X. Chen, and L. Yang. Integration of control system design and implementation over the Internet using the Jini technology. Software—Practice and Experience, 33(12):1151–1175, October 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yang:1991:USC

[YF91] Cui-Qing Yang and Raphael A. Finkel. Utility servers in Charlotte. Software— Practice and Experience, 21(5):429–441, May 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yerramalla:2006:VAN

[YFC06] Sampath Yerramalla, Edgar Fuller, and Bojan Cukic. A validation approach for neural network-based online adaptive systems. Software—Practice and Experience, 36(11–12):1209–1225, September/October 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yang:1997:OOC

[YH97] Feng-Chao Yang and Yu-Kuen Ho. An objectoriented cooperative distributed problem solving shell with groupware management ability. Software—Practice and Experience, 27(11):1307–1334, November 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=7272;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=7272&
PLACEBO=IE.pdf.

Yu:2020:RSI

[YHGC20] Xiao Yu, Congcong Han, Junqi Guo, and Yinghe Chen. The role of superior image composition in children’s analogical reasoning. Software—Practice and Experience, 50 (11):2082–2094, November 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yau:2006:SSA

[YHGY06] Stephen S. Yau, Dazhi Huang, Haishan Gong, and Yisheng Yao. Support for situation awareness in trustworthy ubiquitous computing application software. Software—Practice and Experience, 36(9): 893–921, July 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yin:2022:SEM

[YHK^{+}22] Xiaohua Yin, Zhiqiu Huang, Shuanglong Kan, Guohua Shen, Zhe Chen, Yang Liu, and Fei Wang. SafeOSL: Ensuring memory safety of C via ownership-based intermediate language. Software—Practice and Experience, 52(5):1114–1142, May 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yi:2012:PSL

[Yi12] Qing Yi. POET: a scripting language for applying parameterized sourceto-source program transformations. Software— Practice and Experience, 42(6):675–706, June 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yip:1982:ICP

[Yip82] C. K. Yip. Implementation of a CORAL66 programming environment. Software—Practice and Experience, 12(5):411–431, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yip:1984:PGS

[Yip84] C. K. Yip. The Pascal graphics system. Software —Practice and Experience, 14(2):101–118, February 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yehudai:1995:TRP

[YL95] Amiram Yehudai and Fernando Libedinsky. On tuning recursive procedures. Software—Practice and Experience, 25(9):997– 1002, September 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yu:2005:MXD

[YLM^{+}05] Yijun Yu, Jianguo Lu, John Mylopoulos, Weiwei Sun, Jing-Hao Xue, and Erik H. D’Hollander. Making XML document markup international. Software—Practice and Experience, 35(1):1–14, January 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Youn:2011:FGB

[YLP^{+}11] Jonghee M. Youn, Jongwon Lee, Yunheung Paek, Jongeun Lee, Hanno Scharwaechter, and Rainer Leupers. Fast graphbased instruction selection for multi-output instructions. Software—Practice and Experience, 41(6):717– 736, May 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Youn:2023:DSA

[YLR23] Dongjun Youn, Sungho Lee, and Sukyoung Ryu. Declarative static analysis for multilingual programs using CodeQL. Software—Practice and Experience, 53(7):1472–1495, July 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yang:2005:LMJ

[YME05] Byung-Sun Yang, Soo-Mook Moon, and Kemal Ebcioğlu. Lightweight monitors for the Java virtual machine. Software —Practice and Experience, 35(3):281–299, March 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yin:2016:PAI

[YMH16] Jie Yin, Chao Ma, and Shi-Min Hu. PAST: accurate instrumentation on fully optimized program. Software—Practice and Experience, 46(3):341–360, March 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Youssef:2017:WGE

[YMY17] Ahmed Youssef, Muthucumaru Maheswaran, and Lamis Youssef. Wireless GINI: an educational platform for hosting virtual wireless networks. Software —Practice and Experience, 47(1):21–59, January 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

You:2015:EIT

[YOH15] Ilsun You, Marek R.

Ogiela, and Myunggwon Hwang. Editorial: Intelligent technologies and applications for big data analytics. Software—Practice and Experience, 45(8): 1019–1021, August 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yata:2007:EDM

[YOM^{+}07] Susumu Yata, Masaki Oono, Kazuhiro Morita, Masao Fuketa, and Jun ichi Aoe. An efficient deletion method for a minimal prefix double array. Software —Practice and Experience, 37(5):523–534, April 25, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yoo:1996:CAA

[Yoo96] H. Chuck Yoo. Comparative analysis of asynchronous I/O in multithreaded UNIX. Software—Practice and Experience, 26(9):987–997, September 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16832.

Young:1981:ISL

[You81] S. J. Young. Improving the structure of large Pascal programs. Software—Practice and Experience, 11(9): 913–927, September 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yu:2022:PSD

[YQL22] Haisheng Yu, Heng Qi, and Keqiu Li. A powerful software-defined cyberphysical system to expand CPS adoption. Software —Practice and Experience, 52(4):904–916, April 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yaung:1992:LAP

[YR92] Alan T. Yaung and Tzvi Raz. Linkage analysis of processes. Software— Practice and Experience, 22(10):849–862, October 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yousefipour:2018:ECA

[YRJ18] Amin Yousefipour, Amir Masoud Rahmani, and Mohsen Jahanshahi. Energy and cost-aware virtual machine consolidation in cloud computing. Software—Practice and Experience, 48(10): 1758–1774, October 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yussupov:2022:SBM

[YSBL22] Vladimir Yussupov, Jacopo Soldani, Uwe Breitenbücher, and Frank Leymann. Standardsbased modeling and deployment of serverless function orchestrations using BPMN and TOSCA. Software—Practice and Experience, 52(6):1454–1495, June 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yan:1995:PMV

[YSM95] Jerry Yan, Sekhar Sarukkai, and Pankaj Mehra. Performance measurement, visualization and modeling of parallel and distributed programs using the AIMS toolkit. Software—Practice and Experience, 25(4):429– 461, April 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yu:1996:DCP

[Yu96]  Tong Lai Yu.  Data

compression for PC software distribution. Software—Practice and Experience, 26(11):1181–1195, November 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?
                                        ID=16769.

Yuval:1975:GRT

[Yuv75]  G. Yuval.  Gathering

run-time statistics without black magic. Software—Practice and Experience, 5(1):105–108, January/March 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yuval:1977:CH

[Yuv77a]  G. Yuval.  CRAY-1

FOR 6000/7000 HACK-ERS. Software—Practice and Experience, 7(3):427– 428, June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yuval:1977:YRR

[Yuv77b] G. Yuval. Is your register really necessary? Software—Practice and Experience, 7(2):295, March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yuval:1977:UCR

[Yuv77c] G. Yuval. The utility of the CDC 6000 registers. Software—Practice and Experience, 7(4):535– 536, July 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yuval:1978:YAS

[Yuv78] G. Yuval. Yet another software profile. Software —Practice and Experience, 8(4):495–496, July 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yuval:1979:FTG

[Yuv79a] G. Yuval. Faster than goto. Software—Practice and Experience, 9(9):791, September 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yuval:1979:IMB

[Yuv79b] G. Yuval. An interpreterminded benchmark. Software—Practice and Experience, 9(6):509–510, June 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yang:2000:DDI

[YWN^{+}00] Jiong Yang, Wei Wang, Silvia Nittel, Richard Muntz, and Vince Busam. Dynamo: design, implementation, and evaluation of cooperative persistent object management in a local area network. Software—Practice and Experience, 30(4):419–448, April 10, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
71004142/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=71004142&PLACEBO=IE.
pdf.

Yu:2012:NWM

[YWT^{+}12] Zhiwei Yu, Chaokun Wang, Clark Thomborson, Jianmin Wang, Shiguo Lian, and Athanasios V. Vasilakos. A novel watermarking method for software protection in the cloud. Software—Practice and Experience, 42(4):409– 430, April 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yang:2024:NBM

[YYLJ24] Yun-Tai Yang, Hong-Bao Ye, Yi-Shan Li, and Hewijin Christine Jiau. A noise blocking methodology for effective collaboration. Software—Practice and Experience, 54(5):875– 895, May 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yang:2011:CAD

[YYSG11] Yanqin Yang, Haijin Yan, Zili Shao, and Minyi Guo. Compiler-assisted dynamic scratch-pad memory management with space overlapping for embedded systems. Software—Practice and Experience, 41(7):737– 752, June 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yu:2022:EIA

[YZL^{+}22] Sui Yu, Lichen Zhang, Peng Li, Lixia Li, Bin Yan, and Zhipeng Cai. Efficient interest-aware data dissemination in mobile opportunistic networks. Software —Practice and Experience, 52(4):917–928, April 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yang:2012:SQR

[YZW^{+}12] Xiaohu Yang, Liping Zhao, Xinyu Wang, Ye Wang, Jie Sun, and Albert Jerry Cristoforo. Satisfying quality requirements in the design of a partition-based, distributed stock trading system. Software—Practice and Experience, 42(2):131– 157, February 2012. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yang:2007:RUL

[YZYL07] Jin-Min Yang, Da-Fang Zhang, Xue-Dong Yang, and Wen-Wei Li. Reliable user-level rollback recovery implementation for multithreaded processes on windows. Software—Practice and Experience, 37(3):331– 346, March 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhao:2007:AFO

[ZA07] Peng Zhao and José Nelson Amaral. Ablego: a function outlining and partial inlining framework. Software —Practice and Experience, 37(5):465–491, April 25, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zamboni:2003:ESS

[Zam03] Diego Zamboni. Editorial: Security Software. Software —Practice and Experience, 33(5):395–396, April 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zelkowitz:1974:OSP

[ZB74] Marvin V. Zelkowitz and William G. Bail. Optimization of structured programs. Software—Practice and Experience, 4(1):51– 57, January/March 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ziafat:2018:OSV

[ZB18] Hassan Ziafat and Seyed Morteza Babamir. Optimal selection of VMs for resource task scheduling in geographically distributed clouds using fuzzy c-mean and MOLP. Software— Practice and Experience, 48(10):1820–1846, October 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zese:2023:WAR

[ZB23] Riccardo Zese and Elena Bellodi. A web application for reasoning on probabilistic description logics knowledge bases. Software—Practice and Experience, 53(9):1741–1762, September 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zendra:2001:CAG

[ZC01] Olivier Zendra and Dominique Colnet. Coping with aliasing in the GNU Eiffel Compiler implementation. Software—Practice and Experience, 31(6): 601–613, May 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
78003104/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=78003104&PLACEBO=IE.
pdf.

Zhang:2002:ATC

[ZC02] Jian Zhang and S. C. Cheung. Automated test case generation for the stress testing of multimedia systems. Software—Practice and Experience, 32(15):1411–1435, December 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/
                                        99018097/START;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?
                                        ID=99018097&PLACEBO=IE.
                                        pdf.

Zhang:2003:MDF

[ZC03] Jia Zhang and Jen-Yao Chung. Mockup-driven fast-prototyping methodology for Web application development. Software— Practice and Experience, 33(13):1251–1272, November 10, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zini:2022:PCR

[ZCCB22] Matteo Zini, Giorgiomaria Cicero, Daniel Casini, and Alessandro Biondi. Profiling and controlling I/Orelated memory contention in COTS heterogeneous platforms. Software—Practice and Experience, 52(5):

1095–1113, May 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zheng:2006:MMC

[ZCN06] Yongjie Zheng, Alvin T. S.

Chan, and Grace Ngai. MCL: a MobiGATE coordination language for highly adaptive and reconfigurable mobile middleware. Software—Practice and Experience, 36(11– 12):1355–1380, September/ October 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhao:2013:DAT

[ZCO13] Dan Zhao, Shauvik Roy Choudhary, and Alessandro Orso. Developing analysis and testing plug-ins for modern IDEs: an experience report. Software —Practice and Experience, 43(4):465–478, April 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zobel:1995:FAM

[ZD95] Justin Zobel and Philip W.

Dart. Finding approximate matches in large lexicons. Software—Practice and Experience, 25(3):331– 345, March 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zdun:2007:SPS

[Zdu07] Uwe Zdun. Systematic pattern selection using pattern language grammars and design space analysis. Software—Practice and Experience, 37(9):983– 1016, July 25, 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2017:RRP

[ZDY^{+}17] Weishan Zhang, Pengcheng Duan, Laurence T Yang, Feng Xia, Zhongwei Li, Qinghua Lu, Wenjuan Gong, and Su Yang. Resource requests prediction in the cloud computing environment with a deep belief network. Software—Practice and Experience, 47(3):473–488, March 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zelkowitz:1972:PMI

[Zel72] Marvin V. Zelkowitz. PIT: a macro-implemented implementation language. Software—Practice and Experience, 2(4):337–346, October/December 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zelkowitz:1977:ESP

[Zel77] Marvin V. Zelkowitz. Effects of structured programming on PL/I programmers. Software— Practice and Experience, 7(6):793–795, November 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zelkowitz:1980:CSR

[Zel80] Marvin V. Zelkowitz. A case study in rapid prototyping. Software—Practice and Experience, 10 (12):1037–1042, December 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2006:CHD

[ZG06] Youtao Zhang and Rajiv Gupta. Compressing heap data for improved memory performance. Software—Practice and Experience, 36(10):1081–1111, August 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhu:2024:FLF

[ZGB24] Wuji Zhu, Mohammad Goudarzi, and Rajkumar Buyya. FLight: a lightweight federated learning framework in edge and fog computing. Software —Practice and Experience, 54(5):813–841, May 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2007:LFC

[ZGG07] Xiangyu Zhang, Neelam Gupta, and Rajiv Gupta. Locating faulty code by multiple points slicing. Software—Practice and Experience, 37(9):935–961, July 25, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2023:TBA

[ZGGL23] Weizhe Zhang, Wenchao Gu, Cuiyun Gao, and Michael R. Lyu. A transformer-based approach for improving app review response generation. Software—Practice and Experience, 53(2):438–454, February 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2023:LSP

[ZGL^{+}23] Sen Zhang, Xiao-Wei Guo, Chao Li, Yi Liu, Sijiang Fan, Ran Zhao, and Canqun Yang. A large scale parallel fluid-structure interaction computing platform for simulating structural responses to a detonation shock. Software —Practice and Experience, 53(1):211–240, January 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zheng:1991:DDT

[ZH91] Yi Zheng and Jim Hague.

DMT — a demonstration tool. Software—Practice and Experience, 21(9):949– 961, September 1991. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zastre:2001:EE

[ZH01] Michael Zastre and R. Nigel Horspool. Exploiting exceptions. Software— Practice and Experience, 31(12):1109–1123, October 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/
85008991/START;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=85008991&PLACEBO=IE.
pdf.

Zhou:2003:CCB

[Zho03] Neng-Fa Zhou. CGLIB — a constraint-based graphics library. Software— Practice and Experience, 33(13):1199–1216, November 10, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhou:2019:CFS

[ZHO^{+}19] Huan Zhou, Yang Hu, Xue Ouyang, Jinshu Su, Spiros Koulouzis, Cees de Laat, and Zhiming Zhao. CloudsStorm: a framework for seamlessly programming and controlling virtual infrastructure functions during the DevOps lifecycle of cloud applications. Software— Practice and Experience, 49(10):1421–1447, October 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhu:2014:ICC

[ZHZ^{+}14] Jianlin Zhu, Jin Huang, Daicui Zhou, Federico Carminati, Guoping Zhang, and Qiang He. Identifying composite crosscutting concerns through semisupervised learning. Software—Practice and Experience, 44(12):1525–1545, December 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhu:2017:MVW

[ZHZ17] Yi Zhu, Zhiqiu Huang, and Hang Zhou. Modeling and verification of Web services composition based on model transformation. Software—Practice and Experience, 47(5): 709–730, May 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhong:2022:ICS

[ZHZL22] Chenxing Zhong, Huang Huang, He Zhang, and Shanshan Li. Impacts, causes, and solutions of architectural smells in microservices: an industrial investigation. Software—Practice and Experience, 52(12):2574–2597, December 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zimmer:1990:RS

[Zim90] J. A. Zimmer. Restructuring for style. Software —Practice and Experience, 20(4):365–389, April 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2015:ACI

[ZJY^{+}15] Yanmei Zhang, Shujuan Jiang, Guan Yuan, Xiaolin Ju, and Hongchang Zhang. An approach of class integration test order determination based on test levels. Software—Practice and Experience, 45(5):657– 687, May 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhao:2022:FC

[ZKB^{+}22a] Zhiming Zhao, Spiros Koulouzis, Riccardo Bianchi, Siamak Farshidi, Zeshun Shi, Ruyue Xin, Yuandou Wang, Na Li, Yifang Shi, Joris Timmermans, and W. Daniel Kissling. Featured cover. Software— Practice and Experience, 52(9):i, September 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhao:2022:NVN

[ZKB^{+}22b] Zhiming Zhao, Spiros Koulouzis, Riccardo Bianchi, Siamak Farshidi, Zeshun Shi, Ruyue Xin, Yuandou Wang, Na Li, Yifang Shi, Joris Timmermans, and W. Daniel Kissling. Notebook-as-a-VRE (NaaVRE): From private notebooks to a collaborative cloud virtual research environment. Software—Practice and Experience, 52(9):1947–1966, September 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2021:EAS

[ZKW21] Michael Zhang, Chandra Krintz, and Rich Wolski. Edge-adaptable serverless acceleration for machine learning Internet of Things applications. Software—Practice and Experience, 51(9):1852–1867, September 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhou:2021:SHC

[ZKZ^{+}21] Junlong Zhou, Angeliki Kritikakou, Dakai Zhu, Jose L. Martinez Lastra, and Shiyan Hu. Software and hardware co-design for sustainable cyber-physical systems. Software—Practice and Experience, 51 (11):2105–2107, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zwaenepoel:1984:PRP

[ZL84] Willy Zwaenepoel and Keith A. Lantz. Perseus: retrospective on a portable operating system. Software —Practice and Experience, 14(1):31–48, January 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2008:VTB

[ZLG08] Hua Zhang, Joohan Lee, and Ratan Guha. VCluster: a thread-based Java middleware for SMP and heterogeneous clusters with thread migration support. Software—Practice and Experience, 38(10):1049– 1071, August ??, 2008. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2018:RRA

[ZLTX18] Yifei Zhang, Yue Li, Tian Tan, and Jingling Xue. Ripple: Reflection analysis for Android apps in incomplete information environments. Software—Practice and Experience, 48(8): 1419–1437, August 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2577.

Zhang:2011:MPT

[ZLWG11] Pengcheng Zhang, Wenrui Li, Dingsheng Wan, and Lars Grunske. Monitoring of Probabilistic Timed Property Sequence Charts. Software—Practice and Experience, 41(7):841– 866, June 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2024:OCU

[ZLX^{+}24] Haotong Zhang, Weiwei Lin, Rong Xie, Shenghai Li, Zhiyan Dai, and James Z. Wang. An optimal container update method for edge-cloud collaboration. Efficient Management of Microservicebased Systems and Applications: Software: Practice and Experience, 54(4):617– 634, April 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zou:2018:MFR

[ZLY18] Quan Zou, Guoqing Li, and Wenyang Yu. MapReduce functions to remote sensing distributed data processing — global vegetation drought monitoring as example. Software—Practice and Experience, 48 (7):1352–1367, July 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2578.

Zhang:2019:RNO

[ZLZ^{+}19] Fei Zhang, Guangming Liu, Bo Zhao, Xiaoming Fu, and Ramin Yahyapour. Reducing the network overhead of user mobility-induced virtual machine migration in mobile edge computing. Software—Practice and Experience, 49(4):673– 693, April 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2021:SOO

[ZLZ^{+}21] Jie Zhang, Xiang Li, Xuyun Zhang, Yuan Xue, Gautam Srivastava, and Wanchun Dou. Service offloading oriented edge server placement in smart farming. Software—Practice and Experience, 51 (12):2540–2557, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2023:PSP

[ZLZ^{+}23] Jian Zhang, Jie Liu, Naichun Zhou, Jing Tang, Xie He, and Jianqiang Chen. ParTransgrid: a scalable parallel preprocessing tool for unstructuredgrid cell-centered computational fluid dynamics applications. Software—Practice and Experience, 53(1): 6–26, January 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zobel:1995:ACF

[ZM95] Justin Zobel and Alistair Moffat. Adding compression to a full-text retrieval system. Software—Practice and Experience, 25(8):891– 903, August 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zaplata:2013:DFC

[ZML13] Sonja Zaplata, Matthias Meiners, and Winfried Lamersdorf. Designing future-context-aware dynamic applications with structured context prediction. Software—Practice and Experience, 43(10): 1185–1204, October 2013.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2022:SR

[ZML22] Hantao Zhang, Baoluo Meng, and Yiwen Liang. Sort race. Software—Practice and Experience, 52(8): 1867–1878, August 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhu:2018:ESS

[ZNWS18] Xiaoyan Zhu, Binbin Niu, E. James Whitehead Jr., and Zhongbin Sun. An empirical study of software change classification with imbalance data-handling methods. Software—Practice and Experience, 48 (11):1968–1999, November 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zorn:1993:MCC

[Zor93] Benjamin G. Zorn. The measured cost of conservative garbage collection. Software—Practice and Experience, 23(7):733– 756, July 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zuniga-Prieto:2018:DRC

[ZPGHIA18] Miguel Zúñiga-Prieto, Javier González-Huerta, Emilio Insfran, and Silvia Abrahão.

Dynamic reconfiguration of cloud application architectures. Software—Practice and Experience, 48(2): 327–344, February 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

https://onlinelibrary.
wiley.com/doi/abs/10.
1002/spe.2457.

Zunino:2007:BSS

[ZPSC07] C. Zunino, I. Pompili, A. Sanna, and L. Ciminiera. A brokerage system for solar data archives. Software—Practice and Experience, 37(8): 881–896, July 10, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zolfaghari:2021:RCD

[ZPSH21] Behrouz Zolfaghari, Reza M. Parizi, Gautam Srivastava, and Yoseph Hailemariam. Root causing, detecting, and fixing flaky tests: State of the art and future roadmap. Software —Practice and Experience, 51(5):851–867, May 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhu:2024:NIB

[ZQWL24] Xiaoyan Zhu, Tian Qiu, Jiayin Wang, and Xin Lai. A novel instance-based method for cross-project just-in-time defect prediction. Software—Practice and Experience, 54(6): 1087–1117, June 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zorzo:1999:UCA

[ZRX^{+}99] A. F. Zorzo, A. Romanovsky, J. Xu, B. Randell, R. J. Stroud, and I. S. Welch. Using coordinated atomic actions to design safety-critical systems: a production cell case study. Software—Practice and Experience, 29(8):677– 697, July 10, 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?
ID=62004390;   http://
www3.interscience.wiley.
com/cgi-bin/fulltext?
ID=62004390&PLACEBO=IE.
pdf.

Zada:2023:OSE

[ZSAM23] Islam Zada, Sara Shahzad, Shaukat Ali, and Raja Majid Mehmood. OntoSuSD: Software engineering approaches integration ontology for sustainable software development. Software—Practice and Experience, 53(2):283–317, February 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zimmermann:2005:SEY

[ZSFY05] Roger Zimmermann, Cyrus Shahabi, Kun Fu, and Shu-Yuen Didi Yao. Scalability evaluation of the Yima streaming media architecture. Software—Practice and Experience, 35(4):345– 359, April 10, 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhong:2023:SBC

[ZSH^{+}23] Yi Zhong, Mengyu Shi, Jiawei He, Chunrong Fang, and Zhenyu Chen. Securitybased code smell definition, detection, and impact quantification in Android. Software—Practice and Experience, 53 (11):2296–2321, November 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhou:2021:BBB

[ZSL21] Ao Zhou, Qibo Sun, and Jinglin Li. BCEdge: Blockchain-based resource management in D2Dassisted mobile edge computing. Software—Practice and Experience, 51(10): 2085–2102, October 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zolfaghari:2022:EAV

[ZSRR22] Rahmat Zolfaghari, Amir Sahafi, Amir Masoud Rahmani, and Reza Rezaei. An energy-aware virtual machines consolidation method for cloud computing: Simulation and verification. Software—Practice and Experience, 52(1):194– 235, January 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhu:2024:RAD

[ZTC24] Lulai Zhu, Damian Andrew Tamburri, and Giuliano Casale. RADF: Architecture decomposition for function as a service. Efficient Management of Microservice-based Systems and Applications: Software: Practice and Experience, 54(4):566–594, April 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2022:UCT

[ZWAL22] Shaoyang Zhang, Lian Wen, Sajid Anwer, and Baoxing Liu. Using composition trees to matching software requirements — an external agency’s approach to support software acquisition. Software—Practice and Experience, 52(6):1402–1425, June 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zheng:2021:GES

[ZWC21] Zibin Zheng, Shangguang Wang, and Rodrigo N. Calheiros. Guest editorial: Special issue on blockchain and decentralized applications. Software—Practice and Experience, 51(10):1985–1986, October 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zou:2022:SSH

[ZWF^{+}22] Xiaomin Zou, Fang Wang, Dan Feng, Feiyu Yang, Mengya Lei, and Chaojie Liu. SPHT: a scalable and high-performance hashing scheme for persistent memory. Software—Practice and Experience, 52(7): 1679–1697, July 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zeng:2017:NSD

[ZWKX17] Lingfang Zeng, Yang Wang, Kenneth B. Kent, and Ziliang Xiao. Naplus: a software distributed shared memory for virtual clusters in the cloud. Software—Practice and Experience, 47(9):1201–1220, September 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2014:DIT

[ZWML14] Wanfeng Zhang, Lizhe Wang, Yan Ma, and Dingsheng Liu. Design and implementation of task scheduling strategies for massive remote sensing data processing across multiple data centers. Software —Practice and Experience, 44(7):873–886, July 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2022:MVN

[ZWQC22] Peiying Zhang, Chao Wang, Zeyu Qin, and Haotong Cao. A multidomain virtual network embedding algorithm based on multiobjective optimization for Internet of Drones architecture in Industry 4.0. Software—Practice and Experience, 52(3):710–728, March 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhu:2015:APL

[ZWSS15] Xiaoyan Zhu, E. James Whitehead, Jr., Caitlin Sadowski, and Qinbao Song. An analysis of programming language statement frequency in C, C++, and Java source code. Software—Practice and Experience, 45(11):1479–1495, November 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhou:2021:NTP

[ZWX^{+}21] Jian Zhou, Haoming Wang, Fu Xiao, Xiaoyong Yan, and Lijuan Sun. Network traffic prediction method based on echo state network with adaptive reservoir. Software—Practice and Experience, 51 (11):2238–2251, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhao:2017:UAR

[ZXT^{+}17] Jiaqi Zhao, Changlong Xue, Xinlin Tao, Shugong Zhang, and Jie Tao. Using adaptive resource allocation to implement an elastic MapReduce framework. Software—Practice and Experience, 47(3):349– 360, March 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zeng:2017:TCE

[ZXW^{+}17] Lingfang Zeng, Shijie Xu, Yang Wang, Kenneth B. Kent, David Bremner, and Chengzhong Xu. Toward cost-effective replica placements in cloud storage systems with QoSawareness. Software— Practice and Experience, 47(6):813–829, June 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2024:PMJ

[ZXX24] Qiang Zhang, Lei Xu, and Baowen Xu. Python meets JIT compilers: a simple implementation and a comparative evaluation. Software—Practice and Experience, 54(2):225–256, February 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhou:2020:EAC

[ZYF20] Ziyi Zhou, Huiqun Yu, and Guisheng Fan. Effective approaches to combining lexical and syntactical information for code summarization. Software—Practice and Experience, 50 (12):2313–2336, December 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zeng:2020:DHC

[ZYW^{+}20] Feng Zeng, Lan Yao, Baoling Wu, Wenjia Li, and Lin Meng. Dynamic human contact prediction based on naive Bayes algorithm in mobile social networks. Software—Practice and Experience, 50 (11):2031–2045, November 2020. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Zhu:2022:JTD

[ZYW^{+}22] Xiaoyan Zhu, Chenyu Yan, E. James Whitehead, Jr., Binbin Niu, Lei Zhu, and Long Pan. Just-in-time defect prediction for software hunks. Software—Practice and Experience, 52(1):130– 153, January 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2012:TBN

[ZYYC12] Gaofeng Zhang, Yun Yang, Dong Yuan, and Jinjun Chen. A trust-based noise injection strategy for privacy protection in cloud. Software—Practice and Experience, 42(4):431– 445, April 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhao:2011:BPD

[ZZ11] Xulin Zhao and Ying Zou. A business process-driven approach for generating software modules. Software—Practice and Experience, 41(10):1049–1071, September 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2017:DLS

[ZZC^{+}17] Weishan Zhang, Dehai Zhao, Zhi Chai, Laurence T. Yang, Xin Liu, Faming Gong, and Su Yang. Deep learning and SVMbased emotion recognition from Chinese speech for smart affective services. Software—Practice and Experience, 47(8): 1127–1138, August 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhong:2024:DCC

[ZZH^{+}24] Chenxing Zhong, He Zhang, Huang Huang, Zhikun Chen, Chao Li, Xiaodong Liu, and Shanshan Li. DOMICO: Checking conformance between domain models and implementations. Efficient Management of Microservicebased Systems and Applications: Software: Practice and Experience, 54(4):595– 616, April 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhan:2021:FPG

[ZZJ21] Jinyu Zhan, Xingzhi Zhou, and Wei Jiang. Field programmable gate arraybased all-layer accelerator with quantization neural networks for sustainable cyber-physical systems. Software—Practice and Experience, 51 (11):2203–2224, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhu:2024:IRI

[ZZJZ24] Kun Zhu, Nana Zhang, Changjun Jiang, and Dandan Zhu. IMDAC: a robust intelligent software defect prediction model via multi-objective optimization and end-to-end hybrid deep learning networks. Software—Practice and Experience, 54(2):308– 333, February 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zaidan:2017:NDW

[ZZKA17] B. B. Zaidan, A. A. Zaidan, H. Abdul. Karim, and N. N. Ahmad. A new digital watermarking evaluation and benchmarking methodology using an external group of evaluators and multi-criteria analysis based on ‘large-scale data’. Software—Practice and Experience, 47(10): 1365–1392, October 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhao:2021:SBP

[ZZL^{+}21] Juan Zhao, Min Zhu, Xiaoyong Li, Zhenyu Huang, Jincai Li, and Junqiang Song. Solving Boolean polynomial systems by parallelizing characteristic set method for cyber-physical systems. Software—Practice and Experience, 51 (11):2143–2167, November 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhao:2024:UOD

[ZZL24] Ziyan Zhao, Li Zhang, and Xiaoli Lian. Usefulness of open domain model for identifying missing software requirements concepts. Software—Practice and Experience, 54(3): 437–464, March 2024. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhao:2024:DSI

[ZZLL24] Ziyan Zhao, Li Zhang, Xiaoli Lian, and Heyang Lv. DRIP: Segmenting individual requirements from software requirement documents. Software—Practice and Experience, 54(5):842– 874, May 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhao:2021:MSC

[ZZT^{+}21] Haitao Zhao, Yinyang Zhu, Jiawen Tang, Zhe Han, and Gagangeet Singh Aujla. Message-sensing classified transmission scheme based on mobile edge computing in the Internet of Vehicles. Software—Practice and Experience, 51 (12):2501–2518, December 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhou:1993:ULS

[ZZWD93] Songnian Zhou, Xiaohu Zheng, Jingwen Wang, and Pierre Delisle. Utopia: a load sharing facility for large, heterogeneous distributed computer systems. Software—Practice and Experience, 23 (12):1305–1336, December 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).